A Complete Bibliography of Publications in J.UCS: Journal of Universal Computer Science

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

01 July 2023 Version 1.35

Title word cross-reference

(α, β, γ, δ) [RHM15]. (E)T0L [CVPS95]. 120 [Mar04]. 2 [HFOB08]. 3 [CLCC10, HFOB08, IRMK12, uRLFH^{+}13, MM99, MGT14, Pel14, RPR11, SVFR15, SHK10, SE09, SBS15, TG10, VCB08]. 360^{∘} [BFF11]. 4 [Fod06, Mar04]. < [SMFM05]. > [SMFM05]. _{1}^{1} [Gom08]. ^{2} [EK00, HL03, dIVG^{+}06]. _{α} [FK16]. _{n} [Ior00]. α [FK16, LM07]. C[0; 1] [JW13]. C^{0} [ZG05]. d [RK97]. H [PY00, SDÖ^{+}12]. k [CT08b, KKB12, NNT16, VJTJ07, ZZ00]. N [Hon97, CT08b, Gon06, Ior00, PGT09]. O(\sqrt{n}) [CE06]. P [Alh04, BG04, CG04, CFSC04, FF04, Fon00, HNP98, IS04, Lan10, LZM04, MVPP02, PI04, PY00, Pau07, PJRC04, PP99]. {p, q} [Mar06a]. S [ZZ00]. T [GHNT97, YS05, DG07]. T_{0} [Wei10]. T_{2} [Wei10]. x^{v} [Gon06]. Z/p [Mue04].

-algebraic [Hon97]. -Boxes [ZZ00]. -cell [Mar04]. -completeness [Lan10]. -Confidence [SDÖ^{+}12]. -decomposition [YS05]. -Depletion [GHNT97]. -families [FK16]. -GD-Sat [VJTJ07]. -norm [DG07]. -out-of- [CT08b]. -Person [PGT09]. -SAT [MM99]. -Selective [HNP98]. -Social [RHM15]. -Stable [LM07]. -Systems [PP99]. -th [RK97]. -th-Order [ZZ00]. -Time [Fon00]. -valued [Ior00].

.NET [MI04, MABS05].

’01 [TM01a, TM01b]. ’02 [TM02a, TM02b]. ’03 [Toc03].

1 [RS01a]. 1.0 [PLSF08]. 10th [BdVG06]. 11th [BM07]. 1394 [MS00]. 1st [CMZZ07].

2 [RS01b]. 2.0 [BC11, BARR09, BRW03, GL11d, KR11, PCLCC11, PLSF08, SC14, vZdlH12]. 2000 [EGG^{+}01]. 2002 [Toc02]. 2005-2013 [SZWdP14]. 2008 [SAKAM11]. 21st [KM01].

3 [CFF^{+}13]. 3A [ESG10]. 3D [VDSF98]. 3D-Display [VDSF98].

5.0 [IdFC05].

60th [CSY02b, CI05]. 65th [CSZ07].

7 [LKB^{+}02, SAKAM11]. 7th [IFd03].

802.11 [YLW^{+}14]. 8th [Lin04b].

9th [ML05].

= [Gaß09].

AAL [HBFV13]. Abilities [Ozd13]. Absence [Kri97]. Absorptive [MRP14]. Abstract [Ara97, BG97, Boe97c, BRS00b, BG01, BS01, BP08, DDG97, DGL03, GRS08, GK97, GS97, Kir09, OL08, SA97, Sch09a, STW09, Sch01c, SV08, SN01, WTA01, Win97b, ZSK09]. Abstraction [Aic01, DSLO04, SMSdB05]. Abstractions [For07]. Abstracts [SSS12]. Academia [GGP08b, LMA^{+}14, SZWdP14]. Academic [CM09, DL15, MC07]. Accelerate [AH04]. Acceleration [LL97]. Accelerometer [PS12]. Acceptance [HGIPCPPM11, Kro13, MX05, TTB13]. Accepting [DMM07, MVM00]. Access [ASW^{+}03, BDPSNG97, CCM09, GMP^{+}13, Güt08, Her09, Hul08, LPP96, MM07, MP09, NML09, NOGG^{+}13, STVT07, SSV02, XMLC13, dLH08]. Accessibility [BSP^{+}13, CSAC^{+}15, DD13, GMP^{+}13, Kim10, LN08, Poo03]. Accessible [ASHT^{+}16, AVA08, BRAS^{+}12, CIM14, GGMM^{+}13, SGLM16, WCH14]. according [IGS08]. Accumulation [CDCH09]. Accuracy [GOM^{+}13, NO98, TNRGCP^{+}13]. Accurate [LS95, RSW04, SHK10, DMMM95]. Achievements [Kar13, MM15]. Achieving [BV07, GL11a, HBFV13, MSM07]. ACIS [CMZZ07]. Acknowledgment [GPCZ^{+}13]. ACO [JNdMM12]. ACO-based [JNdMM12]. Acoustic [uRLFH^{+}13]. Acquisition [ARS16, SKHK14]. across [AY12, HMW08, HLNA^{+}12, KHLAP12]. Action [Ara03, BM13, CM03, DL15, MCM07, MMD12, YMP08]. Actions [BBP08, Kri97, MSM07]. Activation [RMGCGCF08, TL11]. Active [AGG^{+}08, HM00, PI04, Pos98]. ActiveTM [MB09]. Activities [CRLNAR05, CJH12, DBAB12, FMA^{+}05, GPCZ^{+}13, Güt08, HLNA^{+}12, HG11, IMR^{+}12]. Activity [AGO^{+}13, DBBS08, Dus05, GVRT^{+}10, OCW13, SA14]. Activity-Aware [GVRT^{+}10]. Activity-Based [Dus05]. Actor [DF05, SK13, SdBM05]. Acts [DGBM08]. Acute [Faz06]. Acyclic [Wat02]. Ad [EMZB14, HHHX09, HJZ07, KTJ05, LKK08, NOP08, VVM^{+}06, CAD^{+}06]. Ad-Hoc [KTJ05, NOP08, VVM^{+}06, CAD^{+}06]. Adaptability [FSELC13]. Adaptable [CC08b, FSELC13, VAH07]. Adaptation [AMA^{+}14, CSC08, CMP08, CCP11, GHHE^{+}08, IS10, Kel08, KJKS14, LSG^{+}14, MCG14, ONRV08, PRCRARLN10, PTNMC08, PB04, Sob05]. Adapted [SBPR15]. Adapting [CGLdMAC14, GRCGK14, PLSF08, SSBS08]. Adaptive [ARS16, BGP08, BS08, BSB09, CL95, CM98, CC08b, CCS10, COBP^{+}14, DB03, EDA14, Flo04, GDW10, GBCA12, GRGPL08, GSZ15, GSPK08, GSMBFPK10, HMW08, HC08, HBFV13, JP07, KWH03, LGAP11, LT13, MP10, NdMM12, NS05, Ret08a, SBGI14, SSAB^{+}13, SMMC10, SZZM10, TLJ11, ZSG14, dG15]. Adders [RK97]. Adding [CM00, HC08]. Additional [Ano07, Tak06]. Additive [CHPHV10, CSY02a, ÉK02, Gan04]. Address [BAML07]. Addressing [TT98]. ADDS [SFVFMN04]. Adjoint [BD05, JMP06]. Administration [SBGI14]. Admissible [Mul00]. Adoption [BS12a, BARR09, TTB13, vBK08]. Ads [MCMMAP^{+}14]. Adult [HJVK15]. Adults [Sch01b]. Advanced [ACM16, Car98a, GMdMC12, LMA^{+}14, MS11, Sch99]. Advances [Aki09, BBL13, CSY02b, CCS10, FMR09, GB10, JK12a, Lin08a, Lin08b, LZZK14, NR12, VRGPSP07, WP15, YMA15]. Advantage [AR04, ES05]. Adverbs [POJB08]. Adversarial [Kou09]. Advertising [DZ08, LVS13]. Advice [CL95, CBBT07, HNP98]. Advices [RMM^{+}08]. AEH [VBO08]. AES [PBTW07]. Affect [ASAAASJ16, CPCLSAGC11, HAFS15, LGZ01]. Affective [LGGGC14]. Affectively [FDC^{+}13]. Affects [Kar13]. Affine [Mes02, PPG95]. Affordances [GJP^{+}12b, KR11]. African [BKL12]. Again [LM94b]. Against [FL10, FFK04, HLC08]. Age [ABPS95, Hol96, NSL96]. Aged [HA10]. Agencies [SDLM14]. Agent [ADMdB09, BGP07, BB08a, BBP08, CS10, ČSŽ09, CYL11, FGS98, FHJ^{+}99, GVRT^{+}10, GSP04, KJZJ08, KOW01, KJ10, KDKDN08, KZ08, KL09a, LKK08, LS07, LEC11, MK12, NKS^{+}09, OCW13, OO08, PKP08, PT09, PB05, PZDH09, PPJ08, Ros05, Sat10, SH09, SOO97, Sto03, TEK08, LST14, PT09]. Agent-Based [BB08a, KDKDN08, PPJ08, GVRT^{+}10, LS07, OO08, Sat10, TEK08]. Agent-mediated [OCW13]. Agent-Oriented [FHJ^{+}99, KOW01].

Agents

[ABPS95, BBGV07, CO08, Dud08, GRGN13, JRR16, Kat05, ML02, MTK97, VO02]. Aggregation [BQBW03, GH08, LGAP11, LT09]. Agile [LULGFC13, SFP12]. Aging [Coo06]. Agreement [CT08a, Rad14]. Aided [CLCC10, Reb96]. Air [HA13]. Airborne [RMZ15]. Aksharas [MR11]. ALCQI [RH10]. Alert [GRGN13, Yan05]. Algebra [BB08a, DLL14, Hav05, Mar98, RBB06, Tru10]. Algebraic [ACA^{+}16, DF00, GMB08, Hon96, Kud99, MJ13, Pop98, Ste00, Hon97]. algebraicness [Hon99]. Algebras [BB10a, Cet00, CBO05, GL00, GK13, Ior00, Ior07, Ior08, JRW10, KSU02, Rat00, Spi05]. Algorithm [AC05, AK09, AR95, BDGW96, BM11, CT04, CE06, CW09, CW00, DXZL07, DL99, EACGFK13, Gir05, GGS08, GCC16, GCVRSPGP07, Gra98, HKK13, HHH^{+}02, HT06, HCK11, JS16, KPdF06, LLLL99, LVV10, LWG14, LHC^{+}13, LD06, LA07, NAK08, NZM09, NR08b, OFCB08, PB14, PBTW07, PKSR09, PPG95, QZ07, QGT^{+}14, RHM15, RS00, SBGI14, SESMT10, Sar05, SH09, TIL08, TWW07, TLJ11, WHCL09, WXZL15, Wat02, YS05, ZZH^{+}12, dSLMW08, DMMM95]. Algorithmic [Cal96b, Cha05, Mar06a, Mar96, Svo96]. Algorithms [AHT09, AT10b, And97, Bai12, BMM^{+}09, BCC^{+}06, BMV12, BC16, BCS15, CAGMPGdAS13, CAJ06, CS02, DT07, EC00, FOSS99, GYY09, HI00, Hem06, JNdMM12, LdSM08, MS01, PTO^{+}12, PKP08, Pav95, QQ11, RR06b, SKP08, SWY09, UDC97, VF11]. Alias [NI03]. Aligment [KJ10]. Align [LAAPVGMM15, SMV08]. Aligned [HLS15]. Aligning [MKI^{+}12]. Alignment [dLMVGM13, RVC12, Wol99a, Wol99b, Wol00]. Alleviate [BD00]. Allocation [AAM14, CAS^{+}13, PKP08, Şte95, VUT^{+}08]. Allocators [DM10]. Almost [AP09, dTU04]. Alpha [BCM12]. Alter [vB96]. Alternate [MM98]. Alternating [Dru06]. Alternation [Gup01]. Alternative [BCNR07, CCHdCN08, GD14]. Alternatives [KCK10]. Amalgamation [IAS16]. Ambient [AGGH08, BAR06, BdI10, FJP06, GMC08b, HME^{+}06, HB13, JRO10, dIdSGZ10, dIB13, PANS13, VVM^{+}06]. Ambients [PP99]. Ambiguity [vZG11]. AMCC [OCRPdlMG07]. America [Lin11b]. American [GCG08]. AmI [FG10, dIVG^{+}06]. Among [Jun10c, BR03, HKKvP08, KCK10, Pel14, PMP02, YYZ^{+}09]. Amplification [LZLW13]. Amplitude [Svo95]. Analog [IS10, WZC07]. Analysing [BS12a, Faz06, Fod06, GMK05]. Analysis [AA16a, AGMT10, AGA12, ASAAASJ16, AdGCD^{+}15, BHK10, BD06, BR97, Ber10, BDhKB09, Bör02, BISZ08, BCR09, CBRH12, CWTT11, CKPK13, CLC04, Cla05, DH10b, DT07, DJJN09, DBAB12, ESM08, EHEH05, FF08, FBCMR15, FP06, GSW04, HSD^{+}14, IO04, JL16, KSK^{+}16, KL02, KK13b, Kim10, Krä98, KL09a, Laf04, LRR04, LKHL09, LLCN09, LVS13, LL12, LLZ16, MSSY95, MGT14, MPF04, ND08, OMO10, PA12, PRAT09, PCKJ11, PS12, PO04, Puc10, RS00, SKSN07, Shi11, SZS12, SSGS10, SSS12, SK04, TCK^{+}01, TCW12, VPF09, VSGP05, VBO08, WKSD^{+}11, ZTN^{+}15, ZBKK12]. Analytical [TBVRGLD15]. Analytics [CRB15, DL15, EKW^{+}15, GXC^{+}15, LNH^{+}15, OAR^{+}14, PJ15, RLMS13, VKW15, dlFVPHB15]. Analyze [HGMT08]. Analyzing [BDN05, BvdTV09, CFMP15, CSA10, KKK16, MMB08, Ruf01, TB16, dCH11, dH00]. Android [JL16]. Angles [KD11]. Animated [DCR^{+}07]. Animations [LMO01]. ANN [BCG^{+}09, LWH09]. Annealing [FSSPLG^{+}13]. Annotated [LKMS08]. Annotation [BE98, LKB^{+}02, OHYJ16, RPCA15, SUKG13]. Annotations [CKdL08, dIdSGZ10, MRGF14, SLK11]. Anonymity [FM95]. Anonymization [RHM15]. Anonymous [FP06, LWY09]. Answering [SLD^{+}16]. Anti [KDKDN08]. Anti-Crisis [KDKDN08]. Antiport [CG04]. AOM [MG14]. AOP [CBF05, TF09, VM13]. apartness [Vît05]. Aperiodic [CK95]. APIs [MI07]. APOSDLE [SKH^{+}10]. App [AHEAS^{+}15, MV15, RKH15]. Append [LLS05]. Append-only [LLS05]. Application [Aïm98, And96, BFF99, BM05, BFN05, Cap05, CFMP15, DF99, GL11a, GJP^{+}12b, Güt08, Has02, INK09, KJZJ08, KKB12, KP00, Kul07, LHK^{+}13, LLCN09, LGL09, LSX12, LULGFC13, LMRG14, MBC13, MOG^{+}10, ML98, Mes02, MNS^{+}12, MPL11, PS09, QQ11, SFVFMN04, SA09, TJS^{+}13, TSE^{+}15, TD96, TEC^{+}07, TLJ11, TKT07, WKSD^{+}11, XWGS09]. Application-Level [LMRG14]. Applications [AA16b, Ada06, ACM16, AS07, AVA08, BAZ14, Ber10, BdGFMT14, BG01, BSPO11, BPS11, CPMVG13, DB03, DPZ08, EMGB^{+}12, FSMP07, FJP06, GSW04, GPFL12, GTGT10, GL11c, GL11d, HKTV06, HM00, HMM00, HBI98, HMMP10, JL09, JK12a, KSK^{+}16, KKM08, KH12, Kon03, KTKP09, LSK06, LM94a, LGGGC14, LS10, LMRG14, LGP10, Mar95, MG14, MT02, MHH98, NR12, Ngu09, PT09, PMRO08, PLG^{+}08, PD99, Pop07, SLPL14, Spa08, SCLM03, TFG06, TGLP10, URG^{+}13, VBVNHdDSL12, VRGPSP07, VCB08, VC13, VTGA13, WL13, YJY14, YSP09, ZGC^{+}08, ZABB07]. Applied [CR12, FOAB08, GPCPL12, GPVILN13, GRGPL08, MS00, MCC13, Poo03, SMFM05, HBFV13]. Apply [Die96]. Applying [BdGFMT14, CMSE09, FML13, GR02, HB00, HCWA03, Hor04, KKTT16, LULGFC13, MLHCGB16, MHH98, RRR10, VBVNHdDSL12]. Approach [AGMT10, Aic01, ASAAASJ16, ALHM^{+}14, AdO11, APDC08, ACP06, BGMR^{+}16, BMa11, BdS13, BBC02, BS08, BHC05, BHS^{+}06, BM11, BMMM14, CGFSHG09, CBF05, CRMLN^{+}07, CRB15, CS10, CBR^{+}05, CTM10, CLC04, CS14, CS09, CG09, CDD^{+}03, CDD^{+}04, CP01, DM07, DXZL07, Dus05, EDA14, FGB^{+}14, FWT11, FAT^{+}13, FBSEGP15, Fro02, GDW10, GLCV08, GV00, GAMP10, HCBB15, HMW08, HH03, HAI13, JBH^{+}10, KDKB07, KHN99, Kon03, KDKDN08, KP00, KMN16, LGAP11, Laf04, LCHD12, LVV10, LPSF10, LWY09, MYCA11, Ma12, MYC14, MM98, Mar06a, MG14, MOS^{+}13, MDY10, MMS11, Mos05, MGAVF10, Nag06, NVB12, Ngu09, OCW13, OBO09, PTO^{+}12, PS09, PMLL09, PCKJ11, PCLT15, RPR11, RMGT09, RGRR15, RGHH97, SR10, dPSZPVLR^{+}16, SSAB^{+}13, SBPR15, SW04, SFVFMN04, SMV08, SVV10, Tar12, TE06]. Approach [Tra13, URG^{+}13, VJ09, WK05, XMZbL10, ZWH10, ZG97, dOBGH^{+}14, vP05]. Approaches [ARQH14, CCHdCN08, GD14, GOM^{+}13, Gär99, Gio98, HLHD^{+}07, KJKS14, MLX10, XLMR10]. Approval [FZ00]. Approximate [AHT09, And96, CNQ04, DMS05, Fen95b, KKB12, KHN99, Rat00]. Approximating [GD14]. Approximation [DXZL07, GH08, NR08b]. Approximations [Ang98]. Apps [OEK16]. APS [Dud08]. APT [CCYK15]. Arabic [ASAAASJ16, AZMA15, BZA08, DA13, SSSS06]. Arbiters [DM10]. Arcade [FKK^{+}10]. Architectural [DSLO04, JST11, PF11, PR06]. Architecture [BGK02, BdGFMT14, BEH^{+}05, CCP^{+}07, DRRGdP07, dAFL07, For97, GVRT^{+}10, GPL13, GHS06, HMB09, HMW08, IMR^{+}12, Kap95, KAG00, KCKL10, LGMM^{+}13, LKMS08, LSG^{+}14, MJGS12, MPF^{+}16, MK12, NBGS06, Oli12, PdCdlTR06, Riz15, RGPK15, SMGMT09, SW04, TNM09, TSCY01, VC13, WZC07, ZAB^{+}08]. Architectures [AFL08, ACL95, BP09, ESM08, FKO14, GG08, GPFL12, GN00, HM99, JGM^{+}13, MN00, OMO10, PRCCS13, POB11a, RVW07, RdL08, WO09, dAO13]. Archives [ASW^{+}03, LKMS08, Wac02]. Archiving [CS03]. Area [Kuh03, NVB12]. Areas [CSZ07]. Argumentation [AT13, PGT09]. Ariadne [SD97]. ARIES [Kir09]. ARIES-based [Kir09]. Arithmetic [AGO^{+}13, AR01, CC96, LS95, Mar95, Mes02, NK95, Pop95]. ARLearn [TKK^{+}12]. ARM [TTB09]. Armed [Pet09]. arranged [Oli01]. Array [VTRN12]. Arrhythmias [ZHG^{+}06]. Arrow [MTK97]. Art [HSR10, Toc02, Toc04]. Artefact [NdCFB08]. Artefacts [dMTS^{+}14]. Articulation [JL09]. Artificial [CAGMPGdAS13, LHZS12, PA12, VV12, YLL^{+}07]. Arts [SVFR15]. ASCII [SS08]. ASIC [Hei96]. ASIPs [GHM04]. ASM [Boe97b, Bör02, EGG^{+}01, GR01, Mea97, Sch01a, Sch08a, ZG97]. ASM-Approach [ZG97]. ASM-Based [GR01]. ASMs [AFL08, BM97, Str97]. Aspect [ASAAASJ16, APDC08, CBBT07, FRD14, LRS^{+}11, MSA13, MDO^{+}09, PTNMC08, PRCCS13, PF11, PZ03, Tan08, URG^{+}13, VM13]. Aspect-Based [ASAAASJ16]. Aspect-Oriented [LRS^{+}11, MDO^{+}09, PTNMC08, PRCCS13, PF11, URG^{+}13, MSA13]. AspectJ [ARRB14, DB03, FR04, PHPP06, RMM^{+}08, TT08]. AspectJ-based [ARRB14]. AspectLua [BV07, CBF05]. Aspects [AFK01, BGBA10, BY97, CPHC11, FJP06, HL96, KMM14, dOMdAL^{+}08, Mau97i, PHJ^{+}08, Rad14, Ric05, SDJ99, TF09, ZD09]. Assembling [CMZZ07]. Assertive [CBR^{+}05]. Assessing [BCM12, Cam98, CE11, JBH^{+}10, LdSM08, MNF^{+}13, MM15]. Assessment [ARS16, Aïm98, Car98b, CdSCSSA16, GXC^{+}15, HCBB15, HLS15, HMHGR15, HAI13, Hop98, LRB16, LMMPFVI14, LLSA13, Ma12, NKS^{+}09, SVFR15, SDÖ^{+}12, SA03]. Assessments [HSFE12]. Assigned [GYY09]. Assignment [ANdMM08, CDD^{+}04, CDD^{+}07, FA06, LA07, MTB^{+}08, Man97, Rad96, YLW^{+}14]. Assignments [Ist07, SWY09]. Assimilation [NHH06]. Assistance [Coo06, TAL08, VK03, ZHG^{+}06]. Assistant [LV95, PT13]. Assistants [PT09]. Assisted [Bic15, FGS98, HB13, dIdSGZ10]. Assisting [GVRT^{+}10, GFT09]. Associate [WZZ^{+}09]. Association [AK09, GAMP10, NNT16, VTHM16]. Associative [OP15, SKH^{+}10]. Assurance [ATSJ05, PB05]. Astrocyte [Pau07]. Astrocyte-Like [Pau07]. Asymptotic [Leu07]. Asynchronous [BMV12, Bur05, GÁVCNC14, GFBR08, Kri97, MRO05, Tddd03]. AT-HOME [vZdlH12]. Athens [MRK^{+}98]. ATM [HHH^{+}02]. Atomic [ES05, FF07, MM06, JLRW05]. Atomicity [AC05, BJ05a, BJ05b, Bur05, CJ07, Kie05a, MSM07, MR05]. Attachment [QF12]. Attack [ABAL09, Che09, CCYK15, Yan05]. Attacks [AHT09, CDP13, CT08a, DGlS12, HSFE12, HLC08, Kou09, PHJ^{+}08, YKA16]. Attempt [Ior07, Ior08]. Attention [BVV^{+}10, DIKL14, LBG07, Sch02c]. Attitude [UT16]. Attitudes [HGIPCPPM11, OT16, UHÖD15]. Attribute [LYLX15, XMLC13, ZQQ15]. Attribute-Based [XMLC13, ZQQ15]. Attributes [PF11, Yon11, Zgr07]. Attribution [KD01]. Auctions [BB08a]. Audience [HPE14]. Audio [CS05b, uRLFH^{+}13, SdBC13]. Audiovisual [ASW^{+}03]. Audit [CLC04]. Auditing [LCDP15, GMP^{+}13]. Augment [BBIC13, CCP^{+}07]. Augmented [CS07, DLL16, DAk13, GHO10, HFIBJ13, HBFV13, KH12, MCMMAP^{+}14, TKK^{+}12]. AULA [PMRO08, SMFM05]. Authentication [AP09, BR97, BDhKB09, HLCL11, NLLJ12, Shi11, Vle14, WLLL09].

Author [POJB08]. Authoring [AGG^{+}08, CC08b, CCS10, DDSS05, FMA^{+}05, GDW10, GSPK08, GSMBFPK10, HC08, IRMK12, MOMGSRFM07, OCB^{+}10, OB95, PRCRARLN10, SMMC10, Ste08]. Authority [ZQQ15]. Authorization [SWY09]. Auto [KM07]. AutoCAD [GL11a]. Autoconfiguration [BAML07]. Automata [AT10a, BC16, BvZH09, CL97, CS00, CSY02b, CFJS15, Cha05, CR07, Cv99, Dru06, Géc02, Kar02, MM99, Mar00b, MK99, Mat99, Sal02, STW09, SSSS10, Wat02]. Automata-Defined [STW09]. Automated [BGP07, CdSCSSA16, CWTT11, CDD^{+}07, CCMP08, LMMPFVI14, MK12, Rie02, Rus07, SMSdB05, vKL04]. Automatic [AG06, AP05, BQBW03, Cer97, CCHdCN08, GR01, GPLV13, GGDBP^{+}08, GSS99, Gir05, HC08, OHYJ16, QF12, QL12, RMGT09, SBCJ03, Vle14, dTU04, GHHE^{+}08]. Automatically [BGBA10, GMB11, PTL^{+}09, SS03]. Automating [DBAB12]. Automation [CJO^{+}13, GXC^{+}15, MPR^{+}08, NuR05]. Automaton [BAPG03, MSSY95]. Automorphism [Har00]. Autonomic [ZDE14]. Autonomous [CBR^{+}05, NH09, SM02b]. Avalanche [ZZ95]. Avatar [AT07]. Avatar-based [AT07]. AVC [LLYC12]. Average [DL99]. Avoidance [GG08, Gan04, HCH^{+}09]. Award [NH03]. Aware [AWGS04, CCP11, DMCM14, EDA14, FKB^{+}15, FWT11, FAT^{+}13, GVRT^{+}10, GO14, JK10b, Jun10c, LPSF10, LS10, LGP10, OOHS06, SBGI14, SW04, SSM11, SKSP09, SLJC08, SMP^{+}11, TSE^{+}15, VVM^{+}06, WLtHN14, AGG^{+}08, GKZ05, HCBB15, Sat10, TGLP10, VDLG10]. Awareness [DLL16, GRCGK14, HBFV13, IS10, MDO^{+}09, MNDRF10, WP03]. Axiomatic [Gol05, KCK10]. Axiomatization [DOM10].

B [BCD13, SBCD15, Yeu04]. B2C [Ros05]. Back [LdSM08, ZG97, dSLMW08, LEJ^{+}08]. Back-Ends [ZG97]. Back-to-Front [LdSM08, dSLMW08]. Background [BZA08, DGK^{+}99, Gär99]. Bad [PHPP06, Tom03]. Baffletext [SSSS06]. Bagatelle [Oli01]. Balance [Pop07]. Balanced [Lep98]. Balancing [BMV12, HCH^{+}09, HJZ07, LWY11]. Banach [Hav05]. Band [BCM12]. Bandwidth [SCW08, SLT08]. Bandwidths [SCW08]. Bank [Bai12, Has02]. Banking [KW10]. Banks [BKL12]. Barriers [Hul08, LK01]. Barter [BCNR07]. Base [Bec03, FC03, Kuz04, Ach06]. Based [AA16a, Ada06, Aic01, ASAAASJ16, AZMA15, AK09, AGO^{+}13, ARS^{+}08, AGGH08, ACB02, Azz10, BGP07, BB08a, BD06, BHH^{+}06, BS03, BCG^{+}09, BDL^{+}06, BARB12, BCNR07, BE98, CGFSHG09, CHPHV10, CJO^{+}13, CCH06, CI12, CLM10, CKdL08, CDD^{+}03, CCMP08, CS02, CVK97, Cv99, DTG10, DG07, DXZL07, Dus05, Dvo00, DSRR03, DM10, EIH08, ESB04, EMGB^{+}12, FL14, Fro02, FLF^{+}14, GR01, GRGN13, GBHA12, GGP08a, GGMM^{+}13, GHM04, GB03, HPE14, HKS96, HR06, HIGM13, HNJ^{+}10, JRR16, Jun05a, Jun08, KNLS00, KJZJ08, KY10, KKB12, KTL^{+}11, KDKDN08, Kri97, KMN16, Kwo97, LGAP11, LF16, LDO^{+}12, LdPK^{+}14, LZM04, LST14, Lin09, LRS^{+}11, LLSA13, LLZ16, MH96, MMEdP12, MGM^{+}08, Mau97i, MR08, MOS^{+}13, MMS11, MUSA03, MGPB07, MP09, NPC^{+}09, NLLJ12, NWDX09, NW04, OPP09, OBO09, PT13]. Based [PEPP08, PT09, PRS95, PKSR09, PD04, PB04, PPJ08, PF15, QZ07, RTB13, RBLR02, RP98, Saf08, SAA08, SNAF07, SBTH04, SdBM05, Sob05, SA11, SKL08, SRR04, TWW07, TCW12, VPF09, VdSdMC08, VTGA13, VdR09, Vle14, WSL07, WHCL09, WPL98, WK05, WMJ^{+}07, XMLC13, XHP^{+}09, XLMR10, YLL^{+}07, YX10, YLW^{+}14, Yeu04, ZG05, ZZH^{+}12, ZQQ15, dH04, dlELR09, ASH11, ASTL07, AHSN01, AO04, AdO11, ACA^{+}16, ARRB14, AT07, AH04, AY12, AFP^{+}13, Bar03, BMa11, BRR99, Bjø01, BS08, CAGMPGdAS13, CNQ04, CXB12, CAD^{+}06, Cam98, CM09, CYL11, Che09, CB12, CDD^{+}04, CDD^{+}07, CCS10, CCP11, DSM13, DDS04, DM04, DZBB^{+}12, DHC11, ESG10, Esp06, FGS98, FL10, FZ00, FMT^{+}15, GFO12, GMHGRG^{+}13, GVRT^{+}10, GRS08, GLSD11, GAMP10, GHS06, GCC16, HM01, HBF10, HH03, HKK13, HP15]. based [IPCVC12, JM15, JV11, JL16, JGL08, JGM10, JNdMM12, KK13a, KDKB07, Kim10, KH12, Kir09, Köh09, KOW01, KJ10, KTKP09, KO99, Kro13, LN02, LNML03, LS07, LGMM^{+}13, LA03b, LGES11, LSX12, LNH^{+}15, MWM10, MJGS12, MLX10, MMB08, NSMBACBG12, Nav09, NDAM09, NdMM12, NZCG05, NS05, NSNK05, NSF^{+}10, OCW13, ONRV08, OO08, PTO^{+}12, PRB^{+}11, PLB14, PS09, PZJ09, PMLL09, PCLT15, PR06, QF12, RPR11, RSW04, dPSZPVLR^{+}16, Sat10, SMMC10, SCK^{+}09, Sha11, SH09, SBCJ03, SKHK14, Smy00, SKH14, SvRvdV^{+}13, SvRS14, Sto03, Tar12, TE06, TEK08, Tra13, VAH07, VJ09, VVM^{+}06, VSML03, WXZL15, WL13, WZZ^{+}09, WD02, WMSH09, ZWH10, ZTN^{+}15, ZDE14, vZdlH12, Hop98]. Bases [Bos09, CCHdCN08, Her02]. Basic [OK98]. Basis [BNA15, RH10]. Battles [LTY^{+}16]. Bayes [Alm06, LEJ^{+}08]. Bayesian [BMM^{+}09, GSMBFPK10, KK13b]. BCK [Ior07, Ior08]. BCTCS [DG00]. BDD [AMS04, CNQ04]. BDD-based [CNQ04]. BDDs [DM10]. BDI [MK12]. Be [Pos01, KSdV09]. Beating [PJH12]. Become [SGLM16]. Becoming [Abr07]. before [YYZ^{+}09]. Behavior [AY00, BST09, Bor07, HB11, IPCVC12, KCKL10, MCC13, ODSO11, RVC12]. Behavioral [APR05, Dru12, Dru13, GH08, GK06, MM06, Mar06b]. Behaviors [JKKW16]. Behaviour [EGDG09, KK13b, PS12, RPCA15, TTB09]. Behavioural [Ber06, CSC08, CR07, DF00]. Behind [HOPN11]. Being [IGS08]. Belgian [RLL^{+}10]. Belief [KK13b]. Believe [LM94b]. Benchmarking [Vie03]. Benefits [BdGFMT14, Hal07, NH03]. Benzenoids [BR07]. Berlin [BDL^{+}06]. Best [BDM15, BKL12]. Better [AS14]. Between [BCG^{+}09, CSY02a, Ior00, uRLFH^{+}13, SZZ95, Ara03, AGK^{+}10, CNRM03, DTG10, EMGB^{+}12, Her05, HMSS01, KM07, Luk08, dOMdAL^{+}08, Sch08b, Sch05c, TBL15, VLE12, Yeu04]. Beyond [CL07, Fra98, JRO10, SS08, SKL08, Toc04, SLD^{+}16]. Bias [RB08, PU97]. Bibliography [CS03]. Bibliometric [KL02]. Biconnected [AR95]. Bifurcation [BCG^{+}09]. Big [JCB15, LAHZ^{+}15, TBVRGLD15, VPB15, YMA15]. BigBatch [dOMdAL^{+}08]. Bilateral [CT08a]. Binarization [BSB09, SKP08]. Binarize [dSLMW08]. Binary [CWTT11, KM95, Sar05, Ver08]. Binding [ARRB14]. Bins [PJRC04]. Bio [BCM12, LWS11, NR12]. Bio-Inspired [BCM12, LWS11, NR12]. Biochemical [DH10b]. Biological [AM96, BD06, KDKB07, Lav96]. Biologically [RAC10]. Biomedical [TRR06]. Biomes [CT16]. Biometrics [SMMM13]. Biometry [MMD12]. Bipartite [LM07]. Birthday [CSY02b, CI05, CSZ07]. Bisimulations [dFER06]. Bit [Bos08a, LKK08]. Bit-Complexity [Bos08a]. Bit-map [LKK08]. Bitstream [dLH08]. Black [ABB14, YKA16]. Black-box [ABB14]. Blended [DD13]. Blending [GFBR08]. Blind [SdBC13]. Bloch [Ret08b]. Block [FL10]. Block-based [FL10]. Blocking [GN00, AC05]. Blocks [BBC02, FC03, KC08]. Blogosphere [ZBKK12]. Blowup [RWZ09]. BLTI [FCM^{+}12]. Bluetooth [JFL^{+}13]. Board [PdlCBKN14, Ret08a]. Bodies [Mar98, Wur10]. Book [Bjø01]. Bookmarks [Jun05a]. Boole [Mar00a]. Boolean [Bol06, Dun96, DL99, FH00, Ist07, OS98, RR06a, Rud04]. Boost [ZD09]. Boosting [KK13a]. Boosting-based [KK13a]. Bootstrapping [Lar01]. Border [FTARR05b, LKT10, MGMS12]. Borders [VLE12]. Bosch [Kuh03]. Bot [MPF^{+}16]. Botnet [KSK^{+}16, VV12]. Bound [HYC^{+}05, KWH03, SH10]. Bounded [BR05, BSHI99, BM96, CNQ04, HHH^{+}02]. Bounded-Weight [BSHI99]. Bounds [AV07, Akr09, Dvo97, Man97, PŞ95, Pav95, Şte05]. box [ABB14]. Boxes [NdMM06b, ZZ96, ZZ00]. Boys [BBM12]. BPEL [BFN05, DMCM14]. BPM [DKL10]. Brain [BDL^{+}06]. Brain-Computer [BDL^{+}06]. Brainstorming [AFP^{+}13, PPJ04]. Branch [Gra98]. Branching [DOM10, FH06, Roj96]. Brand [KKTT16, WWD15]. Brasilian [IFd03]. Brazilian [BdVG06, BM07, Lin04b, ML05, VMdCJ08, SSdS^{+}11]. Breadth [APM04]. Break [FBSEGP15]. Breakthroughs [BMUF14, CRC04]. Brick [BvZH09]. Bridge [CNRM03]. Bridges [TBL15, CI05]. Bridging [Gio98, Mar02a, OBO09]. Brief [Blõ06]. Bringing [GJAB95]. Broadband [Hul08]. Broadcast [ES05]. Broadcasting [ZZH^{+}12]. BROCA [Car95]. Broker [LB98]. Brouwer [Sch05c]. Brouwerian [Rat05]. Browsers [EMGB^{+}12]. Browsing [APJK09, INK09, Jun05a]. BT [MS11]. Budget [SJ13]. Build [dARGSB11]. Building [ATOFF98, BCFM05, BBC02, CJO^{+}13, CH07, DHO98, FC03, HBI98, Kom02, MOG^{+}10, TBL15, Vie03, WLKW11, Cam98]. built [BEH^{+}05]. Bus [CSAC^{+}15, OFCB08, ZDE14]. Business [BARR09, BST09, BFN05, CXB12, CPMVG13, DMCM14, GMK05, HA03, JV05, JJ12, LAHZ^{+}15, Ram01, SBAZ11, SMV08, SL05, WL13]. Business-driven [CXB12]. BYOD [HPE14]. Bytecode [MI05].

C [GCVRSPGP07, Mea97, MI07, Oli01, Puc10, Sch01c, dOLC^{+}07]. Cache [BS06, KAG00, WTA01]. Caching [HJZ07, KP95, NAK08]. CafeOBJ [OF13]. Calculating [SLN16]. Calculation [RB06, SP16]. Calculi [SAB99]. Calculus [Aic01, BBP95, Bar03, CSF99, CD13, Kri97, Poo03, PS04, CR00b]. Call [Ano07, RKJ16, TKSL05]. Call-Centre [TKSL05]. Camera [BSB09, CDR^{+}09, Lin09, TG10, VPF09]. Camera-Based [Lin09, VPF09]. Camera-Captured [BSB09]. CAMMD [OOHS06]. Campus [DBB13, PGSAP14, YKD^{+}08]. Can [CS04, GMP^{+}13, KSdV09, LM01]. Canetti [HK15]. Canonical [BB10a, Ish00]. Cantor [Her97]. Capabilities [GRHMM^{+}15, JFL^{+}13]. Capability [SS15]. Capacity [MRP14, YLW^{+}14]. Capital [CNRM03, CPFSdAS12, KL02, Liy02, Pre04, Vie03]. CAPTCHA [SSSS06]. Capture [TWH00]. Captured [BSB09]. Capturing [BRS00b, dFBNGS^{+}14, GMK05, MS10, PEPP08]. Card [BH14, Kim12]. Cardiac [ZHG^{+}06]. Cardiovascular [SSAB^{+}13]. Care [GK06, HB13, dIdSGZ10, JFL^{+}13]. Caring [BE11]. Case [AC05, AR01, BS12a, BB08b, BS03, BU13, Bic15, BDM15, BRS00b, BG00a, BG00b, Bor07, Bur05, CGLdMAC14, CM07, CMSE09, CBG04, DT09, DRA^{+}04, DAk13, DKD^{+}13, DL99, DGBM08, DBAB12, GB03, GJP^{+}12a, GRCGK14, HA10, HB00, HLK09b, JV11, KCK10, KW10, KP00, uRLFH^{+}13, LRR04, LNML03, LA03a, LULGFC13, MTB^{+}08, MRK^{+}98, MM03, Mat04, MT99, MRP14, MV15, MS11, MP06, PMRO08, PGSAP14, PGDD15, QZB^{+}00, RH10, Rah99, RF12, RAS15, SGLM16, SLD^{+}16, SA97, SZWdP14, SF00, Ste00, Str97, TWH00, TTB09, Træ08, VRGPSP05, VdR09, WL13, CM09]. Case-Based [BS03, GB03, LNML03]. Case-Study [WL13]. Cases [dFBNGS^{+}14, Dor95, LSX12, VdR09]. Catalog [VM13]. Catalogue [MRGF14]. Catalytic [NXSA12]. Catalytic-Like [NXSA12]. Categorical [Lin04c, SDÖ^{+}12]. Categories [GLS00]. Categorisation [ADS98]. Categorization [Yan05]. Category [WZZ^{+}09]. Catenation [Sal10]. CAUCE [TGLP10]. Cauchy [Nag06]. Causality [KL09b, Zou06]. Caused [Gün02]. CAVIAr [MP10]. CBR [Hef04]. CC [CGD^{+}12]. CC-LO [CGD^{+}12]. CCSL [RTJ01]. CDMA [CAJ06]. CDN [NAK08]. CE [Hul08]. Ceilings [Dun96]. Cell [Mea97, WZC07, Mar04]. Cells [MMP15]. Cellular [BvZH09, Cha05, MM99, Mar00b]. Center [PB14, TLJ11]. Centered [CIM14, KD02]. Centrality [Jun08]. Centralization [BCZ04]. centralized [MSA13]. Centre [TKSL05]. Century [KM01]. Ceramic [ACR11]. Cerebellar [JL09]. CERIF [LJLCR13]. Certificate [WMSH09]. Certificate-based [WMSH09]. Certificateless [HLC08]. Certified [NZCG05]. CETL [MC07]. Chain [ACR11, PPJ08, SKSP09]. Chainable [Ilj09]. Chains [Ric05]. Chaitin [CG96a, CN97, Gro00, ŞM96]. Challenge [CGD^{+}12, Mat02, RBB06, WB07]. Challenges [ABM^{+}06, BSPO11, BPS11, BAR06, BRO08, Buf01, CRC04, GPA08, Hop98, KA97, Lin11a, MR12, RBLR02, RSFMJ12, RA06, YSP09]. Change [ABAL09, DBBS08, Die10, SA14]. Changeable [JS16]. Changes [LLS05, NPB06]. Channel [ACA^{+}16, YLW^{+}14]. Channels [WLKW11]. Chaos [FFK04]. Chaotic [NWDX09]. Character [CH07, TJ15]. Characterisation [BRH^{+}08, SS09a, Zou06]. Characteristics [APJK09, GCC16, KWH03, LF98, SZZ95, ZZ95]. Characters [RR11]. Chasqui [NSFVH05]. Chat [GÁVCNC14, HLK09a, PLB14]. Chat-based [PLB14]. Chatrooms [PLB14]. Chats [CIM14]. Chauffeur [SHH10]. Chebyshev [BB98]. Checking [ALHM^{+}14, AH04, BRF^{+}09, CNQ04, CL07, DTG10, Dru13, EKP03, FGBS14, GMB11, Gle03, HNS07, Hor04, RSVR01, SMSdB05, SV08, TLR09, Win97b, Yeu04]. Checklist [dMTS^{+}14]. Checkpointing [LMRG14, SV05]. Checkpoints [RMGT09]. Chemical [TKD^{+}09, TWW07]. Chemistry [HMM00]. CHESS [LKHL09]. CHESS-64 [LKHL09]. Childhood [Faz06]. Children [DA13, GCL^{+}13, MBC13, PZLAS^{+}13, RSP^{+}14]. Chinese [SZWdP14, WBS12]. Chip [ACP06, FPSFCG07, KASN08, KU00, KKH12, MPPS95, NdMM12, PTO^{+}12]. Chip-Multiprocessors [KU00]. Choice [DS10, HNJ^{+}10, MR11, Pet12]. Choices [JJ12]. Choose [BDM15]. Chording [SHH10]. Choreographies [CVM11, RVC12]. Choreography [Arr07]. Christian [Len00]. Chronic [CCH06, GFT09]. Chronicle [PS09]. Chronicle-based [PS09]. Church [Ban97b, Lin04c]. CIAM [MROH08, PMRO08]. Cipher [LKHL09]. Ciphers [DGlS12]. Ciphertext [CS07, XMLC13]. Ciphertext-Policy [XMLC13]. Circuit [BCC^{+}06, LWC^{+}04]. Circuit-SAT [LWC^{+}04]. Circuital [WZC07]. Circuits [Bai05, PBTW07, TH99]. Circularly [Ilj09]. Circulation [SKSN07]. Cities [CMMP16, PGDD15, SC14]. Citizenship [She96]. City [KDKDN08, MCMMAP^{+}14, PGDD15, SCS13, VDBNR98]. Clarity [BE11]. Class [Bai12, MR05, Müh96, NuR05, NNT16, Pop98, RTJ01, SSGS10, Yeu04]. Classes [CP02, Fon00, GW03, Gol05, HK95]. Classical [BB10a, KO99]. Classification [ARQH14, BPHN06, BMGMF08, CWTT11, CLM10, FMB^{+}11, Gär99, HPC10, HFOB08, HW10, KK13a, LDO^{+}12, Ngu09, Pal15, PO11, RSP^{+}14, dPSZPVLR^{+}16, SE09, SKH14, TSE^{+}15, XHP^{+}09]. Classifier [Alm06]. Classifiers [LCC11, MR11, OJSB08]. Classify [MVMRULS12]. Classifying [Azz10, LN02]. ClassON [GRCGK14]. Classroom [BHC05, Car96, DBB13, HG11, PB04, TLS12, Viv96, ZABB07]. Classroom-Based [PB04]. Classrooms [MN96, UHÖD15]. Client [GHS06]. Client-Server [GHS06]. CLIM [Moo08]. Clinical [dFBNGS^{+}14, SSBS08]. Clock [PF15]. Clock-Skew-Based [PF15]. Clocked [CGFSHG09]. Clos [Ver08]. Closed [CVFN07, Sch10, Wei08]. Closure [HHY02, Kwo97]. Closure-Based [Kwo97]. Closures [Sal10]. Cloud [ASAIN14, CYL11, DHC11, FML13, FZAP13, IAS16, JBBH13, Kro13, LGMM^{+}13, LYLX15, LCDP15, LAHZ^{+}15, MSTW11, Ozd13, Pal15, PCS^{+}13, RSM^{+}13, RLMS13, Riz15, TKP11, Vle14, XNKG15, XMLC13, BS12a, CPFSdAS12, GPFL12, Kim12, RMFM12]. Cloud-Based [Vle14, Kro13]. Clouds [LdPK^{+}14, MSA13, THS11]. CLR [MI04, MI05]. Cluster [FSSPLG^{+}13, JP07, dOMdAL^{+}08, RTB13]. Cluster-Based [RTB13]. Clustering [AMC^{+}12, CA14, HT06, LALS08, LVV10, PCKJ11, QZ07, QF12, SDÖ^{+}12, SGS13, TNRGCP^{+}13, WC09]. Clusters [Hri02, Ist07]. CML [CB04, dOLC^{+}07]. Co [CAS^{+}13, DM07, Has02, Ilj09, Nd05, SNAF07, TLS12, TJ15, DBAB12, GL11a]. Co-Allocation [CAS^{+}13]. Co-AutoCAD [GL11a]. Co-Design [Nd05, SNAF07]. Co-Designing [TLS12]. Co-evolution [DM07]. Co-Lab [DBAB12]. Co-occurrence [TJ15]. Co-operative [Has02]. Co-r.e [Ilj09]. Coaching [FML13]. Coalgebraic [KAM03, RTJ01]. Coating [Pos98]. COBIT [CI12]. CoCharNet [TJ15]. COCV [KZ03]. Code [CCYK15, Dor07, FKO14, FBCMR15, Fra98, HSFE12, dSJPM14, KR03a, LdPK^{+}14, Mea97, NI03, Raj07]. Codes [ACA^{+}16, BSHI99, CT97, DLR97, Fen15, Gün96, GHNT97, IDS02, RMGT09, VSGP05]. Codifiable [AMVM01]. Coding [ACA^{+}16, CS05b, LLYC12, MS03]. Coequalisers [Pal05]. Coercing [Ban97b]. Coercion [DRS06]. Coexistence [CVFN07]. Coffein [Thi00]. Cognitive [Aïm98, Cañ08, GGP08b, Kar13, Kat05, SKHK14, WKSD^{+}11, ZTN^{+}15, vdV08]. Cognitive-dissonance [Aïm98]. Coherence [DF00]. Coincidence [SS09a]. Coinduction [Ber10]. Collaboration [BM13, CE11, GL11a, GFBR08, GlRBdSG11, GL11b, GGB^{+}08, HAI13, HA13, LWY09, MR08, MC07, NOP08, VMFO14]. Collaborations [HNJ^{+}10, LNH^{+}15]. Collaborative [AT13, ACR11, AB09, AGGH08, AFP^{+}13, BAZ14, BMG^{+}05, CGD^{+}12, CM09, CPHC11, CGP^{+}07a, DLL16, DDSS05, DDJ^{+}11, FSdRSS11, HMW08, HLNA^{+}12, HOPN11, HG11, IMR^{+}12, IRMK12, IBN^{+}11, Jun05a, Kah01, KOW01, KSdV09, KR11, LST14, Luk08, MPG13, MNDRF10, MRO05, PRBLAP^{+}13, PJO15, PSVOVI07, PMRO08, PCKJ11, RS11, RPR11, SSAB^{+}13, SBAZ11, SE09, SL96, SLJC08, SLPL14, Sut01, TRR06, TR10, TLS12, Tom01, VBP^{+}11, WLKW11, WSF08]. Collation [DH10a]. Collect [ZG05]. Collecting [GGS08]. Collection [CSY02b, CI05, CSZ07]. Collections [BMGMF08, LKMS08]. Collective [CAS^{+}13, JN08, Jun10a, KPV^{+}11, KD05, Nal10, NCH16]. Colleges [Len00]. Collisionful [BPSN97]. Color [CM11]. Colorability [RR06a, RR06b]. Coloring [Bod01]. Colour [CDR^{+}09]. Coloured [SBCD15]. Column [Güt12a, Güt12b, Güt12c, Güt12d, Güt12e, Güt12f, Güt13a, Güt13b, Güt13c, Güt13d, Güt13e, Güt13f, Güt14d, Güt14a, Güt14b, Güt14c, Güt15a, Güt15b, Güt15c, Güt15d, Güt15e, Güt16a, Güt16b, Mau94, Mau95a, Mau95b, Mau95c, Mau95d, Mau95e, Mau95f, Mau95g, Mau95h, Mau95i, Mau95j, Mau95k, Mau95l, Mau96b, Mau96c, Mau96d, Mau96e, Mau96f, Mau96g, Mau96h, Mau96i, Mau96j, Mau96k, Mau96l, Mau96m, Mau97a, Mau97b, Mau97c, Mau97d, Mau97e, Mau97f, Mau97g, Mau97h, Mau98a, Mau98b, Mau98c, Mau98d, Mau98e, Mau98f, Mau98g, Mau98h, Mau99a, Mau99h, Mau99i, Mau99b, Mau99c, Mau99d, Mau99e, Mau99f, Mau99g, Mau00a, Mau00b, Mau00c, Mau00d, Mau00e, Mau00f, Mau01a, Mau01b, Mau01c, Mau01d, Mau01e, Mau02a, Mau02b, Mau02c, Mau02d, Mau02e, Mau02f, Mau02g, Mau03b, Mau03c, Mau03d, Mau04a, Mau04b, Mau05, Mau06d]. Column [Mau06e, Mau06a, Mau06b, Mau06c, Mau07d, Mau07a, Mau07b, Mau07c, Mau08a, Mau08b, Mau08c, Mau09a, Mau09b, Mau09c, Mau09d, Mau10a, Mau10b, Mau10c, Mau10d, Mau11a, Mau11b, Mau11c, Mau11d]. Combating [VV12]. Combinational [AH04, TH99]. Combinations [PB05]. Combinatoric [Mar04]. Combinatorics [CSZ07]. Combinators [Lin04c]. Combined [DS03, XCJ13]. Combining [BS96, CL95, OJSB08, PLB14, RGPK15, SEK13]. Commerce [HGIPCPPM11, BS12b, TSDP07]. Common [PGT09, YYZ^{+}09, Lei08]. Communicating [BCG^{+}99, GV00, SH96]. Communication [Alh04, AT07, BE11, Bur05, Cam98, DM07, Epp04, FMS12, GFBR08, GlRBdSG11, Gün96, HM99, KKK16, KTJ05, KF10a, KF10b, LZGC09, LGZ01, MGM^{+}08, MTK97, OEK16, PKP08, RS02, SLT08, WLKW11]. Communications [GBCA12, Kil08, KD05, ND08]. Communicative [NN07]. Communities [dMBHR15, CMSE09, CBG04, DDS04, Dro04, KFK05, KJ10, KM06, Lin04a, MPL11, NW04, OAR^{+}14, Pre04, WF12, ZBKK12]. Community [AM11, FLF^{+}14, JK10b, LNHZ09, MNF^{+}13, SK08]. Community-Based [FLF^{+}14]. Commutativity [MSM07]. Comonads [UV05]. Compact [LW08]. Company [TTB13, BC11]. Comparative [BG98, DJJN09, dOMdAL^{+}08]. Compare [SGB^{+}13]. Comparing [ABB14, GOM^{+}13, GLD^{+}12, LNML03]. Comparison [AV07, BCG^{+}09, BKL12, GLSD11, HPC10, KBF^{+}11, Lav96, Mat04, MSSV14, OBO09, PV95, SGS13, XWGS09]. Compass [PKSR09]. Compatibility [Win97a]. Compatible [CD10]. Compensating [BFN05]. Compensation [DZ08]. Competence [Bec03, DSAFW07, HCWA03, HP15, RN03, Tar12]. Competence-based [HP15]. Competences [GRC15, MLHCGB16]. Competencies [LA03a, LA03b, RN03, dKR03]. Competency [JV11]. Competitive [AR04, Ano07]. Compilation [CBBT07, KR03b, Kwo97]. Compile [CPS07]. Compile-time [CPS07]. Compiler [BD00, Gle03, KZ03, ZG97, dSC06a, KZ03]. Compilers [ZPFG03, dSC06b]. Compiling [EGG^{+}01, Goo01, MABS05, Sch01c]. Complement [RK97]. Complementarity [Lip00]. Complementary [dCPUH^{+}07, RR11, PS97]. Complements [Vît05]. Complete [FKS^{+}04, GK13]. Completeness [Bul95, IK97, RR06a, ECHS10, Lan10]. Completions [BB10a]. Complex [ARS00, GNP05, LNH^{+}15, MRP14, SS09a, SKHK14, TKD^{+}09, TE06, ZG05]. Complexity [AHT09, Akr09, BHK10, BH02, Bos08a, BISZ08, BCR09, CFSC04, CP02, DDS10, DJJN09, GXC^{+}15, GKK^{+}02, Hem06, Jun01, KD05, LLYC12, RWZ09, vZG11]. Complexity-Grounded [GXC^{+}15]. Component [AFP^{+}13, BB04, CBR^{+}05, IPCVC12, LRS^{+}11, ONRV08, PMLL09, RP08, Sch02a, SdBM05]. Component-Based [LRS^{+}11, SdBM05, IPCVC12, PMLL09]. Components [APR05, Bar03, CCMP08, DR04, FKO14, FVIG12, FJP06, Meh02, POB11a, RdL08, WO09, ZAB^{+}08, dAO13]. Componentwise [Rex98]. Compose [PCLT15]. Composing [ASH11]. Composite [LWL10, OCW13]. Composition [AdO11, AF04, CSC08, CBR^{+}05, CDD^{+}07, CBO05, CCP11, FVIG12, LASL12, MYC14, MM06, PSVOVI07, TF09, XCJ13]. Compositional [AK05, KF10a, SMSdB05]. Compositionally [OF13]. Compositions [EMZB14]. Compound [BQBW03, vSO11]. Comprehension [CL95, RKH15]. Compression [CVK97, Cv99, Fen95a, Fen97, Gün96, Log04, Log06, QGT^{+}14, Wol99a, Wol99b, Wol00]. Compromise [MSHN06]. Computability [BHK10, Bos08b, Bra02, BD05, BISZ08, BCR09, CS00, CI05, CG09, CDF97, Ilj10, Pau13, Sko08, Spa08, TY09, YTM05]. Computable [Ilj09, JW13, LW08, Ret08b, Wei08, WG09, Wei10, MYT09]. Computably [Her97]. Computation [CPS07, DF99, LDO^{+}12, LO98, RTL05, Roj96, RA06, SI00, Ski97, DMMM95]. Computational [AGA12, Aur01, BCM12, BH02, DH10b, DMS05, GTGT10, HGS^{+}08, Jun01, KK06, LLH03, Lip00, Mac96, NH09, NCH16, PRBLAP^{+}13, RA06]. Computations [CJ98, DF05, FPLS03, FA06, HMM00, KKH12, KK10]. Computer [AWGS04, BMUF14, BDL^{+}06, BMG^{+}05, BRO08, BCDK97, CMS94, Cam98, CSW^{+}08, CJ07, Dom01, Dvo00, DSRR03, GPA08, KP01, KNLS00, LV95, LGZ01, LG08, MSC03, MNS^{+}12, MN96, PF15, PD99, Pop95, RB08, SHH10, SCS13, SAA08, SRR04, Vai00, VV06, VAS05, WLKW11, YLL^{+}07]. Computer-Based [Dvo00, DSRR03, KNLS00, SRR04, Cam98]. Computer-Human [AWGS04]. Computer-Mediated [LGZ01, WLKW11]. Computer-Supported [MN96, MSC03]. Computerized [Car95]. Computers [ACL95, BFMSP05, FMLNF07, Roj96, Svo95]. Computing [AV07, ABCP02, BZM^{+}10, BZ09, BHC05, BAR06, CW12, DH10a, DP99, Die10, FML13, FZAP13, GPFL12, GFT09, HWN02, HCK11, JK10b, KDGH09, KK10, Kuz04, LUR16, LKHL09, Lip10, LSG^{+}14, dIB13, LGP10, LS95, LAHZ^{+}15, NR12, NS06, NC04, PSVOVI07, PPJ04, RMFM12, SEK13, SS09b, SIIJ09, SLPL14, TKP11, Ted09, TGLP10, VRGPSP07, VMFO14, VO10, Wol99a, XMLC13, Zim01, von98]. Concatenation [DDS10]. Concept [APM04, CL95, CR04, CJ07, GSW04, HM00, HSD^{+}14, LSX12, dLMVGM13, Mau96a, PD04, PO04, RS02, RN03, SW13, YX10]. Concepts [GR02, GBHA12, MLHCGB16, TD96, Tom95]. Conceptual [AHPSCDK14, Arr07, BGBA10, dMBHR15, Bor07, CXB12, DT12, FGSW14, FL14, KS10, KBN14, Ma10, MSTW12, MS05, OP15, SLJC08, Ski00, Tha10, VO10]. Conceptualization [HLHD^{+}07, NSFVH05]. Concerns [CRMLN^{+}07, LRS^{+}11, URG^{+}13]. Concurrency [Gro09]. Concurrent [AC05, BRF^{+}09, FPLS03, dAFL07, SRI08]. Condition [Vît05]. Conditional [Cam98, CVPS95, Roj96]. Conditions [BH00, DRRGdP07, GWW05, dFCC07]. Conference [BHS^{+}06, Müh96, RMF^{+}98, TM01a, TM01b]. Confidence [SDÖ^{+}12]. Configurable [RMGCGCF08]. Configuration [BBIC13, BDM15, Die10, FG10]. Configure [Hel07]. Conflict [Gan04, HNJ^{+}10, PBB08, SW09]. Conflicts [HME^{+}06, PBB07, YYZ^{+}09]. Confluence [AT97]. Conformance [DTG10, LKZK10]. Confusion [HLK09a]. Congestion [HCH^{+}09]. Congestion-Avoidance [HCH^{+}09]. Conical [KNSN07]. Conjecture [vP05]. Conjunctions [POJB08]. Connected [Luk08]. Connecting [FTARR05a, KB06, Ois98]. Connection [KO99]. Connection-based [KO99]. Connectionist [RAC10]. Connections [CCS00, DMM07, Ior00]. Connectivity [Tat07]. Connector [Bai05]. Connex [MM96, ŞM96]. Cons [von98]. Consensus [HN07, HNJ^{+}10, Sob05, Zgr07]. Consensus-Based [Sob05]. Consent [RY09]. Consequence [VCB08]. Consequences [Tom03]. Consideration [CLCC10]. Considerations [LA03a, PT13, TJS^{+}13]. Considering [Bur05, RHM15]. Consistency [ALHM^{+}14, ASAIN14, CHPHV10, WTA01, Yeu04]. Consistent [SV05, Tru10]. Consolidation [Wur10]. Constant [CS07, Fen15, KR03a, Lan98, Ret08b]. Constants [Sal02, Sko08]. Constrained [CS10, CT04, Str97, NOGG^{+}13]. Constraint [ABPS95, BD06, Fal10, FH06, Ist07, PRW95, Rat00, RVW07]. Constraints [BR03, CVM11, CS10, DS03, Dru06, KBN14, LEC11, LWL10, PLBG13, RRB03]. Construct [RRR10]. Constructed [Her09, KPV^{+}11, ZD09]. Constructing [DT07, GLD^{+}12, RS05, Wat02, ZZ96]. Construction [AC05, AK05, CCHdCN08, GÁVCNC14, HSSM^{+}04, Ish00, LSK06, RMF^{+}98, RS00, Shu97, Thi00, THS11, VSGP05, YWD08, ZG97]. Constructions [CD13]. Constructive [Bar05, BB10a, Ber05, BB10b, Ber10, BY97, BH08, Cla05, CS05a, Ish97, Krä98, Leš09, Mac01, Mos05, Rat05, Ric05, SH06, Spi05, vP05]. Constructively [BB09, Sch05b].

Constructivist [MHLB12]. Constructivity [CI05, Şte96]. Constructor [GFO12]. Constructor-based [GFO12]. Consume [KM06]. Consumer [CBRH12, KKTT16]. Contained [DOS95]. Containing [Tru10]. Content [AMA^{+}14, ACB02, CWTT11, Fel01, HM01, JM15, LZ09, PLB14, SKL08, SJ13, SKH^{+}10, THS11, VAH07]. Content-Based [ACB02, HM01, JM15, PLB14]. Contents [CGLdMAC14, CRLNAR05]. Context [ADS98, BCA^{+}10, BH02, BHS^{+}06, BVV^{+}10, CGLdMAC14, CBNDR10, CLVM09, CLM10, CCP11, DMCM14, EMGB^{+}12, FWT11, FDR^{+}15, GMC08b, GO14, HCBB15, HBF10, HBFV13, HI99, Jun10b, Jun10c, LKT10, LSX12, LPSF10, LS10, LGP10, MH98, MOG^{+}10, Meh02, MNDRF10, OOHS06, Sat10, SHK10, SW04, SMP^{+}11, SKH14, TGLP10, VDLG10, VVM^{+}06]. Context-Aware [CCP11, FWT11, GO14, Jun10c, LPSF10, LS10, LGP10, OOHS06, VVM^{+}06, HCBB15, Sat10, TGLP10, VDLG10]. Context-Awareness [HBFV13]. Context-based [LSX12]. Context-Free [BH02, HI99, FDR^{+}15]. Context-Oriented [GMC08b]. Context-Sensitive [CLVM09, Meh02]. Contexts [DCS09, LZZK14, Mar07]. Contextual [ESG10, GFBR08, NBGS06]. Contextualized [BVV^{+}10, SBMD10]. Continued [Les95]. Continuity [Ber05, Ish97, SBAZ11, Sch05b, ZG05]. Continuous [LLS05, Ois98, SR00]. Contract [NSMBACBG12, RMMLBLGS09]. Contract-based [NSMBACBG12]. Contracts [AF04]. Contrapositive [Pet12]. Contribution [STFM12]. Contributions [BR07]. Control [BDPSNG97, BRS00b, BG00a, BG00b, Cap05, CS04, CJ98, EHEH05, HMW08, HB00, KKK^{+}14, KP00, MT99, MM07, MP09, NML09, NOGG^{+}13, Pau07, PS04, QZB^{+}00, RVC12, SU01, SF00, TJS^{+}13, TYSY09, TNM09, TWH00, XMLC13, dH00]. Control-Supervisory [TNM09]. Controlled [ABB14, CS09, DT09, MZ12, PRT^{+}08]. Controller [GRHMM^{+}15, Ste00]. Controllers [CA14, JL09, TSCY01]. Controlling [Tan08]. Convergence [ELFAR15]. Converging [ES03]. Conversation [CH07, SKSN07]. Conversational [DGBM08, PT09, SKSN07]. Conversion [LKMS08]. Conversions [BR03, TY09]. Converter [MUF03]. Convex [Mar98]. Cooking [OHYJ16]. Cooperation [BvdTV09, BH00, FP95, HHHX09, MKA11, MN96, PS97]. Cooperative [CLC09, IPCVC12, JR02, LGL09]. Coordinate [BBIC13]. Coordinated [BS11, FR04, ZDE14]. Coordinating [APR05, LWS11]. Coordination [BAPG03, SD97]. Coproduction [FFK04]. Copula [ND08]. Copyright [TSDP07]. Core [GSdSB16, KKH12, NDAM09, Par09, RMGCGCF08]. Cores [PTO^{+}12]. Coroutines [dRI04]. Corporate [KA97, MPF^{+}16, TBVRGLD15]. Corpus [GBHA12, RPCA15]. Correct [CP01, GPSV03, Pop95, ZG97]. Correcting [BSHI99, DLR97]. Correction [Kon02, PJN13]. Correctness [Ara97, Bel08, Bur05, FF07, Gle03, RK97, RSVR01]. Correlated [ND08]. Correlation [AGK^{+}10, LWC^{+}04, OMO10]. Correlator [dCPUH^{+}07]. Cost [ASH11, CCMP08, LKP11, PA12, SCW08, WLtHN14, dLH08]. Cost-Aware [WLtHN14]. Cost-Sensitive [LKP11]. Costs [DF05]. COTS [CCMP08]. Could [SGLM16]. Countable [Pet12]. Countermeasures [MGCGCG12]. Counting [dAFL07, Gup01, Lin03, LdL07, Sch05a]. Countries [GCG08, LM01]. Coupled [BAZ14]. Coupling [AO04].

Coupling-based [AO04]. Course [GGP08a, HR06, LB98, RAS15, WBS12]. Courses [AMUFVI09, BM13, BEPT14, CRB15, EACGFK13, GSPK08, GRCGK14, LMMRV^{+}15, OP15, Pel14, PCLT15, RGPK15, SS07, VBO08, dlFVPHB15]. Courseware [CDD^{+}07, GSZ15, MRK^{+}98, MH96, TR98]. Coverage [GR01, SCT09]. Covering [BCHM12, ECHS10]. Coverings [HN07]. CPU [SH10]. CPU-bound [SH10]. CQL [Buz06]. Craig [RG00]. Crawling [SR10, VdSdMC08]. Create [GMdMC12, HLHD^{+}07]. ‘Create-by-Reuse’ [HLHD^{+}07]. Created [Pel14]. Creating [AKM07, HAS^{+}07, IRMK12, Ret08a]. Creation [Ach06, ABCP02, BM03b, Kul05, NW04, VAH07]. Creativity [BFF11, DDJ^{+}11, MR08, SMP^{+}11, WP15]. Credible [LS95]. CREDO [CR04]. Crime [HFOB08, LLCN09]. Crisis [KDKDN08, LM01]. Criteria [CDBZ09, GR01, IO04, MTB^{+}08, SCT09, RLL^{+}10]. Criterion [NZM09, ZZ95]. Critical [GRHMM^{+}15, SBCD15]. CRMDatabase [TBVRGLD15]. Croatian [AR04]. Cropping [Fod06]. Cross [BC11, CS03, EGK^{+}12, Ern11, FBCMR15, LKT10, SHZ^{+}10, VLE12]. Cross-Border [LKT10]. Cross-company [BC11]. Cross-Disciplinary [CS03]. Cross-Industrial [Ern11]. Cross-Language [FBCMR15]. Crossing [MGMS12, Ste00]. CrossMDA [APDC08]. Crowd [GVT11]. CRWL [CP06]. Cryptanalysis [CGFSHG09, HK14]. Crypto [MPPS95]. Cryptographic [BDPSNG97, BNA15, JT05, Nd05, OS98, PZDH09, SZZ95, ZZ95]. Cryptographically [ZZ96]. Cryptography [CSW^{+}08, JR96, KK06, SG96]. Cryptosystem [NSNK05]. Cryptosystems [Aki09]. Crystal [HI00]. Crystallization [HNYO04]. CSC [MUF03]. CSCL [BKH^{+}13, VGBLGS^{+}08]. CSCW [DGBM08, PT09]. CSCWD [BSPO11, BPS11, YSP09]. CSP [SS03, Yeu04]. CTML [WSF08]. Cube [KD05]. Cubes [CK95]. CUBICA [MA13]. Cultural [TSDP07, vBK08]. Culture [KWH03, Kuh03, RvS12, Ste08]. Culture-Bound [KWH03]. Culture-sensitive [Kuh03]. Current [FBSEGP15, MR12]. Curricula [CDG14]. Curriculum [HMM00, TLS12]. Curve [AM96]. Curves [AHRH08, ACA^{+}16, DXZL07, GOF05]. Custom [NR08a]. Customer [CNRM03, SL10]. Customers [JMKT12]. customised [MB09]. Customization [HSSM^{+}04, LB98, SEK13, dAO13]. Cut [KHN99, YLW^{+}14]. CVEs [MGM^{+}08]. Cyber [CDCH09]. Cyberspace [WWD15, vB96]. Cycle [BM97, HNYO04]. Cycles [BI08]. Cyclic [dAFL07, JS16, Lin03, LdL07]. Cyclical [MUF03]. Cyprus [UHÖD15].

D [EK00, CLCC10, HFOB08, IRMK12, uRLFH^{+}13, Mar04, MGT14, Pel14, RPR11, SVFR15, SHK10, SE09, SBS15, TG10, VCB08]. D0L [Hon99]. DAA [DR10]. Dagstuhl [BJ05a]. Dagstuhl-Seminar [BJ05a]. Dale [KD11]. Dangers [KB06]. Danish [BARR09]. Dashboard [PJ15]. Data [Ada06, ARG05, ARFT05, AM96, AAGU97, BCM08, Bor07, BM96, BBP08, CSFFM12, CSAC^{+}15, CMMP16, CAR08, CCS00, DDJ^{+}11, Dru12, EK00, FKB^{+}15, Faz06, FTARR05a, FTARR05b, GKZ05, GM05, GOF05, GSS99, Gir05, GCC16, GC14, Gün96, HMA^{+}05, IAS16, JFZ09, KS05, KDGH09, KL09b, LSR10, LAM12, LDSG09, LLS05, LYLX15, LCDP15, Log04, Log06, dIdSGZ10, MBA12, MN14, MHA^{+}15, MNL13, MPPS95, MUF03, NHH06, NDAM09, NCH16, PRBLAP^{+}13, PD10, PCLT15, PS12, RHM15, RC07, RRM^{+}12, RA06, RJB10, SA10, SVK^{+}15, SBCJ03, SHZ^{+}10, SDLM14, STBFM09, SS15, SVV10, Suz06, TFMDM10, TL11, Tra13, VAS05, VGAPGS^{+}15, VBB13, VTHM16, WC09, WKXL05, XNKG15, ZSK09, Zim01, dH04, LKB^{+}02, JCB15, LAHZ^{+}15, TBVRGLD15, VPB15, YMA15]. Data-Dependent [RC07]. Data-Intensive [BCM08]. Data-on-Tag [dIdSGZ10]. Data-oriented [Ada06]. Database [ACB02, CCP^{+}07, GSW97, HKTV06, HL09, Heg10, Köh09, MBA12, PRAT09, SW10, Was98a, Was98b, YJY14]. Databases [AK09, ASAIN14, BDPSNG97, Buz06, FPT10, HPC10, HT13, Kul05, LHC^{+}13, PO04, Sch09a]. Dataflow [KAG00, KF10a]. Datasets [RMZ15, SDÖ^{+}12]. Days [SA14, ZDI10]. DCM [PKSR09]. DDBSs [AAM14]. DEA [ASH11]. Deadlock [PFS07]. Deal [CHPHV10]. Dealing [dLMVGM13, ODSO11]. Debugging [EGDG09, OCB^{+}10, dlELR09]. Decades [Lin11b]. Decentralized [BBdOR14]. Decidable [HI99]. Deciding [PTL^{+}09]. Decision [AMBP04, AR01, CVSM11, CM09, CCH06, CSF99, CW12, Dvo97, FTARR05a, GM05, GVT11, Hon95, KK10, LEJ^{+}08, MLX10, MDY10, Pop05, RSP^{+}14, RLL^{+}10, SA10, SLD^{+}16, TB16, TNRGCP^{+}13, XLMR10, YX10]. Decision-making [GVT11]. Decision-Related [KK10]. Decision-support [SLD^{+}16]. Decisions [DLL14]. Declarative [CGP07b, KD01]. Decloning [JS16]. Decodability [Kon02]. Decoders [GHNT97]. Decomposition [Ito02, MK99, Rud04, YS05]. Decompositions [Sal10]. Decoupled [VC13]. Decreasing [PPG95]. Dedicated [CS00, KP01, Lav96, ZHG^{+}06]. Dedication [Lom07]. DEDS [Cap05]. Deduction [BFF99, CM07, HMSR99, HNS07]. Defeasible [PGT09]. Defence [CCYK15]. Define [dMBHR15]. Defined [Hon95, IDS02, STW09, PRB^{+}11]. Defining [BST09, DGBM08, KWH03, KAM03, MM06, MHLB12, PLBG13, RGP10, dOBGH^{+}14]. Definition [ASHT^{+}16, BMa11, CVFN07, EGG^{+}01, Mos05, Rad14, YWD08]. Definitions [TBS08]. Deformation [BS11]. Degree [GÁVCNC14]. Degrees [FK16, HAS^{+}07]. Delay [Kon02]. Delivering [GDW10, LS10, MH96]. Delivery [LST14, NZCG05, PSS^{+}13, QZYL11, SW04]. Demand [DL99, INK09, KÜ10, May02]. Demand-Driven [DL99, INK09]. Demands [Fel01]. Demise [Odl94]. Democracy [SM02a]. Denial [Che09, LM03a]. Denial-of-Service [Che09]. Denotational [TBS08]. Densest [AC07]. Density [RMZ15]. Denumerable [Cre09]. Deontic [KJ10]. Dependability [EK99, MT99, MSF99]. Dependence [SNAF07]. Dependencies [CCM09, HL09, Pop05, Sch05a, Zgr07]. Dependency [CL08]. Dependent [OFCB08, RC07]. Dependently [Xi03]. Depletion [GHNT97]. Deployment [CGPAP13, DH10a, HA10, LWY11]. Derick [Jür10, JMSY10, Sal10]. Derivation [CKdL08, GPSV03]. Derivations [Was98a]. Deriving [HN07, MCC13]. Describe [AHPSCDK14]. Described [Świ07]. Description [Ad03, Ano99, Cer97, KTL^{+}11, MMP15, MS05, PR06, QZB^{+}00]. Descriptional [GKK^{+}02, vZG11]. Descriptions [APR05, CDD^{+}03, GLSD11]. Design [ACR11, ARS16, Aic01, AHPSCDK14, ACM16, ARS^{+}08, AS07, Ban96, BHH^{+}06, BHRS03, BGP08, BCC^{+}06, BvdTV09, Bör02, BJMBA15, BTD^{+}07, BJ05a, BQV14, CVM11, CMM01, CFMP15, CLC09, DM10, EHEH05, FZAP13, FP05, Fro02, GLCV08, Gio98, GSdSB16, GGMM^{+}13, GSPK08, GJAB95, HL96, Hei96, HLHD^{+}07, JST11, JJ12, JGW11, KCK10, KWH03, KD02, Köh09, LN02, Mai05, MFG13, Mat02, MPPS95, MOG^{+}10, MGPB07, MSF99, Nd05, NdMM08, NKS^{+}09, OCRPdlMG07, PK98, QC12, RMF^{+}98, RS01a, RS01b, Sch06, SH96, She05, SKHK14, SJ13, TIL08, TJS^{+}13, TCS^{+}03, TFG06, TT98, Thi00, TAL08, TBVRGLD15, TLJ11, UDC97, VLE12, WLLL09, WLKW11, WBS12, WKSD^{+}11, ZSK09, dSC06a, vdV08, SNAF07, DGBM08]. Design-Oriented [VLE12]. Designing [BRO08, CGP^{+}07a, HBT12, HVCA12, LM10, NSL96, PTO^{+}12, PMRO08, STBFM09, TLS12]. Designs [EKP03, NdMM12]. Desynchronisation [BCM12]. Detail [UCM13]. Detailed [ARQH14]. Detailing [LdSM08]. Detect [LM03a, PLB14, TH99]. Detecting [JFZ09, PHPP06, PS12, YYZ^{+}09, ZDI10]. Detection [BDL^{+}06, BAML07, BMMM14, Che09, CCYK15, CAJ06, Dru12, DGL03, FBCMR15, GPLV13, HKK13, JT05, KSK^{+}16, KASN08, KD11, KMN16, LKP11, LMMRV^{+}15, RKJ16, RK97, SW13, Sto99, Tin16, VV12]. Detectors [SW13]. Determinant [DF99]. Determination [OT16, UHÖD15]. Determine [vD05]. Determining [Lan98, MX05, OEK16, Wie08, Zgr07]. Determinism [Alh04, Gup01]. Deterministic [And97, BC16, CL97, CFJS15, Lan10, RR06b, Wat02]. Develop [JBBH13, VBVNHdDSL12, VTGA13, ZGC^{+}08]. Developer [GlRBdSG11, JL16]. Developing [BAZ14, BPHN06, BdGFMT14, CSA10, DB03, GVRT^{+}10, GSP04, GHO10, HPE14, JK10b, LGGGC14, MBA12, MBC13, OEK16, Ozd13, RFMLP10, SBAZ11, SKHK14]. Development [ACR11, AHSN01, AGGH08, AVA08, AFP^{+}13, BP09, BBL13, BC11, BM97, Bör02, BQV14, Cañ08, CPSAGPGC12, DSAFW07, Fro02, FJP06, GMHGRG^{+}13, GPA08, GNP05, GMC^{+}08a, GHM04, HLP^{+}13, HVCA12, HH03, HP15, Hul08, Kar13, Kie05a, KY10, Kul07, LRR04, Lar01, LCHD12, LGZ01, LASL12, LNHZ09, LULGFC13, MPM12, MH02, MMM12b, MM12, MROH08, MKI^{+}12, Müh96, PJ15, PLG^{+}08, PTNMC08, PRCCS13, SdBC13, SFP12, Sha11, SFVFMN04, SA09, SKHK14, SF00, SBCD15, TGLP10, TPC^{+}12, UHÖD15, WKS^{+}99, dCVM12, Bec03]. Developments [BG01, DGK^{+}99, FMLNF07, Pob11b, SBG^{+}12]. Device [LPSF10]. Devices [BCHM12, CGLdMAC14, DH10a, GP10, HLP^{+}13, HLNA^{+}12, HG11, OOHS06, Pos98, QGT^{+}14, RMGCGCF08]. DFA [BDGW96]. Diagnosability [PS09]. Diagnoses [HGMT08]. Diagnosis [CLCC10, FMA^{+}05, KASN08, RSP^{+}14, SIIJ09]. Diagram [AFP^{+}13, CMM01, CT16, MCC13]. Diagram-based [AFP^{+}13]. Diagrams [Are02, Dvo97, Gan04, MM06, Pop05]. Dialog [Træ08]. Dialogs [KM07]. Dialogue [RPCA15]. Diameter [SLN16]. Diamodl [Træ08]. Didactic [PT13]. Difference [Ars97, vZG11]. Differences [Jéz95]. Different [Bjø01, BRAS^{+}12, GOM^{+}13, Her02, KBF^{+}11, KJL09, MHLB12, RdKO11, WLKW11]. Differentiable [ML98]. Differential [APNA12, CG09, Fen95a, JS16, Wan95]. Differently [VJ09]. Difficult [LM03b]. Difficulties [DIKL14]. Difficulty [ZZ96]. Diffie [HLCL11]. Digit [NK95, Fen95b]. Digital [BRH^{+}08, BBM12, CFMP15, CGP^{+}07a, DXZL07, DKM04, Fel01, HBT12, HLNA^{+}12, HT01, IS10, IAS16, KKM08, Kul07, LKMS08, ML95, NVB12, NSL96, QGT^{+}14, SdOB09, Sch09b, SV05, TSDP07, Yon11]. dimensional [GOF05, Sch10]. Dimensions [KKK16, ML02]. Diminishing [HLK09a]. Direct [BR03, Sch05c]. Directed [Mar95, NLLJ12, Ozd13]. Directory [CS09, LVV10]. Dirichlet [BY97]. Disabled [WCH14]. Disambiguation [MNL13].

Disaster [DH10a]. Disciplinary [CS03]. Discipline [Nav09]. DisCo [AKP01]. Disconnected [Luk08]. Discontinuous [Ang98]. Discourse [Sut01]. Discovering [CBRH12, ELS04, LT13, Suz06, TCK^{+}01, TGEM07, ZD09]. Discovery [AdI16, AK09, BQBW03, CAD^{+}06, LKK08, LWH09, May02, PTL11, ZSG14]. Discrete [ACAMM15, DZBB^{+}12, DXZL07, PS09, Smy00, Tab07]. Discretionary [BDPSNG97]. Discriminative [LHK^{+}13]. Discussion [GÁVCNC14, SCS13, vdV08]. Diseases [CCH06, GFT09, SSAB^{+}13]. Disentangling [TBS08]. Disjunctive [Her96, Sta02]. Disk [SV05]. Dispatch [DB12]. Display [VDSF98]. Displaying [IGS08]. Displays [TJS^{+}13]. Disruptive [HCPdASY14]. dissonance [Aïm98]. Distance [BKH^{+}13, CGFSHG09, CRB15, DGK^{+}99, GD14, HKS96, HOS96, Has02, HL96, Hol96, KMR96, PPG95, RKH15, Rex98, SS07]. Distances [CSY02a, CSY02a]. Distortion [CS05b]. Distributed [ACAMM15, AAGU97, AR95, Ara97, BAZ14, BEH08, BMV12, BBC02, CT04, CE06, Che09, EGK^{+}12, FPLS03, FR04, FHJ^{+}99, GWG96, HVCA12, JCB15, Kap95, KP95, Kri99, Kul05, LCHD12, LHC^{+}13, LAAPVGMM15, MMM12b, MM12, PLBG13, Sch02a, SL09, SSV02, SLPS98, UDC97, UFF12, VPB15]. Distribution [BP09, BCZ04, NAK08, Nag06, PLBG13, PD99, Zim01]. Distributions [HT06, PMLL09]. Diversity [BDN05, BM03b]. Divide [BBM12]. Division [DF99, Dor95, Fen95b, GS12]. DNP [Gaß10]. Do [LM94b, LGZ01, Sch02c, Ver10]. Docs [FCM^{+}12]. Docs4Learning [FCM^{+}12]. Document [BMGMF08, BSB09, BNCGD^{+}11, BVG08, CWTT11, CL08, LDO^{+}12, Lin08a, Lin08b, Lin11a, Lin11b, LKMS08, PTL^{+}09, SUKG13, SFVFMN04, SH11, VL14, dSLMW08].

Document-Oriented [SFVFMN04].

Documents

[AGA12, BMMM14, HM00, JST11, Kol05, LdSM08, Lin09, NML09, RdKO11, RGHH97, STW09, Ski97, WD02, dCH11]. Does [SA14, IKC14]. Domain [CW09, CHH16, ESM08, FRD14, GMHGRG^{+}13, KDKB07, Kom02, KHG10, LRS^{+}11, MB09, MMS08, MG14, RY09, Sar05, SEK13, Træ08, VGAPGS^{+}15, VSML03, WSF08]. Domain-customised [MB09]. Domain-Oriented [SEK13]. Domain-Specific [ESM08, LRS^{+}11, MG14]. Domains [DGBM08, FH06, LA03a]. Domatic [RR06a, RR06b]. Dominant [NL10]. Dortmund [TD96]. Dots [KB06]. Douglas [CI05]. Down [JMEL10, HA03]. Drag [HFIBJ13, MPG13, WHD04]. Drag-and-Drop [WHD04]. Draw [BR05]. Drift [SW13]. Drive [KKK^{+}14]. Driven [Alm06, BdGFMT14, CJH12, CKdL08, DL99, EK00, FLF^{+}14, GLCV08, GO14, HPE14, INK09, KCKL10, KF10a, KF10b, LRI03, LdPK^{+}14, LPSF10, MYC14, MR08, MMS11, PRCCS13, SMGMT09, SFP12, SMV08, SJ13, SVV10, VAPM12, dOBGH^{+}14, APDC08, BP09, CXB12, GLD^{+}12, MP10, TGLP10]. Driver [PS12]. Drives [Hul08]. Drop [HFIBJ13, WHD04]. Dropout [CRB15]. DS [CAJ06, MP09]. DS/CDMA [CAJ06]. DSP [RSW04]. DSR [JP07]. DT0L [Hon01]. Dual [FDC^{+}13, JW13]. Dual-Modal [FDC^{+}13]. Duplicate [BAML07]. Duquenne [Kuz04]. Duration [CSF99, LGZ01]. during [BS11, GCL^{+}13, HSD^{+}14, LTY^{+}16]. DVB [LZ09]. DVB/GPRS [LZ09]. DWCMM [SS15]. Dynamic [AY00, AdO11, AK09, BS08, BCG^{+}14, Buz06, CBF05, CCM09, CBR^{+}05, CT16, CW00, DGK^{+}99, GM05, HHHX09, JK10b, LWS11, LLCN09, LA03a, LLS05, LSG^{+}14, LLZ16, Mau03a, MP09, PRCRARLN10, PTNMC08, PPJ08, SLK11, SZZM10, SCW08, SCLM03, XNKG15, ZSK09]. Dynamical [CJ98, Ois98]. Dynamically [GCVRSPGP07]. Dynamics [BPC04, Dit02, LCC^{+}12]. Dyslexic [KM13].

E-Assessment [HMHGR15]. e [?]aulaSancho:2005:SWT. e-Barter [BCNR07]. e-Commerce [HGIPCPPM11, TSDP07, BS12b]. e-Consent [RY09]. e-Document [VL14]. e-Documents [NML09]. E-Goods [NZCG05]. e-Government [GR14, RC10, SVV10, VL14, SVV10]. E-Health [MHA^{+}15]. E-Learning [QL13, MPF^{+}16, PGSAP14, ATGP09, AMA^{+}14, ASHT^{+}16, FMLNF07, GGP08a, LT13, LM03b, MPG13, dPPRRGSSPP15, PB04, Ret08a, ES03, ELS04, SMFM05, Hel07, HAI13, MNS^{+}12, QL12, SW04, Yon11]. E-LOTOS [CM00]. ‘E-Okul’ [Tüf13]. e-Passports [NML09]. E-speranto [JST11]. Early [ALHM^{+}14, FWS^{+}11]. Easing [GHHE^{+}08]. Easy [Aur01]. Eclectic [Mac01]. Eclipse [Træ08]. eco [CVPS95]. eco-grammar [CVPS95]. Economic [DB12, FHH08, ND08]. Economics [GÁVCNC14]. Economy [Joh01, LA03b]. Ecosystem [Kil08]. Edge [HOS96]. Edge-Flipping [HOS96]. Edit [CGFSHG09]. Editor [Mar07, Car96, GGMM^{+}13, Güt12a, Güt12b, Güt12c, Güt12d, Güt12e, Güt12f, Güt13a, Güt13b, Güt13c, Güt13d, Güt13e, Güt13f, Güt14d, Güt14a, Güt14b, Güt14c, Güt15a, Güt15b, Güt15c, Güt15d, Güt15e, Güt16a, Güt16b, Mau94, Mau95a, Mau95b, Mau95c, Mau95d, Mau95e, Mau95f, Mau95g, Mau95h, Mau95i, Mau95j, Mau95k, Mau95l, Mau96b, Mau96c, Mau96d, Mau96e, Mau96f, Mau96g, Mau96h, Mau96i, Mau96j, Mau96k, Mau96l, Mau96m, Mau97a, Mau97b, Mau97c, Mau97d, Mau97e, Mau97f, Mau97g, Mau97h, Mau98a, Mau98b, Mau98c, Mau98d, Mau98e, Mau98f, Mau98g, Mau98h, Mau99a, Mau99h, Mau99i, Mau99b, Mau99c, Mau99d, Mau99e, Mau99f, Mau99g, Mau00a, Mau00b, Mau00c, Mau00d, Mau00e, Mau00f, Mau01a, Mau01b, Mau01c, Mau01d, Mau01e, Mau02a, Mau02b, Mau02c, Mau02d, Mau02e, Mau02f, Mau02g, Mau03b, Mau03c, Mau03d, Mau04a]. Editor [Mau04b, Mau05, Mau06d, Mau06e, Mau06a, Mau06b, Mau06c, Mau07d, Mau07a, Mau07b, Mau07c, Mau08a, Mau08b, Mau08c, Mau09a, Mau09b, Mau09c, Mau09d, Mau10a, Mau10b, Mau10c, Mau10d, Mau11a, Mau11b, Mau11c, Mau11d]. Editorial [Ban07, Pre12]. Edmodo [Bic15]. Education [BU13, BRAS^{+}12, BBM12, BKH^{+}13, BFMSP05, CRB15, CR12, DD13, DGK^{+}99, Dom01, FMLNF07, Flo04, FPS^{+}12, GPCPL12, GPVILN13, HBT12, HL96, Hol96, HR06, IKC14, KM13, KGK12, LHZS12, LUR16, LB98, LZZK14, LMA^{+}14, MHLB12, OT16, PCS^{+}13, PMAM14, PS06, PCLT15, RB08, RSM^{+}13, Riz15, SVFR15, Ste08, SLPS98, TE06, UT16, VV06, VRGPSP05]. Educational [ABFJ06, BGP08, BMG^{+}05, BE98, BQV14, CRMLN^{+}07, Car98a, Car98b, CFMP15, CGVRGPSP07, CSA10, DA13, DKD^{+}13, FML13, GPL13, GLD^{+}12, Gio98, GJP^{+}12b, HL96, LLSA13, LZZK14, Mar07, MOMGSRFM07, MFG13, MH98, MP10, MHH98, MCMMAP^{+}14, OEK16, PZLAS^{+}13, PRCRARLN10, PvW16, PCLT15, RCGBS13, RGRR15, SdOB09, SBG^{+}12, vZdlH12]. Educational-driven [GLD^{+}12]. Educative [GMdMC12]. Educators [AHPSCDK14, VGBLGS^{+}08]. Eduquito [SdOB09]. edX [SGLM16]. Effect [AWGS04, HOPN11, Tin16]. Effective [BB10a, Bos09, CG09, HA03, LWG14, LZZK14, Sch09c, SCW08, YTM05]. Effectiveness [ASH11, BCG98, Gio98, IK97, MRK^{+}98, Ozd13]. Effects [BM03a, BNCGD^{+}11, DB12, GCL^{+}13, WLKW11]. Efficacy [RF12, SBS15]. Efficiency [ASH11]. Efficient [AGO^{+}13, AR95, ACB02, CT08b, DMS05, EMZB14, FKB^{+}15, Fon00, Fon01, FKS^{+}04, HJZ07, HCK11, KC08, KHN99, LLLL99, LCDP15, LSV06, NAK08, Nd05, NSNK05, PJRC04, dCPUH^{+}07, PO11, QZ07, QF12, RKJ16, RSVR01, SUKG13, STVT07, WC09, XMLC13, YWD08, dSLMW08, dL03]. Efficiently [THS11]. Effort [GS12, MSC03, TR98, Til01]. EFL [KWC01, UHÖD15]. Efremovic [Vît05]. Egos [vB96]. eGovernment [SR10]. EHW [DM07]. Eigensystem [CW09]. Eigenvalues [Pav95]. Elderly [GVRT^{+}10]. eLearning [AMC^{+}12, BFMSP05, Flo04, GDW10, HCWA03, UP04]. Election [FP06, MS00]. Elections [LTY^{+}16, LCC^{+}12]. Electronic [AG06, BG98, FZ00, Goe95, HMM00, MS94, MT99, Müh96, MN96, NSMBACBG12, RMF^{+}98, Sch96, TEK08]. Element [FF04, GK13]. Elementary [Tez13, WG09, DMMM95]. Elements [Hav05, MC00, Sha11, TKT07]. Elevation [QGT^{+}14]. ELG [Ret08a]. Eligibilities [Pau13]. Eliminating [BI08, FA06]. Elliptic [LO98]. ELLMTAS [UHÖD15]. Embedded [Ad03, BRR99, CFF^{+}13, GHM04, HLS15, JT05, KCKL10, RP08, TTB09]. Embedding [CGD^{+}12, CD10, MCMMAP^{+}14]. Embodied [PT09]. Emergency [SC14]. Emerging [FMB^{+}11, KM06, Kul07]. EMERGO [SKHK14]. EMI [dIVG^{+}06]. Emotion [LGMM^{+}13]. Emotion-based [LGMM^{+}13]. Emotions [AMYH14, CPCLSAGC11, FDC^{+}13, LMMRV^{+}15, PZLAS^{+}13]. EmotionsOnto [LGGGC14]. Empathy [Pre04]. Empirical [AA16b, AS14, CPHC11, FP12, KW10, LASL12, LLZ16, STFM12, VDSF98]. Employee [LA03a]. Employing [LWY09, PK98]. Employment [LLM02]. Empowerment [CGD^{+}12]. Enabled [BBIC13, CWT^{+}15, Die10, LVV10, PRAT09, DR10, RLMS13]. Enabling [BZM^{+}10, GR14, GVT11, LAHZ^{+}15, NOGG^{+}13, PRCRARLN10, RC10, dIVG^{+}06, vSO11]. Enclosing [Neh98, Rih98]. Enclosure [ML98]. Encoders [GHNT97]. Encryption [AP09, HLC08, NZCG05, XNKG15, ZQQ15]. Encyclopedia [Mül02]. End [GHHA10, JL16, LASL12, RvS12, SKH14, SKH12]. End-to-End [SKH12]. End-User [GHHA10, LASL12]. End-users [RvS12]. Ended [HAI13]. Endocrine [Koó06]. Ends [ZG97]. Energy [AGO^{+}13, HJZ07, JFL^{+}13, KCK10, LZM04, NOGG^{+}13, QZ07, RLL^{+}10, YJY14]. Energy-Based [LZM04]. Energy-constrained [NOGG^{+}13]. Energy-Efficient [QZ07]. Enforcement [HHHX09, LF16]. Engagement [LNHZ09, Pel14]. Engine [GRS08, KB06, QC12]. Engineering [AGMT10, AFK01, AFP04, Bjø01, BG00b, CIM14, CDG14, CPCLSAGC11, CRC04, DSAFW07, FKO14, HAS^{+}07, HSR10, HSD^{+}14, IAS16, JBBH13, KPV^{+}11, Lin08a, Lin08b, Lin11a, LULGFC13, Mat02, MS10, Nal10, Nav09, OL08, PCLT15, RAS15, RGHH97, SEK13, TKD^{+}09, Ted09, TE06, UFF12, VAPM12, VK03, WKSD^{+}11, dOBGH^{+}14]. Engines [BCG98, FMT^{+}15, GWG96, MB09]. English [ARS16, BB08a, MV15, RAS15, VKW15]. Enhance [DIKL14]. Enhanced [AdI16, AMA^{+}14, KJZJ08, LZZK14, Ma10, dLMVGM13, MHA^{+}15, PA12, RJB10, VLE12, QC12]. Enhancement [HJZ07, HG11, LD06]. Enhancements [FKO14, Güt08]. Enhancing [BSP^{+}13, BKH^{+}13, HGS^{+}08, KPV^{+}11, MLHCGB16, SBS15, WCH14]. Enough [ODSO11]. Enriched [GLS00]. Enriching [GBHA12]. Ensemble [KLT13, LHK^{+}13, Tra13, ZC09, QL13].

Ensembles [SW13]. Ensure [Gle03]. Ensuring [TT00]. Enter [Rus07]. Enterprise [AUN04, CPHC11, HT06, JV05, LGES11, MRP14, MS11, VMA14, vKL04]. Enterprises [AR04, PCC14, SPRP09]. Entertainment [Bür08, LNHZ09]. Entity [JM15, THJ16, VO10]. Entrainment [KM07]. Entrepreneurship [FPS^{+}12]. Entropy [JM10, SESMT10, Sim07]. Enum [BHH^{+}06]. Enum-Based [BHH^{+}06]. Enumerable [Ars97]. Enumeration [FK16, FDR^{+}15]. Enviroments [SM02b]. Environment [AMR^{+}14, BRW03, Car95, CH07, CYL11, CTM10, CPR01, CGPAP13, DLL16, DZBB^{+}12, DDJ^{+}11, DOOJ95, ESB04, FMA^{+}05, GRHMM^{+}15, HCK11, IRMK12, KKK16, KOW01, KWC01, Mau96a, PANS13, PGDD15, RMFM12, Ret08a, RSW04, RSM^{+}13, RLMS13, RRM^{+}12, Ruf01, SdOB09, SSGS10, SLPS98, Viv96, VDBNR98, dBdd04, PS97, WHD04]. Environmental [TM00]. Environments [AB09, BZM^{+}10, BBdOR14, CJH12, CGP^{+}07a, DBAB12, FML13, FDC^{+}13, FWS^{+}11, GHHE^{+}08, GPCAC11, GPL13, GMdMC12, HM99, HMW08, HME^{+}06, HAI13, IMR^{+}12, JGL08, KÜ10, KKTZ09, KDGH09, Kro13, KJKS14, uRLFH^{+}13, LKHL09, LGMM^{+}13, LAAPVGMM15, MFG13, MGT14, NOGG^{+}13, QZYL11, QFB^{+}14, RS11, RPR11, Riz15, RP98, SVK^{+}15, SvRS14, SA03, SSBS08, TGLP10, Tom01, VLE12, VVM^{+}06, WSF08]. Epidemia [ASAIN14]. Epilepsy [RSP^{+}14]. Epistemic [MJ13]. ePortfolios [LAAPVGMM15]. EQF [GLSD11]. EQF-based [GLSD11]. Equation [Jéz95]. Equational [RG00, dR05]. Equations [DDS10, NO98, Pop98, Rud04]. Equilibria [KK10, Pau10]. Equipment [CWT^{+}15]. Equity [KKTT16]. Equivalence [AH04, BRF^{+}09, Hon01, Sch05c]. Equivalences [dFER06]. Equivalent [CR07]. Equivalents [Ber05]. Era [Die96].

Erasure [VSGP05]. EREW [Lep95]. Ergonomics [Cañ08, vdV08]. Error [ATV98, BSHI99, DLR97, FR04, Jéz95, Kon02, Krä98, QQ11]. Error-Correcting [BSHI99, DLR97]. Error-Correction [Kon02]. Errors [ACL95, BB98]. ESFP [TT00]. Establish [AdGCD^{+}15]. Establishing [SSS12]. Estate [Tra13]. Estimating [GH08, OPP09, PMLL09]. Estimation [ACL95, CC08a, DRRGdP07, Fen95b, HYC^{+}05, LD06, Nag06, PA12, PKSR09, SHK10, Świ07, TR98, dlFVPHB15]. eTandem [WBS12]. Etiquette [Pre04]. ETOL [Fer96]. EU4ALL [BRAS^{+}12]. Europe [BBM12]. European [FPS^{+}12, Has02, Tom03]. eval [Nør10]. Evaluate [BCCH11, PZLAS^{+}13, PvW16]. Evaluates [FDC^{+}13]. Evaluating [ASAAASJ16, GWW05, Gio98, MRK^{+}98, MH98, PJO15, PKP08, Reb96, SSSS10, WP15]. Evaluation [Ach06, AHSN01, BB98, BdS13, BG98, BM05, CPLPW15, Cañ08, CPRT05, CGPAP13, DAk13, DL99, EK99, EHEH05, FPSFCG07, GGPTdP11, GOM^{+}13, GSdSB16, GGP08b, HTHW12, Hen98, JFL^{+}13, JGM^{+}13, KHN99, KTL^{+}11, KZ08, Kro13, LMMRV^{+}15, LCC11, LL12, MN14, MBC12, MOG^{+}10, MHH98, NSMBACBG12, OP15, PMAM14, QQ11, QL12, RdKO11, RP98, RMZ15, SAKAM11, SKHK14, SKP08, TWH00, TAL08, TEK08, Tüf13, WPL98, dSC05, dAO13, CEK15]. Evaluations [VDSF98]. Event [BCM12, DZBB^{+}12, GO14, LRI03, Tab07, PS09]. Event-Driven [GO14, LRI03]. Event-related [BCM12]. Events [FRD14, PS12]. Ever [Pos01]. Everyday [MCMMAP^{+}14]. Evidence [GS12, LKT10, LAAPVGMM15]. Evolution [Alh04, APNA12, DSCT10, HNYO04, JS16, MA95, NSFVH05, SKA08, TFMDM10, VAPM12, DM07]. Evolution-Communication [Alh04].

Evolutionary [AMR^{+}14, ANdMM08, BCC^{+}06, CAJ06, DMM07, NdMM08, PTO^{+}12, SESMT10, Tra13]. Evolutionary-based [PTO^{+}12]. evolutions [LMA^{+}14]. Evolvable [NdMM06a]. Evolving [CC07, KS10, LALS08, SA10]. EX [TH99]. EX-OR [TH99]. Exact [BPHN06, KF10a, KF10b, Les95, RR06a, SLN16, WP03]. Exact-Four-Colorability [RR06a]. Examination [Tak06]. Examining [DA13, TT98]. Example [Dit02, Dru13, FG03, MA13, VDSF98, VDBNR98, dKR03]. Examples [FTARR05b, MVMRULS12]. Examplets [GB03]. Exams [SBTH04]. Exception [FF07, SIIJ09]. Exceptions [DMCM14, Suz06]. Excerpt [Rad14]. Exchange [AMBP04, DLY08, EMGB^{+}12, JL08, KJL09, SG02, WLKW11]. Exchanges [VUT^{+}08]. Exclusion [CE06, BSP^{+}13]. Execution [BJ05a, HSFE12, KAG00, PRCRARLN10, PMLL09]. Exercise [GCL^{+}13]. Exercises [QL12]. Existence [Bos09, Gal98, Ois98]. Existing [PK98]. Expansion [JR02]. Expansions [DF99]. Expect [DKM04, Sch02b]. Expected [DMCM14, SHZ^{+}10]. Experience [ASS13, CC08a, Cos08, DAk13, FC03, HMHGR15, IBN^{+}11, Kil08, uRLFH^{+}13, Lei10, MS10, SST07, Sch02b, Sha11, VGGSBLAP12, WCH14]. Experiences [BPC04, BFMSP05, Flo04, GPCZ^{+}13, MS05, RS03, dSC06b]. Experiencing [PdlCBKN14]. Experiment [GJAB95]. Experimental [BCM08, CE11, KTL^{+}11, VK03, dSC05]. Experimenting [CCMP08]. Experiments [ABB14, DLL14, FWS^{+}11, MS01]. Expert [CDBZ09, HSFE12, JL16, May02, MPF04]. Expertise [AM11, WP03]. Experts [FFK04, PJN13]. Explanation [RH10, SH09]. Explanations [Alm06]. Explicit [KKH12, Pre97, SLJC08, OAR^{+}14]. Exploitation [Sch02b]. Exploiting [CR04, GLSD11, GLD^{+}12, PRW95, Ros05, YJY14]. Exploration [FTARR05a, GT01, KS05, PB05, RTB13, SGB^{+}13, Tra13]. Exploratory [CB12]. Explore [HMA^{+}05]. Exploring [GXC^{+}15, Gre08, HGIPCPPM11, JJ12, Pel14, SRI08, WWD15, ZABB07]. Exponential [RR06b]. Exponential-Time [RR06b]. Exponentially [PPG95]. Exposed [VJ09]. Exposure [ASS13, HKKvP08]. Expressibility [Gom08]. Expressing [MCG14, MP06]. Expression [TG10, ZTX^{+}07]. Expressions [MN14, SY99, SSSS10, Sou99]. Expressiveness [HL09]. Extended [CT16, FL14, HK15, PGT09, QQ11, TBVRGLD15]. Extenders [Sar05]. Extendible [CT97, Gro00]. Extending [HSD^{+}14, LMRG14, MCG14, NDAM09, Pal15, PLG^{+}08, VO10]. Extensible [HCD10, HBI98, MI07, TT08, Win97a, dBdd04, BRH^{+}08]. Extension [AR01, Buz06, Gro00, JP07, MI07, RGP10, RFMLP10, YWD08]. Extensions [CS07, MMM^{+}12a, MGMS12, Ros99]. Extentions [Mes02]. Extracting [TJ15, Wac02]. Extraction [BCG^{+}09, CDR^{+}09, Gre08, GC14, HPB12, Jun05a, KC08, NdCFB08, SBPR15, SBMD10, TL11, TEC^{+}07, VV15, WMJ^{+}07, Yan05]. Extractors [XZ00]. Extrinsic [KSR16]. Eyes [SHH10]. Eyes-Free [SHH10].

Facebook [BU13]. Facial [TG10]. Facilitating [AMBP04, Epp04]. Facilitation [CE11]. Factor [HCPdASY14]. factorial [Dit02]. Factorization [HT13]. Factors [BM03a, CLM10, DM04, FP05, MX05, PŞ95, Pel14, Pet09, SK08, vBK08]. Factory [MB09]. Factual [MNS^{+}12]. Faculty [Len00]. Failure [Pet09, SH06]. Failures [ND08]. Fainting [GPLV13]. Fair [DLY08, RMMLBLGS09]. Falls [GPLV13]. Families [SK13, FK16]. Family [DGlS12, PJRC04, TD96]. Famous [Sko08].

Fan [DS10, Sch05c]. Farm [LNML03]. Fashion [PTL11]. Fast [AT10b, CFF^{+}13, Log04, PQ99, Wat02, YS05, DMMM95]. Fat [KD05]. Fault [AG06, BRR99, Gär99, KASN08, Log06, LMRG14, Sch99, WXZL15]. Fault-Tolerant [Gär99]. Faults [KASN08, MSF99, Tab07]. FBT [Hor04]. Fears [CDCH09]. Feature [AA16a, AZMA15, ARRB14, BMGMF08, FWT11, GMB08, GMB11, GBP^{+}08, GCC16, LWS11, LKP11, RdKO11, dPSZPVLR^{+}16, SL10, YKA16, dMTS^{+}14, dOBGH^{+}14]. Feature-Driven [dOBGH^{+}14]. Feature-Opinion [SL10]. Featured [PLG^{+}08]. Features [PLB14, SH11]. Federation [BVV^{+}10]. Feedback [CGFSHG09, HT13, MJG15]. Femtocell [CWT^{+}15]. Femtocell-Enabled [CWT^{+}15]. Festschrift [CS00]. Fibonacci [MS03]. Field [PCS^{+}13]. Fighting [BSP^{+}13]. File [HFIBJ13, KP95, QF12]. Files [SBPR15]. Film [Wac02]. Filter [AT10b, Bai12, CLCC10, Dru13, Sch08b, UDC97]. Filtered [DMM07]. Filtering [AAGU97, CM09, PJO15, PCKJ11, SOO97]. Filters [IKM03]. Finalizers [LI05]. Financial [Dru12, JFZ09]. Finder [BGK02]. Finding [Bos08a, CS02, FF04, LHK^{+}13, NO98, Pau10, PMP02, Tru10]. Fine [Bra02, Kol05, LN08, LL12, MYT09, YTM05]. Fine-computable [MYT09]. Fine-Grained [Kol05, LL12]. Fingerprint [LD06]. Fingerprinting [DCS09, VSGP05]. Finite [AT10a, ANdMM08, BB08a, Cal96a, CC96, Cv99, Dru06, Jéz95, Kar02, Kon02, MSSY95, Pop07, Sal02, Sar05, TY09, Wat02]. Finite-Delay [Kon02]. Finite-Fuzzy-Automaton [MSSY95]. Finite-time [TY09]. Firmness [Hav05]. First [APM04, BTD^{+}07, DLL14, DOM10, GOM^{+}13, MC07, MR05, MGT14, NK95, RA06, HCK11]. First-Class [MR05]. First-order [DOM10]. First-Person [GOM^{+}13, MGT14]. Fitness [GYY09].

Fitted [ZG05]. Fitting [AM96]. Five [SGS13]. Fixed [AGO^{+}13, GHNT97]. Fixed-Length [GHNT97]. Fixed-Point [AGO^{+}13]. Flexibilities [LWG14]. Flexibility [dTU04]. Flexible [ARRB14, BZ09, CRMLN^{+}07, HJVK15, HBI98, HR06, MTB^{+}08, OBO09, SD97]. Flipping [HOS96]. Floating [ATOFF98, DF99, Har07]. Floating-Point [Har07]. Flooding [HHH^{+}02]. FLOP [LMMPFVI14]. FLOSS [GS12]. Flow [BM11, CLC04, Gir05, Jun05c, LM03a, LNHZ09, MTK97, NZM09, PRT^{+}08, PRBLAP^{+}13, PdlCBKN14, RVC12]. Fly [OB95, Hor04]. FNR [Pre12]. Focus [FGSW14, LBG07, Rad96]. Focused [SSV02]. Fold [AGK^{+}10]. Folding [FSSPLG^{+}13, HI00, NR08b]. Forces [DBBS08, Pet09]. Forcing [DG07]. Forecasting [KÜ10, NWDX09]. Foreground [BZA08]. Foreground/Background [BZA08]. Foreign [PSVOVI07, TBL15]. Forensic [IAS16, NVB12]. Forensics [SV05, YLL^{+}07]. FOREST [KP00]. Forests [Tom97]. Foreword [BMMM95, Mul98b, Mul98a]. form [NdCFB08]. Formal [AC05, Abr07, AFK01, AS07, Ban07, BR97, BRS00a, Bjø01, Bol06, CSY02b, CDD^{+}03, CS05a, Dvo00, DSRR03, EGG^{+}01, FL14, GSW04, Gär99, Hal07, HK14, Hei07, Hel07, HNS07, HSD^{+}14, KNLS00, KASS03, KKK^{+}14, KAM03, KSU02, KP97a, LI05, LDSG09, LT09, Ma12, MS00, Mat99, MdCRMP14, NuR05, Pal05, PR06, PS04, PO04, Rad14, Rus07, ST05, SRR04, VLE12, Vel04, YMP08, YWD08, TBL15]. Formality [CM00]. Formalizing [BB08a, Bür08, GSW97]. Formally [GPSV03, Pop95]. Format [GHNT97, MNF^{+}13]. Formation [POR10, SvRvdV^{+}13, SvRS14]. Formative [HMHGR15]. Formulae [DL99]. Formulation [IKM03]. Formulations [AHRH08]. Forth [PS04]. Fortran [NI03]. Forum [Pre97, Pre12]. Forward [Sch01a]. Forwarding [SH10]. Foster [FPS^{+}12]. Fostering [LNHZ09, MH02, RS02]. Foundation [ESM08, FZ00]. Foundations [JR96, KT10, MSC03]. Four [HK95, RR06a]. Fourier [Ang98]. Fourth [JJ12]. FPGA [PBTW07]. FPGAs [GCVRSPGP07, OCRPdlMG07, de 00]. FPT [ECHS10]. FPTAS [LEC11]. Fractals [Ev99]. Fractions [Les95]. Fragmentation [AAM14]. Frama [Puc10]. Frama-C [Puc10]. Frame [LWH09]. Frame-Relayed [LWH09]. Framework [AMA^{+}14, AHPSCDK14, AT10b, ASHT^{+}16, AAGU97, Are02, Arr07, BCA^{+}10, BBIC13, dMBHR15, BCCH11, Bor07, BKL12, CSFFM12, CPLPW15, ČSŽ09, DSM13, Dud08, FBSEGP15, GR02, GRC15, Hef04, HLHD^{+}07, HCD10, HNJ^{+}10, INK09, KJ10, KL09a, LT09, LLCN09, LGES11, LRS^{+}11, MSA13, MBA12, MGNDAM12, MMD12, NuR05, ONRV08, OCB^{+}10, PFS07, Par09, PTNMC08, QL13, SLJC08, SW09, SKH14, SJ13, VAH07, VV15, VBVNHdDSL12, VTGA13, Wol00, WLtHN14, ZTN^{+}15, dIVG^{+}06, vZdlH12, Ram01, Træ08, CEK15]. Frameworks [ESM08, MSSV14, RMFM12]. Fred [CP01]. Fredkin [LZM04]. Free [AP09, Azz10, BR05, BJ97, BH02, Gro00, HI99, KBF^{+}11, RdKO11, SHH10, FDR^{+}15]. Free-Extendible [Gro00]. FreeBSD [GS12]. Freedom [BM96]. French [WBS12]. Frequency [CFJS15, CW09, KTL^{+}11, LGZ01]. Frequency-Domain [CW09]. Friendly [LMMPFVI14, PZJ09]. Front [LdSM08, dSLMW08]. Frontiers [RA06]. Frustration [EACGFK13]. Full [HL09, HA13, LKHL09, Mue04]. Full-Round [LKHL09]. Fun [Rho10]. Function [BPSN97, CJZ13, FKS^{+}04, GW03, KMN16, PPG95, RMGCGCF08, YTM05]. Function-Complete [FKS^{+}04].

Functional [GPSV03, HL09, KASN08, MABS05, Rad14, RTL05, RB07, SH06, Sch05a, SS00, She05, Tur04, CTM10]. Functional-Logic [SH06]. Functionalities [TBVRGLD15]. Functionality [PK98]. functionbased [RS00]. Functions [BDPSNG97, Bol06, CLVM09, Cla05, CPS07, Dun96, FH00, Fon00, Her97, KSY97, LO98, ML98, MYT09, OS98, PQ99, RR08, Rud04, SZZ95, Wei08, ZZ95, DMMM95]. Fundamental [Ban96]. Further [Lin08b, CCHdCN08]. Fusion [GWG96, Jun10b, MP08, MLX10, NHH06, OMO10, RR11, XLMR10]. Future [AKM07, Car96, COBP^{+}14, Dom01, Fel01, FHH08, Fra98, IKC14, KP01, LUR16, LRB16, MO03, SGLM16, Viv96]. Futures [MMiF^{+}08]. Fuzziness [DHP03, WF12]. Fuzzy [ASH11, Ach06, CHPHV10, CPRT05, CA14, CCHdCN08, Fal10, JKKW16, KÜ10, KCK10, LHC^{+}13, LA07, MBA12, MSSY95, RMGCGCF08, TWW07, Tra13, YX10]. Fuzzy-DEA [ASH11].

G [AKM95, PV95, Sch96]. G9 [PGSAP14]. Gabor [CLCC10]. GAC [ZZ95]. GADYM [TIL08]. Galois [CCS00, ZS04]. Game [CB12, DLL16, FPS^{+}12, Kul05, MMS11, Pel14, PD04, VDBNR98]. Game-Based [PD04, CB12]. Games [ARS16, AY12, Bür08, CGP^{+}07a, PZLAS^{+}13, PvW16, PGT09, Ret08a, SKHK14, SCW08]. Gamification [HCPdASY14]. Gaming [RAS15]. Gap [OBO09, Sch02c]. Gate [LZM04]. Gates [SDJ99, TH99]. Gateway [LWY11]. Gathering [SHZ^{+}10]. Gaussian [Dru13, KMN16]. Gaze [JGL08]. Gaze-based [JGL08]. Gazetteers [Rie02]. GD [VJTJ07]. Gelfand [CS05a]. Gender [BBM12]. Gene [BMM^{+}09, ZTX^{+}07]. General [AT10b, KK10, MMEdP12, MG14, NML09, OK98]. Generalization [ACAMM15]. Generalized [Cv99, Gün96, Mar06b, Sch01a].

Generalizing [PU97]. Generated [BGBA10, BD00, MNS^{+}12, SS03, SKL08, GC14]. Generating [CP01, Rad14]. Generation [AKP01, AKMS94, BHQW02, BS08, CVM11, CYL11, CT16, CDCH09, CBO05, FHH08, Gir05, Hri02, KD01, LZZK14, LMA^{+}14, LAHZ^{+}15, MJGS12, MSHN06, MS94, MDY10, OHYJ16, PD99, Sal02, Sch96, Sch02a, SLJC08, Sto99, XZSS09]. Generative [DSAFW07, SMGMT09]. Generator [PBB08, TEC^{+}07]. Generic [Bod01, LGMM^{+}13, LKMS08, MBA12, Pau99, RBB06]. Genetic [ACP06, Bai12, BMM^{+}09, CS02, EC00, FOSS99, Gir05, GYY09, LL97, LLLL99, LVV10, LWG14, TIL08, TWW07, TLJ11, UDC97]. Geo [CAGMPGdAS13, GHS06]. Geo-based [GHS06]. Geo-Recommender [CAGMPGdAS13]. Geometric [ACA^{+}16, SBCJ03, SMK^{+}04, Ukk10]. Geometrical [BCG^{+}09]. Geometrically [Ma10]. Geometry [Aur01, Smy00, Wan95]. German [RA06, FZ00, NH03]. Germany [Bec03]. Gesturing [HA13]. Getting [FCM^{+}12]. Girls [BBM12]. Given [BCCH11, HL03, GWW05]. Glimpses [Dom01]. Global [ATV98, AMS04, AH04, CPSAGPGC12, dFER06, HP15, Köh09, KNSN07, LHZS12, LHC^{+}13, ML98, Mes02, MM12, She96, ZZ95, dTU04]. Globally [HHH^{+}02]. GLORP [Lei08]. GmbH [Kuh03]. Go [PTL11, Pob11b, MPRS95]. GO-like [MPRS95]. Goal [AGMT10, CJH12]. Goal-Driven [CJH12]. Goal-oriented [AGMT10]. Goals [PGT09, RBLR02]. Going [SLD^{+}16]. Gong [BPSN97]. Good [Odl94]. Goods [JKKW16, NZCG05, Sch09b]. Google [FCM^{+}12]. Gopher [MA95]. Gossip [VSGP05]. Governance [Pob11b, RMFM12]. Government [GR14, MSSV14, Pau13, RC10, SG96, SDLM14, SVV10, VL14]. Governmental [GVT11]. GPRS [LZ09]. GPS [HZZ^{+}12]. GPU [JJL08]. Grade [FSELC13]. Graded [LST07]. Graduate [HAS^{+}07]. Grained [Kol05, LL12]. Grammar [Arr07, BH00, FP95, MVM00, MZ12, PO11, CVPS95]. Grammars [DT09, PRS95]. Grand [WB07]. Grant [FZ00]. Granularity [BR03]. Granules [Sch02d]. Graph [APM04, AC07, AR95, BR07, Bod01, CGFSHG09, CT04, CDF97, DSCT10, HKK13, KD01, KASS03, PO11, RR06a, SNAF07, SGS13, VJ09, WZZ^{+}09]. Graph-Based [CGFSHG09, HKK13, VJ09]. Graph-Grammar [PO11]. Graph-theoretical [BR07]. Graphical [GGMM^{+}13, NS06, SS08]. Graphics [Fel01]. Graphs [DSLO04, Her09, LNML03, LST07, LM07, PB14, SLN16, ZZ07]. Gravi [HMA^{+}05]. Gravitation [Lan98]. Gray [Dor07, Lom07]. greedy [Man97]. Greek [HA10, LTY^{+}16]. Grid [CAS^{+}13, CPCLSAGC11, JN08, LVV10, RFMLP10, SMK^{+}04, TCS^{+}03, ZZ07, dOMdAL^{+}08]. Grid-Enabled [LVV10]. Grids [Sto02]. Grounded [GXC^{+}15]. Grounding [Kat05, Liy02, RRM^{+}12]. Group [CVSM11, CS14, GBCA12, Har00, KTJ05, Mue04, POR10, RLL^{+}10, SE09, SLJC08, SDJ99, dKR03]. Group-Aware [SLJC08]. Groups [GÁVCNC14, GFT09, MX05, SG02, XNKG15]. Groupware [DPZ08, GMC^{+}08a, GGB^{+}08, MDO^{+}09, MROH08, PLG^{+}08, SL96]. Growing [Her97]. Growth [MR14]. GSM [CDP13]. GTFOF [MBA12]. Guaranteed [NO98]. Guaranteeing [GGDBP^{+}08]. Guest [Ban07, Pre12, Car96]. GUI [RMGCGCF08, VK03, WKXL05]. Guidance [Sut01, VJ09]. Guide [SDLM14]. Guided [LWC^{+}04, Ram01, TSE^{+}15]. Guigues [Kuz04]. Gurevich [Boe97c, DDG97]. Gyrolayout [UCM13].

H [Loo06]. H.264 [LLYC12]. H.264/AVC [LLYC12]. Halpern [Leu07]. Halting [Svo95]. Hamiltonicity [ZZ07]. Hammerstein [CW09]. Hamming [Fen15]. Hand [BSB09]. Hand-Held [BSB09]. handed [VCB08]. Handel [GCVRSPGP07]. Handel-C [GCVRSPGP07]. Handheld [Mar07, PSVOVI07, RAS15]. Handling [FF07, MUF03]. Handover [GGDBP^{+}08]. Hands [DGN13]. Handset [SKH14]. Handset-based [SKH14]. Handwriting [FOAB08]. Handwritten [MR11, PO11, RR11]. Haptic [SdBC13]. hardness [Bod01]. Hardware [AG06, Ad03, Lav96, MJS13, Nd05, NdMM06a, NdMM06b, ON97, dCPUH^{+}07, SNAF07, She05]. Hardware/Software [SNAF07]. Hash [BPSN97, BDPSNG97, Mue04, RR08]. Hashing [Lep98, PQ99]. Haskell [Ad03, RTL06, VFC03]. Haskell# [dL03]. Hausdorff [LW08, Sta05]. Hazards [TH99]. HC [CSFFM12]. HCI [GCG08]. HD0L [Hon02]. Head [PRT^{+}08, TG10, LLLL99]. Health [ABM^{+}06, CSFFM12, Coo06, GK06, MHA^{+}15, Rob06]. Healthcare [MYY06, RY09, SLD^{+}16, TFG06, vZdlH12]. Heap [RS00]. Heat [Jéz95]. Heavy [HVM00]. Heights [PŞ95]. Held [BSB09]. Hellman [HLCL11]. Helmut [CSY02b]. Help [Hen98, LM01]. Helping [KHG10, LAAPVGMM15]. Helsinki [LKT10]. Hénon [Gal98]. Hermann [KP01]. Hermitian [ACA^{+}16]. HEROs [LRR04]. Heterogeneous [ASW^{+}03, BMV12, BARB12, BS96, Jun10c, Kou09, LZ09, POR10, Rat00, VBB13]. Heterogeneous-Group [POR10]. Heuristic [AMS04, BCS15, GGP08b, RMGT09, XMZbL10]. Heuristics [FKS^{+}04, POR10]. Heyting [KSU02]. HIBE [CS07]. Hidden [Dru12, FMS12, Mar06b, PSS^{+}13, VdSdMC08]. Hierarchical [BAPG03, HN07, HHH^{+}02, JP07, LGAP11, LYLX15, MR11, PRAT09, PO11, TYSY09, TSE^{+}15]. Hierarchies [Mar02a]. Hierarchy [BG97, FK16, FFK04, HHH98, KKH12, RR06a, TKSL05, ZD09]. High [Bör02, Fen95b, GOF05, LS95, LZZK14, MPPS95, Pel14, PSS07, SAA08, Zim01]. High-dimensional [GOF05]. High-level [PSS07]. High-radix [Fen95b]. High-Speed [MPPS95]. Higher [BRAS^{+}12, BBM12, DD13, Flo04, HBT12, HR06, JM10, KM13, MHLB12, PS06, RB07, SVFR15]. Higher-Order [RB07]. Highly [BBL13, HMA^{+}05, HHHX09, KSY97]. HIKS [DGK^{+}99]. Hill [KD11]. Hill-and-Dale [KD11]. Historians [BVG08]. Historical [BZA08, Lar01]. Histories [Kom02]. History [BVG08, Kar13]. History-Sciences [BVG08]. HMAC [RR08]. HME [Mar07]. HMMs [EIH08]. Hoc [HHHX09, HJZ07, KTJ05, LKK08, CAD^{+}06, EMZB14, NOP08, VVM^{+}06]. HOL [BRW03]. HOL-Z [BRW03]. HOL4 [HK14]. Hold [IKC14]. Hole [YKA16]. Home [GHHE^{+}08, GK06, Hei96, HB13, JFL^{+}13, MYY06, SW09, vZdlH12, DOOJ95, vZdlH12]. Home-based [vZdlH12]. Home-care [JFL^{+}13]. homeML [MNF^{+}13]. Homes [OCW13]. Homogeneous [Her09]. Homography [SCK^{+}09]. Homomorphic [Aki09]. Honor [AFK01]. Honour [CSY02b, CI05, CSZ07]. Horn [FH00]. Hot [Toc03]. HTML [Kol05]. HTML5 [BAZ14]. Huffman [Tom97]. Human [AWGS04, AT10b, AT07, BMUF14, BRO08, CNRM03, CPHC11, CPFSdAS12, FP05, GPA08, GYY09, HPB12, Kim10, LST14, LG08, LK01, MGM^{+}08, MN96, NPB06, SCK^{+}09, UP04]. Human-agent [LST14]. Human-Computer [BMUF14, GPA08]. Hunt [KY10]. HWOES [LLM02]. Hybrid [BGMR^{+}16, BCHM12, BZA08, CS14, CM98, GGS08, HIGM13, IMR^{+}12, KLT13, Kie05b, LWH09, LA07, MJG15, NBGS06, NZM09, OFCB08, Sob05]. Hygiene [CD10].

Hygiene-Compatible [CD10]. Hyper [Sto99, AKM95, PV95, Sch96]. Hyper-G [AKM95, PV95, Sch96]. Hyper-Tableaux [Sto99]. Hyperbolic [MM99, Mar00b, Mar02b, MS03, Mar04, Mar06a, SS09b, UCM13]. Hypercube [Har00]. Hyperlink [Fro02]. Hyperliterate [NSL96]. Hypermedia [ARS16, AKMS94, BGP08, BE98, CC08b, De 96, DOS95, DOOJ95, DHO98, FGS98, GR02, GRGPL08, GSMBFPK10, HC08, LT13, LM94a, MRK^{+}98, Mat02, MP10, MHH98, RP98, SMMC10, SH09, Ski00, TD96, WD02, Toc02]. Hypermedia-Based [RP98]. Hyperspace [TT98]. Hypertext [De 96, MH96, SA11, TT98]. Hypertexts [Meh02]. Hyperwave [LLM02]. Hypothesis [Sch05c].

I-Know [TM02a, TM02b, Toc03]. I.a [Ior07]. I.b [Ior08]. Ibero [GCG08]. Iceberg [HOPN11]. ICT [AHSN01, GMHGRG^{+}13, HA10, MHLB12, PGDD15, RF12]. ICTs [DIKL14]. ID [CS07]. IDE [FRD14]. IDEA [GCVRSPGP07, LGES11]. Ideal [Her09]. Ideals [BH08, CDF97, SS09a]. Identification [BHC05, CW09, Fon00, HK15, HL03, KM13, PF15, VSML03, WXZL15]. Identifying [CHH16, CA14, LA03a, SSdS^{+}11]. Identity [DH10a, MJGS12, PSS^{+}13, Yon11]. Identity-based [MJGS12]. Identity-Hidden [PSS^{+}13]. Idioms [ARRB14]. IDS [Yan05]. IEEE [MS00, Pop95]. IEEE802.16e [HLCL11]. if [PTL^{+}09]. Ignoring [ZDI10]. II [Ban97a, HMSS01, NSFVH05]. II. [Boe97b]. Illumination [DRRGdP07]. Illustrative [FG03]. I’m [SM04]. Image [ACB02, BvZH09, CM11, Cv99, FL10, HKTV06, HKL^{+}06, LN08, SH11, WH08, WWD15]. Images [BSB09, BNCGD^{+}11, CDR^{+}09, CVK97, Hon96, VJ09, dSLMW08]. Imagesemantics [SKL08]. Immersive [CH07, GOM^{+}13, IKC14, KGK12, PMAM14, QFB^{+}14]. Immune [CAGMPGdAS13, NZM09, VV12, YLL^{+}07, ZTX^{+}07]. Impact [ABAL09, BVG08, CPSAGPGC12, CE11, GK06, LTY^{+}16, LM94a, MNF^{+}13, MNS^{+}12, SH10, SA14, SZWdP14]. Impacts [WWD15]. Impedance [Neh98]. Impending [Odl94]. Imperfect [MTK97]. Implement [ASHT^{+}16, PRT^{+}08]. Implementation [Ad03, Ara97, Bai12, BHH^{+}06, CJO^{+}13, CXB12, CGPAP13, DLL14, DS08, Fro02, GCVRSPGP07, HM00, IdFC05, JGW11, Moo08, OCRPdlMG07, dCPUH^{+}07, PBTW07, Pop95, RS02, Rho10, SH06, SDLM14, TT08, TBVRGLD15, UDC97, VFC03, WZC07, WD02, YKD^{+}08]. Implementations [BCNR07, Gle03, dOMdAL^{+}08]. Implementing [BGP07, BEH^{+}05, CGP07b, FR04, Lep95, RMGCGCF08, UP04]. Implication [GNP05]. Implications [HI00, LF98, MO03]. Implicit [CP02, HT13, MJG15, NW04, Rih98]. Implicitly [Bol06]. Importance [AdGCD^{+}15, VO10]. Impossibility [MSHN06]. Impression [LGZ01]. Improve [CGLdMAC14, CS04, HMHGR15, JR02, LS07, MMB08, PFS07, STFM12]. Improved [LEC11, LYLX15, PKP08, WSL07]. Improvement [BPHN06, HCPdASY14, SR00]. Improving [AHSN01, BCG98, CSFFM12, CNQ04, FF07, FMS12, GMK05, HPB12, HLCL11, JS16, KJKS14, Mau97i, MGPB07, Ozd13, Reb96, RR06b, TNRGCP^{+}13, TEC^{+}07, VBO08, VL14, WKS^{+}99]. IMS [BTD^{+}07, FCM^{+}12, GSPK08]. In-Service [AHSN01]. in-World [PMAM14]. Inc. [Bec03]. Inc./Germany [Bec03]. Incentive [dMBHR15]. Inclusion [CR00a, PGSAP14, SdOB09]. Inclusions [CG09]. Inclusive [DD13, SGLM16, SCS13]. Incomplete [CHPHV10, CL97, HN07, MRGF14, Pop05].

Incompleteness [CPC00]. Incomputable [Pau10]. Inconsistency [Ngu05]. Incorporation [FSELC13]. Increase [PBTW07]. Increasing [CS05b]. Incremental [AK09, BARB12, BQV14, Dud08, FTARR05b, FDR^{+}15, SF00, dH04]. Incrementally [EKP03]. Indecomposable [Ev99]. Independence [AC07, Ban97a]. Independent [CDBZ09, EIH08, HKL^{+}06]. Index [PSS^{+}13]. Indexing [BNCGD^{+}11, Rad96]. Indicators [dPPRRGSSPP15]. indiGo [DRA^{+}04]. Individual [AT13, CNRM03, CDD^{+}04, GYY09, RN03, RPR11]. Individualized [CJH12]. Indonesia [KW10]. Indoor [SM02b]. Induced [JMEL10]. Induction [Ber10]. Inductive [CCHdCN08]. Industrial [CPMVG13, Ern11, KA97, SFP12, TSCY01]. Industry [Bür08, LH03, TKD^{+}09, VGCPAH16]. Inequality [CG96a]. Inertial [PKSR09]. Inexact [And96]. Inference [BCA^{+}10, BMM^{+}09, VFC03]. Infinitary [Cre09]. Infinite [Cal96a, CT16, FDR^{+}15, Mar02a, Tru10]. Influence [SP16]. Influences [ASS13]. Info [BSP^{+}13]. Info-exclusion [BSP^{+}13]. Informal [GPCZ^{+}13, GGB^{+}08, VLE12]. Informatics [ABM^{+}06, BVG08, PS06, TM00, VV06]. InFormation [dSBGdAdLM08, ATSJ05, AY00, AP09, And96, ABPS95, AKM95, ARS00, AUN04, AdGCD^{+}15, BDM15, BKK^{+}08, BEH^{+}05, BCCH11, BS96, BP97, BG98, CCCP08, CHPHV10, CP15, CXB12, Cal96b, CS03, CR04, CYL11, CPHC11, CLM10, CKPK13, CDR^{+}09, CS09, CS02, Dru12, DS08, EMGB^{+}12, FMR09, GLCV08, GlRBdSG11, Gre08, Güt08, HKS96, HW10, HTHW12, HBF10, HMA^{+}05, Hol96, HGMT08, JM15, JR02, JM10, KC08, Kap95, KDKB07, KSR16, KM06, KTKP09, KMM14, Kul05, LM03a, LJLCR13, LST14, LLZ16, Mar96, MLX10, MSC03, MRP14, MR12, MMM12b, MKI^{+}12, Mue95, MTK97, Nag06, NPB06, NHH06, NBGS06, PB05, PJH12, Rad96, RAWW05, RvS12, RMFM12, Rie02, RSFMJ12, Sch01b, Sch02c, Sin06, SCLM03, TR10, TEC^{+}07, TSDP07, VMA14, WLKW11, Wol99a, Wol99b, Wol00, Wur10, XLMR10, YX10, YW13]. Information [dCVM12, dKR03, vKL04, Svo96]. Information-Theoretical-Based [CS02]. Informative [KC08]. Informer [TEK08, SOO97]. Infrastructure [BBdOR14, LS10, NML09, TKSL05]. Infrastructures [LAHZ^{+}15, Mai05, SL05]. Inhabited [Pet12, SW09]. Inheritance [DRS06]. Inhibitors [IS04]. Initializing [HT13]. Initiatives [HCPdASY14, MPM12]. Injection [BRR99, MSF99]. Innova [FSELC13]. Innovation [LMA^{+}14, NL10, RS11, SBRS11, SBG^{+}12, VBP^{+}11]. Innovations [MR12]. Innovative [GDW10, Uzu13]. Inoteroperability [Kon03]. Input [PA12, VCB08]. Inquiry [HB11, KY10, MOS^{+}13]. Inquiry-Based [KY10, MOS^{+}13]. Insecurity [LCDP15]. Insertion [RMGT09]. Insight [CBRH12]. Insights [BBM12, PdlCBKN14]. Inspection [MM12]. Inspections [dMTS^{+}14]. Inspired [ANdMM08, BCM12, CAGMPGdAS13, LWS11, NR12, NZM09]. Inspiring [JRO10]. Instance [JR02]. Instances [LWS11]. Instantiated [RR08]. Institute [JV11]. Institutional [LKT10]. Institutions [GPCZ^{+}13, GMP^{+}13, Mar06b, Ros99]. Instruction [Reb96, SAA08]. Instructional [DSAFW07, MGPB07, Uzu13]. Insulated [FZT13]. Intangible [CDCH09]. Integer [PŞ95]. Integers [GN10]. Integral [FF08, MYT09]. Integrate [FP05, MPF^{+}16]. Integrated [CPLPW15, Dro04, Fal10, HKS96, HMM00, LCHD12, LWG14, LHC^{+}13, SSGS10, UHÖD15, WKS^{+}99, Bec03]. Integrating [BFF99, BM97, BMG^{+}05, CLC09, DDSS05, GMC^{+}08a, GMdMC12, HNS07, Jun10b, dIdSGZ10, NOP08, PJN13, RJB10, SL05, Træ08, WKXL05]. Integration [AR04, BKK^{+}08, Bos08a, BBP08, CLC04, CM98, DKD^{+}13, DJJN09, FHJ^{+}99, GO14, GJP^{+}12a, HMSR99, HA10, HA03, KFK05, KTKP09, LF05, LAAPVGMM15, ME03, NDAM09, Pau99, SA10, VBB13, ZSG14, dTU04]. Integrative [LWH09, RN03]. Integrity [Kap95, Sue10]. Intellectual [CPFSdAS12, KL02, Liy02]. Intelligence [AGGH08, BAR06, BdI10, Cai10, CN08a, FJP06, GGP08a, GBP^{+}08, JRO10, JN08, Jun10a, LHZS12, dIB13, NN07, NH09, NCH16, PANS13]. Intelligent [CN08a, CN08b, CWTT11, GWG96, HME^{+}06, HAFS15, JNS09, JL09, JCB15, KKTZ09, KHG10, LGL09, LEJ^{+}08, NdMM08, RY09, VVM^{+}06, VUT^{+}08, VK03, ZMAS10]. Intensive [BCM08, FP05, GMK05, KKH12]. Intent [ZSG14]. Intention [SSM11, WWD15]. Intention-Aware [SSM11]. Intentions [HGIPCPPM11]. Inter [CE11, MLHCGB16]. Inter-methodological [MLHCGB16]. Inter-Organizational [CE11]. Interaction [AF04, BMUF14, BRO08, BEPT14, CFMP15, CBNDR10, GPA08, HAI13, IPCVC12, JGL08, Kom02, LG08, LM15, Luk08, MGM^{+}08, MFG13, MUF03, MX05, PEPP08, RPR11, RMM^{+}08, SHH10, SE09, SAB99]. Interaction-oriented [SAB99]. Interactions [BZ09, Kim10]. Interactive [BBL13, BJMBA15, CH07, CDBZ09, DBBS08, DZ08, Epp04, FZAP13, FTARR05a, GMC^{+}08a, GHS06, GYY09, GPSV03, Has02, HMA^{+}05, KSdV09, LNHZ09, LKB^{+}02, OHYJ16, PJN13, PSS07, RMF^{+}98, SLK11, TEK08]. Interactivity [ASS13, CGD^{+}12]. Interconnection [BB04]. Intercultural [DM04]. Interdisciplinary [BM03b, CG96b, Müh96]. Interest [CHH16, FLF^{+}14, SBMD10].

Interest-Driven [FLF^{+}14]. Interesting [Suz06]. Interests [Jun05a, YYZ^{+}09]. Interface [BDL^{+}06, BCCH11, Cañ08, CEK15, DA13, HOPN11, KD02, MR08, PT09, PLG^{+}08, SH96, Sha11, vdV08]. Interfaces [AWGS04, BCFM05, CSC08, GLCV08, HVM00, IPCVC12, MROH08, PMRO08, PLBG13, PLSF08, SLJC08, Sob05, SMV08]. Interference [LdSM08, LWY11, Mul00, dSLMW08]. Interleaved [dCPUH^{+}07]. Intermediate [IK97, LF16, Vel05]. Intermediation [TR10]. InterMod [LULGFC13]. Internal [Heg10, Kat05]. Internals [Rho10]. International [AUN04, CMZZ07, CR12, GPCPL12, GPVILN13, LJLCR13, TM01a, TM01b, TPC^{+}12]. Internet [Bor07, COBP^{+}14, DMCM14, DMG07, FZ00, FHH08, Fra98, HW10, HTHW12, HR06, Kim10, KWC01, Len00, LM94b, LVS13, LAHZ^{+}15, MJGS12, MA95, Mue95, NLLJ12, Pos01, QC12, Rie02, SZWdP14, SV95, She96, SM02a, Tez13, ZNX^{+}12]. Interoperability [AMC^{+}12, GPFL12, GRC15, LJLCR13, LVV10, MSSV14, MMD12, RLMS13, Riz15]. Interoperable [CBN^{+}06, MHA^{+}15]. InterOrganizational [WKS^{+}99]. Interpolation [AM96, RG00]. Interpreter [ESM08, HCD10, KKB12, MM96]. Interpreters [KASS03]. Interpreting [JST11]. Interpretive [JMP06]. Interrelationships [Pel14]. Interruption [ZMAS10]. Interval [FH06, GMS03, Hor04, Mar95, MKP98, Pop98, STVT07, Tru10, Zgr07]. Intervals [Mar98, MMEdP12, Sch02d, Tru10, WFOC98]. Intractability [Kuz04]. Intraframe [LLYC12]. Introduce [LM03b]. Introducing [CPR06, HMHGR15, LKT10]. Introduction [Blõ06, Boe97a, Boz99, DRA^{+}04, DR04, Hef04, Mar96, Wol99b]. Introductory [Pel14]. Intruders [AAJR05].

Intrusion [JT05, KMN16, RKJ16]. Intrusive [SV05]. Intuitionistic [Gol05]. Invariance [HW97]. Invariant [HI00]. Invariants [OF13, Sch08a, Tab07]. Invasive [CBR^{+}05]. Inventory [CLC04]. Inverse [Neh98]. Invertible [And97]. Investigating [AGK^{+}10, BJ05b, Kar13, VBP^{+}11]. Investigation [AA16b, CCHdCN08, IAS16, JR96, uRLFH^{+}13, SM04]. Investigations [CD13]. Investment [RRR10, Wie08]. Invisible [Car00]. Invocation [PdP^{+}04, Tddd03]. Involving [HLCL11]. Ioan [CSZ07]. IoT [CJH12]. IP [RMGCGCF08, SH10]. iPad [DGN13]. IPCity [Ano07]. IPD [Bec03]. IQM3 [CCCP08]. Irish [Mac01]. Irreducibility [Cha05]. Irregularly [CGFSHG09]. Isabelle [Pau99]. ISDN [LPP96]. Islands [CAR08]. Isolate [Tab07]. Isolation [XMZbL10]. Isometries [Ilj10]. Isomorphism [SAA08]. Issue [AFK01, BHRS03, Boe97b, BG01, CS00, CSY02b, DG00, Dvo00, DSRR03, EK99, GALR02, HMSR99, IFd03, KP01, KU00, KZ03, LA03b, Lin04a, Lin04b, Mat99, Mau03a, RS01a, RS01b, RS03, RA06, TM01a, TM01b, Toc02, TM02a, TM02b, Toc03, Mul98b, Mul98a]. Issues [Ban96, Car98b, CDP13, DPZ08, FHH08, Goo01, Kri99, Pal15, Rob06, TT98, UP04]. IT-based [TE06]. Itai [FP06]. Item [CAGMPGdAS13, KTL^{+}11]. Items [LLS05, PJRC04]. Iterated [Cla05, Rah99]. Iteration [Kud99]. Iterations [Leu07]. Iterative [BvdTV09]. ITIL [CI12]. ITP [CPR06]. IV [Ior00]. IWIM [BAPG03].

J [Mau03a]. J.UCS [AFK01, BHRS03, BG01, CS00, CSY02b, DG00, Dvo00, DSRR03, EK99, GALR02, HMSR99, IFd03, KP01, KU00, KZ03, LA03b, Lin04a, Lin04b, Mat99, MS94, RS01a, RS01b, RS03, RA06, TM01a, TM01b, Toc02, TM02a, TM02b, Toc03]. J2ME [DB03].

Japanese [RA06]. Japanese-German [RA06]. Japlo [Esp06]. JAVA [dSC05, DR10, Esp06, Fra98, NC04, ONRV08, SS07, Tddd03, dSC06b, dBdd04, von98]. Java-based [ONRV08]. JavaScript [CCYK15]. Jbook [BB08b]. Jim [Lom07]. JIT [dSC05]. Jobs [Joh01]. Join [FA06, KL09b]. Join-Set [FA06]. joins [BST09]. Joint [Epp04, SBCD15, WZZ^{+}09]. Journal [KKM08, Kul07, MS94, CMS94]. Journals [Odl94]. JPlag [PMP02]. JUCS [Boe97b]. Judgements [WC08]. Jürgensen [CSY02b]. Just [dSC06b, Ver10]. Just-In-Time [dSC06b]. Juxtaposing [MGAVF10].

KADD [FMA^{+}05]. Kalman

[AT10b, Dru13]. Kan [Ros99]. Kannada [MR11, RR11]. Karatsuba [KPdF06]. Karp [HNP98]. Kenzo [DRS06]. Kerberos [BR97]. Kernel [CLVM09, KMN16, Puc10, TF09]. Key [Cam98, CT08a, FZT13, GBCA12, HLC08, JL08, KJL09, XNKG15, ZBKK12, DGlS12]. Key-Insulated [FZT13]. Key-Share [CT08a]. Keyboard [BH14]. Keyboard-Card [BH14]. Keys [Was98a, Was98b]. Keywords [TGEM07, WMJ^{+}07]. KGC [HLC08]. KILT [MR14]. Kit [FG03]. KM [BCZ04, FP05, Hef04, Pet09]. KMDL [GMK05]. KNN [HKK13]. Know [LM94b, TM01a, TM01b, TM02a, TM02b, Toc03]. Knowledge [ASH11, Ach06, ASTL07, AR04, AMBP04, Arr07, BS03, BQBW03, BE11, BEH^{+}05, BM03a, BBC02, BM03b, BCZ04, BS96, BP97, BM05, Cam98, CN08a, CN08b, CNRM03, CWTT11, CLC04, CBG04, DDS04, DAd03, Dus05, ES03, Epp04, Ern11, ELFAR15, FP05, FFK04, GÁVCNC14, GL11b, GMK05, Güt08, HT01, HH03, HA03, HNYO04, HMSS01, JV05, JNS09, Joh01, JN08, KD02, Kom02, KPV^{+}11, KA97, Kuh03, KHG10, LRR04, LLH03, LKK08, LM01, LKT10, LA03b, LH03, LGES11, LF05, Liy02, LNHZ09, LK01, LdP11, Mai05, MTB^{+}08, MPM12, MS10, MT02, May02, MHA03, MH02, MMS11, MUSA03, MPF04, MMB08, MS11, Nal10, Ngu05, NH09, NH03, NW04, OCB^{+}10, PRB^{+}11, Pet09, PD04, PPP^{+}11, Rad01, Ram01, RAWW05, RS03, RS02, Rie02, RGHH97, SKSN07, SST07, Sch01b]. Knowledge [SE09, Sch02c, Sch03, SG02, SD97, Sto03, Sto02, SL05, SSV02, SSS12, TCS^{+}03, TL11, Thi00, TCK^{+}01, Til01, TKSL05, TM00, TM01a, TM01b, TM02a, TM02b, Toc03, Toc04, Tom01, THS11, VdR09, Wac02, WK05, WP03, ZWH10, ZD09, dTR03]. Knowledge-Attention-Gap [Sch02c]. Knowledge-base [Ach06]. Knowledge-based [HH03, LA03b, LGES11]. Knowledge-building [Cam98]. Knowledge-Intensive [FP05, GMK05]. Knuth [Ukk10]. Koblitz [AHRH08]. König [Sch05c]. Korea [HLK09b]. Korrigan [CPR01]. Kraft [CG96a, Gro00]. Kraft-Chaitin [CG96a, Gro00]. Krawczyk [HK15]. Kudo [Tak06]. Kuwait [DA13].

LaaS [PCS^{+}13]. Lab [AAAK15, CDG14, DBAB12, LKT10]. label [KK13a]. Labelled [DSLO04]. Labelling [JM15]. Laboratories [OAR^{+}14, PCS^{+}13]. Lag [Wie08]. Laha [PD99]. LALR [PBB07, PBB08]. LAN [LPP96]. Landau [BB10b]. Landing [JL09]. Landmarks [ZS04]. Language [AMUFVI09, AJBTEB06, AAAK15, Ad03, BCFM05, Bic15, Bra15, CPLPW15, CP02, DDS10, DCR^{+}07, EIH08, FRD14, FPLS03, FBCMR15, GRS08, KPS96, Kat05, KT10, LRB16, Lan10, LLSA13, Ma10, MMdMGM06, MSC03, MPR^{+}08, Mos05, OL08, PSVOVI07, PR06, RRB03, RTL05, RAS15, RAC10, RGHH97, RTJ01, Sal10, Sch06, SdBM05, TBL15, VKW15, WKSD^{+}11, WKTL01, Wol00, YMP08, dR05, dOLC^{+}07, KAM03]. Language-Independent [EIH08]. Languages [AMVM01, BRH^{+}08, Bol06, BH02, CSY02b, CS03, CGP07b, FDR^{+}15, HBF10, HI99, HHY02, Ito02, Kri99, KSU02, KP97b, LF06, MM98, MOMGSRFM07, Mat99, Mau03a, MCG14, MABS05, MI07, SY99, Sal10, Sta05, SGS13, dSC06a, DS08, BdVG06, BM07, IFd03, Lin04b, ML05]. Large [AR04, BMGMF08, CVSM11, HBT12, KS05, NHH06, QQ11, SLN16, Ski00, Sta02, VO10, WF12, Wur10]. Large-Scale [Ski00, WF12, HBT12]. LaSca [CVSM11]. Laser [MSF99]. LATE [Mau96a]. Latency [BD00, FMT^{+}15, LF98, PZJ09]. Latency/Power [FMT^{+}15]. Latent [BNCGD^{+}11, FBCMR15, HKKvP08]. Latin [Lin11b]. Lattice [Cet00, FH06, YX10, ZS04]. Lattice-Valued [YX10]. Lattices [APM04, CR04, HI00, Neg05, Sim07, ZQQ15]. Law [Pob11b, dFCC07]. Laws [GMB08]. Layer [EGK^{+}12, SHZ^{+}10, VL14, WH08]. Layout [CL08, MAT08, UCM13, VPF09]. Lazy [dAFL07, Lin03, PB05, RB07]. LBS [YKD^{+}08]. LCF [KM95]. LCP [CTM10]. LCP-Nets [CTM10]. Leader [FP06, MS00]. Leadership [AHSN01]. LeadFlow4LD [PRBLAP^{+}13]. Leakage [CKPK13]. Learing [Sut01]. Learned [BC11, BRAS^{+}12, CPLPW15, VGGSBLAP12]. Learner [Bra15, DGN13, GPCZ^{+}13, GGP08a, LT13, TSE^{+}15]. Learners [AHEAS^{+}15, AAAK15, HJVK15, HAI13, SdBC13]. Learning [AWGS04, ATGP09, Aïm98, AMA^{+}14, ARN04, AAAK15, ASHT^{+}16, AMBP04, AY12, BDGW96, BZ09, BBdOR14, BT08, BEH08, BGP08, Bic15, BDM15, BM13, BDL^{+}06, BARB12, BS08, BRAS^{+}12, BAP^{+}16, Bra15, BMG^{+}05, BTD^{+}07, BVV^{+}10, CGD^{+}12, CGLdMAC14, CPLPW15, CRLNAR05, CIM14, CRB15, CdSCSSA16, CJH12, CDBZ09, CGP^{+}07a, CGPAP13, DZBB^{+}12, DBB13, DD13, DL15, DDSS05, DIKL14, Dud08, DBAB12, EKW^{+}15, FDC^{+}13, FMLNF07, FTARR05b, FMA^{+}05, Flo04, Fon01, FWS^{+}11, FG03, GM05, GPCAC11, GPCZ^{+}13, GGP08a, GSdSB16, GXC^{+}15, GRC15, GMdMC12, GSPK08, Güt08, HKKvP08, HMW08, HBT12, Has01, Has02, Hel07, HLHD^{+}07, HAFS15, HMHGR15, HAI13, Hop98, HG11, IMR^{+}12, JBBH13, JBH^{+}10, JGW11, KWH03, KKK16, KLT13, KY10, KHLAP12, KOW01, KPV^{+}11, Kro13, KJKS14, KR11, LT13, LHK^{+}13, LM03b, LGMM^{+}13]. Learning [LBG07, Liy02, LLSA13, LAAPVGMM15, LZZK14, LMA^{+}14, MPG13, MVMRULS12, MHA^{+}15, ML95, Mau96a, MHLB12, MNDRF10, MOS^{+}13, MNS^{+}12, MMS11, MRO05, MM15, MC07, MMiF^{+}08, MS05, MX05, OAR^{+}14, Ozd13, PRBLAP^{+}13, PANS13, PJ15, dPPRRGSSPP15, PD04, PB04, QL12, QL13, QFB^{+}14, RAWW05, RKH15, Ret08a, RS11, RAS15, RLMS13, RGPK15, RP98, SVFR15, dPSZPVLR^{+}16, SGLM16, SVK^{+}15, SLD^{+}16, Sch01b, SE09, SW04, SL96, SH96, SS07, SA09, SBG^{+}12, SMMM13, SvRvdV^{+}13, SvRS14, SA03, TSE^{+}15, VV15, VLE12, VGAPGS^{+}15, Viv96, VKW15, WPL98, WMJ^{+}07, WCH14, Yon11, dlFVPHB15, AMR^{+}14, CB12, ES03, ELS04, MPF^{+}16, PGSAP14, SMFM05, LF05]. Learning-Aware [TSE^{+}15]. Learning-Based [WMJ^{+}07]. Least [HCK11]. Lectures [LHZS12]. Legacy [LAM12, VGGSBLAP12, dSBGdAdLM08]. Legal [BKL12, Kom02, Pau13]. LEGO [BvZH09]. Lemma [Sch05c]. LemmaGen [JMEL10]. Lemmata [Kud99]. Lemmatisation [JMEL10]. Lempel [Fen95a, Log04]. Length [Fen15, GHNT97]. Lessons [BC11, BRAS^{+}12, CPLPW15, VGGSBLAP12]. Let [AMYH14, dARGSB11]. lets [dIVG^{+}06]. Leukaemia [Faz06]. Level [BCA^{+}10, Bör02, BAP^{+}16, CPRT05, KNLS00, LMRG14, Ngu05, Rad14, SNAF07, UCM13, WH08, Kim12, PSS07]. Level-of-Detail [UCM13]. Levels [FM95, HLCL11, dTU04]. Leveraging [BC11, HA10, MJG15, OAR^{+}14]. Lexical [DCS09, MSSY95, PO04]. Lexicographic [KM95]. Librarian [Sto03]. Libraries [DGK^{+}99, DKM04, Fel01, HT01, ML95, NSL96]. Library [FC03]. LiDAR [RMZ15]. Life [BM97, PD99, SE09]. Lifelong [BRAS^{+}12, PD04]. Lifetime [SHZ^{+}10]. LIFT [dSBGdAdLM08]. Light [BRS00b, BG00a, BG00b, HB00, HW10, KP00, KJL09, QZB^{+}00, SR10, SF00, TWH00, dH00]. Light-Control [TWH00]. Light-Weight [KJL09, SR10]. Lightweight [EMZB14, TT08]. Like [NXSA12, Pau07, BBM12, CJZ13, CVK97, MPRS95, NPB06]. Limit [Sch08b]. Limitations [CCM09, FPSFCG07, FBSEGP15]. Limited [Fer96, GKK^{+}02]. Limiting [YTM05]. Limits [Cha96]. Lindenmayerian [HK95]. Line [ALHM^{+}14, BM11, FKO14, FWT11, FG10, JGM^{+}13, KD11, RP08, SEK13, dMTS^{+}14, Dru06, DT07, Hen98, LS95, OJSB08, RS00]. Linear [Akr09, And97, BG97, CGFSHG09, CPC00, DLR97, Han10, HK14, HNP98, Ish97, LVS13, NO98, Neg05, NR08b, SSSS10, Was98a, Gra98]. Linear-Nondeterminism [HNP98]. Linear-Sized [HNP98]. Lines [ABB14, BB08b, JGM10, LM10, LWS11, NKS^{+}09]. Linguistic [CHPHV10, CTM10, RLL^{+}10, vB96, PLB14]. Link [BE98, FGS98]. Link-based [FGS98]. Linked [DDJ^{+}11, PD10, PCLT15, VGAPGS^{+}15]. Linking [CSAC^{+}15, GAMP10]. Links [AKM07, SLT08]. Linux [SH10, TTB09]. Liouville [Neh98]. Lipton [HNP98]. Liquidization [HNYO04]. LISP [MM96, Cos08, DS08, Lei08, Lei10, NR08a, Rho10]. List [TTB09]. Listen [AMYH14].

Listening [RKH15]. Lite [WKXL05]. Lite-Weight [WKXL05]. Literacy [MLHCGB16]. Literature [GSZ15, PvW16, RB08, WMJ^{+}07]. Liveness [Zou06]. Living [HB13, dIdSGZ10, LKT10]. LMS [FCM^{+}12]. Load [BMV12, HF01, HCH^{+}09, HJZ07, LWY11, NAK08, NWDX09]. Local [APNA12, NVB12]. Localisation [MPF04]. Locality [FKB^{+}15]. Localization [AGK^{+}10]. Localized [PS12]. Locally [LW08]. Locating [AAJR05, PB14, SR10]. Location [AY12, WTA01]. Location-based [AY12]. Locomotion [BS11, MGT14]. Locomotive [CH07]. Log [SBPR15, VBO08]. Logic [BCDK97, CS00, CI05, CP06, Cre09, DG07, DOM10, DS03, Dru06, Hor04, Ior07, Ior08, IK97, JMP06, KJ10, KR03b, Kwo97, LKZK10, MLX10, MGMS12, MTK97, NL10, ONRV08, PR06, RK97, RMGCGCF08, SH06, Sch09a, SS00, SV08, SN01, SDJ99, TH99, VV06, XLMR10, dSC06a, dH04, dFCC07]. Logic-based [KJ10, MLX10]. Logical [RRB03]. Logics [GFO12, GMS03, Ish00, KO99, Mar06b]. Logistic [BS12b]. Long [BM05, RLL^{+}10]. Long-Term [BM05, RLL^{+}10]. Longest [CS10]. Longest-Path [CS10]. Lookahead [FKS^{+}04]. Looking [LRB16]. Loop [Nør10, SBCJ03, Sto99]. Loop-Detection [Sto99]. Loops [RC07]. Loss [Odl94]. Lossless [Log04]. Lost [TT98]. LOTOS [CM00]. Low [EIH08, FK16, JFL^{+}13, LLYC12, MWM10, SV05, SJ13]. Low-Budget [SJ13]. Low-Intrusive [SV05]. LP [MKP98]. LSB [CS05b]. LSE [LCC11]. Lua [HBI98, IdFC05, MI04, MI05, SRI08, dRI04]. LuaInterface [MI04]. LuaTS [LRI03]. Lucas [AFK01, Mac01]. Luggage [TLJ11]. Lukasiewicz [Ior00, Sta05]. Lukasiewicz-Moisil [Ior00]. Lymphoid [Faz06]. m [CIM14, CB12, Fod06]. m-Learning [CIM14, CB12]. MAC [ARQH14]. Machine [BDL^{+}06, BS01, CR00b, GV00, GK97, KLT13, MCC13, OL08, dPSZPVLR^{+}16, WTA01, WMJ^{+}07, MM96, ŞM96]. Machines [Ara97, ANdMM08, BCG^{+}99, BG97, Boe97c, BRS00b, BG01, BP08, CLVM09, DDG97, GRS08, GKK^{+}02, GS97, Kir09, LLLL99, OL08, SA97, Sch09a, STW09, Sch01c, SV08, SN01, Win97b, ZSK09]. Macro [ARS^{+}08, CD10]. Macros [CD10]. Made [dL03]. MaF [MGNDAM12]. Magic [AJBTEB06]. Magnetic [PKSR09]. Mailing [TTB09]. Mainstream [Rus07]. Maintaining [Kap95, WP15]. Maintenance [DAd03, LSV06, TR98, dH04]. Make [ES03, GPCZ^{+}13]. Makespan [NZM09]. Making [AMBP04, MLX10, SA10, XLMR10, YX10, GVT11]. Malicious [CCYK15, HLC08]. MALL [MV15]. Man [PHJ^{+}08]. Man-in-the-Middle [PHJ^{+}08]. Managed [PMAM14]. Management [ASTL07, ARN04, APDC08, Arr07, ABM^{+}06, BCA^{+}10, Bec03, BS03, BEH^{+}05, BM03a, BBC02, BP97, Buf01, CCCP08, CSFFM12, CAD^{+}06, CMMP16, CCP^{+}07, CB12, CPFSdAS12, CDD^{+}07, Dus05, ES03, FSdRSS11, GMHGRG^{+}13, GPCAC11, GBCA12, Güt08, HT01, HH03, HCWA03, HA03, HNYO04, JV05, JV11, JNS09, JGM10, KD02, KDKDN08, KA97, Kuh03, LRR04, LLH03, LM01, LA03b, LF05, Liy02, LK01, MYC14, MTB^{+}08, MPM12, MMM^{+}12a, MT02, MNDRF10, MPF^{+}16, NH09, NBGS06, NH03, PPJ08, Rad01, Ram01, RBLR02, RS03, RN03, Rob06, SBAZ11, Sch02c, Sch03, SG02, SC14, SW09, SA09, Sto03, Tar12, TFMDM10, Til01, TM00, Toc04, TTB13, Tom03, Tom01, TSDP07, WL13, WK05, WLtHN14, ZMAS10, dKR03, ZDE14, TM01a, TM01b, TM02a, TM02b, Toc03]. Manager [NH03]. Managing [ACR11, AAGU97, BCZ04, CG96b, FZ00, Güt13c, KHG10, LA03b, LH03, MBA12, MS10, QGT^{+}14, SPRP09, SSV02, VdR09, dTR03, Güt12a, Güt12b, Güt12c, Güt12d, Güt12e, Güt12f, Güt13a, Güt13b, Güt13d, Güt13e, Güt13f, Güt14d, Güt14a, Güt14b, Güt14c, Güt15a, Güt15b, Güt15c, Güt15d, Güt15e, Güt16a, Güt16b, Mau94, Mau95a, Mau95b, Mau95c, Mau95d, Mau95e, Mau95f, Mau95g, Mau95h, Mau95i, Mau95j, Mau95k, Mau95l, Mau96b, Mau96c, Mau96d, Mau96e, Mau96f, Mau96g, Mau96h, Mau96i, Mau96j, Mau96k, Mau96l, Mau96m, Mau97a, Mau97b, Mau97c, Mau97d, Mau97e, Mau97f, Mau97g, Mau97h, Mau98a, Mau98b, Mau98c, Mau98d, Mau98e, Mau98f, Mau98g, Mau98h, Mau99a, Mau99h, Mau99i, Mau99b, Mau99c, Mau99d, Mau99e, Mau99f, Mau99g, Mau00a, Mau00b, Mau00c, Mau00d, Mau00e, Mau00f, Mau01a, Mau01b, Mau01c, Mau01d]. Managing [Mau01e, Mau02a, Mau02b, Mau02c, Mau02d, Mau02e, Mau02f, Mau02g, Mau03b, Mau03c, Mau03d, Mau04a, Mau04b, Mau05, Mau06d, Mau06e, Mau06a, Mau06b, Mau06c, Mau07d, Mau07a, Mau07b, Mau07c, Mau08a, Mau08b, Mau08c, Mau09a, Mau09b, Mau09c, Mau09d, Mau10a, Mau10b, Mau10c, Mau10d, Mau11a, Mau11b, Mau11c, Mau11d]. MANET [AdI16, NM07]. Manifesto [JLRW05]. Manipulation [AMS04, Pop05, VCB08]. Manipulator [LL97]. Manufacturing [BCS15, LGL09]. Manuscripts [AGA12, BZA08]. Many [Din97]. Map [BHQW02, BM05, CFMP15, Gal98, SI00, LKK08]. Mapping [AGA12, ACP06, AMVM01, BARB12, BTD^{+}07, CL95, CLC04, GOM^{+}13, MG14, MC07, NW04, PTO^{+}12, SZWdP14]. Mappings [BBP08, Ish97, Leu07]. Maps [GRGN13, LNML03]. MareMonstrum [STFM12]. Margins [TNM09]. Marker [KH12]. Market [CO08, MMiF^{+}08, SCW08, SLT08, Tra13, ZDI10]. Marketplace [GJAB95]. Marketplaces [Ros05]. Markov [LLZ16, Ric05, Şte95]. Markovian [KHN99]. Markup [BVG08, SLK11]. Mash [KKM08]. Mash-ups [KKM08]. Mashup [FVIG12, INK09]. Mashups [BCG^{+}14, Kul07]. Masses [BKK^{+}08]. Massive [BEPT14, EACGFK13, GRCGK14, LMMRV^{+}15, OP15, PCLT15, RGPK15, RA06]. Master [SKSP09]. Mastering [LK01]. Match [BS03, CDD^{+}03, She05]. Matching [EACGFK13, Han10, LSX12, MGNDAM12, PPG95, SSSS10, Ukk10, Xi03]. Matchings [BR07]. Material [PSF98, SW04]. Materialization [CSA10]. Math [CSA10]. Mathematical [FHJ^{+}99, GRGPL08, Koó06, vD05]. Mathematics [Cha96, CS05a, Ish97, RP98, Mac01]. Mathuria [Tak06]. Matrices [BNA15, LM07]. Matrix [AMVM01, HT13, Pav95, TH99]. Matroids [LST07]. Maturity [CCCP08, SPRP09, SS15]. Maurer [KP01]. Max [YLW^{+}14]. Max-Cut [YLW^{+}14]. Maximal [PB14, VTHM16]. Maximality [BH08]. Maximum [FF04, YLW^{+}14]. May [STFM12]. MC [BZ09]. MC-Supporter [BZ09]. MDA [AGMT10, BRF^{+}09, PEPP08, STBFM09]. MDD [GAMP10, TTB13]. Me [dARGSB11, AMYH14, SBRS11]. Mean [AT10b]. Meaning [ST05]. Means [Ara97, CSFFM12, CNQ04, FML13, BMM^{+}09, RCGBS13]. Measure [ASH11, CPCLSAGC11, Fon01, Sta05]. Measurement [CI12, FG03, OMO10, Pau09, PKSR09, Sue10, WFOC98, ZWH10]. Measures [CFSC04, RGP10, VV15]. Measuring [RF12, XWGS09]. MECCANO [BBIC13]. Mechanical [TIL08]. Mechanics [Wan95]. Mechanism [AR01, Bur05, FOAB08, HME^{+}06, HCH^{+}09, LM10, OO08]. Mechanisms [ACR11, dMBHR15, CMSE09, Die10, GL11d, LWS11, NML09]. Media [AMA^{+}14, Bür08, DIKL14, FP12, GGMM^{+}13, KMR96, KKTT16, LTY^{+}16, LMA^{+}14, LAHZ^{+}15, MLHCGB16, SZWdP14, Tin16]. Median [CS02]. Mediated [LGZ01, OCW13, WLKW11]. Mediating [Car95, KW10]. Mediation [BCG^{+}14]. Mediator [BBGV07]. Medical [HKL^{+}06, HGMT08, KBF^{+}11, MNS^{+}12, OOHS06]. Medicine [LEJ^{+}08]. Medieval [AGA12, BVG08]. ‘MEDIP [HKL^{+}06]. ‘MEDIP-Platform [HKL^{+}06]. Medium [SPRP09]. Medium-sized [SPRP09]. mEducator [DKD^{+}13]. Meeting [Güt08, Lin11a, YYZ^{+}09]. Meetings [ZAB^{+}08]. Meets [TKK^{+}12, KZ03]. MEM [CW00]. MEM-Slicing [CW00]. Member [Tom03]. Membership [Bol06]. Membrane [ABCP02, DP99, NS06, NC04, PPJ04, ZZH^{+}12]. Membranes [PI04]. Memories [KA97, Shu97]. Memory [BD00, Gio98, HCD10, JR02, KM13, KKH12, KNSN07, MM96, ŞM96, WTA01, dKR03]. Memory-Intensive [KKH12]. Mental [Dit02, Kom02]. Menus [BH14]. Merge [LM10]. Merging [AT13, SMMC10]. Mesh [Lep95, LWY11, WXZL15, YLW^{+}14]. Message [CT04, HF01, PSS^{+}13, RMGT09]. Message-Minimizing [HF01]. Message-Optimal [CT04]. Message-Passing [RMGT09]. Messaging [LGZ01]. Meta [CXB12, Kar13, KJZJ08, LSR10, LKB^{+}02]. Meta-Cognitive [Kar13]. Meta-Data [LSR10, LKB^{+}02]. Meta-Model [CXB12]. Metaclasses [FOAB08]. Metadata [ARN04, ACB02, BAP^{+}16, BVV^{+}10, Duv01, LVV10, MGPB07, STVT07, SKL08, SM04]. Metadirectory [FVIG12]. Metaheuristic [BCHM12, dG15]. Metaheuristics [MAT08]. MetaJ [dBdd04]. Metal [LH03]. Metamodel [GRS08]. Metamodel-based [GRS08]. Metamodeling [ESB04, IPCVC12]. Metaphors [SH96].

Metaprogramming [dBdd04]. Metasynthesis [CN08a]. Method [AT13, APJK09, BT08, BCG^{+}09, Bör02, BJMBA15, BS12b, Cap05, Che09, CSF99, HB00, HSR10, Hon95, Jun01, Kie05b, Kwo97, LM03a, LHK^{+}13, LKT10, LdSM08, MPF04, NNT16, OF13, Ois98, PZLAS^{+}13, PRBLAP^{+}13, POR10, PdP^{+}04, PLSF08, RGP10, RdKO11, SZZM10, SBCD15, TH99, Tddd03, THJ16, VL14, Wan95, WH08, WZZ^{+}09, XWGS09, dlFVPHB15]. Methodological [FSdRSS11, FBSEGP15, FG03, LA03a, MLHCGB16]. Methodologies [AG06, DJJN09, GR14, GSP04, NdMM12, RC10, RGPK15, SR00, UFF12]. Methodology [BZA08, FP05, GPCZ^{+}13, Hel07, KÜ10, LULGFC13, MROH08, MS01, PKP08, PBB07, SNAF07, ZPFG03]. Methods [Abr07, AV07, ACAMM15, AS07, Ban07, BD06, BCC^{+}06, BEH^{+}05, BZA08, Faz06, Hal07, Hei07, HTHW12, HT13, JCB15, KLT13, LVS13, MS00, MHLB12, ML98, Rih98, Rus07, dPSZPVLR^{+}16, STVT07, Ste00, Suz06, Ver08, VO10]. Metric [BB10a, Dru06, Fal10, Ilj09, Leš09, SLN16, SDÖ^{+}12, Sim07, WSL07]. Metric-Entropy [Sim07]. Metrics [CI12, EGDG09, MHH98, MR14, SFP12, TR98]. Metropolitan [HZZ^{+}12]. Mexico [HAS^{+}07]. MFFP [LHC^{+}13]. MFFP-tree [LHC^{+}13]. mHaskell [RTL05]. Microblog [TCW12]. Microblogging [MS11]. Microprocessor [BRR99, OCRPdlMG07]. Microprocessor-based [BRR99]. Microstructure [CO08]. Microworlds [Tom95]. Middle [PHJ^{+}08, WPL98]. Middleware [CMMP16, FG10, LSG^{+}14, MDO^{+}09, MSSV14, PRCRARLN10]. Middleware-Oriented [MSSV14]. Migration [CE06, KL09a]. Mihajlo [JV11]. MINCE [AMS04]. Mind [WP15]. Minimal [CL97, KHN99, RR06a, Wat02]. Minimalist [GJAB95]. Minimisation [BC16]. Minimising [EACGFK13]. Minimization [Köh09, LWY11]. Minimizing [HF01]. Mining [ARFT05, AK09, AP05, BGMR^{+}16, BHQW02, Bor07, CdSCSSA16, CCS00, FTARR05a, GKZ05, HGMT08, Jun05b, KDKB07, LLS05, LHC^{+}13, RKJ16, RP08, SP16, SL10, SZS12, Suz06, TB16, VGCPAH16, VC13, VTHM16, WKXL05, Yan05, ZTX^{+}07]. ministerial [PS97]. Minor [CDF97, Din97]. MIRACLE [MSC03, SCLM03]. Mirror [Bai12]. Mirroring [PRB^{+}11]. Misbehaviour [MGCGCG12]. Mismatch [GG08]. Mismatching [CSC08]. Missile [MCC13]. Misuse [VdR09]. Mitigate [BNCGD^{+}11]. Mixed [GLS00, GRHMM^{+}15, MA13, PdlCBKN14]. Mixed-Critical [GRHMM^{+}15]. MLab [AAAK15]. Mnemonics [SHH10]. mobiDIÁK [ABFJ06]. Mobile [AMA^{+}14, AHEAS^{+}15, AMR^{+}14, AAAK15, AY12, BZ09, BBIC13, Bic15, Bra15, CGLdMAC14, CPLPW15, Cai10, CS10, CDG14, CGPAP13, DAk13, DIKL14, EMZB14, FPLS03, Fra98, GP10, HCBB15, HOPN11, HHHX09, Jun10b, KSK^{+}16, KK13b, Kim10, KH12, KW10, LKK08, LEC11, LZ09, LS10, LULGFC13, LGP10, LZZK14, MM15, NOP08, OOHS06, OAR^{+}14, OO08, OEK16, Pal15, PP99, Pob11b, QGT^{+}14, RTL05, RKH15, RFMLP10, SBAZ11, Sat10, SG02, SZWdP14, SM02b, TBL15, UHÖD15, VBVNHdDSL12, VVM^{+}06, VV12, XNKG15, YJY14, Yon11, ZMAS10, ZABB07, vBK08, JFL^{+}13, RTL06]. Mobile-Enabled [BBIC13]. Mobile-Web [Kim10]. Mobility [GGDBP^{+}08, Gün02, HZZ^{+}12, HA13, RTL06]. Moby [OD03]. Moby/RT [OD03]. Modal [BGBA10, Cre09, FDC^{+}13]. Model [AAM14, ASH11, APDC08, Arr07, BS12a, BAPG03, BB04, BCM12, BdGFMT14, BCD13, BE98, BQV14, CCCP08, CNQ04, CXB12, CBN^{+}06, CPRT05, CS07, CBNDR10, CM11, CSA10, CKdL08, CL07, CCP11, DT12, DSCT10, DMS05, ESM08, EMGB^{+}12, FPT10, FBSEGP15, FL14, GH08, GLCV08, GMB11, GXC^{+}15, GGMM^{+}13, HK15, HC08, HAFS15, HBF10, Hor04, HZZ^{+}12, HIGM13, HLC08, Ish00, JJ12, JL09, JK10b, KASS03, KY10, KCKL10, Kir09, KBN14, Kom02, KJL09, LT09, LPSF10, LASL12, LLZ16, Ma10, MYCA11, MBA12, MN14, Mar07, MMEdP12, MGM^{+}08, MSSY95, MG14, MS10, MT02, Meh02, MP10, MR08, MR14, MMS11, MGPB07, ND08, NBGS06, NWDX09, OBO09, PEPP08, PRCRARLN10, PRCCS13, PO11, PS04, RSVR01, SMGMT09, SMK^{+}04, SMSdB05, Ski00, SV08, SW13, SDLM14, SMV08, SJ13, SS15]. Model [Sto99, SK04, TGLP10, TKSL05, TPC^{+}12, TLR09, Tra13, VAPM12, VJTJ07, WTA01, Win97b, ZTX^{+}07, dLH08, dMTS^{+}14, vBK08]. Model-Based [GGMM^{+}13, MR08]. Model-Checking [SV08, TLR09]. Model-Driven [CKdL08, GLCV08, KCKL10, LPSF10, PRCCS13, SMGMT09, SMV08, SJ13, VAPM12, APDC08, MP10, TGLP10]. Modeling [Ada06, AY00, Ara03, BS06, BH01, BJ97, CRMLN^{+}07, CJO^{+}13, ČSŽ09, DRS06, FWT11, HR06, JKKW16, KNLS00, KT10, KKK^{+}14, KSdV09, LNML03, Mai05, ME03, NSMBACBG12, PF11, RY09, SK13, SHK10, Sch99, SBCJ03, SZS12, SKH12, SK04, TWW07, TNM09, Træ08, WKSD^{+}11, WSF08, dAO13, dOLC^{+}07]. Modelling [BEH^{+}05, BS08, BFN05, CSAC^{+}15, CRLNAR05, DKL10, DH10b, FGSW14, Fod06, GGPTdP11, HLK09b, KS10, MSTW12, MMS08, MOMGSRFM07, MMS11, NVB12, PA12, PKP08, PD99, PO04, Tha10, TTB09, VCB08, VO02, WFOC98, vB96]. Models [ARN04, AGGH08, Bai05, BGBA10, BARR09, BST09, BCNR07, CdSCSSA16, DTG10, DZ08, Die96, Dit02, EMZB14, EGG^{+}01, ELS04, Flo04, FMT^{+}15, GOM^{+}13, GMB08, GMC^{+}08a, Gol05, dARGSB11, KCKL10, Koó06, LDSG09, PLG^{+}08, QGT^{+}14, RRB03, SSM11, SMSdB05, SSGS10, TD96, WL13, Yeu04]. ModelSec [SMGMT09]. Modern [AGA12, GHS06, Lin04a, NS05]. Modes [VDSF98, WLKW11]. Modification [DDG97]. Modified [ACAMM15]. Modular [CB04, DMMM95, HLP^{+}13, PdCdlTR06, SdBM05, VBVNHdDSL12]. Modularity [CS04, NKS^{+}09]. Modularizing [BB08b]. Module [BDN05, HNS07, MSA13, Win97a]. Module-centralized [MSA13]. Modules [IBN^{+}11]. Modulo [Gon06]. MOF [ESB04]. MOF-Based [ESB04]. Moisil [Ior00]. Molecular [Mac96]. Moments [JM10]. MOMIS [VBB13]. Monadic [CDF97, GL00, MP08]. Monitoring [AG06, CKPK13, Coo06, DHC11, DS03, Dru06, GRGN13, GK06, JFL^{+}13, LGAP11, MSA13, SVV10, TYSY09, ZHG^{+}06]. Monitors [JT05]. Monkeys [CG09]. Monoidal [DG07]. Monopoly [KB06]. Monotone [Dun96, FH00]. MONSTR [BBP95, Ban96, Ban97a, Ban97b]. Montages [KP97b]. MOOC [FBSEGP15]. MOOCs [AHPSCDK14, AHEAS^{+}15, SGLM16]. Moodle [BTD^{+}07]. Morphological [MMP15]. Morris [Ukk10]. Motifs [LHK^{+}13]. Motion [DRRGdP07, EC00, KKK^{+}14, ME03]. Motivation [RKH15, STFM12]. Motor [GRHMM^{+}15, KKK^{+}14]. Mounting [LLLL99]. Mouse [PRT^{+}08]. Movements [PRT^{+}08]. Movies [TJ15]. Moving [Lep98, vdV08]. MP [CSA10]. MPEG [HM01, LKB^{+}02]. MPEG-7 [LKB^{+}02]. MS [EHEH05]. MS-RTCP [EHEH05]. MSB [NK95]. MSB-First [NK95]. MSparc [MN00]. mStar [SLPS98]. MTAC [For97]. Multi [AT10b, ACP06, ADMdB09, BCA^{+}10, BCC^{+}06, BM99, ČSŽ09, DLY08, DM10, EMZB14, dAFL07, GRGN13, GSP04, HJRW97, IO04, JMP06, KK13a, KKH12, KZ08, LLLL99, LWG14, LPSF10, Lin04c, LS10, MKS09, MAT08, MK12, NKS^{+}09, PTO^{+}12, PKP08, PT09, PB05, Pel14, RSW04, RLL^{+}10, SDÖ^{+}12, SW09, SZS12, TJS^{+}13, VL14, WH08, WHCL09, Wei08, ZC09, ZQQ15]. Multi-adjoint [JMP06]. Multi-Agent [ADMdB09, ČSŽ09, GSP04, KZ08, NKS^{+}09, PKP08, PT09, PB05]. Multi-Agents [GRGN13]. Multi-Authority [ZQQ15]. Multi-Combinators [Lin04c]. Multi-Core [KKH12]. Multi-Criteria [IO04, RLL^{+}10]. Multi-Device [LPSF10]. Multi-DSP [RSW04]. Multi-Functions [Wei08]. Multi-head [LLLL99]. Multi-Human [AT10b]. Multi-label [KK13a]. Multi-Layer [VL14, WH08]. Multi-Level [BCA^{+}10]. Multi-Objective [BCC^{+}06, EMZB14, MAT08, SDÖ^{+}12, ACP06, LWG14, PTO^{+}12, WHCL09]. Multi-Processor [dAFL07]. Multi-Purpose [LS10]. Multi-Relational [SZS12]. Multi-represented [Wei08]. Multi-Secret [BM99]. Multi-Selectivity [HJRW97]. Multi-Strategy [MK12]. Multi-sub-swarm [ZC09]. Multi-Terminal [DM10]. Multi-user [DLY08, Pel14, SW09, TJS^{+}13]. Multiagent [TYSY09]. Multicriteria [CFMP15]. Multidevice [PLSF08]. Multidimensional [HAI13]. Multidisciplinary [DH10b]. Multifaceted [LWH09, MJG15]. Multilayer [RMGCGCF08, ZC09]. Multilevel [FMS12, ZDE14]. Multilingual [dFBNGS^{+}14, JST11, JMEL10]. Multimedia [AFP04, Azz10, BG98, HM01, HKS96, HMMP10, IBN^{+}11, Jun10a, Kol05, Kou09, LM94b, LZGC09, LZ09, LLYC12, MSC03, MH98, Mül02, QZYL11, SLD^{+}16, SHZ^{+}10, WLLL09, YLZW10]. Multimedia-Based [Azz10, HKS96]. Multimodal [MYY06, MMD12, PANS13, SMMM13]. Multipath [SKH12]. Multiple [Alm06, GWG96, HFIBJ13, HLK09a, LWS11, ML02, PJN13, SCK^{+}09, Wol99a, Wol99b, Wol00, XMZbL10, dG15]. Multiplication [AHRH08, MJS13]. Multiplier [MJS13]. Multiply [BAPG03]. Multiprocessors [KU00]. Multirate [Bai12]. MultiResolution [KS05]. Multisecret [DLR97]. Multisecret-Sharing [DLR97]. Multithreaded [For97, GN00, KU00, SU01]. Multithreading [BD00, MN00]. Multitrail [CRB15]. Multiuser [CAJ06]. Multivariate [ML98]. Municipal [LTY^{+}16, SC14]. Muscle [BS11]. Museum [LN08]. MUSIC [LZGC09]. Mutual [CE06]. MUVEs [STFM12]. MV [Ior00]. my [GMP^{+}13]. MyLearningMentor [AHEAS^{+}15].

Nabuco [Lin11b]. Naive [FDR^{+}15, Alm06]. Name [Kim12]. Name-card [Kim12]. Named [JM15, THJ16]. Names [BQBW03]. NASDAQ [DBBS08]. Nash [KK10, Pau10]. National [MRK^{+}98]. Natural [CM07, MGMS12, RAC10, RGHH97, Wol00, dlELR09]. Naturals [Pet12]. Navigable [THS11]. Navigation [AMR^{+}14, CJH12, RTB13, SdBC13]. Near [SHH10]. Nearly [Bos08a]. Necessary [Roj96]. Need [BARR09, Bür08, Pre97]. Needs [SdOB09]. Negated [BH00]. Negative [MVMRULS12, Pre97, Pre12]. Negotiating [Shu97]. Negotiation [BGP07, BEH08, IO04, MK12, OPP09, RMMLBLGS09]. Negotiations [LS07]. Neighborhood [ACAMM15, GMS03, HKTV06, LA07, MGMS12]. Net [DT09, Jun01, She96]. Nets [AY00, CC07, DT09, FFK04, IDS02, RPR11, SBCD15, TEK08, TLR09, Zou06, CTM10]. Network [AKM95, ACP06, BBGV07, BMM^{+}09, BvdTV09, CSAC^{+}15, CW09, CM11, EK00, HHHX09, HHH^{+}02, Jun05c, JK12a, KASN08, LDO^{+}12, LLYC12, LCZ^{+}12, Log04, LSV06, LNH^{+}15, MJGS12, MAT08, MPF04, NdMM12, OFCB08, PZDH09, PZJ09, PCKJ11, RHM15, SLD^{+}16, SHZ^{+}10, SKH^{+}10, SZS12, TWW07, Tru10, Vie03, VVM^{+}06, WZC07, WP15, XHP^{+}09, Yan05]. Network-based [VVM^{+}06]. Network-on-Chip [ACP06, KASN08, NdMM12]. Network1 [Jun08]. Networked [KKK^{+}14, KP95, MPL11, SBRS11]. Networking [BU13, CJH12, JKKW16, JK12b, Kim12, SA14, SC14, WWD15]. Networks [AHT09, AMBP04, BBIC13, Bic15, BMV12, CAD^{+}06, CPHC11, CWT^{+}15, CHH16, CL08, DKL10, DH10b, DMM07, EMZB14, Fal10, FSdRSS11, FP05, FMS12, FLF^{+}14, GSMBFPK10, HKKvP08, HCH^{+}09, Jun10b, KK13b, KTJ05, KF10a, Kou09, LGAP11, LKK08, LZ09, LCZ^{+}12, LWY11, LZZK14, MWM10, MdCRMP14, MH02, MC07, MMB08, ND08, NVB12, NPC^{+}09, OP15, Ozd13, Pal15, PTO^{+}12, PA12, PdCdlTR06, QZ07, Saf08, SESMT10, SHZ^{+}10, SBG^{+}12, SP16, SvRvdV^{+}13, SSdS^{+}11, SSS12, Tat07, TR10, TCW12, TB16, TJ15, VAS05, WF12, WXZL15, YLW^{+}14, CMZZ07]. Networks-on-Chip [PTO^{+}12]. Neumann [Roj96]. Neural [BS01, CW09, CM11, FGS98, Log04, Log06, MMP15, PA12, Pau07, TWW07, WZC07, XHP^{+}09]. Neutral [MVMRULS12]. New-Product [ACR11]. News [ASAAASJ16, AAGU97]. Newspapers [BARR09]. Newton [Lan98, SZZM10]. Next [CDCH09, LZZK14, LMA^{+}14, LAHZ^{+}15, MJGS12, MS94, PD10, Sch96]. NFAs [vZG11]. NIKVision [MBC13]. Nimbus [TWH00]. NMAC [RR08]. NMAC/HMAC [RR08]. no [BCG^{+}99]. NoC [BARB12, JNdMM12]. Node [HJZ07].

Nodes [BBC02, CBG04, PdCdlTR06]. Noisy [Świ07]. Non [AC05, And97, BH14, CBR^{+}05, CTM10, Cre09, Gra98, HGIPCPPM11, KH12, KO99, MMP15, OAR^{+}14, OO08, Rad14, RGRR15, RMMLBLGS09, TBL15]. Non-blocking [AC05]. Non-classical [KO99]. Non-Denumerable [Cre09]. Non-Deterministic [And97]. Non-explicit [OAR^{+}14]. Non-formal [TBL15]. Non-Functional [Rad14, CTM10]. Non-Invasive [CBR^{+}05]. Non-linear [Gra98]. Non-Marker [KH12]. Non-Photorealistic [MMP15]. Non-repudiation [OO08, RMMLBLGS09]. Non-shoppers [HGIPCPPM11]. Non-Standard [BH14]. Non-traditional [RGRR15]. Nonaka [Sch03]. Nonblocking [Gro09]. Nondeterminism [Gup01, HNP98]. Nondeterministic [Mul00, VF11]. Nonexpansive [Leu07]. Nonhomomorphicity [ZZ00]. Noninterference [LF16]. Nonlinear [KSY97, NO98, SI00, SZZM10]. Nonlinearity [SZZ95]. Nonperfect [OK98]. Nonrandom [Her05]. Nonstrict [MABS05]. norm [DG07]. Normal [Her02, ND08]. Normative [ADMdB09]. North [UHÖD15]. Note [BSHI99, CFSC04, CDF97, Fen15, HNP98, Kuh03, Lan10, RK97, Sch10]. Notes [Bra02, Vai00, VPF09]. Notion [Kwo97]. Notions [Bos08b, BH08]. Novel [AAM14, AAAG95, BM05, KMN16, MJGS12, RKJ16, TIL08, VAS05, WH08, ZC09, ZZH^{+}12]. NP [Bod01, ECHS10, Gaß09, Mac96]. NP-completeness [ECHS10]. NP-hardness [Bod01]. NQL [GK13]. NQL-Complete [GK13]. Number [AC07, Gon06, Han10, Hon95, MSHN06, RR06a, RR06b, Was98b]. numbering [AR95]. Numbers [CN97, Her96, Her02, MGMS12, RK97]. Numeric [Log04]. Numerical [FTARR05b, Gal98, Gra98, NHH06, Ois98].

O [AO04]. O-O [AO04]. Oberon [KP97a]. Obesity [dLMVGM13]. Object [BDPSNG97, CBN^{+}06, CM03, CLC09, DF00, DGL03, EKP03, FR04, Fro02, GMC08b, GHM04, HKKvP08, KH12, LN02, MMdMGM06, MCM07, NuR05, NR08a, NSF^{+}10, RRB03, Tom95, dFCC07]. Object-based [LN02]. Object-Oriented [BDPSNG97, CM03, CLC09, DF00, FR04, GHM04, MCM07, NuR05, NR08a, dFCC07, Fro02]. Objective [BCC^{+}06, EMZB14, MAT08, SDÖ^{+}12, ACP06, LWG14, PTO^{+}12, WHCL09]. Objectives [LAAPVGMM15]. Objects [Ara03, BAZ14, BRH^{+}08, BBIC13, CdSCSSA16, CBO05, DDSS05, DOOJ95, DHO98, FPLS03, GSdSB16, GHO10, LdL07, MC07, NSFVH05, vD05]. Oblivious [CT08b]. Observability [BJ05b]. Observable [LF16, VO02]. Observation [Ber06, BRS00a, NHH06]. Observation-oriented [BRS00a]. Observational [GlRBdSG11]. Observations [CJZ13, HSFE12, Świ07]. Obstacles [SL96]. Obstruction [CDF97]. Obstructions [Din97]. Obtaining [HVCA12]. Occupational [GLD^{+}12]. occurrence [TJ15]. OCL [MN14]. OCP [NBGS06]. OCR [BNCGD^{+}11, RdKO11]. ODEDialect [CGP07b]. ODEs [Rih98]. ODL [MRK^{+}98]. ODR [PCLCC11]. OE [Kuh03]. OE-sales [Kuh03]. OER [TPC^{+}12]. Off [BCZ04, OJSB08, ACL95, Jéz95, RS00]. Off-line [OJSB08, RS00]. Office [FZ00, MTB^{+}08]. Offline [RR11]. Offs [GFBR08]. Offshore [MPM12]. Okul’ [Tüf13]. OLAP [BdGFMT14, LT09]. OLAP-OLTP [LT09]. Old [Fel01, MS01]. OLSR [BAML07, NM07]. OLTP [LT09]. Omega [Her96, CN97]. Omega-Words [Her96]. On-Chip-Pipelining [KKH12]. On-line [Dru06, Hen98, LS95]. On-the-fly [Hor04]. One [Lan10, SCS13, She96, LKT10, Viv96]. One-way [Lan10]. Online [BM13, GRGPL08, HAS^{+}07, HMHGR15, JFZ09, JMKT12, JK12b, KASN08, LCC^{+}12, LLM02, LMMRV^{+}15, LLS05, LM15, MX05, MR11, OP15, PJH12, PCLT15, RR11, RGPK15, SGLM16, SL10, Ste08, WKTL01]. only [LLS05]. Ontological [Arr07, HBF10, HBFV13, MBA12, MGPB07]. Ontologies [AT13, DDSS05, FMB^{+}11, FSMP07, FL14, GGPTdP11, KPV^{+}11, PCLCC11, QQ11, RCGBS13, Ros05, SSBS08, VSML03]. Ontology [Arr07, BS03, CXB12, CYL11, CHH16, CDD^{+}03, CGP07b, DDS04, DHC11, DJJN09, GBHA12, JV11, KDKB07, KJ10, LSR10, LSX12, LGGGC14, MMS08, MGNDAM12, MUSA03, NDAM09, OCW13, RBLR02, SST07, Sha11, Sto03, Tar12, VAH07, XCJ13, XWGS09, ZWH10]. Ontology-Based [BS03, CDD^{+}03, MUSA03, RBLR02, CXB12, DDS04, DHC11, JV11, KDKB07, NDAM09, OCW13, Sha11, Sto03, Tar12, VAH07]. Ontoolcole [VGBLGS^{+}08]. OntoShare [DDS04]. OntoUML [BGBA10]. OO [GHM04]. OO-ASIPs [GHM04]. Open [BFF11, Cal96b, CVFN07, DDJ^{+}11, EIH08, Fra98, FWS^{+}11, GR02, HAI13, IAS16, KPV^{+}11, LMMRV^{+}15, MMD12, OP15, Pal15, PB05, PCLT15, QFB^{+}14, RGPK15, SVK^{+}15, SDLM14, SvRS14, TTB09, TSCY01, Pel14]. Open-Ended [HAI13]. Open-Vocabulary [EIH08]. OpenCourseWare [TPC^{+}12]. Opening [GPCAC11]. OpenMP [LMRG14]. Openness [DD13]. OpenRISC [MJS13]. OpenSim [PMAM14]. Operating [MBC12, NS06, Sue10, WKXL05]. Operation [HHY02, LH03, NXSA12, PRS95]. Operational [JMP06, XCJ13]. Operational/Interpretive [JMP06]. Operationalization [MOMGSRFM07]. operative [Has02]. Operator [JW13, Spi05]. Operators [BD05, SS09b, dH04]. Opinion [BGMR^{+}16, KKTT16, KSR16, SL10, VGCPAH16]. Opinions [MVMRULS12, Tez13]. Opportunistic [FLF^{+}14, Pal15]. Opportunities [Hop98, RBLR02, RSFMJ12, SL96]. Optical [PRT^{+}08]. Optima [KK10]. Optimal [AHT09, BDGW96, BEH08, Bos08a, Cer97, CT04, HHH^{+}02, MAT08, NdMM06b, RRR10, TKF06]. Optimally [Lep95]. Optimisation [EMZB14, TCS^{+}03]. Optimised [Bai12]. Optimistic [DLY08]. Optimization [BMM^{+}09, BS12b, CCM09, DB12, EGK^{+}12, JNdMM12, KZ03, LKP11, LA07, LWY11, ML98, Mes02, NAK08, NdMM08, NI03, OFCB08, SESMT10, SHZ^{+}10, WHCL09, XZSS09, ZC09, ZZH^{+}12]. Optimizations [DR10, SKSP09, dSC06b]. Optimized [CBBT07, KKH12, KR03b]. Optimizing [MTB^{+}08, ZPFG03, dSC06a]. Optimum [CT97, Tom97]. OR-joins [BST09]. Oracles [Gaß09]. Oral [MV15]. Orbits [Gal98, Ois98]. Orchestrating [HLNA^{+}12]. Orchestration [QL12]. Order [CDF97, Din97, FSELC13, Gan04, HPB12, HHH98, RB07, Ste00, ZZ00, DOM10]. Ordered [HN07, Sal10, Vai00]. Ordering [AMS04]. ORE [OCB^{+}10]. Organic [HNYO04]. Organisational [FP05, FG03, Liy02, Shu97, dKR03]. Organisations [Vie03]. Organization [Has01, HA10, IBN^{+}11, WKS^{+}99]. Organizational [CE11, CBG04, TKD^{+}09, dTR03]. Organizations [CSFFM12, Lin04a, MMB08]. Organizing [DAd03]. Orientation [CM09, LD06, PKSR09, VDBNR98]. Oriented [AFL08, BDPSNG97, BEH^{+}05, BCG^{+}14, CBR^{+}05, CM03, CLC09, CBBT07, DF00, FGSW14, FR04, FHJ^{+}99, FLF^{+}14, GPFL12, GMC08b, GHM04, Has02, KFK05, LWL10, LRS^{+}11, MDO^{+}09, MCM07, MSSV14, NuR05, NR08a, NOP08, PTNMC08, PRCCS13, PF11, QZYL11, RRB03, SEK13, SIIJ09, SFVFMN04, TM02a, TM02b, Tom95, URG^{+}13, VLE12, VDSF98, WK05, XZSS09, XCJ13, YLZW10, ZFS98, dCVM12, dFCC07, Ada06, AGMT10, BRS00a, Fro02, KOW01, MSA13, MHA03, SAB99, ZGC^{+}08]. Origins [Bör02]. ORM [Lei08]. ORPMS [DHC11]. Orthogonal [DDS10, GOF05]. Orthography [AGA12]. OTS [OF13]. OTS/CafeOBJ [OF13]. Our [Gün02, dSC06b]. Outcome [MNS^{+}12, MX05]. Outcomes [BRAS^{+}12, CR12, FSELC13, GPCPL12, GPVILN13, GRC15, TPC^{+}12]. Outerplanar [PB14]. Outline [Tra00]. Outsourcing [JJ12]. Overflow [RK97]. Overlapping [BJ97, WF12, YLW^{+}14]. Overlaps [Ist07]. Overload [Sch02c]. Overloading [CVFN07]. Overview [BM03b, LM94a, MR12, Wol99b]. Overweight [dLMVGM13]. OWL [GGPTdP11, Rad14, SST07].

P [EGK^{+}12, Gaß09, Gaß10, MPK04, NXSA12, VTRN12]. P2P [BBP08, CW12, LWY09]. PA [Vai00]. PA-Ordered [Vai00]. Package [LM10]. Packaged [CPSAGPGC12]. Packing [FOSS99, PJRC04]. Page [KC08]. Pages [BCM08, GC14]. Paillier [NSNK05]. Pair [SL09]. Pairs [Sim07]. Palm [BDhKB09]. Panels [VTGA13]. Paperfolding [Cer97]. Papers [BdVG06, BM07, CMZZ07, CSY02b, CI05, CSZ07]. Paradigm [BAZ14, CBR^{+}05, Kul07]. Paradox [LMA^{+}14]. Parallel [AHRH08, BDGW96, FKB^{+}15, JL08, SKA08, Ski97, UDC97, dL03]. Parallelism [PRW95]. Parameter [CA14, Nag06, Świ07]. Parameterized [BLS01, CT16]. Parameters [AdGCD^{+}15, LKP11, Win97a, dLH08]. Parametric [Mat04, TLR09]. Parametrization [LL97]. Parent [Tez13]. Pareto [KK10, NdMM06b, WHCL09]. Pareto-Based [WHCL09]. Pareto-Optimal [NdMM06b]. Parikh [AMVM01, Hon97, MPRS95]. Parsable [MVM00]. Parser [PBB08]. Parsing [SS08, Wol00]. Part [CR00a, Boe97b, Ior07, Ior08, RS01a, RS01b]. Partial [CT04, GT01, Lin04c, WC08]. Partially [GCVRSPGP07, Vai00]. Partially-Ordered [Vai00]. Participants [YYZ^{+}09]. Participating [AHEAS^{+}15]. Participation [GVT11]. Participative [DRA^{+}04]. Particle [DB12, DXZL07, LA07, ZC09, ZZH^{+}12]. Partition [Fer96]. Partition-Limited [Fer96]. Partitioning [CW00]. Partitions [APM04, CS02, HN07]. Partner [Ano07]. Pascal [MKP98]. Passing [RMGT09]. Passive [CC08a]. Passports [NML09]. Password [Shi11]. Passwords [KJL09]. Past [dH04]. Patents [SV95]. Path [BGK02, Bor07, CS10, JBH^{+}10, MPPS95, NLLJ12, Str97]. Path-Finder [BGK02]. Pathfinder [OP15]. Paths [SLT08]. Pathways [VUT^{+}08]. Patient [ZFS98, MBA12]. Patients [GFT09]. Pattern [BP09, BDhKB09, BCC^{+}06, BJMBA15, CSA10, Dru12, LLCN09, LKZK10, MPL11, RKJ16, Sch06, Sha11, SSSS10, Ukk10, Ver10, Xi03, XZSS09]. Pattern-driven [BP09]. Pattern-Matching [SSSS10]. Pattern-Oriented [XZSS09]. Patterns [CO08, HI00, HL03, JFZ09, LWH09, LM15, MCM07, MP06, PRB^{+}11, PK98, RCGBS13, TL11, ZAB^{+}08]. Pavelka [GL00]. Pay [PTL11]. Pay-As-You-Go [PTL11]. Payment [DMG07, OO08]. PC [ON97, OT16]. PCs [LRB16, MO03]. PDAs [MRO05, ZAB^{+}08]. PDE [SS09b, WHCL09]. PDE-PEDA [WHCL09]. PDFs [HPB12].

PEDA [WHCL09]. Pedagogical [CM07, CD13, RP98, WBS12]. Pedagogy [DA13, VLE12]. Peer [EACGFK13, PZJ09, QF12]. Peer-to-Peer [PZJ09, QF12]. Peers [SBRS11]. Pentagonal [Mar02b]. People [FPS^{+}12, GRGN13, SdOB09, TM02a]. People-Oriented [TM02a]. Perceived [uRLFH^{+}13]. Perceiving [CBNDR10]. Perception [RF12, VCB08]. Perceptions [Len00, Ozd13]. Perceptive [GHHE^{+}08]. Perceptron [RMGCGCF08]. Perfect [BR07, BS03, JR96, She05]. Performance [BS06, CGLdMAC14, DR10, DL99, EHEH05, FPSFCG07, FSdRSS11, GN00, HVM00, HMHGR15, JS16, KAG00, KZ08, Man97, MBC12, NSMBACBG12, PJO15, PRAT09, PJ15, PBTW07, RSW04, SAKAM11, SAA08, SZWdP14, TEC^{+}07, YJY14, Zim01]. Performance-Energy [YJY14]. Perhaps [Vel05]. Periodic [Gal98]. Peripheral [WP03]. Perlman [HNS07]. Permanent [LdL07]. Permission [JL16, SWY09]. Permission-based [JL16]. Permission-Role [SWY09]. Permutability [Neg05]. Persian [SSSS06]. Persian/Arabic [SSSS06]. Persistency [AT97]. Person [GOM^{+}13, MGT14, PGT09, ZBKK12]. Personal [BZM^{+}10, CKPK13, CGPAP13, GPCAC11, HT01, HH03, Kro13, LV95, MOS^{+}13, NPC^{+}09, RLMS13, RJB10, SLD^{+}16]. Personalisation [PB04]. Personalised [GDW10]. Personality [AWGS04]. Personality-Aware [AWGS04]. Personalization [NPC^{+}09, SMFM05]. Personalized [APJK09, BVV^{+}10, CHH16, ESG10, FGB^{+}14, HCBB15, HB11, HIGM13, YLZW10]. Personnel [Bec03]. Perspective [BEH08, BRS00a, BdI10, CPFSdAS12, DL15, Dro04, FPT10, GHO10, Hul08, JBBH13, JJ12, NL10, SL96, SIIJ09, Ver08]. Perspectives [HNYO04, JL16, MPM12, MC07, OO08]. Persuasive [GFT09]. Perturbation [BB98, FSSPLG^{+}13]. Pervasive [ABM^{+}06, BZM^{+}10, BDhKB09, HCK11, KDGH09, LKHL09, MYY06, QZYL11, Rob06, XHP^{+}09]. Peter [AFK01]. Petri [CC07, DT09, IDS02, RPR11, SBCD15, TEK08, TLR09, Zou06]. PETs [BKH^{+}13]. PEWS [BCFM05, MP06]. Phase [VJTJ07]. Phasetransition [NPB06]. Phasetransition-like [NPB06]. Phone [vBK08]. Photographed [PTL^{+}09]. Photorealistic [MMP15]. PhyMEL [PdlCBKN14]. PhyMEL-WS [PdlCBKN14]. Physical [AY12, DLL16, HLNA^{+}12, HA13, KHLAP12]. Physically [AAJR05, PdlCBKN14]. Physics [CC96, Spa08]. Pi [BBP95, CR00b, PS04]. Pi-Calculus [BBP95, PS04, CR00b]. Pick [HFIBJ13]. PICTAC [CBNDR10]. Picture [Kri99]. Pictures [IGS08]. Piecemeal [GJP^{+}12a]. Pilot [Bec03, BRAS^{+}12, SFP12]. PINE [Saf08]. Pipeline [CGVRGPSP07]. Pipeline-scheduling [CGVRGPSP07]. Pipelining [KKH12, RC07]. Pirate [VSGP05]. Pitfalls [PF15]. Pizzarotti [CBG04]. PKDS [HLCL11]. PKI [OO08]. PLA [JRR16]. Place [Coo06, Jun01, Tab07]. Place/Transition [Jun01]. Place/Transition-Net [Jun01]. Placement [Tab07]. Placing [BJ97]. Plagiarism [MKZ06]. Plagiarisms [PMP02]. Plane [MM99, Mar00b, Mar02b, MS03, Mar06a]. Planner [LL97]. Planning [CPSAGPGC12, EC00, LLLL99, LGL09, LWG14, LST14, LCZ^{+}12, SKSP09]. Plans [Bjø01]. Plant [BPHN06]. Plastic [SLJC08]. Plasticity [SLJC08]. Platform [BCG^{+}14, Die10, GMHGRG^{+}13, GO14, JNdMM12, dIdSGZ10, MKS09, MABS05, PZDH09, PT10, SKHK14, SMMM13, SK04, VPB15, WKTL01, WCH14, YLZW10, ZFS98, HKL^{+}06]. Platforms [BQV14, DR10, Die10, JKKW16, Jun10b, LZZK14, Pal15, dPPRRGSSPP15, SBMD10, SEK13]. Plausible [RAC10]. Played [IGS08]. PLAYER [FPS^{+}12]. Players [BMV12, GGMM^{+}13]. Playful [PANS13]. PLE [MPL11]. PLEASERS [FC03]. POCA [HT06]. PoEML [CRMLN^{+}07, PRCRARLN10]. Poetry [AMYH14]. Point [ATOFF98, AGO^{+}13, DF99, Har07, KWH03, RMZ15, Tab07, Ukk10, DB12]. Pointer [Win97a]. Points [BCG^{+}09, LL12, SMK^{+}04]. Points-to [LL12]. Policy [CPRT05, JL16, MCG14, RLL^{+}10, XMLC13]. Polygonal [DXZL07]. Polygons [AAAG95, BR05]. Polyhexes [BR07]. Polymorphic [CCYK15, VFC03]. Polynomial [CP02, FDR^{+}15, Gon06, Gra98, HJRW97, HHH98, MM99, PŞ95]. Polynomial-Time [HJRW97]. Polynomials [AV07, Akr09, Şte96, Şte05]. Polypodic [Boz99]. Polytime [CPS07]. Popular [RR08]. Population [BG04, BARB12, JS16]. Population-Based [BARB12]. Portable [HG11]. Portal [ABFJ06, HW10, HTHW12, Kim12, MB09, RJB10]. Portals [DDJ^{+}11, GHS06]. Portlets [DR04]. Portugal [Tom03]. Portuguese [GMP^{+}13]. Pose [SHK10, TG10]. Poset [DSLO04]. Position [MWM10]. Position-based [MWM10]. Positioning [JGW11, YKD^{+}08]. Positive [AV07, Akr09, Hav05, Hem99, MVMRULS12, Ozd13, Şte05, XMZbL10]. Positivity [ST05]. Possible [SM02a]. Possibly [Tru10]. post [Sch03]. post-Nonaka [Sch03]. Potential [CR04, Hol96, KR11, LNH^{+}15, RS00]. Potential-functionbased [RS00]. Potentials [BKH^{+}13]. Poverty [LM01]. Poverty-stricken [LM01]. Power [ABCP02, DP99, FMT^{+}15, Hem99, HK95, Hon02, KPS96, MWM10, MVPP02, NWDX09, XMZbL10, Hon97, Hon99]. Powerful [MT02, Sto99]. Powering [CO08]. Practical [AR01, GAMP10, HH03, Kel08, RRM^{+}12, VFC03]. Practicality [PK98]. Practice [Abr07, BKL12, CMSE09, DDS04, Dro04, KA97, Lin04a, MV15, MCMMAP^{+}14, Pre04]. Practices [GlRBdSG11, KT10, MPL11, PRB^{+}11, QFB^{+}14, SL09]. Practitioner [HL09]. Pragmatic [JBH^{+}10, MGAVF10, PPP^{+}11]. PRAM [For97, Lep98]. Pratt [Ukk10]. Pre [BEH^{+}05, Vît05]. Pre-apartness [Vît05]. Pre-built [BEH^{+}05]. Precedence [CT04]. Precise [BFN05]. Precision [CC96, VCB08]. Predation [PLB14]. Predicate [SV08]. Predicated [KR03a]. Predicting [JFZ09]. Prediction [Bor07, CRB15, CLVM09, RSW04, RAC10, WLtHN14]. Predictive [TNM09]. Predictor [Log04]. Predictors [Log06]. Preference [IO04]. Preferences [CFMP15, CTM10, FGB^{+}14]. Prefix [CT97, Gro00]. Prefix-Free [Gro00]. Preparing [Die96]. Preprocessing [Jun05b, WC09]. Prerequisite [HCWA03]. Presburger [AR01]. Preschool [ARS16]. Presence [DLL16]. Presentation [BH14, JMKT12, Kwo97, Moo08, Thi00]. Presentations [SS08]. Preservation [BRH^{+}08, Yon11]. Preserving [PSS^{+}13, Sch08a, TY09, YWD08, LCDP15]. Pressure [Spa08]. Prevent [MGCGCG12]. Prevention [JT05]. Price [BGP07, BEH08, SCW08, de 00]. Pricing [SCW08]. Primary [BAP^{+}16, RF12, Sin06]. Prime [Mue04, MPRS95]. Primes [MJS13]. Primitive [HI99, Mue04]. Primitivity [HHY02]. Principles [DS10, Rat05, RP98, vD05]. Print [BG98, Nør10]. Printed [BMMM14, RdKO11]. PRISM [FP06]. Privacy [BZM^{+}10, BKH^{+}13, KDGH09, LZLW13, LCDP15, MKI^{+}12, PSS^{+}13, RHM15, SG96, TFG06, YWD08, Yon11, YMA15, vSO11]. Privacy-Preserving [YWD08, LCDP15]. Prize [GGS08]. Probabilistic [Bai05, Bos08b, CFSC04, GSMBFPK10, Wol99b, ZS04]. Probability [AK09, PMLL09, Svo95]. Probability-Based [AK09]. Problem [ACP06, BY97, DBAB12, EC00, FOSS99, FSSPLG^{+}13, GRGPL08, GGS08, GBP^{+}08, GK13, HF01, Hon01, Kat05, Kul05, LLH03, LJ00, Neh98, NR08b, QZB^{+}00, RR06b, Sch02c, SA03, Str97, TT98, VUT^{+}08, WPL98]. Problem-Based [WPL98]. Problem-solving [DBAB12]. Problematic [Tom03]. Problems [BCC^{+}06, Cal96b, DB12, ECHS10, Fel01, HI99, Ist07, LA07, Mac96, RWZ09, RR06a, TIL08, ZZH^{+}12, dG15, MKP98]. Procedure [CWT^{+}15, WC09]. Procedures [AR01]. Proceedings [BJ05a, RMF^{+}98]. Process [Aïm98, ARS^{+}08, BB08a, BMa11, BEH^{+}05, BST09, BHC05, BQV14, CHPHV10, Car00, CJH12, DRA^{+}04, Dru13, dFER06, FG10, dARGSB11, GHO10, HSSM^{+}04, HCPdASY14, HP15, IKM03, KFK05, LALS08, LRR04, LWG14, LLZ16, dLMVGM13, MHA03, MM12, NSMBACBG12, PFS07, PJN13, RBB06, RVC12, Şte95, TKD^{+}09, VK03, WL13, WKSD^{+}11, WK05, WLtHN14, ZGC^{+}08]. Process-Oriented [KFK05, MHA03]. Processes [CSFFM12, Cam98, DMCM14, GMK05, JV05, LCHD12, Liy02, MJ13, MMB08, RS05, SH10, SL96, SD97, SMV08, SL05, TWW07, TKSL05, VL14, WKS^{+}99]. ProcessGene [WL13]. Processing [AMUFVI09, CN08a, CN08b, FKB^{+}15, HKTV06, HKL^{+}06, JCB15, KKB12, Kir09, Kri99, KL09b, Lin09, Lin11b, NPB06, Ngu05, NCH16, PS12, WZC07]. Processor [dAFL07, Kah01, SAA08]. Processors [DMM07, KU00, KKH12, MJS13, SU01]. Procrastination [dPPRRGSSPP15].

Product

[ABB14, ACR11, ALHM^{+}14, BB08b, BC11, CKdL08, FKO14, FWT11, FG10, JMKT12, JGM10, JGM^{+}13, LM10, LWS11, NKS^{+}09, RP08, RGHH97, SEK13, dMTS^{+}14]. Production [ARS^{+}08, BCS15, HL96, Mea97, SJ13, Wol00, ZTN^{+}15]. Productivity [ASH11, Ern11, GS12, GL11b]. Products [BG98, Géc02]. Professional [FML13, RS03, Sch96]. Professor [AFK01, CS00, KP01]. Profile [HR06, Tar12]. Profiles [CDD^{+}04, DD13, GAMP10, Kul05, NPC^{+}09, SBMD10, TSE^{+}15]. Profiling [CS14, MV15, SOO97]. PROFIsafe [MM03]. Profiting [Vie03]. ProFusion [GWG96]. Program [FF07, Gle03, KD01, LMMPFVI14, RB06]. Programmable [FPSFCG07, ON97]. Programmes [DSAFW07, Sto02]. Programming [AG06, BdVG06, BM07, BMG^{+}05, CLC09, Esp06, FPLS03, For07, GHHA10, GMC08b, IFd03, KP97b, Kwo97, Lin04b, LF06, LKZK10, ML05, NO98, Pel14, QL12, RBB06, RVW07, SH06, SL09, SS07, She05, SRI08, SGS13, Tom95, Tur04, Ver08, Wei08, WHD04, ZTX^{+}07, dSC06a, dL03, dFCC07, VMdCJ08]. Programs [AO04, AdMGPV06, CP06, CBBT07, CP01, GPSV03, JMP06, LF16, MP08, PSS07, PMP02, RS05, SS08, SS00, SBCJ03]. Progress [KK06]. Progressing [SVV10]. Progressive [BT08]. Project [ATOFF98, Ano07, Bec03, CG96b, CB12, DRA^{+}04, DHC11, GPCZ^{+}13, GS12, LHZS12, MCC13, SvRvdV^{+}13, SvRS14, BRAS^{+}12, DKD^{+}13, FSELC13, FPS^{+}12, GLD^{+}12, HKL^{+}06]. Project-based [SvRvdV^{+}13, SvRS14]. Projection [HFOB08, TPC^{+}12]. Projects [AMC^{+}12, ASHT^{+}16, BM05, CR12, GMHGRG^{+}13, GPCPL12, GPVILN13, OPP09, Sto02, TKD^{+}09, TTB09, VRGPSP05]. Prolog [Ara97]. Promise [Car98a]. Promote [IS10, LVV10]. Promoters [IS04]. Promoters/Inhibitors [IS04]. Promoting [LJLCR13]. Promotion [TPC^{+}12]. Proof [Ara97, Bod01, BRW03, HNS07, OF13, PT13, RH10, Sch05c, SAB99]. Proof-Theoretical [RH10]. Proofs [CP01, RK97]. Propagating [Fer96]. Propagation [Jéz95, KR03a, LLZ16, PJH12, Rat00, SZZ95, Tak03]. Properties [AdMGPV06, Ber06, BCD13, HW97, OK98, OS98, Tak03]. Property [Ban97b]. Proposal [AHPSCDK14, CPLPW15, FSdRSS11, FL14, GPL13, GCC16, GMdMC12, HBF10, RVW07, TEK08]. Proposals [FZ00, VGAPGS^{+}15]. Propositional [CM07, GMS03, HL09, MGMS12]. Pros [von98]. Prospects [KA97, Sto02]. Protecting [LZ09, Pos98]. Protection [KDGH09, TSDP07]. Protein [FSSPLG^{+}13, HI00]. Proteins [AGK^{+}10]. Protocol [DH10a, EHEH05, GBCA12, HNS07, JT05, KTJ05, MWM10, MM03, PSS^{+}13, RMMLBLGS09, Tak06, WTA01]. Protocols [BLS01, Bel08, BH00, CT08a, HJZ07, Kon03, Laf04, MdCRMP14, PZDH09, SKH12]. ProtoMon [JT05]. Prototype [YKD^{+}08]. Prototypes [DAk13]. Prototyping [BQV14, MR08, ON97, RSW04]. Prototyping-Driven [MR08]. Provably [DMS05, NSNK05]. Prover [Pau99]. Provide [EDA14, LMRG14]. Provided [PGDD15]. Providing [LGAP11, MKS09, Pav95, RH10, TSDP07]. Proving [Ber06, KO99, Wan95]. Provision [GK06, MR05, NPC^{+}09, SCW08, VVM^{+}06]. Provisioning [HJVK15, HHH^{+}02, YLZW10]. Proximity [ZABB07]. Proxy [AT07, BS06]. Prune [Gra98]. Pruning [VSML03, ZC09]. Pruning-based [VSML03]. Pseudo [Cet00, FH00]. Pseudo-Boolean [FH00].

Pseudo-Wajsberg [Cet00].

Pseudorandom [MSHN06]. Pseudosimple [Bul95]. PSL [MM98]. PSO [AZMA15, LWH09]. PSO-ANN [LWH09]. PSO-Based [AZMA15]. Psycho [PLB14]. Psycho-linguistic [PLB14]. Psychological [APJK09]. Public [BM96, CSAC^{+}15, HA10, HLC08, LRR04, LCDP15, PGDD15, SDLM14, TJS^{+}13, XNKG15]. Public-Key [XNKG15]. Publication [LNH^{+}15, RB08]. Publications [Jür10]. Publishing [Goe95, KMR96, MS94, Sch96]. Pulse [PJH12]. Pupin [JV11]. Purchase [WWD15]. Purchases [JKKW16]. Pure [BM13, MZ12]. Purely [RTL05]. Purpose [CGVRGPSP07, LS10, NML09]. Purposes [LMO01, vSO11]. Pushdown [CFJS15]. Putting [TCS^{+}03]. Putzmeister [Bec03]. Puzzlement [Car98a]. Puzzles [HGS^{+}08]. PVS [Are02, LDSG09, Tra00, dH00]. Py [DGlS12]. Py-Family [DGlS12].

Qcb [Sch10, Sch09c]. Qcb-spaces [Sch10, Sch09c]. QKD [LZLW13]. QL [GK13]. QoE [LVS13, SMMC10]. QoE-based [SMMC10]. QoP [TFG06]. QoS [ASTL07, AdO11, CI12, Kim12, LWL10, SIIJ09, VVM^{+}06, dLH08]. QoS-based [AdO11]. QoS-Security [CI12]. Quadratic [Akr09, FH00, LA07]. Quadrature [Bai12, Bos08a, Lan98]. Qualifications [GLSD11, GLD^{+}12]. Qualitative [DS08, Fal10, Gio98, MMEdP12, MGAVF10, PRW95, SVFR15]. Qualitative-Metric [Fal10]. Quality [BPHN06, BQV14, CCCP08, CSFFM12, CdSCSSA16, CLM10, CC08a, GWW05, GS12, JBH^{+}10, Kil08, KJKS14, uRLFH^{+}13, MM15, MGPB07, NM07, PB05, PF11, RRR10, SGB^{+}13, TFG06, dCH11]. Quality-Investment [RRR10]. Quality-of-Experience [uRLFH^{+}13]. Quanta [SKSN07]. Quantified [Rat00]. Quantitative [Gio98, LdSM08]. Quantum [ANdMM08, KK06, NZM09, Svo95, Svo96]. Quantum-Inspired [ANdMM08, NZM09]. Quasi [Géc02, Sch10, CSY02a]. Quasi-Distances [CSY02a]. Quasi-Products [Géc02]. Quasi-zero-dimensional [Sch10]. Queries [BDM15, Kie05b, MHA^{+}15, Saf08, STVT07, SCT09]. Query [CCM09, HHH98, JR02, Jun08, KF10a, KF10b, KL09b, Ma10, MMdMGM06, PRAT09, VF11]. Query-Driven [KF10a, KF10b]. Querying [NNT16]. Question [Gaß09, HPE14, SLD^{+}16]. Question-Answering [SLD^{+}16]. Question-Driven [HPE14]. Questions [Aur01, HAI13, IRMK12]. Queue [AC05, CE06]. Queuing [FMT^{+}15]. Quick [NNT16]. QUM [SGB^{+}13]. Quo [BP08]. Quorum [ES05]. Quotient [Fen95b, Pal05]. Quotient-digit [Fen95b]. QVT [MYCA11, STBFM09].

R [LMA^{+}14, LCZ^{+}12]. R-evolutions [LMA^{+}14]. r.e [Ilj09]. Radix [LS95, Fen95b]. RADS [NAK08]. RAHIM [LGAP11]. Railroad [Ste00]. RAM [SMK^{+}04]. RAMs [BG97]. Random [Cal95, Dru13, HW97, Her05, Ist07, OS98, VJTJ07]. Randomized [Hem06, RR06b]. Randomness [Blõ06, BM99, RA06, Şte96]. range [DMMM95]. RankFeed [Kie05b]. Ranking [Fen97, GLSD11, TKT07, WC08, APJK09, FGB^{+}14]. Rapid [HLP^{+}13, IS10, MOG^{+}10, RSW04]. Rate [HSFE12, HCK11, KM07]. Rates [Leu07]. Rational [Cla05]. Rationally [ÉK02]. Rationals [KM95]. RAVEN [Ruf01]. Raw [RMM^{+}08]. RBAC [Ma12, MP09]. RBSAD [Oli12]. RCPSP [JRR16]. RDBMS [HVM00]. RDBMS-WWW [HVM00]. RDF [TL11, TKT07]. Re [APJK09, FGB^{+}14, FBCMR15, TT98, BV07]. RE-AspectLua [BV07]. Re-Examining [TT98]. Re-ranking [APJK09, FGB^{+}14].

Re-Use [FBCMR15]. Reachability [GN10, Tat07]. Reactive [CMM01, CS04, HJZ07, LRI03]. Read [Nør10]. Read-eval-print [Nør10]. Readers [ASAAASJ16]. Readiness [NVB12, QFB^{+}14]. Reading [CL95, HPB12, SS00]. Real [AdMGPV06, AT07, Bai12, BS11, BNA15, CP01, DRRGdP07, DDJ^{+}11, DHC11, EGK^{+}12, EHEH05, EGDG09, GL11a, Her96, HZZ^{+}12, Kim10, KCKL10, LZLW13, MN00, OD03, ODSO11, OL08, PMLL09, RSVR01, Ruf01, SF00, Ste00, SSBS08, Tra13, XZ00, ZHG^{+}06]. Real-Time [AdMGPV06, EHEH05, EGDG09, KCKL10, MN00, OD03, ODSO11, OL08, PMLL09, RSVR01, Ruf01, SF00, Ste00, AT07, Bai12, DRRGdP07, DHC11, GL11a, Kim10, LZLW13, ZHG^{+}06]. Real-valued [BNA15]. Real-World [SSBS08]. Realisability [Ber10]. Realising [Hal07, KR11]. Realistic [KP97b]. Reality [DLL16, DAk13, GCL^{+}13, HFIBJ13, HBFV13, KM13, KH12, LAHZ^{+}15, MA13, MOG^{+}10, MCMMAP^{+}14, PdlCBKN14, TKK^{+}12]. Realization [CW09, HW10]. Reals [Gaß10]. Realtime [BvZH09]. Rearranging [BB09]. Reasoning [And96, AK05, BS03, Cam98, Car95, DHP03, DS08, Fal10, Gro09, GALR02, GALR03, GB10, LNML03, MLX10, PGT09, Poo03, RH10, SA97, Sch02d, Tak03, VdR09, Wol99b, ML02]. Reasons [vSO11]. Rebeca [SMSdB05]. Recognising [KC08]. Recognition [AGO^{+}13, BT08, EIH08, FOAB08, GPCZ^{+}13, JM15, KH12, MR11, OCW13, RR11, THJ16, VJ09]. Recommendation [GGP08a, GCC16, Jun05c, Kie05b, LS07, PCKJ11, PJN13, SBGI14, SKH^{+}10, TCW12, VPB15, WZZ^{+}09, SBRS11]. Recommendations [BKL12, CHH16, MKS09, SLD^{+}16]. Recommender [AM11, CAGMPGdAS13, CP15, CS14, ESG10, HIGM13, KTL^{+}11, LGMM^{+}13, PJO15, RGRR15, SMP^{+}11, VDLG10]. Recommenders [MJG15]. Recommending [BAP^{+}16, DDJ^{+}11, LNH^{+}15]. Reconciling [Dus05, HNJ^{+}10]. Reconfigurable [CFF^{+}13, FPSFCG07, GCVRSPGP07, Oli12, SK13, VRGPSP07]. Reconfigurable/Programmable [FPSFCG07]. Reconfiguration [PBTW07]. Reconstruction [Shu97]. Recordings [Fen97]. Recoverable [NZCG05]. Recovering [PK98]. Recovery [FR04, GSW97]. Rectangles [BJ97]. Rectangular [MS03]. Rectilinear [ECHS10]. Recurrence [SW13]. Recurrent [CW09, TWW07]. Recursion [VFC03, YTM05]. Recursive [CP02, GS97]. Recursively [Ars97]. Recycling [MH02]. Redials [GGDBP^{+}08]. Redistribution [HF01]. Reduced [CS05b]. Reducibilities [CN97]. Reducing [MM15]. Reduction [CRB15, Jun01, DMMM95]. Reductions [Hem99]. Redundant [FPT10, FA06]. Redwood [WHD04]. Reentrancy [Tan08]. Refactorings [GMB11, MRGF14, VM13]. Reference [CPRT05, dAFL07, HMB09, Lin03, LdL07, Win97a]. References [MS11]. Referential [Kap95]. Refinement [Aic01, MM07, RGP10, Sch08a]. Refinements [CMM01, FGBS14, Mar06b, Sch01a]. Refining [Mea97]. Reflections [FP12, GPFL12]. Reflective [CCP^{+}07, dIVG^{+}06]. Regard [Tez13]. Region [LKT10, Smy00]. Region-based [Smy00]. Regions [Tak03]. Registers [CGFSHG09, PSS07]. Registration [CWT^{+}15]. Regression [ACAMM15]. Regular [HI99, Ito02, KPS96, MS03, ST05, Sou99]. Regularity [Leu07]. Regulatory [BMM^{+}09]. Reification [CBG04]. Reinforcement [LBG07, MMiF^{+}08, MS05].

Reinvention [Car00]. Related

[BNCGD^{+}11, CSZ07, DGlS12, KK10, Kri99, dPPRRGSSPP15, Sal10, Vai00, BCM12]. Related-key [DGlS12]. Relating [dFBNGS^{+}14]. Relation [ESG10, Kat05, Świ07, Was98a, Zou06]. Relation-based [ESG10]. Relational [BB04, FPT10, HPC10, MYCA11, SZS12, VV15, Was98b]. Relations [AY00, Ara03, FPT10, GLS00, LN02, MYCA11, RAC10, ST05, VV15]. Relationship [uRLFH^{+}13, Sch08b, SZZ95]. Relationships [AH04, BCM08, GS12, LLCN09]. Relative [PRT^{+}08, PJO15]. Relativizations [Gaß09]. Relativized [Gaß10]. Relativizing [GW03]. Relaxing [Bur05]. Relayed [LWH09]. Release [CPSAGPGC12, Tak06]. Relevance [HM01, WC08, ZD09]. Reliability [KHN99, KMM14, UT16]. Reliable [Die10, LO98]. RelView [BH01]. Remarks [Fer96, IDS02]. Remote [HMW08, HSFE12, HA13, OAR^{+}14, PdP^{+}04, Tddd03, WLKW11]. Removal [NdCFB08]. Remove [LdSM08, MRGF14]. Removing [PBB07, dSLMW08]. Rendering [MMP15]. Renewable [KCK10]. Reo [Bai05]. Repeated [ASS13, BDM15]. Replacement [BCS15]. Replicated [FMT^{+}15]. Replication [AAM14, ES05]. Reporting [Die10, WLtHN14]. Repositories [BS96, MPL11]. Repository [GSdSB16, HKKvP08, MSC03]. Reposting [LLZ16]. Represent [Dit02]. Representation [CS05a, DGL03, GL00, HFOB08, Heg10, JW13, KM95, LSR10, LCC11, LW08, LM15, PRBLAP^{+}13, SUKG13]. Representational [Sut01]. Representations [AT10a, ACAMM15, BQBW03, FL14, Hel07]. Represented [Bol06, Bos08b, Wei08]. Representing [ASH11, GHNT97, LN02, Pau09]. repudiation [OO08, RMMLBLGS09]. Reputation [CMSE09, LDO^{+}12]. Request [Jun05b]. Requests [CAS^{+}13]. Requirement [AGMT10, CIM14]. Requirements [BdS13, BRS00b, BG00b, BKL12, CPCLSAGC11, CCMP08, DSM13, FG03, GWW05, HB00, Hei07, HVCA12, HSD^{+}14, IO04, PEPP08, SF00, TWH00, dOBGH^{+}14]. Rescue [CDBZ09]. Research [ACM16, BM13, BFMSP05, CR12, Cos08, FZ00, FP12, GMHGRG^{+}13, GPA08, GPCPL12, GPVILN13, HWN02, IBN^{+}11, KGK12, LRR04, Lei10, LJLCR13, LZZK14, LMA^{+}14, MNF^{+}13, RB08, STFM12, SZWdP14, SSS12, TRR06, TPC^{+}12, VRGPSP05, WP15, SL96]. Residue [Mue04]. Resilience [Gre08]. Resilient [LWY09, KSY97]. Resolution [EIH08, PBB08]. Resolving [WXZL15, ZSG14]. Resource [CXB12, CS10, GKZ05, LSV06, MNL13, MKS09, PKP08, VUT^{+}08, WZZ^{+}09, XCJ13]. Resource-aware [GKZ05]. Resource-Oriented [XCJ13]. Resources [BAP^{+}16, DKD^{+}13, FPLS03, GKK^{+}02, GMdMC12, HBT12, PCLT15, RGRR15]. Responds [FDC^{+}13]. Response [DH10a, HPE14, KM07, LVS13]. responsible [vB96]. Responsive [FWS^{+}11]. RESTifying [VGGSBLAP12]. Restoration [CM11]. Restricted [GN10, HK15, KPS96]. Restricting [KB06]. Restructuring [GSS99]. Results [BCG98, DGlS12, ECHS10, FGB^{+}14, FWS^{+}11, NH03, Pau09, Pre97, Pre12, RTB13, RR08, SDÖ^{+}12, Spi05]. Retinal [BCG^{+}09]. Retrials [GGDBP^{+}08]. Retrieval [ACB02, AdGCD^{+}15, BCG98, BDM15, BCCH11, BvZH09, dSBGdAdLM08, CP15, CR04, CYL11, DMS05, FL10, HM01, HKTV06, JM15, JR02, KDKB07, KSR16, KBF^{+}11, LKB^{+}02, MUSA03, Mue95, Rie02, SMK^{+}04, SH11, SKL08, WC08]. Reusable [CP01, MC07]. Reuse [BFF99, BMa11, BV07, BdS13, FKO14, GHM04, GB03, POB11a, RdL08, SBCJ03, WO09, dAO13, HLHD^{+}07]. Reuse-based [BMa11]. Reversible [KPdF06, SDJ99]. Review [AY12, CDP13, EACGFK13, GSZ15, HSR10, MHLB12, PvW16, RMFM12, RVW07, VGAPGS^{+}15]. Reviews [AA16a, BG98, dPSZPVLR^{+}16, SL10]. Revisited [CG96a, Loo06, WMSH09]. Revisiting [Ver10]. Revitalization [RP08]. Revocation [XMLC13]. Rewriting [CB04, CP06, LF16, MPK04]. Rewriting-Based [LF16]. RFD [RRR10]. RFID [BHS^{+}06, dIdSGZ10, VTGA13]. RIAs [LPSF10]. Rich [GSPK08]. Richly [HMSS01]. Richness [BM03b]. Riddance [Odl94]. Riders [Azz10]. Ridge [LD06]. Riemann [FF08]. Riesz [JW13, LW08]. Right [SG96]. Rights [TSDP07]. Rigorous [Gal98, Pav95]. Ring [Gaß10]. Rings [FP06]. Ripple [JMEL10]. Ripple-Down [JMEL10]. Risk [Ma12, MYC14, MMM^{+}12a, ND08, SFP12, SKSP09, TTB13, dTR03]. Risk-Aware [SKSP09]. Risk-Driven [MYC14, SFP12]. Risks [BS12a]. RITA [CEK15]. Road [Saf08, VJ09, XHP^{+}09]. Roaming [YWD08]. Robert [Kuh03]. Robot [AMR^{+}14]. Robots [SM02b]. Robust [CDR^{+}09, DRRGdP07, LGAP11, PTO^{+}12, PPG95, VJ09]. Robustness [CS05b, TNM09]. ROC [OJSB08]. ROC-space [OJSB08]. Rodeh [FP06]. Role [Bic15, DGN13, JM15, JV05, Kom02, KW10, MRP14, MP09, PB04, RKH15, Sto03, SWY09, RLMS13]. ROLE-enabled [RLMS13]. Roles [JGW11, RMF^{+}98]. Room [Müh96]. Root [XMZbL10]. Roots [AV07, Akr09, Gon06, Mue04, Şte05]. Rosetta [KAM03]. Rosser [Ban97b, Lin04c]. Rotating [SHK10]. Rotation [PQ99]. Rotation-Symmetric [PQ99]. Rough [Ngu09]. Round [ACL95, CJZ13, Jéz95, LKHL09]. Round-off [ACL95, Jéz95]. Rounding [BB98]. Route [HHH^{+}02]. Routes [Ara03, JNdMM12]. Routing [ARQH14, CS10, FLF^{+}14, HYC^{+}05, HHH^{+}02, HCH^{+}09, HJZ07, JP07, LEC11, LWY09, MWM10, MdCRMP14, NM07, PTO^{+}12, de 00, dG15]. RSA [Gon06, NZCG05]. RSA-based [NZCG05]. RSAb [MPPS95]. RT [OD03]. RTCP [EHEH05]. Rudeanu [CS00]. Rule [AK09, ABCP02, BGP07, CCHdCN08, Dru13, Esp06, FTARR05b, NNT16, Pob11b, ZTX^{+}07]. Rule-Based [BGP07, Esp06]. Ruler [NR08b]. Rules [Bos08a, CG04, CJ98, FTARR05a, JMEL10, KPS96, LHC^{+}13, MVPP02, MDY10, Neg05, NNT16, NXSA12, OCB^{+}10, PI04, VTHM16, ZTN^{+}15]. Run [CSC08, ZTN^{+}15]. Run-time [CSC08, ZTN^{+}15]. Runnable [ELFAR15]. Running [EGG^{+}01, MI05]. Runtime [GRHMM^{+}15, PBTW07, Sue10]. RUX [PLSF08]. RUX-Method [PLSF08].

S [DS08]. S-languages [DS08]. SaaS [JJ12]. Safe [GRHMM^{+}15, PTNMC08, MMdMGM06]. Safer [FAT^{+}13]. Safety [BCD13]. sales [Kuh03]. SAT [AMS04, AH04, CNQ04, FKS^{+}04, LWC^{+}04, MM99, Vel04, VJTJ07]. SAT-based [AH04, CNQ04]. Satisfaction [Ist07, PRW95, SS07]. Satisfiability [GK13, RR06b]. Satisfying [Ist07]. SBLP [BdVG06, BM07, VMdCJ08]. Scaffolding [MHLB12]. Scalability [Mac96, VC13]. Scalable [EHEH05, HLS15, Kap95, KS05]. Scalar [AHRH08]. Scale [CVSM11, GXC^{+}15, NHH06, OEK16, Ski00, UHÖD15, UT16, WF12, HBT12]. Scanned [PTL^{+}09]. Scattering [HW10]. SCBS [Vie03]. Scenario [BCCH11, ČSŽ09, DTG10, KTL^{+}11, SKHK14, SGS13, Tru10]. Scenario-based [SKHK14]. Scenarios [BS08, ES03, GH08, Hel07, HBFV13, JBBH13, NOP08, PRCRARLN10, RLL^{+}10, SVFR15, TBL15, Yan05]. Scene [HFOB08, SHK10]. Schauder [Bos09]. Scheduled [KAG00]. Scheduling [BARB12, CT04, HCK11, LJ00, LWG14, SAA08, SM02b, CGVRGPSP07]. Schema [BBP08, FC03, STW09, TFMDM10, TKT07, Was98a, Was98b]. Schemas [VO10]. Schematron [KBN14]. Scheme [EHEH05, FH06, FLF^{+}14, LYLX15, LCDP15, LWY09, NLLJ12, QF12, SLJC08, VAS05, WZC07, ZQQ15, ZSG14, LZLW13, dR05]. Schemes [ACA^{+}16, BM99, CT08b, DLR97, Man97, OK98, Shi11]. Schoenhage [DDG97]. Scholarly [LAM12, Odl94]. School [CM09, GMP^{+}13, Mac01, OT16, Pel14, Ret08a, Sin06, Tez13, TLS12, WPL98]. Schools [FZAP13, LM03b, RF12]. Schoolteachers [AHSN01]. Science [BCDK97, CJ07, Dom01, GTGT10, JK10a, KP01, RB08, RA06, TLS12, Vai00, VV06, vdV08, CMS94]. Sciences [BVG08]. Scientific [AM11, Car95, HTHW12, LZZK14, SSdS^{+}11, SZS12, WMJ^{+}07, von98]. Scoping [Kwo97]. Score [dlFVPHB15]. Scores [OF13]. SCR [HB00]. Scripted [PRCRARLN10]. Scripting [HMW08, MI04]. Scripts [MI05, SE09]. SD [NDAM09]. SD-Core [NDAM09]. SDL [EGG^{+}01, GK97]. SDL-2000 [EGG^{+}01]. SDLMAS [ČSŽ09]. SeAAS [HMB09]. SEAL [HMSS01]. SEAL-II [HMSS01]. Seamless [GGDBP^{+}08, Luk08, Ozd13]. Search [APM04, APNA12, AMS04, BBGV07, BCG98, BCC^{+}06, FGB^{+}14, FKS^{+}04, FMT^{+}15, GWG96, JJL08, JNdMM12, KJZJ08, KB06, LdPK^{+}14, LWH09, Mue04, QC12, QF12, RTB13, SAB99, VGBLGS^{+}08, VGGSBLAP12, Wol99a, Wol99b, Wol00]. Search-oriented [SAB99]. Searchable [XNKG15]. Searching [Kie05b, KJZJ08, Wit08, ZD09]. Seasonal [Faz06]. SeCA [BS12a]. Second [AKMS94, AUN04, CDF97, PPJ04, RA06, Sch02a, SE09, Ste00]. Second-Order [Ste00]. Secondary [BAP^{+}16, CM09, OT16, TLS12]. Secrecy [Blõ06, JR96]. Secret [BM99, OK98]. Secure [AP09, ACA^{+}16, BdGFMT14, CAD^{+}06, HK15, HLC08, KTJ05, MMM^{+}12a, MKI^{+}12, Nd05, NSNK05, PO11, Pos01, RFMLP10, RMMLBLGS09, SKSP09, STBFM09, VL14, WLLL09]. Secured [VAS05]. Securing [CS09, SBTH04]. Security [ATSJ05, AA16b, ACM16, BS12a, Bel08, CSFFM12, CPRT05, CDP13, CI12, CSW^{+}08, CWT^{+}15, DSM13, FMR09, HMB09, HLCL11, IAS16, JL16, KDGH09, Laf04, LKHL09, LZGC09, LZLW13, MMM^{+}12a, MR12, MKI^{+}12, NML09, Par09, PHJ^{+}08, QC12, RvS12, RMFM12, RSFMJ12, SMGMT09, SPRP09, SFP12, Shi11, UFF12, VdR09, XHP^{+}09, YWD08, Yon11, YW13, YMA15]. Security-enhanced [QC12]. See [AMYH14]. Seeking [GlRBdSG11, PCLT15]. Segmentation [BM11, FL10, KC08, KD11, LALS08, WH08]. Segmentation-based [FL10]. Segments [FTARR05a]. SELaKT [MPF04]. Selectable [BCG98]. Selected [BdVG06, BM07, CMZZ07]. Selecting [GP10, HNJ^{+}10]. Selection [ASTL07, AZMA15, BJMBA15, BMGMF08, CHPHV10, CCMP08, FTARR05b, GBP^{+}08, GCC16, KCK10, LKP11, LWL10, MK12, dPSZPVLR^{+}16, SW13, TJS^{+}13, YKA16]. Selective [CS07, Has02, HNP98, INK09]. Selective-ID [CS07]. Selectivity [HJRW97]. Self [AMR^{+}14, BD05, CMZZ07, CFF^{+}13, CC07, FAT^{+}13, HCBB15, MCG14, OCRPdlMG07, Ozd13, PB14, Pau13, RF12, SBS15, ZDE14]. Self-Adaptation [MCG14]. Self-Adjoint [BD05]. Self-Assembling [CMZZ07]. Self-Assessment [HCBB15]. Self-Aware [FAT^{+}13]. Self-directed [Ozd13]. Self-Efficacy [RF12, SBS15].

Self-Evolving [CC07]. Self-learning [AMR^{+}14]. Self-management [ZDE14]. Self-Reconfigurable [CFF^{+}13]. Self-Service [Pau13]. Self-stabilizing [PB14]. Self-Timed [OCRPdlMG07]. Semantic [ATGP09, ASHT^{+}16, BBGV07, BNCGD^{+}11, CJO^{+}13, CTM10, CDD^{+}03, CDD^{+}04, CDD^{+}07, DSM13, FMB^{+}11, FBCMR15, GMHGRG^{+}13, GLSD11, GSdSB16, HGMT08, JM15, Jun05a, Jun05b, JN08, Jun08, KKB12, KJKS14, LST14, LKB^{+}02, MMM^{+}12a, Ngu05, OHYJ16, OBO09, RGRR15, SR10, TCS^{+}03, VGBLGS^{+}08, VGGSBLAP12, VBB13, WL13, ZTN^{+}15, ZD09, dTU04, DKD^{+}13, CLC09, CPMVG13, GL11c, Hef04, HNJ^{+}10, Nal10, OCB^{+}10, RRM^{+}12, SGB^{+}13, SMFM05, SA11, VAH07]. Semantic-based [CDD^{+}04, CDD^{+}07]. Semantically [AdI16, LKMS08, RJB10]. SemanticMiner [MUSA03]. Semantics [Are02, ADMdB09, Ban97a, Ban97b, BAPG03, BRF^{+}09, BCM08, Bür08, CM03, CB04, DG07, GLD^{+}12, GK97, KF10a, KF10b, KAM03, LI05, MCM07, TBS08, Tra00, TGEM07, XCJ13, YMP08, dlELR09]. Semantized [KJZJ08]. Semi [AP05, GHHE^{+}08, LSR10, LHK^{+}13, MUF03, THJ16]. Semi-Automatic [AP05, GHHE^{+}08]. Semi-Structured [LSR10, MUF03]. Semi-Supervised [LHK^{+}13, THJ16]. Semiautomatic [Hri02]. Seminar [BJ05a]. Seminars [SH96]. Semiringal [GLS00]. Semirings [ÉK02, Vai00]. Semistructured [LDSG09]. Sense [MNL13]. Sensemaking [ZABB07]. Sensitive [CLVM09, CS09, LKP11, Meh02, Kuh03]. Sensitivity [PA12]. Sensor [LGAP11, LLYC12, Log06, MWM10, MdCRMP14, MAT08, PZDH09, PdCdlTR06, QZ07, SHZ^{+}10]. Sensors [SHK10]. Sentiment [AA16a, ASAAASJ16, CBRH12, RPCA15, dPSZPVLR^{+}16]. Separation [BZA08, CRMLN^{+}07, Gaß10, PI04, Wei10]. Sequence [Are02, BD06, GR01, HKTV06, Hon01, Lav96, LWH09, SSGS10]. Sequences [BH01, Cer97, HW97, Her05, dCPUH^{+}07, RC07, Sta02]. Sequencing [GSPK08, GSMBFPK10]. Sequential [Ish97, NHH06, Şte95, YTM05]. Sergiu [CS00]. Serial [Gün96, NK95]. Series [BB98, BB09, DS03, Dru06, Faz06, GWW05, HK95, Hon96, Hon02, NWDX09, Hon97, Hon99]. Serious [Pel14, SKHK14]. Server [GHS06, SAKAM11, GJP^{+}12a]. Serverless [AHT09]. Servers [BS06]. Service [AA16a, ASTL07, AdI16, AHSN01, AFL08, AdO11, AF04, BCFM05, BGMR^{+}16, BP09, BKH^{+}13, BCG^{+}14, Buf01, CAD^{+}06, CBR^{+}05, Che09, DT12, DKL10, EMZB14, FGSW14, GPFL12, JJ12, Jun10b, JK10a, KK13b, Kim12, KW10, LM03a, LWS11, LWL10, MYC14, MDO^{+}09, MKA11, NL10, NOP08, NM07, NPC^{+}09, PCS^{+}13, Pau13, QZYL11, Rad14, RvS12, Riz15, RVC12, SLD^{+}16, SEK13, SIIJ09, SW09, SvRvdV^{+}13, SM02b, VAPM12, WK05, XCJ13, YLZW10, ZGC^{+}08, ZSG14, ZDE14, dCVM12, dG15, LLM02]. Service-Dominant [NL10]. Service-Oriented [BCG^{+}14, CBR^{+}05, FGSW14, NOP08, SEK13, SIIJ09, YLZW10, dCVM12, ZGC^{+}08]. Serviceability [Kim12]. Services [BCA^{+}10, BHH^{+}06, BHS^{+}06, Buf01, CVM11, CMMP16, CI12, CS09, COBP^{+}14, EDA14, EMGB^{+}12, FHJ^{+}99, GG08, GPFL12, GFBR08, GSdSB16, GO14, GJP^{+}12a, HMB09, HMMP10, JBBH13, Jun10c, JK12b, KKTZ09, KTKP09, MSTW12, MSM07, MCG14, MP06, PTL11, PD10, PS09, PGDD15, PPP^{+}11, QL12, Ram01, RLMS13, RRM^{+}12, SR10, SBGI14, Sat10, SC14, TR10, TSDP07, VBVNHdDSL12, Vle14, ZFS98, dLH08]. Sessions [CGD^{+}12, DDJ^{+}11, MP09]. Set [CGP07b, CK95, FA06, GCC16, KK10, NNT16, PMP02, Rat05, SST07, Sta02, VTHM16, WH08, WXZL15]. Sets [AJBTEB06, Ars97, Bul95, CDF97, DHO98, Gro00, HNP98, Her97, Hon95, Ilj09, KS05, KK10, dCPUH^{+}07, Pet12, Pop07, RdKO11, RA06, Vît05, Wei08]. Setting [DLY08]. Settings [Ret08a, RCGBS13]. Seventy [JMSY10]. Several [ACL95, Gaß09, KJKS14, Sko08]. Sevices [BCNR07]. SEWASIE [BBGV07]. ShanghAI [LHZS12]. Shannon [Fen97]. Shape [KM06]. Shapes [MGAVF10]. Shaping [FGS98]. Sharable [RGHH97]. Share [CT08a, MPG13]. Shared [Gio98, GGB^{+}08, MPG13]. Shared-Workspace [GGB^{+}08]. Sharing [BM99, CS03, DDS04, DLR97, HFIBJ13, Kom02, LYLX15, MHA^{+}15, MKS09, MPL11, MS11, NW04, OK98, QF12, RAWW05, SBMD10, SE09, SM04, WP03, XNKG15, ZWH10]. Shark [SG02]. SHARP [GRGPL08]. Sheet [MM98]. Shift [AT10b, CGFSHG09, Lip00]. Shifting [MJS13]. Shifts [Spa08]. Shooters [GOM^{+}13]. Shop [NZM09]. shoppers [HGIPCPPM11]. Shopping [FKK^{+}10]. Short [NWDX09]. Short-term [NWDX09]. Shortcut [MP08]. Shortcuts [BH14]. Shortest [Str97]. Should [dSJPM14]. Showing [BdGFMT14]. Shrink [Ev99]. Shuffle [Ito02, NSNK05, NXSA12, PRS95, VTRN12]. sic [Mes02]. Siemens [Ram01]. SIFT [SH11]. Sign [Mar00a, VJ09]. Signal [LWC^{+}04, WZC07]. Signaling [HLNA^{+}12]. Signals [AH04, UV05]. Signature [NSMBACBG12, NZCG05, OJSB08]. Signatures [DR10, PO11, WMSH09]. Signcryption [FZT13]. Signed [Tat07]. Signing [RMMLBLGS09]. Silence [Tin16]. Silhouette [CVK97]. Silhouette-like [CVK97]. Sim [QFB^{+}14, Pel14]. SimCon [MOG^{+}10]. Similar [RKJ16, SA10]. Similarity [LNH^{+}15, NSF^{+}10, VV15, WSL07, XWGS09]. Similarity-based [LNH^{+}15]. SIMOLA [CPLPW15]. Simple [BR05, EMGB^{+}12, MSSY95, Wat02, dL03]. Simplification [Hon02]. Simplify [DH10a]. Simply [Her02]. Simpson [CJ98]. Simulated [FSSPLG^{+}13, SW13, SA03]. Simulating [HPB12, LZM04, PY00]. Simulation [And97, DH10b, DZBB^{+}12, DCR^{+}07, For97, GRS08, HW10, MOG^{+}10, NHH06, PKP08, PRW95, Sch01a, SV08, TKD^{+}09]. Simulations [BB98, CDBZ09, GDW10, Lep98]. Simulator [CGVRGPSP07, MMiF^{+}08, NC04, PdlCBKN14]. Simultaneous [SU01]. Simultaneously [CDR^{+}09]. Single [FA06, Mar02b, TG10]. Singularity [Rex98]. Site [CLVM09]. Sites [BU13, BRAS^{+}12, LN08]. Situated [ARN04, CPLPW15]. Situation [JK10b]. Situation-Aware [JK10b]. Situational [DLL16, HSR10]. Situations [LM01]. Size [CS07, Dvo97, JS16, SCW08]. Sized [HNP98, SPRP09]. Skeleton [AAAG95]. Skew [BMMM14, KD11, PF15]. Skill [Bec03, CDD^{+}07, HH03, HCWA03, SA03]. Skills [CDD^{+}03, GDW10, LA03b, PD04, RBLR02, SdBC13, SKHK14]. Skipjack [CJZ13]. Skipjack-like [CJZ13]. SKYWare [ELFAR15]. Slack [HCK11]. slender [Hon97]. Slicing [CW00, RB06, RB07]. Sliding [KTJ05, KL09b]. Small [GXC^{+}15, Len00, MX05, SPRP09, TTB13]. Smart [BCA^{+}10, BBIC13, CMMP16, GHHE^{+}08, GHHA10, Jun10c, MOG^{+}10, NOGG^{+}13, OCW13, PGDD15, SW09, TLS12, XHP^{+}09]. Smartphone [AGO^{+}13]. Smartphone-Based [AGO^{+}13]. SmartSocial [SP16]. SMDM [RGP10]. Smells [PHPP06]. SMS [HKK13]. SNAP [ATOFF98]. Snippets [WHD04]. Snort [SAKAM11]. SOA [CTM10, DKL10, HMB09]. Social [AGG^{+}08, BZ09, BU13, Bic15, BvdTV09, CHH16, DIKL14, FP12, GÁVCNC14, HKKvP08, JKKW16, Jun05c, Jun08, Jun10b, JK12a, JK12b, KKTT16, Kim12, LTY^{+}16, LDO^{+}12, LCC^{+}12, LZZK14, LMA^{+}14, LAHZ^{+}15, MKS09, MOS^{+}13, MPF04, OAR^{+}14, Ozd13, PCKJ11, PJH12, PD04, Pre04, RHM15, SBGI14, SESMT10, SBMD10, SA14, SZWdP14, SC14, SP16, SvRvdV^{+}13, SSdS^{+}11, SZS12, TCW12, TB16, Tin16, TJ15, Vie03, WWD15, ZBKK12]. Social-Aware [SBGI14, AGG^{+}08]. Socially [KPV^{+}11, ZD09]. Society [ARS00, Gün02, LdP11, MO03, Sch01b, Sin06]. Socio [DKD^{+}13, FHH08, LKT10]. Socio-Economic [FHH08]. Socio-Institutional [LKT10]. Socio-semantic [DKD^{+}13]. Sofic [Spa08]. Soft [BM03a, GDW10, HMSS01, MSF99, VTHM16]. Software [ABB14, AFK01, ALHM^{+}14, ACM16, AFP04, AK05, BFF99, Bar03, BMa11, BB08b, Bjø01, BM97, BPC04, CMP08, Car98b, CB12, CL07, CPCLSAGC11, CPFSdAS12, CPSAGPGC12, CRC04, DAd03, Dud08, ESM08, ELFAR15, FKO14, FWT11, Fod06, Fro02, FG10, GG08, GNP05, GlRBdSG11, GB03, HKL^{+}06, HW10, HCPdASY14, HCD10, HP15, JBBH13, JJ12, JGM10, JGM^{+}13, Kar13, KD02, Kel08, Kie05a, KCKL10, KKK^{+}14, LM10, LCHD12, LWS11, LdP11, MPM12, MS10, MM12, Mül02, Nd05, OPP09, PA12, PTNMC08, PRCCS13, POB11a, Raj07, RP08, RdL08, SNAF07, SFP12, Sch02b, SV95, TSCY01, VK03, WO09, Wie08, WPL98, dAO13, dMTS^{+}14]. Software/Hardware [Nd05]. SoLo [Oli01]. Solution [BCD13, Jéz95, MM99, PPJ08, RRM^{+}12, Sch96, SS09b]. Solutions [Aur01, CG09, FML13, Fel01, HLHD^{+}07, JV05, LM03b, Mac96, NO98, Neh98, Pop98, Rih98, VUT^{+}08]. Solve [EC00, GBP^{+}08]. Solved [FOSS99]. Solver [LWC^{+}04]. Solving [DB12, GRGPL08, Gra98, HME^{+}06, JRR16, LLH03, LA07, MKP98, Rat00, SA03, ZZH^{+}12, DBAB12]. SOM [LVV10]. Some [Aur01, Bra02, GPFL12, IDS02, Kud99, KMM14, LM03b, MC00, MKP98, OK98, Pav95, Vai00, XLMR10, ZDI10, AHT09]. Songs [IGS08]. Sonography [CLCC10]. SOPHIE [Arr07]. SOS [HLNA^{+}12]. Source [AR04, FKO14, FBCMR15, IAS16, JP07, MKS09, MMD12, PB05, TTB09]. Sources [MUF03, Rie02, SA10, VBB13, VO02, dTU04]. South [BKL12]. SP [CJZ13]. SP/SPS [CJZ13]. Space [BJ97, Hav05, Köh09, LRI03, Leš09, MTB^{+}08, MM99, Mar04, PB05, YTM05, CP02, OJSB08, DGBM08]. SPACE-DESIGN [DGBM08]. Spaces [AGG^{+}08, AY12, BEH^{+}05, Bos08b, GHHA10, HLNA^{+}12, Ilj09, Jun10c, KHLAP12, KPV^{+}11, KNSN07, Lip00, LW08, Pal05, RAWW05, ST05, Sch08b, Sch09c, Sto02, Vît05, Sch10]. Spam [HKK13, LKP11, VV12]. Spanish [CM09, GGP08b, PGDD15, dPSZPVLR^{+}16]. Spanish-Speaking [GGP08b]. Spanning [DT07, HNS07]. Sparse [LST07]. Spartan [CFF^{+}13]. Spartan-3 [CFF^{+}13]. Spatial [Ara03, CDR^{+}09, Fal10, GALR02, GALR03, GB10, LN02, LLCN09, ML02, Poo03, Saf08, SH96, SA11, URG^{+}13, VCB08, VDBNR98]. Spatial-reasoning [ML02]. Spatial-Temporal [LLCN09]. Spatio [DSCT10, MMEdP12]. Spatio-Temporal [MMEdP12, DSCT10]. Spatiotemporal [DMS05, KL09b]. Speaking [DA13, GGP08b]. Special [AFK01, BHRS03, Boe97b, BG01, CS00, CSY02b, DT09, Dor95, DG00, Dvo00, DSRR03, EK99, GALR02, HMSR99, IFd03, JFZ09, KP01, KU00, KZ03, LA03b, Lin04a, Lin04b, Mat99, Mau03a, RS01a, RS01b, RS03, RA06, SdOB09, TM01a, TM01b, Toc02, TM02a, TM02b, Toc03, Mul98b, Mul98a]. Specializers [NR08a]. Specific [BDL^{+}06, ESM08, FRD14, FGSW14, LRS^{+}11, MG14]. Specification [BRS00a, DSM13, DF00, Fro02, Gär99, JBH^{+}10, KASS03, KAM03, KP97a, MC00, Mea97, MR08, MM07, OD03, RTJ01, Sou99, SRR04, WTA01, Zim01, dR05]. Specifications [ALHM^{+}14, Ber06, BRW03, CM03, DS03, Dvo00, DSRR03, EGG^{+}01, Hor04, KP97b, NuR05, PU97, SS03]. Specify [RPR11]. Specifying [Hei07, RRB03, RCGBS13, Ste00]. Spectral [AC07]. Spectrum [BD05, FK16]. Speculation [SU01]. Speculative [CW00]. Speech [KM07]. Speed [MPPS95, VCB08]. speranto [JST11]. Spiking [Pau07]. Spiral [ATGP09, HC08, Tin16]. SPL [FGBS14]. Splice [CLVM09]. Splicing [KPS96, Rah99]. Spline [Ang98]. Spline-Fourier [Ang98]. Splittings [Ars97]. Spooler [MPF^{+}16]. Spot [HMSS01]. Spots [Toc03]. Spotting [RdKO11]. Spreadsheets [For07]. Spring [VM13]. Spring/AOP [VM13]. SPS [CJZ13]. SQL [PRAT09, SCT09]. SQL/XML [PRAT09]. Square [MYT09]. Squaring [MJS13]. st [AR95]. st-numbering [AR95]. Stability [Kar02, Kou09, NKS^{+}09]. stabilizing [PB14]. Stable [LM07]. Stack [And97, Lan10, UDC97]. Stacked [ACAMM15, CL08]. Stage [Log04]. Stakeholder [TTB13]. Standard [BH14, CDP13, CI12, FZ00, HLC08, KJL09]. Standards [BS08, BQV14, CSA10, Duv01, Fra98]. Standards-based [BS08]. Standpoint [DKL10]. Starting [KWH03]. Starting-Point [KWH03]. Starts [HA03]. State [Ara97, ANdMM08, BB08a, Bar03, BG97, Boe97c, BRS00b, BG01, BP08, DDS10, DDG97, GRS08, GK97, GCG08, GS97, Hav05, HSR10, Kir09, MCC13, OL08, PdlCBKN14, SA97, Sch09a, STW09, Sch01c, SV08, SN01, Toc02, WTA01, Win97b, Yeu04, ZSK09, Toc04]. State-based [Bar03]. State-of-the-Art [HSR10, Toc04]. Statecharts [MM03, Tra00]. States [BDL^{+}06, STW09]. Static [AMS04, FA06, Puc10, SLN16, SSGS10]. Statistical [AdGCD^{+}15, Che09, SW13, Spa08]. Statistics [Les95]. Steganographic [WP15]. Steganography [CS05b, FMS12]. Steiner [DT07, GGS08]. STEM [IBN^{+}11]. Step [BTD^{+}07, GBP^{+}08]. Stepping [Pre04, SZZM10]. Stepping-Stones [Pre04]. Stereoscopic [HFOB08, SJ13]. Stochastic [EGDG09, SKA08]. Stone [CS05a]. Stones [Pre04]. Stopwatch [TLR09]. Storage [DDG97, Köh09, LCDP15, XNKG15]. Stores [GJP^{+}12b]. Story [dARGSB11]. STP [HNS07]. Strategies [ABB14, APNA12, BCS15, CW12, Has02, IS10, JRW10, KBF^{+}11, LGL09, PZ03, RR11, SKA08, SMMC10]. Strategy [BCHM12, CM98, HA03, IKM03, JRR16, JMKT12, MK12, OBO09]. Strategy-Based [OBO09]. Stream [BCG^{+}99, DGlS12, FKB^{+}15, TFMDM10, Tra13]. Streaming [GWW05, PZJ09]. Streams [ARG05, FTARR05b, GKZ05, GM05, JFZ09, Jun05b, KL09b, LLS05, THJ16, Yan05]. Strengths [KM13]. stricken [LM01]. String [Cal95]. Strong [BM03a, CN97, RTL06, ZZ96]. Structural [AH04, BS03, DGL03, GNP05, HWN02, KBN14, KZ08, LKZK10, PK98, SCT09, TF09, XCJ13]. Structure [BCM08, Cet00, CJZ13, HMM00, IPCVC12, Kat05, LYLX15, MUF03, SST07, TKF06, VdSdMC08, Zou06]. Structure-Based [VdSdMC08]. Structured [GC14, HMA^{+}05, HMSS01, Kwo97, LSR10, MUF03, PSF98, PSS07, SH09, Ski97, THS11, WD02]. Structures [BGBA10, Boz99, BM96, Fro02, Gaß09, Her09, HCWA03, Ilj10, KFK05, Kud99, LN02, Vai00]. Structuring [CL08, MC00, MHA03]. Student [CGLdMAC14, ELS04, EACGFK13, HR06, Kar13, Kro13, OT16, PJ15, Pel14, RKH15, STFM12, SCS13, SBS15, SS07, Tez13, UHÖD15]. Students [FP12, KM13, RAS15, WPL98]. Studies [DRA^{+}04, Gal98, SKH14, VRGPSP05, VK03]. Study [AC05, AS14, AR01, BS12a, BB08b, BB10b, BC11, BU13, Bic15, BCM08, BRS00b, BG00a, Bor07, BE98, CGLdMAC14, CM09, CPHC11, CB12, CPSAGPGC12, CE11, CMSE09, CBG04, DLL16, DAk13, DJJN09, DGBM08, DBAB12, GlRBdSG11, GJP^{+}12a, HA10, HB00, JV11, KP00, LA03a, LASL12, LULGFC13, MRK^{+}98, MM03, MdCRMP14, Mat04, MT99, MV15, MS11, OEK16, PFS07, PMRO08, Par09, Pau13, QZB^{+}00, RH10, RF12, RAS15, SLD^{+}16, SFP12, SA97, SC14, SF00, Ste00, Str97, Sut01, TWH00, WL13, BG00b]. Studying [WLKW11]. Sturm [Neh98]. Style [MC00, MM98]. Styles [BM13, LT13, SA09]. sub [ZC09]. Subalgebras [BB10a]. Subcellular [AGK^{+}10]. Subdivision [CM98]. Subgraph [AC07, SAA08]. Subgroup [AP05]. Submission [FZ00]. Subrecursive [Sko08]. Subroutines [MKP98]. Subsets [Sch10]. Substitution [NdMM06b, ZZ96]. Substrings [Han10]. Substructural [Ish00]. Success [HSFE12, HLK09b, Pet09, RGPK15, SK08, Til01]. Successful [GJAB95, Hef04]. Succinct [AT10a]. Suitability [BCCH11]. Suitable [LKHL09]. Suite [MNF^{+}13, WL13]. Summability [BB10b]. Summaries [RPCA15]. Summarization [AZMA15, SL10]. Sums [XMZbL10]. Sun [dSC06b]. Superdistribution [Sch09b]. Superorganisms [Rad01]. Supervised [LHK^{+}13, THJ16]. Supervisory [TNM09]. Supply [PPJ08, SKSP09]. Support [ATGP09, AHEAS^{+}15, ARS^{+}08, BFF11, CVFN07, CVSM11, CM09, CCH06, CLVM09, Coo06, FKS^{+}04, GFT09, GFBR08, GL11b, GPSV03, HKKvP08, HVCA12, KDKB07, KY10, KOW01, KSdV09, LM10, LEJ^{+}08, LST14, LMRG14, MMM^{+}12a, ML95, MOG^{+}10, MR08, MRO05, MKI^{+}12, NOP08, PJ15, PZDH09, PMAM14, RSP^{+}14, RLL^{+}10, SMP^{+}11, SMMM13, SvRvdV^{+}13, TFMDM10, TR10, TB16, VAPM12, WC09, WHD04, dMTS^{+}14, vKL04, SLD^{+}16]. Supported [MN96, OT16, UT16, MSC03]. Supporter [BZ09]. Supporting [AVA08, BZ09, CRMLN^{+}07, CDG14, GP10, GGB^{+}08, HPE14, HA13, LASL12, LS10, MYCA11, MDO^{+}09, MOMGSRFM07, MHLB12, NW04, POR10, PBB08, PLG^{+}08, PSF98, SA10, SD97, TF09, VGBLGS^{+}08]. Suprema [Bar05]. Surface [LLLL99]. Surfaces [AT10a, ZG05]. Surjective [Her97]. Surrogate [EMZB14]. Surveillance [CO08]. Survey [ARQH14, AFP04, DH10b, HM99, HPC10, Kel08, MS00, MKZ06, Pal15, RR06a, UFF12]. Suspending [Ban97a]. Sustain [RKH15]. Sustainability [Joh01, Jun10c]. Sustainable [KNSN07, dIB13, MH02, MPF04]. SVC [PGSAP14]. SVC-G9 [PGSAP14]. SVM [NWDX09, WSL07]. Swarm [DB12, DXZL07, GGP08a, GBP^{+}08, LA07, ZZH^{+}12, ZC09]. SWEBOK [Nav09]. SWEBOK-based [Nav09]. Swedish [Pet09]. Sweep [DT07]. Sweep-line [DT07]. SWePT [BGMR^{+}16]. SWET [SGB^{+}13]. Switch [GN00, KASN08]. Switching [SI00]. Switzerland [BBM12]. Sylvester [vP05]. Symbol [BT08, Fen97]. Symbolic [Gra98, Laf04, Mau03a, PR06]. Symbolic-Numerical [Gra98]. Symmetric [PQ99, SS09b, vZG11]. Symmetries [DGL03]. Symmetry [Jun01]. Symport [MVPP02, CG04]. Symport/Antiport [CG04]. Symposia [RA06]. Symposium [BdVG06, BM07, IFd03, Lin04b, ML05, VMdCJ08]. Synchronisation [BCM12]. Synchronisation/Desynchronisation [BCM12]. Synchronization [CS04, DTG10, Kar02, MYCA11, PFS07, SY99, Sal02, SG02]. Synchronized [BH02].

Synchronizing [MHA^{+}15]. Synchronous [ANdMM08, GFBR08, KWC01, MPG13]. Synchronously [MX05]. Synergy [Rob06]. Synopsis [BG00a]. Synsets [Hri02]. Syntactic [MMM^{+}12a, OBO09]. Syntax [AMUFVI09, Wol00]. Synthesis [Cap05, GHM04, LCC11, TKF06]. SySML [WKSD^{+}11]. System [AM11, AHSN01, AY00, AAAK15, AKM95, Ano99, ARS00, AS07, AT07, BGP07, BS11, BR97, BH01, BHRS03, BPHN06, BDM15, Bör02, BBP08, BCS15, BJ05a, CAGMPGdAS13, CJO^{+}13, CXB12, CFF^{+}13, CVSM11, CM09, CCH06, CSW^{+}08, CW09, CLCC10, CWTT11, CW12, CCMP08, CD10, DDS04, DRS06, DHC11, DMG07, EKP03, EIH08, ESG10, FZAP13, FDC^{+}13, FPSFCG07, GHHA10, GRGN13, GRGPL08, GMC08b, GHM04, HPE14, HKL^{+}06, HLP^{+}13, HMHGR15, HIGM13, JST11, JL09, KJZJ08, KY10, KM07, KTL^{+}11, KNSN07, LN02, LZLW13, LMMPFVI14, MRK^{+}98, MRP14, MUF03, MYY06, MCC13, MMM12b, MRO05, MR05, NBGS06, NKS^{+}09, OL08, PRAT09, QZYL11, RKJ16, RS01a, RS01b, RSP^{+}14, RGRR15, RFMLP10, RY09, SNAF07, SG02, SH09, SMP^{+}11, Sue10, TYSY09, TRR06, TBVRGLD15, Tra13, TSDP07, Tüf13, VGGSBLAP12, VTRN12, Vie03]. System [VBB13, VdR09, WP03, XHP^{+}09, YLL^{+}07, YKD^{+}08, ZHG^{+}06, ZMAS10, dH00, dTU04, dKR03, dlELR09]. Systematic [Aic01, BRH^{+}08, BdS13, CLC04, JGM10, JGM^{+}13, Nør10, PvW16, RMFM12, Sut01]. Systems [Alh04, And96, AKMS94, AGGH08, ADMdB09, AUN04, Azz10, BCG^{+}99, BRF^{+}09, BBL13, BRR99, BGP08, BG04, BH00, CP15, CBN^{+}06, CAS^{+}13, CMM01, CN08a, CN08b, CCP^{+}07, CG04, ČSŽ09, CR00a, CS04, CS14, CM07, CGP07b, CFSC04, CCS10, DT12, DM04, DGBM08, Dus05, Dvo00, DSRR03, ES05, EGK^{+}12, EGDG09, ELS04, FMR09, Fer96, FGSW14, FF08, FR04, Flo04, Fod06, FF04, dAFL07, FP95, GPCAC11, GVRT^{+}10, GLCV08, Gär99, GNP05, GV00, GMC^{+}08a, GSP04, Gra98, HMSR99, Hel07, HOPN11, HGMT08, Hon95, Hop98, HT06, IS04, JR02, JNS09, JK10b, KWH03, Kap95, KNLS00, KKK^{+}14, KP95, Koó06, KZ08, KHG10, LT13, LS07, LM94a, LGMM^{+}13, LZM04, LJLCR13, LRS^{+}11, Loo06, Ma12, MVM00, MVPP02, MBC12, MHA^{+}15, MZ12, MR12, MPF^{+}16, MKI^{+}12]. Systems [MPK04, NS06, NO98, NdMM08, NdMM12, NS05, NH09, NXSA12, Ois98, OD03, ONRV08, ODSO11, OO08, Ozd13, PJO15, PI04, PKP08, PB05, PY00, Pau07, PS09, PJRC04, PMLL09, PP99, PJH12, PJN13, PR06, Pop07, QF12, RSW04, RSFMJ12, SI00, SR00, Sch02a, SSM11, SS09b, SZZM10, Sob05, Ste00, Sto03, SBCD15, SRR04, Świ07, Tab07, TFMDM10, TE06, TEK08, UFF12, VMA14, VMFO14, VDLG10, VV12, WKXL05, WKSD^{+}11, WC08, YLL^{+}07, ZDE14, dCVM12, dKR03, vKL04, CVPS95].

t [KSY97, Gün96]. T-Codes [Gün96]. t-resilient [KSY97]. Table [Mue04, NdCFB08, Suz06]. Table-form [NdCFB08]. Tableau [Pau99, SAB99]. Tableaux [Sto99]. tabled [CVPS95]. Tables [BI08]. Tablet [LRB16, OT16, UT16]. Tabu [JJL08]. Tabular [For07]. Tactics [PdP^{+}04]. Tag [LdPK^{+}14, dIdSGZ10, MKS09, QF12, THS11]. Tagger [TEC^{+}07]. Tagging [CBNDR10, GPCZ^{+}13, PJH12]. Tags [KJZJ08]. Tailoring [Kah01]. Tallinn [LKT10]. TAM [LLSA13]. Tamil [MR11]. Tangible [MBC13]. Target [GP10]. Targeting [GOM^{+}13]. Task [AGGH08, CDD^{+}04, CDD^{+}07, SSM11, VBP^{+}11, VCB08, VDSF98, WSF08]. Tasks [AGG^{+}08, DGBM08, HA13, MRO05, SA10, SM02b, TR10]. Taxi [HZZ^{+}12]. Taxicab [CCD03]. Taxonomies [ARS^{+}08]. Taxonomy [BC16, GMC^{+}08a, PCC14, QQ11]. tccp [AdMGPV06]. Teach [CL95]. Teacher [Kro13, RF12]. Teachers [Die96, FML13, IBN^{+}11, LAAPVGMM15, Sin06, WKTL01]. Teaching [Bjø01, De 96, GMP^{+}13, HKS96, KMR96, Kar13, LMO01, Len00, ML95, Mau96a, PSF98, QFB^{+}14, Rad96, Tom95, TBL15, VDBNR98, VKW15, FSELC13]. Teal [HCH^{+}09]. Teal-Time [HCH^{+}09]. Team [FP95, Has02, JRR16, SvRvdV^{+}13, SvRS14, SM02b]. Team-Oriented [Has02]. Teams [CG96b, GS12, LCHD12, LST14]. Technical [MRK^{+}98]. Technique [CA14, SBCJ03, SKP08]. Techniques [AK05, BFF99, Bjø01, CPRT05, CKdL08, DH10b, HPC10, HGMT08, Kel08, LSR10, MC00, RP08, TNRGCP^{+}13]. Technological [FBSEGP15, FP05, QFB^{+}14, TE06]. Technologies [BSP^{+}13, Car98a, CPMVG13, GR14, GHS06, Hef04, HLS15, LdP11, LAHZ^{+}15, NS05, Par09, PCS^{+}13, RC10, Rob06, SMFM05, Uzu13, VGCPAH16, YSP09]. Technology [AMA^{+}14, AB09, BP97, Bra15, BFMSP05, BHS^{+}06, CR12, CPHC11, DA13, Die96, Dro04, FMLNF07, GPCPL12, GPVILN13, GK06, HKS96, Hol96, JGW11, KHLAP12, LUR16, LZZK14, MHA^{+}15, PEPP08, TM02b, TBL15, UP04, VLE12, WKS^{+}99, dSC05]. Technology-Enhanced [VLE12]. Technology-Oriented [TM02b]. Tele [JFL^{+}13, MYY06, VRGPSP05, ZHG^{+}06]. Tele-assistance [ZHG^{+}06]. Tele-Education [VRGPSP05]. Tele-Home [MYY06]. Tele-mobile [JFL^{+}13]. Telecommunication [BCA^{+}10, Buf01]. Telecommunications [Gün02]. Telediagnosis [SSAB^{+}13]. Telematic [ZFS98]. Telemedical [HLP^{+}13]. Telemedicine [SZWdP14]. Telemetry [Log04, Log06]. Telephony [uRLFH^{+}13]. Television [IS10]. Tell [dARGSB11]. Template [GC14]. Template-generated [GC14]. Templates [PF11]. Temporal [BR03, DOM10, DS03, Dru06, Dru12, Dru13, DHP03, DS08, EGDG09, GMS03, GALR02, GALR03, GB10, HMA^{+}05, KR03b, LLCN09, MBA12, MMEdP12, PR06, RKJ16, STVT07, dH04, DSCT10]. Temporal-Fuzzy [MBA12]. Ten [Boe97c, CG09, Lar01]. Tendencies [FMR09, GNP05]. Term [Ber06, BM05, GK13, KD01, NWDX09, RLL^{+}10]. Terminal [DM10]. Termination [TT00]. Terminology [SBPR15]. Terms [Tüf13]. Terrain [CT16]. territories [MPRS95]. Terrorism [CDCH09]. Test [Aic01, Che09, GR01, Gir05, IRMK12, KM13, KW10, LCHD12, LdPK^{+}14, NuR05, Tab07]. Test-Design [Aic01]. Test-Driven [LdPK^{+}14]. TESTAF [NuR05]. Testing [ABB14, AO04, Bol06, GR01, Gir05, GGP08b, Kon03, LRB16, MPPS95, NuR05, Nør10, OCB^{+}10, SCT09, Vel04]. Text [AZMA15, Bjø01, BHQW02, BM11, CKPK13, CDR^{+}09, Fen95a, Fen97, HPB12, HGMT08, KD11, KDKB07, LALS08, QF12, SHH10, SUKG13, SOO97, THJ16, VV15, VC13, VTHM16]. Texts [GBHA12, KBF^{+}11]. Textual [BCCH11]. Texture [FL10]. Textured [AT10a]. th [RK97, ZZ00]. Their [Win97a, ASH11, Aur01, BCM08, BH00, BVG08, FP06, FSMP07, GNP05, LTY^{+}16, MMP15, MYT09, QQ11, TD96]. Thematic [LALS08, RAC10]. Theorem [Ber05, BB10b, GL00, Gro00, IK97, KO99, Sch09c, Vel05, Sch05c]. Theorems [BB08b, BG97, CS05a, DS10, Lin04c, Wan95]. Theoretic [FPT10]. Theoretical [BCNR07, CS02, HYC^{+}05, RH10, SDJ99, BR07]. Theoretically [AP09]. Theories [KS10]. Theory [Abr07, Aïm98, Cal96b, DS08, FMT^{+}15, GSW04, GL11c, Hem06, JGW11, KKTT16, MJ13, Mar96, NR12, Nag06, Rat05, SH06, Tha10, VTHM16, ZS04, Svo96]. Theory-based [FMT^{+}15]. Thesauri [Rie02]. Thesaurus [LdPK^{+}14].

Thesaurus-Based [LdPK^{+}14]. Things [BCG^{+}14, DMCM14, LAHZ^{+}15, NLLJ12, QC12, ZNX^{+}12]. Think [LM94b]. thinking [SM04]. though [BZ09]. Thoughts [Bjø01]. Thousand [CG09]. Thread [CW00, LJ00]. Threads [Lep98]. Threat [ABAL09]. Three [BGK02, DRA^{+}04, Hon01, Shi11]. Three-Tier [BGK02]. Throttle [MT99]. Tier [BGK02]. Tietze [Sch09c]. Tile [ACR11, Mar02b]. Tiling [Mar02b, Mar04]. Tilings [MS03, Mar06a]. Time [AdMGPV06, And97, BS11, BG97, CVM11, CS10, CPC00, CP02, DF05, DOM10, DDJ^{+}11, DS03, Dru06, EGK^{+}12, EHEH05, EGDG09, Faz06, FDR^{+}15, Fon00, GWW05, HJRW97, HCH^{+}09, HCK11, KCKL10, Köh09, LEC11, Loo06, MN00, ME03, NWDX09, NR08b, OD03, OFCB08, ODSO11, OL08, PMLL09, PD99, RSVR01, RR06b, Ruf01, SZZM10, SF00, Ste00, Tak06, dSC06b, dG15, AT07, Bai12, CSC08, CPS07, DRRGdP07, DHC11, GL11a, Kim10, LZLW13, TY09, ZTN^{+}15, ZHG^{+}06]. Time-based [Kim10]. Time-Dependent [OFCB08]. Time-Release [Tak06]. Time-Series [Dru06]. Time-space [CP02]. Time-varying [Loo06]. Timed [MK99, OCRPdlMG07, OL08, PS09, SV08]. Timed-event [PS09]. Timelines [ATGP09]. TIPTOE [GLD^{+}12]. Tissue [MPK04]. Tolerance [LMRG14]. Tolerant [Gär99, Log06]. Tomescu [CSZ07]. Tomorrow [LMO01]. Too [Din97]. Tool [AGA12, CKdL08, CPR06, DSM13, DBB13, DS08, FG03, GÁVCNC14, GGPTdP11, GLSD11, GRCGK14, GPSV03, HVCA12, Hor04, Kar13, KHG10, LMRG14, MOG^{+}10, MS05, OD03, SSAB^{+}13, SBAZ11, SV05, SSGS10, TAL08, dSBGdAdLM08]. Tool-based [DSM13]. Tooling [MHLB12]. Toolkit [FP05]. Tools [BHRS03, BMG^{+}05, BVG08, EGDG09, FZ00, GR14, GJP^{+}12a, LSR10, Mar00b, Mat04, NCH16, PSVOVI07, PK98, RS01a, RS01b, RC10, SGB^{+}13, SR00, SL09, SMP^{+}11, SBG^{+}12, VGBLGS^{+}08, vB96, PS97]. Toolset [FGBS14, LB98, AKP01]. Top [HA03, KKB12, NNT16]. Top- [KKB12, NNT16]. Top-down [HA03]. Topic [BHQW02]. Topological [RWZ09, Spa08, ZZ07]. Topologies [LSV06]. Topology [ABAL09, CS05a, KD05, LKK08, ME03, Pal05, WG09, Wei10]. Total [SCS13, Tur04]. Totality [SCS13]. Touch [CBNDR10]. Tour [Ram01]. Tourism [WCH14]. Tourist [FKK^{+}10]. Toy [MM96]. ToyLisp [ŞM96]. Traces [HZZ^{+}12]. Tracing [VSGP05]. Track [OMO10]. Track-To-Track [OMO10]. Tracking [AT10b, Azz10, GOM^{+}13, SVK^{+}15, SCK^{+}09, TG10]. Trade [BCZ04, GFBR08]. Trade-Off [BCZ04]. Trade-Offs [GFBR08]. Tradeoff [FMT^{+}15]. Tradeoffs [YJY14]. Trader [FAT^{+}13]. Trading [CO08, FAT^{+}13, SLT08]. Traditional [Odl94, RGRR15]. Traffic [Che09, GH08, HCH^{+}09, KDKDN08, LF98, PS12]. Tragic [Odl94]. TRAILER [GPCZ^{+}13]. Trails [GR02]. TrailTRECer [GR02]. Training [AHSN01, DM04, ES03, FSELC13, Mau97i, MPF^{+}16, PMAM14, PD04, RKH15]. Trainings [ARS^{+}08]. Trajectories [VTRN12]. Trajectory [LL97]. Transaction [Kir09]. Transactional [ASAIN14, HCD10]. Transactions [BFN05, DMG07]. Transclusions [Kol05, KM01]. Transfer [BM05, CT08b, LKT10, MH02, MPF04, RN03, VAS05]. Transference [LRR04]. Transformation [CCP11, DSLO04, Jun08, KASS03, KM06, MYCA11, MP10, PEPP08, VBB13]. Transformational [Gär99]. Transformations [BRF^{+}09, CR07, SW10, SBCJ03, SAB99, SK04]. Transgenetic [GGS08]. Transit [BEH08]. Transition [Jon01, Luk08, PR06, VJTJ07]. Transition-Net [Jun01]. Transitive [Ban97b]. Translates [HL03]. Translation [BBP95, CGP07b, KBN14, MI05, ZPFG03]. Transmission [CLCC10, PZJ09]. Transparency [RN03]. Transparent [GL11a]. Transposition [GD14]. Trapezoidal [CJ98]. Traps [PF15]. Traversals [CNQ04]. Treasure [KY10]. Treat [Ior07, Ior08]. Tree [BC16, Géc02, GGS08, HNS07, PZJ09, Sar05, Sch99, UCM13, WZZ^{+}09, XWGS09, LHC^{+}13]. Trees [AMUFVI09, BR05, DT07, GM05, Lep95, RRR10, Rah99, Sch05a, TNRGCP^{+}13]. Treeworld [Ski00]. Trending [BMUF14]. Trends [BRO08, CSY02b, JK10a, KGK12, LUR16, LG08, VGCPAH16, ZDI10]. Triangulations [HOS96]. Trigger [GWW05]. Tropos [MMM^{+}12a]. Trust [CAD^{+}06, KW10, LGAP11, LWL10, Pre04, SLD^{+}16, WLKW11]. Trust-Oriented [LWL10]. Trusted [DH10a, Die10, Lip10, PT10]. Tube [BM05]. Tubes [Ede01]. Tuning [LN08, Vel04]. Tuple [LRI03]. Turing [Hem99]. Turkey [KCK10]. Turkish [Tüf13, UHÖD15]. Tutorial [CPR06]. Tutoring [KHG10]. Tutors [HAFS15, KHG10]. TV [CFMP15, HL96, LZ09, SJ13]. Tweeters [DBB13]. Tweets [AS14]. TwiSNER [THJ16]. Twitter [CBRH12, CKPK13, DBB13, KSR16, dIdSGZ10, THJ16]. Two [APNA12, BPC04, BCNR07, CR00a, GBP^{+}08, GK13, KNLS00, Lin11b, Log04, Mar02a, Mau97i, PS97, VCB08, BRAS^{+}12, CCHdCN08, DT07]. Two-Element [GK13]. Two-handed [VCB08]. Two-Level [KNLS00]. Two-Stage [Log04]. Type [AAAG95, Hri02, MMdMGM06, MS03, Sou99, VFC03, Win97a]. Type-safe [MMdMGM06]. Typed [FPLS03, Xi03]. Types [MHLB12, Moo08, RMM^{+}08, VO10, Win97a]. Typicality [FL14]. Typicality-Based [FL14]. Typology [NL10]. TYPUS [MR14].

U [YKD^{+}08]. U-campus [YKD^{+}08]. U.S. [SC14]. Ubiquitous [BBdOR14, BHC05, BAR06, CJH12, CDCH09, GHHA10, JK10b, KKK16, Kim10, LSG^{+}14, LGP10, PMRO08, Par09, TFG06, TGLP10, WKXL05, YLZW10]. UCA [SCS13]. UCL [Lei08]. UCL-GLORP [Lei08]. UCS [Mau03a]. Ultra [CLCC10, EIH08]. Ultra-Sonography [CLCC10]. UML [Ada06, Are02, CM00, GAMP10, JGM10, KCKL10, LM10, MM06, MM03, MCC13, RH10, RFMLP10, SSGS10, Tra00, YMP08, Yeu04]. UML-Based [Ada06, JGM10]. Unambiguity [Hon95]. Unavoidable [ELFAR15]. Unbounded [Cal96a]. Uncertain [Pop05]. Uncertainties [AM96]. Uncertainty [DHP03, MLX10, MNDRF10, Rat00, WFOC98, XLMR10, YX10]. Uncolorability [RR06a]. Unconstrained [BSB09, Mes02]. Uncovering [LCC^{+}12]. Undecidability [MGMS12]. Undecidable [HI99]. Underestimate [Sch02c]. Undergraduate [HMM00]. Underlying [JJ12]. Understand [KKTT16]. Understanding [ES05, JK12b, KW10, Liy02, MR14, PJH12, SL09, TCW12, VMFO14]. Underutilised [BKH^{+}13]. Uneven [KD11]. Unfolding [JMP06]. Unhygienic [CD10]. Unification [SDÖ^{+}12, Wol99a, Wol99b, Wol00]. Unified [ASW^{+}03, KT10, Mau96a]. Uniform [Ber05]. Unifying [CJ07]. UniGame [PD04]. Union [Tom03]. Uniquely [MVM00]. Uniqueness [Sch05c]. Unit [MYT09, Nør10, PKSR09]. Unitarily [Ior07, Ior08]. UNITE [SBS15]. Units [CRLNAR05, CRMLN^{+}07, NOP08]. Universal [AR01, CMS94, Leš09, Roj96]. Universe [Cha05]. Universities [Len00, PD04]. University [DO01, FP12, LRR04, MRK^{+}98, UHÖD15, HAS^{+}07]. Unknown [AMR^{+}14, CT08a, CW09]. Unportable [Rho10]. Unstructured [HMSS01, QF12]. Unsupervised [GC14, Suz06]. UOWHF [Sar05]. Update [JRW10, Köh09, MHA^{+}15]. Updates [GT01, STW09]. Updating [GRHMM^{+}15]. upon [ASH11, SLK11]. Upper [HYC^{+}05]. ups [KKM08]. Urgent [BARR09]. URL [APJK09]. Urysohn [Leš09, Sch09c]. Usability [AFP^{+}13, GGPTdP11, GGP08b, HTHW12, Kar13, Kim10, LRB16, LULGFC13, MGT14, PHJ^{+}08, TT98, Tüf13, UT16]. Usage [Jun05b, KD02, KK13b, MRP14, NSF^{+}10, OEK16, vBK08]. Usage-based [NSF^{+}10]. Usage-Centered [KD02]. Use [BU13, BQBW03, Bra15, BNCGD^{+}11, CCP^{+}07, CM00, Die96, FBCMR15, HBT12, KASS03, Len00, MH98, MS05, PBTW07, RMF^{+}98, SGB^{+}13, STFM12, SBS15, SD97, Tez13, WPL98]. Used [ATGP09, DAd03, KSdV09]. Useful [CC96]. User [APJK09, ASS13, BDL^{+}06, BE98, BVV^{+}10, CIM14, CFMP15, CWT^{+}15, CHH16, CS09, DAk13, DA13, EDA14, FGB^{+}14, FDC^{+}13, Flo04, GWW05, GHHA10, GLCV08, GGDBP^{+}08, GL11b, HOPN11, HT06, IPCVC12, JKKW16, JL16, JJ12, Jun05a, KWH03, LASL12, LVS13, LMMPFVI14, MR08, MROH08, NOGG^{+}13, NPC^{+}09, PRB^{+}11, PMRO08, PLBG13, PLG^{+}08, PMAM14, PLSF08, SBMD10, SW04, SG02, SLJC08, Sha11, Sob05, SKH14, SMV08, SKL08, SSV02, TGEM07, DLY08, Pel14, SW09, TJS^{+}13, CEK15]. User-Aware [EDA14]. User-Centered [CIM14]. User-defined [PRB^{+}11]. User-Friendly [LMMPFVI14]. User-Generated [SKL08]. User-given [GWW05]. Users [GP10, HKKvP08, HB11, KM07, PT09, TCW12, RvS12]. Uses [SH96]. USF [Par09]. USF-PAS [Par09]. Using [AMR^{+}14, Ad03, ARS^{+}08, AH04, BS12a, BB08a, Bai12, BDPSNG97, BGBA10, BCG98, BCM12, BHQW02, CH07, CE06, CJ98, CM11, CR07, CGP^{+}07a, CPCLSAGC11, CS02, DB03, DGK^{+}99, Dit02, Dru06, EMZB14, EC00, FGB^{+}14, FBCMR15, FF04, GÁVCNC14, Gir05, Gle03, GCVRSPGP07, Hef04, HLNA^{+}12, HL03, HSD^{+}14, IO04, JBBH13, JS16, KCK10, KC08, KASN08, KKB12, KSR16, Kir09, KHN99, KDKDN08, KNSN07, KMN16, KR11, LALS08, LLLL99, LKP11, LS07, LPSF10, LVS13, Log04, LKZK10, MCM07, MTB^{+}08, MSM07, MRO05, MAT08, NAK08, NO98, NZCG05, Oli12, Pal15, PRT^{+}08, PA12, PT09, PMRO08, PSS^{+}13, PCS^{+}13, POJB08, PCKJ11, PPG95, Rad14, RAWW05, Ret08a, Rho10, Rie02, SESMT10, SGB^{+}13, SLD^{+}16, Sch01a, STW09, SH11, STBFM09, Str97, SLPS98, Tab07, TNRGCP^{+}13, TEK08, THS11]. Using [UDC97, VTHM16, ZAB^{+}08, APJK09, AT10b, AK09, AGO^{+}13, AP05, BZA08, CLVM09, CT16, CDR^{+}09, DB12, DDJ^{+}11, FZAP13, FJP06, GRGN13, HK14, HGS^{+}08, HAI13, HG11, KH12, LNML03, LKB^{+}02, MOG^{+}10, MP10, Mue04, NuR05, PLSF08, Puc10, Ste00, SKH^{+}10, TG10, TJ15, VV15, VdR09, VV12]. Usual [Fro02]. Utility [FLF^{+}14, RHM15]. Utility-Oriented [FLF^{+}14]. Utilization [HB11, HGMT08, Kim12, Pet09]. Utilization-level [Kim12]. Utilizing [Rob06]. Utterances [KM07]. UX [DAk13].

V [Ban97b]. Vadis [BP08]. Valid [RGP10]. Validated [ATOFF98]. Validating [BGBA10, Hei07, OCB^{+}10, dOBGH^{+}14]. Validation [AUN04, EK99, MP10, MSF99, SBPR15, STW09, ZPFG03]. Validity [UT16]. Value [CCD03, SKH12, Vel05]. Valued [AS14, YX10, BNA15, Ior00]. Values [Akr09, KF10a, KF10b, LF16, SS09a, Zgr07]. Valve [DB12]. Valve-point [DB12]. VANET [ARQH14]. VANETs [MGCGCG12, MKA11]. Variability [HSD^{+}14, JGM10]. Variable [AMS04, ASAIN14, Fen15, GHNT97, LA07].

Variable-Length [Fen15, GHNT97]. Variable-Ordering [AMS04]. Variables [DMCM14, RLL^{+}10]. Variant [Bod01, EGK^{+}12, FP95, NSNK05]. Variants [Hon01]. Variations [FP06]. Various [AV07, CT16, LWG14, dTU04]. Varying [DRRGdP07, Loo06]. Vascular [BDhKB09]. VBSME [Oli12]. VCA [QZ07]. VCC [ALHM^{+}14]. VDL [Jon01]. VDM [Jon01, Oli01]. VDMTools(R) [Lar01]. VDTNs [PSS^{+}13]. Vector [CLVM09, WC09]. Vectors [Ara03]. Vehicle [XHP^{+}09, dG15]. Vehicular [HCH^{+}09]. Vein [BDhKB09]. Velocity [DBBS08]. Vendor [CPFSdAS12, MPM12]. Verb [JM15]. Verifiable [NZCG05, NSNK05]. Verification [AS07, ADMdB09, AUN04, BLS01, BHRS03, BCD13, BCG^{+}14, CP06, Gär99, Har07, KKK^{+}14, KZ03, LDSG09, LD06, LKZK10, MM03, Mat04, NI03, Ois98, OD03, OJSB08, POJB08, RS01a, RS01b, RVC12, Ruf01, SK13, SNAF07, Sch01a, SBCJ03, SMSdB05, SdBM05, TSCY01, Vel04, WTA01, dFCC07, dMTS^{+}14, WB07]. Verified [Lan98]. Verifying [AdMGPV06, Bur05, Hei07, Ste00]. Versioned [MMdMGM06]. Versions [Gaß10]. versus [BCZ04, CVPS95, PK98]. Vertical [AAM14]. Vertices [AR95]. via [Bor07, CBRH12, GOF05, KS05, KL09b, Lep98, Pel14, PJH12, RTB13, Rud04, SSS12, Yon11, ZC09]. Video [HL96, LVS13, MM15, OHYJ16, PZLAS^{+}13, RMF^{+}98, SLK11, STVT07, Wac02, ZTN^{+}15]. Videogaming [SdBC13]. View [Ern11, HP15, KB06, Mac01, SCK^{+}09]. Viewes [RTB13]. Viewpoint [Nav09]. Views [BCG98, Heg10, XLMR10]. VIKAMINE [AP05]. VIMM [Sue10]. Virtual [AY12, dMBHR15, CT16, CDG14, DO01, FDC^{+}13, GS12, GMdMC12, GCL^{+}13, IRMK12, IBN^{+}11, JKKW16, JGL08, KM13, KHLAP12, KJKS14, uRLFH^{+}13, Lin04a, LLSA13, LAHZ^{+}15, MOG^{+}10, MPF^{+}16, MGT14, MCMMAP^{+}14, NSFVH05, PRT^{+}08, PdlCBKN14, PGSAP14, PT10, PCC14, RS11, RPR11, Ros05, SVFR15, SdOB09, SE09, SH96, SBS15, Tom01, VBP^{+}11, VDBNR98, DDS04]. Virtuality [TKK^{+}12]. Virtualized [Sue10]. Visible [GPCZ^{+}13]. Vision [BdI10, FPSFCG07, JRO10, ZTN^{+}15]. Visions [RS03]. Visited [KTL^{+}11]. Visitor [MCM07, Sch06, Ver10]. Visual [ARFT05, AP05, BGBA10, BNA15, CS03, CE11, DCR^{+}07, FTARR05a, GGPTdP11, KC08, KS05, LSR10, LM15, NPB06, SHH10, SCLM03, TCK^{+}01, ZS04, dlFVPHB15]. Visualisation [KL02]. Visualise [HSD^{+}14]. Visualization [ATV98, AMUFVI09, BHS^{+}06, BM05, DBBS08, Epp04, GOF05, HBF10, HMA^{+}05, Jun10a, LSR10, POR10, Pop05, TKT07]. Visualizations [HLK09a]. Visualizing [Jun05c, Wac02, dCH11]. VIVO [LJLCR13]. VLIW [For97]. VOC [ZPFG03]. Vocabulary [EIH08]. Voice [MNS^{+}12]. Voice-Over [MNS^{+}12]. Voices [PSS07]. Volunteer [VMFO14]. Voronoi [CT16]. Voting [QZ07]. Voting-Based [QZ07]. vs [BM96]. Vulnerabilities [AA16b, LM03a].

WAIS [Mue95]. Wait [BM96]. Wait-Freedom [BM96]. Wajsberg [Cet00]. Walk [GN10]. Walkthroughs [GGP08b]. WAM [SA97]. Wang [CK95]. Warehouse [SS15, ZSK09]. Warehouses [STBFM09, dH04]. Warehousing [MSTW11]. Warming [YYZ^{+}09]. Warming-up [YYZ^{+}09]. was [PTL^{+}09]. Water [BM11]. Watermarking [HPC10]. Wave [JJ12, PD10]. Wavelets [LSK06]. Waves [Ang98]. Way [Ukk10, Lan10]. Weak [BI08, CR00a, FF08, HL09, Sch08b]. Weaknesses [BPSN97]. Wearable [HLNA^{+}12]. Weaving [LAM12, LRS^{+}11, PZ03]. WEB [Ada06, BCG98, AA16b, ASTL07, AGMT10, AHSN01, AdO11, ARS^{+}08, AVA08, BCFM05, BGMR^{+}16, BC11, BARR09, BCM08, BCNR07, CVM11, CM09, CCH06, CI12, CG96b, CLC09, CLM10, CPMVG13, CAR08, CCS10, De 96, DM04, DZBB^{+}12, DCS09, DR04, DDJ^{+}11, EDA14, EMGB^{+}12, FMB^{+}11, FVIG12, GG08, GHS06, GO14, GL11c, GL11d, GC14, HM99, HPE14, Hef04, HBI98, HB11, IAS16, JST11, Jun05a, Jun05b, KC08, KJZJ08, Kim10, KY10, KOW01, KM06, KTKP09, KB06, KMM14, KJKS14, KR11, LB98, LAM12, LN08, LF98, MM98, MNL13, Mau97i, MSM07, MCG14, MPR^{+}08, MP06, Nav09, NS05, OCB^{+}10, PTL11, PSVOVI07, PD10, PS09, PCLCC11, PLSF08, RTB13, RRM^{+}12, RJB10, SGB^{+}13, SMFM05, Sch96, SBTH04, SH09, SC14, SA09, Sob05, TT98, THS11]. Web [TSDP07, URG^{+}13, VAH07, VdSdMC08, VTGA13, WZZ^{+}09, Wit08, ZGC^{+}08, ZD09, dCVM12]. Web-Based [CCH06, HPE14, Mau97i, SBTH04, AHSN01, CCS10, DM04, DZBB^{+}12, KY10, KOW01, KTKP09, NS05, SH09]. Web-Decision [CM09]. Web-Services [PS09]. WebA [TAL08]. WebAnima [PT09]. Weblog [HLK09b, SK08]. Weblogs [RAWW05]. WebQuest [PFS07]. Webs [FGS98, Sto02]. Website [ASS13, GMP^{+}13, GGP08b, IBN^{+}11]. Websites [GMP^{+}13, TAL08]. Week [PPJ04]. Weekend [SA14]. Weight [BSHI99, KJL09, WKXL05, SR10]. Weighted [AT10a, Bos08a, Cv99]. Weights [Fen15]. Well [LM94b]. weSPOT [MOS^{+}13]. WFA [CVK97]. Wheelchair [PdlCBKN14]. Whiteboard [VPF09]. Whiteboards [KSdV09]. Who [Duv01, vB96, SdBC13]. Wide [MM98, TT98]. Widget [GJP^{+}12a, Kro13]. Widget- [Kro13]. Widgets [GJP^{+}12b]. Wiener [CW09]. Wiki [HNJ^{+}10, MMB08, SA11]. Wiki-based [MMB08]. Wikipedia [KKB12]. Wikipedia-Based [KKB12]. Wikis [Nal10]. Will [LM94b, Pos01]. WiMAX [LCZ^{+}12]. WiMAX-R [LCZ^{+}12]. Window [KTJ05, KL09b]. Windows [dG15, MBC12, SH10, SAKAM11]. Wireless [AAJR05, BCHM12, CMZZ07, GBCA12, HM99, Kon03, LGAP11, LZ09, LLYC12, LSV06, LWY11, MWM10, MdCRMP14, NOGG^{+}13, NVB12, OO08, PKP08, PZDH09, PdCdlTR06, SHZ^{+}10, WXZL15, YLW^{+}14]. within [AA16b, AMBP04, BH01, Dru12, FML13, FCM^{+}12, FP05, HW10, Lin04a, MNF^{+}13, RLMS13, SHZ^{+}10, VK03, ZBKK12]. without [BJ97, Kie05b]. WLAN [CLCC10]. Wonders [Car00]. Wood [Jür10, JMSY10, Sal10]. Wookie [GJP^{+}12a]. Word [Car96, HHY02, Kah01, RdKO11]. WordNet [Hri02]. Words [CSY02a, EIH08, Her96, HI99, Mar02a, PS04, MPRS95, Kah01]. Work [ES03, Ern11, GlRBdSG11, GL11b, Lep95, LA03a, Mai05, Man97, SD97, TCS^{+}03, WKSD^{+}11, FCM^{+}12]. Work-greedy [Man97]. Work-Optimally [Lep95]. Worker [GL11b]. Workers [ASH11, MTB^{+}08, dG15]. Workflow [Dus05, GLCV08, MPR^{+}08, MP06, TKF06, XZSS09]. Workforce [VUT^{+}08]. Workgroups [SSdS^{+}11]. Working [KBF^{+}11, SCLM03]. Workload [HVM00]. Workplace [Car00, MX05]. Workshop [AUN04, CMZZ07]. Workspace [GGB^{+}08]. Workspaces [MPG13]. Workstations [EK00]. World [BGBA10, PdlCBKN14, PMAM14, SBS15, She96, SSBS08, VBP^{+}11, XZ00, MM98, TT98]. WorldOfQuestions [IRMK12]. Worlds [CVFN07, GMdMC12, IRMK12, LLSA13, MPF^{+}16, MCMMAP^{+}14]. Wrapped [Ede01]. Wrapping [CAR08]. Writing [Car95, KD11, KWC01, OF13, WBS12, WKTL01]. WS [DMCM14, PdlCBKN14].

WS-BPEL [DMCM14]. WWW

[Cai95, GBHA12, HVM00, Hop98, MH96, PV95, PSF98, Rad96, Reb96]. WWW-Aided [Reb96]. WWW-Based [Hop98].

X [BCG^{+}99, GV00, dTU04]. X-Global [dTU04]. X-Machine [GV00]. X-Machines [BCG^{+}99]. XEN [Puc10]. XII [VMdCJ08]. XML [BVG08, KBN14, LKB^{+}02, MN14, PRAT09, RRB03, STW09, SW10, WD02, dCH11, dTU04]. XML-based [WD02]. XML-Enabled [PRAT09]. XOCL [RRB03]. XP [SH10]. XPath [AJBTEB06, SSSS10].

YAP [dSC06a]. Year [MC07]. Years [Boe97c, HA10, JMSY10, Lar01]. Yosida [CS05a]. Young [FPS^{+}12].

Z [BRW03, MC00]. Z-Specifications [BRW03]. zero [Sch10]. Ziv [Fen95a, Log04]. Zoning [FOAB08]. Zoom [RTB13]. Zoom-Based [RTB13]. ZRTP [HGS^{+}08, PHJ^{+}08]. ZZ [HL03].

References

Abirami:2016:FBS

[AA16a] A. M. Abirami and A. Askarunisa. Feature based sentiment analysis for service reviews. J.UCS: Journal of Universal Computer Science, 22(5):650– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                                         www.jucs.org/jucs_22_
                                                         5/feature_based_sentiment_
                                                         analysis.

Abunadi:2016:EIS

[AA16b] I. Abunadi and M. Alenezi.

An empirical investigation of security vulnerabilities within Web applications. J.UCS: Journal of Universal Computer Science, 22(4):537– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_22_
4/an_empirical_investigation_
of.

Aichholzer:1995:NTS

[AAAG95] Oswin Aichholzer, Franz Aurenhammer, David Alberts, and Bernd Gärtner. A novel type of skeleton for polygons. J.UCS: Journal of Universal Computer Science, 1(12): 752–761, December 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/a_novel_type_of_
skeleton_for_polygons.

Alotaibi:2015:MML

[AAAK15] H. M. Alotaibi, R. A.

Alamer, and H. S. Al-Khalifa. MLab: A mobile language learning lab system for language learners. J.UCS: Journal of Universal Computer Science, 21(10):1307–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_21_10/mlab_
                                        a_mobile_language.

Amati:1997:FFN

[AAGU97] G. Amati, D. D. Aloisi, V. Giannini, and F. Ubaldini. A framework for filtering news and managing distributed data. J.UCS: Journal of Universal Computer Science, 3(8):1007–1021, August 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs_3_8/a_framework_
                                        for_filtering;   internal&
                                        sk=05460486.

Adelstein:2005:PLW

[AAJR05] F. Adelstein, P. Alla, R. Joyce, and G. G. Richard, III. Physically locating wireless intruders. J.UCS: Journal of Universal Computer Science, 11(1):4– 19, January 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_11_1/physically_
                                        locating_wireless_intruders.

Abdalla:2014:NVF

[AAM14] H. I. Abdalla, A. A. Amer, and H. Mathkour. A novel vertical fragmentation, replication and allocation model in DDBSs. J.UCS: Journal of Universal Computer Science, 20(10):1469–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
10/a_novel_vertical_
fragmentation.

Antunes:2009:CTE

[AB09] P. Antunes and R. O.

Briggs. Collaborative technology and environments. J.UCS: Journal of Universal Computer Science, 15(16):3079–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15;
 http://www.jucs.org/
jucs_15_16#;   http:
//www.jucs.org/jucs_
15_16/collaborative_
technology_and_environments.

Abdelouahab:2009:TCA

[ABAL09] M. A. Abdelouahab, A. Bouabdallah, M. Achemlal, and S. Laniepce. The topology change attack: Threat and impact. J.UCS: Journal of Universal Computer Science, 15(2):465– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15_
                                        2/the_topology_change_
                                        attack.

Accioly:2014:CEC

[ABB14] P. Accioly, P. Borba, and R. Bonifacio. Controlled experiments comparing black-box testing strategies for software product lines. J.UCS: Journal of Universal Computer Science, 20(5):615–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_20_5/controlled_
                                        experiments_comparing_
                                        black.

Arroyo:2002:MCP

[ABCP02] F. Arroyo, A. Baranda, J. Castellanos, and G. Paun. Membrane computing: The power of (rule) creation. J.UCS: Journal of Universal Computer Science, 8(3):369– 381, March 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        3/the_power_of_rule.

Antal:2006:MEP

[ABFJ06] P. Antal, N. Bátfai, I. Fazekas, and P. Jeszenszky. The mobiDIÁK educational portal. J.UCS: Journal of Universal Computer Science, 12(9):

1118–1127, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_9/the_mobidiak_educational_
portal.

Augusto:2006:PHM

[ABM^{+}06] J. C. Augusto, N. D.

Black, H. G. McAllister, P. J. McCullagh, and C. D. Nugent. Pervasive health management: New challenges for health informatics. J.UCS: Journal of Universal Computer Science, 12(1):1–5, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12;
 http://www.jucs.org/
jucs_12_1#;   http://
www.jucs.org/jucs_12_
1/pervasive_health_management_
new.

Andreoli:1995:CAI

[ABPS95] J. M. Andreoli, U. M.

Borghoff, R. Pareschi, and J. H. Schlichter. Constraint agents for the information age. J.UCS: Journal of Universal Computer Science, 1(12): 762–789, December 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
constraint_agents_for_
the_information_age.

Abrial:2007:FMT

[Abr07]  J.-R. Abrial.  Formal

methods: Theory becoming practice. J.UCS: Journal of Universal Computer Science, 13(5):619– 628, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        13_5/formal_methods_
                                        theory_becoming.

Abrial:2005:FCN

[AC05] J.-R. Abrial and D. Cansell. Formal construction of a non-blocking concurrent queue algorithm (a case study in atomicity). J.UCS: Journal of Universal Computer Science, 11(5):744–770, May 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_11_5/formal_
                                        construction_of_a.

Andersen:2007:SDS

[AC07] R. Andersen and S. M. Cioaba. Spectral densest subgraph and independence number of a graph. J.UCS: Journal of Universal Computer Science, 13(11):1501–1513, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_

11/spectral_densest_
subgraph_and.

Alzubi:2016:SCC

[ACA^{+}16] O. A. Alzubi, T. M.

Chen, J. A. Alzubi, H. Rashaideh, and N. Al-Najdawi. Secure channel coding schemes based on algebraic–geometric codes over Hermitian curves. J.UCS: Journal of Universal Computer Science, 22(4):552–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
22_4/secure_channel_
coding_schemes.

Allende-Cid:2015:DNR

[ACAMM15] H. Allende-Cid, H. Allende, R. Monge, and C. Moraga. Discrete neighborhood representations and modified stacked generalization methods for distributed regression. J.UCS: Journal of Universal Computer Science, 21(6):842– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
6/discrete_neighborhood_
representations_and.

Atnafu:2002:ECB

[ACB02] S. Atnafu, R. Chbeir, and L. Brunie. Efficient content-based and metadata retrieval in image database. J.UCS: Journal of Universal Computer Science, 8(6):613– 622, June 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
6/efficient_content_
based_and.

Achs:2006:CEF

[Ach06]  Á. Achs.  Creation

and evaluation of fuzzy knowledge-base. J.UCS: Journal of Universal Computer Science, 12(9): 1087–1103, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_9/creation_and_evaluation_
of.

Asserrhine:1995:ERE

[ACL95] J. Asserrhine, J. Chesneaux, and J. Lamotte. Estimation of round-off errors on several computers architectures. J.UCS: Journal of Universal Computer Science, 1(7):454–468, July 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
estimation_of_round_
off_errors_on_several_
computers.

Aljawarneh:2016:ARS

[ACM16] S. A. Aljawarneh, F. Cena, and A. Maatuk. Advanced research on software security design and applications. J.UCS: Journal of Universal Computer Science, 22(4):453–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_22;   http:
//www.jucs.org/jucs_
22_4/advanced_research_
on_software.

Ascia:2006:MOG

[ACP06] G. Ascia, V. Catania, and M. Palesi. A multi-objective genetic approach to mapping problem on network-on-chip. J.UCS: Journal of Universal Computer Science, 12(4):370–394, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
4/a_multi_objective_
genetic.

Agost:2011:MMC

[ACR11] M.-J. Agost, P. Company, and F. Romero. Managing mechanisms for collaborative newproduct development in the ceramic tile design chain. J.UCS: Journal of Universal Computer Science, 17(2):224– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_17_
                                        2/managing_mechanisms_
                                        for_collaborative.

Alves:2003:IEH

[Ad03] N. M. M. Alves and S. de Mello Schneider. Implementation of an embedded hardware description language using Haskell. J.UCS: Journal of Universal Computer Science, 9(8):795– 812, August 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        8/implementation_of_
                                        an_embedded.

Adamko:2006:UBM

[Ada06]  A. Adamkó.  UML-

based modeling of dataoriented WEB applications. J.UCS: Journal of Universal Computer Science, 12(9):1104–1117, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12_
                                        9/uml_based_modeling_
                                        of.

Ayter:2015:SAE

[AdGCD^{+}15] J. Ayter, d. G. Chifu, S. Déjean, C. Desclaux, and J. Mothe. Statistical analysis to establish the importance of information retrieval parameters. J.UCS: Journal of Universal Computer Science, 21(13):1767–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
13/statistical_analysis_
to_establish.

Aguilera:2016:SES

[AdI16] U. Aguilera and D. López de Ipiña. A semantically enhanced service discovery for MANET. J.UCS: Journal of Universal Computer Science, 22(7):896–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
22_7/a_semantically_
enhanced_service.

Astefnoaei:2009:SVN

[ADMdB09] L. Aştefnoaei, M. Dastani, J.-J. Meyer, and F. S. de Boer. On the semantics and verification of normative multi-agent systems. J.UCS: Journal of Universal Computer Science, 15(13):2629–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15_
                                        13/on_the_semantics_
                                        and.

Alpuente:2006:VRT

[AdMGPV06] M. Alpuente, M. del Mar Gallardo, E. Pimentel, and A. Villanueva. Verifying realtime properties of tccp programs. J.UCS: Journal of Universal Computer Science, 12(11): 1551–1573, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        12_11/verifying_real_
                                        time_properties.

AlvaresdeOliveira:2011:QBA

[AdO11] F. G. Alvares de Oliveira Jr. and J. M. Parente de Oliveira. QoS-based approach for dynamic Web service composition. J.UCS: Journal of Universal Computer Science, 17(5):712–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_17_5/qos_based_
                                        approach_for.

Attardi:1998:CC

[ADS98] G. Attardi, Marco S. Di, and D. Salvi. Categorisation by context. J.UCS: Journal of Universal Computer Science, 4(9):719–736, September 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_9/categorisation_
by_context.

Andrade:2004:CCS

[AF04] L. F. Andrade and J. L.

Fiadeiro. Composition contracts for service interaction. J.UCS: Journal of Universal Computer Science, 10(4):375–390, April 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
4/composition_contracts_
for_service.

Aichernig:2001:FAS

[AFK01] B. K. Aichernig, B. Fröhlich, and A. Kerschbaumer. Formal aspects of software engineering — J.UCS special issue in honor of Professor Peter Lucas. J.UCS: Journal of Universal Computer Science, 7(8):629–630, August 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
8/formal_aspects_of_
software.

Altenhofen:2008:ASO

[AFL08] M. Altenhofen, A. Friesen, and J. Lemcke. ASMs in service oriented architectures. J.UCS: Journal of Universal Computer Science, 14(12):2034–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_14_12/asms_
in_service_oriented.

Amor:2004:SMS

[AFP04] M. Amor, L. Fuentes, and M. Pinto. A survey of multimedia software engineering. J.UCS: Journal of Universal Computer Science, 10(4):473–498, April 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
4/a_survey_of_multimedia.

Azevedo:2013:DUD

[AFP^{+}13] D. Azevedo, B. Fonseca, H. Paredes, S. Lukosch, J. Janeiro, and R. O. Briggs. On the development and usability of a diagram-based collaborative brainstorming component. J.UCS: Journal of Universal Computer Science, 19(7):873– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_

7/on_the_development_
and.

Abraham:2006:APM

[AG06] A. Abraham and C. Grosan.

Automatic programming methodologies for electronic hardware fault monitoring. J.UCS: Journal of Universal Computer Science, 12(4):408– 431, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_12_4/automatic_
programming_methodologies_
for.

Ahmad:2012:CAM

[AGA12] M. Ahmad, S. Gruner, and M. T. Afzal. Computational analysis of medieval manuscripts: a new tool for analysis and mapping of medieval documents to modern orthography. J.UCS: Journal of Universal Computer Science, 18(20):2750–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_20/computational_
analysis_of_medieval.

Arroyo:2008:ASA

[AGG^{+}08] R. F. Arroyo, M. Gea, J. L. Garrido, P. A. Haya, and R. M. Carro. Authoring social-aware tasks on active spaces. J.UCS: Journal of Universal Computer Science, 14(17):2840–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_17/authoring_
social_aware_tasks.

Arroyo:2008:DAI

[AGGH08] R. F. Arroyo, M. Gea, J. L. Garrido, and P. A. Haya. Development of ambient intelligence systems based on collaborative task models. J.UCS: Journal of Universal Computer Science, 14(9):1545–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_9/development_of_
ambient_intelligence.

Assfalg:2010:ICB

[AGK^{+}10] J. Aßfalg, J. Gong, H.- P. Kriegel, A. Pryakhin, T. Wei, and A. Zimek. Investigating a correlation between subcellular localization and fold of proteins. J.UCS: Journal of Universal Computer Science, 16(5):604– ??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_

16_5/investigating_a_
correlation_between.

Aguilar:2010:MAG

[AGMT10] J. A. Aguilar, I. Garrigós, J.-N. Mazón, and J. Trujillo. An MDA approach for goal-oriented requirement analysis in Web engineering. J.UCS: Journal of Universal Computer Science, 16(17):2475–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
17/an_mda_approach_for.

Anguita:2013:EES

[AGO^{+}13] D. Anguita, A. Ghio, L. Oneto, X. Parra, and J. L. Reyes-Ortiz. Energy efficient Smartphonebased activity recognition using fixed-point arithmetic. J.UCS: Journal of Universal Computer Science, 19(9):1295–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_19_9/energy_
efficient_smartphone_
based.

Arora:2004:UGS

[AH04] R. Arora and M. S. Hsiao.

Using global structural relationships of signals to accelerate SAT-based combinational equivalence checking. J.UCS: Journal of Universal Computer Science, 10(12): 1597–1628, December 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_10_12/using_
global_structural_relationships.

Alario-Hoyos:2015:MMA

[AHEAS^{+}15] C. Alario-Hoyos, I. Estévez-Ayres, M. Pérez Sanagustín, D. Leony, and C. Delgado Kloos. MyLearningMentor: A mobile app to support learners participating in MOOCs. J.UCS: Journal of Universal Computer Science, 21(5):735–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
21_5/my_learning_mentor_
a.

Alario-Hoyos:2014:PCF

[AHPSCDK14] C. Alario-Hoyos, M. Pérez-Sanagustín, D. Cormier, and C. Delgado-Kloos. Proposal for a conceptual framework for educators to describe and design MOOCs. J.UCS: Journal of Universal Computer Science, 20(1):6– ??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_

20_1/proposal_for_a_
conceptual.

Ahmadi:2008:PFS

[AHRH08] O. Ahmadi, D. Hankerson, and F. Rodríguez-Henríquez. Parallel formulations of scalar multiplication on Koblitz curves. J.UCS: Journal of Universal Computer Science, 14(3):481– 504, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
3/parallel_formulations_
of_scalar.

Akahori:2001:DEW

[AHSN01] K. Akahori, H. Horiguschi, K. Suzuki, and M. Nambu. Development and evaluation of Web-based in-service training system for improving the ICT leadership of schoolteachers. J.UCS: Journal of Universal Computer Science, 7(3):211–225, March 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_7_3/development_
and_evaluation_of.

Aguirre:2009:OSN

[AHT09] C. Aguirre, R. Huerta, and L. Tsimring. Optimal serverless networks attacks, complexity and some approximate algorithms. J.UCS: Journal of Universal Computer Science, 15(14):2747–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
14/optimal_serverless_
networks_attacks.

Aichernig:2001:TDT

[Aic01] B. K. Aichernig. Testdesign through abstraction — a systematic approach based on the refinement calculus. J.UCS: Journal of Universal Computer Science, 7(8): 710–735, August 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_7_8/test_design_
through_abstraction.

Aimeur:1998:AAC

[Aïm98]  E. Aïmeur.  Applica-

tion and assessment of cognitive-dissonance theory in the learning process. J.UCS: Journal of Universal Computer Science, 4(3):216– 247, March 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_3/application_

and_assessment_of; internal&
sk=0C220489.

Almendros-Jimenez:2006:MSX

[AJBTEB06] J. M. Almendros-Jiménez, A. Becerra-Terón, and F. J. Enciso-Baños. Magic sets for the XPath language. J.UCS: Journal of Universal Computer Science, 12(11):1651–1678, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
11/magic_sets_for_the.

Arbab:2005:CCR

[AK05] F. Arbab and J. N.

Kok. Compositional construction and reasoning techniques for software. J.UCS: Journal of Universal Computer Science, 11(10):1577–1579, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11;
http://www.jucs.org/
jucs_11_10#;   http:
//www.jucs.org/jucs_
11_10/compositional_
construction_and_reasoning.

Amornchewin:2009:MDD

[AK09] R. Amornchewin and W. Kreesuradej. Mining dynamic databases using probability-based incremental association rule discovery algorithm. J.UCS: Journal of Universal Computer Science, 15(12):2409–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_12/mining_dynamic_
databases_using.

Akinwande:2009:AHC

[Aki09] M. Akinwande. Advances in homomorphic cryptosystems. J.UCS: Journal of Universal Computer Science, 15(3):506– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
3/advances_in_homomorphic_
cryptosystems.

Andrews:1995:HGN

[AKM95] K. Andrews, F. Kappe, and H. Maurer. The Hyper-G network information system. J.UCS: Journal of Universal Computer Science, 1(4): 206–220, April 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
4/the_hyper_g_network.

Afzal:2007:CLF

[AKM07] M. T. Afzal, N. Kulathuramaiyer, and H. Maurer. Creating links into the future. J.UCS: Journal of Universal Computer Science, 13(9):1234– 1245, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_9/creating_links_
into_the.

Andrews:1994:SGH

[AKMS94] K. Andrews, F. Kappe, H. Maurer, and K. Schmaranz. On second generation hypermedia systems. J.UCS: Journal of Universal Computer Science, 0 (0):127–136, November 15, 1994. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_0_
0/on_second_generation_
hypermedia.

Aaltonen:2001:DTN

[AKP01] T. Aaltonen, M. Katara, and Pitkänen. DisCo Toolset — the new generation. J.UCS: Journal of Universal Computer Science, 7(1):3–18, January 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
1/disco_toolset_the_
new.

Akritas:2009:LQC

[Akr09] A. G. Akritas. Linear and quadratic complexity bounds on the values of the positive roots of polynomials. J.UCS: Journal of Universal Computer Science, 15(3):523– ??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_3/linear_and_
quadratic_complexity.

Alhazov:2004:DEC

[Alh04]  A. Alhazov.  On de-

terminism of evolutioncommunication P systems. J.UCS: Journal of Universal Computer Science, 10(5):502–508, May 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
5/on_determinism_of_
evolution.

Alferez:2014:CCE

[ALHM^{+}14] M. Alférez, R. E. Lopez-Herrejón, A. Moreira, V. Amaral, and A. Egyed. Consistency checking in early software product line specifications — the VCC approach. J.UCS: Journal of Universal Computer Science, 20(5): 640–??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_

20_5/consistency_checking_
in_early.

Almonayyes:2006:MED

[Alm06] A. Almonayyes. Multiple explanations driven naïve Bayes classifier. J.UCS: Journal of Universal Computer Science, 12(2):127–139, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_2/multiple_explanations_
driven_naive.

Akyildiz:1996:CFI

[AM96] Y. Akyildiz and S. Markov.

Curve fitting and interpolation of biological data under uncertainties. J.UCS: Journal of Universal Computer Science, 2(2):59–69, February 18, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/curve_
fitting_and_interpolation_
of_biological_data_under_
uncertainties.

Afzal:2011:ERS

[AM11] M. T. Afzal and H. Maurer. Expertise recommender system for scientific community. J.UCS: Journal of Universal Computer Science, 17 (11):1529–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_11/expertise_
recommender_system_for.

Alamri:2014:MCA

[AMA^{+}14] A. Alamri, G. Muhammad, A. A. Al Elaiwi, K. N. Al-Mutib, and M. S. Hossain. Media content adaptation framework for technology enhanced mobile elearning. J.UCS: Journal of Universal Computer Science, 20(15): 2016–??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
20_15/media_content_
adaptation_framework.

Apostolou:2004:FKE

[AMBP04] D. Apostolou, G. Mentzas, K. Baraboutis, and S. Papadopoulou. Facilitating knowledge exchange and decision making within learning networks. J.UCS: Journal of Universal Computer Science, 10(3):205–226, March 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
3/facilitating_knowledge_
exchange_and.

Alier:2012:CPE

[AMC^{+}12] M. Alier, E. Mayol, M. J.

Casañ, J. Piguillem, J. W. Merriman, M. Á. Conde, F. J. García-Peñalvo, W. Tebben, and C. Severance. Clustering projects for eLearning interoperability. J.UCS: Journal of Universal Computer Science, 18(1):106–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_1/clustering_
projects_for_eLearning.

Algabri:2014:SLM

[AMR^{+}14] M. Algabri, H. Mathkour, H. Ramdane, M. Alsulaiman, and K. Al-Mutib. Self-learning mobile robot navigation in unknown environment using evolutionary learning. J.UCS: Journal of Universal Computer Science, 20(10):1459–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_20_10/self_
learning_mobile_robot.

Aloul:2004:MSG

[AMS04] F. A. Aloul, I. L. Markov, and K. A. Sakallah. MINCE: a static global variable-ordering heuristic for SAT search and BDD manipulation. J.UCS:

Journal of Universal Computer Science, 10 (12):1562–1596, December 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
10_12/mince_a_static_
global.

Almeida-Martinez:2009:VST

[AMUFVI09] F. J. Almeida-Martínez, J. Urquiza-Fuentes, and J. Á. Velázquez-Iturbide. Visualization of syntax trees for language processing courses. J.UCS: Journal of Universal Computer Science, 15(7):1546– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
7/visualization_of_syntax_
trees.

Atanasiu:2001:CLP

[AMVM01] A. Atanasiu, C. Martin-Vide, and A. Mateescu. Codifiable languages and the Parikh matrix mapping. J.UCS: Journal of Universal Computer Science, 7(9):783– 793, September 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_7_9/codifiable_
languages_and_the.

Arellano:2014:LML

[AMYH14] D. Arellano, C. Manresa-Yee, and V. Helzle. Let me listen to poetry, let me see emotions. J.UCS: Journal of Universal Computer Science, 20(7):1006– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
7/let_me_listen_to.

Andreeva:1996:IIS

[And96]  P. T. Andreeva.  In-

exact information systems and its application to approximate reasoning. J.UCS: Journal of Universal Computer Science, 2(2):70–76, February 18, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/inexact_
information_systems_
and_its_application_
to_approximate_reasoning.

Andersen:1997:LTS

[And97] N. Andersen. Linear time simulation of invertible non-deterministic stack algorithms. J.UCS: Journal of Universal Computer Science, 3(3):148– 171, March 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/

                                        jucs_3_3/linear_time_
                                        simulation.

Araujo:2008:QIE

[ANdMM08] M. P. M. Araujo, N. Nedjah, and L. de Macedo Mourelle. Quantum-inspired evolutionary state assignment for synchronous finite state machines. J.UCS: Journal of Universal Computer Science, 14 (15):2532–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_15/quantum_inspired_
                                        evolutionary_state.

Anguelov:1998:SFA

[Ang98]  R. Anguelov.  Spline-

Fourier approximations of discontinuous waves. J.UCS: Journal of Universal Computer Science, 4(2):110–113, February 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_2/spline_
                                        fourier_approximations_
                                        of.

Anonymous:1999:SD

[Ano99] Anonymous. System description. J.UCS: Journal of Universal Computer Science, 5(3):??, March 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
3/system_description.

Anonymous:2007:ICC

[Ano07]  Anonymous.  IPCity:

Competitive call for an additional project partner. J.UCS: Journal of Universal Computer Science, 13(1):??, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
ujs/jucs/info/announcements_
and_advertisements/jucs_
13_01/ipcity_competitive_
call.pdf.

Alexander:2004:CBT

[AO04] R. T. Alexander and J. Offutt. Couplingbased testing of O-O programs. J.UCS: Journal of Universal Computer Science, 10(4):391– 427, April 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
10_4/coupling_based_
testing_of.

Atzmueller:2005:SAV

[AP05] M. Atzmueller and F. Puppe.

Semi-automatic visual subgroup mining using VIKAMINE. J.UCS: Journal of Universal Computer Science, 11 (11):1752–1765, ???? 2005.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        11_11/semi_automatic_
                                        visual_subgroup.

Alomair:2009:ITS

[AP09] B. Alomair and R. Poovendran. Information theoretically secure encryption with almost free authentication. J.UCS: Journal of Universal Computer Science, 15(15):2937–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15_
                                        15/information_theoretically_
                                        secure_encryption.

Alves:2008:CMD

[APDC08] M. Pitanga Alves, P. F. Pires, F. C. Delicato, and M. L. M. Campos. Cross-MDA: a model-driven approach for aspect management. J.UCS: Journal of Universal Computer Science, 14(8):1314– 1343, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_8/crossmda_a_
                                        model_driven.

Ahmed:2009:PUR

[APJK09] S. Ahmed, S. Park, J. J. Jung, and S. Kang. A personalized URL re-ranking method using psychological user browsing characteristics. J.UCS: Journal of Universal Computer Science, 15(4):926– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
4/a_personalized_url_
reranking.

Abello:2004:BFS

[APM04] J. Abello, A. J. Pogel, and L. Miller. Breadth first search graph partitions and concept lattices. J.UCS: Journal of Universal Computer Science, 10(8):934–954, August 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
8/breadth_first_search_
graph.

Ali:2012:TLS

[APNA12] M. Ali, M. Pant, A. K.

Nagar, and C. W. Ahn. Two local search strategies for differential evolution. J.UCS: Journal of Universal Computer Science, 18(13):1853–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/

                                        jucs_18_13/two_local_
                                        search_strategies.

Amaro:2005:CBD

[APR05] S. Amaro, E. Pimentel, and A. M. Roldán. Coordinating behavioral descriptions of components. J.UCS: Journal of Universal Computer Science, 11(10):1676–1694, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_11_
                                        10/coordinating_behavioral_
                                        descriptions_of.

Aranha:1995:EDA

[AR95] R. F. M. Aranha and C. Pandu Rangan. An efficient distributed algorithm for st-numbering the vertices of a biconnected graph. J.UCS: Journal of Universal Computer Science, 1(9): 633–651, September 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        an_efficient_distributed_
                                        algorithm_for_st_numbering_
                                        the_vertices_of_a_biconnected_
                                        graph.

Armando:2001:PEM

[AR01] A. Armando and S. Ranise. A practical extension mechanism for decision procedures: the case study of universal Presburger arithmetic. J.UCS: Journal of Universal Computer Science, 7(2): 124–140, February 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_7_2/a_practical_
extension_mechanism.

Alfirevic:2004:KIS

[AR04] N. Alfirevic and D. Racic.

Knowledge integration as a source of competitive advantage in large Croatian enterprises. J.UCS: Journal of Universal Computer Science, 10(6): 712–722, June 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_10_6/knowledge_
integration_as_a.

Araujo:1997:CPD

[Ara97] L. Araujo. Correctness proof of a distributed implementation of Prolog by means of abstract state machines. J.UCS: Journal of Universal Computer Science, 3(5):568– 602, May 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs_3_5/correctness_
proof;   http://www.

                                        jucs.org/jucs_3_5/correctness_
                                        proof; internal&sk=05460486.

Araki:2003:AVM

[Ara03] J. Araki. Action vectors: Modeling spatial relations between objects and routes. J.UCS: Journal of Universal Computer Science, 9(9):1046– 1072, September 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_9_9/action_
                                        vectors_modeling_spatial.

Aredo:2002:FSU

[Are02] D. B. Aredo. A framework for semantics of UML sequence diagrams in PVS. J.UCS: Journal of Universal Computer Science, 8(7):674–697, July 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_8_7/a_framework_
                                        for_semantics.

Aguilar-Ruiz:2005:VDM

[ARFT05] J. S. Aguilar-Ruiz and F. J. Ferrer-Troyano. Visual data mining. J.UCS: Journal of Universal Computer Science, 11 (11):1749–1751, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_

11;   http://www.jucs.
org/jucs_11_11#;   http:
//www.jucs.org/jucs_
11_11/visual_data_mining.

Aguilar-Ruiz:2005:DS

[ARG05] J. S. Aguilar-Ruiz and J. Gama. Data streams. J.UCS: Journal of Universal Computer Science, 11(8):1349–1352, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11;
http://www.jucs.org/
jucs_11_8#;   http://
www.jucs.org/jucs_11_
8/data_streams.

Allert:2004:SMM

[ARN04] H. Allert, C. Richter, and W. Nejdl. Situated models and metadata for learning management. J.UCS: Journal of Universal Computer Science, 10(1):4–13, January 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
10_1/situated_models_
and_metadata.

Ajmal:2014:CVM

[ARQH14] S. Ajmal, A. Rasheed, A. Qayyum, and A. Hasan. Classification of VANET MAC, routing and approaches a detailed survey. J.UCS: Journal of Universal Computer Science, 20(4):462–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
4/classification_of_
vanet_mac.

Arroyo:2007:OGS

[Arr07]  S. Arroyo.  Ontology

and grammar of the SO-PHIE choreography conceptual framework — an ontological model for knowledge management. J.UCS: Journal of Universal Computer Science, 13(9):1157–1183, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
9/ontology_and_grammar_
of.

Andrade:2014:FFB

[ARRB14] R. Andrade, H. Rebelo, M. Ribeiro, and P. Borba. Flexible feature binding with AspectJ-based idioms. J.UCS: Journal of Universal Computer Science, 20(5):692– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
5/flexible_feature_binding_
with.

Arslanov:1997:DSR

[Ars97]  A. Arslanov.  Differ-

ence splittings of recursively enumerable sets. J.UCS: Journal of Universal Computer Science, 3(3):172–184, March 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_3_3/difference_
splittings_of_recursively.

Antoniou:2000:ISC

[ARS00] I. Antoniou, M. Reeve, and V. Stenning. The information society as a complex system. J.UCS: Journal of Universal Computer Science, 6(3): 272–288, March 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
6_3/the_information_
society_as.

Aqqal:2008:UTS

[ARS^{+}08] A. Aqqal, C. Rensing, R. Steinmetz, N. Elkamoun, and A. Berraissoul. Using taxonomies to support the macro design process for the production of Web based trainings. J.UCS: Journal of Universal Computer Science, 14(10):1763–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
10/using_taxonomies_
to_support.

Agudo:2016:DAA

[ARS16] J. E. Agudo, M. Rico, and H. Sánchez. Design and assessment of adaptive hypermedia games for English acquisition in preschool. J.UCS: Journal of Universal Computer Science, 22(2):161– ??, ???? 2016. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_22_2/design_and_
assessment_of_adaptive.

Arbab:2007:AFM

[AS07] F. Arbab and M. Sirjani. Applications of formal methods to system design and verification. J.UCS: Journal of Universal Computer Science, 13(13):1970–1971, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_13;   http:
//www.jucs.org/jucs_
13_13#;   http://www.
jucs.org/jucs_13_13/
applications_of_formal_
methods.

Alwagait:2014:WTB

[AS14] E. Alwagait and B. Shahzad.

When are tweets better valued? an empirical study. J.UCS: Journal of Universal Computer Science, 20(10):1511–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_20_10/when_are_
tweets_better.

Al-Smadi:2016:ABS

[ASAAASJ16] M. Al-Smadi, M. Al-Ayyoub, H. Al-Sarhan, and Y. Jararweh. An aspect-based sentiment analysis approach to evaluating Arabic news affect on readers. J.UCS: Journal of Universal Computer Science, 22(5):630– ??, ???? 2016. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
22_5/an_aspect_based_
sentiment.

Arrieta-Salinas:2014:EVC

[ASAIN14] I. Arrieta-Salinas, J. E.

Armendáriz-Iñigo, and J. Navarro. Epidemia: Variable consistency for transactional cloud databases. J.UCS: Journal of Universal Computer Science, 20(14):1876–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_20_14/epidemia_

                                        variable_consistency_
                                        for.

Abdoli:2011:RCF

[ASH11] A. Abdoli, J. Shahrabi, and J. Heidary. Representing a composing fuzzy-DEA model to measure knowledge workers productivity based upon their efficiency and cost effectiveness. J.UCS: Journal of Universal Computer Science, 17 (10):1390–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        17_10/representing_a_
                                        composing_fuzzy.

Amado-Salvatierra:2016:TSD

[ASHT^{+}16] H. R. Amado-Salvatierra, J. R. Hilera, S. O. Tortosa, R. Hernández Rizzardini, and N. Piedra. Towards a semantic definition of a framework to implement accessible e-learning projects. J.UCS: Journal of Universal Computer Science, 22(7):921–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_22_7/towards_a_
                                        semantic_definition.

Al-Shamaileh:2013:WIR

[ASS13] O. Al-Shamaileh and A. Sutcliffe. Website interactivity and repeated exposure, what influences user experience? J.UCS: Journal of Universal Computer Science, 19(8):1123– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
8/website_interactivity_
and_repeated.

Adamopoulou:2007:WSS

[ASTL07] P. Adamopoulou, E. Sakkopoulos, A. Tsakalidis, and M. Lytras. Web service selection based on QoS knowledge management. J.UCS: Journal of Universal Computer Science, 13(9):1138–1156, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
9/web_service_selection_
based.

Avrithis:2003:UAH

[ASW^{+}03] Y. Avrithis, G. Stamou, M. Wallace, F. Marques, P. Salembier, X. Giro, W. Haas, H. Vallant, and M. Zufferey. Unified access to heterogeneous audiovisual archives. J.UCS: Journal of Universal Computer Science, 9(6): 510–519, June 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        6/unified_access_to_
                                        heterogeneous.

Aoto:1997:PC

[AT97] T. Aoto and Y. Toyama. Persistency of confluence. J.UCS: Journal of Universal Computer Science, 3(11):1134–1147, November 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_3_11/
                                        persistency_of_confluence.

Arita:2007:RTH

[AT07] D. Arita and R.-I. Taniguchi. Real-time human proxy: An avatar-based communication system. J.UCS: Journal of Universal Computer Science, 13(2): 161–176, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        13_2/real_time_human_
                                        proxy.

Albert:2010:SRT

[AT10a] J. Albert and G. Tischler. On succinct representations of textured surfaces by weighted finite automata. J.UCS: Journal of Universal Computer Science, 16(5):586– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
5/on_succinct_representations_
of.

Ali:2010:GFM

[AT10b] A. Ali and K. Terada.

A general framework for multi-human tracking using Kalman filter and fast mean shift algorithms. J.UCS: Journal of Universal Computer Science, 16(6):921–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_6/a_general_
framework_for.

Agnol:2013:MCA

[AT13] J. M. Hauagge Dall Agnol and C. A. Tacla. A method for collaborative argumentation in merging individual ontologies. J.UCS: Journal of Universal Computer Science, 19(12):1808–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_12/a_method_
for_collaborative.

Aguilar:2009:SST

[ATGP09] D. A. G. Aguilar, R. Therón, and F. J. García-Peñalvo. Semantic spiral timelines used as support for elearning. J.UCS: Journal of Universal Computer Science, 15(7):1526– ??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_7/semantic_spiral_
timelines_used.

Al-Twaijry:1998:SPB

[ATOFF98] H. A. Al-Twaijry, S. F. Oberman, S. T. Fu, and M. J. Flynn. The SNAP project: Building validated floating point. J.UCS: Journal of Universal Computer Science, 4(2):99–109, February 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_2/the_
snap_project_building.

Abraham:2005:IAS

[ATSJ05] A. Abraham, J. Thomas, S. Sanyal, and L. Jain. Information assurance and security. J.UCS: Journal of Universal Computer Science, 11(1):1–3, January 28, 2005. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
 http://www.jucs.org/
jucs_11_1/information_
assurance_and_security.

Aid:1998:GEV

[ATV98] R. Aïd, L. Testard, and G. Villard. Global error visualization. J.UCS: Journal of Universal Computer Science, 4(2): 90–98, February 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_2/global_
error_visualization.

Augusto:2004:SIW

[AUN04] J. C. Augusto and U. Ultes-Nitsche. Second international workshop on verification and validation of enterprise information systems. J.UCS: Journal of Universal Computer Science, 10(11):1495– 1497, November 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
 http://www.jucs.org/
jucs_10_11/second_international_
workshop_on.

Aurenhammer:2001:CGS

[Aur01] F. Aurenhammer. Computational geometry — some easy questions and their recent solutions. J.UCS: Journal of Universal Computer Science, 7(5):338–354, May 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_7_
                                        5/computational_geometry_
                                        some_easy.

Akritas:2007:CVM

[AV07] A. G. Akritas and P. S. Vigklas. A comparison of various methods for computing bounds for positive roots of polynomials. J.UCS: Journal of Universal Computer Science, 13(4):455–467, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        13_4/a_comparison_of_
                                        various.

Arrue:2008:SDA

[AVA08] M. Arrue, M. Vigo, and J. Abascal. Supporting the development of accessible Web applications. J.UCS: Journal of Universal Computer Science, 14(16):2699–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_16/supporting_the_
                                        development_of.

Abrahamian:2004:EPA

[AWGS04] E. Abrahamian, J. Weinberg, M. Grady, and C. M. Stanton. The effect of personalityaware computer-human interfaces on learning. J.UCS: Journal of Universal Computer Science, 10(1):27–37, January 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_10_1/the_effect_
of_personality.

Allain:2000:MIS

[AY00] L. Allain and P. Yim.

Modeling information system behavior with dynamic relations nets. J.UCS: Journal of Universal Computer Science, 6(11):1109–1130, November 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
11/modeling_information_
system_behavior.

Avouris:2012:RML

[AY12] N. Avouris and N. Yiannoutsou. A review of mobile location-based games for learning across physical and virtual spaces. J.UCS: Journal of Universal Computer Science, 18(15):2120–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_15/a_review_
of_mobile.

Al-Zahrani:2015:PBF

[AZMA15] A. M. Al-Zahrani, H. Mathkour, and H. Abdalla. PSO-based feature selection for Arabic text summarization. J.UCS: Journal of Universal Computer Science, 21(11): 1454–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_21_11/pso_based_
                                        feature_selection.

Azzedin:2010:CTF

[Azz10] F. A. Azzedin. Classifying and tracking free riders in multimedia-based systems. J.UCS: Journal of Universal Computer Science, 16(10):1368–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_16_
                                        10/classifying_and_tracking_
                                        free.

Baier:2005:PMR

[Bai05] C. Baier. Probabilistic models for Reo connector circuits. J.UCS: Journal of Universal Computer Science, 11(10): 1718–1748, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_

11_10/probabilistic_
models_for_reo.

Baicher:2012:RTI

[Bai12] G. Singh Baicher. Realtime implementation of a class of optimised multirate quadrature mirror filter bank using genetic algorithms. J.UCS: Journal of Universal Computer Science, 18(13): 1871–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_13/real_time_
implementation_of.

Boudjit:2007:DAD

[BAML07] S. Boudjit, C. Adjih, P. Mühlethaler, and A. Laouiti. Duplicate address detection and autoconfiguration in OLSR. J.UCS: Journal of Universal Computer Science, 13(1):4–31, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_13_1/duplicate_
address_detection_and.

Banach:1996:MFI

[Ban96] R. Banach. MONSTR I — fundamental issues and the design of MON-STR. J.UCS: Journal of Universal Computer Science, 2(4):164–216, April 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_2_4/monstr_
                                        i_fundamental_issues.

Banach:1997:MIS

[Ban97a]  R. Banach.  MON-

STR II — suspending semantics and independence. J.UCS: Journal of Universal Computer Science, 3(7):756– 802, July 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs_3_7/monstr_
                                        ii_suspending_semantics;
                                         internal&sk=05460486.

Banach:1997:MVT

[Ban97b] R. Banach. MONSTR V — transitive coercing semantics and the Church–Rosser property. J.UCS: Journal of Universal Computer Science, 3(12):1283–1336, December 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_3_12/
                                        monstr_v_transitive_
                                        coercing.

Banach:2007:FMG

[Ban07]  R. Banach.  Formal

methods guest editorial.

J.UCS: Journal of Universal Computer Science, 13(5):593–601, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_5/formal_methods.

Bozo:2016:MRP

[BAP^{+}16] J. Bozo, R. Alarcon, M. Peralta, T. Mery, and V. Cabezas. Metadata for recommending primary and secondary level learning resources. J.UCS: Journal of Universal Computer Science, 22(2):197– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_22_
2/metadata_for_recommending_
primary.

Banach:2003:MHA

[BAPG03] Richard Banach, Farhab Arbab, George A. Papadopoulos, and John R. W. Glauert. A multiply hierarchical automaton semantics for the IWIM coordination model. J.UCS: Journal of Universal Computer Science, 9(1):2– 33, January 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/

                                        jucs_9_1/a_multiply_
                                        hierarchical_automaton.

Barbosa:2003:TCS

[Bar03]  L. S. Barbosa.  To-

wards a calculus of statebased software components. J.UCS: Journal of Universal Computer Science, 9(8):891–909, August 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        8/towards_a_calculus_
                                        of.

Baroni:2005:CS

[Bar05] M. A. Baroni. Constructive suprema. J.UCS: Journal of Universal Computer Science, 11 (12):1865–1877, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        11_12/constructive_suprema.

Bravo:2006:UCA

[BAR06] J. Bravo, X. Alamán, and T. Riesgo. Ubiquitous computing and ambient intelligence: New challenges for computing. J.UCS: Journal of Universal Computer Science, 12(3):233–235, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12;

http://www.jucs.org/
jucs_12_3#;   http://
www.jucs.org/jucs_12_
3/ubiquitous_computing_
and_ambient.

Bolanos:2012:MSH

[BARB12] F. Bolanos, J. E. Aedo, F. Rivera, and N. Bagherzadeh. Mapping and scheduling in heterogeneous NoC through populationbased incremental learning. J.UCS: Journal of Universal Computer Science, 18(7):901–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
7/mapping_and_scheduling_
in.

Bjorn-Andersen:2009:WAD

[BARR09] N. Bjørn-Andersen, L. Bloch Rasmussen, and S. Rasmussen. Web 2.0 adoption by Danish newspapers — urgent need for new business models? J.UCS: Journal of Universal Computer Science, 15(3):692–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_3/web_20_adoption_
by.

Baloian:2014:DDC

[BAZ14] N. Baloian, D. Aguirre, and G. Zurita. Developing distributed collaborative applications with HTML5 under the coupled objects paradigm. J.UCS: Journal of Universal Computer Science, 20(13): 1712–??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
20_13/developing_distributed_
collaborative_applications.

Barrio:1998:PSR

[BB98] R. Barrio and J. C. Berges. Perturbation simulations of rounding errors in the evaluation of Chebyshev series. J.UCS: Journal of Universal Computer Science, 4(6):561– 573, June 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_6/perturbation_
simulations_of_rounding;
internal&sk=0C220489.

Barbosa:2004:RMC

[BB04] M. A. Barbosa and L. S. Barbosa. A relational model for component interconnection. J.UCS: Journal of Universal Computer Science, 10(7):808–823, July 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
7/a_relational_model_
for.

Badica:2008:FAB

[BB08a] A. Badica and C. Badica. Formalizing agentbased English auctions using finite state process algebra. J.UCS: Journal of Universal Computer Science, 14(7):1118– 1135, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
7/formalizing_agent_
based_english.

Batory:2008:MTS

[BB08b] D. Batory and E. Börger.

Modularizing theorems for software product lines: The Jbook case study. J.UCS: Journal of Universal Computer Science, 14(12):2059–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
12/modularizing_theorems_
for_software.

Berger:2009:RSC

[BB09] J. Berger and D. S.

Bridges. Rearranging series constructively. J.UCS: Journal of Universal Computer Science, 15(17):3160–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_15_17/rearranging_
                                        series_constructively.

Bauer:2010:CES

[BB10a] A. Bauer and J. Blanck. Canonical effective subalgebras of classical algebras as constructive metric completions. J.UCS: Journal of Universal Computer Science, 16 (18):2496–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_16_18/canonical_
                                        effective_subalgebras_
                                        of.

Berger:2010:CSL

[BB10b] J. Berger and D. S. Bridges. A constructive study of Landau’s summability theorem. J.UCS: Journal of Universal Computer Science, 16 (18):2523–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        16_18/a_constructive_
                                        study_of.

Bonifacio:2002:KNB

[BBC02] M. Bonifacio, M. Bouquet, and R. Cuel. Knowledge nodes: the building blocks of a distributed approach to knowledge management. J.UCS: Journal of Universal Computer Science, 8(6):652– 661, June 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
8_6/knowledge_nodes_
the_building.

Barbosa:2014:DIU

[BBdOR14] J. L. V. Barbosa, D. N. Ferrari Barbosa, J. Machado de Oliveira, and S. Andrade Rabello, Jr. A decentralized infrastructure for ubiquitous learning environments. J.UCS: Journal of Universal Computer Science, 20(12): 1649–??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
20_12/a_decentralized_
infrastructure_for.

Beneventano:2007:SNM

[BBGV07] D. Beneventano, S. Bergamaschi, F. Guerra, and M. Vincini. The SE-WASIE network of mediator agents for semantic search. J.UCS: Journal of Universal Computer Science, 13(12):1936–1969, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_

                                        12/the_sewasie_network_
                                        of.

Bernardos:2013:MME

[BBIC13] A. M. Bernardos, L. Bergesio, J. Iglesias, and J. R. Casar. MECCANO: a mobile-enabled configuration framework to coordinate and augment networks of smart objects. J.UCS: Journal of Universal Computer Science, 19(17):2503–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_19_17/meccano_a_
                                        moblie_enabled.

Benavent:2013:ADH

[BBL13] A. Peñalver Benavent, F. Botella Beviá, and J. A. Gallud Lázaro. Advances in the development of highly interactive systems. J.UCS: Journal of Universal Computer Science, 19(8):1023– ??, ???? 2013. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        19_8/advances_in_the_
                                        development.

Botturi:2012:BLG

[BBM12] L. Botturi, C. Bramani, and S. McCusker. Boys are like girls: Insights in the gender digital divide in higher education in Switzerland and Europe. J.UCS: Journal of Universal Computer Science, 18(3):353– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
3/boys_are_like_girls.

Banach:1995:TPC

[BBP95] R. Banach, J. Balazs, and G. Papadopoulos. A translation of the picalculus into MONSTR. J.UCS: Journal of Universal Computer Science, 1(6):339–398, June 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/A_Translation_of_
the_Pi-Calculus_Into_
MONSTR.

Brzykcy:2008:SMA

[BBP08] G. Brzykcy, J. Bartoszek, and T. Pankowski. Schema mappings and agents’ actions in P2P data integration system. J.UCS: Journal of Universal Computer Science, 14(7):1048–1060, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_14_7/schema_
mappings_and_agents.

Bertoni:2011:LWN

[BC11] M. Bertoni and K. Chirumalla. Leveraging Web 2.0 in new product development: Lessons learned from a cross-company study. J.UCS: Journal of Universal Computer Science, 17(4):548–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_17_4/leveraging_
                                        web_20_in.

Bjorklund:2016:TMA

[BC16] J. Björklund and L. Cleophas. A taxonomy of minimisation algorithms for deterministic tree automata. J.UCS: Journal of Universal Computer Science, 22(2):180–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_22_2/a_taxonomy_
                                        of_minimisation.

Baladron:2010:MLC

[BCA^{+}10] C. Baladrón, A. Cadenas, J. Aguiar, B. Carro, and A. Sánchez-Esguevillas. Multi-level context management and inference framework for smart telecommunication services. J.UCS: Journal of Universal Computer Science, 16(15):1973–??, ???? 2010. CODEN ????

ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16_15/multi_
level_context_management.

Biondi:2006:MOE

[BCC^{+}06] T. Biondi, A. Ciccazzo, V. Cutello, S. D’Antona, G. Nicosia, and S. Spinella. Multi-objective evolutionary algorithms and pattern search methods for circuit design problems. J.UCS: Journal of Universal Computer Science, 12(4):432–449, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_4/multi_objective_
evolutionary_algorithm.

Bonnel:2011:FEI

[BCCH11] N. Bonnel, M. Chevalier, C. Chrisment, and G. Hubert. A framework to evaluate interface suitability for a given scenario of textual information retrieval. J.UCS: Journal of Universal Computer Science, 17(6):831–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_6/a_framework_to_
evaluate.

Bon:2013:SMB

[BCD13] P. Bon and S. Collart-Dutilleul. From a solution model to a B model for verification of safety properties. J.UCS: Journal of Universal Computer Science, 19(1):2– ??, ???? 2013. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_1/from_a_solution_
model_to.

Bridges:1997:LCS

[BCDK97] D. S. Bridges, C. S. Calude, M. J. Dinneen, and B. Khoussainov. Logic in computer science. J.UCS: Journal of Universal Computer Science, 3(11):??, November 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs;   http://
medoc.springer.de:8000/
jucs/jucs_3_11/logic_
in_computer_science;
 internal&sk=0156049D/
jucs_3_11.

Ba:2005:PNL

[BCFM05] C. Ba, M. A. Carrero, M. Halfeld Ferrari, and M. A. Musicante. PEWS: a new language for building Web service interfaces. J.UCS: Journal of Universal Computer Science, 11(7):1215–1233, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
7/pews_a_new_language.

Berenci:1998:IEW

[BCG98] E. Berenci, C. Carpineto, and V. Giannini. Improving the effectiveness of WEB search engines using selectable views of retrieval results. J.UCS: Journal of Universal Computer Science, 4(9):737– 747, September 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_9/improving_
the_effectiveness_of.

Balanescu:1999:CSX

[BCG^{+}99] T. Balanescu, A. J.

Cowling, H. Georgescu, M. Gheorghe, M. Holcombe, and C. Vertan. Communicating stream X-machines systems are no more than X-machines. J.UCS: Journal of Universal Computer Science, 5(9):494–507, September 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
9/communicating_stream_
x_machines.

Bevilacqua:2009:CBG

[BCG^{+}09] V. Bevilacqua, L. Cariello, M. Giannini, G. Mastronardi, V. Santarcangelo, R. Scaramuzzi, and A. Troccoli. A comparison between a geometrical and an ANN based method for retinal bifurcation points extraction. J.UCS: Journal of Universal Computer Science, 15(13):2608–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
13/a_comparison_between_
a.

Brogi:2014:DVM

[BCG^{+}14] A. Brogi, J. Cubo, L. González, E. Pimentel, and R. Ruggia. Dynamic verification of mashups of service-oriented things through a mediation platform. J.UCS: Journal of Universal Computer Science, 20(8):1049–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
8/dynamic_verification_
of_mashups.

Bajuelos:2012:HMS

[BCHM12] A. L. Bajuelos, S. Canales, G. Hernández, and M. Martins. A hybrid metaheuristic strategy for covering with wireless devices. J.UCS: Journal of Universal Computer Science, 18(14):1906–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_14/a_hybrid_
metaheuristic_strategy.

Blanco:2008:SSD

[BCM08] L. Blanco, V. Crescenzi, and P. Merialdo. Structure and semantics of data-intensive Web pages: An experimental study on their relationships. J.UCS: Journal of Universal Computer Science, 14(11):1877–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_11/structure_
and_semantics_of.

Bhattacharya:2012:AAB

[BCM12] B. Sen Bhattacharya, D. Coyle, and L. P. Maguire. Assessing alpha band event-related synchronisation/desynchronisation using a bio-inspired computational model. J.UCS: Journal of Universal Computer Science, 18 (13):1888–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_13/assessing_
alpha_band_event.

Bravetti:2007:TBM

[BCNR07] M. Bravetti, A. Casalboni, M. Núñez, and I. Rodriguez. From theoretical e-barter models to two alternative implementations based on Web sevices. J.UCS: Journal of Universal Computer Science, 13(13):2035–2075, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_
                                        13/from_theoretical_
                                        ebarter_models.

Brattka:2009:CCA

[BCR09] V. Brattka, P. Collins, and R. Rettinger. Computability and complexity in analysis. J.UCS: Journal of Universal Computer Science, 15(6):1143– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15;
                                         http://www.jucs.org/
                                        jucs_15_06/computability_
                                        and_complexity_in;
                                        http://www.jucs.org/
                                        jucs_15_6#.

Bucki:2015:HAM

[BCS15] R. Bucki, B. Chramcov, and P. Suchánek. Heuristic algorithms for manufacturing and replacement strategies of the production system.

J.UCS: Journal of Universal Computer Science, 21(4):503–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_4/heuristic_
algorithms_for_manufacturing.

Bonifacio:2004:MKT

[BCZ04] M. Bonifacio, P. Camussone, and C. Zini. Managing the KM tradeoff: Knowledge centralization versus distribution. J.UCS: Journal of Universal Computer Science, 10(3):162–175, March 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
10_3/managing_the_km_
trade.

Beyls:2000:CGM

[BD00] K. E. Beyls and E. H.

D’Hollander. Compiler generated multithreading to alleviate memory latency. J.UCS: Journal of Universal Computer Science, 6(10):968–993, October 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
10/compiler_generated_
multithreading_to.

Brattka:2005:CSS

[BD05] V. Brattka and R. Dillhage. Computability of the spectrum of self-adjoint operators. J.UCS: Journal of Universal Computer Science, 11(12):1884–1900, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_11_
                                        12/computability_of_
                                        the_spectrum.

Bavarian:2006:CBM

[BD06] M. Bavarian and V. Dahl. Constraint based methods for biological sequence analysis. J.UCS: Journal of Universal Computer Science, 12(11):1500– 1520, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12_
                                        11/constraint_based_
                                        methods_for.

Blanco:2014:SBA

[BdGFMT14] C. Blanco, I. García-Rodríguez de Guzmán, E. Fernández-Medina, and J. Trujillo. Showing the benefits of applying a model driven architecture for developing secure OLAP applications. J.UCS: Journal of Universal Computer Science, 20(2):79– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
2/showing_the_benefits_
of.

Balcazar:1996:OPA

[BDGW96] José L. Balcázar, Josep Díaz, Ricard Gavaldà, and Osamu Watanabe. An optimal parallel algorithm for learning DFA. J.UCS: Journal of Universal Computer Science, 2(3):97–112, March 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
an_optimal_parallel_
algorithm_for_learning_
dfa.

Bhattacharyya:2009:VPA

[BDhKB09] D. Bhattacharyya, P. Das, T. h. Kim, and S. K. Bandyopadhyay. Vascular pattern analysis towards pervasive palm vein authentication. J.UCS: Journal of Universal Computer Science, 15(5): 1081–??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_5/vascular_pattern_
analysis_towards.

Bravo:2010:AIV

[BdI10] J. Bravo and D. López de Ipiña. Ambient intelligence vision: a perspective. J.UCS: Journal of Universal Computer Science, 16(12):1478–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16;   http:
//www.jucs.org/jucs_
16_12/ambient_intelligence_
vision_a;   http://www.
jucs.org/jucsrssfeed-
issue.

Blankertz:2006:BBC

[BDL^{+}06] B. Blankertz, G. Dornhege, S. Lemm, M. Krauledat, G. Curio, and K.- R. Müller. The Berlin brain-computer interface: Machine learning based detection of user specific brain states. J.UCS: Journal of Universal Computer Science, 12(6):581– 607, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_12_6/the_berlin_
brain_computer.

Bigot:2015:LCB

[BDM15] A. Bigot, S. Déjean, and J. Mothe. Learning to choose the best system configuration in information retrieval: the case of repeated queries. J.UCS: Journal of Universal Computer Science, 21(13):1726–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_13/learning_
to_choose_the.

Bergel:2005:AMD

[BDN05] A. Bergel, S. Ducasse, and O. Nierstrasz. Analyzing module diversity. J.UCS: Journal of Universal Computer Science, 11(10):1613–1644, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_11_10/analyzing_
module_diversity.

Baraani-Dastjerdi:1997:UCH

[BDPSNG97] A. Baraani-Dastjerdi, J. Pieprzyk, R. Safavi-Naini, and J. R. Getta. Using cryptographic hash functions for discretionary access control in object-oriented databases. J.UCS: Journal of Universal Computer Science, 3(6):730– 753, June 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs_3_6/using_
cryptographic_functions_
for; internal&sk=05460486.

Benitti:2013:ESA

[BdS13] F. Barreto Vavassori Benitti and R. C. da Silva.

Evaluation of a systematic approach to requirements reuse. J.UCS: Journal of Universal Computer Science, 19(2):254–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_2/evaluation_
of_a_systematic.

Bigonha:2006:SPS

[BdVG06] M. A. Silva Bigonha and A. de Vasconcellos Garcia. Selected papers from SBLP 2006: The 10th Brazilian Symposium on Programming Languages. J.UCS: Journal of Universal Computer Science, 12(7):762–763, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_12;   http:
//www.jucs.org/jucs_
12_7#; http://www.jucs.
org/jucs_12_7/sblp_2006.

Brusilovsky:1998:SUM

[BE98] P. Brusilovsky and J. Eklund. A study of user model based link annotation in educational hypermedia. J.UCS: Journal of Universal Computer Science, 4(4):429– 448, April 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_4/a_
study_of_user;   internal&
sk=0C220489.

Bischof:2011:CCK

[BE11] N. Bischof and M. J. Eppler. Caring for clarity in knowledge communication. J.UCS: Journal of Universal Computer Science, 17(10):1455–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
10/caring_for_clarity_
in.

Beck:2003:SCM

[Bec03] S. Beck. Skill and competence management as a base of an Integrated Personnel Development (IPD) — a pilot project in the Putzmeister, Inc./ Germany. J.UCS: Journal of Universal Computer Science, 9(12):1381– 1387, December 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_9_12/skill_and_
competence_management.

Boehm:2005:MIP

[BEH^{+}05] K. Boehm, W. Engelbach, J. Haertwig, M. Wilcken, and M. Delp. Modelling and implementing prebuilt information spaces. architecture and methods for process oriented knowledge management. J.UCS: Journal of Universal Computer Science, 11(4):605–633, April 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_11_4/modelling_
                                        and_implementing_pre.

Barth:2008:OTP

[BEH08] D. Barth, L. Echabbi, and C. Hamlaoui. Optimal transit price negotiation: The distributed learning perspective. J.UCS: Journal of Universal Computer Science, 14(5):745– 765, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_5/optimal_transit_
                                        price_negotiation.

Bella:2008:WCS

[Bel08] G. Bella. What is correctness of security protocols? J.UCS: Journal of Universal Computer Science, 14(12):2083–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_14_12/what_
                                        is_correctness_of.

Bry:2014:IMC

[BEPT14] F. Bry, M. Ebner, A. Pohl, and B. Taraghi. Interaction in massive courses. J.UCS: Journal of Universal Computer Science, 20(1):1–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_20;   http:
//www.jucs.org/jucs_
20_1/interaction_in_
massive_courses.

Berger:2005:CEU

[Ber05]  J. Berger.  Construc-

tive equivalents of the uniform continuity theorem. J.UCS: Journal of Universal Computer Science, 11(12):1878–1883, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
12/constructive_equivalents_
of_the.

Berregeb:2006:PPB

[Ber06]  N. Berregeb.  Prov-

ing properties for behavioural specifications with term observation. J.UCS: Journal of Universal Computer Science, 12(10):1413–1425, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12_
                                        10/proving_properties_
                                        for_behavioural.

Berger:2010:RIC

[Ber10] U. Berger. Realisability for induction and coinduction with applications to constructive analysis. J.UCS: Journal of Universal Computer Science, 16(18):2535–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_16_
                                        18/realisability_for_
                                        induction_and.

Baar:1999:IDT

[BFF99] T. Baar, B. Fischer, and D. Fuchs. Integrating deduction techniques in a software reuse application. J.UCS: Journal of Universal Computer Science, 5(3):52– 72, March 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_5_3/integrating_
                                        deduction_techniques_
                                        in.

Brocco:2011:OCS

[BFF11] M. Brocco, F. Forster, and M. R. Frieß. 360^{∘} open creativity support. J.UCS: Journal of Universal Computer Science, 17(12):1673–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_12/360_degree_
open_creativity.

Bravo:2005:CER

[BFMSP05] J. Bravo, B. Fernández-Manjón, and J. M. Sánchez-Pérez. Computers and education: Research and experiences in eLearning technology. J.UCS: Journal of Universal Computer Science, 11(9):1454–1457, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11;
 http://www.jucs.org/
jucs_11_9#;   http://
www.jucs.org/jucs_11_
9/computers_and_education_
research.

Butler:2005:PMC

[BFN05] M. Butler, C. Ferreira, and M. Y. Ng. Precise modelling of compensating business transactions and its application to BPEL. J.UCS: Journal of Universal Computer Science, 11(5):712–743, May 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_

                                        5/precise_modelling_
                                        of_compensating.

Blass:1997:LTH

[BG97] A. Blass and Y. Gurevich. The linear time hierarchy theorems for abstract state machines and RAMs. J.UCS: Journal of Universal Computer Science, 3(4):247– 278, April 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs_3_4/linear_
                                        time_hierarchy;   http:
                                        //www.jucs.org/jucs_
                                        3_4/linear_time_hierarchy;
                                         internal&sk=05460486.

Buckner:1998:CEP

[BG98] K. Buckner and M. Gillham. A comparative evaluation of print and electronic reviews of multimedia information products. J.UCS: Journal of Universal Computer Science, 4(3):248– 258, March 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_3/a_
                                        comparative_evaluation_
                                        of; internal&sk=0C220489.

Borger:2000:LCC

[BG00a] E. Börger and R. Gotzhein. The light control case study: a synopsis. J.UCS:

Journal of Universal Computer Science, 6(7): 582–585, July 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
7/the_light_control_
case.

Borger:2000:REL

[BG00b] E. Börger and R. Gotzhein.

Requirements engineering — The Light Control Case Study. J.UCS: Journal of Universal Computer Science, 6(7):??, July 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
7/requirements_engineering_
the_light.

Borger:2001:ASM

[BG01] E. Börger and U. Glässer.

Abstract state machines 2001: New developments and applications

— J.UCS special issue. J.UCS: Journal of Universal Computer Science, 7(11):914–916, November 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
7_11/abstract_state_
machines_2001.

Bernardini:2004:PS

[BG04] F. Bernardini and M. Gheorghe. Population P systems. J.UCS: Journal of Universal Computer Science, 10(5):509– 539, May 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_10_5/population_
p_systems.

Benevides:2010:VMA

[BGBA10] A. Botti Benevides, G. Guizzardi, B. F. Bastos Braga, and J. P. Andrade Almeida. Validating modal aspects of OntoUML conceptual models using automatically generated visual world structures. J.UCS: Journal of Universal Computer Science, 16(20):2904–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
20/validating_modal_
aspects_of.

Barley:2002:ATT

[BGK02] M. Barley, H. W. Guesgen, and G. Karl. An architecture for a three-tier path-finder. J.UCS: Journal of Universal Computer Science, 8(8):739– 750, August 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
8_8/an_architecture_
for_a.

Baca-Gomez:2016:WSS

[BGMR^{+}16] Y. R. Baca-Gomez, A. Martinez, P. Rosso, H. Estrada, and D. I. Hernandez Farias. Web service SWePT: A hybrid opinion mining approach. J.UCS: Journal of Universal Computer Science, 22(5):671–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
22_5/web_service_swept_
a.

Badica:2007:IRB

[BGP07] C. Badica, M. Ganzha, and M. Paprzycki. Implementing rule-based automated price negotiation in an agent system. J.UCS: Journal of Universal Computer Science, 13(2):244–266, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
2/implementing_rule_
based_automated.

Berlanga:2008:LDA

[BGP08] A. J. Berlanga and F. J.

García-Peñalvo. Learning design in adaptive educational hypermedia systems. J.UCS: Journal of Universal Computer Science, 14(22):3627–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_22/learning_
                                        design_in_adaptive.

Bordihn:2000:GSN

[BH00] H. Bordihn and M. Holzer. Grammar systems with negated conditions in their cooperation protocols. J.UCS: Journal of Universal Computer Science, 6(12):1165–1184, December 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        6_12/grammar_systems_
                                        with_negated.

Berghammer:2001:MSW

[BH01] R. Berghammer and T. Hoffmann. Modeling sequences within the RelView system. J.UCS: Journal of Universal Computer Science, 7(2):107–123, February 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_7_
                                        2/modeling_sequences_
                                        within_the.

Bordihn:2002:CCS

[BH02] H. Bordihn and M. Holzer.

On the computational complexity of synchronized context-free languages. J.UCS: Journal of Universal Computer Science, 8(2):119– 140, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
2/on_the_computational_
complexity.

Bridges:2008:CNM

[BH08] D. S. Bridges and R. S.

Havea. Constructive notions of maximality for ideals. J.UCS: Journal of Universal Computer Science, 14(22):3648–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
22/constructive_notions_
of_maximality.

Berman:2014:KCM

[BH14] B. Berman and J. P.

Hourcade. Keyboardcard menus: A new presentation of non-standard shortcuts. J.UCS: Journal of Universal Computer Science, 20(7):986– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_

                                        7/keyboard_card_menus_
                                        a.

Bravo:2005:UCC

[BHC05] J. Bravo, R. Hervás, and G. Chavira. Ubiquitous computing in the classroom: An approach through identification process. J.UCS: Journal of Universal Computer Science, 11(9):1494– 1504, ???? 2005. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_11_9/ubiquitous_
                                        computing_in_the.

Benyo:2006:DIE

[BHH^{+}06] B. Benyó, M. F. Hatwágner, T. Heckenast, K. Kovács, Á. Varga, and N. Varjasi. Design and implementation of enum-based services. J.UCS: Journal of Universal Computer Science, 12(9):1128–1138, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_12_9/design_
                                        and_implementation_of.

Bauer:2010:CCA

[BHK10] A. Bauer, P. Hertling, and K.-I. Ko. Computability and complexity in analysis. J.UCS: Journal of Universal Computer Science, 16(18):2495–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16;   http:
//www.jucs.org/jucs_
16_18/computability_
and_complexity_in;
http://www.jucs.org/
jucsrssfeed-issue.

Bohm:2002:TMG

[BHQW02] K. Böhm, G. Heyer, U. Quasthoff, and C. Wolff. Topic map generation using text mining. J.UCS: Journal of Universal Computer Science, 8(6): 623–633, June 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
6/topic_map_generation_
using.

Berghammer:2003:JUS

[BHRS03] R. Berghammer, D. Haneberg, W. Reif, and G. Schellhorn. J.UCS special issue on tools for system design and verification. J.UCS: Journal of Universal Computer Science, 9(2):86–87, February 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/free;
http://www.jucs.org/
jucs_9_2/tools_for_system_
design.

Bravo:2006:VSC

[BHS^{+}06] J. Bravo, R. Hervás, I. Sánchez, G. Chavira, and S. Nava. Visualization services in a conference context: An approach by RFID technology. J.UCS: Journal of Universal Computer Science, 12(3):270– 283, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12_
                                        3/visualization_services_
                                        in_a.

Barros:2008:ECW

[BI08] A. Barros and R. Ierusalimschy. Eliminating cycles in weak tables. J.UCS: Journal of Universal Computer Science, 14 (21):3481–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_21/eliminating_cycles_
                                        in_weak.

Bicen:2015:RSL

[Bic15] H. Bicen. The role of social learning networks in mobile assisted language learning: Edmodo as a case study. J.UCS: Journal of Universal Computer Science, 21(10): 1297–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_10/the_role_
of_social.

Brattka:2008:CCA

[BISZ08] V. Brattka, H. Ishihara, M. Schröder, and N. Zhong. Computability and complexity in analysis. J.UCS: Journal of Universal Computer Science, 14(6):800, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14;
http://www.jucs.org/
jucs_14_6#;   http://
www.jucs.org/jucs_14_
6/computability_and_
complexity_in.

Bernard:1997:FSM

[BJ97] M. Bernard and F. Jacquenet.

Free space modeling for placing rectangles without overlapping. J.UCS: Journal of Universal Computer Science, 3(6): 703–720, June 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs_3_6/free_space_
modeling_for;   internal&
sk=05460486.

Burton:2005:ASD

[BJ05a] J. Burton and C. B.

Jones. Atomicity in system design and execution (Proceedings of Dagstuhl-Seminar 04181). J.UCS: Journal of Universal Computer Science, 11(5): 634–635, May 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
 http://www.jucs.org/
jucs_11_5/atomicity_
in_system_design.

Burton:2005:IAO

[BJ05b] J. Burton and C. B. Jones. Investigating atomicity and observability. J.UCS: Journal of Universal Computer Science, 11(5):661–686, May 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
5/investigating_atomicity_
and_observability.

Bouassida:2015:IDP

[BJMBA15] N. Bouassida, S. Jamoussi, A. Msaed, and H. Ben-Abdallah. An interactive design pattern selection method. J.UCS: Journal of Universal Computer Science, 21 (13):1746–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
21_13/an_interactive_
design_pattern.

Bjorner:2001:TSE

[Bjø01] D. Bjørner. On teaching software engineering based on formal techniques — thoughts about and plans for — a different software engineering text book. J.UCS: Journal of Universal Computer Science, 7(8):641– 667, August 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_7_8/on_teaching_
software_engineering.

Bourimi:2013:PCS

[BKH^{+}13] M. Bourimi, D. Kesdogan, M. Heupel, D. e. D. I. Abou-Tair, and N. Lambropoulos. PETs at CSCL service: Underutilised potentials for privacy enhancing distance education. J.UCS: Journal of Universal Computer Science, 19(7):912– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
7/PETs_at_CSCL_service.

Blythe:2008:IIM

[BKK^{+}08] J. Blythe, D. Kapoor, C. A. Knoblock, K. Lerman, and S. Minton. Information integration for the masses. J.UCS: Journal of Universal Computer Science, 14(11): 1811–??, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_11/information_integration_
                                        for_the.

Botha:2012:FCB

[BKL12] C.-L. Botha, E. Kritzinger, and M. Loock. A framework for the comparison of best practice recommendations and legal requirements for South African banks. J.UCS: Journal of Universal Computer Science, 18(6): 845–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        18_6/a_framework_for_
                                        the.

Blomer:2006:RSB

[Blõ06] J. Blõmer. Randomness and secrecy — a brief introduction. J.UCS: Journal of Universal Computer Science, 12(6):654– 671, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        12_6/randomness_and_
                                        secrecy_a.

Baukus:2001:VPP

[BLS01] K. Baukus, Y. Lakhnech, and K. Stahl. Verification of parameterized protocols. J.UCS: Journal of Universal Computer Science, 7(2):141– 158, February 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
7_2/verification_of_
parameterized_protocols.

Brit:1996:WFV

[BM96] Hagit Brit and Shlomo Moran. Wait-freedom vs. bounded wait freedom in public data structures. J.UCS: Journal of Universal Computer Science, 2(1):2–19, January 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
2_1/wait_freedom_vs_
bounded.

Boerger:1997:IAS

[BM97] E. Boerger and L. Mearelli.

Integrating ASMs into the software development life cycle. J.UCS: Journal of Universal Computer Science, 3(5):603– 665, May 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs_3_5/integrating_
asm;   http://www.jucs.
org/jucs_3_5/integrating_
asm; internal&sk=05460486.

Blundo:1999:RMS

[BM99] C. Blundo and B. Masucci. Randomness in multi-secret sharing schemes. J.UCS: Journal of Universal Computer Science, 5(7):367– 389, July 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_5_7/randomness_
                                        in_multi_secret.

Bohinc:2003:SES

[BM03a] T. Bohinc and S. Markham. The strong effects of the soft factors of knowledge management. J.UCS: Journal of Universal Computer Science, 9(7): 581–591, July 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        7/the_strong_effects_
                                        of.

Bonifacio:2003:RDK

[BM03b] M. Bonifacio and A. Molani. The richness of diversity in knowledge creation: An interdisciplinary overview. J.UCS: Journal of Universal Computer Science, 9(6): 491–500, June 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_

9_6/the_richness_of_
diversity.

Burkhard:2005:TMV

[BM05] R. A. Burkhard and M. Meier. Tube map visualization: Evaluation of a novel knowledge visualization application for the transfer of knowledge in long-term projects. J.UCS: Journal of Universal Computer Science, 11(4):473–494, April 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
4/tube_map_visualization_
evaluation.

Bigonha:2007:SPS

[BM07] R. S. Bigonha and M. Musicante. Selected papers from SBLP 2007: The 11th Brazilian Symposium on Programming Languages. J.UCS: Journal of Universal Computer Science, 13(6):679– 681, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13;
http://www.jucs.org/
jucs_13_6#; http://www.
jucs.org/jucs_13_6/sblp_
2007.

Brodic:2011:NAW

[BM11] D. Brodić and Z. Milivojević. A new approach to water flow algorithm for text line segmentation. J.UCS: Journal of Universal Computer Science, 17(1):30– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
1/a_new_approach_to.

Blagojevic:2013:CLS

[BM13] M. Blagojević and M. Milosević. Collaboration and learning styles in pure online courses: an action research. J.UCS: Journal of Universal Computer Science, 19(7):984– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
7/collaboration_and_
learning_styles.

Barreto:2011:SPD

[BMa11] A. Silva Barreto, L. G. P. Murta, and A. R. avalcanti da Rocha. Software process definition: a reuse-based approach. J.UCS: Journal of Universal Computer Science, 17(13):1765–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_13/software_
process_definition_a.

Bravo:2005:IET

[BMG^{+}05] C. Bravo, M. J. Marcelino, A. Gomes, M. Esteves, and A. J. Mendes. Integrating educational tools for collaborative computer programming learning. J.UCS: Journal of Universal Computer Science, 11(9):1505–1517, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
9/integrating_educational_
tools_for.

Brank:2008:FSC

[BMGMF08] J. Brank, D. Mladeni, M. Grobelnik, and N. Mili-Frayling. Feature selection for the classification of large document collections. J.UCS: Journal of Universal Computer Science, 14(10):1562–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
10/feature_selection_
for_the.

Bevilacqua:2009:BGR

[BMM^{+}09] V. Bevilacqua, G. Mastronardi, F. Menolascina, P. Pannarale, and G. Romanazzi. Bayesian gene regulatory network inference optimization by means of genetic algorithms. J.UCS: Journal of Universal Computer Science, 15(4):826– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15_
                                        4/bayesian_gene_regulatory_
                                        network.

Bajard:1995:F

[BMMM95] J.-C. Bajard, D. Michelucci, J.-M. Moreau, and J.- M. Muller. Foreword. J.UCS: Journal of Universal Computer Science, 1(7):436–438, July 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        Foreword.

Brodic:2014:ASD

[BMMM14] D. Brodić, C. A. B. Mello, C. A. Maluckov, and Z. N. Milivojevic. An approach to skew detection of printed documents. J.UCS: Journal of Universal Computer Science, 20(4):488– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_20_
                                        4/an_approach_to_skew.

Baldassarri:2014:TBH

[BMUF14] S. Baldassarri, J. A. Macías, and J. Urquiza-Fuentes. Trending breakthroughs in human-computer interaction. J.UCS: Journal of Universal Computer Science, 20(7):941– ??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_20;   http:
//www.jucs.org/jucs_
20_7/trending_breakthroughs_
in_human.

Bittencourt:2012:DLB

[BMV12] L. F. Bittencourt, F. K.

Miyazawa, and A. L. Vignatti. Distributed load balancing algorithms for heterogeneous players in asynchronous networks. J.UCS: Journal of Universal Computer Science, 18(20):2771–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
20/distributed_load_
balancing_algorithms.

Buckley:2015:RVV

[BNA15] N. Buckley, A. K. Nagar, and S. Arumugam. On real-valued visual cryptographic basis matrices. J.UCS: Journal of Universal Computer Science, 21(12):1536–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        21_12/on_real_valued_
                                        visual.

Bulcao-Neto:2011:ULS

[BNCGD^{+}11] R. F. Bulcão-Neto, J. A. Camacho-Guerrero, M. Dutra, Á. Barreiro, J. Parapar, and A. A. Macedo. The use of latent semantic indexing to mitigate OCR effects of related document images. J.UCS: Journal of Universal Computer Science, 17(1):64–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_17_1/the_use_of_
                                        latent.

Bodlaender:2001:GNH

[Bod01] H. L. Bodlaender. A generic NP-hardness proof for a variant of graph coloring. J.UCS: Journal of Universal Computer Science, 7(12):1114– 1124, December 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_7_12/a_generic_
                                        np_hardness.

Boerger:1997:I

[Boe97a] E. Boerger. Introduction. J.UCS: Journal of Universal Computer Science, 3(5):414, May 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_3_
5/intro.

Boerger:1997:JSA

[Boe97b] E. Boerger. JUCS special ASM issue. Part II. J.UCS: Journal of Universal Computer Science, 3(5):414–415, May 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_volume_
3;   http://medoc.springer.
de:8000/jucs_3_5/intro;
internal&sk=05460486;
 internal&sk=05460486#
toc_3_5.

Boerger:1997:TYG

[Boe97c] E. Boerger. Ten years of Gurevich’s abstract state machines. J.UCS: Journal of Universal Computer Science, 3(4):230– 232, April 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_volume_
3;   http://medoc.springer.
de:8000/jucs_3_4/introduction;
 http://www.jucs.org/
jucs_3_4/introduction;
internal&sk=05460486;
 internal&sk=05460486#
toc_3_4.

Bollig:2006:TMF

[Bol06] B. Bollig. Testing membership in formal languages implicitly represented by Boolean functions. J.UCS: Journal of Universal Computer Science, 12(6):710–724, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12_
                                        6/testing_membership_
                                        in_formal.

Borger:2002:ODA

[Bör02] E. Börger. The origins and the development of the ASM method for high level system design and analysis. J.UCS: Journal of Universal Computer Science, 8(1):2–74, January 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        1/the_origins_and_the.

Borzemski:2007:IPB

[Bor07] L. Borzemski. Internet path behavior prediction via data mining: Conceptual framework and case study. J.UCS: Journal of Universal Computer Science, 13(2):287– 316, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
2/internet_path_behavior_
prediction.

Bosserhoff:2008:BCF

[Bos08a]  V. Bosserhoff.  The

bit-complexity of finding nearly optimal quadrature rules for weighted integration. J.UCS: Journal of Universal Computer Science, 14(6):938– 955, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
6/the_bit_complexity_
of.

Bosserhoff:2008:NPC

[Bos08b] V. Bosserhoff. Notions of probabilistic computability on represented spaces. J.UCS: Journal of Universal Computer Science, 14(6):956–995, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_6/notions_of_
probabilistic_computability.

Bosserhoff:2009:EES

[Bos09]  V. Bosserhoff.  On

the effective existence of Schauder bases. J.UCS: Journal of Universal Computer Science, 15(6): 1145–??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        15_6/on_the_effective_
                                        existence.

Bozapalidis:1999:IPS

[Boz99] S. Bozapalidis. An introduction to polypodic structures. J.UCS: Journal of Universal Computer Science, 5(9):508– 520, September 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_5_
                                        9/an_introduction_to_
                                        polypodic.

Borghoff:1997:ITK

[BP97] U. Borghoff and R. Pareschi. Information technology for knowledge management. J.UCS: Journal of Universal Computer Science, 3(8):835–842, August 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs_3_8/information_
                                        technology_for_knowledge;
                                        http://medoc.springer.
                                        de:8000/jucs_volume_
                                        3; internal&sk=05460486;
                                         internal&sk=05460486#
                                        toc_3_8.

Borger:2008:QVA

[BP08] E. Börger and A. Prinz. Quo vadis abstract state machines? J.UCS: Journal of Universal Computer Science, 14(12): 1921–??, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14;   http://www.jucs.
org/jucs_14_12#;   http:
//www.jucs.org/jucs_
14_12/quo_vadis_abstract_
state.

Barrett:2009:DPD

[BP09] R. Barrett and C. Pahl.

Distribution pattern-driven development of service architectures. J.UCS: Journal of Universal Computer Science, 15(11): 2166–??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_11/distribution_pattern_
driven_development.

Boronat:2004:TES

[BPC04] A. Boronat, J. Pérez, and I. Carsí, J. Á. Ramos. Two experiences in software dynamics. J.UCS: Journal of Universal Computer Science, 10(4):428–453, April 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
4/two_experiences_in_
software.

Berke:2006:DEQ

[BPHN06] J. Berke, Z. Polgár, Z. Horváth, and T. Nagy. Developing on exact quality and classification system for plant improvement. J.UCS: Journal of Universal Computer Science, 12(9):1154– 1164, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_12_9/developing_
                                        on_exact_quality.

Borges:2011:CNA

[BPS11] M. R. S. Borges, J. A. Pino, and W. Shen. CSCWD: New applications and challenges. J.UCS: Journal of Universal Computer Science, 17(14):1884–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_17;   http:
                                        //www.jucs.org/jucs_
                                        17_14/CSCWD_new_applications_
                                        and;   http://www.jucs.
                                        org/jucsrssfeed-issue.

Bakhtiari:1997:WGC

[BPSN97] S. Bakhtiari, J. Pieprzyk, and R. Safavi-Naini. On the weaknesses of Gong’s collisionful hash function. J.UCS: Journal of Universal Computer Science, 3(3):185– 196, March 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_3_
3/hash_function.

Biemann:2003:ADA

[BQBW03] C. Biemann, U. Quasthoff, K. Böhm, and C. Wolff. Automatic discovery and aggregation of compound names for the use in knowledge representations. J.UCS: Journal of Universal Computer Science, 9(6):530– 541, June 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
6/automatic_discovery_
and_aggregation.

Butter:2014:IPM

[BQV14] M. Careaga Butter, M. G. Badilla Quintana, and E. Sepulveda Valenzuela. Incremental prototyping model for the development of educational platforms: a process of design and quality standards. J.UCS: Journal of Universal Computer Science, 20(10):1407–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
10/incremental_prototyping_
model_for.

Bella:1997:FAK

[BR97] G. Bella and E. Riccobene. Formal analysis of the Kerberos authentication system. J.UCS: Journal of Universal Computer Science, 3(12):1337– 1381, December 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_3_12/
                                        formal_analysis_of_the.

Bettini:2003:DGC

[BR03] C. Bettini and S. Ruffini. Direct granularity conversions among temporal constraints. J.UCS: Journal of Universal Computer Science, 9(9):1123– 1136, September 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        9/direct_granularity_
                                        conversions_among.

Bagheri:2005:HDF

[BR05] A. Bagheri and M. Razzazi. How to draw free trees inside bounded simple polygons. J.UCS: Journal of Universal Computer Science, 11(6): 804–829, ???? 2005. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        11_6/how_to_draw_free.

Balaban:2007:PMP

[BR07] A. T. Balaban and M. Randić. Perfect matchings in polyhexes, or recent graph-theoretical contributions to benzenoids. J.UCS: Journal of Universal Computer Science, 13(11):1514– 1539, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
11/perfect_matchings_
in_polyhexes.

Brattka:2002:SNF

[Bra02] V. Brattka. Some notes on fine computability. J.UCS: Journal of Universal Computer Science, 8(3):382–395, March 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_8_3/some_notes_
on_fine.

Bradley:2015:MLL

[Bra15] L. Bradley. The mobile language learner — use of technology in language learning. J.UCS: Journal of Universal Computer Science, 21(10):1269–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_

10/the_mobile_language_
learner.

Boticario:2012:ALL

[BRAS^{+}12] J. G. Boticario, A. Rodriguez-Ascaso, O. C. Santos, E. Raffenne, L. Montandon, D. Roldán, and F. Buendía. Accessible lifelong learning at higher education: Outcomes and lessons learned at two different pilot sites in the EU4ALL Project. J.UCS: Journal of Universal Computer Science, 18(1):62–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_1/accessible_
lifelong_learning_at.

Barbosa:2009:CSE

[BRF^{+}09] P. Barbosa, F. Ramalho, J. Figueiredo, A. Júnior, A. Costa, and L. Gomes. Checking semantics equivalence of MDA transformations in concurrent systems. J.UCS: Journal of Universal Computer Science, 15(11):2196–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_11/checking_
semantics_equivalence_
of.

Becker:2008:SCO

[BRH^{+}08] C. Becker, A. Rauber, V. Heydegger, J. Schnasse, and M. Thaller. Systematic characterisation of objects in digital preservation: The eXtensible characterisation languages. J.UCS: Journal of Universal Computer Science, 14(18):2936–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
18/systematic_characterisation_
of_objects.

Bravo:2008:DHC

[BRO08] C. Bravo, M. A. Redondo, and M. Ortega. Designing the human computer interaction: Trends and challenges. J.UCS: Journal of Universal Computer Science, 14(9):1388– ??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14;
 http://www.jucs.org/
jucs_14_9#;   http://
www.jucs.org/jucs_14_
9/designing_the_human_
computer.

Benso:1999:FIE

[BRR99] A. Benso, M. Rebaudengo, and M. S. Reorda. Fault injection for embedded microprocessor-based systems. J.UCS: Journal of Universal Computer Science, 5(10):693– 711, October 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        5_10/fault_injection_
                                        for_embedded.

Beynon:2000:FSO

[BRS00a] M. Beynon, J. Rungrattanaubol, and J. Sinclair. Formal specification from an observation-oriented perspective. J.UCS: Journal of Universal Computer Science, 6 (4):407–421, April 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_6_4/formal_
                                        specification_from_an.

Borger:2000:CRA

[BRS00b] E. Börger, E. Riccobene, and J. Schmid. Capturing requirements by abstract state machines: The light control case study. J.UCS: Journal of Universal Computer Science, 6(7):597– 620, July 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_6_
                                        7/capturing_requirements_
                                        by_abstract.

Brucker:2003:HZP

[BRW03] Achim D. Brucker, Frank Rittinger, and Burkhart Wolff. HOL-Z 2.0: a proof environment for Zspecifications. J.UCS: Journal of Universal Computer Science, 9(2): 152–172, February 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_9_2/hol_z_2.

Borghoff:1996:CKH

[BS96] U. Borghoff and J. Schlichter.

On combining the knowledge of heterogeneous information repositories. J.UCS: Journal of Universal Computer Science, 2(7):514–532, July 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
2_7/on_combining_the_
knowledge.

Borger:2001:NAM

[BS01] E. Börger and D. Sona. A neural abstract machine. J.UCS: Journal of Universal Computer Science, 7(11):1006–1023, November 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_

                                        11/a_neural_abstract_
                                        machine.

Bergmann:2003:SCB

[BS03] R. Bergmann and M. Schaaf. Structural case-based reasoning and ontologybased knowledge management: a perfect match? J.UCS: Journal of Universal Computer Science, 9(7):608–626, July 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_9_7/structural_
                                        case_based_reasoning.

Berczes:2006:PMP

[BS06] T. Bérczes and J. Sztrik. Performance modeling of proxy cache servers. J.UCS: Journal of Universal Computer Science, 12(9):1139–1153, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12_
                                        9/performance_modeling_
                                        of_proxy.

Boticario:2008:SBM

[BS08] J. G. Boticario and O. C. Santos. A standardsbased modelling approach for dynamic generation of adaptive learning scenarios. J.UCS: Journal of Universal Computer Science, 14(17):2859–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
17/a_standards_based_
modelling.

Baldassarri:2011:CSR

[BS11] S. Baldassarri and F. Seron.

Coordinated system for real time muscle deformation during locomotion. J.UCS: Journal of Universal Computer Science, 17(3):349– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
3/coordinated_system_
for_real.

Baars:2012:ASR

[BS12a] T. Baars and M. Spruit.

Analysing the security risks of Cloud adoption using the SeCA model: a case study. J.UCS: Journal of Universal Computer Science, 18(12): 1662–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_12/analysing_
the_security_risks.

Bucki:2012:MLO

[BS12b] R. Bucki and P. Suchanek.

The method of logistic optimization in Ecommerce. J.UCS: Journal of Universal Computer Science, 18(10): 1238–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_10/the_method_of_
logistic.

Bukhari:2009:ABU

[BSB09] S. S. Bukhari, F. Shafait, and T. M. Breuel. Adaptive binarization of unconstrained hand-held camera-captured document images. J.UCS: Journal of Universal Computer Science, 15 (18):3343–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_18/adaptive_
binarization_of_unconstrained.

Bent:1999:NBW

[BSHI99] R. Bent, M. Schear, L. Hemaspaandra, and G. Istrate. A note on bounded-weight errorcorrecting codes. J.UCS: Journal of Universal Computer Science, 5(12): 817–827, December 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_5_12/a_note_on_
bounded.

Barroso:2013:TEA

[BSP^{+}13] J. Barroso, F. E. Sandnes, H. Paredes, L. Hadjileontiadis, and P. Martins. Technologies for enhancing accessibility and fighting info-exclusion. J.UCS: Journal of Universal Computer Science, 19(18):2637–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_19;   http:
//www.jucs.org/jucs_
19_18/technologies_for_
enhancing_accessibility.

Borges:2011:CAC

[BSPO11] M. R. S. Borges, W. Shen, J. A. Pino, and S. F. Ochoa. CSCWD: Applications and challenges. J.UCS: Journal of Universal Computer Science, 17(2):162–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_17;   http:
//www.jucs.org/jucs_
17_2/cscwd_applications_
and_challenges;   http:
//www.jucs.org/jucsrssfeed-
issue.

Borger:2009:DBJ

[BST09] E. Börger, O. Sörensen, and B. Thalheim. On defining the behavior of OR-joins in business process models. J.UCS: Journal of Universal Computer Science, 15(1):3– ??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_1/on_defining_the_
behavior.

Barrat:2008:PLM

[BT08] S. Barrat and S. Tabbone. A progressive learning method for symbol recognition. J.UCS: Journal of Universal Computer Science, 14(2):224– 236, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
2/a_progressive_learning_
method.

Burgos:2007:FSM

[BTD^{+}07] D. Burgos, C. Tattersall, M. Dougiamas, H. Vogten, and R. Koper. A first step mapping IMS learning design and moodle. J.UCS: Journal of Universal Computer Science, 13(7):924– 931, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
7/a_first_step_mapping.

Bicen:2013:USN

[BU13] H. Bicen and H. Uzunboylu. The use of social networking sites in education: a case study of Facebook. J.UCS: Journal of Universal Computer Science, 19(5):658– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
5/the_use_of_social.

Buford:2001:TSS

[Buf01] J. Buford. Telecommunication services and service management challenges. J.UCS: Journal of Universal Computer Science, 7(5):355– 365, May 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
5/telecommunication_
services_and_service.

Bulitko:1995:CPS

[Bul95] Vadim Bulitko. On completeness of pseudosimple sets. J.UCS: Journal of Universal Computer Science, 1(2):151– 154, February 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
2/on_completeness_of_
pseudosimple.

Burton:2005:RAV

[Bur05] J. Burton. Relaxing atomicity and verifying correctness: Considering the case of an asynchronous communication mechanism. J.UCS: Journal of Universal Computer Science, 11(5):771–802, May 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_11_
                                        5/relaxing_atomicity_
                                        and_verifying.

Burger:2008:NFM

[Bür08] T. Bürger. The need for formalizing media semantics in the games and entertainment industry. J.UCS: Journal of Universal Computer Science, 14(10):1775–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_10/the_need_
                                        for_formalizing.

Buza:2006:ECD

[Buz06]  A. Buza.  Extension

of CQL over dynamic databases. J.UCS: Journal of Universal Computer Science, 12(9):1165– 1176, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_12_9/extension_
of_cql_over%20.

Batista:2007:RAA

[BV07] T. Batista and M. Vieira.

RE-AspectLua — achieving reuse in AspectLua. J.UCS: Journal of Universal Computer Science, 13(6):786–805, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_6/re_aspectlua_achieving_
reuse.

Boella:2009:ACI

[BvdTV09] G. Boella, L. van der Torre, and S. Villata. Analyzing cooperation in iterative social network design. J.UCS: Journal of Universal Computer Science, 15(13):2676–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_13/analyzing_
cooperation_in_iterative.

Burkard:2008:IHT

[BVG08] B. Burkard, G. Vogeler, and S. Gruner. Informatics for historians: Tools for medieval document XML markup, and their impact on the historysciences. J.UCS: Journal of Universal Computer Science, 14(2):193– 210, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_2/informatics_for_
historians_tools.

Butoianu:2010:UCP

[BVV^{+}10] V. Butoianu, P. Vidal, K. Verbert, E. Duval, and J. Broisin. User context and personalized learning: a federation of contextualized attention metadata. J.UCS: Journal of Universal Computer Science, 16(16):2252–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
16/user_context_and_
personalized.

Botha:2009:RLB

[BvZH09] L. Botha, L. van Zijl, and M. Hoffmann. Realtime LEGO brick image retrieval with cellular automata. J.UCS: Journal of Universal Computer Science, 15(14): 2765–??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_14/realtime_lego_
brick_image.

Bridges:1997:CAD

[BY97] D. Bridges and W. Yuchuan.

Constructive aspects of the Dirichlet problem. J.UCS: Journal of Universal Computer Science, 3(11):1148–1161, November 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_3_11/
constructive_aspects_
of_the.

Baloian:2009:MSF

[BZ09] N. Baloian and G. Zurita. MC-supporter: Flexible mobile computing supporting learning though social interactions. J.UCS: Journal of Universal Computer Science, 15(9):1833–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
9/mc_supporter_flexible_
mobile.

Boussellaa:2008:MSF

[BZA08] W. Boussellaa, A. Zahour, and A. Alimi. A methodology for the separation of foreground/ background in Arabic historical manuscripts using hybrid methods. J.UCS: Journal of Universal Computer Science, 14(2):

284–298, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_2/a_methodology_for_
                                        the.

Bagues:2010:EPP

[BZM^{+}10] S. Alcalde Bagüés, A. Zeidler, I. R. Matias, C. Klein, and C. Fernandez Valdivielso. Enabling personal privacy for pervasive computing environments. J.UCS: Journal of Universal Computer Science, 16(3):341– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_16_
                                        3/enabling_personal_
                                        privacy_for.

Chenaina:2014:IFC

[CA14] T. Chenaina and A. Alraddadi. Identifying fuzzy controllers parameter by fuzzy clustering technique. J.UCS: Journal of Universal Computer Science, 20(2):107– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_20_
                                        2/identifying_fuzzy_
                                        controllers_parameters.

Campo:2006:SSD

[CAD^{+}06] C. Campo, F. Almenárez, D. Díaz, C. García-Rubio, and A. M. López. Secure service discovery based on trust management for ad-hoc networks. J.UCS: Journal of Universal Computer Science, 12(3):340–356, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_3/secure_service_
discovery_based.

Cabanas-Abascal:2013:IBG

[CAGMPGdAS13] A. Cabanas-Abascal, E. García-Machicado, L. Prieto-González, and A. de Amescua Seco. An item based geo-recommender system inspired by artificial immune algorithms. J.UCS: Journal of Universal Computer Science, 19(13):2013–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
13/an_item_based_geo-
recommender.

Cailliau:1995:AW

[Cai95]  R. Cailliau.  About

WWW. J.UCS: Journal of Universal Computer Science, 1(4):221– 231, April 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/about_
                                        www.

Cai:2010:MI

[Cai10] Y. Cai. Mobile intelligence. J.UCS: Journal of Universal Computer Science, 16(12):1650–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_16_
                                        12/mobile_intelligence.

Ciriaco:2006:DCM

[CAJ06] F. Ciriaco, T. Abrão, and P. J. E. Jeszensky. DS/ CDMA multiuser detection with evolutionary algorithms. J.UCS: Journal of Universal Computer Science, 12(4):450– 480, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12_
                                        4/ds_cdma_multiuser_
                                        detection.

Calude:1995:WRS

[Cal95] C. Calude. What is a random string. J.UCS: Journal of Universal Computer Science, 1(1):48–66, January 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        1_1/what_is_a_random/
                                        html/abstract.html.

Calude:1996:FUI

[Cal96a] C. Calude. The finite, the unbounded and the infinite. J.UCS: Journal of Universal Computer Science, 2(5):242– 244, May 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_2_5/the_finite_
the_unbounded.

Calude:1996:AIT

[Cal96b] Cristian Calude. Algorithmic information theory: Open problems. J.UCS: Journal of Universal Computer Science, 2(5):439–441, May 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
algorithmic_information_
theory_open_problems.

Campos:1998:CRK

[Cam98]  M. Campos.  Condi-

tional reasoning: a key to assessing computerbased knowledge-building communication processes. J.UCS: Journal of Universal Computer Science, 4(4):404–428, April 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_4/conditional_

                                        reasoning_a_key; internal&
                                        sk=0C220489.

Canas:2008:CEI

[Cañ08]  J. Cañas.  Cognitive

ergonomics in interface development evaluation. J.UCS: Journal of Universal Computer Science, 14(16):2630–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_16/cognitive_
                                        ergonomics_in_interface.

Capkovic:2005:ADC

[Cap05] F. Capkovic. An application of the DEDS control synthesis method. J.UCS: Journal of Universal Computer Science, 11(2):303–326, February 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_11_
                                        2/an_application_of_
                                        the.

Carlson:1995:BCE

[Car95]  P. Carlson.  BROCA:

a computerized environment for mediating scientific reasoning through writing. J.UCS: Journal of Universal Computer Science, 1(8):571– 590, August 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
a_computerized_environment_
for_mediating.

Carlson:1996:WGE

[Car96]  P. Carlson.  A word

from the Guest Editor: “Classroom of the Future”. J.UCS: Journal of Universal Computer Science, 2(10):660–662, October 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/a_word_
from_the_guest_editor.

Carlson:1998:AET

[Car98a] P. Carlson. Advanced educational technologies

— promise and puzzlement. J.UCS: Journal of Universal Computer Science, 4(3):210– 215, March 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_3/advanced_
educational_technologies_
promise; internal&sk=
0C220489.

Carlson:1998:AIE

[Car98b] Patricia A. Carlson. Assessment issues for educational software. J.UCS: Journal of Universal Computer Science, 4(4): 326–329, April 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_4/assessment_
                                        issues_for_educational;
                                        http://medoc.springer.
                                        de:8000/jucs/jucs_volume_
                                        4;   internal&sk=0C220489;
                                        internal&sk=0C220489#
                                        toc_4_4.

Carlson:2000:WIW

[Car00] P. A. Carlson. Wonders of the invisible workplace: IT and process reinvention. J.UCS: Journal of Universal Computer Science, 6(3):256–271, March 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_6_3/wonders_
                                        of_the_invisible.

Corchuelo:2008:WWD

[CAR08] R. Corchuelo, J. L. Arjona, and D. Ruiz. Wrapping Web data islands. J.UCS: Journal of Universal Computer Science, 14(11):1808–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_14;   http:
                                        //www.jucs.org/jucs_
                                        14_11#;   http://www.
                                        jucs.org/jucs_14_11/
                                        wrapping_web_data_islands.

Cankar:2013:CAC

[CAS^{+}13] M. Cankar, M. Artac, M. Sterk, U. Lotric, and B. Slivnik. Coallocation with collective requests in grid systems. J.UCS: Journal of Universal Computer Science, 19(3):282– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
3/coallocation_with_
collective_requests.

Chalub:2004:MRS

[CB04] F. Chalub and C. Braga.

A modular rewriting semantics for CML. J.UCS: Journal of Universal Computer Science, 10(7): 789–807, July 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_10_7/a_modular_
rewriting_semantics.

Chua:2012:ESG

[CB12] A. Y. Chua and R. S.

Balkunje. An exploratory study of game-based Mlearning for software project management. J.UCS: Journal of Universal Computer Science, 18 (14):1933–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_14/an_exploratory_
study_of.

Cordeiro:2007:OCA

[CBBT07] E. S. Cordeiro, R. S. Bigonha, M. A. S. Bigonha, and F. Tirelo. Optimized compilation of around advice for aspect oriented programs. J.UCS: Journal of Universal Computer Science, 13(6):753–766, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_13_6/optimized_
                                        compilation_of_around.

Cacho:2005:ADA

[CBF05] N. Cacho, T. Batista, and F. Fernandes. AspectLua — a dynamic AOP approach. J.UCS: Journal of Universal Computer Science, 11(7):1177–1197, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_11_7/aspect_
                                        lua_a_dynamic.

Cuel:2004:KNR

[CBG04] R. Cuel, M. Bonifacio, and M. Grosselle. Knowledge nodes: the reification of organizational communities. the pizzarotti case study. J.UCS: Journal of Universal Computer Science, 10(3):227–234, March 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
10_3/knowledge_nodes_
the_reification.

Calsavara:2006:OMI

[CBN^{+}06] A. Calsavara, A. Borges, L. Nunes, D. Variani, and C. Kolb. An object model for interoperable systems. J.UCS: Journal of Universal Computer Science, 12(7):885– 902, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
7/an_object_model_for.

Chavira:2010:PMP

[CBNDR10] G. Chavira, J. Bravo, S. W. Nava-Díaz, and J. C. Rolón. PIC-TAC: a model for perceiving touch interaction through tagging context. J.UCS: Journal of Universal Computer Science, 16(12):1577–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
12/pictac_a_model_for.

Cruz:2005:AOG

[CBO05] A. M. Cruz, L. S. Barbosa, and J. N. Oliveira. From algebras to objects: Generation and composition. J.UCS: Journal of Universal Computer Science, 11(10):1580–1612, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
10/from_algebras_to_
objects.

Cao:2005:NIA

[CBR^{+}05] F. Cao, B. R. Bryant, R. R. Raje, A. M. Olson, M. Auguston, W. Zhao, and C. C. Burt. A non-invasive approach to assertive and autonomous dynamic component composition in the service-oriented paradigm. J.UCS: Journal of Universal Computer Science, 11(10): 1645–1675, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_10/a_non_invasive_
approach.

Chamlertwat:2012:DCI

[CBRH12] W. Chamlertwat, P. Bhattarakosol, T. Rungkasiri, and C. Haruechaiyasak. Discovering consumer insight from Twitter via sentiment analysis. J.UCS: Journal of Universal Computer Science, 18(8): 973–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_8/discovering_consumer_
insight_from.

Chaitin-Chatelin:1996:FPA

[CC96] F. Chaitin-Chatelin. Is finite precision arithmetic useful for physics? J.UCS: Journal of Universal Computer Science, 2(5):380–395, May 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
is_finite_precision_
arithmetic_useful_for_
physics.

Capra:2007:SEP

[CC07] L. Capra and W. Cazzola. Self-evolving Petri nets. J.UCS: Journal of Universal Computer Science, 13(13):2002–2034, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_13_13/self_
evolving_petri_nets.

Collange:2008:PEQ

[CC08a] D. Collange and J.-L.

Costeux. Passive estimation of quality of experience. J.UCS: Journal of Universal Computer Science, 14(5):625– 641, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        5/passive_estimation_
                                        of_quality.

Cristea:2008:AAA

[CC08b] A. Cristea and R. Carro. Authoring of adaptive and adaptable hypermedia. J.UCS: Journal of Universal Computer Science, 14(17):2756–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_14;   http:
                                        //www.jucs.org/jucs_
                                        14_17#;   http://www.
                                        jucs.org/jucs_14_17/
                                        authoring_of_adaptive_
                                        and.

Caballero:2008:IIQ

[CCCP08] I. Caballero, A. Caro, C. Calero, and M. Piattini. IQM3: Information quality management maturity model. J.UCS: Journal of Universal Computer Science, 14(22):3658–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_14_22/iqm3_
                                        information_quality_
                                        management.

Calude:2003:WVT

[CCD03] C. S. Calude, E. Calude, and M. J. Dinneen. What is the value of taxicab (6)? J.UCS: Journal of Universal Computer Science, 9(10):1196–1203, October 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
10/what_is_the_value.

Chang:2006:WBD

[CCH06] C.-C. Chang, C.-S. Cheng, and Y.-S. Huang. A Web-based decision support system for chronic diseases. J.UCS: Journal of Universal Computer Science, 12(1):115–125, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_12_1/a_web_
based_decision.

Cintra:2008:ACF

[CCHdCN08] M. E. Cintra, H. A. Camargo, E. R. Hruschka, and M. do Carmo Nicoletti. Automatic construction of fuzzy rule bases: a further investigation into two alternative inductive approaches. J.UCS: Journal of Universal Computer Science, 14(15): 2456–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_15/automatic_
construction_of_fuzzy.

Cali:2009:DQO

[CCM09] A. Calì, D. Calvanese, and D. Martinenghi. Dynamic query optimization under access limitations and dependencies. J.UCS: Journal of Universal Computer Science, 15(1):33– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15_
                                        1/dynamic_query_optimization_
                                        under.

Cortellessa:2008:EAS

[CCMP08] V. Cortellessa, I. Crnkovic, F. Marinelli, and P. Potena. Experimenting the automated selection of COTS components based on cost and system requirements. J.UCS: Journal of Universal Computer Science, 14(8):1228–1255, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        8/experimenting_the_
                                        automated_selection.

Carvalho:2007:URA

[CCP^{+}07] N. Carvalho, A. Correia, Jr., J. Pereira, L. Rodrigues, R. Oliveira, and S. Guedes. On the use of a reflective architecture to augment database management systems. J.UCS: Journal of Universal Computer Science, 13(8):1110– 1135, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
8/on_the_use_of.

Cubo:2011:CAC

[CCP11] J. Cubo, C. Canal, and E. Pimentel. Contextaware composition and adaptation based on model transformation. J.UCS: Journal of Universal Computer Science, 17(5):777–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_5/context_aware_composition_
and.

Cristofor:2000:GCD

[CCS00] D. Cristofor, L. Cristofor, and D. A. Simovici. Galois connections and data mining. J.UCS: Journal of Universal Computer Science, 6(1):60–73, January 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
1/galois_connections_
and_data.

Cristea:2010:AAA

[CCS10] A. I. Cristea, R. Carro, and C. D. Stewart. Advances in authoring of adaptive Web-based systems. J.UCS: Journal of Universal Computer Science, 16(19):2754–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16;
 http://www.jucs.org/
jucs_16_19/advances_
in_authoring_of;   http:
//www.jucs.org/jucsrssfeed-
issue.

Choi:2015:PMJ

[CCYK15] J. Choi, C. Choi, I. You, and P. Kim. Polymorphic malicious JavaScript code detection for APT attack defence. J.UCS: Journal of Universal Computer Science, 21(3):369– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
3/polymorphic_malicious_
javascript_code.

Costanza:2010:EHC

[CD10] P. Costanza and T. D’Hondt. Embedding hygiene-compatible macros in an unhygienic macro system. J.UCS: Journal of Universal Computer Science, 16 (2):271–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/

jucs_16_2/embedding_
hygiene_compatible_macros.

Colson:2013:IPC

[CD13] L. Colson and V. Demange. Investigations on a pedagogical calculus of constructions. J.UCS: Journal of Universal Computer Science, 19(6):729–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_6/investigations_
on_a_pedagogical.

Chu:2009:ILI

[CDBZ09] T.-Q. Chu, A. Drogoul, A. Boucher, and J.-D. Zucker. Interactive learning of independent experts’ criteria for rescue simulations. J.UCS: Journal of Universal Computer Science, 15(13):2701–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
13/interactive_learning_
of_independent.

Chu:2009:NGT

[CDCH09] H.-C. Chu, D.-J. Deng, H.-C. Chao, and Y.-M. Huang. Next generation of terrorism: Ubiquitous cyber terrorism with the accumulation of all intangible fears. J.UCS: Journal of Universal Computer Science, 15(12):2373–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_15_12/next_
generation_of_terrorism.

Colucci:2003:FAO

[CDD^{+}03] S. Colucci, T. Di Noia, E. Di Sciascio, F. M. Donini, M. Mongiello, and M. Mottola. A formal approach to ontologybased semantic match of skills descriptions. J.UCS: Journal of Universal Computer Science, 9(12):1437–1454, December 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
12/a_formal_approach_
to.

Colucci:2004:SBA

[CDD^{+}04] S. Colucci, T. Di Noia, E. Di Sciascio, F. M. Donini, M. Mongiello, and G. Piscitelli. Semanticbased approach to task assignment of individual profiles. J.UCS: Journal of Universal Computer Science, 10(6):723– 730, June 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_

10_6/semantic_based_
approach_to.

Colucci:2007:SBS

[CDD^{+}07] S. Colucci, T. Di Noia, E. Di Sciascio, F. M. Donini, and A. Ragone. Semantic-based skill management for automated task assignment and courseware composition. J.UCS: Journal of Universal Computer Science, 13(9):1184–1212, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_9/semantic_based_
skill_management.

Courcelle:1997:NCG

[CDF97] B. Courcelle, R. G.

Downey, and M. R. Fellows. A note on the computability of graph minor obstruction sets for monadic second order ideals. J.UCS: Journal of Universal Computer Science, 3(11):1194– 1198, November 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_3_11/
a_note_on_the.

Colace:2014:MVL

[CDG14] F. Colace, M. De Santo, and L. Greco. A “Mobile Virtual Lab” for supporting engineering curricula. J.UCS: Journal of Universal Computer Science, 20(15):2054–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
15/a_mobile_virtual_
lab.

Cattaneo:2013:SIA

[CDP13] G. Cattaneo, G. De Maio, and U. Ferraro Petrillo. Security issues and attacks on the GSM standard: a review. J.UCS: Journal of Universal Computer Science, 19(16):2437–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_16/security_
issues_and_attacks.

Chowdhury:2009:RET

[CDR^{+}09] S. P. Chowdhury, S. Dhar, K. Rafferty, A. K. Das, and B. Chanda. Robust extraction of text from camera images using colour and spatial information simultaneously. J.UCS: Journal of Universal Computer Science, 15(18):3325–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
18/robust_extraction_
of_text.

Cechinel:2016:MMA

[CdSCSSA16] C. Cechinel, S. da Silva Camargo, M.-Á. Sicilia, and S. Sánchez-Alonso. Mining models for automated quality assessment of learning objects. J.UCS: Journal of Universal Computer Science, 22(1):94– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_22_
1/mining_models_for_
automated.

Chaudhuri:2006:DME

[CE06] P. Chaudhuri and T. Edward. An O(\sqrt{n}) distributed mutual exclusion algorithm using queue migration. J.UCS: Journal of Universal Computer Science, 12(2):140– 159, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
2/an_0_square_root.

Comi:2011:AIV

[CE11] A. Comi and M. J. Eppler.

Assessing the impact of visual facilitation on interorganizational collaboration: An experimental study. J.UCS: Journal of Universal Computer Science, 17(10):1430–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_17_10/assessing_
                                        the_impact_of.

Charfi:2015:RUI

[CEK15] S. Charfi, H. Ezzedine, and C. Kolski. RITA: a useR Interface evaluaTion frAmework. J.UCS: Journal of Universal Computer Science, 21(4):526–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_21_4/rita_a_
                                        user_interface.

Cerny:1997:ODA

[Cer97] A. Cerny. Optimal description of automatic paperfolding sequences. J.UCS: Journal of Universal Computer Science, 3(10):1085–1099, October 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_3_10/
                                        optimal_description_
                                        of_automatic.

Ceterchi:2000:LSP

[Cet00] R. Ceterchi. The lattice structure of pseudo-Wajsberg algebras. J.UCS:

Journal of Universal Computer Science, 6(1): 22–38, January 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_6_1/the_lattice_
structure_of.

Canto:2013:FSR

[CFF^{+}13] E. Cantó, M. Fons, F. Fons, M. López, and R. Ramos. Fast selfreconfigurable embedded system on Spartan-3. J.UCS: Journal of Universal Computer Science, 19(3):301–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_3/fast_self_
reconfigurable_embedded.

Calude:2015:DFP

[CFJS15] C. S. Calude, R. Freivalds, S. Jain, and F. Stephan. Deterministic frequency pushdown automata. J.UCS: Journal of Universal Computer Science, 21 (12):1563–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
21_12/deterministic_
frequency_pushdown_automata.

Chagas:2015:AMI

[CFMP15] D. Almeida Chagas, E. Sucupira Furtado, C. Loiola Brito Maia, and P. Rogerio Pinheiro. Analyzing the multicriteria of the interaction design of an educational map application for digital TV from user preferences. J.UCS: Journal of Universal Computer Science, 21(11): 1470–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_11/analyzing_
the_multicriteria_of.

Cordon-Franco:2004:NCM

[CFSC04] A. Cordón-Franco and F. Sancho-Caparrini. A note on complexity measures for probabilistic P systems. J.UCS: Journal of Universal Computer Science, 10(5):559– 566, May 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_10_5/a_note_on_
complexity.

Calude:1996:KCI

[CG96a] Cristian Calude and Cristian Grozea. Kraft-Chaitin inequality revisited. J.UCS: Journal of Universal Computer Science, 2(5):306– 310, May 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
5/kraft_chaitin_inequality_
revisited.

Chinowsky:1996:MIP

[CG96b] P. Chinowsky and R. Goodman. Managing interdisciplinary project teams through the Web. J.UCS: Journal of Universal Computer Science, 2(9):597–609, September 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
9/managing_interdisciplinary_
project_teams.

Cavaliere:2004:SSA

[CG04] M. Cavaliere and D. Genova. P systems with Symport/Antiport of rules. J.UCS: Journal of Universal Computer Science, 10(5):540–558, May 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_10_5/p_systems_
with_symport.

Collins:2009:ECS

[CG09] P. Collins and D. S.

Graça. Effective computability of solutions of differential inclusions: The ten thousand monkeys approach. J.UCS: Journal of Universal Computer Science, 15(6): 1162–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_6/effective_
computability_of_solutions.

Caballe:2012:CEI

[CGD^{+}12] S. Caballe, D. Ganan, I. Dunwell, A. Pierri, and T. Daradoumis. CC-LO: Embedding interactivity, challenge and empowerment into collaborative learning sessions. J.UCS: Journal of Universal Computer Science, 18(1):25–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_1/cc_lo_embedding_
interactivity.

Caballero-Gil:2009:GBA

[CGFSHG09] P. Caballero-Gil, A. Fúster-Sabater, and C. Hernández-Goya. Graph-based approach to the edit distance cryptanalysis of irregularly clocked linear feedback shift registers. J.UCS: Journal of Universal Computer Science, 15(15):2981–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.

org/jucs_15_15/graph_
based_approach_to.

Cabot:2014:ALC

[CGLdMAC14] A. Garcia Cabot, E. Garcia-Lopez, L. de Marcos, and J. Abraham-Curto. Adapting learning contents to mobile devices and context to improve students’ learning performance: A case study. J.UCS: Journal of Universal Computer Science, 20(15):2032–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_20_15/adapting_
learning_contents_to.

Collazos:2007:DCL

[CGP^{+}07a] C. A. Collazos, L. A.

Guerrero, J. A. Pino, S. F. Ochoa, and G. Stahl. Designing collaborative learning environments using digital games. J.UCS: Journal of Universal Computer Science, 13(7): 1022–1032, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_13_7/designing_
collaborative_learning_
environments.

Corcho:2007:OSD

[CGP07b] O. Corcho and A. Gómez-Pérez. ODEDialect: a set of declarative languages for implementing ontology translation systems. J.UCS: Journal of Universal Computer Science, 13(12):1805–1834, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
12/odedialect_a_set_
of.

Conde:2013:IDE

[CGPAP13] M. Á. Conde, F. J. García-Peñalvo, M. Alier, and J. Piguillem. The implementation, deployment and evaluation of a mobile personal learning environment. J.UCS: Journal of Universal Computer Science, 19(7):854– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
7/the_implementation_
deployment_and.

Chaves-Gonzalez:2007:PSS

[CGVRGPSP07] J. M. Chaves-González, M. A. Vega-Rodríguez, J. A. Gómez-Pulido, and J. M. Sánchez-Pérez. Pipeline-scheduling simulator for educational purpose. J.UCS: Journal of Universal Computer Science, 13(7):959–969, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
7/pipeline_scheduling_
simulator_for.

Chan:2007:BIC

[CH07] R. Chan and J. Hoshino.

Building immersive conversation environment using locomotive interactive character. J.UCS: Journal of Universal Computer Science, 13(2):149– 160, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
2/building_immersive_
conversation_environment.

Chaitin:1996:LM

[Cha96]  G. J. Chaitin.  The

limits of mathematics. J.UCS: Journal of Universal Computer Science, 2(5):270–305, May 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_2_5/the_limits_
of_mathematics.

Chaitin:2005:AIC

[Cha05] G. Chaitin. Algorithmic irreducibility in a cellular automata universe. J.UCS: Journal of Universal Computer Science, 11(12):1901–1903, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_11_
                                        12/algorithmic_irreducibility_
                                        in_a.

Chen:2009:NDM

[Che09] C.-L. Chen. A new detection method for distributed denial-of-service attack traffic based on statistical test. J.UCS: Journal of Universal Computer Science, 15(2):488– ??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        15_2/a_new_detection_
                                        method.

Chen:2016:DOS

[CHH16] R.-C. Chen, Hendry, and C.-Y. Huang. A domain ontology in social networks for identifying user interest for personalized recommendations. J.UCS: Journal of Universal Computer Science, 22(3):319– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_22_
                                        3/a_domain_ontology_
                                        in.

Cabrerizo:2010:SPB

[CHPHV10] F. J. Cabrerizo, R. Heradio, I. J. Pérez, and E. Herrera-Viedma. A selection process based on additive consistency to deal with incomplete fuzzy linguistic information. J.UCS: Journal of Universal Computer Science, 16(1):62– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
1/a_selection_process_
based.

Calude:2005:CCL

[CI05] C. S. Calude and H. Ishihara. Constructivity, computability, and logic. A collection of papers in honour of the 60th birthday of Douglas Bridges. J.UCS: Journal of Universal Computer Science, 11(12):1863–1864, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11;
 http://www.jucs.org/
jucs_11_12#;   http:
//www.jucs.org/jucs_
11_12/constructivity_
computability_and_logic.

Charuenporn:2012:QSM

[CI12] P. Charuenporn and S. Intakosum. QoS-security metrics based on ITIL and COBIT standard for measurement Web services. J.UCS: Journal of Universal Computer Science, 18(6):775– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_18_
                                        6/qos_security_metrics_
                                        based.

Calvo:2014:UCR

[CIM14] R. Calvo, A. Iglesias, and L. Moreno. User-centered requirement engineering for accessible chats in mlearning. J.UCS: Journal of Universal Computer Science, 20(7):964– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_20_
                                        7/user_centered_requirement_
                                        engineering.

Chesneaux:1998:DCC

[CJ98] J. M. Chesneaux and F. Jézéquel. Dynamical control of computations using the trapezoidal and Simpson’s rules. J.UCS: Journal of Universal Computer Science, 4(1):2–10, January 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_1/dynamical_
                                        control_of_computations.

Coleman:2007:AUC

[CJ07] J. W. Coleman and C. B.

Jones. Atomicity: a unifying concept in computer science. J.UCS: Journal of Universal Computer Science, 13(8):1042– 1043, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13;
 http://www.jucs.org/
jucs_13_8#;   http://
www.jucs.org/jucs_13_
8/atomicity_a_unifying_
concept.

Chen:2012:GDP

[CJH12] J. Chen, Q. Jin, and R. Huang. Goal-driven process navigation for individualized learning activities in ubiquitous networking and IoT environments. J.UCS: Journal of Universal Computer Science, 18(9):1132– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
9/goal_driven_process_
navigation.

Caffarel:2013:IBA

[CJO^{+}13] J. Caffarel, S. Jie, J. Olloqui, R. Martínez, and A. Santamaría. Implementation of a building automation system based on semantic modeling. J.UCS: Journal of Universal Computer Science, 19(17):2543–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_19_
                                        17/implementation_of_
                                        a_building.

Cui:2013:OSL

[CJZ13] T. Cui, C. Jin, and G. Zhang. Observations of skipjack-like structure with SP/SPS round function. J.UCS: Journal of Universal Computer Science, 19(16):2453–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_19_
                                        16/observations_of_skipjack_
                                        like.

Culik:1995:ASW

[CK95] Karel Culik II and Jarkko Kari. An aperiodic set of Wang cubes. J.UCS: Journal of Universal Computer Science, 1(10): 675–686, October 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        an_aperiodic_set_of_
                                        wang_cubes.

Cirilo:2008:PDT

[CKdL08] E. Cirilo, U. Kulesza, and C. J. Pereira de Lucena. A product derivation tool based on modeldriven techniques and annotations. J.UCS: Journal of Universal Computer Science, 14(8):1344– 1367, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_8/a_product_
derivation_tool.

Choi:2013:TAM

[CKPK13] D. Choi, J. Kim, X. Piao, and P. Kim. Text analysis for monitoring personal information leakage on Twitter. J.UCS: Journal of Universal Computer Science, 19(16):2472–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
16/text_analysis_for_
monitoring.

Carlson:1995:CCM

[CL95] P. Carlson and V. Larralde. Combining concept mapping and adaptive advice to teach reading comprehension. J.UCS: Journal of Universal Computer Science, 1(3):156– 161, March 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
combining_concept_mapping_

                                        and_adaptive_advice_
                                        to_teach_reading_comprehension.

Calude:1997:MDI

[CL97] E. Calude and M. Lipponen. Minimal deterministic incomplete automata. J.UCS: Journal of Universal Computer Science, 3(11):1180– 1193, November 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_3_11/
                                        minimal_deterministic_
                                        incomplete_automata.

Clarke:2007:MCS

[CL07] E. M. Clarke and F. Lerda. Model checking: Software and beyond. J.UCS: Journal of Universal Computer Science, 13(5):639– 649, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        13_5/model_checking_
                                        software_and.

Chidlovskii:2008:SDN

[CL08] B. Chidlovskii and L. Lecerf. Stacked dependency networks for layout document structuring. J.UCS: Journal of Universal Computer Science, 14 (18):2998–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_18/stacked_dependency_
networks_for.

Clark:2005:CAI

[Cla05]  J. Clark.  Construc-

tive analysis of iterated rational functions. J.UCS: Journal of Universal Computer Science, 11(12):1904–1931, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
12/constructive_analysis_
of_iterated.

Choy:2004:SAK

[CLC04] S. Y. Choy, W. B. Lee, and C. F. Cheung. A systematic approach for knowledge audit analysis: Integration of knowledge inventory, mapping and knowledge flow analysis. J.UCS: Journal of Universal Computer Science, 10(6):674–682, June 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
6/a_systematic_approach_
for.

Chiu:2009:ISW

[CLC09] P.-H. Chiu, C.-C. Lo, and K.-M. Chao. Integrating Semantic Web and object-oriented programming for cooperative design. J.UCS: Journal of Universal Computer Science, 15(9):1970– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
9/integrating_semantic_
web_and.

Chen:2010:GFA

[CLCC10] W.-M. Chen, C.-H. Lo, H.-C. Chao, and C.-C. Chang. Gabor filter aided 3D ultra-sonography diagnosis system with WLAN transmission consideration. J.UCS: Journal of Universal Computer Science, 16(10):1327–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16_10/gabor_
filter_aided_3d.

Choi:2010:WCC

[CLM10] J. Choi, G. Lee, and J. Moon. Web context classification based on information quality factors. J.UCS: Journal of Universal Computer Science, 16(16):2232–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_

16/web_context_classification_
based.

Chen:2009:SSP

[CLVM09] Y. Chen, F. Liu, B. Vanschoenwinkel, and B. Manderick. Splice site prediction using support vector machines with contextsensitive kernel functions. J.UCS: Journal of Universal Computer Science, 15(13):2528–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
13/splice_site_prediction_
using.

Cools:1998:HSS

[CM98] R. Cools and B. Maerten.

A hybrid subdivision strategy for adaptive integration. J.UCS: Journal of Universal Computer Science, 4(5):486– 500, May 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_5/a_
hybrid_subdivision_strategy;
internal&sk=0C220489.

Clark:2000:ULA

[CM00] R. Clark and A. Moreira. Use of E-LOTOS in adding formality to UML. J.UCS: Journal of Universal Computer Science, 6(11):1071–1087, November 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_6_
                                        11/use_of_e_lotos.

Carvilhe:2003:OOA

[CM03] C. Carvilhe and M. A. Musicante. Objectoriented action semantics specifications. J.UCS: Journal of Universal Computer Science, 9(8): 910–934, August 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        9_8/object_oriented_
                                        action_semantics.

Colson:2007:PND

[CM07] L. Colson and D. Michel. Pedagogical natural deduction systems: the propositional case. J.UCS: Journal of Universal Computer Science, 13 (10):1396–1410, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        13_10/pedagogical_natural_
                                        deduction_systems.

Castellano:2009:WDS

[CM09] E. J. Castellano and L. Martínez. A Webdecision support system based on collaborative filtering for academic orientation. Case study of the Spanish secondary school. J.UCS: Journal of Universal Computer Science, 15(14):2786–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_15_14/a_web_
decision_support.

Chickerur:2011:CIR

[CM11] S. Chickerur and A. K.

M. Color image restoration using neural network model. J.UCS: Journal of Universal Computer Science, 17(1):107–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_17_1/color_
image_restoration_using.

Cansell:2001:DRD

[CMM01] D. Cansell, D. Mery, and S. Merz. Diagram refinements for the design of reactive systems. J.UCS: Journal of Universal Computer Science, 7(2):159–174, February 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_7_2/diagram_
refinements_for_the.

Carafoli:2016:DMM

[CMMP16] L. Carafoli, F. Mandreoli, R. Martoglia, and W. Penzo. A data management middleware for ITS services in smart cities. J.UCS: Journal of Universal Computer Science, 22(2):228– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_22_
                                        2/a_data_management_
                                        middleware.

Canal:2008:SA

[CMP08] C. Canal, J. M. Murillo, and P. Poizat. Software adaptation. J.UCS: Journal of Universal Computer Science, 14(13): 2107–??, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14;   http://www.jucs.
                                        org/jucs_14_13#;   http:
                                        //www.jucs.org/jucs_
                                        14_13/software_adaptation%
                                        20.

Calude:1994:JUC

[CMS94] C. Calude, H. Maurer, and A. Salomaa. Journal of Universal Computer Science. J.UCS: Journal of Universal Computer Science, 0(0):109– 117, November 15, 1994. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
journal_of_universal_
computer_science.

Cruz:2009:ARM

[CMSE09] C. C. P. Cruz, C. L. R.

Motta, F. M. Santoro, and M. Elia. Applying reputation mechanisms in communities of practice: a case study. J.UCS: Journal of Universal Computer Science, 15(9):1886– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
9/applying_reputation_
mechanisms_in.

Calinescu:2007:SPA

[CMZZ07] G. Calinescu, I. I. Mandoiu, A. Zelikovsky, and M. Zimand. Selected papers from the 1st ACIS International Workshop on Self-Assembling Wireless Networks. J.UCS: Journal of Universal Computer Science, 13(1):1–3, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13;
 http://www.jucs.org/
jucs_13_1#;   http://
www.jucs.org/jucs_13_
1/self_assembling_wireless_
networks.

Calude:1997:CON

[CN97] C. S. Calude and A. Nies. Chaitin Omega numbers and strong reducibilities. J.UCS: Journal of Universal Computer Science, 3(11):1162–1166, November 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_3_11/
                                        chaitin_omega_numbers_
                                        and.

Cao:2008:IMK

[CN08a] L. Cao and N. T. Nguyen. Intelligence metasynthesis and knowledge processing in intelligent systems. J.UCS: Journal of Universal Computer Science, 14(14):2256–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        14/intelligence_metasynthesis_
                                        and_knowledge.

Cao:2008:KPI

[CN08b] L. Cao and N. T. Nguyen. Knowledge processing in intelligent systems. J.UCS: Journal of Universal Computer Science, 14(14):2255–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.

org/jucs_14;   http:
//www.jucs.org/jucs_
14_14#;   http://www.
jucs.org/jucs_14_14/
knowledge_processing_
in_intelligent%20%20.

Cabodi:2004:ISB

[CNQ04] G. Cabodi, S. Nocco, and S. Quer. Improving SATbased bounded model checking by means of BDD-based approximate traversals. J.UCS: Journal of Universal Computer Science, 10(12): 1693, December 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_10_12/improving_
sat_based_bounded.

Cegarra-Navarro:2003:IKB

[CNRM03] J. G. Cegarra-Navarro and B. Rodrigo-Moya. Individual knowledge as a bridge between human and customer capital. J.UCS: Journal of Universal Computer Science, 9(12):1469–1486, December 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_9_12/individual_
knowledge_as_a.

Cao:2008:MMP

[CO08] L. Cao and Y. Ou. Market microstructure patterns powering trading and surveillance agents. J.UCS: Journal of Universal Computer Science, 14(14):2288–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_14/market_microstructure_
patterns_powering.

Cubo:2014:ASF

[COBP^{+}14] J. Cubo, G. Ortiz, J. Boubeta-Puig, H. Foster, and W. Lamersdorf. Adaptive services for the future Internet. J.UCS: Journal of Universal Computer Science, 20(8):1046–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_20;   http:
//www.jucs.org/jucs_
20_8/adaptive_services_
for_the.

Cook:2006:HMA

[Coo06]  D. J. Cook.  Health

monitoring and assistance to support aging in place. J.UCS: Journal of Universal Computer Science, 12(1):15– 29, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
1/health_monitoring_
and_assistance.

Costanza:2008:LRE

[Cos08]  P. Costanza.  Lisp:

Research and experience. J.UCS: Journal of Universal Computer Science, 14(20):3279–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14;
http://www.jucs.org/
jucs_14_20#;   http:
//www.jucs.org/jucs_
14_20/lisp_research_
and_experience.

Crossley:2001:FAG

[CP01] J. Crossley and I. Poernomo. Fred: An approach to generating real, correct, reusable programs from proofs. J.UCS: Journal of Universal Computer Science, 7(1):71–88, January 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
1/fred_an_approach_to.

Covino:2002:IRL

[CP02] E. Covino and G. Pani.

An implicit recursive language for the polynomial time-space complexity classes. J.UCS: Journal of Universal Computer Science, 8(1):75– 84, January 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_8_1/an_implicit_
recursive_language.

Cleva:2006:VCP

[CP06] J. M. Cleva and I. Pita. Verification of CRWL programs with rewriting logic. J.UCS: Journal of Universal Computer Science, 12(11):1594–1617, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
11/verification_of_crwl_
programs.

Cacheda:2015:IRR

[CP15] F. Cacheda and J. Parapar. Information retrieval and recommender systems. J.UCS: Journal of Universal Computer Science, 21(13):1706–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_21;   http:
//www.jucs.org/jucs_
21_13/information_retrieval_
and_recommender.

Caporaso:2000:ILT

[CPC00] S. Caporaso, G. Pani, and E. Covino. Incompleteness in linear time. J.UCS: Journal of Universal Computer Science, 6(12):1185–1196, December 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
12/incompleteness_in_
linear_time.

Colomo-Palacios:2011:UAG

[CPCLSAGC11] R. Colomo-Palacios, C. Casado-Lumbreras, P. Soto-Acosta, and A. Garcia-Crespo. Using the affect grid to measure emotions in software requirements engineering. J.UCS: Journal of Universal Computer Science, 17(9):1281– ??, ???? 2011. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_9/using_the_
affect_grid.

Colomo-Palacios:2012:HIC

[CPFSdAS12] R. Colomo-Palacios, E. Fernandes, M. Sabbagh, and A. de Amescua Seco. Human and intellectual capital management in the Cloud: Software vendor perspective. J.UCS: Journal of Universal Computer Science, 18(11): 1544–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_18_11/human_and_
                                        intellectual_capital.

Cheikhrouhou:2011:ESH

[CPHC11] N. Cheikhrouhou, M. Pouly, C. Huber, and A. Choudhary. An empirical study on human and information technology aspects in collaborative enterprise networks. J.UCS: Journal of Universal Computer Science, 17(2):203– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_17_
                                        2/an_empirical_study_
                                        on.

Cacchione:2015:PIE

[CPLPW15] A. Cacchione, E. Procter-Legg, S. Abbas Petersen, and M. Winter. A proposal for an integrated evaluation framework for mobile language learning: Lessons learned from SIMOLA — situated mobile language learning. J.UCS: Journal of Universal Computer Science, 21(10):1248–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_21_
                                        10/a_proposal_for_an.

Colomo-Palacios:2013:IBA

[CPMVG13] R. Colomo-Palacios, T. Moser, and R. Valencia-García. Industrial and business applications of Semantic Web technologies. J.UCS: Journal of Universal Computer Science, 19(13):1868–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_19;   http:
//www.jucs.org/jucs_
19_13/industrial_and_
business_applications.

Choppy:2001:KE

[CPR01] Christine Choppy, Pascal Poizat, and Jean-Claude Royer. The Korrigan environment. J.UCS: Journal of Universal Computer Science, 7(1):19–36, January 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
1/the_korrigan_environment.

Clavel:2006:IIT

[CPR06] M. Clavel, M. Palomino, and A. Riesco. Introducing the ITP tool: a tutorial. J.UCS: Journal of Universal Computer Science, 12(11):1618–1650, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
11/introducing_the_itp_
tool.

Casola:2005:RMS

[CPRT05] V. Casola, R. Preziosi, M. Rak, and L. Troiano. A reference model for security level evaluation: Policy and fuzzy techniques. J.UCS: Journal of Universal Computer Science, 11(1):150–174, January 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
1/a_reference_model_
for.

Covino:2007:CTC

[CPS07] E. Covino, G. Pani, and D. Scrimieri. Compiletime computation of polytime functions. J.UCS: Journal of Universal Computer Science, 13(4): 468–478, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_4/compile_time_computation_
of.

Colomo-Palacios:2012:SIG

[CPSAGPGC12] R. Colomo-Palacios, P. Soto-Acosta, F. J. García-Peñalvo, and Á. García-Crespo. A study of the impact of global software development in packaged software release planning. J.UCS: Journal of Universal Computer Science, 18(19):2646–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
19/a_study_of_the.

Cazanescu:2000:WIS

[CR00a] V. E. Cazanescu and G. Rosu. Weak inclusion systems: Part two. J.UCS: Journal of Universal Computer Science, 6(1):5–21, January 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
1/weak_inclusion_system_
part.

Ciobanu:2000:PCM

[CR00b] G. Ciobanu and M. Rotaru. A pi-calculus machine. J.UCS: Journal of Universal Computer Science, 6(1):39–59, January 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
1/a_pi_calculus_machine.

Carpineto:2004:EPC

[CR04] C. Carpineto and G. Romano. Exploiting the potential of concept lattices for information retrieval with CREDO. J.UCS: Journal of Universal Computer Science, 10(8):985–1013, August 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
10_8/exploiting_the_
potential_of.

Ciobanu:2007:ETA

[CR07] G. Ciobanu and S. Rudeanu. Equivalent transformations of automata by using behavioural automata. J.UCS: Journal of Universal Computer Science, 13(11):1540–1549, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
11/equivalent_transformations_
of_automata.

Cantoni:2012:OIR

[CR12] L. Cantoni and I. Rega. Outcomes of international research projects on technology applied to education. J.UCS: Journal of Universal Computer Science, 18(3):312– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18;

http://www.jucs.org/
jucs_18_3/rethinking_
education_in_the;   http:
//www.jucs.org/jucsrssfeed-
issue.

Cambruzzi:2015:DPR

[CRB15] W. Cambruzzi, S. J.

Rigo, and J. L. V. Barbosa. Dropout prediction and reduction in distance education courses with the learning analytics multitrail approach. J.UCS: Journal of Universal Computer Science, 21(1):23– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
1/dropout_prediction_
and_reduction.

Corchuelo:2004:BCS

[CRC04] R. Corchuelo and A. Ruiz-Cortés. Breakthroughs and challenges in software engineering. J.UCS: Journal of Universal Computer Science, 10(4):303– 305, April 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
http://www.jucs.org/
jucs_10_4/breakthroughs_
and_challenges_in.

Cresswell:2009:NDI

[Cre09] M. J. Cresswell. Nondenumerable infinitary modal logic. J.UCS: Journal of Universal Computer Science, 15(1):63– ??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_1/non_denumerable_
infinitary_modal.

Caeiro-Rodriguez:2005:CAM

[CRLNAR05] M. Caeiro-Rodríguez, M. Llamas-Nistal, and L. Anido-Rifón. From contents to activities: Modelling units of learning. J.UCS: Journal of Universal Computer Science, 11(9):1458– 1469, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
9/from_contents_to_activities.

Caeiro-Rodriguez:2007:SMF

[CRMLN^{+}07] M. Caeiro-Rodríguez, M. J. Marcelino, M. Llamas-Nistal, L. Anido-Rifón, and A. J. Mendes. Supporting the modeling of flexible educational units PoEML: a separation of concerns approach. J.UCS: Journal of Universal Computer Science, 13(7):980–990, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/

jucs_13_7/supporting_
the_modeling_of.

Calude:2000:ALC

[CS00] C. S. Calude and G. Stefanescu. Automata, logic, and computability: J.UCS special issue dedicated to Professor Sergiu Rudeanu Festschrift. J.UCS: Journal of Universal Computer Science, 6(1):1–2, January 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
1/automata_logic_and_
computability.

Cristofor:2002:FMP

[CS02] D. Cristofor and D. Simovici.

Finding median partitions using informationtheoretical-based genetic algorithms. J.UCS: Journal of Universal Computer Science, 8(2): 153–172, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_8_2/finding_
median_partitions_using.

Camhy:2003:CDB

[CS03] D. G. Camhy and R. Stubenrauch. A cross-disciplinary bibliography on visual languages for information sharing and archiving. J.UCS: Journal of Universal Computer Science, 9(4):369–396, April 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_9_4/a_cross_
disciplinary_bibliography.

CerschiSeceleanu:2004:SCI

[CS04] C. Cerschi Seceleanu and T. Seceleanu. Synchronization can improve reactive systems control and modularity. J.UCS: Journal of Universal Computer Science, 10(10):1429–1468, October 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
10/synchronization_can_
improve_reactive.

Coquand:2005:FTC

[CS05a] T. Coquand and B. Spitters. Formal topology and constructive mathematics: the Gelfand and Stone–Yosida representation theorems. J.UCS: Journal of Universal Computer Science, 11 (12):1932–1944, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_12/formal_topotoly_
and_constructive.

Cvejic:2005:IRL

[CS05b] N. Cvejic and T. Seppänen.

Increasing robustness of LSB audio steganography by reduced distortion LSB coding. J.UCS: Journal of Universal Computer Science, 11(1):56– 65, January 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_11_1/increasing_
robustness_of_lsb.

Chatterjee:2007:CSC

[CS07] S. Chatterjee and P. Sarkar.

Constant size ciphertext HIBE in the augmented selective-ID model and its extensions. J.UCS: Journal of Universal Computer Science, 13(10): 1367–1395, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_10/constant_size_
ciphertext_hibe.

Claycomb:2009:UCA

[CS09] W. Claycomb and D. Shin.

A user controlled approach for securing sensitive information in directory services. J.UCS: Journal of Universal Computer Science, 15 (15):2999–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        15_15/a_user_controlled_
                                        approach.

Camponogara:2010:MAR

[CS10] E. Camponogara and R. Boveto Shima. Mobile agent routing with time constraints: a resource constrained longest-path approach. J.UCS: Journal of Universal Computer Science, 16(3):372–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_16_3/mobile_
                                        agent_routing_with.

Christensen:2014:HAG

[CS14] I. Christensen and S. Schiaffino. A hybrid approach for group profiling in recommender systems. J.UCS: Journal of Universal Computer Science, 20(4):507– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_20_
                                        4/a_hybrid_approach_
                                        for.

Choi:2010:DAM

[CSA10] N. Choi, I.-Y. Song, and Y. An. Developing and analyzing the MP (Materialization Pattern) model for math educational standards. J.UCS: Journal of Universal Computer Science, 16(17):2394–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_17/developing_and_
analyzing_the.

Caceres:2015:MLA

[CSAC^{+}15] P. Cáceres, A. Sierra-Alonso, C. E. Cuesta, B. Vela, and J. M. Cavero. Modelling and linking accessibility data in the public bus network. J.UCS: Journal of Universal Computer Science, 21(6):777–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_6/modelling_
and_linking_accessibility.

Camara:2008:CRT

[CSC08] J. Cámara, G. Salaün, and C. Canal. Composition and run-time adaptation of mismatching behavioural interfaces. J.UCS: Journal of Universal Computer Science, 14(13):2182–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
13/composition_and_run_
time.

Chetcuti-Sperandio:1999:DMD

[CSF99] N. Chetcuti-Sperandio and L. Fariñas del Cerro. A decision method for duration calculus. J.UCS: Journal of Universal Computer Science, 5(11): 743–764, November 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_5_11/a_decision_
                                        method_for.

Caballero:2012:HTF

[CSFFM12] I. Caballero, L. E. Sánchez, A. Freitas, and E. Fernández-Medina. HC+: Towards a framework for improving processes in health organizations by means of security and data quality management. J.UCS: Journal of Universal Computer Science, 18(12):1703–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_18_12/hc%2B_
                                        towards_a_framework.

Chen:2008:CCS

[CSW^{+}08] L. Chen, W. Susilo, H. Wang, D. S. Wong, E. Dawson, X. Lai, M. Mambo, A. Miyaji, Y. Mu, D. Pointcheval, B. Preneel, and N. Smart. Cryptography in computer system security. J.UCS: Journal of Universal Computer Science, 14(3):314–317, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_14;   http:
//www.jucs.org/jucs_
14_3#;   http://www.
jucs.org/jucs_14_3/cryptography_
in_computer_system.

Calude:2002:ADQ

[CSY02a] C. S. Calude, K. Salomaa, and S. Yu. Additive distances and Quasi-Distances between words. J.UCS: Journal of Universal Computer Science, 8(2):141–152, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
2/additive_distances_
and_quasi.

Calude:2002:ATA

[CSY02b] C. S. Calude, K. Salomaa, and S. Yu. Advances and trends in automata and formal languages: a collection of papers in honour of the 60th birthday of Helmut Jürgensen — J.UCS special issue. J.UCS: Journal of Universal Computer Science, 8(2):117– 118, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/free;
                                         http://www.jucs.org/
                                        jucs_8_2/advances_and_
                                        trends_in.

Calude:2007:CRA

[CSZ07] C. S. Calude, G. Stefanescu, and M. Zimand. Combinatorics and related areas: a collection of papers in honour of the 65th birthday of Ioan Tomescu. J.UCS: Journal of Universal Computer Science, 13(11):1498–1500, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13;
                                         http://www.jucs.org/
                                        jucs_13_11#;   http:
                                        //www.jucs.org/jucs_
                                        13_11/combinatorics_
                                        and_related_areas.

Cavrak:2009:SSM

[ČSŽ09] I. Čavrak, A. Stranjak, and M. Žagar. SDL-MAS: a scenario modeling framework for multi-agent systems. J.UCS: Journal of Universal Computer Science, 15(4):898– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15_
                                        4/sdlmas_a_scenario_
                                        modeling.

Calude:1997:OEP

[CT97] C. S. Calude and I. Tomescu.

Optimum extendible prefix codes. J.UCS: Journal of Universal Computer Science, 3(11):1167– 1179, November 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_3_11/
optimum_extendible_prefix_
codes.

Chaudhuri:2004:MOD

[CT04] P. Chaudhuri and H. Thomson. A message-optimal distributed graph algorithm: Partial precedence constrained scheduling. J.UCS: Journal of Universal Computer Science, 10(2):106–119, February 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
2/a_message_optimal_
distributed.

Chen:2008:BUK

[CT08a] L. Chen and Q. Tang. Bilateral unknown key-share attacks in key agreement protocols. J.UCS: Journal of Universal Computer Science, 14(3):416– 440, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        3/bilateral_unknown_
                                        key_share.

Chu:2008:EOT

[CT08b] C.-K. Chu and W.-G. Tzeng. Efficient k-out-of-n oblivious transfer schemes. J.UCS: Journal of Universal Computer Science, 14(3):397– 415, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        3/efficient_k_out_of.

Choros:2016:PDG

[CT16] K. Choroś and J. Topolski. Parameterized and dynamic generation of an infinite virtual terrain with various biomes using extended Voronoi diagram. J.UCS: Journal of Universal Computer Science, 22(6):836– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_22_
                                        6/parameterized_and_
                                        dynamic_generation.

Chatel:2010:LNL

[CTM10] P. Châtel, I. Truck, and J. Malenfant. LCP-Nets: a linguistic approach for non-functional preferences in a semantic SOA environment.

J.UCS: Journal of Universal Computer Science, 16(1):198–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_1/lcp_nets_a_
linguistic.

Culik:1999:GWF

[Cv99] K. Culik II and P. C. von Rosenberg. Generalized weighted finite automata based image compression. J.UCS: Journal of Universal Computer Science, 5(4):227–242, April 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_5_4/generalized_
weighted_finite_automata.

Camarao:2007:OCW

[CVFN07] C. Camarão, C. Vasconcellos, L. Figueiredo, and J. Nicola. Open and closed worlds for overloading: a definition and support for coexistence. J.UCS: Journal of Universal Computer Science, 13(6):874–890, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_6/open_and_closed_
worlds.

Culik:1997:CSL

[CVK97] Karel Culik II, Vladimir Valenta, and Jarkko Kari. Compression of silhouette-like images based on WFA. J.UCS: Journal of Universal Computer Science, 3(10):1100–1113, October 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_3_10/
compression_of_silhouette_
like.

Cambronero:2011:DGW

[CVM11] M. E. Cambronero, V. Valero, and E. Martínez. Design and generation of Web services choreographies with time constraints. J.UCS: Journal of Universal Computer Science, 17(13):1800–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
13/design_and_generation_
of.

Csuhaj-Varju:1995:CTE

[CVPS95] Erzsebet Csuhaj-Varju, Gheorghe Păun, and Arto Salomaa. Conditional tabled eco-grammar systems versus (E)T0L systems. J.UCS: Journal of Universal Computer Science, 1(5):252– 268, May 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
conditional_tabled_eco_
grammar_systems.

Carvalho:2011:LLS

[CVSM11] G. Carvalho, A. S. Vivacqua, J. M. Souza, and S. P. J. Medeiros. LaSca: a large scale group decision support system. J.UCS: Journal of Universal Computer Science, 17(2):261– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
2/lasca_a_large_scale.

Codrescu:2000:DST

[CW00] L. Codrescu and D. S.

Wills. On dynamic speculative thread partitioning and the MEM-slicing algorithm. J.UCS: Journal of Universal Computer Science, 6(10):908– 927, October 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_6_10/on_dynamic_
speculative_thread.

Chen:2009:HWR

[CW09] Y.-C. Chen and J.-S.

Wang. A Hammerstein– Wiener recurrent neural network with frequencydomain eigensystem realization algorithm for unknown system identification. J.UCS: Journal of Universal Computer Science, 15(13):2547–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_13/a_hammerstein_
wiener_recurrent.

Chmaj:2012:DSP

[CW12] G. Chmaj and K. Walkowiak. Decision strategies for a P2P computing system. J.UCS: Journal of Universal Computer Science, 18(5):599– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
5/decision_strategies_
for_a.

Chen:2015:SUE

[CWT^{+}15] C.-M. Chen, T.-Y. Wu, R. Tso, M. Mambo, and M.-E. Wu. On the security of a user equipment registration procedure in femtocell-enabled networks. J.UCS: Journal of Universal Computer Science, 21(3):406– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
3/on_the_security_of.

Chiang:2011:ISA

[CWTT11] T.-A. Chiang, C.-Y. Wu, C. V. Trappey, and A. J. C. Trappey. An intelligent system for automated binary knowledge document classification and content analysis. J.UCS: Journal of Universal Computer Science, 17(14):1991–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_14/an_intelligent_
system_for.

Cai:2012:COB

[CXB12] H. Cai, B. Xu, and F. Bu. A conceptual ontology-based resource meta-model towards business-driven information system implementation. J.UCS: Journal of Universal Computer Science, 18(17): 2493–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_17/a_conceptual_ontology_
based.

Chang:2011:OBA

[CYL11] Y.-S. Chang, C.-T. Yang, and Y.-C. Luo. An ontology based agent generation for information retrieval on cloud environment. J.UCS: Journal of Universal Computer Science, 17(8):1135– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
8/an_ontology_based_
agent.

Dhir:2013:EEU

[DA13] A. Dhir and A. Alsumait. Examining the educational user interface, technology and pedagogy for Arabic speaking children in Kuwait. J.UCS: Journal of Universal Computer Science, 19(7):1003–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_7/examining_
the_educational_user.

Dias:2003:OKU

[DAd03] Márcio Greyk Batista Dias, Nicolas Anquetil, and Käthia Maríal de Oliveira. Organizing the knowledge used in software maintenance. J.UCS: Journal of Universal Computer Science, 9(7):641–658, July 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_9_7/organizing_
the_knowledge_used.

Formiga:2007:NAC

[dAFL07] A. de Araújo Formiga and R. Dueire Lins. A new architecture for concurrent lazy cyclic reference counting on multi-processor systems. J.UCS: Journal of Universal Computer Science, 13(6):817–829, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_13_6/a_new_
architecture_for.

Dhir:2013:CSU

[DAk13] A. Dhir and M. Alkahtani. A case study on user experience (UX) evaluation of mobile augmented reality prototypes. J.UCS: Journal of Universal Computer Science, 19(8):1175– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
8/a_case_study_on.

deAlmeida:2013:SCA

[dAO13] E. Santana de Almeida and F. Oquendo. Software components, architectures and reuse modeling, customization and evaluation. J.UCS: Journal of Universal Computer Science, 19(2):183–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_19;   http:
//www.jucs.org/jucs_
19_2/software_components_
architectures_and.

Goncalves:2011:LMT

[dARGSB11] J. C. de A. R. Gonçalves, F. M. Santoro, and F. Araujo Baião. Let me tell you a story — on how to build process models. J.UCS: Journal of Universal Computer Science, 17(2):276– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
2/let_me_tell_you.

DantasdeSouza:2003:DAJ

[DB03] A. D. Dantas de Souza and P. Borba. Developing adaptive J2ME applications using AspectJ. J.UCS: Journal of Universal Computer Science, 9(8):935–955, August 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_9_8/developing_
adaptive_j2me_applications.

Deep:2012:SED

[DB12] K. Deep and J. C.

Bansal. Solving economic dispatch problems with valve-point effects using particle swarm optimization. J.UCS: Journal of Universal Computer Science, 18(13):1842–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
13/solving_economic_
dispatch_problems.

Duque:2012:AAP

[DBAB12] R. Duque, L. Bollen, A. Anjewierden, and C. Bravo. Automating the analysis of problemsolving activities in learning environments: the Co-Lab case study. J.UCS: Journal of Universal Computer Science, 18 (10):1279–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_10/automating_the_
analysis_of.

Dhir:2013:TCT

[DBB13] A. Dhir, K. Buragga, and A. A. Boreqqah. Tweeters on campus: Twitter a learning tool in classroom? J.UCS: Journal of Universal Computer Science, 19(5):672– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_19_
                                        5/tweeters_on_campus_
                                        twitter.

Dao:2008:NVF

[DBBS08] H. T. Dao, A. Bazinet, R. Berthier, and B. Shneiderman. NASDAQ velocity and forces: An interactive visualization of activity and change. J.UCS: Journal of Universal Computer Science, 14(9):1391–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_9/nasdaq_velocity_
                                        and_forces.

deOliveira:2004:MEE

[dBdd04] A. A. de Oliveira, T. H. Braga, M. de Almeida Maia, and R. da Silva Bigonha. MetaJ: An extensible environment for metaprogramming in Java. J.UCS: Journal of Universal Computer Science, 10(7):872–891, July 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_10_7/metaj_
                                        an_extensible_environment.

daCruz:2011:VAQ

[dCH11] D. da Cruz and P. R.

Henriques. Visualizing and analyzing the quality of XML documents. J.UCS: Journal of Universal Computer Science, 17(1):126–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_1/visualizing_and_
analyzing_the.

Perez:2007:HIE

[dCPUH^{+}07] M. del Carmen Peréz, J. Ureña, Á. Hernández, C. De Marziani, A. Jiménez, and W. P. Marnane. Hardware implementation of an efficient correlator for interleaved complementary sets of sequences. J.UCS: Journal of Universal Computer Science, 13(3):388– 406, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
3/hardware_implementation_
of_an.

DiIorio:2007:VLA

[DCR^{+}07] V. O. Di Iorio, D. P.

Coura, L. V. Reis, M. Oikawa, and C. R. Junior. A visual language for animated simulation. J.UCS: Journal of Universal Computer Science, 13(6):767– 785, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_
                                        6/a_visual_language_
                                        for.

DiLecce:2009:FLC

[DCS09] V. Di Lecce, M. Calabrese, and D. Soldo. Fingerprinting lexical contexts over the Web. J.UCS: Journal of Universal Computer Science, 15(4):805–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        15_4/fingerprinting_
                                        lexical_contexts_over.

deCastro:2012:SOD

[dCVM12] V. de Castro, J. M. Vara, and E. Marcos. Service-oriented development of Web information systems. J.UCS: Journal of Universal Computer Science, 18(17):2474–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_18_
                                        17/service_oriented_
                                        development_of.

Dias:2013:BIL

[DD13] S. B. Dias and J. A. Diniz. From blended to inclusive learning: Accessibility, profiles, openness, and higher education. J.UCS: Journal of Universal Computer Science, 19(18):2722–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_19_18/from_
blended_to_inclusive.

Dexter:1997:GAS

[DDG97] S. Dexter, P. Doyle, and Y. Gurevich. Gurevich abstract state machines and Schoenhage storage modification machines. J.UCS: Journal of Universal Computer Science, 3(4):279– 303, April 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs_3_4/schoenhage_
machines;   http://www.
jucs.org/jucs_3_4/gurevich_
abstract_state_machines;
 internal&sk=05460486.

Dolog:2011:ROL

[DDJ^{+}11] P. Dolog, F. Durao, K. Jahn, Y. Lin, and D. Kjaersgaard Peitersen. Recommending open linked data in creativity sessions using Web portals with collaborative real time environment. J.UCS: Journal of Universal Computer Science, 17(12):1690–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_17_
                                        12/recommending_open_
                                        linked_data.

Davies:2004:OOB

[DDS04] J. Davies, A. Duke, and Y. Sure. OntoShare — an ontology-based knowledge sharing system for virtual communities of practice. J.UCS: Journal of Universal Computer Science, 10(3):262–283, March 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_10_3/ontoshare_
                                        an_ontology_based.

Daley:2010:OCL

[DDS10] M. Daley, M. Domaratzki, and K. Salomaa. Orthogonal concatenation: Language equations and state complexity. J.UCS: Journal of Universal Computer Science, 16(5):653– ??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_16_5/orthogonal_
                                        concatenation_language_
                                        equations.

Dodero:2005:IOC

[DDSS05] J. M. Dodero, P. Díaz, A. Sarasa, and I. Sarasa. Integrating ontologies into the collaborative authoring of learning objects. J.UCS: Journal of Universal Computer Science, 11(9):1568–1576, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
9/integrating_ontologies_
into_the.

DeBra:1996:THH

[De 96] P. M. E. De Bra. Teaching hypertext and hypermedia through the Web. J.UCS: Journal of Universal Computer Science, 2(12):797–804, December 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
12/teaching_hypertext_
and_hypermedia.

deDinechin:2000:PRF

[de 00] F. de Dinechin. The price of routing in FP-GAs. J.UCS: Journal of Universal Computer Science, 6(2):227– 239, February 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
2/the_price_of_routing.

Daumas:1999:DFP

[DF99] M. Daumas and C. Finot.

Division of floating point expansions with an application to the computation of a determinant. J.UCS: Journal of Universal Computer Science, 5(6):323–338, June 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_5_6/division_of_
floating_point.

Diaconescu:2000:BCO

[DF00] R. Diaconescu and K. Futatsugi. Behavioural coherence in object-oriented algebraic specification. J.UCS: Journal of Universal Computer Science, 6(1):74–96, January 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_6_1/behavioural_
coherence_in_object.

DiSanto:2005:TCA

[DF05] M. Di Santo and F. Frattolillo. Time costs in actor computations. J.UCS: Journal of Universal Computer Science, 11(6):850–873, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_11_6/time_costs_
in_actor.

Bulcao-Neto:2014:CRM

[dFBNGS^{+}14] R. de Freitas Bulcão-Neto, J. A. Camacho Guerrero, P. Schor, A. Stanquini Lopes, M. Branquinho Dutra, and A. Alaniz Macedo. Capturing and relating multilingual clinical cases. J.UCS: Journal of Universal Computer Science, 20(9):1154– ??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_20_9/capturing_
and_relating_multilingual.

deFreitas:2007:LPV

[dFCC07] L. de Freitas, M. Caraciolo, and M. Cornélio. Logic programming for verification of objectoriented programming law conditions. J.UCS: Journal of Universal Computer Science, 13(6):721– 736, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
6/logic_programming_
for_verification.

Escrig:2006:PEG

[dFER06] D. de Frutos Escrig and C. G. Rodríguez. Process equivalences as global bisimulations. J.UCS: Journal of Universal Computer Science, 12 (11):1521–1550, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        12_11/process_equivalences_
                                        as_global.

Dunne:2000:JUS

[DG00] P. E. Dunne and A. Gibbons. J.UCS special issue on BCTCS. J.UCS: Journal of Universal Computer Science, 6(4):405– 406, April 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        6_4/special_issue_on_
                                        bctcs.

Diaconescu:2007:FSM

[DG07] D. Diaconescu and G. Georgescu. On the forcing semantics for monoidal t-norm based logic. J.UCS: Journal of Universal Computer Science, 13(11):1550–1572, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        13_11/on_the_forcing_
                                        semantics.

deGrancy:2015:AMV

[dG15] G. Senarclens de Grancy. An adaptive metaheuristic for vehicle routing problems with time windows and multiple service workers. J.UCS: Journal of Universal Computer Science, 21(9):1143– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
9/an_adaptive_metaheuristic_
for.

Duque:2008:DTD

[DGBM08] R. Duque, J. Gallardo, C. Bravo, and A. J. Mendes. Defining tasks, domains and conversational acts in CSCW systems: the SPACE-DESIGN case study. J.UCS: Journal of Universal Computer Science, 14(9):1463–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_9/defining_tasks_
domains_and.

Dietinger:1999:DBL

[DGK^{+}99] T. Dietinger, Ch. Gütl, B. Knögler, D. Neussl, and K. Schmaranz. Dynamic background libraries — new developments in distance education using HIKS. J.UCS: Journal of Universal Computer Science, 5(1): 2–10, January 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_

                                        1/dynamic_background_
                                        libraries_new.

Dugat:2003:ARO

[DGL03] V. Dugat, P. Gambarotto, and Y. Larvor. Abstract representation of object and structural symmetries detection. J.UCS: Journal of Universal Computer Science, 9(9):1008– 1029, September 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        9/abstract_representation_
                                        of_object.

Ding:2012:NRR

[DGlS12] L. Ding, J. Guan, and W. l. Sun. New results of related-key attacks on all Py-family of stream ciphers. J.UCS: Journal of Universal Computer Science, 18(12): 1741–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        18_12/new_results_of_
                                        related.

Dhir:2013:RIH

[DGN13] A. Dhir, N. M. Gahwaji, and G. Nyman. The role of the iPad in the hands of the learner. J.UCS: Journal of Universal Computer Science, 19(5):706– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
5/the_role_of_the.

deGroot:2000:ALC

[dH00] A. de Groot and J. Hooman.

Analyzing the light control system with PVS. J.UCS: Journal of Universal Computer Science, 6(7):621–649, July 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
7/analyzing_the_light_
control.

deAmo:2004:IMD

[dH04] S. de Amo and M. Halfeld Ferrari Alves. Incremental maintenance of data warehouses based on past temporal logic operators. J.UCS: Journal of Universal Computer Science, 10(9):1035–1064, September 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
9/incremental_maintenance_
of_data.

Danner:2010:TCI

[DH10a] P. Danner and D. Hein.

A trusted computing identity collation protocol to simplify deployment of new disaster response devices. J.UCS: Journal of Universal Computer Science, 16(9):1139–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_9/a_trusted_
computing_identity.

Decraene:2010:MSC

[DH10b] J. Decraene and T. Hinze. A multidisciplinary survey of computational techniques for the modelling, simulation and analysis of biochemical networks. J.UCS: Journal of Universal Computer Science, 16(9):1152– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
/a_multidisciplinary_
survey_of.

Dong:2011:OOB

[DHC11] H. Dong, F. Khadeer Hussain, and E. Chang. ORPMS: An ontologybased real-time project monitoring system in the cloud. J.UCS: Journal of Universal Computer Science, 17(8):1161– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
8/orpms_an_ontology_
based.

Duval:1998:BHO

[DHO98] E. Duval, K. Hendrikx, and H. Olivié. Building hypermedia with objects and sets. J.UCS: Journal of Universal Computer Science, 4(5):501– 521, May 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_5/building_
hypermedia_with_objects;
internal&sk=0C220489.

Dubois:2003:FUT

[DHP03] Didier Dubois, Allel HadjAli, and Henri Prade. Fuzziness and uncertainty in temporal reasoning. J.UCS: Journal of Universal Computer Science, 9(9):1168–1194, September 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
9/fuzziness_and_uncertainty_
in.

Lopez-de-Ipina:2013:TSC

[dIB13] D. López de Ipiña and J. Bravo. Towards sustainable computing through ambient intelligence. J.UCS: Journal of Universal Computer Science, 19(17):2486–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_19;   http:
                                        //www.jucs.org/jucs_
                                        19_17/towards_sustainable_
                                        computing_through.

Lopez-de-Ipina:2010:AAL

[dIdSGZ10] D. López de Ipiña, I. Díaz de Sarralde, and J. García-Zubia. An ambient assisted living platform integrating RFID data-on-tag care annotations and Twitter. J.UCS: Journal of Universal Computer Science, 16(12):1521–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_16_
                                        12/an_ambient_assisted_
                                        living.

Diem:1996:PTU

[Die96] R. Diem. Preparing teachers to use and apply technology: New models for a new era. J.UCS: Journal of Universal Computer Science, 2(10):711– 716, October 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_2_10/preparing_
                                        teachers_to_use.

Dietrich:2010:RPC

[Die10]  K. Dietrich.  On re-

liable platform configuration change reporting mechanisms for trusted computing enabled platforms. J.UCS: Journal of Universal Computer Science, 16(4):507– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
4/on_reliable_platform_
configuration.

Drigas:2014:IML

[DIKL14] A. S. Drigas, R.-E. Ioannidou, G. Kokkalia, and M. Lytras. ICTs, mobile learning and social media to enhance learning for attention difficulties. J.UCS: Journal of Universal Computer Science, 20(10):1499–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_20_10/icts_
mobile_learning_and.

Dinneen:1997:TMM

[Din97]  M. J. Dinneen.  Too

many minor order obstructions. J.UCS: Journal of Universal Computer Science, 3(11):1199– 1206, November 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:

                                        8000/jucs/jucs_3_11/
                                        too_many_minor_order.

Dittrich:2002:MMR

[Dit02] G. Dittrich. Mental models to represent dynamics — using the example “factorial”. J.UCS: Journal of Universal Computer Science, 8(10):957–964, October 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        10/mental_models_to_
                                        represent.

deIpina:2006:ERF

[dIVG^{+}06] D. López de Ipiña, J. I. Vázquez, D. García, J. Fernández, I. García, D. Sainz, and A. Almeida. EMI^{2}lets: a reflective framework for enabling AmI. J.UCS: Journal of Universal Computer Science, 12(3):297–314, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_12_3/emi2lets_a_
                                        reflective_framework.

Duong:2009:CAO

[DJJN09] T. H. Duong, G. S. Jo, J. J. Jung, and N. T. Nguyen. Complexity analysis of ontology integration methodologies: a comparative study. J.UCS: Journal of Universal Computer Science, 15(4):877–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_4/complexity_
analysis_of_ontology.

Dietze:2013:SSI

[DKD^{+}13] S. Dietze, E. Kaldoudi, N. Dovrolis, D. Giordano, C. Spampinato, M. Hendrix, A. Protopsaltis, D. Taibi, and H. Q. Yu. Socio-semantic integration of educational resources — the case of the mEducator Project. J.UCS: Journal of Universal Computer Science, 19(11):1543–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_19_11/socio_
semantic_integration_
of.

Danylevych:2010:SNM

[DKL10] O. Danylevych, D. Karastoyanova, and F. Leymann. Service networks modelling: An SOA & BPM standpoint. J.UCS: Journal of Universal Computer Science, 16(13):1668–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_

                                        13/service_networks_
                                        modelling_an.

Dreher:2004:WWE

[DKM04] H. Dreher, H. Krottmaier, and H. Maurer. What we expect from digital libraries. J.UCS: Journal of Universal Computer Science, 10(9):1110–1122, September 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        10_9/what_we_expect_
                                        from.

deVasconcelos:2003:OMI

[dKR03] J. B. de Vasconcelos, C. Kimble, and A. Rocha. Organisational memory information systems an example of a group memory system for the management of group competencies. J.UCS: Journal of Universal Computer Science, 9(12):1410– 1427, December 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        9_12/organisational_
                                        memory_information_systems.

Dunne:1999:ACP

[DL99] P. E. Dunne and P. H. Leng. The average case performance of an algorithm for demand-driven evaluation of Boolean formulae. J.UCS: Journal of Universal Computer Science, 5(5):288– 306, May 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
5_5/the_average_case_
performance.

deCarvalho:2003:HPP

[dL03] F. H. de Carvalho, Jr. and R. D. Lins. Haskell#: Parallel programming made simple and efficient. J.UCS: Journal of Universal Computer Science, 9(8): 776–794, August 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
9_8/haskell_parallel_
programming_made.

Dix:2015:LAA

[DL15] A. Dix and J. Leavesley.

Learning analytics for the academic: An action perspective. J.UCS: Journal of Universal Computer Science, 21(1):48– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
1/learning_analytics_
for_the.

delaEncina:2009:DSB

[dlELR09] A. de la Encina, L. Llana, and F. Rubio. A debugging system based on natural semantics. J.UCS: Journal of Universal Computer Science, 15 (14):2836–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_14/a_debugging_system_
based.

de-la-Fuente-Valentin:2015:VAM

[dlFVPHB15] L. de-la Fuente-Valentín, A. Pardo, F. López Hernández, and D. Burgos. A visual analytics method for score estimation in learning courses. J.UCS: Journal of Universal Computer Science, 21(1):134– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
1/a_visual_analytics_
method.

deLope:2008:CMB

[dLH08] L. Rodríguez de Lope and K. D. Hackbarth. Cost model for bitstream access services with QoS parameters. J.UCS: Journal of Universal Computer Science, 14(5):653–672, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/

jucs_14_5/cost_model_
for_bitstream.

Danylenko:2014:DAI

[DLL14] A. Danylenko, J. Lundberg, and W. Löwe. Decisions: Algebra, implementation, and first experiments. J.UCS: Journal of Universal Computer Science, 20(9):1174– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
9/decisions_algebra_
implementation_and.

Datcu:2016:CGS

[DLL16] D. Datcu, S. Lukosch, and H. Lukosch. A collaborative game to study presence and situational awareness in a physical and an augmented reality environment. J.UCS: Journal of Universal Computer Science, 22(2):247– ??, ???? 2016. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
22_2/a_collaborative_
game_to.

Martinez-Villasenor:2013:EPC

[dLMVGM13] M. de Lourdes Martinez-Villaseñor and M. González-Mendoza. An enhanced process of concept alignment for dealing with overweight and obesity. J.UCS: Journal of Universal Computer Science, 19(9):1315– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
9/an_enhanced_process_
of.

Ding:1997:LMS

[DLR97] C. Ding, T. Laihonen, and A. Renvall. Linear multisecret-sharing schemes and error-correcting codes. J.UCS: Journal of Universal Computer Science, 3(9):1023–1036, September 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_3_9/linear_
multisecret_sharing.

Dodis:2008:OFE

[DLY08] Y. Dodis, P. J. Lee, and D. H. Yum. Optimistic fair exchange in a multiuser setting. J.UCS: Journal of Universal Computer Science, 14(3):318– 346, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_3/optimistic_fair_
exchange_in.

Denman-Maier:2004:IFW

[DM04] E. Denman-Maier. Intercultural factors in Webbased training systems. J.UCS: Journal of Universal Computer Science, 10(1):90–104, January 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
1/intercultural_factors_
in_web.

Damavandi:2007:CEC

[DM07] Y. Baleghi Damavandi and K. Mohammadi. Coevolution for communication: An EHW approach. J.UCS: Journal of Universal Computer Science, 13(9):1300– 1308, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_9/coevolution_for_
communication_an.

Dvorak:2010:DAA

[DM10] V. Dvorák and P. Mikusek.

Design of arbiters and allocators based on multiterminal BDDs. J.UCS: Journal of Universal Computer Science, 16 (14):1826–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/

                                        jucs_16_14/design_of_
                                        arbiters_and.

Bezerra:2015:CFD

[dMBHR15] J. de Melo Bezerra, C. Massaki Hirata, and D. Randall. A conceptual framework to define incentive mechanisms for virtual communities. J.UCS: Journal of Universal Computer Science, 21(9):1107– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_21_
                                        9/a_conceptual_framework_
                                        to.

Domingos:2014:ITA

[DMCM14] D. Domingos, F. Martins, C. Cândido, and R. Martinho. Internet of Things aware WS-BPEL business processes context variables and expected exceptions. J.UCS: Journal of Universal Computer Science, 20(8):1109– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_20_
                                        8/internet_of_things_
                                        aware.

Duric:2007:IPS

[DMG07] Z. Durić, O. Marić, and D. Gasević. Internet payment system: a new payment system for Internet transactions. J.UCS: Journal of Universal Computer Science, 13(4):479– 503, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
4/internet_payment_system_
a.

Dragoi:2007:ANE

[DMM07] C. Dragoi, F. Manea, and V. Mitrana. Accepting networks of evolutionary processors with filtered connections. J.UCS: Journal of Universal Computer Science, 13 (11):1598–1614, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_13_11/accepting_
networks_of_evolutionary.

Daumas:1995:MRR

[DMMM95] Marc Daumas, Christophe Mazenc, Xavier Merrheim, and Jean-Michel Muller. Modular range reduction: a new algorithm for fast and accurate computation of the elementary functions. J.UCS: Journal of Universal Computer Science, 1(3):162–175, March 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_1_3/modular_
                                        range_reduction.

Delis:2005:PEC

[DMS05] V. Delis, C. Makris, and S. Sioutas. A provably efficient computational model for approximate spatiotemporal retrieval. J.UCS: Journal of Universal Computer Science, 11(6):830–849, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_11_6/a_provably_
                                        efficient_computational.

deMello:2014:VSP

[dMTS^{+}14] R. Maiani de Mello, E. Nogueira Teixeira, M. Schots, C. M. Lima Werner, and G. Horta Travassos. Verification of software product line artefacts: A checklist to support feature model inspections. J.UCS: Journal of Universal Computer Science, 20(5):720– ??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        20_5/verification_of_
                                        software_product.

Datta:2001:TVU

[DO01] A. Datta and T. Ottmann. Towards a virtual university. J.UCS: Journal of Universal Computer Science, 7(10):870–885, October 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
10/towards_a_virtual_
university.

deOliveira:2014:DVF

[dOBGH^{+}14] R. Pereira de Oliveira, D. Blanes, J. Gonzalez-Huerta, E. Insfran, S. Abrahão, S. Cohen, and E. Santana de Almeida. Defining and validating a feature-driven requirements engineering approach. J.UCS: Journal of Universal Computer Science, 20(5):666– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
5/defining_and_validating_
a.

deOliveira:2007:CCM

[dOLC^{+}07] F. de Oliveira, Jr., R. Lima, M. Cornelio, S. Soares, P. Maciel, R. Barreto, M. Oliveira, Jr., and E. Tavares. CML: C modeling language. J.UCS: Journal of Universal Computer Science, 13(6):682–700, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        13_6/cml_c_modeling_
                                        language.

Domik:2001:GFC

[Dom01] G. Domik. Glimpses into the future of computer science education. J.UCS: Journal of Universal Computer Science, 7(5): 366–378, May 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_7_
                                        5/glimpses_into_the_
                                        future.

Doder:2010:AFO

[DOM10] D. Doder, Z. Ognjanovi, and Z. Markovi. An axiomatization of a firstorder branching time temporal logic. J.UCS: Journal of Universal Computer Science, 16(11): 1439–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        16_11/an_axiomatization_
                                        of_a.

Mattos:2008:CAB

[dOMdAL^{+}08] G. de Oliveira Mattos, A. de Araújo, R. Dueire Lins, F. H. de Carvalho Júnior, and F. M. Junqueira Martins. Comparative aspects between the cluster and Grid implementations of Big-Batch. J.UCS: Journal of Universal Computer Science, 14(18):3031–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
18/comparative_aspects_
between_the.

Duval:1995:HEH

[DOOJ95] E. Duval, H. Olivie, P. O’Hanlon, and D. Jameson. HOME: An environment for hypermedia objects. J.UCS: Journal of Universal Computer Science, 1(5):269– 291, May 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
home_an_environment_
for_hypermedia_objects.

Doran:1995:SCD

[Dor95]  R. W. Doran.  Spe-

cial cases of division. J.UCS: Journal of Universal Computer Science, 1(3):176–194, March 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_1_3/special_
cases_of_division.

Doran:2007:GC

[Dor07] R. W. Doran. The Gray code. J.UCS: Journal of Universal Computer Science, 13(11):1573–1597, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_
                                        11/the_gray_code.

Duval:1995:CH

[DOS95] E. Duval, H. Olivie, and N. Scherbakov. Contained hypermedia. J.UCS: Journal of Universal Computer Science, 1(10): 687–705, October 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        contained_hypermedia.

Dassow:1999:PMC

[DP99] J. Dassow and G. Paun. On the power of membrane computing. J.UCS: Journal of Universal Computer Science, 5(2): 33–51, February 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_5_
                                        2/on_the_power_of.

Paule-Ruiz:2015:PRI

[dPPRRGSSPP15] M. del Puerto Paule-Ruiz, M. Riestra-Gonzalez, M. Sánchez-Santillan, and J. R. Pérez-Pérez. The procrastination related indicators in e-learning platforms. J.UCS: Journal of Universal Computer Science, 21(1):7–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
1/the_procrastination_
related_indicators.

Salas-Zarate:2016:SCS

[dPSZPVLR^{+}16] M. del Pilar Salas-Zarate, M. A. Paredes-Valverde, J. Limon-Romero, D. Tlapa, and Y. Baez-Lopez. Sentiment classification of Spanish reviews: An approach based on feature selection and machine learning methods. J.UCS: Journal of Universal Computer Science, 22(5):691–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_22_5/sentiment_
classification_of_spanish.

Dimitriadis:2008:GIA

[DPZ08] Y. A. Dimitriadis, J. A.

Pino, and I. Zigurs. Groupware: Issues and applications. J.UCS: Journal of Universal Computer Science, 14(1): 1–3, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14;
 http://www.jucs.org/
jucs_14_1#;   http://
www.jucs.org/jucs_14_

                                        1/groupware_issues_and_
                                        applications.

Diaz:2004:PWC

[DR04] O. Díaz and J. J. Rodríguez. Portlets as Web components: an introduction. J.UCS: Journal of Universal Computer Science, 10(4):454– 472, April 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        10_4/portlets_as_web_
                                        components.

dAmorim:2005:ESS

[dR05] M. d’Amorim and G. Rosu. An equational specification for the Scheme language. J.UCS: Journal of Universal Computer Science, 11(7):1327–1348, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_11_
                                        7/an_equational_specification_
                                        for.

Dietrich:2010:POD

[DR10] K. Dietrich and F. Röck. Performance optimizations for DAA signatures on Java enabled platforms. J.UCS: Journal of Universal Computer Science, 16(4):519– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
4/performance_optimizations_
for_daa.

Decker:2004:PPI

[DRA^{+}04] B. Decker, J. Rech, K.- D. Althoff, A. Klotz, E. Leopold, and A. Voss. Participative process introduction: Three case studies from the indiGo project. J.UCS: Journal of Universal Computer Science, 10(3):186– 204, March 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
10_3/_participative_
process_introduction.

deMoura:2004:CL

[dRI04] A. L. de Moura, N. Rodriguez, and R. Ierusalimschy. Coroutines in Lua. J.UCS: Journal of Universal Computer Science, 10(7):910–925, July 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
7/coroutines_in_lua.

Droschl:2004:CPI

[Dro04] G. Droschl. Communities of practice: An integrated technology perspective. J.UCS: Journal of Universal Computer Science, 10(3):284– 293, March 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
10_3/communities_of_
practice_an.

Diaz:2007:RTA

[DRRGdP07] J. Díaz, E. Ros, R. Rodriguez-Gomez, and B. del Pino. Real-time architecture for robust motion estimation under varying illumination conditions. J.UCS: Journal of Universal Computer Science, 13(3):363–376, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_13_3/real_time_
architecture_for.

Dominguez:2006:MIC

[DRS06] C. Domínguez, J. Rubio, and F. Sergeraert. Modeling inheritance as coercion in the Kenzo system. J.UCS: Journal of Universal Computer Science, 12(12):1701–1730, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_12_12/modeling_
inheritance_as_coercion.

Drusinsky:2006:LMM

[Dru06]  D. Drusinsky.  On-

line monitoring of metric temporal logic with time-series constraints using alternating finite automata. J.UCS: Journal of Universal Computer Science, 12(5):482– 498, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
5/on_line_monitoring_
of.

Drusinsky:2012:BTP

[Dru12] D. Drusinsky. Behavioral and temporal pattern detection within financial data with hidden information. J.UCS: Journal of Universal Computer Science, 18(14): 1950–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_14/behavioral_and_
temporal_pattern.

Drusinsky:2013:BTR

[Dru13] D. Drusinsky. Behavioral and temporal rule checking for Gaussian random process — a Kalman filter example. J.UCS: Journal of Universal Computer Science, 19(15): 2198–??, ???? 2013. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_

                                        19_15/behavioral_and_
                                        temporal_rule.

Drusinsky:2003:MTL

[DS03] D. Drusinsky and M. T. Shing. Monitoring temporal logic specifications combined with time series constraints. J.UCS: Journal of Universal Computer Science, 9(11):1261– 1276, November 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_9_11/monitoring_
                                        temporal_logic_specification.

Durand:2008:TRA

[DS08] I. A. Durand and S. R. Schwer. A tool for reasoning about qualitative temporal information: the theory of S-languages with a Lisp implementation. J.UCS: Journal of Universal Computer Science, 14(20):3282–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        20/a_tool_for_reasoning.

Diener:2010:CPF

[DS10] H. Diener and P. Schuster. On choice principles and fan theorems. J.UCS: Journal of Universal Computer Science, 16(18):2556–??, ????

2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_18/on_choice_
principles_and.

Dodero:2007:GIE

[DSAFW07] J. M. Dodero, S. Sánchez-Alonso, and D. Frosch-Wilke. Generative instructional engineering of competence development programmes. J.UCS: Journal of Universal Computer Science, 13(9): 1213–1233, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_13_9/generative_
instructional_engineering_
of.

Brito:2008:LLI

[dSBGdAdLM08] K. dos Santos Brito, V. C. Garcia, E. Santana de Almeida, and S. R. de Lemos Meira. LIFT — a Legacy In-Formation Retrieval Tool. J.UCS: Journal of Universal Computer Science, 14(8):1256–1284, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_14_8/lift_a_
legacy_information.

daSilva:2005:EEJ

[dSC05] A. Faustino da Silva and V. Santos Costa. An experimental evaluation of JAVA JIT technology. J.UCS: Journal of Universal Computer Science, 11(7):1291–1310, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_7/an_experimental_
evaluation_of.

daSilva:2006:DYC

[dSC06a] A. F. da Silva and V. S. Costa. The design of the YAP compiler: An optimizing compiler for logic programming languages. J.UCS: Journal of Universal Computer Science, 12(7):764– 787, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
7/the_design_of_the.

daSilva:2006:OEO

[dSC06b] A. F. da Silva and V. S. Costa. Our experiences with optimizations in Sun’s Java just-in-time compilers. J.UCS: Journal of Universal Computer Science, 12(7):788– 810, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_

12_7/our_experiences_
with_optimizations.

DelMondo:2010:GMS

[DSCT10] G. Del Mondo, J. G.

Stell, C. Claramunt, and R. Thibaud. A graph model for spatio-temporal evolution. J.UCS: Journal of Universal Computer Science, 16(11):1452–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
11/a_graph_model_for.

Junior:2014:WSC

[dSJPM14] L. L. Nunes da Silva Junior, A. Plastino, and L. G. P. Murta. What should I code now? J.UCS: Journal of Universal Computer Science, 20(5):797–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
20_5/what_should_i_code.

daSilva:2008:NEA

[dSLMW08] J. M. Monte da Silva, R. Dueire Lins, F. M. J. Martins, and R. Wachenchauzer. A new and efficient algorithm to binarize document images removing back-to-front interference. J.UCS: Journal of Universal Computer Science, 14(2):299– 313, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        2/a_new_and_efficient.

Denford:2004:AAT

[DSLO04] M. Denford, A. Solomon, J. Leaney, and T. O’Neill. Architectural abstraction as transformation of poset labelled graphs. J.UCS: Journal of Universal Computer Science, 10(10):1408–1428, October 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        10_10/architectural_
                                        abstraction_as_transformation.

Daramola:2013:TBS

[DSM13] O. Daramola, G. Sindre, and T. Moser. A tool-based semantic framework for security requirements specification. J.UCS: Journal of Universal Computer Science, 19(13):1940–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_19_
                                        13/a_tool_based_semantic.

Dvorak:2003:FSC

[DSRR03] V. Dvorak, M. Sveda, C. Rattray, and J. W.

Rozenblit. Formal specifications of computer-based systems — J.UCS special issue. J.UCS: Journal of Universal Computer Science, 9(11):1258– 1260, November 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
 http://www.jucs.org/
jucs_9_11/formal_specifications_
of_computer.

Dumitrescu:2007:ATS

[DT07] A. Dumitrescu and C. D.

Tóth. Analysis of two sweep-line algorithms for constructing spanning trees and Steiner trees. J.UCS: Journal of Universal Computer Science, 13(11):1615–1627, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_13_11/analysis_
of_two_sweep.

Dassow:2009:PNC

[DT09] J. Dassow and S. Turaev.

Petri net controlled grammars: the case of special Petri nets. J.UCS: Journal of Universal Computer Science, 15(14): 2808–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/

                                        jucs_15_14/petri_net_
                                        controlled_grammars.

Dahanayake:2012:CMI

[DT12] A. Dahanayake and B. Thalheim. A conceptual model for IT service systems. J.UCS: Journal of Universal Computer Science, 18(17):2452–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_18_
                                        17/a_conceptual_model_
                                        for.

Dang:2010:CCB

[DTG10] D.-H. Dang, A.-H. Truong, and M. Gogolla. Checking the conformance between models based on scenario synchronization. J.UCS: Journal of Universal Computer Science, 16(17):2293–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_16_17/checking_
                                        the_conformance_between.

deLandaFarias:2003:MOR

[dTR03] Luciana de Landa Farias, Guilherme H. Travassos, and Ana Regina Rocha. Managing organizational risk knowledge. J.UCS: Journal of Universal Computer Science, 9(7):670– 681, July 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
7/managing_organizational_
risk_knowledge.

deMeo:2004:XGS

[dTU04] P. de Meo, G. Terracina, and D. Ursino. X-Global: a system for the “almost automatic” and semantic integration of XML sources at various flexibility levels. J.UCS: Journal of Universal Computer Science, 10(9):1065– 1109, September 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_10_9/x_global_a_
system.

Dudek:2008:AFI

[Dud08] D. Dudek. The APS framework for incremental learning of software agents. J.UCS: Journal of Universal Computer Science, 14(14):2263–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
14/the_aps_framework_
for.

Dunne:1996:CMB

[Dun96] P. E. Dunne. Ceilings of monotone Boolean functions. J.UCS: Journal of Universal Computer Science, 2(7):533– 548, July 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
ceilings_of_monotone_
boolean_functions.

Dustdar:2005:RKM

[Dus05] S. Dustdar. Reconciling knowledge management and workflow management systems: The activity-based knowledge management approach. J.UCS: Journal of Universal Computer Science, 11(4):589–604, April 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
4/reconciling_knowledge_
management_and.

Duval:2001:MSW

[Duv01]  E. Duval.  Metadata

standards: What, who & why. J.UCS: Journal of Universal Computer Science, 7(7):591– 601, July 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
7/metadata_standards_
what_who.

Dvorak:1997:BSD

[Dvo97]  V. Dvorak.  Bounds

on size of decision diagrams. J.UCS: Journal of Universal Computer Science, 3(1):2–22, January 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_3_
1/bounds_on_size.

Dvorak:2000:JUS

[Dvo00] V. Dvorak. J.UCS special issue on formal specifications of computer-based systems. J.UCS: Journal of Universal Computer Science, 6(11): 1052–1053, November 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_6_11/j_ucs_
special_issue.

Dong:2007:APA

[DXZL07] F. Dong, R. Xiao, Y. Zhong, and Y. Liu. An approach to polygonal approximation of digital curves based on discrete particle swarm algorithm. J.UCS: Journal of Universal Computer Science, 13(10):1449–1461, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_

                                        13_10/an_approach_to_
                                        polygonal.

Dickinger:2008:CMI

[DZ08] A. Dickinger and S. Zorn. Compensation models for interactive advertising. J.UCS: Journal of Universal Computer Science, 14(4):557–565, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        4/compensation_models_
                                        for_interactive.

Despotovic-Zrakic:2012:WBE

[DZBB^{+}12] M. Despotović-Zrakić, D. Barać, Z. Bogdanović, B. Jovanić, and B. Radenković. Web-based environment for learning discrete event simulation. J.UCS: Journal of Universal Computer Science, 18(10): 1259–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_18_10/web_based_
                                        environment_for.

Estevez-Ayres:2013:APR

[EACGFK13] I. Estévez-Ayres, R. M. Crespo-García, J. A. Fisteus, and C. Delgado Kloos. An algorithm for peer review matching in massive courses for minimising students’ frustration. J.UCS: Journal of Universal Computer Science, 19(15):2173–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
15/an_algorithm_for_
peer.

Eldershaw:2000:UGA

[EC00] C. Eldershaw and S. Cameron.

Using genetic algorithms to solve the motion planning problem. J.UCS: Journal of Universal Computer Science, 6(4): 422–432, April 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
4/using_genetic_algorithms_
to.

Estivill-Castro:2010:NCF

[ECHS10] V. Estivill-Castro, A. Heednacram, and F. Suraweera. NP-completeness and FPT results for rectilinear covering problems. J.UCS: Journal of Universal Computer Science, 16(5): 622–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_5/np_completeness_
and_ftp.

ElHog:2014:UAA

[EDA14] C. El Hog, R. Ben Djemaa, and I. Amous. A user-aware approach to provide adaptive Web services. J.UCS: Journal of Universal Computer Science, 20(7):944–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_20_7/a_user_
                                        aware_approach.

Edelsbrunner:2001:WT

[Ede01] H. Edelsbrunner. 180 wrapped tubes. J.UCS: Journal of Universal Computer Science, 7(5): 379–399, May 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_7_
                                        5/180_wrapped_tubes.

Entrialgo:2009:TSM

[EGDG09] J. Entrialgo, J. García, J. L. Díaz, and D. F. García. Tools and stochastic metrics for debugging temporal behaviour of real-time systems. J.UCS: Journal of Universal Computer Science, 15(8):1563–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_15_8/tools_and_
                                        stochastic_metrics.

Eschbach:2001:FDS

[EGG^{+}01] R. Eschbach, U. Glässer, R. Gotzhein, M. von Lövis, and A. Prinz. Formal definition of SDL-2000 — compiling and running SDL specifications as ASM models. J.UCS: Journal of Universal Computer Science, 7(11):1024–1049, November 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
11/formal_definition_
of_sdl.

Elias:2012:VDP

[EGK^{+}12] S. Elias, V. Gokul, K. Krithivasan, M. Gheorghe, and G. Zhang. A variant of distributed P systems for real time cross layer optimization. J.UCS: Journal of Universal Computer Science, 18(13):1760–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_13/a_variant_
of_distributed.

Elramly:2005:ADP

[EHEH05] N. A. Elramly, A. S.

Habib, O. S. Essa, and H. M. Harb. Analysis, design, and performance evaluation of MS-RTCP: More scalable scheme for the real-time control protocol. J.UCS: Journal of Universal Computer Science, 11(6):874– 897, ???? 2005. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        11_6/analysis_design_
                                        and_performance.

Einsele:2008:LIO

[EIH08] F. Einsele, R. Ingold, and J. Hennebert. A languageindependent, open-vocabulary system based on HMMs for recognition of ultra low resolution words. J.UCS: Journal of Universal Computer Science, 14 (18):2982–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_18/a_language_independent_
                                        open.

Echtle:1999:JUS

[EK99] K. Echtle and J. Keller. J.UCS special issue on dependability evaluation and validation. J.UCS: Journal of Universal Computer Science, 5 (10):632, October 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_5_10/j_ucs_
                                        special_issue.

Evripidou:2000:DDN

[EK00] P. Evripidou and C. Kyriacou. Data driven network of workstations (D^{2}NOW). J.UCS: Journal of Universal Computer Science, 6(10):1015– 1033, October 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
10/data_driven_network_
of.

Esik:2002:RAS

[ÉK02] Z. Ésik and W. Kuich.

Rationally additive semirings. J.UCS: Journal of Universal Computer Science, 8(2):173– 183, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_8_2/rationally_
additive_semirings.

Ehrich:2003:COS

[EKP03] H.-D. Ehrich, M. Kollmann, and R. Pinger. Checking object system designs incrementally. J.UCS: Journal of Universal Computer Science, 9(2):106– 119, February 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
9_2/checking_object_
system_designs.

Ebner:2015:LA

[EKW^{+}15] M. Ebner, Kinshuk, D. Wohlhart, B. Taraghi, and V. Kumar. Learning analytics. J.UCS: Journal of Universal Computer Science, 21(1):1–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_21;   http:
                                        //www.jucs.org/jucs_
                                        21_1/learning_analytics.

Exman:2015:SUC

[ELFAR15] I. Exman, J. Llorens, A. Fraga, and J. M. Alvarez-Rodríguez. SKY-Ware: The unavoidable convergence of software towards runnable knowledge. J.UCS: Journal of Universal Computer Science, 21(11):1405–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_21_
                                        11/skyware_the_unavoidable_
                                        convergence.

Esposito:2004:DSM

[ELS04] F. Esposito, O. Licchelli, and G. Semeraro. Discovering student models in e-learning systems. J.UCS: Journal of Universal Computer Science, 10(1):47–57, January 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
1/discovering_student_
models_in.

Espada:2012:SMB

[EMGB^{+}12] J. Pascual Espada, O. Sanjuán Martínez, B. C. Pelayo G-Bustelo, J. M. Cueva Lovelle, and P. Ordoñez de Pablos. A simple model based on Web services to exchange context information between Web browsers and Web applications. J.UCS: Journal of Universal Computer Science, 18(11): 1410–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_11/a_simple_model_
based.

Efstathiou:2014:EMO

[EMZB14] D. Efstathiou, P. McBurney, S. Zschaler, and J. Bourcier. Efficient multi-objective optimisation of service compositions in mobile ad hoc networks using lightweight surrogate models. J.UCS: Journal of Universal Computer Science, 20(8): 1089–??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_

                                        20_8/efficient_multi_
                                        objective_optimisation.

Eppler:2004:FKC

[Epp04]  M. J. Eppler.  Facil-

itating knowledge communication through joint interactive visualization. J.UCS: Journal of Universal Computer Science, 10(6):683–690, June 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_10_
                                        6/facilitating_knowledge_
                                        communication_through.

Erne:2011:WPK

[Ern11] R. Erne. What is productivity in knowledge work?

— a cross-industrial view. J.UCS: Journal of Universal Computer Science, 17(10):1367–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_17_10/what_
                                        is_productivity_in.

Efimova:2003:CKM

[ES03] L. Efimova and J. Swaak. Converging knowledge management, training and e-learning: Scenarios to make it work. J.UCS: Journal of Universal Computer Science, 9(6):571–578, June 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_9_6/converging_
knowledge_management_
training.

Ekwall:2005:RUA

[ES05] R. Ekwall and A. Schiper.

Replication: Understanding the advantage of atomic broadcast over quorum systems. J.UCS: Journal of Universal Computer Science, 11(5): 703–711, May 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_5/replication_understanding_
the_advantage.

Emerson:2004:MBM

[ESB04] M. J. Emerson, J. Sztipanovits, and T. Bapty. A MOF-based metamodeling environment. J.UCS: Journal of Universal Computer Science, 10 (10):1357–1382, October 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_10_10/a_mof_
based_metamodeling.

ElHelou:2010:PCR

[ESG10] S. El Helou, C. Salzmann, and D. Gillet. The 3A personalized, contextual and relationbased recommender system. J.UCS: Journal of Universal Computer Science, 16(16):2179–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
16/the_3A_personalized_
contextual.

Edwards:2008:MIF

[ESM08] G. Edwards, C. Seo, and N. Medvidovic. Model interpreter frameworks: a foundation for the analysis of domain-specific software architectures. J.UCS: Journal of Universal Computer Science, 14(8):1182–1210, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
8/model_interpreter_
frameworks_a.

Espak:2006:JRB

[Esp06] M. Espák. Japlo: Rulebased programming on Java. J.UCS: Journal of Universal Computer Science, 12(9):1177– 1189, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_12_9/japlo_rule_
based_programming.

Ewert:1999:SIF

[Ev99] S. Ewert and A. van der Walt. Shrink indecomposable fractals. J.UCS: Journal of Universal Computer Science, 5(9):521–531, September 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_5_9/shrink_
indecomposable_fractals.

French:2006:ERJ

[FA06] A. French and J. N. Amaral. Eliminating redundant join-set computations in static single assignment. J.UCS: Journal of Universal Computer Science, 12(8):1007– 1019, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
8/eliminating_redundant_
join_set.

Falda:2010:SRI

[Fal10] M. Falda. Spatial reasoning with integrated qualitative-metric fuzzy constraint networks. J.UCS: Journal of Universal Computer Science, 16 (11):1390–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_

16_11/spatial_reasoning_
with_integrated.

Fernandez:2013:SAT

[FAT^{+}13] J. Martínez Fernández, J. C. Augusto, G. Trombino, R. Seepold, and N. Martinez Madrid. Selfaware trader: A new approach to safer trading. J.UCS: Journal of Universal Computer Science, 19(15):2292–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
15/self_aware_trader_
a.

Fazekas:2006:ADC

[Faz06] M. Fazekas. Analysing data of childhood acute lymphoid leukaemia by seasonal time series methods. J.UCS: Journal of Universal Computer Science, 12(9):1190–1195, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
9/analysing_data_of_
childhood%20.

Flores:2015:CLS

[FBCMR15] E. Flores, A. Barrón-Cedeño, L. Moreno, and P. Rosso. Cross-language source code re-use detection using latent semantic analysis. J.UCS: Journal of Universal Computer Science, 21(13): 1708–??, ???? 2015. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
21_13/cross_language_
source_code.

Fidalgo-Blanco:2015:MAT

[FBSEGP15] Á. Fidalgo-Blanco, M. L.

Sein-Echaluce, and F. J. García-Peñalvo. Methodological approach and technological framework to break the current limitations of MOOC model. J.UCS: Journal of Universal Computer Science, 21(5):712–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
21_5/methodological_
approach_and_technological.

Feldmann:2003:EBS

[FC03] R. L. Feldmann and R. Carbon. Experience base schema building blocks of the PLEASERS library. J.UCS: Journal of Universal Computer Science, 9(7):659– 669, July 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
9_7/experience_base_
schema_building.

Forment:2012:DGG

[FCM^{+}12] M. Alier Forment, M. J. Casany, E. Mayol, J. Piguillem, N. Galanis, F. J. García-Peñalvo, and M. Á. Conde. Docs4Learning: Getting Google Docs to work within the LMS with IMS BLTI. J.UCS: Journal of Universal Computer Science, 18(11): 1483–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        18_11/docs4Learning_
                                        getting_google_docs.

Feidakis:2013:DMS

[FDC^{+}13] M. Feidakis, T. Daradoumis, S. Caballe, J. Conesa, and D. Gañán. A dualmodal system that evaluates user’s emotions in virtual learning environments and responds affectively. J.UCS: Journal of Universal Computer Science, 19(11):1638–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_19_
                                        11/a_dual_modal_system.

Florencio:2015:NIE

[FDR^{+}15] C. C. Florêncio, J. Daenen, J. Ramon, J. Van den Bussche, and D. Van Dyck. Naive infinite enumeration of contextfree languages in incremental polynomial time. J.UCS: Journal of Universal Computer Science, 21(7):891–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
21_7/naive_infinite_
enumeration_of.

Fellner:2001:GCD

[Fel01] Dieter W. Fellner. Graphics content in digital libraries: Old problems, recent solutions, future demands. J.UCS: Journal of Universal Computer Science, 7(5):400– 409, May 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
7_5/graphics_content_
in_digital.

Fenwick:1995:DZL

[Fen95a]  P. Fenwick.  Differ-

ential Ziv–Lempel text compression. J.UCS: Journal of Universal Computer Science, 1 (8):591–602, August 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/differential_ziv_
lempel_text_compression.

Fenwick:1995:HRD

[Fen95b] P. Fenwick. High-radix division with approximate quotient-digit estimation. J.UCS: Journal of Universal Computer Science, 1(1):2–22, January 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/high_
                                        radix_division_with/
                                        html/abstract.html.

Fenwick:1997:SRT

[Fen97]  P. Fenwick.  Symbol

ranking text compression with Shannon recordings. J.UCS: Journal of Universal Computer Science, 3(2):70–85, February 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_3_2/symbol_
                                        ranking_text_compression.

Fenwick:2015:NVL

[Fen15]  P. Fenwick.  A note

on variable-length codes with constant Hamming weights. J.UCS: Journal of Universal Computer Science, 21(9):1136– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_21_
                                        9/a_note_on_variable.

Fernau:1996:RPP

[Fer96]  H. Fernau.  Remarks

on propagating partitionlimited ETOL systems. J.UCS: Journal of Universal Computer Science, 2(11):745–755, November 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_2_11/remarks_on_
propagating_partition.

Fontana:2004:FME

[FF04] F. Fontana and G. Franco.

Finding the maximum element using P systems. J.UCS: Journal of Universal Computer Science, 10(5):567–580, May 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
5/finding_the_maximum_
element.

Fetzer:2007:IPC

[FF07] C. Fetzer and P. Felber. Improving program correctness with atomic exception handling. J.UCS: Journal of Universal Computer Science, 13(8):1047–1072, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_

                                        8/improving_program_
                                        correctness_with.

Ferreira:2008:RIW

[FF08] F. Ferreira and G. Ferreira. The Riemann integral in weak systems of analysis. J.UCS: Journal of Universal Computer Science, 14(6):908– 937, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        6/the_riemann_integral_
                                        in.

Fuhr:2004:AHC

[FFK04] D. Fuhr and F. Fuchs-Kittowski. Against hierarchy and chaos knowledge coproduction in nets of experts. J.UCS: Journal of Universal Computer Science, 10(3):176–185, March 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_10_
                                        3/against_hierarchy_
                                        and_chaos.

Fuglseth:2003:TKM

[FG03] A. M. Fuglseth and K. Groenhaug. A tool kit for measurement of organisational learning: Methodological requirements and an illustrative example. J.UCS: Journal of Universal Computer Science, 9(12):1487– 1499, December 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
12/a_tool_kit_for.

Fuentes:2010:CPS

[FG10] L. Fuentes and N. Gámez.

Configuration process of a software product line for AmI middleware. J.UCS: Journal of Universal Computer Science, 16 (12):1592–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_12/configuration_
process_of_a.

Fathy:2014:ARR

[FGB^{+}14] N. Fathy, T. F. Gharib, N. Badr, A. S. Mashat, and A. Abraham. A personalized approach for re-ranking search results using user preferences. J.UCS: Journal of Universal Computer Science, 20(9):1232–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
20_9/a_personalized_
approach_for.

Ferreira:2014:TCS

[FGBS14] F. Ferreira, R. Gheyi, P. Borba, and G. Soares. A toolset for checking SPL refinements. J.UCS: Journal of Universal Computer Science, 20(5):587– ??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_20_5/a_toolset_
for_checking.

Faro:1998:LBS

[FGS98] A. Faro, D. Giordano, and C. Santoro. Link-based shaping of hypermedia webs assisted by a neural agent. J.UCS: Journal of Universal Computer Science, 4(7):630– 651, July 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_7/link_
based_shaping_of;   internal&
sk=0C220489.

Ferrarotti:2014:CMS

[FGSW14] F. Ferrarotti, G. Grossmann, K.-D. Schewe, and Q. Wang. Conceptual modelling with specific focus on service-oriented systems. J.UCS: Journal of Universal Computer Science, 20(3):254– ??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_20;   http:

//www.jucs.org/jucs_
20_3/conceptual_modelling_
with_specific.

Foldes:2000:MHQ

[FH00] S. Foldes and P. L. Hammer. Monotone, Horn and quadratic pseudo-Boolean functions. J.UCS: Journal of Universal Computer Science, 6(1):97– 104, January 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
1/monotone_horn_and_
quadratic.

Fernandez:2006:ICB

[FH06] A. J. Fernández and P. M. Hill. An interval constraint branching scheme for lattice domains. J.UCS: Journal of Universal Computer Science, 12(11):1466–1499, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
11/an_interval_constraint_
branching.

Fiedler:2008:SEI

[FHH08] M. Fiedler, K. D. Hackbarth, and H. Hlavacs. Socio-economic issues in future generation Internet. J.UCS: Journal of Universal Computer Science, 14(5):612–614, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14;
 http://www.jucs.org/
jucs_14_5#;   http://
www.jucs.org/jucs_14_
5/socio_economic_issues_
in.

Franke:1999:AOI

[FHJ^{+}99] A. Franke, S. M. Hess, Ch. G. Jung, M. Kohlhase, and V. Sorge. Agentoriented integration of distributed mathematical services. J.UCS: Journal of Universal Computer Science, 5(3):156– 225, March 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
3/agent_oriented_integration_
of.

Fuentes:2006:DAI

[FJP06] L. Fuentes, D. Jiménez, and M. Pinto. Development of ambient intelligence applications using components and aspects. J.UCS: Journal of Universal Computer Science, 12(3):236–251, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_3/development_of_
ambient_intelligence.

Faizrahmanov:2016:ESH

[FK16] M. Faizrahmanov and I. Kalimullin. The enumeration spectrum hierarchy of α-families and low_{α} degrees. J.UCS: Journal of Universal Computer Science, 22(7):943– ??, ???? 2016. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
22_7/the_enumeration_
spectrum_hierarchy.

Falt:2015:TEL

[FKB^{+}15] Z. Falt, M. Krulis, D. Bednárek, J. Yaghob, and F. Zavoral. Towards efficient locality aware parallel data stream processing. J.UCS: Journal of Universal Computer Science, 21(6):816– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
6/towards_efficient_
locality_aware.

Fleischer:2010:TSA

[FKK^{+}10] R. Fleischer, T. Kamphans, R. Klein, E. Langetepe, and G. Trippen. The tourist in the shopping arcade. J.UCS: Journal of Universal Computer Science, 16(5):676– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
5/the_tourist_in_the.

Fantinato:2014:SCA

[FKO14] M. Fantinato, U. Kulesza, and F. Oquendo. Software components, architectures and reuse: Software product line engineering and source code enhancements. J.UCS: Journal of Universal Computer Science, 20(5):583–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_20;   http:
//www.jucs.org/jucs_
20_5/software_components_
architectures_and.

Franco:2004:FCL

[FKS^{+}04] J. Franco, M. Kouril, J. Schlipf, S. Weaver, M. Dransfield, and W. M. Vanfleet. Functioncomplete lookahead in support of efficient SAT search heuristics. J.UCS: Journal of Universal Computer Science, 10 (12):1655–1692, December 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
12/function_complete_
lookahead_in.

Fauzi:2010:BBA

[FL10] M. F. A. Fauzi and P. H. Lewis. Block-based against segmentationbased texture image retrieval. J.UCS: Journal of Universal Computer Science, 16(3):402–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16_3/block_
based_against_segmentation.

Frixione:2014:TEM

[FL14] M. Frixione and A. Lieto. Towards an extended model of conceptual representations in formal ontologies: A typicality-based proposal. J.UCS: Journal of Universal Computer Science, 20(3):257–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_20_3/towards_an_
extended_model.

Fu:2014:UOR

[FLF^{+}14] X. Fu, W. Li, G. Fortino, P. Pace, G. Aloi, and W. Russo. A utilityoriented routing scheme for interest-driven communitybased opportunistic networks. J.UCS: Journal of Universal Computer Science, 20(13):1829–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_20_13/a_utility_
                                        oriented_routing.

Flor:2004:EAU

[Flo04] T. Flor. Experiences with adaptive user and learning models in eLearning systems for higher education. J.UCS: Journal of Universal Computer Science, 10(1):58–72, January 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_10_
                                        1/experiences_with_adaptive_
                                        user.

Flinn:1995:LA

[FM95] B. Flinn and H. Maurer. Levels of anonymity. J.UCS: Journal of Universal Computer Science, 1(1):35–47, January 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_1_1/levels_
                                        of_anonymity/html/abstract.
                                        html.

Ferrero:2005:ADL

[FMA^{+}05] B. Ferrero, M. Martín, A. Alvarez, M. Urretavizcaya, and I. Fernández-Castro. Authoring and diagnosis of learning activities with the KADD ET environment. J.UCS:

Journal of Universal Computer Science, 11(9): 1530–1542, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_11_9/authoring_
and_diagnosis_of.

Fahad:2011:TCW

[FMB^{+}11] M. Fahad, N. Moalla, A. Bouras, M. Abdul Qadir, and M. Farukh. Towards classification of Web ontologies for the emerging semantic Web. J.UCS: Journal of Universal Computer Science, 17(7):1021–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_7/towards_classification_
of_web.

Fardoun:2013:APS

[FML13] H. M. Fardoun, A. Mashat, and S. R. López. Applying professional solutions within the educational environments by means of cloud computing: Coaching for teachers. J.UCS: Journal of Universal Computer Science, 19(12):1703–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_12/applying_

                                        professional_solutions_
                                        within.

Fernandez-Manjon:2007:CEN

[FMLNF07] B. Fernández-Manjón, M. Llamas-Nistal, and R.- Á. Fernández. Computers in education: New developments in e-learning technology. J.UCS: Journal of Universal Computer Science, 13(7):920– 923, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13;
                                         http://www.jucs.org/
                                        jucs_13_7#;   http://
                                        www.jucs.org/jucs_13_
                                        7/computers_in_education_
                                        new.

Fernandez-Medina:2009:SIS

[FMR09] E. Fernández-Medina and A. Rodríguez. Security in information systems: New advances and tendencies. J.UCS: Journal of Universal Computer Science, 15(15):2912–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_15;   http:
                                        //www.jucs.org/jucs_
                                        15_15#;   http://www.
                                        jucs.org/jucs_15_15/
                                        security_in_information_
                                        systems.

Fraczek:2012:MSI

[FMS12] W. Fraczek, W. Mazurczyk, and K. Szczypiorski. Multilevel steganography: Improving hidden communication in networks. J.UCS: Journal of Universal Computer Science, 18(14):1967–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
14/multilevel_steganography_
improving_hidden.

Freire:2015:QTB

[FMT^{+}15] A. Freire, C. Macdonald, N. Tonellotto, I. Ounis, and F. Cacheda. Queuing theory-based latency/ power tradeoff models for replicated search engines. J.UCS: Journal of Universal Computer Science, 21(13):1790–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
21_13/queuing_theory_
based_latency.

Freitas:2008:MZM

[FOAB08] C. O. A. Freitas, L. S.

Oliveira, S. B. K. Aires, and F. Bortolozzi. Metaclasses and zoning mechanism applied to handwriting recognition. J.UCS: Journal of Universal Computer Science, 14(2): 211–223, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_2/metaclasses_and_
                                        zoning_mechanism.

Fodor:2006:SMA

[Fod06] N. Fodor. 4M — software for modelling and analysing cropping systems. J.UCS: Journal of Universal Computer Science, 12(9):1196– 1207, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        12_9/4m_software_for_
                                        modelling%20.

Fontani:2000:EIC

[Fon00] S. Fontani. Efficient identification of classes of P - time functions. J.UCS: Journal of Universal Computer Science, 6(8): 759–780, August 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_6_
                                        8/efficient_identification_
                                        of_classes.

Fontani:2001:EML

[Fon01] S. Fontani. Efficient measure learning. J.UCS: Journal of Universal Computer Science, 7 (9):794–815, September 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
9/efficient_measure_
learning.

Forsell:1997:MMV

[For97] M. Forsell. MTAC — a multithreaded VLIW architecture for PRAM simulation. J.UCS: Journal of Universal Computer Science, 3(9):1037– 1055, September 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_3_9/mtac_
a_multithreaded_vliw.

Forster:2007:PTS

[For07] C. H. Q. Forster. Programming through spreadsheets and tabular abstractions. J.UCS: Journal of Universal Computer Science, 13(6):806– 816, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
6/programming_through_
spreadsheets_and.

Franck-Oberaspach:1999:PPS

[FOSS99] G. Franck-Oberaspach, D. B. Schweiger, and K. Svozil. A packing problem, solved by genetic algorithms. J.UCS: Journal of Universal Computer Science, 5(8):464– 470, August 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
8/a_packing_problem_
solved.

Freund:1995:VTC

[FP95] Rudolf Freund and Gheorghe Păun. A variant of team cooperation in grammar systems. J.UCS: Journal of Universal Computer Science, 1(2):105– 130, February 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
a_variant_of_team_cooperation_
in_grammar_systems.

Forzi:2005:MTI

[FP05] T. Forzi and M. Peters. A methodology and a toolkit that integrate technological, organisational, and human factors to design KM within knowledge-intensive networks. J.UCS: Journal of Universal Computer Science, 11(4):495–525, April 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
4/a_methodology_and_a.

Fokkink:2006:VIR

[FP06] W. Fokkink and J. Pang.

Variations on Itai–Rodeh leader election for anonymous rings and their analysis in PRISM. J.UCS: Journal of Universal Computer Science, 12(8): 981–1006, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_12_8/variations_
on_itai_rodeh.

Ferri:2012:USS

[FP12] P. Ferri and A. Pozzali. University students and social media: Reflections from an empirical research. J.UCS: Journal of Universal Computer Science, 18(3):377–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_3/university_
students_and_social.

Figueira:2003:DTC

[FPLS03] A. R. Figueira, H. Paulino, L. Lopes, and F. Silva. Distributed typed concurrent objects: a programming language for distributed computations with mobile resources. J.UCS: Journal of Universal Computer Science, 9(8):745–760, August 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
8/distributed_typed_
concurrent_objects.

Fonseca:2012:PEP

[FPS^{+}12] B. Fonseca, Â. Pereira, R. Sanders, V. Barracho, U. Lapajne, M. Rus, M. Rahe, A. Mostert, T. Klein, V. Bojovic, S. Bo jak, L. Morgado, Z. Bo jak, J. Carvalho, I. Duarte, A. Casaramona, A. Soraci, H. Paredes, P. Martins, R. Gonçalves, P. Neves, R. Rodrigues Nunes, J. Lima, and J. Varajão. PLAYER — a European Project and a game to foster entrepreneurship education for young people. J.UCS: Journal of Universal Computer Science, 18(1):86–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_18_1/player_
a_european_project.

Fernandez-Perez:2007:PEL

[FPSFCG07] J. Fernández-Pérez, F. J. Sánchez-Fernández, and R. Carmona-Galán. Performance evaluation and limitations of a vision system on a reconfigurable/programmable chip. J.UCS: Journal of Universal Computer Science, 13(3):440–453, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
3/performance_evaluation_
and_limitations.

Ferrarotti:2010:RRR

[FPT10] F. A. Ferrarotti, A. L.

Paoletti, and J. M. Turull Torres. Redundant relations in relational databases: a model theoretic perspective. J.UCS: Journal of Universal Computer Science, 16(20):2934–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_20/redundant_
relations_in_relational.

Filho:2004:ICE

[FR04] F. C. Filho and C. M. F.

Rubira. Implementing coordinated error recovery for distributed objectoriented systems with AspectJ. J.UCS: Journal of Universal Computer Science, 10(7):843–858, July 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
7/implementing_coordinated_
error_recovery.

Franz:1998:OSB

[Fra98] M. Franz. Open standards beyond Java: On the future of mobile code for the Internet. J.UCS: Journal of Universal Computer Science, 4(5):522– 533, May 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_5/open_
                                        standards_beyond_java;
                                         internal&sk=0C220489.

Fabry:2014:DSA

[FRD14] J. Fabry, R. Robbes, and M. Denker. DIE: A domain specific aspect language for IDE events. J.UCS: Journal of Universal Computer Science, 20(2):135–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_20_2/die_a_
                                        domain_specific.

Fronk:2002:OOA

[Fro02] A. Fronk. An objectoriented approach to design, specification, and implementation of hyperlink structures based on usual software development. J.UCS: Journal of Universal Computer Science, 8(10):892– 912, October 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_8_10/an_object_
oriented_approach.

Ferreira:2011:PMC

[FSdRSS11] R. Pinto Ferreira, J. Neves Silva, F. do Rocio Strauhs, and A. L. Soares. Performance management in collaborative networks: a methodological proposal. J.UCS: Journal of Universal Computer Science, 17(10):1412–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
10/performance_management_
in_collaborative.

Fidalgo:2013:TIP

[FSELC13] Á. Fidalgo, M. L. Sein-Echaluce, D. Lerís, and O. Castañeda. Teaching Innova Project: the incorporation of adaptable outcomes in order to grade training adaptability. J.UCS: Journal of Universal Computer Science, 19(11):1500–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_11/teaching_
innova_project_the.

Freitas:2007:OTA

[FSMP07] F. Freitas, H. Stuckenschmidt, A. Malucelli, and H. S. Pinto. Ontologies and their applications. J.UCS: Journal of Universal Computer Science, 13(12):1801–1804, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13;
                                         http://www.jucs.org/
                                        jucs_13_12#;   http:
                                        //www.jucs.org/jucs_
                                        13_12/ontologies_and_
                                        their_applications.

Frausto-Solis:2013:CPS

[FSSPLG^{+}13] J. Frausto-Solís, M. Sánchez-Pérez, E. Liñan-García, J. P. Sánchez-Hernández, and M. Ramachandran. Cluster perturbation simulated annealing for protein folding problem. J.UCS: Journal of Universal Computer Science, 19(15):2207–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_19_15/cluster_perturbation_
                                        simulated_annealing.

Ferrer-Troyano:2005:CSV

[FTARR05a] F. J. Ferrer-Troyano, J. S. Aguilar-Ruiz, and J. C. Riquelme. Connecting segments for visual data exploration and interactive mining of decision rules. J.UCS: Journal of Universal Computer Science, 11(11):1835–1848, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
11/connecting_segments_
for_visual.

Ferrer-Troyano:2005:IRL

[FTARR05b] F. J. Ferrer-Troyano, J. S.

Aguilar-Ruiz, and J. C. Riquelme. Incremental rule learning and border examples selection from numerical data streams. J.UCS: Journal of Universal Computer Science, 11(8):1426– 1439, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
8/incremental_rule_learning_
and.

Fernandez-Villamor:2012:MWC

[FVIG12] J. I. Fernández-Villamor, C. Á. Iglesias, and M. Garijo. A metadirectory of Web components for mashup composition. J.UCS: Journal of Universal Computer Science, 18(17):2407–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_18_
                                        17/a_metadirectory_of_
                                        web.

Friedrich:2011:ERE

[FWS^{+}11] M. Friedrich, M. Wolpers, R. Shen, C. Ullrich, R. Klamma, D. Renzel, A. Richert, and B. von der Heiden. Early results of experiments with responsive open learning environments. J.UCS: Journal of Universal Computer Science, 17(3):451–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_17_3/early_
                                        results_of_experiments.

Fernandes:2011:AFM

[FWT11] P. Fernandes, C. Werner, and E. Teixeira. An approach for feature modeling of context-aware software product line. J.UCS: Journal of Universal Computer Science, 17(5):807–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        17_5/an_approach_for_
                                        feature.

Fellner:2000:ESM

[FZ00] D. W. Fellner and M. Zens. Electronic submission, managing and approval of grant proposals at the German Research Foundation based on standard Internet and office tools. J.UCS: Journal of Universal Computer Science, 6(3):356–366, March 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_6_3/electronic_
submission_managing_
and.

Fardoun:2013:IDS

[FZAP13] H. M. Fardoun, B. Zafar, A. H. Altalhi, and A. Paules. Interactive design system for schools using cloud computing. J.UCS: Journal of Universal Computer Science, 19(7):950– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
7/interactive_design_
system_for.

Fan:2013:KIS

[FZT13] J. Fan, Y. Zheng, and X. Tang. Key-insulated signcryption. J.UCS: Journal of Universal Computer Science, 19 (10):1351–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_10/key_insulated_
signcryption.

Galias:1998:RNS

[Gal98] Z. Galias. Rigorous numerical studies of the existence of periodic orbits for the Hénon map. J.UCS: Journal of Universal Computer Science, 4(2):114–124, February 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_2/rigorous_
                                        numerical_studies_of.

Guesgen:2002:JUS

[GALR02] H. W. Guesgen, F. D. Anger, G. Ligozat, and R. V. Rodríguez. J.UCS special issue on spatial and temporal reasoning. J.UCS: Journal of Universal Computer Science, 8(8):737–738, August 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/free;
                                        http://www.jucs.org/
                                        jucs_8_8/spatial_and_
                                        temporal_reasoning.

Guesgen:2003:STR

[GALR03] H. W. Guesgen, F. D. Anger, G. Ligozat, and R. V. Rodríguez. Spatial and temporal reasoning. J.UCS: Journal of Universal Computer Science, 9(9):984– 985, September 28, 2003.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
http://www.jucs.org/
jucs_9_9/spatial_and_
temporal_reasoning.

Giachetti:2010:LUM

[GAMP10] G. Giachetti, M. Albert, B. Marín, and O. Pastor. Linking UML and MDD through UML profiles: a practical approach based on the UML association. J.UCS: Journal of Universal Computer Science, 16(17):2353–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
17/linking_uml_and_mdd.

Ganter:2004:CAA

[Gan04] B. Ganter. Conflict avoidance in additive order diagrams. J.UCS: Journal of Universal Computer Science, 10(8):955–966, August 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
8/conflict_avoidance_
in_additive.

Gartner:1999:TAS

[Gär99] F. C. Gärtner. Transformational approaches to the specification and verification of fault-tolerant systems: Formal background and classification. J.UCS: Journal of Universal Computer Science, 5(10):668–692, October 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
10/transformational_
approaches_to_the.

Gassner:2009:ORP

[Gaß09] C. Gaßner. Oracles and relativizations of the P =? NP question for several structures. J.UCS: Journal of Universal Computer Science, 15(6):1186– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
6/oracles_and_relativizations_
of.

Gassner:2010:SRV

[Gaß10] C. Gaßner. The separation of relativized versions of P and DNP for the ring of the reals. J.UCS: Journal of Universal Computer Science, 16(18): 2563–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_18/the_separation_
of_relativized.

Garcia-Alvarez:2014:CTS

[GÁVCNC14] M. T. García-Álvarez, L. Varela-Candamio, and I. Novo-Corti. Chat as a tool for social knowledge construction using asynchronous discussion groups in economics degree. J.UCS: Journal of Universal Computer Science, 20(10):1443–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
10/chat_as_a_tool.

Grabert:2003:CBR

[GB03] M. Grabert and D. Bridge.

Case-based reuse of software examplets. J.UCS: Journal of Universal Computer Science, 9(7): 627–640, July 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
7/case_based_reuse_of.

Guesgen:2010:AST

[GB10] H. W. Guesgen and M. Bhatt. Advances in spatial and temporal reasoning. J.UCS: Journal of Universal Computer Science, 16(11):1388–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16;   http:

                                        //www.jucs.org/jucs_
                                        16_11#;   http://www.
                                        jucs.org/jucs_16_11/
                                        advances_in_spatial_
                                        and.

Gharout:2012:AGK

[GBCA12] S. Gharout, A. Bouabdallah, Y. Challal, and M. Achemlal. Adaptive group key management protocol for wireless communications. J.UCS: Journal of Universal Computer Science, 18(6):874–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        18_6/adaptive_group_
                                        key_management.

Gharib:2012:EOC

[GBHA12] T. F. Gharib, N. Badr, S. Haridy, and A. Abraham. Enriching ontology concepts based on texts from WWW and corpus. J.UCS: Journal of Universal Computer Science, 18(16):2234–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_18_16/enriching_
                                        ontology_concepts_based.

Gomez:2008:TSS

[GBP^{+}08] Y. Gómez, R. Bello, A. Puris, M. M. García, and A. Nowe. Two step swarm intelligence to solve the feature selection problem. J.UCS: Journal of Universal Computer Science, 14(15): 2582–??, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_15/two_step_swarm_
intelligence.

Grigalis:2014:USD

[GC14] T. Grigalis and A. Cenys.

Unsupervised structured data extraction from template-generated Web pages. J.UCS: Journal of Universal Computer Science, 20(2):169– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
2/unsupervised_structured_
data_extraction.

Goswami:2016:PRF

[GCC16] S. Goswami, A. Chakrabarti, and B. Chakraborty. A proposal for recommendation of feature selection algorithm based on data set characteristics. J.UCS: Journal of Universal Computer Science, 22(6):760–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/

                                        jucs_22_6/a_proposal_
                                        for_recommendation.

Granollers:2008:SHI

[GCG08] T. Granollers, C. A. Collazos, and M. P. González. The state of HCI in Ibero–American countries. J.UCS: Journal of Universal Computer Science, 14(16):2599–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        16/the_state_of_hci_
                                        in.

Guixeres:2013:EVR

[GCL^{+}13] J. Guixeres, L. Cantero, E. Lurbe, J. Saiz, M. Alcañiz, A. Cebolla, P. Escobar, R. Baños, C. Botella, J. F. Lison, and J. Alvarez. Effects of virtual reality during exercise in children. J.UCS: Journal of Universal Computer Science, 19(9):1199–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_19_9/effects_of_
                                        virtual_reality.

Granado-Criado:2007:DPR

[GCVRSPGP07] J. M. Granado-Criado, M. A. Vega-Rodríguez, J. M. Sánchez-Pérez, and J. A. Gómez-Pulido. A dynamically and partially reconfigurable implementation of the IDEA algorithm using FPGAs and Handel-C. J.UCS: Journal of Universal Computer Science, 13(3):407– 418, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
3/a_dynamically_and_
partially.

Galvao:2014:AAA

[GD14] G. R. Galvão and Z. Dias.

On alternative approaches for approximating the transposition distance. J.UCS: Journal of Universal Computer Science, 20(9):1259–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
20_9/on_alternative_
approaches_for.

Gaffney:2010:ADP

[GDW10] C. Gaffney, D. Dagger, and V. Wade. Authoring and delivering personalised simulations — an innovative approach to adaptive eLearning for soft skills. J.UCS: Journal of Universal Computer Science, 16(19): 2780–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        16_19/authoring_and_
                                        delivering_personalised.

Gecseg:2002:QPT

[Géc02] F. Gécseg. On quasiproducts of tree automata. J.UCS: Journal of Universal Computer Science, 8(2):184– 192, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        2/on_quasi_products_
                                        of.

Geyer:2008:TOB

[GFBR08] W. Geyer, R. S. Silva Filho, B. Brownholtz, and D. F. Redmiles. The trade-offs of blending synchronous and asynchronous communication services to support contextual collaboration. J.UCS: Journal of Universal Computer Science, 14(1):4–26, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        1/the_trade_offs_of.

Gaina:2012:CBL

[GFO12] D. Gaina, K. Futatsugi, and K. Ogata. Constructor-based logics. J.UCS: Journal of Universal Computer Science, 18(16):2204–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
16/constructor_based_
logics.

Gasca:2009:ASG

[GFT09] E. Gasca, J. Favela, and M. Tentori. Assisting support groups of patients with chronic diseases through persuasive computing. J.UCS: Journal of Universal Computer Science, 15(16): 3081–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_16/assisting_
support_groups_of.

Gacek:2008:MAW

[GG08] C. Gacek and C. Gamble.

Mismatch avoidance in Web services software architectures. J.UCS: Journal of Universal Computer Science, 14(8):1285– 1313, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
8/mismatch_avoidance_
in_web.

Gutwin:2008:SIC

[GGB^{+}08] C. Gutwin, S. Greenberg, R. Blum, J. Dyck, K. Tee, and G. McEwan. Supporting informal collaboration in shared-workspace groupware. J.UCS: Journal of Universal Computer Science, 14(9):1411–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_9/supporting_
informal_collaboration_
in.

Gimenez-Guzman:2008:GSM

[GGDBP^{+}08] J. M. Gimenez-Guzman, M. J. Domenech-Benlloch, V. Pla, V. Casares-Giner, and J. Martinez-Bauset. Guaranteeing seamless mobility with user redials and automatic handover retrials. J.UCS: Journal of Universal Computer Science, 14(10):1597–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
10/guaranteeing_seamless_
mobility_with.

Gonzalez-Garcia:2013:MBG

[GGMM^{+}13] M. González-García, L. Moreno, P. Martínez, R. Miñon, and J. Abascal. A modelbased graphical Editor to design accessible media players. J.UCS: Journal of Universal Computer Science, 19(18): 2656–??, ???? 2013. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_18/a_model_based_
graphical.

Gil:2008:LCR

[GGP08a] A.-B. Gil and F. J.

García-Peñalvo. Learner course recommendation in e-learning based on swarm intelligence. J.UCS: Journal of Universal Computer Science, 14(16): 2737–??, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_16/learner_course_
recommendation_in.

Gonzalez:2008:TWU

[GGP08b] M. P. González, T. Granollers, and A. Pascual. Testing Website usability in Spanish-speaking academia through heuristic evaluation and cognitive walkthroughs. J.UCS: Journal of Universal Computer Science, 14(9): 1513–??, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_9/testing_website_
usability_in.

Garcia:2011:UEV

[GGPTdP11] J. García, F. J. García-Peñalvo, R. Therón, and P. Ordoñez de Pablos. Usability evaluation of a visual modelling tool for OWL ontologies. J.UCS: Journal of Universal Computer Science, 17(9):1299–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_9/usability_
evaluation_of_a.

Goldbarg:2008:HTA

[GGS08] E. F. G. Goldbarg, M. C. Goldbarg, and C. C. Schmidt. A hybrid transgenetic algorithm for the prize collecting Steiner tree problem. J.UCS: Journal of Universal Computer Science, 14(15):2491–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_15/a_hybrid_
transgenetic_algorithm.

Garcia:2008:ABM

[GH08] A. E. Garcia and K. D. Hackbarth. Approximation to a behavioral model for estimating traffic aggregation scenarios. J.UCS: Journal of Universal Computer Science, 14(5):731–744, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
5/approximation_to_a_
behavioral.

Garcia-Herranz:2010:TUE

[GHHA10] M. García-Herranz, P. Haya, and X. Alamán. Towards a ubiquitous enduser programming system for smart spaces. J.UCS: Journal of Universal Computer Science, 16(12):1633–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_12/towards_a_
ubiquitous_end.

Garcia-Herranz:2008:ESH

[GHHE^{+}08] M. García-Herranz, P. A.

Haya, A. Esquivel, G. Montoro, and X. Alamán. Easing the smart home: Semi-automatic adaptation in perceptive environments. J.UCS: Journal of Universal Computer Science, 14(9):1529–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_9/easing_the_
smart_home.

Goudarzi:2004:OOE

[GHM04] M. Goudarzi, S. Hessabi, and A. Mycroft. Objectoriented embedded system development based on synthesis and reuse of OO-ASIPs. J.UCS: Journal of Universal Computer Science, 10(9):1123– 1155, September 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
10_9/object_oriented_
embedded_system.

Gunther:1997:RVL

[GHNT97] U. Günther, P. Hertling, R. Nicolescu, and M. Titchener. Representing variable-length codes in fixed-length T -depletion format in encoders and decoders. J.UCS: Journal of Universal Computer Science, 3(11):1207– 1225, November 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_3_11/
representing_variable_
length_codes.

Guerrero:2010:DAO

[GHO10] L. A. Guerrero, H. Horta, and S. F. Ochoa. Developing augmented objects: a process perspective. J.UCS: Journal of Universal Computer Science, 16(12):1612–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
12/developing_augmented_
objects_a.

Gocek:2006:MTC

[GHS06] P. Gocek, M. Hartmann, and H. Schleusener. Modern technologies in clientserver architecture for geo-based interactive Web portals. J.UCS: Journal of Universal Computer Science, 12(9):1208– 1214, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
9/modern_technologie_
in_client.

Giordano:1998:BQQ

[Gio98] D. Giordano. Bridging qualitative and quantitative approaches in evaluating the educational effectiveness of a shared design memory. J.UCS: Journal of Universal Computer Science, 4(4): 349–381, April 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_4/bridging_
qualitative_and_quantitative;
internal&sk=0C220489.

Girgis:2005:ATD

[Gir05] M. R. Girgis. Automatic test data generation for data flow testing using a genetic algorithm.

J.UCS: Journal of Universal Computer Science, 11(6):898–915, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_6/automatic_test_
data_generation.

Gutwin:1995:BMS

[GJAB95] C. Gutwin, M. Jones, K. Adolphe, and P. Brackett. Bringing ITS to the marketplace: a successful experiment in minimalist design. J.UCS: Journal of Universal Computer Science, 1(3):195– 200, March 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/gutwin_
1_3.

Griffiths:2012:WWS

[GJP^{+}12a] D. Griffiths, M. Johnson, K. Popat, P. Sharples, and S. Wilson. The Wookie Widget Server: a case study of piecemeal integration of tools and services. J.UCS: Journal of Universal Computer Science, 18(11):1432–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
11/the_wookie_widget_
server.

Griffiths:2012:EAW

[GJP^{+}12b] D. Griffiths, M. W. Johnson, K. Popat, P. Sharples, and S. Wilson. The educational affordances of widgets and application stores. J.UCS: Journal of Universal Computer Science, 18(16):2252–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
16/the_educational_affordances_
of.

Glaesser:1997:ASM

[GK97] U. Glaesser and R. Karges.

Abstract state machine semantics of SDL. J.UCS: Journal of Universal Computer Science, 3 (12):1382–1414, December 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_3_12/
abstract_state_machine_
semantics.

Glascock:2006:IBM

[GK06] A. P. Glascock and D. M.

Kutzik. The impact of behavioral monitoring technology on the provision of health care in the home. J.UCS: Journal of Universal Computer Science, 12(1):59–79, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_12_1/the_impact_
                                        of_behavioral.

Gorazd:2013:TSP

[GK13] T. A. Gorazd and J. Krzaczkowski. Term satisfiability problem for twoelement algebras is in QL or is NQL-complete. J.UCS: Journal of Universal Computer Science, 19(10):1375–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_19_10/term_
                                        satisfiability_problem_
                                        for.

Goldstine:2002:DCM

[GKK^{+}02] J. Goldstine, M. Kappes, C. M. R. Kintala, H. Leung, A. Malcher, and D. Wotschke. Descriptional complexity of machines with limited resources. J.UCS: Journal of Universal Computer Science, 8(2):193– 234, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        2/descriptional_complexity_
                                        of_machines.

Gaber:2005:RAM

[GKZ05] M. M. Gaber, S. Krishnaswamy, and A. Zaslavsky. Resource-aware mining of data streams. J.UCS: Journal of Universal Computer Science, 11(8):1440–1453, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_8/resource_aware_
mining_of.

Georgescu:2000:RTM

[GL00] G. Georgescu and I. Leustean.

A representation theorem for monadic Pavelka algebras. J.UCS: Journal of Universal Computer Science, 6(1):105– 111, January 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
6_1/a_representation_
theorem_for.

Gao:2011:ATR

[GL11a] L. Gao and T. Lu. Achieving transparent and realtime collaboration in Co-AutoCAD application. J.UCS: Journal of Universal Computer Science, 17(14):1887–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_14/achieving_
transparent_and_real.

Granitzer:2011:KWK

[GL11b] M. Granitzer and S. N. Lindstaedt. Knowledge work: Knowledge worker productivity, collaboration and user support. J.UCS: Journal of Universal Computer Science, 17(10):1365–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_17;   http:
                                        //www.jucs.org/jucs_
                                        17_10/knowledge_work_
                                        knowledge_worker;   http:
                                        //www.jucs.org/jucsrssfeed-
                                        issue.

Granitzer:2011:SWT

[GL11c] M. Granitzer and S. N. Lindstaedt. Semantic Web: Theory and applications. J.UCS: Journal of Universal Computer Science, 17(7):981–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_17;   http:
                                        //www.jucs.org/jucs_
                                        17_7/semantic_web_theory_
                                        and;   http://www.jucs.
                                        org/jucsrssfeed-issue.

Granitzer:2011:WAM

[GL11d] M. Granitzer and S. N. Lindstaedt. Web 2.0: Applications and mechanisms. J.UCS: Journal of Universal Computer Science, 17(4):513–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_17;   http:
//www.jucs.org/jucs_
17_4/web20_applications_
and_mechanisms;   http:
//www.jucs.org/jucsrssfeed-
issue.

Garcia:2008:MDA

[GLCV08] J. Guerrero García, C. Lemaigre, J. M. González Calleros, and J. Vanderdonckt. Model-driven approach to design user interfaces for workflow information systems. J.UCS: Journal of Universal Computer Science, 14 (19):3160–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_19/model_driven_approach_
to.

Gatteschi:2012:ESC

[GLD^{+}12] V. Gatteschi, F. Lamberti, C. Demartini, R. van Wezel, and S. Bettiol. Exploiting semantics for constructing and comparing occupational and educational-driven qualifications: the TIPTOE Project. J.UCS: Journal of Universal Computer Science, 18(1):5–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_18_1/exploiting_
                                        semantics_for_constructing.

Glesner:2003:UPC

[Gle03] S. Glesner. Using program checking to ensure the correctness of compiler implementations. J.UCS: Journal of Universal Computer Science, 9(3): 191–222, March 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        3/using_program_checking_
                                        to.

Goncalves:2011:CIS

[GlRBdSG11] M. K. Gonçalves, l. R. B. de Souza, and V. M. González. Collaboration, information seeking and communication: An observational study of software developers’ work practices. J.UCS: Journal of Universal Computer Science, 17(14): 1913–??, ???? 2011. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        17_14/collaboration_
                                        information_seeking_
                                        and.

Grosu:2000:MRE

[GLS00] R. Grosu, D. Lucanu, and Gh. Stefanescu. Mixed relations as enriched semiringal categories. J.UCS: Journal of Universal Computer Science, 6(1): 112–129, January 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_6_1/mixed_relations_
as_enriched.

Gatteschi:2011:RTE

[GLSD11] V. Gatteschi, F. Lamberti, A. Sanna, and C. Demartini. A ranking tool exploiting semantic descriptions for the comparison of EQF-based qualifications. J.UCS: Journal of Universal Computer Science, 17(7): 1060–??, ???? 2011. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_7/a_ranking_tool_
exploiting.

Gama:2005:LDT

[GM05] J. Gama and P. Medas.

Learning decision trees from dynamic data streams. J.UCS: Journal of Universal Computer Science, 11(8):1353–1366, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
8/learning_decision_
trees_from.

Gheyi:2008:ALF

[GMB08] R. Gheyi, T. Massoni, and P. Borba. Algebraic laws for feature models. J.UCS: Journal of Universal Computer Science, 14(21):3573–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_21/algebraic_
                                        laws_for_feature.

Gheyi:2011:ACF

[GMB11] R. Gheyi, T. Massoni, and P. Borba. Automatically checking feature model refactorings. J.UCS: Journal of Universal Computer Science, 17(5):684–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        17_5/automatically_checking_
                                        feature_model.

Giraldo:2008:TIM

[GMC^{+}08a] W. J. Giraldo, A. I. Molina, C. A. Collazos, M. Ortega, and M. A. Redondo. Taxonomy for integrating models in the development of interactive groupware systems. J.UCS: Journal of Universal Computer Science, 14(19):3142–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_19/taxonomy_
for_integrating_models.

Gonzalez:2008:COP

[GMC08b] S. González, K. Mens, and A. Cádiz. Contextoriented programming with the ambient object system. J.UCS: Journal of Universal Computer Science, 14(20):3307–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
20/context_oriented_
programming_with.

Griol:2012:PCL

[GMdMC12] D. Griol, J. M. Molina, A. S. de Miguel, and Z. Callejas. A proposal to create learning environments in virtual worlds integrating advanced educative resources. J.UCS: Journal of Universal Computer Science, 18(18): 2516–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_18/a_proposal_to_
create.

Garcia-Moreno:2013:SBP

[GMHGRG^{+}13] C. García-Moreno, Y. Hernández-González, M. Á. Rodríguez-García, J. A. Miñarro-Giménez, R. Valencia-García, and A. Almela. A semantic based platform for research and development projects management in the ICT domain. J.UCS: Journal of Universal Computer Science, 19(13):1914–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
13/a_semantic_based_
platform.

Gronau:2005:KCA

[GMK05] N. Gronau, C. Müller, and R. Korf. KMDL — capturing, analysing and improving knowledgeintensive business processes. J.UCS: Journal of Universal Computer Science, 11(4):452– 472, April 28, 2005. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_4/kmdl_capturing_
analysing_and.

Goncalves:2013:CAM

[GMP^{+}13] R. Gonçalves, J. Martins, J. Pereira, V. Santos, and M. Pérez Cota. Can I access my school Website? Auditing accessibility of the Portuguese teaching institutions Websites. J.UCS: Journal of Universal Computer Science, 19(18):2639–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
18/can_i_access_my.

Goranko:2003:PIN

[GMS03] V. Goranko, A. Montanari, and G. Sciavicco. Propositional interval neighborhood temporal logics. J.UCS: Journal of Universal Computer Science, 9(9):1137– 1167, September 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
9/propositional_interval_
neighborhood_temporal.

Gopinath:2000:PSB

[GN00] K. Gopinath and M. K. Krishna Narasimhan. Performance of switch blocking on multithreaded architectures. J.UCS: Journal of Universal Computer Science, 6(10):928– 947, October 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
6_10/performance_of_
switch_blocking.

Ginzboorg:2010:RRW

[GN10] P. Ginzboorg and V. Niemi.

Reachability in restricted walk on integers. J.UCS: Journal of Universal Computer Science, 16(5): 686–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_5/reachability_in_
restricted_walk.

Gecow:2005:STC

[GNP05] A. Gecow, M. Nowostawski, and M. Purvis. Structural tendencies in complex systems development and their implication for software systems. J.UCS: Journal of Universal Computer Science, 11(2):327– 356, February 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_11_2/structural_
tendencies_in_complex.

Gonzalez:2014:EDI

[GO14] L. González and G. Ortiz. An event-driven integration platform for context-aware Web services. J.UCS: Journal of Universal Computer Science, 20(8):1071– ??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_

20_8/an_event_driven_
integration.

Goetze:1995:EP

[Goe95] D. Goetze. Electronic publishing. J.UCS: Journal of Universal Computer Science, 1(4):232– 234, April 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
electronic_publishing.

Garcia-Osorio:2005:VHD

[GOF05] C. García-Osorio and C. Fyfe. Visualization of high-dimensional data via orthogonal curves. J.UCS: Journal of Universal Computer Science, 11(11):1806–1819, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
11/visualization_of_
high_dimensional.

Goldblatt:2005:ACI

[Gol05] R. Goldblatt. Axiomatic classes of intuitionistic models. J.UCS: Journal of Universal Computer Science, 11(12):1945– 1962, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_

12/axiomaatic_classes_
of.

Gomaa:2008:E

[Gom08] W. Gomaa. Expressibility in _{1}^{1}. J.UCS: Journal of Universal Computer Science, 14(10):1654–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
10/expressibility_in_
sigma.

Garcia:2013:ETA

[GOM^{+}13] A. S. García, A. Olivas, J. P. Molina, J. Martínez, P. González, and D. Martínez. An evaluation of targeting accuracy in immersive first-person shooters comparing different tracking approaches and mapping models. J.UCS: Journal of Universal Computer Science, 19(8):1086– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
8/an_evaluation_of_targeting.

Gonda:2006:NMR

[Gon06] J. Gonda. The number of the modulo n roots of the polynomial x^{v} — x^{v} and the RSA. J.UCS: Journal of Universal Computer Science, 12(9):1215– 1228, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
9/the_number_of_the.

Goos:2001:IC

[Goo01] G. Goos. Issues in compiling. J.UCS: Journal of Universal Computer Science, 7(5):410– 419, May 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
5/issues_in_compiling.

Ghiani:2010:SMU

[GP10] G. Ghiani and F. Paternò. Supporting mobile users in selecting target devices. J.UCS: Journal of Universal Computer Science, 16(15):2019–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
15/supporting_mobile_
users_in.

Garcia-Penalvo:2008:HCI

[GPA08] F. J. García-Peñalvo and J. Abascal. Humancomputer interaction research and development challenges. J.UCS: Journal of Universal Computer Science, 14(16): 2597–??, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14;   http://www.jucs.
org/jucs_14_16#;   http:
//www.jucs.org/jucs_
14_16/human_computer_
interaction_research.

Garcia-Penalvo:2011:OLM

[GPCAC11] F. J. García-Peñalvo, M. Á. Conde, M. Alier, and M. J. Casany. Opening learning management systems to personal learning environments. J.UCS: Journal of Universal Computer Science, 17(9):1222– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
9/opening_learning_management_
systems.

Garcia-Penalvo:2012:OIR

[GPCPL12] F. J. García-Peñalvo, R. Colomo-Palacios, and M. Lytras. Outcomes of international research projects on technology applied to education. J.UCS: Journal of Universal Computer Science, 18(1):1–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_18;   http:
//www.jucs.org/jucs_
18_1/outcomes_of_international_
research;   http://www.

jucs.org/jucsrssfeed-
issue.

Garcia-Penalvo:2013:TPT

[GPCZ^{+}13] F. J. García-Peñalvo, M. Á. Conde, V. Zangrando, A. García-Holgado, A. M. Seoane, M. Alier, N. Galanis, F. Brouns, H. Vogten, D. Griffiths, A. Mykowska, G. Ribeiro-Alves, and M. Minović. TRAILER Project (Tagging, Recognition, Acknowledgment of Informal Learning Experiences) a methodology to make learners’ informal learning activities visible to the institutions. J.UCS: Journal of Universal Computer Science, 19(11):1661–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
11/trailer_project_tagging_
recognition.

Garcia-Penalvo:2012:SRA

[GPFL12] F. J. García-Peñalvo, M. Alier Forment, and M. Lytras. Some reflections about service oriented architectures, Cloud computing applications, services and interoperability. J.UCS: Journal of Universal Computer Science, 18(11):1405–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_18;   http:
                                        //www.jucs.org/jucs_
                                        18_11/some_reflections_
                                        about_service;   http:
                                        //www.jucs.org/jucsrssfeed-
                                        issue.

Garrido:2013:PAE

[GPL13] J. E. Garrido, V. M. R. Penichet, and M. D. Lozano. A proposal of an architecture for educational environments. J.UCS: Journal of Universal Computer Science, 19(7):965–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        19_7/a_proposal_of_an.

Garrido:2013:ADF

[GPLV13] J. E. Garrido, V. M. Penichet, M. D. Lozano, and J. A. F. Valls. Automatic detection of falls and fainting. J.UCS: Journal of Universal Computer Science, 19(8): 1105–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_19_8/automatic_
                                        detection_of_falls.

Guttmann:2003:TSI

[GPSV03] Walter Guttmann, Helmuth Partsch, Wolfram Schulte, and Ton Vullinghs. Tool support for the interactive derivation of formally correct functional programs. J.UCS: Journal of Universal Computer Science, 9(2):173– 188, February 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
9_2/tool_support_for_
the.

Garcia-Penalvo:2013:OIR

[GPVILN13] F. J. García-Peñalvo, J. Á. Velázquez-Iturbide, and M. Llamas-Nistal. Outcomes of international research projects on technology applied to education. J.UCS: Journal of Universal Computer Science, 19(11):1496–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_19;   http:
//www.jucs.org/jucs_
19_11/international_
research_projectsoon.

Gargantini:2001:ABT

[GR01] A. Gargantini and E. Riccobene. ASM-based testing: Coverage criteria and automatic test sequence. J.UCS: Journal of Universal Computer Science, 7(11):1050–1067, November 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_7_
                                        11/asm_based_testing_
                                        coverage.

Gams:2002:TFA

[GR02] E. Gams and S. Reich. The TrailTRECer framework: Applying open hypermedia concepts to trails. J.UCS: Journal of Universal Computer Science, 8(10):913–923, October 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        8_10/the_trailtrecer_
                                        framework_applying.

Gago:2014:MTT

[GR14] J. Santos Gago and L. Anido Rifón. Methodologies, technologies and tools enabling e-government. J.UCS: Journal of Universal Computer Science, 20(11):1522–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_20;   http:
                                        //www.jucs.org/jucs_
                                        20_10/methodologies_
                                        technologies_and_tools.

Granvilliers:1998:SNB

[Gra98]  L. Granvilliers.  A

symbolic-numerical branch and prune algorithm for solving non-linear polynomial systems. J.UCS: Journal of Universal Computer Science, 4 (2):125–146, February 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_2/a_
symbolic_numerical_branch.

Goncalves:2015:IFC

[GRC15] M. J. A. Gonçalves, Á. Rocha, and M. Pérez Cota. Interoperability framework for competences and learning outcomes. J.UCS: Journal of Universal Computer Science, 21(8):1042– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
8/interoperability_framework_
for_competences.

Gutiirrez-Rojas:2014:AAT

[GRCGK14] I. Gutiírrez-Rojas, R. M.

Crespo-García, and C. Delgado Kloos. Adapting an awareness tool for massive courses: the case of ClassON. J.UCS: Journal of Universal Computer Science, 20(1):24– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_20_
                                        1/adapting_an_awareness_
                                        tool.

Gregg:2008:EIE

[Gre08] D. G. Gregg. Exploring information extraction resilience. J.UCS: Journal of Universal Computer Science, 14(11): 1911–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_11/exploring_
                                        information_extraction_
                                        resilience.

Garrido:2013:ASP

[GRGN13] P. Castro Garrido, I. Luque Ruiz, and M. Á. Gómez-Nieto. An alert system for people monitoring based on multi-agents using maps. J.UCS: Journal of Universal Computer Science, 19(9):1257– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_19_
                                        9/an_alert_system_for.

Gil:2008:SOA

[GRGPL08] A.-B. Gil, R. Rodríguez, F. J. García-Peñalvo, and R. López. SHARP online: An adaptive hypermedia system applied to mathematical problem solving. J.UCS: Journal of Universal Computer Science, 14(19):3099–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_14_19/sharp_
online_an_adaptive.

Gutierrez-Rivas:2015:SMC

[GRHMM^{+}15] J. L. Gutiérrez-Rivas, S. Holmbacka, M. Míndez-Macías, W. Lund, S. Lafond, J. Lilius, and J. Díaz-Alonso. Safe motor controller in a mixedcritical environment with runtime updating capabilities. J.UCS: Journal of Universal Computer Science, 21(2):177–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_2/safe_motor_
controller_in.

Grozea:2000:FEP

[Gro00] C. Grozea. Free-extendible prefix-free sets and an extension of the Kraft-Chaitin theorem. J.UCS: Journal of Universal Computer Science, 6(1): 130–135, January 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_6_1/free_extendible_
prefix_free.

Groves:2009:RAN

[Gro09] L. Groves. Reasoning about nonblocking concurrency. J.UCS: Journal of Universal Computer Science, 15(1):72– ??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        15_1/reasoning_about_
                                        nonblocking_concurrency.

Gargantini:2008:MBL

[GRS08] A. Gargantini, E. Riccobene, and P. Scandurra. A metamodel-based language and a simulation engine for abstract state machines. J.UCS: Journal of Universal Computer Science, 14(12):1949–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        12/a_metamodel_based_
                                        language.

Gurevich:1997:RAS

[GS97] Y. Gurevich and M. Spielmann. Recursive abstract state machines. J.UCS: Journal of Universal Computer Science, 3(4): 233–246, April 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs_3_4/recursiv_

abstract;   http://www.
jucs.org/jucs_3_4/recursiv_
abstract; internal&sk=
05460486.

Giaglis:2012:DEP

[GS12] G. M. Giaglis and D. Spinellis. Division of effort, productivity, quality, and relationships in FLOSS virtual teams: Evidence from the FreeBSD project. J.UCS: Journal of Universal Computer Science, 18(19): 2625–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_19/division_
of_effort_productivity.

Gluz:2016:TSR

[GSdSB16] J. C. Gluz, E. L. Silveira, L. R. Jardim da Silva, and J. L. V. Barbosa. Towards a semantic repository for learning objects: Design and evaluation of core services. J.UCS: Journal of Universal Computer Science, 22(1):16– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_22_
1/towards_a_semantic_
repository.

Gutierrez-Santos:2010:APS

[GSMBFPK10] S. Gutierrez-Santos, J. Mayor-Berzal, C. Fernandez-Panadero, and C. Delgado Kloos. Authoring of probabilistic sequencing in adaptive hypermedia with Bayesian networks. J.UCS: Journal of Universal Computer Science, 16(19):2801–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_19/authoring_
of_probabilistic_sequencing.

Gomez-Sanz:2004:MDM

[GSP04] J. Gómez-Sanz and J. Pavón. Methodologies for developing multi-agent systems. J.UCS: Journal of Universal Computer Science, 10(4):359–374, April 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
4/methodologies_for_
developing_multi.

Gutierrez-Santos:2008:ACR

[GSPK08] S. Gutierrez-Santos, A. Pardo, and C. Delgado Kloos. Authoring courses with rich adaptive sequencing for IMS learning design. J.UCS: Journal of Universal Computer Science, 14(17):2819–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
17/authoring_courses_
with_rich.

Ginsburg:1999:ADR

[GSS99] S. Ginsburg, N. C. Shu, and D. A. Simovici. Automatic data restructuring. J.UCS: Journal of Universal Computer Science, 5(4):243– 286, April 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
4/automatic_data_restructuring.

Gurevich:1997:FDR

[GSW97] Y. Gurevich, N. Soparkar, and Ch. Wallace. Formalizing database recovery. J.UCS: Journal of Universal Computer Science, 3(4):320– 340, April 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs_3_4/formalizing_
database;   http://www.
jucs.org/jucs_3_4/formalizing_
database; internal&sk=
05460486.

Ganter:2004:FCA

[GSW04] B. Ganter, G. Stumme, and R. Wille. Formal concept analysis: Theory and applications. J.UCS: Journal of Universal Computer Science, 10(8):

926, August 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/free;
                                         http://www.jucs.org/
                                        jucs_10_8/formal_concept_
                                        analysis_theory.

Grubisic:2015:ACL

[GSZ15] A. Grubisić, S. Stankov, and B. Zitko. Adaptive courseware: A literature review. J.UCS: Journal of Universal Computer Science, 21(9):1168– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_21_
                                        9/adaptive_courseware_
                                        a_literature.

Gurevich:2001:PUE

[GT01] Y. Gurevich and N. Tillmann. Partial updates: Exploration. J.UCS: Journal of Universal Computer Science, 7 (11):917–951, November 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        7_11/partial_updates_
                                        exploration.

Gervasi:2010:CSA

[GTGT10] O. Gervasi, K. Tan, M. Gavrilova, and D. Taniar. Computational science and its applications. J.UCS: Journal of Universal Computer Science, 16(6): 889–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16;   http://www.jucs.
org/jucs_16_06/computational_
science_and_its;   http:
//www.jucs.org/jucs_
16_6#.

Gunther:1996:DCS

[Gün96] U. Günther. Data compression and serial communication with generalized T-codes. J.UCS: Journal of Universal Computer Science, 2(11): 769–795, November 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
data_compression_and_
serial_communiction.

Gunther:2002:NMO

[Gün02] J. Günther. The new mobility of our society caused by telecommunications. J.UCS: Journal of Universal Computer Science, 8(5):456– 481, May 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
5/the_new_mobility_of.

Gupta:2001:DNA

[Gup01]  S. Gupta.  Determin-

ism, nondeterminism, alternation, and counting. J.UCS: Journal of Universal Computer Science, 7(9):816–825, September 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_7_9/determinism_
                                        nondeterminism_alternation_
                                        and.

Gutl:2008:EMI

[Güt08] C. Gütl. Enhancements of meeting information management and application for knowledge access and learning activities. J.UCS: Journal of Universal Computer Science, 14(10):1625–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        10/enhancements_of_meeting_
                                        information.

Gutl:2012:MECa

[Güt12a] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 18(2): 142–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_

18;   http://www.jucs.
org/jucs_18_2/editorial;
http://www.jucs.org/
jucsrssfeed-issue.

Gutl:2012:MECb

[Güt12b] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 18 (5):598–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_18; http://
www.jucs.org/jucs_18_
5/editorial;   http:/
/www.jucs.org/jucsrssfeed-
issue.

Gutl:2012:MECc

[Güt12c] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 18 (10):1236–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_18; http://
www.jucs.org/jucs_18_
10/editorial;   http:
//www.jucs.org/jucsrssfeed-
issue.

Gutl:2012:MECd

[Güt12d] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 18 (14):1905–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_18; http://
www.jucs.org/jucs_18_
14/editorial;   http:
//www.jucs.org/jucsrssfeed-
issue.

Gutl:2012:MECe

[Güt12e] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 18 (16):2203–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18;   http://www.jucs.
org/jucs_18_16/editorial.

Gutl:2012:MECf

[Güt12f] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 18 (20):2747–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_18; http://
www.jucs.org/jucs_18_
20/editorial.

Gutl:2013:MECa

[Güt13a] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 19 (1):1–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_19; http://
www.jucs.org/jucs_19_
1/editorial.

Gutl:2013:MECb

[Güt13b] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 19 (3):281–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_19; http://
www.jucs.org/jucs_19_
3/editorial.

Gutl:2013:MECc

[Güt13c] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 19 (6):728–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_19; http://
www.jucs.org/jucs_19_
6/editorial.

Gutl:2013:MECd

[Güt13d] C. Gütl. Managing Editor’s column. J.UCS:

Journal of Universal Computer Science, 19 (10):1350–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        19_10/editorial.

Gutl:2013:MECe

[Güt13e] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 19 (12):1701–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_19; http://
                                        www.jucs.org/jucs_19_
                                        12/editorial.

Gutl:2013:MECf

[Güt13f] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 19 (15):2172–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_19; http://
                                        www.jucs.org/jucs_19_
                                        15/editorial.

Gutl:2014:MECa

[Güt14a] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 20 (2):77–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_20; http://
www.jucs.org/jucs_20_
2/editorial.

Gutl:2014:MECb

[Güt14b] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 20 (4):461–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_20; http://
www.jucs.org/jucs_20_
4/editorial.

Gutl:2014:MECc

[Güt14c] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 20(9): 1152–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_20; http://
www.jucs.org/jucs_20_
9/editorial.

Gutl:2014:MEC

[Güt14d] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 20 (14):1875–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_20; http://
                                        www.jucs.org/jucs_20_
                                        14/editorial.

Gutl:2015:MECa

[Güt15a] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 21 (2):175–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_21; http://
                                        www.jucs.org/jucs_21_
                                        2/editorial.

Gutl:2015:MECb

[Güt15b] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 21 (4):502–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_21; http://
                                        www.jucs.org/jucs_21_
                                        4/editorial.

Gutl:2015:MECc

[Güt15c] C. Gütl. Managing Editor’s column. J.UCS:

Journal of Universal Computer Science, 21 (7):890–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_21; http://
www.jucs.org/jucs_21_
7/editorial.

Gutl:2015:MECd

[Güt15d] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 21(9): 1106–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_21; http://
www.jucs.org/jucs_21_
9/editorial.

Gutl:2015:MECe

[Güt15e] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 21 (12):1535–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_21; http://
www.jucs.org/jucs_21_
12/editorial.

Gutl:2016:MECa

[Güt16a] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 22 (2):159–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_22; http://
                                        www.jucs.org/jucs_22_
                                        02/editorial.

Gutl:2016:MECb

[Güt16b] C. Gütl. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 22 (7):894–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_22; http://
                                        www.jucs.org/jucs_22_
                                        07/editorial.

Georgescu:2000:NAC

[GV00] H. Georgescu and C. Vertan. A new approach to communicating Xmachine systems. J.UCS: Journal of Universal Computer Science, 6(5): 490–502, May 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_6_
                                        5/a_new_approach_to.

Garcia-Vazquez:2010:ABA

[GVRT^{+}10] J. P. García-Vázquez, M. D. Rodríguez, M. E. Tentori, D. Saldaña, Á. G. Andrade, and A. N. Espinoza. An agent-based architecture for developing activity-aware systems for assisting elderly. J.UCS: Journal of Universal Computer Science, 16(12):1500–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_12/an_agent_
based_architecture.

Garcia:2011:ECP

[GVT11] A. C. B. Garcia, A. S.

Vivacqua, and T. C. Tavares. Enabling crowd participation in governmental decision-making. J.UCS: Journal of Universal Computer Science, 17(14):1931–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_14/enabling_
crowd_participation_
in.

Glasser:2003:RFC

[GW03] C. Glaßer and G. Wechsung. Relativizing function classes. J.UCS: Journal of Universal Computer Science, 9(1):34–50, January 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        1/relativizing_function_
                                        classes.

Gauch:1996:PIF

[GWG96] S. Gauch, G. Wang, and M. Gomez. ProFusion: Intelligent fusion from multiple, distributed search engines. J.UCS: Journal of Universal Computer Science, 2(9):637– 649, September 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_2_
                                        9/profusion_intelligent_
                                        fusion_from.

Gao:2005:ETC

[GWW05] L. Gao, M. Wang, and X. S. Wang. Evaluating trigger conditions on streaming time series with user-given quality requirements. J.UCS: Journal of Universal Computer Science, 11(8):1397–1410, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_11_
                                        8/evaluating_trigger_
                                        conditions_on.

Goggins:2015:LAS

[GXC^{+}15] S. Goggins, W. Xing, X. Chen, B. Chen, and B. Wadholm. Learning analytics at “Small” scale: Exploring a complexitygrounded model for assessment automation. J.UCS: Journal of Universal Computer Science, 21(1): 66–??, ???? 2015. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
21_1/learning_analytics_
at_small.

Gong:2009:IGA

[GYY09] D. Gong, X. Yao, and J. Yuan. Interactive genetic algorithms with individual fitness not assigned by human. J.UCS: Journal of Universal Computer Science, 15 (13):2446–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_13/interactive_genetic_
algorithms_with.

Hofer-Alfeis:2003:EIK

[HA03] J. Hofer-Alfeis. Effective integration of knowledge management into the business starts with a top-down knowledge strategy. J.UCS: Journal of Universal Computer Science, 9(7):719– 728, July 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
7/effective_integration_
of_knowledge.

Hatzikou:2010:LID

[HA10] M. Hatzikou and I.-P. Agiovlasitis. Leveraging ICT deployment and integration in a public organization aged 176 years: a Greek case study. J.UCS: Journal of Universal Computer Science, 16(8):1102–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_16_8/leveraging_
                                        ICT_deployment_and.

Huang:2013:GAS

[HA13] W. Huang and L. Alem. Gesturing in the air: Supporting full mobility in remote collaboration on physical tasks. J.UCS: Journal of Universal Computer Science, 19(8):1158–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_19_8/gesturing_
                                        in_the_air.

Hernandez:2015:MAL

[HAFS15] Y. Hernández, G. Arroyo-Figueroa, and L. E. Sucar. A model of affect and learning for intelligent tutors. J.UCS: Journal of Universal Computer Science, 21(7):912– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
7/a_model_of_affect.

Hoang:2013:AOE

[HAI13] L. P. Hoang and N. Arch-Int. Assessment of openended questions using a multidimensional approach for the interaction and collaboration of learners in E-learning environments. J.UCS: Journal of Universal Computer Science, 19(7):932– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
7/assessment_of_open_
ended.

Hall:2007:RBF

[Hal07] A. Hall. Realising the benefits of formal methods. J.UCS: Journal of Universal Computer Science, 13(5):669–678, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_13_5/realising_
the_benefits_of.

Han:2010:LNM

[Han10] Y.-S. Han. On the linear number of matching substrings. J.UCS: Journal of Universal Computer Science, 16(5):715–??, ???? 2010. CODEN ????

ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_16_5/on_the_
                                        linear_number.

Harary:2000:AGH

[Har00] F. Harary. The automorphism group of a hypercube. J.UCS: Journal of Universal Computer Science, 6(1):136– 138, January 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        6_1/the_automorphism_
                                        group_of.

Harrison:2007:FPV

[Har07] J. Harrison. Floatingpoint verification. J.UCS: Journal of Universal Computer Science, 13(5): 629–638, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        13_5/floating_point_
                                        verification.

Hasebrook:2001:LLO

[Has01] J. Hasebrook. Learning in the learning organization. J.UCS: Journal of Universal Computer Science, 7(6):472–487, June 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/

jucs_7_6/learning_in_
the_learning.

Hasebrook:2002:COI

[Has02] J. P. Hasebrook. Cooperative and interactive distance learning: Application of team-oriented and selective learning strategies in a European bank. J.UCS: Journal of Universal Computer Science, 8(9):834– 847, September 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
8_9/co_operative_and_
interactive.

Heileman:2007:COG

[HAS^{+}07] G. L. Heileman, C. T.

Abdallah, W. Shu, C. G. Christodoulou, and D. Knotts. Creating online graduate engineering degrees at the University of New Mexico. J.UCS: Journal of Universal Computer Science, 13(7):1002– 1011, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_7/creating_online_
graduate_engineering.

Havea:2005:FSS

[Hav05] R. S. Havea. On firmness of the state space and positive elements of a Banach algebra. J.UCS: Journal of Universal Computer Science, 11(12): 1963–1969, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_12/on_firmness_of_
the.

Heitmeyer:2000:ASR

[HB00] C. Heitmeyer and R. Bharadwaj. Applying the SCR requirements method to the light control case study. J.UCS: Journal of Universal Computer Science, 6(7):650– 678, July 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
6_7/applying_the_scr_
requirements.

Holub:2011:IUB

[HB11] M. Holub and M. Bieliková. An inquiry into the utilization of behavior of users in personalized Web. J.UCS: Journal of Universal Computer Science, 17(13):1830–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
13/an_inquiry_into_the.

Hervas:2013:AAL

[HB13] R. Hervás and J. Bravo.

Ambient assisted living: Home care. J.UCS: Journal of Universal Computer Science, 19(9):1197– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
9/ambient_assisted_living_
home.

Hervas:2010:CMB

[HBF10] R. Hervás, J. Bravo, and J. Fontecha. A context model based on ontological languages: a proposal for information visualization. J.UCS: Journal of Universal Computer Science, 16(12):1539–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_12/a_context_
model_based.

Hervas:2013:AAA

[HBFV13] R. Hervás, J. Bravo, J. Fontecha, and V. Villarreal. Achieving adaptive augmented reality through ontological context-awareness applied to AAL scenarios. J.UCS: Journal of Universal Computer Science, 19(9):1334–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_

                                        9/achieving_adaptive_
                                        augmented_reality.

Hester:1998:BFE

[HBI98] A. Hester, R. Borges, and R. Ierusalimschy. Building flexible and extensible Web applications with Lua. J.UCS: Journal of Universal Computer Science, 4(9):748– 762, September 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_9/building_
                                        flexible_and_extensible.

Hartog:2012:DLR

[HBT12] R. J. Hartog, A. J. Beulens, and J. Tramper. Digital learning resources in higher education: Designing for largescale use. J.UCS: Journal of Universal Computer Science, 18(16):2274–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_18_
                                        16/digital_learning_
                                        resources_in.

Hendrix:2008:SMA

[HC08] M. Hendrix and A. Cristea. A spiral model for adding automatic, adaptive authoring to adaptive hypermedia. J.UCS: Journal of Universal Computer Science, 14(17):2799–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
17/a_spiral_model_for.

Harchay:2015:CAA

[HCBB15] A. Harchay, L. Cheniti-Belcadhi, and R. Braham. A context-aware approach for personalized mobile self-assessment. J.UCS: Journal of Universal Computer Science, 21(8):1061–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
21_8/a_context_aware_
approach.

Herzeel:2010:EIF

[HCD10] C. Herzeel, P. Costanza, and T. D’Hondt. An extensible interpreter framework for software transactional memory. J.UCS: Journal of Universal Computer Science, 16(2): 221–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_2/an_extensible_interpreter_
framework.

Huang:2009:LBC

[HCH^{+}09] C.-J. Huang, I.-F. Chen, K.-W. Hu, H.-Y. Shen, Y.-J. Chen, and D.-X. Yang. A load balancing and congestionavoidance routing mechanism for teal-time traffic over vehicular networks. J.UCS: Journal of Universal Computer Science, 15(13):2506–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
13/a_load_balancing_
and.

Hwang:2011:LST

[HCK11] M. Hwang, D. Choi, and P. Kim. Least slack time rate first: an efficient scheduling algorithm for pervasive computing environment. J.UCS: Journal of Universal Computer Science, 17(6):912–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_17_6/least_
slack_time_rate.

Herranz:2014:GDF

[HCPdASY14] E. Herranz, R. Colomo-Palacios, A. de Amescua Seco, and M. Yilmaz. Gamification as a disruptive factor in software process improvement initiatives. J.UCS: Journal of Universal Computer Science, 20(6):885– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
6/gamification_as_a_
disruptive.

Hockemeyer:2003:ACP

[HCWA03] C. Hockemeyer, O. Conlan, V. Wade, and D. Albert. Applying competence prerequisite structures for eLearning and skill management. J.UCS: Journal of Universal Computer Science, 9(12):1428–1436, December 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
12/applying_competence_
prerequisite_structures.

Hefke:2004:FSI

[Hef04] M. Hefke. A framework for the successful introduction of KM using CBR and Semantic Web technologies. J.UCS: Journal of Universal Computer Science, 10(6):731–739, June 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
6/a_framework_for_the.

Hegner:2010:IRD

[Heg10] S. J. Hegner. Internal representation of database views. J.UCS: Journal of Universal Computer Science, 16(20):2956–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_16_20/internal_
                                        representation_of_database.

Heinrichmeyer:1996:ADH

[Hei96] F. Heinrichmeyer. ASIC design at home. J.UCS: Journal of Universal Computer Science, 2(6): 474–483, June 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/ASIC_
                                        design_at_home.

Heitmeyer:2007:FMS

[Hei07] C. L. Heitmeyer. Formal methods for specifying, validating, and verifying requirements. J.UCS: Journal of Universal Computer Science, 13(5):607–618, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        13_5/formal_methods_
                                        for_specifying.

Helic:2007:FRL

[Hel07] D. Helic. Formal representations of learning scenarios: a methodology to configure E-learning systems. J.UCS: Journal of Universal Computer Science, 13(4):504–530, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_13_4/formal_
representations_of_learning.

Hemaspaandra:1999:PPT

[Hem99] E. Hemaspaandra. On the power of positive Turing reductions. J.UCS: Journal of Universal Computer Science, 5(12):828– 832, December 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
12/on_the_power_of.

Hempel:2006:RAC

[Hem06] H. Hempel. Randomized algorithms and complexity theory. J.UCS: Journal of Universal Computer Science, 12(6):746– 761, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_12_6/randomized_
algorithms_and_complexity.

Henin:1998:ELH

[Hen98] I. Henin. Evaluation of on-line help. J.UCS: Journal of Universal Computer Science, 4(4):449– 460, April 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_4/evaluation_
of_on_line;   internal&
sk=0C220489.

Hertling:1996:DOW

[Her96]  P. Hertling.  Disjunc-

tive omega-words and real numbers. J.UCS: Journal of Universal Computer Science, 2(7):549– 568, July 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
disjunctive_omega_words_
and_real_numbers.

Hertling:1997:SFC

[Her97]  P. Hertling.  Surjec-

tive functions on computably growing Cantor sets. J.UCS: Journal of Universal Computer Science, 3(11):1226–1240, November 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_3_11/
surjective_functions_
on_cmputably.

Hertling:2002:SNN

[Her02] Peter Hertling. Simply normal numbers to different bases. J.UCS: Journal of Universal Computer Science, 8(2): 235–242, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_8_2/simply_
normal_numbers_to.

Hertling:2005:NSB

[Her05]  P. Hertling.  Non-

random sequences between random sequences. J.UCS: Journal of Universal Computer Science, 11(12):1970–1985, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_11_12/nonrandom_
sequences_between_random.

Herranz:2009:IHA

[Her09] J. Herranz. Ideal homogeneous access structures constructed from graphs. J.UCS: Journal of Universal Computer Science, 15(14):2881–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
14/ideal_homogeneous_
access_structures.

Haglin:2001:MML

[HF01] D. J. Haglin and R. W.

Ford. The messageminimizing load redistribution problem. J.UCS: Journal of Universal Computer Science, 7(4):

291–306, April 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_7_4/the_message_
minimizing_load.

Herskovic:2013:PDD

[HFIBJ13] V. Herskovic, C. Fuentes, R. Ibarra, and J. Bustos-Jiménez. Pick & drag & drop: Augmented reality for multiple file sharing. J.UCS: Journal of Universal Computer Science, 19(8):1140–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
8/pick_and_drag_and.

Hana:2008:CSR

[HFOB08] R. O. Abu Hana, C. O. Freitas, L. S. Oliveira, and F. Bortolozzi. Crime scene representation (2D, 3D, stereoscopic projection) and classification. J.UCS: Journal of Universal Computer Science, 14(18):2953–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_14_18/crime_
scene_representation_
2d.

Hurtado:2011:ECL

[HG11] C. Hurtado and L. A.

Guerrero. Enhancement of collaborative learning activities using portable devices in the classroom. J.UCS: Journal of Universal Computer Science, 17(2):332–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_2/enhancement_of_
collaborative_learning.

Hernandez-Garcia:2011:EAI

[HGIPCPPM11] Á. Hernández-García, S. Iglesías-Pradas, J. Chaparro-Peláez, and F. Pascual-Miguel. Exploring the attitudes and intentions of non-shoppers in the acceptance of e-commerce. J.UCS: Journal of Universal Computer Science, 17(9):1314–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_9/exploring_
the_attitudes_and.

Holzinger:2008:SIM

[HGMT08] A. Holzinger, R. Geierhofer, F. Mödritscher, and R. Tatzl. Semantic information in medical information systems: Utilization of text mining techniques to analyze medical diagnoses. J.UCS: Journal of Universal Computer Science, 14(22):

3781–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_22/semantic_
                                        information_in_medical.

Hlavacs:2008:EZU

[HGS^{+}08] H. Hlavacs, W. Gansterer, H. Schabauer, J. Zottl, M. Petraschek, T. Hoeher, and O. Jung. Enhancing ZRTP by using computational puzzles. J.UCS: Journal of Universal Computer Science, 14(5):693–716, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_5/enhancing_
                                        zrtp_by_using.

Hiermann:2003:PKB

[HH03] W. Hiermann and M. Höfferer. A practical knowledgebased approach to skill management and personal development. J.UCS: Journal of Universal Computer Science, 9 (12):1398–1409, December 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        12/a_practical_knowledge_
                                        based.

Hemaspaandra:1998:QOP

[HHH98] E. Hemaspaandra, L. A.

Hemaspaandra, and H. Hempel. Query order and the polynomial hierarchy. J.UCS: Journal of Universal Computer Science, 4(6): 574–588, June 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_6/query_
order_and_the;   internal&
sk=0C220489.

Hong:2002:BFR

[HHH^{+}02] D. W.-K. Hong, C. S.

Hong, Y. J. Hyoung, D.- S. Yun, and W.-S. Kim. Bounded flooding routing algorithm for provisioning the globally optimal route in a hierarchical ATM network. J.UCS: Journal of Universal Computer Science, 8(7):698– 721, July 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
8_7/bounded_flooding_
routing_algorithm.

Ho:2009:CEH

[HHHX09] Y. H. Ho, A. H. Ho, K. A. Hua, and F. Xie. Cooperation enforcement in a highly dynamic mobile ad hoc network. J.UCS: Journal of Universal Computer Science, 15(5):1090–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15_
                                        5/cooperation_enforcement_
                                        in_a.

Hsiao:2002:WOC

[HHY02] H. K. Hsiao, C. C. Huang, and S. S. Yu. Word operation closure and primitivity of languages. J.UCS: Journal of Universal Computer Science, 8(2):243– 256, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        2/word_operation_closure_
                                        and.

Horvath:1999:DUP

[HI99] S. Horváth and M. Ito. Decidable and undecidable problems of primitive words, regular and context-free languages. J.UCS: Journal of Universal Computer Science, 5(9):532–541, September 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_5_
                                        9/decidable_and_undecidable_
                                        problems.

Hart:2000:IPC

[HI00] W. E. Hart and S. Istrail. Invariant patterns in crystal lattices: Implications for protein folding algorithms. J.UCS: Journal of Universal Computer Science, 6(6):560– 581, June 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
6/invariant_patterns_
in_crystal.

Hussein:2013:PRS

[HIGM13] W. Hussein, R. M. Ismail, T. F. Gharib, and M. G. M. Mostafa. A personalized recommender system based on a hybrid model. J.UCS: Journal of Universal Computer Science, 19(15): 2224–??, ???? 2013. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_15/a_personalized_
recommender_system.

Hemaspaandra:1997:PTM

[HJRW97] L. A. Hemaspaandra, Z. Jiang, J. Rothe, and O. Watanabe. Polynomialtime multi-selectivity. J.UCS: Journal of Universal Computer Science, 3(3): 197–229, March 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_3_
3/polynomial_time.

Hermans:2015:FPA

[HJVK15] H. Hermans, J. Janssen, H. Vogten, and R. Koper. Flexible provisioning adult learners. J.UCS: Journal of Universal Computer Science, 21(2):206– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_21_
                                        2/flexible_provisioning_
                                        adult_learners.

Hundewale:2007:EEN

[HJZ07] N. Hundewale, S. Jung, and A. Zelikovsky. Energy efficient node caching and load balancing enhancement of reactive ad hoc routing protocols. J.UCS: Journal of Universal Computer Science, 13(1):110–132, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_13_1/energy_
                                        efficient_node_caching.

Honkala:1995:FCL

[HK95] Juha Honkala and Werner Kuich. On four classes of Lindenmayerian power series. J.UCS: Journal of Universal Computer Science, 1(2):131– 135, February 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
2/on_four_classes_of.

Hasan:2014:TFL

[HK14] O. Hasan and S. A.

Khayam. Towards formal linear cryptanalysis using HOL4. J.UCS: Journal of Universal Computer Science, 20(2):193– ??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
20_2/towards_formal_
linear_cryptanalysis.

Hanzlik:2015:RIS

[HK15] L. Hanzlik and M. Kutylowski. Restricted identification secure in the extended Canetti–Krawczyk model. J.UCS: Journal of Universal Computer Science, 21(3):419–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_3/restricted_
identification_secure_
in.

Ho:2013:GBK

[HKK13] T. P. Ho, H.-S. Kang, and S.-R. Kim. Graphbased KNN algorithm for spam SMS detection. J.UCS: Journal of Universal Computer Science, 19(16):2404–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_19_16/graph_
                                        based_knn_algorithm.

Han:2008:ESL

[HKKvP08] P. Han, G. Kortemeyer, B. J. Krämer, and C. von Prümmer. Exposure and support of latent social networks among learning object repository users. J.UCS: Journal of Universal Computer Science, 14(10):1717–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_10/exposure_
                                        and_support_of.

Hajdu:2006:MPI

[HKL^{+}06] A. Hajdu, J. Kormos, Z. Lencse, L. Trón, and M. Emri. The ‘MEDIP-Platform Independent Software System for Medical Image Processing’ Project. J.UCS: Journal of Universal Computer Science, 12(9):1229– 1239, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12_
                                        9/the_medip_platform_
                                        independent.

Halang:1996:IMB

[HKS96] W. A. Halang, B. J. Krämer, and J. Schormann. Integrated multimediabased distance teaching of information technology. J.UCS: Journal of Universal Computer Science, 2(6):443–455, June 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
integrated_multimedia_
based_distance_teaching.

Hajdu:2006:ANS

[HKTV06] A. Hajdu, J. Kormos, T. Tóth, and K. Veréb. Applications of neighborhood sequence in image processing and database retrieval. J.UCS: Journal of Universal Computer Science, 12(9):1240– 1253, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_9/applications_of_
neighborhood_sequence.

Hauff:1996:EVT

[HL96] M. Hauff and W. Laaser.

Educational video and TV in distance education — production and design aspects. J.UCS: Journal of Universal Computer Science, 2(6):456– 473, June 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/

                                        educational_video_and_
                                        TV_in_distance_education.

Honkala:2003:IZU

[HL03] I. Honkala and A. Lobstein. On identification in ZZ^{2} using translates of given patterns. J.UCS: Journal of Universal Computer Science, 9(10):1204–1219, October 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        10/on_identification_
                                        in_zz2.

Hartmann:2009:WFD

[HL09] S. Hartmann and S. Link. Weak functional dependencies: Full propositional expressiveness for the database practitioner. J.UCS: Journal of Universal Computer Science, 15(1):112–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        15_1/weak_functional_
                                        dependencies_full.

Hwang:2008:CPK

[HLC08] Y. H. Hwang, J. K. Liu, and S. S. Chow. Certificateless public key encryption secure against malicious KGC attacks in the standard model. J.UCS: Journal of Universal Computer Science, 14(3):463–480, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_3/certificateless_
public_key_encryption.

Huang:2011:ISL

[HLCL11] Y.-L. Huang, F.-Y. Leu, C.-H. Chiu, and I.-L. Lin. Improving security levels of IEEE802.16e authentication by involving Diffie–Hellman PKDS. J.UCS: Journal of Universal Computer Science, 17(6):891–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_6/improving_
security_levels_of.

Hernandez-Leo:2007:FCA

[HLHD^{+}07] D. Hernández-Leo, A. Harrer, J. M. Dodero, J. I. Asensio-Pérez, and D. Burgos. A framework for the conceptualization of approaches to ‘Create-by-Reuse’ of learning design solutions. J.UCS: Journal of Universal Computer Science, 13(7):991– 1001, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
7/a_framework_for_the.

Holmer:2009:DCC

[HLK09a] T. Holmer, S. Lukosch, and V. Kunz. Diminishing chat confusion by multiple visualizations. J.UCS: Journal of Universal Computer Science, 15(16):3139–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15_
                                        16/diminishing_chat_
                                        confusion_by.

Hwang:2009:MWS

[HLK09b] J. Hwang, Y. Lee, and S. Kim. Modelling Weblog success: Case of Korea. J.UCS: Journal of Universal Computer Science, 15(8):1589–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_15_8/modelling_
                                        weblog_success_case.

Hernandez-Leo:2012:SOC

[HLNA^{+}12] D. Hernández-Leo, R. Nieves, E. Arroyo, A. Rosales, J. Melero, and J. Blat. SOS: Orchestrating collaborative activities across digital and physical spaces using wearable signaling devices. J.UCS: Journal of Universal Computer Science, 18(15):2165–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
15/sos_orchestrating_
collaborative_activities.

Havlik:2013:MSR

[HLP^{+}13] J. Havlik, L. Lhotska, J. Parak, J. Dvorak, Z. Horcik, and M. Pokorny. A modular system for rapid development of telemedical devices. J.UCS: Journal of Universal Computer Science, 19(9):1242–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_9/a_modular_
system_for.

Hernandez-Leo:2015:AES

[HLS15] D. Hernández-Leo and L. Vicent Safont. Aligned, embedded and scalable assessment technologies. J.UCS: Journal of Universal Computer Science, 21(8):997–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_21;   http:
//www.jucs.org/jucs_
21_8/aligned_embedded_
and_scalable.

Hadjiefthymiades:1999:SWA

[HM99] S. Hadjiefthymiades and L. Merakos. A survey of Web architectures for wireless communication environments. J.UCS: Journal of Universal Computer Science, 5(7):390–417, July 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_5_7/a_survey_of_
web.

Heinrich:2000:ADC

[HM00] E. Heinrich and H. Maurer. Active documents: Concept, implementation and applications. J.UCS: Journal of Universal Computer Science, 6(12):1197–1202, December 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
12/active_documents_
concept_implementation.

Haas:2001:MRC

[HM01] W. Haas and H. Mayer. MPEG and its relevance for content-based multimedia retrieval. J.UCS: Journal of Universal Computer Science, 7(6): 530–547, June 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
6/mpeg_and_its_relevance.

Hinum:2005:GII

[HMA^{+}05] K. Hinum, S. Miksch, W. Aigner, S. Ohmann, C. Popow, M. Pohl, and M. Rester. Gravi++: Interactive information visualization to explore highly structured temporal data. J.UCS: Journal of Universal Computer Science, 11(11):1792– 1805, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
11/gravi_interactive_
information_visualization.

Hafner:2009:SRA

[HMB09] M. Hafner, M. Memon, and R. Breu. SeAAS

— a reference architecture for security services in SOA. J.UCS: Journal of Universal Computer Science, 15(15):2916–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
15/seaas_a_reference_
architecture.

Haya:2006:MSC

[HME^{+}06] P. A. Haya, G. Montoro, A. Esquivel, M. García-Herranz, and X. Alamán. A mechanism for solving conflicts in ambient intelligent environments. J.UCS: Journal of Universal Computer Science, 12(3):284–296, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_3/a_mechanism_for_
solving.

Hettiarachchi:2015:IFA

[HMHGR15] E. Hettiarachchi, E. Mor, M. A. Huertas, and A.- E. Guerrero-Roldán. Introducing a formative e-assessment system to improve online learning experience and performance. J.UCS: Journal of Universal Computer Science, 21(8):1001– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
8/introducing_a_formative_
eassessment.

Hessley:2000:IAE

[HMM00] R. K. Hessley, D. L. Morris, Jr., and M. R. Mueller. Integrated applications of electronic structure computations in the undergraduate chemistry curriculum. J.UCS: Journal of Universal Computer Science, 6(3):345– 355, March 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/

jucs_6_3/integrated_
applications_of_electronic.

Hussain:2010:MSA

[HMMP10] S. Hussain, G. Min, J. Ma, and J. Hyuk Park. Multimedia services and applications. J.UCS: Journal of Universal Computer Science, 16(10):1267–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16;   http:
//www.jucs.org/jucs_
16_10#;   http://www.
jucs.org/jucs_16_10/
multimedia_services_
and_applications.

Hahnle:1999:JUS

[HMSR99] R. Hähnle, W. Menzel, P. H. Schmitt, and W. Reif. J.UCS special issue on integration of deduction systems. J.UCS: Journal of Universal Computer Science, 5(3):??, March 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
3/j_ucs_special_issue.

Hotho:2001:SIS

[HMSS01] A. Hotho, A. Maedche, S. Staab, and R. Studer. SEAL-II — the soft spot between richly structured and unstructured knowledge. J.UCS: Journal of Universal Computer Science, 7(7):566– 590, July 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
7/seal_ii_the_soft.

Harrer:2008:RCA

[HMW08] A. Harrer, N. Malzahn, and A. Wichmann. The remote control approach — an architecture for adaptive scripting across collaborative learning environments. J.UCS: Journal of Universal Computer Science, 14(1):148– 173, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
1/the_remote_control_
approach.

Hernes:2007:DCH

[HN07] M. Hernes and N. T. Nguyen. Deriving consensus for hierarchical incomplete ordered partitions and coverings. J.UCS: Journal of Universal Computer Science, 13(2):317–328, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_2/deriving_consensus_
for_hierarchical.

Hwang:2010:SWF

[HNJ^{+}10] D. Hwang, N. T. Nguyen, J. J. Jung, A. Sadeghi-Niaraki, K.-H. Baek, and Y.-S. Han. A Semantic Wiki framework for reconciling conflict collaborations based on selecting consensus choice. J.UCS: Journal of Universal Computer Science, 16(7):1024–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_7/a_semantic_wiki_
framework.

Hemaspaandra:1998:NLN

[HNP98] L. A. Hemaspaandra, C. Nasipak, and K. Parkins. A note on linear-nondeterminism, linear-sized, Karp–Lipton advice for the P -selective sets. J.UCS: Journal of Universal Computer Science, 4(8):670–674, August 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_8/a_
note_on_linear.

Hojjat:2007:IMC

[HNS07] H. Hojjat, H. Nakhost, and M. Sirjani. Integrating module checking and deduction in a formal proof for the Perlman Spanning Tree Protocol (STP). J.UCS: Journal of Universal Computer Science, 13(13): 2076, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
13/integrating_module_
checking_and.

Hori:2004:OPK

[HNYO04] K. Hori, K. Nakakoji, Y. Yamamoto, and J. Ostwald. Organic perspectives of knowledge management: Knowledge evolution through a cycle of knowledge liquidization and crystallization. J.UCS: Journal of Universal Computer Science, 10(3):252–261, March 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
3/organic_perspectives_
of_knowledge.

Holmberg:1996:PDE

[Hol96] B. Holmberg. On the potential of distance education in the age of information technology. J.UCS: Journal of Universal Computer Science, 2(6):484–491, June 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.

org/jucs_2_6/on_the_
potential_of.

Honkala:1995:DMU

[Hon95] Juha Honkala. A decision method for the unambiguity of sets defined by number systems. J.UCS: Journal of Universal Computer Science, 1(9):652– 657, September 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
a_decision_method_for_
the_unambiguity_of_sets.

Honkala:1996:IAS

[Hon96] Juha Honkala. On images of algebraic series. J.UCS: Journal of Universal Computer Science, 2(4):217–223, April 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
4/on_images_of_algebraic.

Honkala:1997:APS

[Hon97]  J. Honkala.  On N-

algebraic Parikh slender power series. J.UCS: Journal of Universal Computer Science, 3 (10):1114–1120, October 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_3_10/
                                        on_n_algebraic_parikh.

Honkala:1999:ADP

[Hon99] J. Honkala. On algebraicness of D0L power series. J.UCS: Journal of Universal Computer Science, 5(1):11–19, January 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        5_1/on_algebraicness_
                                        of_d0l.

Honkala:2001:TVD

[Hon01] J. Honkala. Three variants of the DT0L sequence equivalence problem. J.UCS: Journal of Universal Computer Science, 7(10):886–892, October 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_7_
                                        10/three_variants_of_
                                        the.

Honkala:2002:SHP

[Hon02] J. Honkala. On the simplification of HD0L power series. J.UCS: Journal of Universal Computer Science, 8(12): 1040–1046, December 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_8_12/on_the_
simplification_of.

Hopper:1998:AWB

[Hop98]  M. Hopper.  Assess-

ment in WWW-Based learning systems: Opportunities and challenges. J.UCS: Journal of Universal Computer Science, 4(4):330– 348, April 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_4/assessment_
in_www_based;   internal&
sk=0C220489.

Herskovic:2011:IEB

[HOPN11] V. Herskovic, S. F. Ochoa, J. A. Pino, and A. Neyem. The iceberg effect: Behind the user interface of mobile collaborative systems. J.UCS: Journal of Universal Computer Science, 17(2):183– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
2/the_iceberg_effect_
behind.

Hornos:2004:FTA

[Hor04] M. J. Hornos. FBT: a tool for applying interval logic specifications to on-the-fly model checking. J.UCS: Journal of Universal Computer Science, 10(11):1498–1518, November 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        10_11/fbt_a_tool_for.

Hanke:1996:EFD

[HOS96] S. Hanke, T. Ottmann, and S. Schuierer. The edge-flipping distance of triangulations. J.UCS: Journal of Universal Computer Science, 2(8): 570–579, August 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_2_
                                        8/the_edge_flipping_
                                        distance.

Holtkamp:2015:CBV

[HP15] P. Holtkamp and J. M. Pawlowski. A competencebased view on the global software development process. J.UCS: Journal of Universal Computer Science, 21(11):1385–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_21_
                                        11/a_competence_based_
                                        view.

Hasan:2012:IET

[HPB12] I. Hasan, J. Parapar, and Á. Barreiro. Improving the extraction of text in PDFs by simulating the human reading order. J.UCS: Journal of Universal Computer Science, 18(5):623–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_5/improving_
the_extraction_of.

Halder:2010:WTR

[HPC10] R. Halder, S. Pal, and A. Cortesi. Watermarking techniques for relational databases: Survey, classification and comparison. J.UCS: Journal of Universal Computer Science, 16(21):3164–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
21/watermarking_techniques_
for_relational.

Haintz:2014:DWB

[HPE14] C. Haintz, K. Pichler, and M. Ebner. Developing a Web-based question-driven audience response system supporting BYOD. J.UCS: Journal of Universal Computer Science, 20(1):39– ??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/

                                        jucs_20_1/developing_
                                        a_web_based.

Horvath:2006:CMS

[HR06] L. Horváth and I. Rudas. Course modeling for student profile based flexible higher education on the Internet. J.UCS: Journal of Universal Computer Science, 12(9):1254– 1266, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        12_9/course_modeling_
                                        for_student.

Hristea:2002:SGW

[Hri02]  F. Hristea.  On the

semiautomatic generation of WordNet type synsets and clusters. J.UCS: Journal of Universal Computer Science, 8(12):1047, December 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_8_12/on_the_
                                        semiautomatic_generation.

Huysegoms:2014:UEF

[HSD^{+}14] T. Huysegoms, M. Snoeck, G. Dedene, A. Goderis, and F. Stumpe. Using and extending formal concept analysis to visualise variability during requirements engineering. J.UCS: Journal of Universal Computer Science, 20(6):842–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_20_6/using_and_
extending_formal.

Holm:2012:SRR

[HSFE12] H. Holm, T. Sommestad, U. Franke, and M. Ekstedt. Success rate of remote code execution attacks — expert assessments and observations. J.UCS: Journal of Universal Computer Science, 18(6):732–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_6/success_rate_of_
remote.

Henderson-Sellers:2010:SME

[HSR10] B. Henderson-Sellers and J. Ralyté. Situational method engineering: Stateof-the-art review. J.UCS: Journal of Universal Computer Science, 16(3): 424–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_3/situational_method_
engineering_state.

Henderson-Sellers:2004:PCC

[HSSM^{+}04] B. Henderson-Sellers, M. Serour, T. McBride, C. Gonzalez-Perez, and L. Dagher.

Process construction and customization. J.UCS: Journal of Universal Computer Science, 10(4): 326–358, April 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        10_4/process_construction_
                                        and_customization.

Hicks:2001:PDL

[HT01] D. Hicks and K. Tochtermann. Personal digital libraries and knowledge management. J.UCS: Journal of Universal Computer Science, 7(7): 550–565, July 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        7_7/personal_digital_
                                        libraries_and.

Hsu:2006:PUD

[HT06] P.-Y. Hsu and P.-H. Ting. POCA: a user distributions algorithm in enterprise systems with clustering. J.UCS: Journal of Universal Computer Science, 12(2):160–186, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_12_2/poca_a_
                                        user_distributions.

Hidasi:2013:IMF

[HT13] B. Hidasi and D. Tikk.

Initializing matrix factorization methods on implicit feedback databases. J.UCS: Journal of Universal Computer Science, 19(12):1834–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_12/initializing_
matrix_factorization_
methods.

Hellmers:2012:UEM

[HTHW12] J. Hellmers, J. Thomaschewski, E.-M. Holt, and T. Wriedt. Usability evaluation methods for a scientific Internet information portal. J.UCS: Journal of Universal Computer Science, 18(10):1308–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_10/usability_
evaluation_methods_for.

Hulicki:2008:DBD

[Hul08]  Z. Hulicki.  Drives

and barriers for development of broadband access — CE perspective. J.UCS: Journal of Universal Computer Science, 14(5):717–730, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_5/drives_and_
barriers_for.

Hernandez:2012:ORD

[HVCA12] J. L. Hernández, A. Vizcaíno, I. Caballero, and G. Aranda. Obtaining requirements for designing a tool to support distributed development. J.UCS: Journal of Universal Computer Science, 18(19):2602–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_18_19/obtaining_
                                        requirements_for_designing.

Hadjiefthymiades:2000:PRW

[HVM00] S. Hadjiefthymiades, I. Varouxis, and D. Martakos. Performance of RDBMS-WWW interfaces under heavy workload. J.UCS: Journal of Universal Computer Science, 6(6):538– 559, June 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_6_6/performance_
                                        of_rdbms_www.

Hertling:1997:IPR

[HW97] P. Hertling and Y. Wang. Invariance properties of random sequences. J.UCS: Journal of Universal Computer Science, 3 (11):1241–1249, November 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_3_11/
invariance_properties_
of_random.

Hellmers:2010:CSS

[HW10] J. Hellmers and T. Wriedt.

Classification of software for the simulation of light scattering and realization within an Internet information portal. J.UCS: Journal of Universal Computer Science, 16(9):1176–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_9/classification_
of_software_for.

Hicks:2002:RSC

[HWN02] D. L. Hicks, U. K. Wiil, and P. J. Nürnberg. Research in structural computing. J.UCS: Journal of Universal Computer Science, 8(10):871–880, October 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
10/research_in_structural_
computing.

He:2005:TUB

[HYC^{+}05] F. He, G. Yang, L. Cheng, X. Song, M. Gu, and J. Sun. On theoretical upper bound for routing estimation. J.UCS: Journal of Universal Computer Science, 11(6):916– 925, ???? 2005. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_6/on_theoretical_
upper_bound.

Huang:2012:MTM

[HZZ^{+}12] H. Huang, D. Zhang, Y. Zhu, M. Li, and M.-Y. Wu. A metropolitan taxi mobility model from real GPS traces. J.UCS: Journal of Universal Computer Science, 18(9):1072– ??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_9/a_metropolitan_
taxi_mobility.

Imran:2016:WDA

[IAS16] A. Imran, S. Aljawarneh, and K. Sakib. Web data amalgamation for security engineering: Digital forensic investigation of open source cloud. J.UCS: Journal of Universal Computer Science, 22(4):494– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_22_
4/web_data_amalgamation_
for.

Iskander:2011:MMV

[IBN^{+}11] M. F. Iskander, J. Baker, J. Kobashigawa Nakatsu, S. Y. Lim, and N. Celik. Multimedia modules and virtual organization Website for collaborative research experience for teachers in STEM. J.UCS: Journal of Universal Computer Science, 17(9):1347– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
9/multimedia_modules_
and_virtual.

Ierusalimschy:2005:IL

[IdFC05] R. Ierusalimschy, L. H.

de Figueiredo, and W. Celes. The implementation of Lua 5.0. J.UCS: Journal of Universal Computer Science, 11(7):1159– 1176, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
7/the_implementation_
of_lua.

Ito:2002:SRC

[IDS02] M. Ito, J. Dassow, and R. Stiebe. Some remarks on codes defined by Petri nets. J.UCS: Journal of Universal Computer Science, 8(2):260– 269, February 28, 2002.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        8_2/some_remarks_on_
                                        codes.

Ierusalimschy:2003:BSP

[IFd03] Roberto Ierusalimschy, Lucilia Figueiredo, and Marco T. de Oliveira Valente. 7th Brasilian Symposium on Programming Languages — J.UCS special issue. J.UCS: Journal of Universal Computer Science, 9(8):729, August 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/free;
                                         http://www.jucs.org/
                                        jucs_9_8/7th_brasilian_
                                        symposium_on.

Ibanez:2008:DPA

[IGS08] J. Ibáñez, D. García, and O. Serrano. Displaying pictures according to the songs being played. J.UCS: Journal of Universal Computer Science, 14(19):3114–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        19/displaying_pictures_
                                        according_to.

Ishihara:1997:ECT

[IK97] H. Ishihara and B. Khoussainov. Effectiveness of the completeness theorem for an intermediate logic. J.UCS: Journal of Universal Computer Science, 3(11):1255–1265, November 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_3_11/
effectiveness_of_the_
completeness.

Ibanez:2014:IEW

[IKC14] M. B. Ibáñez, C. Delgado Kloos, and V. Callaghan. Immersive education: What does the future hold? J.UCS: Journal of Universal Computer Science, 20(12):1605–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_20;   http:
//www.jucs.org/jucs_
20_12/immersive_education_
what_does.

Ilmola:2003:FSF

[IKM03] L. Ilmola and A. Kotsalo-Mustonen. Filters in the strategy formulation process. J.UCS: Journal of Universal Computer Science, 9(6):481–490, June 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_9_6/filters_
                                        in_the_strategy.

Iljazovi:2009:CCC

[Ilj09] Z. Iljazovi. Chainable and circularly chainable co-r.e. sets in computable metric spaces. J.UCS: Journal of Universal Computer Science, 15(6):1206–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_15_6/chainable_
                                        and_circularly_chainable.

Iljazovic:2010:ICS

[Ilj10] Z. Iljazović. Isometries and computability structures. J.UCS: Journal of Universal Computer Science, 16(18):2569–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        16_18/isometries_and_
                                        computability_structures.

Ibanez:2012:ACL

[IMR^{+}12] M. B. Ibáñez, D. Maroto, J. J. García Rueda, D. Leony, and C. Delgado Kloos. Architecture for collaborative learning activities in hybrid learning environments. J.UCS: Journal of Universal Computer Science, 18(15):2187–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
15/architecture_for_
collaborative_learning.

Ikeda:2009:AFD

[INK09] S. Ikeda, T. Nagamine, and T. Kamada. Application framework with demand-driven mashup for selective browsing. J.UCS: Journal of Universal Computer Science, 15(10):2109–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_10/application_
framework_with_demand.

In:2004:RNU

[IO04] H. P. In and D. Olson. Requirements negotiation using multicriteria preference analysis. J.UCS: Journal of Universal Computer Science, 10(4):306–325, April 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
4/requirements_negotiation_
using_multi.

Iorgulescu:2000:CBM

[Ior00]  A. Iorgulescu.  Con-

nections between MV_{n} algebras and n-valued Lukasiewicz-Moisil algebras — IV. J.UCS: Journal of Universal Computer Science, 6(1):139– 154, January 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
1/connection_between_
mvn_algebras.

Iorgulescu:2007:BAP

[Ior07] A. Iorgulescu. On BCK algebras — Part I.a: An attempt to treat unitarily the algebras of logic. new algebras. J.UCS: Journal of Universal Computer Science, 13(11):1628– 1654, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
11/on_bck_algebras_part.

Iorgulescu:2008:BAP

[Ior08] A. Iorgulescu. On BCK algebras — Part I.b: An attempt to treat unitarily the algebras of logic. New algebras. J.UCS: Journal of Universal Computer Science, 14(22):3686–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
22/on_bck_algebras.

Iribarne:2012:MSI

[IPCVC12] L. Iribarne, N. Padilla, J. Criado, and C. Vicente-Chicote. Metamodeling the structure and interaction behavior of cooperative component-based user interfaces. J.UCS: Journal of Universal Computer Science, 18 (19):2669–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_19/metamodeling_the_
structure_and.

Ibanez:2012:CTQ

[IRMK12] M. B. Ibáñez, J. J. Garcia Rueda, D. Morillo, and C. Delgado Kloos. Creating test questions for 3D collaborative virtual worlds: the World-OfQuestions authoring environment. J.UCS: Journal of Universal Computer Science, 18 (18):2556–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_18/creating_test_
questions_for.

Ionescu:2004:SPI

[IS04] M. Ionescu and D. Sburlan.

On P systems with Promoters/Inhibitors. J.UCS: Journal of Universal Computer Science, 10(5):

581–599, May 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        10_5/on_p_systems_with.

Iglesias:2010:ADT

[IS10] M. J. Fernández Iglesias and L. M. Álvarez Sabucedo. From analog to digital television; strategies to promote rapid adaptation and awareness. J.UCS: Journal of Universal Computer Science, 16(8):1056–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        16_8/from_analog_to_
                                        digital.

Ishihara:1997:SCL

[Ish97]  H. Ishihara.  Sequen-

tial continuity of linear mappings in constructive mathematics. J.UCS: Journal of Universal Computer Science, 3(11):1250–1254, November 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_3_11/
                                        sequential_continuity_
                                        of_linear.

Ishihara:2000:CMC

[Ish00] H. Ishihara. A canonical model construction for substructural logics. J.UCS: Journal of Universal Computer Science, 6(1):155–168, January 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_6_1/a_canonical_
model_construction.

Istrate:2007:SAR

[Ist07] G. Istrate. Satisfying assignments of random Boolean constraint satisfaction problems: Clusters and overlaps. J.UCS: Journal of Universal Computer Science, 13 (11):1655–1670, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_11/satisfying_assignments_
of_random.

Ito:2002:SDR

[Ito02] M. Ito. Shuffle decomposition of regular languages. J.UCS: Journal of Universal Computer Science, 8(2):257–259, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_8_2/shuffle_
decomposition_of_regular.

Jansen:2013:UCS

[JBBH13] M. Jansen, L. Bollen, N. Baloian, and H. U. Hoppe. Using cloud services to develop learning scenarios from a software engineering perspective. J.UCS: Journal of Universal Computer Science, 19(14):2037–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_19_14/using_
cloud_services_to.

Janssen:2010:ALP

[JBH^{+}10] J. Janssen, A. J. Berlanga, S. Heyenrath, H. Martens, H. Vogten, A. Finders, E. Herder, H. Hermans, J. Melero Gallardo, L. Schaeps, and R. Koper. Assessing the learning path specification: a pragmatic quality approach. J.UCS: Journal of Universal Computer Science, 16(21): 3191–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_21/assessing_
the_learning_path.

Jung:2015:IDP

[JCB15] J. J. Jung, D. Camacho, and C. Badica. Intelligent distributed processing methods for Big Data.

J.UCS: Journal of Universal Computer Science, 21(6):754–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_21;   http:
//www.jucs.org/jucs_
21_6/intelligent_distributed_
processing_methods.

Jezequel:1995:REP

[Jéz95] Fabienne Jézéquel. Roundoff error propagation in the solution of the heat equation by finite differences. J.UCS: Journal of Universal Computer Science, 1(7):469– 483, July 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
1_7/round_off_error_
propagation.

Jara:2013:EBL

[JFL^{+}13] A. J. Jara, D. Fernandez, P. Lopez, M. A. Zamora, A. F. Skarmeta, and L. Marin. Evaluation of Bluetooth low energy capabilities for tele-mobile monitoring in home-care. J.UCS: Journal of Universal Computer Science, 19(9):1219–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_19_9/evaluation_
                                        of_bluetooth_low.

Jiang:2009:ODP

[JFZ09] T. Jiang, Y. Feng, and B. Zhang. Online detecting and predicting special patterns over financial data streams. J.UCS: Journal of Universal Computer Science, 15 (13):2566–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        15_13/online_detecting_
                                        and_predicting.

Jimenez:2008:GBI

[JGL08] J. Jimenez, D. Gutierrez, and P. Latorre. Gaze-based interaction for virtual environments. J.UCS: Journal of Universal Computer Science, 14(19):3085–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_14_19/gaze_
                                        based_interaction_for_
                                        virtual.

Junior:2010:SMV

[JGM10] E. A. Oliveira Junior, I. M. S. Gimenes, and J. C. Maldonado. Systematic management of variability in UML-based software product lines.

J.UCS: Journal of Universal Computer Science, 16(17):2374–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_17/systematic_
management_of_variability.

Junior:2013:SES

[JGM^{+}13] E. A. Oliveira Junior, I. M. S. Gimenes, J. C. Maldonado, P. C. Masiero, and L. Barroca. Systematic evaluation of software product line architectures. J.UCS: Journal of Universal Computer Science, 19(1):25–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_1/systematic_
evaluation_of_software.

Johnson:2011:PTR

[JGW11] M. Johnson, D. Griffiths, and M. Wang. Positioning theory, roles and the design and implementation of learning technology. J.UCS: Journal of Universal Computer Science, 17(9):1329– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
9/positioning_theory_
roles_and.

Joha:2012:DCU

[JJ12] A. Joha and M. Janssen. Design choices underlying the software as a service (SaaS) business model from the user perspective: Exploring the fourth wave of outsourcing. J.UCS: Journal of Universal Computer Science, 18(11):1501–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        18_11/design_choices_
                                        underlying_the.

Janiak:2008:TSG

[JJL08] A. Janiak, W. Janiak, and M. Lichtenstein. Tabu search on GPU. J.UCS: Journal of Universal Computer Science, 14 (14):2416–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_14/tabu_search_on_
                                        gpu.

Jung:2010:RTS

[JK10a] J. J. Jung and C. Koo. Recent trends in service science. J.UCS: Journal of Universal Computer Science, 16(13):1666–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16;   http:
//www.jucs.org/jucs_
16_13/recent_trends_
in_service_science;
http://www.jucs.org/
jucsrssfeed-issue.

Jung:2010:SAC

[JK10b] Y. Jung and M. Kim.

Situation-aware community computing model for developing dynamic ubiquitous computing systems. J.UCS: Journal of Universal Computer Science, 16(15):2139–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_15/situation_
aware_community_computing.

Jung:2012:ASN

[JK12a] J. J. Jung and P. Kazienko.

Advances on social network applications. J.UCS: Journal of Universal Computer Science, 18 (4):454–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_18;   http:
//www.jucs.org/jucs_
18_4/advances_in_social_
network;   http://www.
jucs.org/jucsrssfeed-
issue.

Jung:2012:UOS

[JK12b] J. J. Jung and P. Kazienko. Understanding online social networking services. J.UCS: Journal of Universal Computer Science, 18(8):970–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_18;   http:
                                        //www.jucs.org/jucs_
                                        18_8/understanding_online_
                                        social_networking;
                                        http://www.jucs.org/
                                        jucsrssfeed-issue.

Jankowski:2016:FMU

[JKKW16] J. Jankowski, K. Kolomvatsos, P. Kazienko, and J. Watróbski. Fuzzy modeling of user behaviors and virtual goods purchases in social networking platforms. J.UCS: Journal of Universal Computer Science, 22(3):416– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_22_
                                        3/fuzzy_modeling_of_
                                        user.

Jeong:2008:PKE

[JL08] I. R. Jeong and D. H. Lee. Parallel key exchange. J.UCS: Journal of Universal Computer Science, 14(3):377– 396, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
3/parallel_key_exchange.

Juang:2009:ACM

[JL09] J.-G. Juang and C.- L. Lee. Applications of cerebellar model articulation controllers to intelligent landing system. J.UCS: Journal of Universal Computer Science, 15(13):2586–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
13/applications_of_cerebellar_
model.

Jha:2016:APB

[JL16] A. K. Jha and W. J. Lee.

Analysis of permissionbased security in Android through policy expert, developer, and end user perspectives. J.UCS: Journal of Universal Computer Science, 22(4):459– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_22_
4/analysis_of_permission_
based.

Jones:2005:AM

[JLRW05] C. Jones, D. Lomet, A. Romanovsky, and G. Weikum. The Atomic Manifesto. J.UCS: Journal of Universal Computer Science, 11(5):636– 650, May 28, 2005. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_11_5/the_atomic_
manifesto.

Jurgensen:2010:EHM

[JM10] H. Jürgensen and D. E. Matthews. Entropy and higher moments of information. J.UCS: Journal of Universal Computer Science, 16(5):749– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
5/entropy_and_higher_
moments.

J:2015:CBI

[JM15] B. Antony J and G. S. Mahalakshmi. Content-based information retrieval by named entity recognition and verb semantic role labelling. J.UCS: Journal of Universal Computer Science, 21(13): 1830–??, ???? 2015. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
21_13/content_based_
information_retrieval.

Juric:2010:LML

[JMEL10] M. Jurič, I. Mozetič, T. Erjavec, and N. Lavrač. LemmaGen: Multilingual lemmatisation with induced ripple-down rules. J.UCS: Journal of Universal Computer Science, 16(9):1190–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_9/lemma_gen_
multilingual_lemmatisation.

Jovic:2012:PPS

[JMKT12] M. Jovic, D. Milutinovic, A. Kos, and S. Tomazic. Product presentation strategy for online customers. J.UCS: Journal of Universal Computer Science, 18 (10):1323–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_10/product_presentation_
strategy_for.

Julian:2006:OIU

[JMP06] P. Julián, G. Moreno, and J. Penabad. Operational/ interpretive unfolding of multi-adjoint logic programs. J.UCS: Journal of Universal Computer Science, 12(11):1679–1699, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12_
                                        11/operational_interpretive_
                                        unfolding_of.

Jurgensen:2010:SYD

[JMSY10] H. Jürgensen, H. Maurer, A. Salomaa, and S. Yu. Seventy years Derick Wood. J.UCS: Journal of Universal Computer Science, 16(5):577– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_16;
                                         http://www.jucs.org/
                                        jucs_16_5#;   http://
                                        www.jucs.org/jucs_16_
                                        5/seventy_years_derick_
                                        wood.

Jung:2008:CIS

[JN08] J. J. Jung and N. Thanh Nguyen. Collective intelligence for semantic and knowledge grid. J.UCS: Journal of Universal Computer Science, 14(7):1016–1019, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14;
                                         http://www.jucs.org/
                                        jucs_14_7#;   http://
                                        www.jucs.org/jucs_14_
                                        7/collective_intelligence_
                                        for_semantic.

Junior:2012:ABA

[JNdMM12] L. Silva Junior, N. Nedjah, and L. de Macedo Mourelle. ACO-based algorithms for search and optimization of routes in NoC platform. J.UCS: Journal of Universal Computer Science, 18(7):917–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_7/aco_based_
algorithms_for.

Jo:2009:KMI

[JNS09] K.-H. Jo, N. T. Nguyen, and E. Szczerbicki. Knowledge management for intelligent systems. J.UCS: Journal of Universal Computer Science, 15 (13):2445–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15;   http://www.jucs.
org/jucs_15_13#;   http:
//www.jucs.org/jucs_
15_13/knowledge_management_
for_intelligent.

Johnston:2001:SJK

[Joh01] P. Johnston. Sustainability and jobs in the knowledge economy. J.UCS: Journal of Universal Computer Science, 7(6): 498–506, June 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_7_
                                        6/sustainability_and_
                                        jobs_in.

Jones:2001:TVV

[Jon01]  C. B. Jones.  The

transition from VDL to VDM. J.UCS: Journal of Universal Computer Science, 7(8):631–640, August 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_7_
                                        8/the_transition_from_
                                        VDL.

Jaddi:2007:AHE

[JP07] F. Jaddi and B. Paillassa. An adaptive hierarchical extension of DSR: The cluster source routing. J.UCS: Journal of Universal Computer Science, 13(1):32– 55, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_
                                        1/an_adaptive_hierarchical_
                                        extension.

Jurgensen:1996:TFC

[JR96] H. Jürgensen and L. Robbins. Towards foundations of cryptography: Investigation of perfect secrecy. J.UCS: Journal of Universal Computer Science, 2(5):347–379, May 28, 1996. CODEN ????

ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_2_5/towards_
foundations_of_cryptography.

Jeribi:2002:ICM

[JR02] L. Jéribi and B. Rumpler. Instance cooperative memory to improve query expansion in information retrieval systems. J.UCS: Journal of Universal Computer Science, 8(6):591– 601, June 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
6/instance_cooperative_
memory_to.

Jose:2010:AIB

[JRO10] R. José, H. Rodrigues, and N. Otero. Ambient intelligence: Beyond the inspiring vision. J.UCS: Journal of Universal Computer Science, 16(12):1480–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
12/ambient_intelligence_
beyond_the.

Jedrzejowicz:2016:PBS

[JRR16] P. Jedrzejowicz and E. Ratajczak-Ropel. PLA based strategy for solving RCPSP by a team of agents. J.UCS: Journal of Universal Computer Science, 22(6):856–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_22_6/pla_based_
strategy_for.

Johnson:2010:AUS

[JRW10] M. Johnson, R. Rosebrugh, and R. Wood. Algebras and update strategies. J.UCS: Journal of Universal Computer Science, 16(5):729– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
5/algebras_and_update_
strategies.

Jedrzejowicz:2016:IPD

[JS16] P. Jedrzejowicz and A. Skakovski. Improving performance of the differential evolution algorithm using cyclic decloning and changeable population size. J.UCS: Journal of Universal Computer Science, 22(6):874–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_22_6/improving_
performance_of_the.

Jakus:2011:ADS

[JST11] G. Jakus, J. Sodnik, and S. Tomažič. The architectural design of a system for interpreting multilingual Web documents in Esperanto. J.UCS: Journal of Universal Computer Science, 17(3):377– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
3/the_architectural_
design_of.

Joglekar:2005:PEM

[JT05] S. P. Joglekar and S. R.

Tate. ProtoMon: Embedded monitors for cryptographic protocol intrusion detection and prevention. J.UCS: Journal of Universal Computer Science, 11(1):83–103, January 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
1/protomon_embedded_
monitors_for.

Junttila:2001:CCP

[Jun01] T. A. Junttila. Computational complexity of the place/transitionnet symmetry reduction method. J.UCS: Journal of Universal Computer Science, 7(4):307– 326, April 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_7_
                                        4/computational_complexity_
                                        of_the.

Jung:2005:CWB

[Jun05a]  J. J. Jung.  Collab-

orative Web browsing based on semantic extraction of user interests with bookmarks. J.UCS: Journal of Universal Computer Science, 11(2):213–228, February 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_11_
                                        2/collaborative_web_
                                        browsing_based.

Jung:2005:SPW

[Jun05b] J. J. Jung. Semantic preprocessing of Web request streams for Web usage mining. J.UCS: Journal of Universal Computer Science, 11(8):1383– 1396, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_11_
                                        8/semantic_preprocessing_
                                        of_web.

Jung:2005:VRF

[Jun05c]  J. J. Jung.  Visu-

alizing recommendation flow on social network.

J.UCS: Journal of Universal Computer Science, 11(11):1780–1791, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
11/visualizing_recommendation_
flow_on.

Jung:2008:QTB

[Jun08] J. J. Jung. Query transformation based on semantic centrality in semantic social network1. J.UCS: Journal of Universal Computer Science, 14(7):1031–1047, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_14_7/query_
transformation_based_
on.

Jung:2010:CIV

[Jun10a]  J. J. Jung.  Collec-

tive intelligence with visualization and multimedia. J.UCS: Journal of Universal Computer Science, 16(7):1004– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16;
 http://www.jucs.org/
jucs_16_7#;   http://
www.jucs.org/jucs_16_
7/collective_intelligence_
with_visualization.

Jung:2010:ISN

[Jun10b] J. J. Jung. Integrating social networks for context fusion in mobile service platforms. J.UCS: Journal of Universal Computer Science, 16(15): 2099–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        16_15/integrating_social_
                                        networks_for.

Jung:2010:SCA

[Jun10c] J. J. Jung. On sustainability of context-aware services among heterogeneous smart spaces. J.UCS: Journal of Universal Computer Science, 16(13):1745–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_16_13/on_sustainability_
                                        of_context.

Jurgensen:2010:DWP

[Jür10]  H. Jürgensen.  Der-

ick Wood’s publications. J.UCS: Journal of Universal Computer Science, 16(5):862–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_16_5/derick_
                                        woods_publications.

Janev:2005:RKM

[JV05] V. Janev and S. Vranes.

The role of knowledge management solutions in enterprise business processes. J.UCS: Journal of Universal Computer Science, 11(4):526–545, April 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
4/the_role_of_knowledge.

Janev:2011:OBC

[JV11] V. Janev and S. Vranes.

Ontology-based competency management: the case study of the Mihajlo Pupin Institute. J.UCS: Journal of Universal Computer Science, 17(7):1089–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_7/ontology_based_
competency_management.

Jafarikhah:2013:RRO

[JW13] T. Jafarikhah and K. Weihrauch.

The Riesz representation operator on the dual of C[0; 1] is computable. J.UCS: Journal of Universal Computer Science, 19(6):750–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_19_6/the_riesz_
                                        representation_operator.

Kuhn:1997:CMK

[KA97] O. Kühn and A. Abecker. Corporate memories for knowledge management in industrial practice: Prospects and challenges. J.UCS: Journal of Universal Computer Science, 3(8):929–954, August 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs_3_8/corporate_
                                        memories_for_knowledge;
                                        internal&sk=05460486.

Kavi:2000:ECP

[KAG00] K. Kavi, J. Arul, and R. Giorgi. Execution and cache performance of the scheduled dataflow architecture. J.UCS: Journal of Universal Computer Science, 6(10):948– 967, October 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_6_10/execution_
                                        and_cache_performance.

Kahler:2001:MTW

[Kah01] H. Kahler. More than WORDs — collaborative tailoring of a word processor. J.UCS: Journal of Universal Computer Science, 7(9):826– 847, September 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
7_9/more_than_words_
collaborative.

Kong:2003:DFC

[KAM03] C. Kong, P. Alexander, and C. Menon. Defining a formal coalgebraic semantics for the Rosetta Specification Language. J.UCS: Journal of Universal Computer Science, 9(11):1322–1349, November 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
11/defining_a_formal_
coalgebraic.

Kappe:1995:SAM

[Kap95] F. Kappe. A scalable architecture for maintaining referential integrity in distributed information systems. J.UCS: Journal of Universal Computer Science, 1(2):84– 104, February 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
a_scalable_architecture_
for_maintaining.

Kari:2002:SSF

[Kar02] J. Kari. Synchronization and stability of finite automata. J.UCS: Journal of Universal Computer Science, 8(2):270– 277, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        8_2/synchronization_
                                        and_stability_of.

Karahoca:2013:MCT

[Kar13]  D. Karahoca.  Meta-

cognitive tool development for history teaching: Investigating how software usability affects student achievements. J.UCS: Journal of Universal Computer Science, 19(5):619–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        19_5/meta_cognitive_
                                        tool_evelopment.

Karimi:2008:ONC

[KASN08] N. Karimi, A. Alaghi, M. Sedghi, and Z. Navabi. Online networkon-chip switch fault detection and diagnosis using functional switch faults. J.UCS: Journal of Universal Computer Science, 14(22):3716–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
22/online_network_on_
chip.

Karsai:2003:UGT

[KASS03] G. Karsai, A. Agrawal, F. Shi, and J. Sprinkle. On the use of graph transformation in the formal specification of model interpreters. J.UCS: Journal of Universal Computer Science, 9(11):1296– 1321, November 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
11/on_the_use_of.

Katarzyniak:2005:LGP

[Kat05] R. P. Katarzyniak. The language grounding problem and its relation to the internal structure of cognitive agents. J.UCS: Journal of Universal Computer Science, 11(2):357–374, February 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
2/the_language_grounding_
problem.

Kulathuramaiyer:2006:RVC

[KB06] N. Kulathuramaiyer and W.-T. Balke. Restricting the view and connecting the dots — dangers of a Web search engine monopoly. J.UCS: Journal of Universal Computer Science, 12(12): 1731–1740, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_12/restricting_the_
view_and.

Kreuzthaler:2011:CDR

[KBF^{+}11] M. Kreuzthaler, M. D. Bloice, L. Faulstich, K.- M. Simonic, and A. Holzinger. A comparison of different retrieval strategies working on medical free texts. J.UCS: Journal of Universal Computer Science, 17(7):1109–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_7/a_comparison_of_
different.

Klimek:2014:TSC

[KBN14] J. Klímek, S. Benda, and M. Necaský. Translation of structural constraints from conceptual model for XML to schematron. J.UCS: Journal of Universal Computer Science, 20(3):277–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_

20_3/translation_of_
structural_constraints.

Kang:2008:RIW

[KC08] J. Kang and J. Choi.

Recognising informative Web page blocks using visual segmentation for efficient information extraction. J.UCS: Journal of Universal Computer Science, 14(11):1893–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
11/recognising_informative_
web_page.

Kahraman:2010:SAR

[KCK10] Cengiz Kahraman, Selcuk Cebi, and İhsan Kaya. Selection among renewable energy alternatives using fuzzy axiomatic design: The case of Turkey. J.UCS: Journal of Universal Computer Science, 16(1):82–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_1/selection_among_
renewable_energy.

Kim:2010:UBM

[KCKL10] J. Kim, J.-Y. Choi, I. Kang, and I. Lee. UML behavior models of realtime embedded software for model-driven architecture. J.UCS: Journal of Universal Computer Science, 16(17):2415–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
17/uml_behavior_models_
of.

Kahl:2001:DTG

[KD01] W. Kahl and F. Derichsweiler. Declarative term graph attribution for program generation. J.UCS: Journal of Universal Computer Science, 7(1):54–70, January 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
7_1/declarative_term_
graph_attribution.

Karner:2002:UCI

[KD02] H. Karner and G. Droschl. Usage-centered interface design for knowledge management software. J.UCS: Journal of Universal Computer Science, 8(6): 634–643, June 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
6/usage_centered_interface_
design.

Kutalek:2005:CCC

[KD05] V. Kutálek and V. Dvorák.

On complexity of collective communications on a fat cube topology. J.UCS: Journal of Universal Computer Science, 11(6):944–961, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
6/on_complexity_of_collective.

Kavallieratou:2011:TLD

[KD11] E. Kavallieratou and F. Daskas. Text line detection and segmentation: Uneven skew angles and hill-and-dale writing. J.UCS: Journal of Universal Computer Science, 17(1):16–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_1/text_line_
detection_and.

Koh:2009:DSP

[KDGH09] B.-S. Koh, M. Denko, S. Gritzalis, and C.-H. Hsu. Data security and privacy protection in pervasive computing environments. J.UCS: Journal of Universal Computer Science, 15(5):967– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15;

                                         http://www.jucs.org/
                                        jucs_15_5#;   http://
                                        www.jucs.org/jucs_15_
                                        5/data_security_and_
                                        privacy.

Khelif:2007:OBA

[KDKB07] K. Khelif, R. Dieng-Kuntz, and P. Barbry. An ontology-based approach to support text mining and information retrieval in the biological domain. J.UCS: Journal of Universal Computer Science, 13(12):1881–1907, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_
                                        12/an_ontology_based_
                                        approach.

Kozlak:2008:ACM

[KDKDN08] J. Koźlak, G. Dobrowolski, M. Kisiel-Dorohinicki, and E. Nawarecki. Anticrisis management of city traffic using agent-based approach. J.UCS: Journal of Universal Computer Science, 14(14):2359–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_14_14/anti_
                                        crisis_management_of.

Kell:2008:SPS

[Kel08] S. Kell. A survey of practical software adaptation techniques. J.UCS: Journal of Universal Computer Science, 14(13): 2110–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_13/a_survey_
of_practical.

Konecny:2010:CSD

[KF10a] M. Konečný and A. Farjudian. Compositional semantics of dataflow networks with query-driven communication of exact values. J.UCS: Journal of Universal Computer Science, 16(18):2629–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_18/compositional_
semantics_of_dataflow.

Konecny:2010:SQD

[KF10b] M. Konečný and A. Farjudian. Semantics of query-driven communication of exact values. J.UCS: Journal of Universal Computer Science, 16(18):2597–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_18/semantics_
of_query_driven.

Kohler:2005:ICP

[KFK05] A. Köhler and F. Fuchs-Kittowski. Integration of communities into process-oriented structures. J.UCS: Journal of Universal Computer Science, 11(3):410–425, March 28, 2005. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        11_3/integration_of_
                                        communities_into.

Kloos:2012:TIE

[KGK12] C. Delgado Kloos, C. Gütl, and F. Kappe. Trends in immersive education research. J.UCS: Journal of Universal Computer Science, 18(18):2514–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_18;
                                         http://www.jucs.org/
                                        jucs_18_18/trends_in_
                                        immersive_education.

Kim:2012:NMB

[KH12] D. Kim and D. Hwang. Non-marker based mobile augmented reality and its applications using object recognition. J.UCS: Journal of Universal Computer Science, 18(20):2832–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
20/non_marker_based_
mobile.

Kyriakou:2010:TMD

[KHG10] P. Kyriakou, I. Hatzilygeroudis, and J. Garofalakis. A tool for managing domain knowledge and helping tutors in intelligent tutoring systems. J.UCS: Journal of Universal Computer Science, 16(19):2841–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
19/a_tool_for_managing.

Kloos:2012:TLA

[KHLAP12] C. Delgado Kloos, D. Hernández-Leo, and J. I. Asensio-Pérez. Technology for learning across physical and virtual spaces. J.UCS: Journal of Universal Computer Science, 18(15):2093–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_18;   http:
//www.jucs.org/jucs_
18_15/technology_for_
learning_across.

Kochs:1999:EAR

[KHN99] H. D. Kochs, H. Hilmer, and T. Nisbach. Efficient approximate reliability evaluation using the Markovian minimal cut approach. J.UCS: Journal of Universal Computer Science, 5(10):644– 667, October 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_5_10/efficient_
approximate_reliability_
evaluation.

Kienzle:2005:ASD

[Kie05a] J. Kienzle. On atomicity and software development. J.UCS: Journal of Universal Computer Science, 11(5):687–702, May 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
5/on_atomicity_and_software.

Kiewra:2005:RRS

[Kie05b] M. Kiewra. RankFeed — recommendation as searching without queries: New hybrid method of recommendation. J.UCS: Journal of Universal Computer Science, 11 (2):229–249, February 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
2/rankfeed_recommendation_
as_searching.

Kilkki:2008:QEC

[Kil08]  K. Kilkki.  Quality

of experience in communications ecosystem. J.UCS: Journal of Universal Computer Science, 14(5):615–624, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_5/quality_of_
experience_in.

Kim:2010:RTA

[Kim10]  Y. B. Kim.  Real-

time analysis of timebased usability and accessibility for human mobile-Web interactions in the ubiquitous Internet. J.UCS: Journal of Universal Computer Science, 16(15):1953–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_15/real_time_
analysis_of.

Kim:2012:ULS

[Kim12] Y. B. Kim. Utilizationlevel and serviceability of a social name-card portal for QoS in a Cloud social networking service. J.UCS: Journal of Universal Computer Science, 18(11):1523–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
11/utilization_level_
and_serviceability.

Kirchberg:2009:UAS

[Kir09]  M. Kirchberg.  Using

abstract state machines to model ARIES-based transaction processing. J.UCS: Journal of Universal Computer Science, 15(1):157–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_1/using_abstract_
state_machines.

Kolaczek:2010:DLB

[KJ10] G. Kolaczek and K. Juszczyszyn. Deontic logic-based framework for ontology aligment in agent communities. J.UCS: Journal of Universal Computer Science, 16(1):178– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
1/deontic_logic_based_
framework.

Kurilovas:2014:SSW

[KJKS14] E. Kurilovas, A. Juskeviciene, S. Kubilinskiene, and S. Serikoviene. Several semantic Web approaches to improving the adaptation quality of virtual learning environments. J.UCS: Journal of Universal Computer Science, 20(10):1418–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
10/several_semantic_
web_approaches.

Kwon:2009:LWK

[KJL09] J. O. Kwon, I. R. Jeong, and D. H. Lee. Lightweight key exchange with different passwords in the standard model. J.UCS: Journal of Universal Computer Science, 15(5):1042–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_15_5/light_
weight_key_exchange.

Kim:2008:AMS

[KJZJ08] C. Kim, J.-Y. Jung, H.- C. Zin, and J. J. Jung. An application of meta search agent system based on semantized tags for enhanced Web searching. J.UCS: Journal of Universal Computer Science, 14(14):2400–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_

                                        14/an_application_of_
                                        meta.

Kawachi:2006:PQC

[KK06] A. Kawachi and T. Koshiba. Progress in quantum computational cryptography. J.UCS: Journal of Universal Computer Science, 12(6):691– 709, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12_
                                        6/progress_in_quantum_
                                        computational.

Kreinovich:2010:CSO

[KK10] V. Kreinovich and B. J. Kubica. From computing sets of optima, Pareto sets, and sets of Nash equilibria to general decision-related set computations. J.UCS: Journal of Universal Computer Science, 16(18): 2657–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        16_18/from_computing_
                                        sets_of.

Kajdanowicz:2013:BBM

[KK13a] T. Kajdanowicz and P. Kazienko. Boostingbased multi-label classification. J.UCS: Journal of Universal Computer Science, 19(4):502– ??, ???? 2013. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_4/boosting_based_
multi_label.

Kekolahti:2013:AMS

[KK13b] P. Kekolahti and J. Karikoski.

Analysis of mobile service usage behaviour with Bayesian belief networks. J.UCS: Journal of Universal Computer Science, 19(3):325– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
3/analysis_of_mobile_
service.

Kim:2012:WBS

[KKB12] J. W. Kim, A. Kashyap, and S. Bhamidipati. Wikipediabased semantic interpreter using approximate top-k processing and its application. J.UCS: Journal of Universal Computer Science, 18(5):650– ??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_5/wikipedia_based_
semantic_interpreter.

Keller:2012:OCP

[KKH12] J. Keller, C. W. Kessler, and R. Hultén. Optimized on-chip-pipelining for memory-intensive computations on multi-core processors with explicit memory hierarchy. J.UCS: Journal of Universal Computer Science, 18 (14):1987–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_14/optimized_
on_chip_pipelining.

Kim:2014:FMV

[KKK^{+}14] Y. Kim, I. Kim, I. Kang, T. Kim, and M. Sung. Formal modeling and verification of motor drive software for networked motion control systems. J.UCS: Journal of Universal Computer Science, 20(14):1903–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
14/formal_modeling_and_
verification.

Karahoca:2016:ACD

[KKK16] D. Karahoca, A. Karahoca, and A. Kurnaz. Analyzing communication dimensions in a ubiquitous learning environment. J.UCS: Journal of Universal Computer Science, 22(1):124–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_22_1/analyzing_
communication_dimensions_
in.

Khan:2008:AMU

[KKM08] M. S. Khan, N. Kulathuramaiyer, and H. Maurer. Applications of mashups for a digital journal. J.UCS: Journal of Universal Computer Science, 14(10):1695–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
10/applications_of_mash_
ups.

Kalampokis:2016:ABE

[KKTT16] E. Kalampokis, A. Karamanou, E. Tambouris, and K. Tarabanis. Applying brand equity theory to understand consumer opinion in social media. J.UCS: Journal of Universal Computer Science, 22(5):709–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
22_5/applying_brand_
equity_theory.

Kim:2009:IES

[KKTZ09] T.-H. Kim, A. Kusiak, D. Taniar, and D. Zhang. Intelligent environments and services. J.UCS: Journal of Universal Computer Science, 15(12):2284–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_15;   http:
//www.jucs.org/jucs_
15_12#;   http://www.
jucs.org/jucs_15_12/
intelligent_environments_
and_services.

Kasztler:2002:BAV

[KL02] A. Kasztler and K.- H. Leitner. Bibliometric analysis and visualisation of intellectual capital. J.UCS: Journal of Universal Computer Science, 8(5):516– 525, May 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
5/bibliometric_analysis_
and_visualisation.

Krol:2009:AMF

[KL09a] D. Król and A. Lupa. Agent migration: Framework for analysis. J.UCS: Journal of Universal Computer Science, 15(4): 941–??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_4/agent_migration_
framework_for.

Kwon:2009:CJQ

[KL09b] O. Kwon and K.-J. Li.

Causality join query processing for data streams via a spatiotemporal sliding window. J.UCS: Journal of Universal Computer Science, 15(12): 2287–??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_12/causality_join_
query_processing.

Kazienko:2013:HEM

[KLT13] P. Kazienko, E. Lughofer, and B. Trawiński. Hybrid and ensemble methods in machine learning. J.UCS: Journal of Universal Computer Science, 19(4):457–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_19;   http:
//www.jucs.org/jucs_
19_4/hybrid_and_ensemble_
methods.

Kornerup:1995:LLB

[KM95] Peter Kornerup and David W.

Matula. LCF: a lexicographic binary representation of the rationals. J.UCS: Journal of Universal Computer Science, 1(7):484–502, July 28, 1995. CODEN ????

ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/a_lexicographic_
                                        binary_representation_
                                        of_the_rationals.

Krottmaier:2001:TC

[KM01] H. Krottmaier and H. Maurer. Transclusions in the 21st Century. J.UCS: Journal of Universal Computer Science, 7(12): 1125, December 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_7_
                                        12/transclusions_in_
                                        the_21st.

Kolbitsch:2006:TWH

[KM06] J. Kolbitsch and H. Maurer. The transformation of the Web: How emerging communities shape the information we consume. J.UCS: Journal of Universal Computer Science, 12(2):187–213, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12_
                                        2/the_transformation_
                                        of_the.

Komatsu:2007:ERU

[KM07] T. Komatsu and K. Morikawa. Entrainment in the rate of utterances in speech dialogs between users and an auto response system. J.UCS: Journal of Universal Computer Science, 13(2):186–198, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
2/entrainment_in_the_
rate.

Kalyvioti:2013:VRT

[KM13] K. Kalyvioti and T. A.

Mikropoulos. A virtual reality test for the identification of memory strengths of dyslexic students in higher education. J.UCS: Journal of Universal Computer Science, 19(18):2698–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
18/a_virtual_reality_
test.

Kulathuramaiyer:2014:SAR

[KMM14] N. Kulathuramaiyer, H. Maurer, and R. Mehmood. Some aspects of the reliability of information on the Web. J.UCS: Journal of Universal Computer Science, 20(9):1284– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
9/some_aspects_of_the.

Kumar:2016:AID

[KMN16] G. R. Kumar, N. Mangathayaru, and G. Narsimha. An approach for intrusion detection using novel Gaussian based kernel function. J.UCS: Journal of Universal Computer Science, 22(4):589–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        22_4/an_approach_for_
                                        intrusion.

Kaderali:1996:MPD

[KMR96] F. Kaderali, H. Müller, and A. Rieke. Media publishing in distance teaching. J.UCS: Journal of Universal Computer Science, 2(6):492– 502, June 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        2_6/media_publishing_
                                        in_distance.

Karsai:2000:TTL

[KNLS00] G. Karsai, G. Nordstrom, A. Ledeczi, and J. Sztipanovits. Towards twolevel formal modeling of computer-based systems. J.UCS: Journal of Universal Computer Science, 6(11):1131–1144, November 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
11/towards_two_level_
formal.

Kubota:2007:SMS

[KNSN07] H. Kubota, S. Nomura, Y. Sumi, and T. Nishida. Sustainable memory system using global and conical spaces. J.UCS: Journal of Universal Computer Science, 13(2):135– 148, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
2/sustainable_memory_
system_using.

Kreitz:1999:CBT

[KO99] Ch. Kreitz and J. Otten.

Connection-based theorem proving in classical and non-classical logics. J.UCS: Journal of Universal Computer Science, 5(3):88–112, March 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_5_3/connection_
based_theorem_proving.

Kohler:2009:GDD

[Köh09]  H. Köhler.  Global

database design based on storage space and update time minimization. J.UCS: Journal of Universal Computer Science, 15(1):195–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        15_1/global_database_
                                        design_based.

Kolbitsch:2005:FGT

[Kol05]  J. Kolbitsch.  Fine-

grained transclusions of multimedia documents in HTML. J.UCS: Journal of Universal Computer Science, 11(6):926– 943, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_11_
                                        6/fine_grained_transclusions_
                                        of.

Komlodi:2002:RIH

[Kom02] A. Komlodi. The role of interaction histories in mental model building and knowledge sharing in the legal domain. J.UCS: Journal of Universal Computer Science, 8(5):557–566, May 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_8_5/the_role_of_
                                        interaction.

Konstantinidis:2002:ECF

[Kon02] S. Konstantinidis. Errorcorrection, and finitedelay decodability. J.UCS: Journal of Universal Computer Science, 8 (2):278–291, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
8_2/error_correction_
and_finite.

Kone:2003:ITA

[Kon03] O. Koné. An inoteroperability testing approach to wireless applications protocols. J.UCS: Journal of Universal Computer Science, 9(10):1220–1243, October 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
10/an_interoperability_
testing_approach.

Koos:2006:MME

[Koó06] I. Koós. Mathematical models of endocrine systems. J.UCS: Journal of Universal Computer Science, 12(9):1267–1277, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
9/mathematical_models_
of_endocrine.

Koukopoulos:2009:SHM

[Kou09] D. Koukopoulos. Stability in heterogeneous multimedia networks under adversarial attacks.

J.UCS: Journal of Universal Computer Science, 15(2):444–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_15_2/stability_
                                        in_heterogeneous_multimedia.

Kojiri:2001:AOS

[KOW01] T. Kojiri, Y. Ogawa, and T. Watanabe. Agentoriented support environment in Web-based collaborative learning. J.UCS: Journal of Universal Computer Science, 7(3): 226–239, March 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_7_
                                        3/agent_oriented_support_
                                        environment.

Klauser:1995:DCN

[KP95] A. Klauser and R. Posch. Distributed caching in networked file systems. J.UCS: Journal of Universal Computer Science, 1(6):399–409, June 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        Distributed_Caching_
                                        in_Networked_File_Systems.

Kutter:1997:FSO

[KP97a] Ph. W. Kutter and A. Pierantonio. The formal specification of Oberon. J.UCS: Journal of Universal Computer Science, 3(5):443– 503, May 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs_3_5/specification_
oberon;   http://www.
jucs.org/jucs_3_5/specification_
oberon; internal&sk=
05460486.

Kutter:1997:MSR

[KP97b] Ph. W. Kutter and A. Pierantonio. Montages specifications of realistic programming languages. J.UCS: Journal of Universal Computer Science, 3(5):416– 442, May 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs_3_5/montages_
specifications;   http:
//www.jucs.org/jucs_
3_5/montages_specifications;
 internal&sk=05460486.

Kronenburg:2000:AFA

[KP00] M. Kronenburg and Ch.

Peper. Application of the FOREST approach to the light control case study. J.UCS: Journal of Universal Computer Science, 6(7):679– 703, July 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_6_
                                        7/application_of_the_
                                        forest.

Kappe:2001:FCS

[KP01] F. Kappe and G. Pail. Future of computer science: J.UCS special issue dedicated to Professor Hermann Maurer. J.UCS: Journal of Universal Computer Science, 7(5):335–337, May 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_7_5/future_
                                        of_computer_science.

Kowada:2006:RKA

[KPdF06] L. A. B. Kowada, R. Portugal, and C. M. Herrera de Figueiredo. Reversible Karatsuba’s algorithm. J.UCS: Journal of Universal Computer Science, 12(5):499–511, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_12_5/reversible_
                                        karatsubas_algorithm.

Kari:1996:PRS

[KPS96] Lila Kari, Gheorghe Păun, and Arto Salomaa. The power of restricted splicing with rules from a regular language. J.UCS: Journal of Universal Computer Science, 2(4):224– 240, April 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
4/the_power_of_restricted.

Kotis:2011:ECK

[KPV^{+}11] K. Kotis, A. Papasalouros, G. Vouros, N. Pappas, and K. Zoumpatianos. Enhancing the collective knowledge for the engineering of ontologies in open and socially constructed learning spaces. J.UCS: Journal of Universal Computer Science, 17(12): 1710–??, ???? 2011. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_12/enhancing_the_
collective_knowledge.

Knoop:2003:CPP

[KR03a] J. Knoop and O. Rüthing.

Constant propagation on predicated code. J.UCS: Journal of Universal Computer Science, 9(8): 829–850, August 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
8/constant_propagation_
on_predicated.

Krebs:2003:OTL

[KR03b] A. Krebs and J. Ruf. Optimized temporal logic compilation. J.UCS: Journal of Universal Computer Science, 9 (2):120–137, February 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        2/optimized_temporal_
                                        logic_compilation.

Kuswara:2011:RPW

[KR11] A. U. Kuswara and D. Richards. Realising the potential of Web 2.0 for collaborative learning using affordances. J.UCS: Journal of Universal Computer Science, 17(2):311–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_17_2/realising_
                                        the_potential_of.

Kramer:1998:CEA

[Krä98]  W. Krämer.  Con-

structive error analysis. J.UCS: Journal of Universal Computer Science, 4(2):147–163, February 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:

8000/jucs/jucs_4_2/constructive_
error_analysis.

Krishnan:1997:ACB

[Kri97] P. Krishnan. An asynchronous calculus based on the absence of actions. J.UCS: Journal of Universal Computer Science, 3(12):1415, December 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_3_12/
an_asynchronous_calculus_
based.

Krishnan:1999:IRD

[Kri99] P. Krishnan. Issues related to distributed processing of picture languages. J.UCS: Journal of Universal Computer Science, 5(9):542– 551, September 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
9/issues_related_to_
distributed.

Kroop:2013:EST

[Kro13] S. Kroop. Evaluation on students’ and teachers’ acceptance of widgetand cloud-based personal learning environments. J.UCS: Journal of Universal Computer Science, 19(14):2150–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        19_14/evaluation_on_
                                        students_and.

Keim:2005:SVD

[KS05] D. A. Keim and J. Schneidewind. Scalable visual data exploration of large data sets via MultiResolution. J.UCS: Journal of Universal Computer Science, 11(11): 1766–1779, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_11_11/scalable_
                                        visual_data_exploration.

Kirchberg:2010:ETC

[KS10] M. Kirchberg and K.-D. Schewe. Evolving theories of conceptual modelling. J.UCS: Journal of Universal Computer Science, 16(20):2902–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_16;
                                         http://www.jucs.org/
                                        jucs_16_20/evolving_
                                        theories_of_onceptual;
                                         http://www.jucs.org/
                                        jucsrssfeed-issue.

Kolfschoten:2009:HIW

[KSdV09] G. L. Kolfschoten, M. Seck, and G.-J. de Vreede. How interactive whiteboards can be used to support collaborative modeling. J.UCS: Journal of Universal Computer Science, 15(16):3126–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
16/how_interactive_whiteboards_
can.

Karim:2016:ADM

[KSK^{+}16] A. Karim, R. Salleh, M. K. Khan, A. Siddiqa, and K.-K. R. Choo. On the analysis and detection of mobile botnet applications. J.UCS: Journal of Universal Computer Science, 22(4):567– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_22_
4/on_the_analysis_and.

Kim:2016:ORT

[KSR16] Y.-S. Kim, Y.-I. Song, and H.-C. Rim. Opinion retrieval for Twitter using extrinsic information. J.UCS: Journal of Universal Computer Science, 22(5):608– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_22_

                                        5/opinion_retrieval_
                                        for_twitter.

Kuich:2002:HAF

[KSU02] W. Kuich, N. Sauer, and F. Urbanek. Heyting algebras and formal languages. J.UCS: Journal of Universal Computer Science, 8(7):722– 736, July 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        8_7/heyting_algebras_
                                        and_formal.

Kurosawa:1997:HNR

[KSY97] K. Kurosawa, T. Satoh, and K. Yamamoto. Highly nonlinear t-resilient functions. J.UCS: Journal of Universal Computer Science, 3(6):721– 729, June 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs_3_6/highly_
                                        nonlinear_t_resilient;
                                         internal&sk=05460486.

Kim:2010:FPU

[KT10] D.-K. Kim and J. Trujillo. Foundations and practices of unified modeling language. J.UCS: Journal of Universal Computer Science, 16(17):2291–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16;   http:
//www.jucs.org/jucs_
16_17/foundations_and_
practices_of;   http:
//www.jucs.org/jucsrssfeed-
issue.

KhorI:2005:SWP

[KTJ05] J. Khor I., J. Thomas, and I. Jonyer. Sliding window protocol for secure group communication in ad-hoc networks. J.UCS: Journal of Universal Computer Science, 11(1):37– 55, January 28, 2005. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_1/sliding_window_
protocol_for.

Kotsis:2009:IIW

[KTKP09] G. Kotsis, D. Taniar, I. Khalil, and E. Pardede. Information integration on Web-based applications and services. J.UCS: Journal of Universal Computer Science, 15(10):2026–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_15;   http:
//www.jucs.org/jucs_
15_10#;   http://www.
jucs.org/jucs_15_10/
information_integration_
on_web.

Konow:2011:VIF

[KTL^{+}11] R. Konow, W. Tan, L. Loyola, J. Pereira, and N. Baloian. A visited item frequency based recommender system: Experimental evaluation and scenario description. J.UCS: Journal of Universal Computer Science, 17(14): 2009–??, ???? 2011. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        17_14/a_visited_item_
                                        frequency.

Keller:2000:JUS

[KU00] J. Keller and T. Ungerer. J.UCS special issue on multithreaded processors and chip-multiprocessors. J.UCS: Journal of Universal Computer Science, 6(10):906–907, October 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_6_10/j_ucs_
                                        special_issue.

Kabak:2010:DFM

[KÜ10] Ö. Kabak and F. Ülengin. A demand forecasting methodology for fuzzy environments. J.UCS: Journal of Universal Computer Science, 16(1):121– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
1/a_demand_forecasting_
methodology.

Kudlek:1999:SAS

[Kud99] M. Kudlek. Some algebraic structures with iteration lemmata. J.UCS: Journal of Universal Computer Science, 5 (9):552–562, September 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
9/some_algebraic_structures_
with.

Kuhn:2003:NCS

[Kuh03] M. Kuhn. A note on culture-sensitive knowledge management in OEsales area of Robert Bosch GmbH. J.UCS: Journal of Universal Computer Science, 9(7):592– 595, July 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
7/a_note_on_culture.

Kulikowski:2005:CIP

[Kul05] J. L. Kulikowski. Creation of information profiles in distributed databases as a game problem. J.UCS: Journal of Universal Computer Science, 11(2):271–284, February 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_11_
                                        2/creation_of_information_
                                        profiles.

Kulathuramaiyer:2007:MEA

[Kul07] N. Kulathuramaiyer. Mashups: Emerging application development paradigm for a digital journal. J.UCS: Journal of Universal Computer Science, 13(4): 531–542, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        13_4/mashups_emerging_
                                        application_development.

Kuznetsov:2004:ICD

[Kuz04] S. O. Kuznetsov. On the intractability of computing the Duquenne– Guigues base. J.UCS: Journal of Universal Computer Science, 10 (8):927–933, August 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_10_8/on_the_
                                        intractability_of.

Koo:2010:TUM

[KW10] C. Koo and Y. Wati. Toward an understanding of the mediating role of “trust” in mobile banking service: An empirical test of Indonesia case. J.UCS: Journal of Universal Computer Science, 16(13):1801–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_13/toward_an_
understanding_of.

Kuo:2001:SEW

[KWC01] C. H. Kuo, D. Wible, and C. L. Chou. A synchronous EFL writing environment for the Internet. J.UCS: Journal of Universal Computer Science, 7(3):240– 253, March 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
3/a_synchronous_efl_
writing.

Kamentz:2003:DCB

[KWH03] E. Kamentz and C. Womser-Hacker. Defining culturebound user characteristics as a starting-point for the design of adaptive learning systems. J.UCS: Journal of Universal Computer Science, 9(7):596– 607, July 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_

                                        9_7/defining_culture_
                                        bound_user.

Kwon:1997:SPC

[Kwo97]  K. Kwon.  A struc-

tured presentation of a closure-based compilation method for a scoping notion in logic programming. J.UCS: Journal of Universal Computer Science, 3(4):341– 376, April 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs_3_4/compilation_
                                        method;   http://www.
                                        jucs.org/jucs_3_4/a_
                                        structured_presentation_
                                        of;   internal&sk=05460486.

Kim:2010:THM

[KY10] D. W. Kim and J. Yao. A treasure hunt model for inquiry-based learning in the development of a Web-based learning support system. J.UCS: Journal of Universal Computer Science, 16 (14):1853–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        16_14/a_treasure_hunt_
                                        model.

Knoop:2003:JUS

[KZ03] J. Knoop and W. Zimmermann. J.UCS special issue on Compiler Optimization meets Compiler Verification (COCV 2002). J.UCS: Journal of Universal Computer Science, 9(3):189– 190, March 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
 http://www.jucs.org/
jucs_9_3/j_ucs_special_
issue.

Krol:2008:SPE

[KZ08] D. Król and M. Zelmozer.

Structural performance evaluation of multi-agent systems. J.UCS: Journal of Universal Computer Science, 14(7):1154– 1178, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_7/structural_
performance_evaluation_
of.

Ley:2003:IEC

[LA03a] T. Ley and D. Albert.

Identifying employee competencies in dynamic work domains: Methodological considerations and a case study. J.UCS: Journal of Universal Computer Science, 9(12):1500– 1518, December 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_

                                        12/identifying_employee_
                                        competencies_in.

Ley:2003:SMM

[LA03b] T. Ley and D. Albert. Skills management — managing competencies in the knowledge-based economy — J.UCS special issue. J.UCS: Journal of Universal Computer Science, 9(12):1370– 1371, December 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/free;
                                         http://www.jucs.org/
                                        jucs_9_12/skills_management_
                                        managing_competencies.

Liu:2007:HFV

[LA07] H. Liu and A. Abraham. An hybrid fuzzy variable neighborhood particle swarm optimization algorithm for solving quadratic assignment problems. J.UCS: Journal of Universal Computer Science, 13(9):1309– 1331, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        13_9/an_hybrid_fuzzy_
                                        variable.

Lozano-Alvarez:2015:HTA

[LAAPVGMM15] A. Lozano-Alvarez, J. I. Asensio-Pérez, G. Vega-Gorgojo, and A. Martínez-Monés. Helping teachers align learning objectives and evidence: Integration of ePortfolios in distributed learning environments. J.UCS: Journal of Universal Computer Science, 21(8):1022– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
8/helping_teachers_align_
learning.

Lafrance:2004:SAA

[Laf04] S. Lafrance. Symbolic approach to the analysis of security protocols. J.UCS: Journal of Universal Computer Science, 10(9):1156–1198, September 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
9/symbolic_approach_
to_the.

Lytras:2015:ETB

[LAHZ^{+}15] M. D. Lytras, W. Al-Halabi, J. X. Zhang, M. Masud, and R. A. Haraty. Enabling technologies and business infrastructures for next generation social media: Big Data, cloud computing, Internet of Things and virtual reality. J.UCS: Journal of Universal Computer Science, 21(11):1379–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_21;   http:
                                        //www.jucs.org/jucs_
                                        21_11/enabling_technologies_
                                        and_business.

Lamprier:2008:UET

[LALS08] S. Lamprier, T. Amghar, B. Levrat, and F. Saubion. Using an evolving thematic clustering in a text segmentation process. J.UCS: Journal of Universal Computer Science, 14(2):178–192, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        2/using_an_evolving_
                                        thematic.

Latif:2012:WSL

[LAM12] A. Latif, M. T. Afzal, and H. Maurer. Weaving scholarly legacy data into Web of data. J.UCS: Journal of Universal Computer Science, 18 (16):2301–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        18_16/weaving_scholarly_
                                        legacy_data.

Lang:1998:VQD

[Lan98] B. Lang. Verified quadrature in determining Newton’s constant of gravitation. J.UCS: Journal of Universal Computer Science, 4(1):16–24, January 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_1/verified_
quadrature_in_determining.

Lange:2010:NCD

[Lan10] K.-J. Lange. A note on the P -completeness of deterministic one-way stack language. J.UCS: Journal of Universal Computer Science, 16(5):795– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
5/a_note_on_the.

Larsen:2001:TYH

[Lar01] P. G. Larsen. Ten years of historical development “Bootstrapping” VDM-Tools(R). J.UCS: Journal of Universal Computer Science, 7(8):692– 709, August 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
8/ten_years_of_historical.

Lizcano:2012:SEU

[LASL12] D. Lizcano, F. Alonso, J. Soriano, and G. López. Supporting end-user development through a new composition model: An empirical study. J.UCS: Journal of Universal Computer Science, 18(2): 143–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_2/supporting_end_
user_development.

Lavenier:1996:DHB

[Lav96]  D. Lavenier.  Dedi-

cated hardware for biological sequence comparison. J.UCS: Journal of Universal Computer Science, 2(2):77–86, February 18, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
dedicated_hardware_for_
biological_sequence_
comparison.

Langenbach:1998:EBT

[LB98] Ch. Langenbach and F. Bodendorf. An education broker toolset for Web course customization. J.UCS: Journal of Universal Computer Science, 4(10):780–791, October 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_10/
an_education_broker_
toolset.

Li:2007:FAR

[LBG07] L. Li, V. Bulitko, and R. Greiner. Focus of attention in reinforcement learning. J.UCS: Journal of Universal Computer Science, 13(9):1246– 1269, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
9/focus_of_attention_
in.

Lopez-Colino:2011:SLC

[LCC11] F. López-Colino and J. Colás. The synthesis of LSE classifiers: From representation to evaluation. J.UCS: Journal of Universal Computer Science, 17(3):399– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
3/the_synthesis_of_lse.

Lee:2012:USD

[LCC^{+}12] J. B. Lee, G. Cabunducan, F. G. C. Cabarle, R. Castillo, and J. A. Malinao. Uncovering the social dynamics of online elections. J.UCS: Journal of Universal Computer Science, 18(4):487– ??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_4/uncovering_the_
social_dynamics.

Liu:2015:IEP

[LCDP15] H. Liu, L. Chen, Z. Davar, and M. R. Pour. Insecurity of an efficient privacypreserving public auditing scheme for cloud data storage. J.UCS: Journal of Universal Computer Science, 21(3):473– ??, ???? 2015. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_3/insecurity_
of_an_efficient.

Leal:2012:IAS

[LCHD12] G. C. Lapasini Leal, A. P. Chaves, E. H. M. Huzita, and M. E. Delamaro. An integrated approach of software development and test processes to distributed teams. J.UCS: Journal of Universal Computer Science, 18 (19):2686–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_19/an_integrated_
approach_of.

Liu:2012:NPW

[LCZ^{+}12] Q. Liu, M. Chen, J. Zhang, B. Shen, and Z. Chu. Network planning for WiMAX-R networks. J.UCS: Journal of Universal Computer Science, 18(9): 1194–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_9/network_planning_
for_wimax.

Liu:2006:ROE

[LD06] L. Liu and T.-S. Dai.

Ridge orientation estimation and verification algorithm for fingerprint enhancement. J.UCS: Journal of Universal Computer Science, 12(10):1426–1438, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
10/ridge_orientation_
estimation_and.

Lins:2007:CRC

[LdL07] R. Dueire Lins, F. Heron de Carvalho Jr., and Z. Dueire Lins. Cyclic reference counting with permanent objects. J.UCS: Journal of Universal Computer Science, 13(6): 830–838, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        13_6/cyclic_reference_
                                        counting_with.

Lee:2012:SNB

[LDO^{+}12] J. Y. Lee, Y. Duan, J. C. Oh, W. Du, H. Blair, L. Wang, and X. Jin. Social network based reputation computation and document classification. J.UCS: Journal of Universal Computer Science, 18(4):532–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        18_4/social_network_
                                        based_reputation.

Lytras:2011:STK

[LdP11] M. Lytras and P. Ordoñez de Pablos. Software technologies in knowledge society. J.UCS: Journal of Universal Computer Science, 17(9):1219–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_17;   http:
                                        //www.jucs.org/jucs_
                                        17_9/software_technologies_
                                        in_knowledge;   http:
                                        //www.jucs.org/jucsrssfeed-
                                        issue.

Lemos:2014:TBT

[LdPK^{+}14] O. A. L. Lemos, A. C. de Paula, G. Konishi, S. Bajracharya, J. Ossher, and C. Lopes. Thesaurusbased tag clouds for test-driven code search. J.UCS: Journal of Universal Computer Science, 20(5):772–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
20_5/thesaurus_based_
tag_clouds.

Lee:2009:FVS

[LDSG09] S. U.-J. Lee, G. Dobbie, J. Sun, and L. Groves. Formal verification of semistructured data models in PVS. J.UCS: Journal of Universal Computer Science, 15(1):241– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
1/formal_verification_
of_semistructured.

Lins:2008:DQM

[LdSM08] R. Dueire Lins, J. M. Monte da Silva, and F. M. Junqueira Martins. Detailing a quantitative method for assessing algorithms to remove back-to-front interference in documents. J.UCS: Journal of Universal Computer Science, 14(2):266–283, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_2/detailing_
                                        a_quantitative_method.

Levner:2011:IFM

[LEC11] E. Levner, A. Elalouf, and T. E. Cheng. An improved FPTAS for mobile agent routing with time constraints. J.UCS: Journal of Universal Computer Science, 17(13):1854–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_17_
                                        13/an_improved_FPTAS_
                                        for.

Leitao:2008:UGO

[Lei08] A. Menezes Leitão. UCL-GLORP — an ORM for Common Lisp. J.UCS: Journal of Universal Computer Science, 14 (20):3333–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_20/ucl_glorp_
                                        an_orm.

Leitao:2010:LRE

[Lei10] A. Menezes Leitão. Lisp: Research and experience. J.UCS: Journal of Universal Computer Science, 16(2):218–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16;   http:
//www.jucs.org/jucs_
16_01/lisp_research_
and_experience;   http:
//www.jucs.org/jucs_
16_2#.

Lindgaard:2008:IDS

[LEJ^{+}08] G. Lindgaard, P. Egan, C. Jones, C. Pyper, M. Frize, R. Walker, C. Boutilier, B. Hui, S. Narasimhan, J. Folkens, and B. Winogron. Intelligent decision support in medicine: back to Bayes? J.UCS: Journal of Universal Computer Science, 14(16):2720–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
16/intelligent_decision_
support_in.

Lennertz:2000:PAI

[Len00] J. Lennertz. Perceptions about Internet use by teaching faculty at small Christian colleges and universities. J.UCS: Journal of Universal Computer Science, 6(3):367– 404, March 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
3/perceptions_about_
internet_use.

Leppanen:1995:IEW

[Lep95] V. Leppanen. On implementing EREW workoptimally on mesh of trees. J.UCS: Journal of Universal Computer Science, 1(1):23–34, January 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        1_1/on_implementing_
                                        erew_work/html/abstract.
                                        html.

Leppanen:1998:BPS

[Lep98] V. Leppänen. Balanced PRAM simulations via moving threads and hashing. J.UCS: Journal of Universal Computer Science, 4(8):675–689, August 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_8/balanced_
                                        pram_simulations_via.

Lester:1995:ESC

[Les95]  David Lester.  Exact

statistics and continued fractions. J.UCS: Journal of Universal Computer Science, 1(7):504– 513, July 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/

exact_statistics_and_
continued_fractions.

Lesnik:2009:CUU

[Leš09] D. Lešnik. Constructive Urysohn universal metric space. J.UCS: Journal of Universal Computer Science, 15(6):1236– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
6/constructive_urysohn_
universal_metric.

Leustean:2007:RAR

[Leu07]  L. Leustean.  Rates

of asymptotic regularity for Halpern iterations of nonexpansive mappings. J.UCS: Journal of Universal Computer Science, 13(11):1680–1691, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_13_11/rates_of_
asymptotic_regularity.

Liu:1998:WTL

[LF98] B. Liu and E. A. Fox.

Web traffic latency: Characteristics and implications. J.UCS: Journal of Universal Computer Science, 4(9):763– 778, September 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_9/web_
                                        traffic_latency_characteristics.

Lindstaedt:2005:IKM

[LF05] S. N. Lindstaedt and J. Farmer. Integration of knowledge management and (e)Learning. J.UCS: Journal of Universal Computer Science, 11(3): 375–377, March 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/free;
                                         http://www.jucs.org/
                                        jucs_11_3/integration_
                                        of_knowledge_management.

Lopez-Fraguas:2006:PL

[LF06] F. J. López-Fraguas. Programming and languages. J.UCS: Journal of Universal Computer Science, 12(11):1464–1465, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_12;   http:
                                        //www.jucs.org/jucs_
                                        12_11#;   http://www.
                                        jucs.org/jucs_12_11/
                                        programming_and_languages.

Lamei:2016:RBE

[LF16] A. Lamei and M. S. Fallah. Rewriting-based enforcement of noninterference in programs with observable intermediate values. J.UCS: Journal of Universal Computer Science, 22(7):956– ??, ???? 2016. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
22_7/rewriting_based_
enforcement_of.

Lozano:2008:NTH

[LG08] M. D. Lozano and J. A.

Gallud. New trends in human computer interaction. J.UCS: Journal of Universal Computer Science, 14(19):3051–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14;
 http://www.jucs.org/
jucs_14_19#;   http:
//www.jucs.org/jucs_
14_19/new_trends_in_
human.

Labraoui:2011:RRA

[LGAP11] N. Labraoui, M. Gueroui, M. Aliouat, and J. Petit. RAHIM: Robust adaptive approach based on hierarchical monitoring providing trust aggregation for wireless sensor networks. J.UCS: Journal of Universal Computer Science, 17(11):1550–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.

                                        org/jucs_17_11/rahim_
                                        robust_adaptive_approach.

Lin:2011:IFK

[LGES11] D. Lin, P. Geißler, S. Ehrlich, and E. Schoop. IDEA: a framework for a knowledge-based enterprise 2.0. J.UCS: Journal of Universal Computer Science, 17(4):515–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_17_4/idea_a_
                                        framework_for.

Lopez-Gil:2014:EOD

[LGGGC14] J.-M. López-Gil, R. Gil, R. García, and C. A. Collazos. EmotionsOnto: an ontology for developing affective applications. J.UCS: Journal of Universal Computer Science, 20(13):1813–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_20_
                                        13/emotions_onto_an_
                                        ontology.

Li:2009:AIS

[LGL09] W. Li, L. Gao, and X. Li. Application of intelligent strategies for cooperative manufacturing planning. J.UCS: Journal of Universal Computer Science, 15(9):1907–??, ???? 2009.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_9/application_of_
intelligent_strategies.

Leony:2013:GAE

[LGMM^{+}13] D. Leony, H. A. Parada Gélvez, P. J. Muñoz-Merino, A. Pardo, and C. Delgado Kloos. A generic architecture for emotion-based recommender systems in cloud learning environments. J.UCS: Journal of Universal Computer Science, 19(14):2075–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_14/a_generic_
architecture_for.

Lozano:2010:MCA

[LGP10] M. D. Lozano, J. A. Gallud, and P. Palanque. Mobile context-aware applications for ubiquitous computing. J.UCS: Journal of Universal Computer Science, 16(15): 1927–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16;   http://www.jucs.
org/jucs_16_15/mobile_
context_aware_applications;

                                         http://www.jucs.org/
                                        jucsrssfeed-issue.

Liu:2001:HDF

[LGZ01] Y. Liu, D. Ginther, and P. Zelhart. How do frequency and duration of messaging affect impression development in computer-mediated communication? J.UCS: Journal of Universal Computer Science, 7(10): 893–913, October 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_7_10/how_do_
                                        frequency_and.

Li:2003:MOK

[LH03] S.-T. Li and H.-C. Hsieh. Managing operation knowledge for the metal industry. J.UCS: Journal of Universal Computer Science, 9(6):472– 480, June 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        6/managing_operation_
                                        knowledge_for.

Lin:2013:IMT

[LHC^{+}13] C.-W. Lin, T.-P. Hong, Y.-F. Chen, T.-C. Lin, and S.-T. Pan. An integrated MFFP-tree algorithm for mining global fuzzy rules from distributed databases. J.UCS: Journal of Universal Computer Science, 19(4):521–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_4/an_integrated_mffp_
tree.

Le:2013:SSE

[LHK^{+}13] T. N. Le, T. B. Ho, S. Kawasaki, T. Kanda, K. Takabayashi, S. Wu, and O. Yokosuka. A semi-supervised ensemble learning method for finding discriminative motifs and its application. J.UCS: Journal of Universal Computer Science, 19(4):563–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_19_4/a_semi_
supervised_ensemble.

Labhart:2012:SLG

[LHZS12] N. Labhart, B. S. Hasler, A. Zbinden, and A. Schmeil. The ShanghAI lectures: a global education project on artificial intelligence. J.UCS: Journal of Universal Computer Science, 18(18):2542–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/

                                        jucs_18_18/the_shanghai_
                                        lectures_a.

Leal:2005:FSF

[LI05] M. Amorim Leal and R. Ierusalimschy. A formal semantics for finalizers. J.UCS: Journal of Universal Computer Science, 11(7):1198–1214, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_11_
                                        7/a_formal_semantics_
                                        for.

Lins:2003:LCR

[Lin03]  R. D. Lins.  Lazy

cyclic reference counting. J.UCS: Journal of Universal Computer Science, 9(8):813–828, August 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

Lindstaedt:2004:VCP

[Lin04a] S. N. Lindstaedt. (virtual) communities of practice within modern organizations — J.UCS special issue. J.UCS: Journal of Universal Computer Science, 10(3):158– 161, March 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/free;
                                         http://www.jucs.org/
                                        jucs_10_3/virtual_communities_
                                        of_practice.

Lins:2004:BSP

[Lin04b] R. D. Lins. 8th Brazilian Symposium on Programming Languages — J.UCS special issue. J.UCS: Journal of Universal Computer Science, 10(7): 749–750, July 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
http://www.jucs.org/
jucs_10_7/brazilian_
symposium_on.

Lins:2004:PCM

[Lin04c] R. D. Lins. Partial categorical multi-combinators and Church–Rosser theorems. J.UCS: Journal of Universal Computer Science, 10(7):769–788, July 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
7/partial_categorical_
multi_combinators.

Lins:2008:ADE

[Lin08a] R. Dueire Lins. Advances in document engineering. J.UCS: Journal of Universal Computer Science, 14(2):174– 177, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14;
http://www.jucs.org/

jucs_14_2#;   http://
www.jucs.org/jucs_14_
2/advances_in_document_
engineering.

Lins:2008:FAD

[Lin08b] R. Dueire Lins. Further advances in document engineering. J.UCS: Journal of Universal Computer Science, 14(18):2909–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_14;   http:
//www.jucs.org/jucs_
14_18#;   http://www.
jucs.org/jucs_14_18/
further_advances_in_
document.

Lins:2009:PCB

[Lin09] R. Dueire Lins. Processing camera-based documents. J.UCS: Journal of Universal Computer Science, 15(18):3158–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_15;   http:
//www.jucs.org/jucs_
15_18#;   http://www.
jucs.org/jucs_15_18/
processing_camera_based_
documents.

Lins:2011:MNC

[Lin11a] R. Dueire Lins. Meeting new challenges in document engineering.

J.UCS: Journal of Universal Computer Science, 17(1):1–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_17;   http:
//www.jucs.org/jucs_
17_1/meeting_new_challenges_
in;   http://www.jucs.
org/jucsrssfeed-issue.

Lins:2011:NTD

[Lin11b] R. Dueire Lins. Nabuco — two decades of document processing in Latin America. J.UCS: Journal of Universal Computer Science, 17(1):151– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
1/nabuco_two_decades_
of.

Lipponen:2000:CCS

[Lip00]  M. Lipponen.  Com-

putational complementarity and shift spaces. J.UCS: Journal of Universal Computer Science, 6(1):169–177, January 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
1/computational_complementarity_
and_shift.

Lipp:2010:TC

[Lip10] P. Lipp. Trusted computing. J.UCS: Journal of Universal Computer Science, 16(4):506–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_16;   http:
                                        //www.jucs.org/jucs_
                                        16_04/trusted_computing;
                                         http://www.jucs.org/
                                        jucs_16_4#.

Liyanage:2002:KIC

[Liy02]  S. Liyanage.  Knowl-

edge and intellectual capital management processes: Grounding knowledge and understanding of organisational learning. J.UCS: Journal of Universal Computer Science, 8(5):526–535, May 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        5/knowledge_and_intellectual_
                                        capital.

Li:2000:TSP

[LJ00] W.-N. Li and J.-F. Jenq. On the thread scheduling problem. J.UCS: Journal of Universal Computer Science, 6(10): 994–1014, October 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_6_10/on_the_
thread_scheduling.

Lezcano:2013:PII

[LJLCR13] L. Lezcano, B. Jörg, B. Lowe, and J. Corson-Rikert. Promoting international interoperability of research information systems: VIVO and CERIF. J.UCS: Journal of Universal Computer Science, 19(12): 1854–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_12/promoting_
international_interoperability_
of.

Lugger:2001:MHB

[LK01] K.-M. Lugger and H. Kraus.

Mastering the human barriers in knowledge management. J.UCS: Journal of Universal Computer Science, 7(6):488– 497, June 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
6/mastering_the_human_
barriers.

Lux:2002:XMI

[LKB^{+}02] Mathias Lux, Werner Klieber, Jutta Becker, Klaus Tochtermann, Harald Mayer, Helmut Neuschmied, and Werner Haas. XML and MPEG-7 for interactive annotation and retrieval using semantic meta-data. J.UCS: Journal of Universal Computer Science, 8(10):965– 984, October 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
10/xml_and_mpeg_7.

Lee:2009:SAF

[LKHL09] C. Lee, J. Kim, S. Hong, and Y.-S. Lee. Security analysis of the full-round CHESS-64 cipher suitable for pervasive computing environments. J.UCS: Journal of Universal Computer Science, 15(5):1007– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
5/security_analysis_
of_the.

Lee:2008:KDA

[LKK08] S. Lee, H. Kim, and C. Kim. A knowledge discovery agent for a topology bit-map in ad hoc mobile networks. J.UCS: Journal of Universal Computer Science, 14(7):1105–1117, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
7/a_knowledge_discovery_
agent.

Llados:2008:GAC

[LKMS08] J. Lladós, D. Karatzas, J. Mas, and G. Sánchez. A generic architecture for the conversion of document collections into semantically annotated digital archives. J.UCS: Journal of Universal Computer Science, 14 (18):2912–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_18/a_generic_
architecture_for.

Lee:2011:CSS

[LKP11] S. M. Lee, D. S. Kim, and J. S. Park. Costsensitive spam detection using parameters optimization and feature selection. J.UCS: Journal of Universal Computer Science, 17(6):944– ??, ???? 2011. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_6/cost_sensitive_
spam_detection.

Lepik:2010:ILL

[LKT10] K.-L. Lepik, M. Krigul, and E. Terk. Introducing Living Lab’s method as knowledge transfer from one socio-institutional context to another: Evidence from Helsinki– Tallinn cross-border region. J.UCS: Journal of Universal Computer Science, 16(8):1089–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
8/introducing_living_
labs_method.

Lu:2010:VSP

[LKZK10] L. Lu, D.-K. Kim, Y. Zhu, and S. Kim. Verification of structural pattern conformance using logic programming. J.UCS: Journal of Universal Computer Science, 16(17): 2455–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_17/verification_of_
structural_pattern.

Lee:1997:GTP

[LL97] Y. D. Lee and B. H. Lee. Genetic trajectory planner for a manipulator with acceleration parametrization. J.UCS: Journal of Universal Computer Science, 3(9):1056–1073, September 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_3_9/genetic_
trajectory_planner_for.

Lundberg:2012:PAF

[LL12] J. Lundberg and W. Lowe.

Points-to analysis: a fine-grained evaluation. J.UCS: Journal of Universal Computer Science, 18(20):2851–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_20/points_to_
analysis_a.

Leong:2009:ADP

[LLCN09] K. Leong, J. Li, S. Chan, and V. Ng. An application of the dynamic pattern analysis framework to the analysis of spatial-temporal crime relationships. J.UCS: Journal of Universal Computer Science, 15(9):1852– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
9/an_application_of_
the.

Lee:2003:KMC

[LLH03] D. T. Lee, G. C. Lee, and Y. W. Huang. Knowledge management for computational problem solving. J.UCS: Journal of Universal Computer Science, 9(6):563– 570, June 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
6/knowledge_management_
for_computational.

Lee:1999:EPA

[LLLL99] W. Lee, S. Lee, B. Lee, and Y. Lee. An efficient planning algorithm for multi-head surface mounting machines using a genetic algorithm. J.UCS: Journal of Universal Computer Science, 5(12):833, December 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
12/an_efficient_planning_
algorithm.

Lennon:2002:HHO

[LLM02] J. Lennon, H. Liu, and H. Maurer. HWOES: a Hyperwave Online Employment Service. J.UCS: Journal of Universal Computer Science, 8(4): 409–422, April 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
4/hwoes_a_hyperwave_
online.

Li:2005:OMC

[LLS05] H.-F. Li, S.-Y. Lee, and M.-K. Shan. Online mining changes of items over continuous appendonly and dynamic data streams. J.UCS: Journal of Universal Computer Science, 11(8):1411–1425, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_11_8/online_
mining_changes_of.

Lorenzo:2013:LLE

[LLSA13] C.-M. Lorenzo, L. Lezcano, and S. Sánchez-Alonso. Language learning in educational virtual worlds — a TAM based assessment. J.UCS: Journal of Universal Computer Science, 19(11): 1615–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_11/language_
learning_in_educational.

Liu:2012:LCH

[LLYC12] X. Liu, J. Liu, K.-Y. Yoo, and H. Cho. Low complexity H.264/AVC intraframe coding for wireless multimedia sensor network. J.UCS: Journal of Universal Computer Science, 18(9):1177– ??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        18_9/low_complexity_
                                        H_264.

Luo:2016:DMR

[LLZ16] G.-X. Luo, Y. Liu, and Z.-Y. Zhang. A dynamic model of reposting information propagation based on empirical analysis and Markov process. J.UCS: Journal of Universal Computer Science, 22(3):360– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_22_
                                        3/a_dynamic_model_of.

Lennon:1994:AHS

[LM94a] J. Lennon and H. Maurer. Applications of hypermedia systems impact: An overview. J.UCS: Journal of Universal Computer Science, 0(0):54–108, November 15, 1994. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/applications_and_
                                        impact.

Lennon:1994:YBY

[LM94b] J. Lennon and H. Maurer. You believe you know what multimedia is? and what Internet will do for you? well … think again! J.UCS: Journal of Universal Computer Science, 0(0):137, November 15, 1994. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_0_0/you_believe_
you_know.

Lennon:2001:CKM

[LM01] J. Lennon and H. Maurer. Can knowledge management help in povertystricken countries and crisis situations. J.UCS: Journal of Universal Computer Science, 7(4): 327–334, April 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
4/can_knowledge_management_
help.

Lafrance:2003:IFM

[LM03a] S. Lafrance and J. Mullins.

An information flow method to detect denial of service vulnerabilities. J.UCS: Journal of Universal Computer Science, 9 (11):1350–1369, November 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
9_11/an_information_
flow_method.

Lennon:2003:WID

[LM03b] J. Lennon and H. Maurer. Why it is difficult to introduce e-learning into schools and some new solutions. J.UCS: Journal of Universal Computer Science, 9(10):1244– 1257, October 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_9_10/why_it_is_
                                        difficult.

Levit:2007:MSB

[LM07] V. E. Levit and E. Mandrescu. Matrices and αstable bipartite graphs. J.UCS: Journal of Universal Computer Science, 13(11):1692–1706, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_13_11/matrices_
                                        and_alpha_stable.

Laguna:2010:USD

[LM10] M. A. Laguna and J. M. Marqués. UML support for designing software product lines: The package merge mechanism. J.UCS: Journal of Universal Computer Science, 16(17):2313–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
17/uml_support_for_designing.

Lucas:2015:VRO

[LM15] M. Lucas and A. Moreira.

A visual representation of online interaction patterns. J.UCS: Journal of Universal Computer Science, 21(11):1496–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_11/a_visual_
representation_of.

Lytras:2014:SMA

[LMA^{+}14] M. D. Lytras, H. I.

Mathkour, H. Abdalla, C. Yanez-Marquez, and P. Ordóñez de Pablos. The social media in academia and education research r-evolutions and a paradox: Advanced next generation social learning innovation. J.UCS: Journal of Universal Computer Science, 20(15):1987–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_20;   http:
//www.jucs.org/jucs_
20_15/the_social_media_
in.

Llana:2014:FUF

[LMMPFVI14] L. Llana, E. Martin-Martin, C. Pareja-Flores, and J. Á. Velázquez-Iturbide. FLOP: A userfriendly system for automated program assessment. J.UCS: Journal of Universal Computer Science, 20(9):1304–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_20_9/flop_a_
user_friendly.

Leony:2015:DEE

[LMMRV^{+}15] D. Leony, P. J. Muñoz-Merino, J. A. Ruipérez-Valiente, A. Pardo, and C. Delgado Kloos. Detection and evaluation of emotions in massive open online courses. J.UCS: Journal of Universal Computer Science, 21(5):638–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_5/detection_
and_evaluation_of.

Lauer:2001:ATP

[LMO01] Tobias Lauer, Rainer Müller, and Thomas Ottmann. Animations for teaching purposes: Now and tomorrow. J.UCS: Journal of Universal Computer Science, 7(5): 420–433, May 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
5/animation_for_teaching_
purposes.

Losada:2014:EAL

[LMRG14] N. Losada, M. J. Martín, G. Rodríguez, and P. González. Extending an applicationlevel checkpointing tool to provide fault tolerance support to OpenMP applications. J.UCS: Journal of Universal Computer Science, 20(9):1351– ??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_20_9/extending_
an_application_level.

LeBer:2002:DOB

[LN02] F. Le Ber and A. Napoli.

The design of an objectbased system for representing and classifying spatial structures and relations. J.UCS: Journal of Universal Computer Science, 8(8):751–773, August 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
8/the_design_of_an.

Leporini:2008:FTI

[LN08] B. Leporini and I. Norscia.

‘fine tuning’ image accessibility for museum Web sites. J.UCS: Journal of Universal Computer Science, 14(19):3250–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
19/fine_tuning_image_
accessibility.

Luong:2015:SBC

[LNH^{+}15] N. T. Luong, T. T. Nguyen, D. Hwang, C. H. Lee, and J. J. Jung. Similarity-based complex publication network analytics for recommending potential collaborations. J.UCS: Journal of Universal Computer Science, 21(6):871–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_6/similarity_
based_complex_publication.

Lohmann:2009:FKF

[LNHZ09] S. Lohmann, J. Niesenhaus, P. Heim, and J. Ziegler. Fostering knowledge flow and community engagement in the development of interactive entertainment. J.UCS: Journal of Universal Computer Science, 15(8):1722–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_8/fostering_
knowledge_flow_and.

LeBer:2003:MCF

[LNML03] F. Le Ber, A. Napoli, J.-L.

Metzger, and S. Lardon. Modeling and comparing farm maps using graphs and case-based reasoning. J.UCS: Journal of Universal Computer Science, 9(9):1073–1095, September 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
9/modeling_and_comparing_
farm.

Luther:1998:RCE

[LO98] W. Luther and W. Otten. Reliable computation of elliptic functions. J.UCS: Journal of Universal Computer Science, 4(1):25–33, January 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_1/reliable_
computation_of_elliptic.

Logeswaran:2004:FTS

[Log04] R. Logeswaran. Fast twostage Lempel–Ziv lossless numeric telemetry data compression using a neural network predictor. J.UCS: Journal of Universal Computer Science, 10(9):1199–1211, September 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        10_9/fast_two_stage_
                                        lempel.

Logeswaran:2006:FTN

[Log06] R. Logeswaran. Fault tolerant neural predictors for compression of sensor telemetry data. J.UCS: Journal of Universal Computer Science, 12(10):1439–1454, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        12_10/fault_tolerant_
                                        neural_predictors.

Lomet:2007:DJG

[Lom07] D. Lomet. Dedication to Jim Gray. J.UCS: Journal of Universal Computer Science, 13(8):1044– 1046, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_
                                        8/dedication_to_jim_
                                        gray.

Loos:2006:TVH

[Loo06] R. Loos. Time-varying H systems revisited. J.UCS: Journal of Universal Computer Science, 12 (10):1455–1463, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_10/time_varying_H_
systems.

Leitold:1996:LAI

[LPP96] H. Leitold, R. Posch, and F. Pucher. LAN access over ISDN. J.UCS: Journal of Universal Computer Science, 2(1):20–33, January 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
2_1/lan_access_over_
isdn.

Linaje:2010:MDC

[LPSF10] M. Linaje, J. C. Preciado, and F. Sánchez-Figueroa. Multi-device context-aware RIAs using a model-driven approach. J.UCS: Journal of Universal Computer Science, 16(15):2038–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16_15/multi_
device_context_aware.

Laborda:2016:LTF

[LRB16] J. García Laborda, T. M.

Royo, and M. Bakieva. Looking towards the future of language assessment: Usability of tablet PCs in language testing. J.UCS: Journal of Universal Computer Science, 22(1):114–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
22_1/looking_towards_
the_future.

Leal:2003:LRE

[LRI03] M. A. Leal, N. Rodriguez, and R. Ierusalimschy. LuaTS — a reactive event-driven tuple space. J.UCS: Journal of Universal Computer Science, 9(8):730– 744, August 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
9_8/luats_a_reactive_
event.

LandetaRodriguez:2004:KMA

[LRR04] J. Landeta Rodríguez, A. Rodríguez Castellanos, and S. Y. Ranguelov. Knowledge management analysis of the research & development & transference process at HEROs: a public university case. J.UCS: Journal of Universal Computer Science, 10(6):702–711, June 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/

jucs_10_6/knowledge_
management_analysis_
of.

Loiret:2011:AOF

[LRS^{+}11] F. Loiret, R. Rouvoy, L. Seinturier, D. Romero, K. Sénéchal, and A. Plsek. An aspect-oriented framework for weaving domainspecific concerns into component-based systems. J.UCS: Journal of Universal Computer Science, 17(5):742– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
5/an_aspect_oriented_
framework.

Lynch:1995:HRL

[LS95] Thomas Lynch and Michael J.

Schulte. A high radix online arithmetic for credible and accurate computing. J.UCS: Journal of Universal Computer Science, 1(7):439–453, July 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
a_high_radix_online_
arithmetic.

Lenar:2007:URI

[LS07] M. Lenar and J. Sobecki.

Using recommendation to improve negotiations in agent-based systems. J.UCS: Journal of Universal Computer Science, 13(2):267–286, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_13_2/using_
recommendation_to_improve.

Lopez:2010:MPI

[LS10] J. M. López and M. Sendín. Multi-purpose infrastructure for delivering and supporting mobile contextaware applications. J.UCS: Journal of Universal Computer Science, 16 (15):2081–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_15/multi_purpose_
infrastructure_for.

Lopes:2014:MAD

[LSG^{+}14] J. Lopes, R. Souza, C. Geyer, C. Costa, J. Barbosa, A. Pernas, and A. Yamin. A middleware architecture for dynamic adaptation in ubiquitous computing. J.UCS: Journal of Universal Computer Science, 20(9):1327–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
9/a_middleware_architecture_
for.

Laszlo:2006:CWA

[LSK06] I. László, F. Schipp, and S. P. Kozaitis. Construction of wavelets and applications. J.UCS: Journal of Universal Computer Science, 12(9):1278– 1291, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_9/construction_of_
wavelets_and.

Lanzenberger:2010:OVT

[LSR10] M. Lanzenberger, J. Sampson, and M. Rester. Ontology visualization: Tools and techniques for visual representation of semi-structured metadata. J.UCS: Journal of Universal Computer Science, 16(7):1036–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
7/ontology_visualization_
tools_and.

Lee:2007:GSG

[LST07] A. Lee, I. Streinu, and L. Theran. Graded sparse graphs and matroids. J.UCS: Journal of Universal Computer Science, 13(11):1671–1679, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        13_11/graded_sparse_
                                        graphs_and.

Lino:2014:SBS

[LST14] N. Lino, C. Siebra, and A. Tate. Semantic based support for planning information delivery in human-agent collaborative teams. J.UCS: Journal of Universal Computer Science, 20(13): 1766–??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        20_13/semantic_based_
                                        support_for.

Lukovszki:2006:REM

[LSV06] T. Lukovszki, C. Schindelhauer, and K. Volbert. Resource efficient maintenance of wireless network topologies. J.UCS: Journal of Universal Computer Science, 12(9):1292– 1311, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12_
                                        9/resource_efficient_
                                        maintenance_of.

Lin:2012:CBO

[LSX12] F. Lin, K. Sandkuhl, and S. Xu. Contextbased ontology matching: Concept and application cases. J.UCS: Journal of Universal Computer Science, 18(9):1093– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
9/context_based_ontology_
matching.

Lenz:2009:FFA

[LT09] H.-J. Lenz and B. Thalheim. A formal framework of aggregation for the OLAP-OLTP model. J.UCS: Journal of Universal Computer Science, 15(1):273–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_1/a_formal_framework_
of.

Lamia:2013:DLS

[LT13] M. Lamia and L. M.

Tayeb. Discovering learner styles in adaptive e-learning hypermedia systems. J.UCS: Journal of Universal Computer Science, 19(11):1522–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
11/discovering_learner_
styles_in.

Lappas:2016:SMB

[LTY^{+}16] G. Lappas, A. Triantafillidou, P. Yannas, A. Kavada, A. Kleftodimos, and O. Vasileiadou. Social media battles: their impact during the 2014 Greek municipal elections. J.UCS: Journal of Universal Computer Science, 22(3):375–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_22_3/social_
media_battles_their.

Lukosch:2008:STB

[Luk08]  S. Lukosch.  Seam-

less transition between connected and disconnected collaborative interaction. J.UCS: Journal of Universal Computer Science, 14(1):59– 87, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
1/seamless_transition_
between_connected.

Losada:2013:AUE

[LULGFC13] B. Losada, M. Urretavizcaya, J.-M. López-Gil, and I. Fernández-Castro. Applying usability engineering in InterMod agile development methodology. A case study in a mobile application. J.UCS: Journal of Universal Computer Science, 19(8):1046– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
8/applying_usability_
engineering_in.

Laborda:2016:FTC

[LUR16] J. García Laborda, H. Uzunboylu, and S. Ross. Future trends in computing technology in education. J.UCS: Journal of Universal Computer Science, 22(1):1–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_22;   http:
//www.jucs.org/jucs_
22_1/future_trends_in_
computing.

Lennon:1995:PCP

[LV95] J. Lennon and A. Vermeer. From personal computer to personal assistant. J.UCS: Journal of Universal Computer Science, 1(6):410– 422, June 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/from_
personal_computer_to_
personal_assistant.

Ljubojevic:2013:AUR

[LVS13] M. Ljubojević, V. Vasković, and D. Starcević. The analysis of the users’ response to the linear Internet video advertising by using QoE methods. J.UCS: Journal of Universal Computer Science, 19(12):1736–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
12/the_analysis_of_the.

Li:2010:SCP

[LVV10] L. Li, V. K. Vaishnavi, and A. Vandenberg. SOM clustering to promote interoperability of directory metadata: a gridenabled genetic algorithm approach. J.UCS: Journal of Universal Computer Science, 16(5):800– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
5/som_clustering_to_
promote.

Lu:2008:CRR

[LW08] H. Lu and K. Weihrauch. Computable Riesz representation for locally compact Hausdorff spaces. J.UCS: Journal of Universal Computer Science, 14(6):845–860, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_6/computable_
riesz_representation_
for.

Lu:2004:SCG

[LWC^{+}04] F. Lu, L.-C. Wang, K.- T. Cheng, J. Moondanos, and Z. Hanna. A signal correlation guided circuit-SAT solver. J.UCS: Journal of Universal Computer Science, 10(12): 1629–1654, December 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_10_12/a_signal_
correlation_guided.

Li:2014:EGA

[LWG14] X. Li, X. Wen, and L. Gao. An effective genetic algorithm for multiobjective integrated process planning and scheduling with various flexibilities in process planning. J.UCS: Journal of Universal Computer Science, 20(14):1926–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
14/an_effective_genetic_
algorithm.

Liou:2009:IDM

[LWH09] S.-W. Liou, C.-M. Wang, and Y.-F. Huang. Integrative discovery of multifaceted sequence patterns by frame-relayed search and hybrid PSO-ANN. J.UCS: Journal of Universal Computer Science, 15(4):742– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
4/integrative_discovery_
of_multifaceted.

Li:2010:TOC

[LWL10] L. Li, Y. Wang, and E.- P. Lim. Trust-oriented composite service selection with QoS constraints. J.UCS: Journal of Universal Computer Science, 16(13): 1720–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_13/trust_oriented_
composite_service.

Lee:2011:BIM

[LWS11] J. Lee, J. Whittle, and O. Storz. Bio-inspired mechanisms for coordinating multiple instances of a service feature in dynamic software product lines. J.UCS: Journal of Universal Computer Science, 17(5):670– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
5/bio_inspired_mechanisms_
for.

Luo:2009:RPA

[LWY09] J. Luo, X. Wang, and M. Yang. A resilient P2P anonymous routing approach employing collaboration scheme. J.UCS: Journal of Universal Computer Science, 15(9):1797–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_9/a_resilient_p2p_
anonymous.

Luo:2011:OGD

[LWY11] J. Luo, W. Wu, and M. Yang. Optimization of gateway deployment with load balancing and interference minimization in wireless mesh networks. J.UCS: Journal of Universal Computer Science, 17(14):2064–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
14/optimization_of_gateway_
deployment.

Liu:2015:ICD

[LYLX15] Z. Liu, H. Yan, Z. Lin, and L. Xu. An improved cloud data sharing scheme with hierarchical attribute structure. J.UCS: Journal of Universal Computer Science, 21(3):454– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
3/an_improved_cloud_
data.

Lian:2009:PMT

[LZ09] S. Lian and Y. Zhang. Protecting mobile TV multimedia content in DVB/GPRS heterogeneous wireless networks. J.UCS: Journal of Universal Computer Science, 15(5):1023–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_5/protecting_
mobile_tv_multimedia.

Lian:2009:MSC

[LZGC09] S. Lian, Y. Zhang, S. Gritzalis, and Y. Chen. Multimedia security in communication (MUSIC). J.UCS: Journal of Universal Computer Science, 15(2):398–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_15;   http:
//www.jucs.org/jucs_
15_2#;   http://www.

jucs.org/jucs_15_2/multimedia_
security_in_communication.

Liu:2013:SRT

[LZLW13] B. Liu, B. Zhao, B. Liu, and C. Wu. A security real-time privacy amplification Scheme in QKD system. J.UCS: Journal of Universal Computer Science, 19(16):2420–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
16/a_security_real_time.

Leporati:2004:SFG

[LZM04] A. Leporati, C. Zandron, and G. Mauri. Simulating the Fredkin gate with energy-based P systems. J.UCS: Journal of Universal Computer Science, 10(5):600–619, May 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_10_5/simulating_
the_fredkin_gate.

Lytras:2014:ASR

[LZZK14] M. Lytras, L. Zhuhadar, J. X. Zhang, and E. Kurilovas. Advances of scientific research on technology enhanced learning in social networks and mobile contexts: Towards high effective educational platforms for next generation education. J.UCS: Journal of Universal Computer Science, 20(10):1402–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_20;   http:
//www.jucs.org/jucs_
20_10/advances_of_scientific_
research.

McCahill:1995:EIG

[MA95] M. P. McCahill and F. X. Anklesaria. Evolution of Internet gopher. J.UCS: Journal of Universal Computer Science, 1(4):235–246, April 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
evolution_of_internet_
gopher.

Ma:2010:GEC

[Ma10]  H. Ma.  A geometri-

cally enhanced conceptual model and query language. J.UCS: Journal of Universal Computer Science, 16(20):2986–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
20/a_geometrically_enhanced_
conceptual.

Ma:2012:FAR

[Ma12] J. Ma. A formal approach for risk assessment in RBAC systems. J.UCS: Journal of Universal Computer Science, 18(17):2432–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
17/a_formal_approach_
for.

Mateu:2013:CEM

[MA13] J. Mateu and X. Alaman. CUBICA: An example of mixed reality. J.UCS: Journal of Universal Computer Science, 19(17):2598–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_17/cubica_an_
example_of.

Monteiro:2005:CNF

[MABS05] M. Monteiro, M. Araújo, R. Borges, and A. Santos. Compiling nonstrict functional languages for the .NET platform. J.UCS: Journal of Universal Computer Science, 11(7):1255–1274, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_11_7/compiling_
non_strict_functional.

MacDonaill:1996:SMC

[Mac96] D. A. MacDónaill. On the scalability of molecular computational solutions to NP problems. J.UCS: Journal of Universal Computer Science, 2(2):87–95, February 18, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_2_
                                        2/on_the_scalability_
                                        of.

MacanAirchinnigh:2001:EVI

[Mac01] M. Mac an Airchinnigh. An eclectic view of the Irish School of Constructive Mathematics, from [Lucas 1978] to [Mac an Airchinnigh 2001]. J.UCS: Journal of Universal Computer Science, 7(8):668–691, August 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_7_
                                        8/an_eclectic_view_of.

Maier:2005:MKW

[Mai05]  R. Maier.  Modeling

knowledge work for the design of knowledge infrastructures. J.UCS: Journal of Universal Computer Science, 11(4): 429–451, April 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_4/modeling_knowledge_
work_for.

Manoharan:1997:BPW

[Man97] S. Manoharan. Bounds on the performance of work-greedy assignment schemes. J.UCS: Journal of Universal Computer Science, 3(7):803– 812, July 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs_3_7/bounds_
on_the_performance;
internal&sk=05460486.

Markov:1995:DIA

[Mar95] Svetoslav Markov. On directed interval arithmetic and its applications. J.UCS: Journal of Universal Computer Science, 1(7):514–526, July 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_1_7/on_directed_
interval_arithmetic.

Markowsky:1996:IAI

[Mar96] George Markowsky. Introduction to algorithmic information theory. J.UCS: Journal of Universal Computer Science, 2(5):245–269, May 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/introduction_to_
                                        algorithmic_information_
                                        theory.

Markov:1998:AIC

[Mar98] S. M. Markov. On the algebra of intervals and convex bodies. J.UCS: Journal of Universal Computer Science, 4(1):34–47, January 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_1/on_
                                        the_algebra_of.

Marcus:2000:USB

[Mar00a] S. Marcus. Under the sign of Boole. J.UCS: Journal of Universal Computer Science, 6(1):3–4, January 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_6_
                                        1/under_the_sign_of.

Margenstern:2000:NTC

[Mar00b] M. Margenstern. New tools for cellular automata in the hyperbolic plane. J.UCS: Journal of Universal Computer Science, 6(12):1226, December 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/

jucs_6_12/new_tools_
for_cellular.

Marcus:2002:BTH

[Mar02a]  S. Marcus.  Bridging

two hierarchies of infinite words. J.UCS: Journal of Universal Computer Science, 8(2):292– 296, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
2/bridging_two_hierarchies_
of.

Margenstern:2002:THP

[Mar02b] M. Margenstern. Tiling the hyperbolic plane with a single pentagonal tile. J.UCS: Journal of Universal Computer Science, 8(2):297–316, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_8_2/tiling_
the_hyperbolic_plane.

Margenstern:2004:THS

[Mar04] M. Margenstern. The tiling of the hyperbolic 4D space by the 120-cell is combinatoric. J.UCS: Journal of Universal Computer Science, 10(9):1212–1238, September 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_10_
                                        9/the_tiling_of_the.

Margenstern:2006:AAA

[Mar06a] M. Margenstern. About an algorithmic approach to tilings {p, q} of the hyperbolic plane. J.UCS: Journal of Universal Computer Science, 12(5): 512–550, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        12_5/about_an_algorithmic_
                                        approach.

Martins:2006:BIR

[Mar06b] M. A. Martins. Behavioral institutions and refinements in generalized hidden logics. J.UCS: Journal of Universal Computer Science, 12(8):1020– 1049, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_12_8/behavioral_
                                        institutions_and_refinements.

Marcelino:2007:HHM

[Mar07] M. J. Marcelino. HME: a handheld model editor for educational contexts. J.UCS: Journal of Universal Computer Science, 13(7):1012– 1021, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_7/hme_a_handheld_
model.

Mateescu:1999:JUS

[Mat99]  A. Mateescu.  J.UCS

special issue on formal languages and automata. J.UCS: Journal of Universal Computer Science, 5(9):493, September 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
5_9/formal_languages_
and_automata.

Mattick:2002:DAC

[Mat02]  V. Mattick.  Design

for all as a challenge for hypermedia engineering. J.UCS: Journal of Universal Computer Science, 8(10):881–891, October 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
10/design_for_all_as.

Matousek:2004:TPV

[Mat04] P. Matousek. Tools for parametric verification. A comparison on a case study. J.UCS: Journal of Universal Computer Science, 10(10): 1469–1494, October 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_10_10/tools_for_
                                        parametric_verification.

Molina:2008:OSN

[MAT08] G. Molina, E. Alba, and E.-G. Talbi. Optimal sensor network layout using multi-objective metaheuristics. J.UCS: Journal of Universal Computer Science, 14(15): 2549–??, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_15/optimal_sensor_
                                        network_layout.

Maurer:1994:MEC

[Mau94] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 0(0): 1–2, November 15, 1994. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_0_
                                        0/editorial.

Maurer:1995:MECa

[Mau95a] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 1(1):1, January 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
1/editorial.

Maurer:1995:MECb

[Mau95b] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 1(2): 83, February 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
2/editorial.

Maurer:1995:MECc

[Mau95c] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 1(3): 155, March 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
3/editorial.

Maurer:1995:MECd

[Mau95d] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 1(4): 205, April 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
4/editorial.

Maurer:1995:MECe

[Mau95e] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 1(5): 251, May 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
5/editorial.

Maurer:1995:MECf

[Mau95f] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 1(6): 338, June 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
6/editorial.

Maurer:1995:MECg

[Mau95g] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 1(7): 435, July 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
7/editorial.

Maurer:1995:MECh

[Mau95h] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 1(8): 570, August 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
8/editorial.

Maurer:1995:MECi

[Mau95i] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 1(9): 632, September 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
9/editorial.

Maurer:1995:MECj

[Mau95j] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 1(10): 674, October 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
10/editorial.

Maurer:1995:MECk

[Mau95k] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 1(11): 718, November 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
11/editorial.

Maurer:1995:MECl

[Mau95l] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 1(12): 751, December 28, 1995.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
12/editorial.

Maurer:1996:LUC

[Mau96a] H. Maurer. LATE: a unified concept for a learning and teaching environment. J.UCS: Journal of Universal Computer Science, 2 (8):580–595, August 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_2_8/late_a_
unified_concept.

Maurer:1996:MECa

[Mau96b] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 2(1):1, January 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
1/editorial.

Maurer:1996:MECb

[Mau96c] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 2(2): 57–58, March 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
2/editorial.

Maurer:1996:MECc

[Mau96d] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 2(3): 96, March 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
3/editorial.

Maurer:1996:MECd

[Mau96e] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 2(4): 163, April 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
4/editorial.

Maurer:1996:MECe

[Mau96f] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 2(5): 241, May 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
5/editorial.

Maurer:1996:MECf

[Mau96g] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 2(6): 442, June 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_2_
                                        6/editorial.

Maurer:1996:MECg

[Mau96h] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 2(7): 513, July 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_2_
                                        7/editorial.

Maurer:1996:MECh

[Mau96i] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 2(8): 569, August 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_2_
                                        8/editorial.

Maurer:1996:MECi

[Mau96j] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 2(9): 596, September 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_2_
                                        9/editorial.

Maurer:1996:MECj

[Mau96k] H. Maurer. Managing Editor’s column. J.UCS:

Journal of Universal Computer Science, 2(10): 659, October 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
10/editorial.

Maurer:1996:MECk

[Mau96l] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 2(11): 744, November 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
11/editorial.

Maurer:1996:MECl

[Mau96m] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 2(12): 796, December 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
12/editorial.

Maurer:1997:MECa

[Mau97a] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 3(1):1, January 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_3_
1/editorial.

Maurer:1997:MECb

[Mau97b] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 3(2): 69, February 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_3_
                                        2/editorial.

Maurer:1997:MECc

[Mau97c] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 3(3): 147, March 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_3_
                                        3/editorial.

Maurer:1997:MECd

[Mau97d] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 3(6): 702, June 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs_3_6/editorial;
                                         http://medoc.springer.
                                        de:8000/jucs_volume_
                                        3; internal&sk=05460486;
                                         internal&sk=05460486#
                                        toc_3_6.

Maurer:1997:MECe

[Mau97e] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 3(7): 754–755, July 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs_3_7/editorial;
 http://medoc.springer.
de:8000/jucs_volume_
3; internal&sk=05460486;
 internal&sk=05460486#
toc_3_7.

Maurer:1997:MECf

[Mau97f] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 3(9): 1022, September 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs;   http://
medoc.springer.de:8000/
jucs/jucs_3_9/editorial;
 internal&sk=0156049D/
jucs_3_9.

Maurer:1997:MECg

[Mau97g] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 3(10): 1084, October 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs;   http://
medoc.springer.de:8000/
jucs/jucs_3_10/editorial;

                                         internal&sk=0156049D/
                                        jucs_3_10.

Maurer:1997:MECh

[Mau97h] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 3(12): 1282, December 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs;   http://
                                        medoc.springer.de:8000/
                                        jucs/jucs_3_12/editorial;
                                         internal&sk=0156049D/
                                        jucs_3_12.

Maurer:1997:TAI

[Mau97i] H. Maurer. On two aspects of improving Webbased training. J.UCS: Journal of Universal Computer Science, 3 (10):1126–1133, October 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_3_10/
                                        on_two_aspects_of.

Maurer:1998:MECa

[Mau98a] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 4(5): 484–485, May 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_5/editorial;
http://medoc.springer.
de:8000/jucs/jucs_volume_
4; internal&sk=0C220489;
internal&sk=0C220489#
toc_4_5.

Maurer:1998:MECb

[Mau98b] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 4(6): 560, June 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_6/editorial;
http://medoc.springer.
de:8000/jucs/jucs_volume_
4; internal&sk=0C220489;
internal&sk=0C220489#
toc_4_6.

Maurer:1998:MECc

[Mau98c] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 4(7): 629, July 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_7/editorial;
http://medoc.springer.
de:8000/jucs/jucs_volume_
4; internal&sk=0C220489;
internal&sk=0C220489#
toc_4_7.

Maurer:1998:MECd

[Mau98d] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 4(8): 669, August 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_8/editorial;
 http://medoc.springer.
de:8000/jucs/jucs_volume_
4#toc_4_8.

Maurer:1998:MECe

[Mau98e] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 4(9): 718, September 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_9/editorial;
 http://medoc.springer.
de:8000/jucs/jucs_volume_
4#toc_4_9.

Maurer:1998:MECf

[Mau98f] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 4(10): 779, October 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_10/
editorial;   http://
medoc.springer.de:8000/
jucs/jucs_volume_4#toc_
4_10.

Maurer:1998:MECg

[Mau98g] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 4(11): 824, November 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_11/
editorial;   http://
medoc.springer.de:8000/
jucs/jucs_volume_4; internal&
sk=01237079;   internal&
sk=01237079#toc_4_11.

Maurer:1998:MECh

[Mau98h] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 4(12): 665–865, December 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_4_12/editorial.

Maurer:1999:MECa

[Mau99a] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 5(1):1, January 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
1/editorial.

Maurer:1999:MECd

[Mau99b] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 5(2): 32, February 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
2/editorial.

Maurer:1999:MECe

[Mau99c] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 5(4): 226, April 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
4/editorial.

Maurer:1999:MECf

[Mau99d] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 5(5): 287, May 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
5/editorial.

Maurer:1999:MECg

[Mau99e] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 5(6): 322, June 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
6/editorial.

Maurer:1999:MECh

[Mau99f] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 5(7): 366, July 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
7/editorial.

Maurer:1999:MECi

[Mau99g] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 5(8): 463, August 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
8/editorial.

Maurer:1999:MECb

[Mau99h] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 5(11): 742, November 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
11/editorial.

Maurer:1999:MECc

[Mau99i] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 5(12): 816, December 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
12/editorial.

Maurer:2000:MECa

[Mau00a]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 6(2):226, February 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/#toc_6_2;   http:
//www.jucs.org/jucs_
6_2/editorial.

Maurer:2000:MECb

[Mau00b] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 6(5): 489, May 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
5/editorial.

Maurer:2000:MECc

[Mau00c] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 6(6): 537, June 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
6/editorial.

Maurer:2000:MECd

[Mau00d] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 6(8): 758, August 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
8/editorial.

Maurer:2000:MECe

[Mau00e] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 6(9): 849, September 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
9/editorial.

Maurer:2000:MECf

[Mau00f] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 6(12): 1164, December 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
12/editorial.

Maurer:2001:MECa

[Mau01a] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 7(3): 210, March 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
3/editorial.

Maurer:2001:MECb

[Mau01b] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 7(4): 290, April 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
4/editorial.

Maurer:2001:MECc

[Mau01c] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 7(9): 782, September 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
9/editorial.

Maurer:2001:MECd

[Mau01d] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 7(10): 869, October 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
10/editorial.

Maurer:2001:MECe

[Mau01e] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 7(12): 1113, December 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
12/editorial.

Maurer:2002:MECa

[Mau02a] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 8(1):1, January 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
1/editorial.

Maurer:2002:MECb

[Mau02b] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 8(3): 368, March 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
http://www.jucs.org/
jucs_8_3/editorial.

Maurer:2002:MECc

[Mau02c] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 8(4):

408, April 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
 http://www.jucs.org/
jucs_8_4/editorial.

Maurer:2002:MECd

[Mau02d] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 8(7): 673, July 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
 http://www.jucs.org/
jucs_8_7/editorial.

Maurer:2002:MECe

[Mau02e] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 8(9): 833, September 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
 http://www.jucs.org/
jucs_8_9/editorial.

Maurer:2002:MECf

[Mau02f] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 8(11): 991, November 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
 http://www.jucs.org/
jucs_8_11/editorial.

Maurer:2002:MECg

[Mau02g] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 8(12): 1039, December 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
 http://www.jucs.org/
jucs_8_12/editorial.

Maurer:2003:JUS

[Mau03a] H. Maurer. J. UCS special issue on dynamic symbolic languages. J.UCS: Journal of Universal Computer Science, 9(4):298– 299, April 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
 http://www.jucs.org/
jucs_9_4/j_ucs_special_
issue.

Maurer:2003:MECa

[Mau03b] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 9(1):1, January 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
 http://www.jucs.org/
jucs_9_1/editorial.

Maurer:2003:MECb

[Mau03c] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 9(5): 397, May 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/free;
                                         http://www.jucs.org/
                                        jucs_9_5/editorial.

Maurer:2003:MECc

[Mau03d] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 9(10): 1195, October 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/free;
                                         http://www.jucs.org/
                                        jucs_9_10/editorial.

Maurer:2004:MECa

[Mau04a] H. Maurer. Managing Editor’s column. J.UCS: Journal of Universal Computer Science, 10(2): 105, February 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/free;
                                         http://www.jucs.org/
                                        jucs_10_2/editorial.

Maurer:2004:MECb

[Mau04b] H. Maurer. Managing Editor’s column. J.UCS:

Journal of Universal Computer Science, 10(9): 1034, September 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
http://www.jucs.org/
jucs_10_9/editorial.

Maurer:2005:MEC

[Mau05]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 11(6):803, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_11;   http:
//www.jucs.org/jucs_
11_6#; http://www.jucs.
org/jucs_11_6/editorial.

Maurer:2006:MECc

[Mau06a]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 12(2):126, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_12;   http:
//www.jucs.org/jucs_
12_2#; http://www.jucs.
org/jucs_12_2/editorial.

Maurer:2006:MECd

[Mau06b]  H. Maurer.  Man-

aging Editor’s column.

J.UCS: Journal of Universal Computer Science, 12(5):481, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_12;   http:
//www.jucs.org/jucs_
12_5#; http://www.jucs.
org/jucs_12_5/editorial.

Maurer:2006:MECe

[Mau06c]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 12(8):980, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_12;   http:
//www.jucs.org/jucs_
12_8#; http://www.jucs.
org/jucs_12_8/editorial.

Maurer:2006:MECa

[Mau06d]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 12(10):1411–1412, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12;
 http://www.jucs.org/
jucs_12_10#;   http:
//www.jucs.org/jucs_
12_10/editorial.

Maurer:2006:MECb

[Mau06e]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 12(12):1700, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_12;   http:
//www.jucs.org/jucs_
12_12#;   http://www.
jucs.org/jucs_12_12/
editorial.

Maurer:2007:MECb

[Mau07a]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 13(4):454, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_13;   http:
//www.jucs.org/jucs_
13_4#; http://www.jucs.
org/jucs_13_4/editorial.

Maurer:2007:MECc

[Mau07b]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 13(5):592, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_13;   http:
//www.jucs.org/jucs_

                                        13_5#; http://www.jucs.
                                        org/jucs_13_5/editorial.

Maurer:2007:MECd

[Mau07c]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 13(9):1136–1137, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_13;   http:
                                        //www.jucs.org/jucs_
                                        13_9#; http://www.jucs.
                                        org/jucs_13_9/editorial.

Maurer:2007:MECa

[Mau07d]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 13(10):1366, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_13;   http:
                                        //www.jucs.org/jucs_
                                        13_10#;   http://www.
                                        jucs.org/jucs_13_10/
                                        editorial.

Maurer:2008:MECa

[Mau08a]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 14(4):505–507, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_14;   http:
//www.jucs.org/jucs_
14_4#; http://www.jucs.
org/jucs_14_4/editorial.

Maurer:2008:MECb

[Mau08b]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 14(10):1560–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_14;   http:
//www.jucs.org/jucs_
14_10#;   http://www.
jucs.org/jucs_14_10/
editorial.

Maurer:2008:MECc

[Mau08c]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 14(22):3625–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_14;   http:
//www.jucs.org/jucs_
14_22#;   http://www.
jucs.org/jucs_14_22/
editorial.

Maurer:2009:MECa

[Mau09a]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 15(3):505–??, ????

2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_15;   http:
                                        //www.jucs.org/jucs_
                                        15_3#; http://www.jucs.
                                        org/jucs_15_3/editorial.

Maurer:2009:MECb

[Mau09b]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 15(8):1562–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_15;   http:
                                        //www.jucs.org/jucs_
                                        15_8#; http://www.jucs.
                                        org/jucs_15_8/editorial.

Maurer:2009:MECc

[Mau09c]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 15(14):2746–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_15;   http:
                                        //www.jucs.org/jucs_
                                        15_14#;   http://www.
                                        jucs.org/jucs_15_14/
                                        editorial.

Maurer:2009:MECd

[Mau09d]  H. Maurer.  Man-

aging Editor’s column.

J.UCS: Journal of Universal Computer Science, 15(17):3158–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_15;   http:
//www.jucs.org/jucs_
15_17#;   http://www.
jucs.org/jucs_15_17/
editorial.

Maurer:2010:MECa

[Mau10a]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 16(3):340–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16;   http:
//www.jucs.org/jucs_
16_3#; http://www.jucs.
org/jucs_16_3/editorial.

Maurer:2010:MECb

[Mau10b]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 16(9):1138–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16;   http:
//www.jucs.org/jucs_
16_9#; http://www.jucs.
org/jucs_16_9/editorial.

Maurer:2010:MECc

[Mau10c]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 16(14):1825–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_16;   http:
                                        //www.jucs.org/jucs_
                                        16_14/editorial;   http:
                                        //www.jucs.org/jucsrssfeed-
                                        issue.

Maurer:2010:MECd

[Mau10d]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 16(21):3163–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_16;   http:
                                        //www.jucs.org/jucs_
                                        16_21/editorial;   http:
                                        //www.jucs.org/jucsrssfeed-
                                        issue.

Maurer:2011:MECa

[Mau11a]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 17(3):348–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_17;   http:

//www.jucs.org/jucs_
17_3/editorial;   http:
//www.jucs.org/jucsrssfeed-
issue.

Maurer:2011:MECb

[Mau11b]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 17(6):830–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_17;   http:
//www.jucs.org/jucs_
17_6/editorial;   http:
//www.jucs.org/jucsrssfeed-
issue.

Maurer:2011:MECc

[Mau11c]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 17(11):1527–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_17;   http:
//www.jucs.org/jucs_
17_11/editorial;   http:
//www.jucs.org/jucsrssfeed-
issue.

Maurer:2011:MECd

[Mau11d]  H. Maurer.  Man-

aging Editor’s column. J.UCS: Journal of Universal Computer Science, 17(13):1764–??, ???? 2011. CODEN ????

ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_17;   http:
                                        //www.jucs.org/jucs_
                                        17_13/editorial;   http:
                                        //www.jucs.org/jucsrssfeed-
                                        issue.

Maybury:2002:KDK

[May02] M. T. Maybury. Knowledge on demand: Knowledge and expert discovery. J.UCS: Journal of Universal Computer Science, 8(5):491– 505, May 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        5/knowledge_on_demand_
                                        knowledge.

Maicher:2009:AFD

[MB09] L. Maicher and B. Bock. ActiveTM — the factory for domain-customised portal engines. J.UCS: Journal of Universal Computer Science, 15(8): 1711–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_15_8/active_tm_
                                        the_factory.

Mahmood:2012:GTF

[MBA12] N. Mahmood, S. M. A. Burney, and K. Ahsan. Generic temporal and fuzzy ontological framework (GTFOF) for developing temporalfuzzy database model for managing Patient’s data. J.UCS: Journal of Universal Computer Science, 18(2):177– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
2/generic_temporal_and_
fuzzy.

Martinovic:2012:PER

[MBC12] G. Martinovic, J. Balen, and B. Cukic. Performance evaluation of recent Windows operating systems. J.UCS: Journal of Universal Computer Science, 18(2):218– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
2/performance_evaluation_
of_recent.

Marco:2013:NDT

[MBC13] J. Marco, S. Baldassarri, and E. Cerezo. NIKVision: Developing a tangible application for and with children. J.UCS: Journal of Universal Computer Science, 19(15):2266–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_19_
                                        15/nik_vision_developing_
                                        a.

MacDonald:2000:SEZ

[MC00] A. MacDonald and D. Carrington. Some elements of Z specification style: Structuring techniques. J.UCS: Journal of Universal Computer Science, 6(12):1203–1225, December 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_6_
                                        12/some_elements_of_z.

Morales:2007:MAC

[MC07] R. Morales and P. Carmichael. Mapping academic collaboration networks: Perspectives from the first year of the reusable learning objects CETL. J.UCS: Journal of Universal Computer Science, 13(7):1033–1041, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_
                                        7/mapping_academic_collaboration_
                                        networks.

Min:2013:DSB

[MCC13] H.-S. Min, S.-M. Chung, and J.-Y. Choi. Deriving system behavior from UML state machine diagram: Applied to missile project. J.UCS: Journal of Universal Computer Science, 19(1):53– ??, ???? 2013. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_1/deriving_system_
behavior_from.

Mezni:2014:EPL

[MCG14] H. Mezni, W. Chainbi, and K. Ghedira. Extending policy languages for expressing the selfadaptation of Web services. J.UCS: Journal of Universal Computer Science, 20(8):1130–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_20_8/extending_
policy_languages_for.

Maidl:2007:UVP

[MCM07] A. Murbach Maidl, C. Carvilhe, and M. A. Musicante. Using visitor patterns in object-oriented action semantics. J.UCS: Journal of Universal Computer Science, 13(6):891–919, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_13_6/using_
visitor_patterns_in.

Munoz-Cristobal:2014:CAE

[MCMMAP^{+}14] J. A. Muñoz-Cristóbal, A. Martínez-Monés, J. I. Asensio-Pérez, S. L. Villagrá-Sobrino, J. E. Hoyos-Torio, and Y. Dimitriadis. City ads: Embedding virtual worlds and augmented reality in everyday educational practice. J.UCS: Journal of Universal Computer Science, 20(12):1670–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_20_12/city_ads_
                                        embedding_virtual.

Mateo:2014:FSR

[MdCRMP14] J. A. Mateo, M. del Carmen Ruiz, H. Maciá, and J. J. Pardo. Formal study of routing protocols for wireless sensor networks. J.UCS: Journal of Universal Computer Science, 20(9):1373–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        20_9/formal_study_of_
                                        routing.

Maciel:2009:SAG

[MDO^{+}09] R. S. P. Maciel, J. M. N. David, M. R. Oei, A. A. de Oliveira Bastos, and L. de Oliveira Menezes. Supporting awareness in groupware through an aspect-oriented middleware service. J.UCS: Journal of Universal Computer Science, 15(9): 1945–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_9/supporting_
awareness_in_groupware.

Mingyi:2010:AGD

[MDY10] Z. Mingyi, L. Danning, and Z. Ying. An approach to generation of decision rules. J.UCS: Journal of Universal Computer Science, 16(1):140– ??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_1/an_approach_to_
generation.

Museros:2003:MMI

[ME03] L. Museros and M. T.

Escrig. Modeling motion by the integration of topology and time. J.UCS: Journal of Universal Computer Science, 9(9):1096–1122, September 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
9/modeling_motion_by_
the.

Mearelli:1997:RAS

[Mea97]  L. Mearelli.  Refining

an ASM specification of the production cell to C++ code. J.UCS: Journal of Universal Computer Science, 3(5):666– 688, May 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs_3_5/production_
                                        cell;   http://www.jucs.
                                        org/jucs_3_5/production_
                                        cell; internal&sk=05460486.

Mehler:2002:CMC

[Meh02]  A. Mehler.  Compo-

nents of a model of context-sensitive hypertexts. J.UCS: Journal of Universal Computer Science, 8(10):924–943, October 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        8_10/components_of_a_
                                        model.

Messine:2002:ESA

[Mes02]  F. Messine.  Exten-

tions [sic] of affine arithmetic: Application to unconstrained global optimization. J.UCS: Journal of Universal Computer Science, 8(11):992– 1015, November 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_8_11/extentions_
of_affine_arithmetic.

Mashat:2013:IDE

[MFG13] A. S. Mashat, H. M. Fardoun, and J. A. Gallud. Interaction design in educational environments. J.UCS: Journal of Universal Computer Science, 19(7):851–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_19;   http:
//www.jucs.org/jucs_
19_7/interaction_design_
in_educational.

Matsumoto:2014:AMD

[MG14] P. Matsumoto and E. Guerra.

An approach for mapping domain-specific AOM applications to a general model. J.UCS: Journal of Universal Computer Science, 20(4):534– ??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
20_4/an_approach_for_
mapping.

Museros:2010:PQA

[MGAVF10] L. Museros, L. González-Abril, F. Velasco, and Z. Falomir. A pragmatic qualitative approach for juxtaposing shapes. J.UCS: Journal of Universal Computer Science, 16(11):1410–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
11/a_pragmatic_qualitative_
approach.

Molina-Gil:2012:CPM

[MGCGCG12] J. Molina-Gil, P. Caballero-Gil, and C. Caballero-Gil. Countermeasures to prevent misbehaviour in VANETs. J.UCS: Journal of Universal Computer Science, 18(6):857– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
6/countermeasures_to_
prevent_misbehaviour.

Martinez:2008:MIC

[MGM^{+}08] D. Martínez, A. S. García, J. Martínez, J. P. Molina, and P. Gonzalez. A model of interaction for CVEs based on the model of human communication. J.UCS: Journal of Universal Computer Science, 14(19):3071–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
19/a_model_of_interaction.

Monica:2012:CUB

[MGMS12] D. Della Monica, V. Goranko, A. Montanari, and G. Sciavicco. Crossing the undecidability border with extensions of propositional neighborhood logic over natural numbers. J.UCS: Journal of Universal Computer Science, 18(20):2798–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_20/crossing_
the_undecidability_border.

Martinez-Gil:2012:MOM

[MGNDAM12] J. Martinez-Gil, I. Navas-Delgado, and J. F. Aldana-Montes. MaF: An ontology matching framework. J.UCS: Journal of Universal Computer Science, 18(2):194– ??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_2/maf_an_ontology_
matching.

Morales:2007:IQT

[MGPB07] E. M. Morales, F. J.

García-Peñalvo, and Á. Barrón. Improving LO quality through instructional design based on an ontological model and metadata. J.UCS: Journal of Universal Computer Science, 13(7):970–979, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_13_7/improving_
lo_quality_through.

Moya:2014:FPL

[MGT14] S. Moya, S. Grau, and D. Tost. First-person locomotion in 3D virtual environments: a usability analysis. J.UCS: Journal of Universal Computer Science, 20(7):1026–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_20_7/first_
person_locomotion_in.

Marshall:1996:DHB

[MH96] A. D. Marshall and S. Hurley. Delivering hypertext based courseware on the WWW. J.UCS: Journal of Universal Computer Science, 2(12):805–828, December 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
delivering_hypertext_
based_courseware_on_
the_www.

McGee:1998:EEM

[MH98] S. McGee and B. Howard.

Evaluating educational multimedia in the context of use. J.UCS: Journal of Universal Computer Science, 4(3):273– 291, March 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_3/evaluating_
educational_multimedia_
in; internal&sk=0C220489.

Milchrahm:2002:KTR

[MH02] E. Milchrahm and A. Hasler.

Knowledge transfer in recycling networks: Fostering sustainable development. J.UCS: Journal of Universal Computer Science, 8(5):546– 556, May 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
5/knowledge_transfer_
in_recycling.

Mertins:2003:POK

[MHA03] K. Mertins, P. Heisig, and K. Alwert. Processoriented knowledge structuring. J.UCS: Journal of Universal Computer Science, 9(6):542– 550, June 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        9_6/process_oriented_
                                        knowledge_structuring.

Masud:2015:SDT

[MHA^{+}15] M. Masud, M. S. Hossain, A. Alamri, A. Almogren, and M. Zakariah. Synchronizing data through update queries in interoperable e-health and technology enhanced learning data sharing systems. J.UCS: Journal of Universal Computer Science, 21(11):1439–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        21_11/synchronizing_
                                        data_through_update.

Mendes:1998:AME

[MHH98] E. Mendes, W. Hall, and R. Harrison. Applying metrics to the evaluation of educational hypermedia applications. J.UCS: Journal of Universal Computer Science, 4(4): 382–403, April 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_4/applying_
                                        metrics_to_the;   internal&
                                        sk=0C220489.

Melero:2012:RCL

[MHLB12] J. Melero, D. Hernández-Leo, and J. Blat. A review of constructivist learning methods with supporting tooling in ICT higher education: Defining different types of scaffolding. J.UCS: Journal of Universal Computer Science, 18(16):2334–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_16/a_review_
of_constructivist.

Mascarenhas:2004:LSN

[MI04] F. Mascarenhas and R. Ierusalimschy. LuaInterface: Scripting the .NET CLR with Lua. J.UCS: Journal of Universal Computer Science, 10(7):892–909, July 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
7/luainterface_scripting_
the_net.

Mascarenhas:2005:RLS

[MI05] F. Mascarenhas and R. Ierusalimschy. Running Lua scripts on the CLR through bytecode translation. J.UCS: Journal of Universal Computer Science, 11(7):1275–1290, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_

                                        7/running_lua_scripts_
                                        on.

Muhammad:2007:CAE

[MI07] H. Muhammad and R. Ierusalimschy. C APIs in extension and extensible languages. J.UCS: Journal of Universal Computer Science, 13(6):839– 853, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_
                                        6/c_apis_in_extension.

Mahrooghi:2013:ATE

[MJ13] H. R. Mahrooghi and R. Jalili. An algebraic theory of epistemic processes. J.UCS: Journal of Universal Computer Science, 19(10):1396–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_19_
                                        10/an_algebraic_theory_
                                        of.

Manzato:2015:LHR

[MJG15] M. G. Manzato, E. B. Santos Junior, and R. Goularte. Leveraging hybrid recommenders with multifaceted implicit feedback. J.UCS: Journal of Universal Computer Science, 21(2):223–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_2/leveraging_
hybrid_recommenders_
with.

Martinez-Julia:2012:NIB

[MJGS12] P. Martinez-Julia and A. F. Gómez-Skarmeta. A novel identity-based network architecture for next generation Internet. J.UCS: Journal of Universal Computer Science, 18(12):1643–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
12/a_novel_identity_
based.

Marin:2013:MSS

[MJS13] L. Marin, A. J. Jara, and A. F. Skarmeta. Multiplication and squaring with shifting primes on OpenRISC processors with hardware multiplier. J.UCS: Journal of Universal Computer Science, 19(16):2368–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_16/multiplication_
and_squaring_with.

Mason:1999:DTA

[MK99] K. Mason and P. Krishnan. Decomposition of timed automata.

J.UCS: Journal of Universal Computer Science, 5(9):574–587, September 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        5_9/decomposition_of_
                                        timed_automata.

Mukun:2012:BAA

[MK12] C. Mukun and M. Y. Kiang. BDI agent architecture for multi-strategy selection in automated negotiation. J.UCS: Journal of Universal Computer Science, 18(10): 1379–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_18_10/BDI_agent_
                                        architecture_for.

Mousannif:2011:CSV

[MKA11] H. Mousannif, I. Khalil, and H. Al Moatassime. Cooperation as a service in VANETs. J.UCS: Journal of Universal Computer Science, 17(8):1202– ??, ???? 2011. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        17_8/cooperation_as_
                                        a_service.

Mouratidis:2012:ASP

[MKI^{+}12] H. Mouratidis, C. Kalloniatis, S. Islam, M.-P. Huget, and S. Gritzalis. Aligning security and privacy to support the development of secure information systems. J.UCS: Journal of Universal Computer Science, 18(12):1608–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
12/aligning_security_
and_privacy.

Mraz:1998:PSS

[MKP98] F. Mráz, M. Kursch, and D. Panuska. Pascal subroutines for solving some problems in interval LP. J.UCS: Journal of Universal Computer Science, 4(2):164– 170, February 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_2/pascal_
subroutines_for_solving.

Memmel:2009:PMS

[MKS09] M. Memmel, M. Kockler, and R. Schirru. Providing multi source tag recommendations in a social resource sharing platform. J.UCS: Journal of Universal Computer Science, 15(3):678–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        15_3/providing_multi_
                                        source_tag.

Maurer:2006:PS

[MKZ06] H. Maurer, F. Kappe, and B. Zaka. Plagiarism — a survey. J.UCS: Journal of Universal Computer Science, 12(8):1050– 1084, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12_
                                        8/plagiarism_a_survey.

Maurer:1995:DLL

[ML95] H. Maurer and J. Lennon. Digital libraries as learning and teaching support. J.UCS: Journal of Universal Computer Science, 1(11):719– 727, November 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        digital_libraries_as_
                                        learning_and_teaching_
                                        support.

Messine:1998:EMM

[ML98] F. Messine and J. L. Lagouanelle. Enclosure methods for multivariate differentiable functions and application to global optimization. J.UCS: Journal of Universal Computer Science, 4(6): 589–603, June 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_6/enclosure_
methods_for_multivariate;
internal&sk=0C220489.

Mitra:2002:SRA

[ML02] D. Mitra and G. Ligozat.

Spatial-reasoning for agents in multiple dimensions. J.UCS: Journal of Universal Computer Science, 8(8):774–792, August 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_8_8/spatial_
reasoning_for_agents.

Musicante:2005:BSP

[ML05] M. A. Musicante and R. M. F. Lima. The 9th Brazilian Symposium on Programming Languages. J.UCS: Journal of Universal Computer Science, 11(7):1115–1116, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_11;   http:
//www.jucs.org/jucs_
11_7#; http://www.jucs.
org/jucs_11_7/sblp_2005.

Marta-Lazo:2016:AIM

[MLHCGB16] C. Marta-Lazo, E. Hergueta-Covacho, and J. A.

Gabelas-Barroso. Applying inter-methodological concepts for enhancing media literacy competences. J.UCS: Journal of Universal Computer Science, 22(1):37– ??, ???? 2016. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
22_1/applying_inter_
methodological_concepts.

Martinez:2010:IFL

[MLX10] L. Martínez, J. Liu, and Y. Xu. Information fusion and logic-based reasoning approaches for decision making under uncertainty. J.UCS: Journal of Universal Computer Science, 16(1):1–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16;   http:
//www.jucs.org/jucs_
16_01/approaches_for_
decision_making;   http:
//www.jucs.org/jucs_
16_1#.

Mitu:1996:TLI

[MM96] B. Mîtu and C. Mîtu. Toy LISP interpreter on a Connex Memory Machine. J.UCS: Journal of Universal Computer Science, 2(5):427– 438, May 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
5/toy_lisp_interpreter_
on.

Marden:1998:PAA

[MM98] P. M. Marden, Jr. and E. V. Munson. PSL: An alternate approach to style sheet languages for the World Wide Web. J.UCS: Journal of Universal Computer Science, 4(10):792–806, October 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_10/
psl_an_alternate_approach.

Margenstern:1999:PSS

[MM99] M. Margenstern and K. Morita. A polynomial solution for 3-SAT in the space of cellular automata in the hyperbolic plane. J.UCS: Journal of Universal Computer Science, 5(9):563– 573, September 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
9/a_polynomial_solution_
for.

Malik:2003:CSV

[MM03] R. Malik and R. Mühlfeld.

A case study in verification of UML statecharts: the PROFIsafe protocol. J.UCS: Journal of Universal Computer Science, 9(2):138– 151, February 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        2/a_case_study_in.

Machado:2006:DAC

[MM06] J. Pereira Machado and P. Blauth Menezes. Defining atomic composition in UML behavioral diagrams. J.UCS: Journal of Universal Computer Science, 12(7):958– 979, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        12_7/defining_atomic_
                                        composition_in.

Mery:2007:SRA

[MM07] D. Méry and S. Merz. Specification and refinement of access control. J.UCS: Journal of Universal Computer Science, 13(8):1073–1093, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_
                                        8/specification_and_
                                        refinement_of.

Mishra:2012:GSI

[MM12] D. Mishra and A. Mishra.

A global software inspection process for distributed software development. J.UCS: Journal of Universal Computer Science, 18(19):2731–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
19/a_global_software_
inspection.

Molnar:2015:ALA

[MM15] A. Molnar and C. Hava Muntean. Assessing learning achievements when reducing mobile video quality. J.UCS: Journal of Universal Computer Science, 21(7):959– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
7/assessing_learning_
achievements_when.

Muller:2008:AWB

[MMB08] C. Müller, B. Meuthrath, and A. Baumgraß. Analyzing Wiki-based networks to improve knowledge processes in organizations. J.UCS: Journal of Universal Computer Science, 14(4):526– 545, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_4/analyzing_wiki_
                                        based_networks.

Milovanovic:2012:IFM

[MMD12] M. Milovanović, M. Minović, and S. Dusan. Interoperability framework for multimodal biometry: Open source in action. J.UCS: Journal of Universal Computer Science, 18(11):1558–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_18_
                                        11/interoperability_
                                        framework_for_multimodal.

Machado:2006:TSV

[MMdMGM06] R. Machado, Á. Freitas Moreira, R. de Matos Galante, and M. Moura Moro. Type-safe versioned object query language. J.UCS: Journal of Universal Computer Science, 12(7): 938–957, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_12_7/type_safe_
                                        versioned_object.

Martinez-Martin:2012:GQS

[MMEdP12] E. Martínez-Martín, M. T. Escrig, and A. P. del Pobil. A general qualitative spatio-temporal model based on intervals. J.UCS: Journal of Universal Computer Science, 18(10):1343–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_10/a_general_
qualitative_spatio.

Moriyama:2008:RLF

[MMiF^{+}08] K. Moriyama, M. Matsumoto, K. i. Fukui, S. Kurihara, and M. Numao. Reinforcement learning on a futures market simulator. J.UCS: Journal of Universal Computer Science, 14(7): 1136–1153, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_7/reinforcement_learning_
on_a.

Matulevicius:2012:SSE

[MMM^{+}12a] R. Matulevicius, H. Mouratidis, N. Mayer, E. Dubois, and P. Heymans. Syntactic and semantic extensions to secure tropos to support security risk management. J.UCS: Journal of Universal Computer Science, 18(6):816– ??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_18_6/syntactic_
                                        and_semantic_extensions.

Mishra:2012:DDI

[MMM12b] A. Mishra, J. Münch, and D. Mishra. Distributed development of information system. J.UCS: Journal of Universal Computer Science, 18(19): 2599–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        18;   http://www.jucs.
                                        org/jucs_18_19/distributed_
                                        development_of_information.

Mendoza:2015:NPR

[MMP15] A. Mendoza, S. Mata, and L. Pastor. Nonphotorealistic rendering of neural cells from their morphological description. J.UCS: Journal of Universal Computer Science, 21(7):935– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_21_
                                        7/non_photorealistic_
                                        rendering_of.

Martin:2008:DMO

[MMS08] B. Martin, A. Mitrovic, and P. Suraweera. ITS domain modelling with ontology. J.UCS: Journal of Universal Computer Science, 14(17):2758–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
17/its_domain_modelling_
with.

Minovic:2011:MKG

[MMS11] M. Minović, M. Milovanović, and D. Starčević. Modelling knowledge and game based learning: Model driven approach. J.UCS: Journal of Universal Computer Science, 17(9):1241–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_9/modelling_
knowledge_and_game.

Muldner:1996:CSH

[MN96] T. Müldner and R. Nicholl.

Computer-supported human cooperation in electronic classrooms. J.UCS: Journal of Universal Computer Science, 2(10): 679–693, October 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
computer_supported_human_
cooperation.

Metzner:2000:MMR

[MN00] A. Metzner and J. Niehaus.

MSparc: Multithreading in real-time architectures. J.UCS: Journal of Universal Computer Science, 6(10):1034–1051, October 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_6_10/msparc_
multithreading_in_real.

Maly:2014:EOE

[MN14] J. Malý and M. Necaský. Evaluation of OCL expressions over XML data model. J.UCS: Journal of Universal Computer Science, 20(3):329–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_20_3/evaluation_
of_ocl_expressions.

Messeguer:2010:CAC

[MNDRF10] R. Messeguer, L. Navarro, P. Damian-Reyes, and J. Favela. Context awareness for collaborative learning with uncertainty management. J.UCS: Journal of Universal Computer Science, 16 (12):1556–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_12/context_awareness_
for_collaborative.

McDonald:2013:AIH

[MNF^{+}13] H. McDonald, C. Nugent, D. Finlay, G. Moore, W. Burns, and J. Hallberg. Assessing the impact of the homeML format and the homeML suite within the research community. J.UCS: Journal of Universal Computer Science, 19(17): 2559–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_17/assessing_
the_impact_of.

Matinfar:2013:WRS

[MNL13] F. Matinfar, M. Nematbakhsh, and G. Lausen. Web resource sense disambiguation in Web of data. J.UCS: Journal of Universal Computer Science, 19(13):1871–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
13/web_resource_sense_
disambiguation.

Minder:2012:CGV

[MNS^{+}12] S. Minder, M. Notari, F. Schmitz, R. Hofer, and U. Woermann. Computer generated voice-over in a medical E-learning application: The impact on factual learning outcome. J.UCS: Journal of Universal Computer Science, 18(3):314– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
3/computer_generated_
voice_over.

Maurer:2003:FPI

[MO03] H. Maurer and R. Oliver. The future of PCs and implications on society. J.UCS: Journal of Universal Computer Science, 9(4):300–308, April 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_9_4/the_future_
of_pcs.

McGlinn:2010:STS

[MOG^{+}10] K. McGlinn, E. O’Neill, A. Gibney, D. O’Sullivan, and D. Lewis. Simcon: a tool to support rapid evaluation of smart building application design using context simulation and virtual reality. J.UCS: Journal of Universal Computer Science, 16(15):1992–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
15/simcon_a_tool_to.

Martinez-Ortiz:2007:SAO

[MOMGSRFM07] I. Martínez-Ortiz, P. Moreno-Ger, J. L. Sierra-Rodríguez, and B. Fernández-Manjón. Supporting the authoring and operationalization of educational modelling languages. J.UCS: Journal of Universal Computer Science, 13(7):938– 947, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_7/supporting_the_
authoring_and.

Moore:2008:ICP

[Moo08] T. Moore. An implementation of CLIM presentation types. J.UCS: Journal of Universal Computer Science, 14(20): 3358–??, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_20/an_implementation_
of_clim.

Mosses:2005:CAL

[Mos05] P. D. Mosses. A constructive approach to language definition. J.UCS: Journal of Universal Computer Science, 11(7):1117– 1134, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
7/a_contructive_approach_
to.

Mikroyannidis:2013:WPS

[MOS^{+}13] A. Mikroyannidis, A. Okada, P. Scott, E. Rusman, M. Specht, K. Stefanov, P. Boytchev, A. Protopsaltis, P. Held, S. Hetzner, K. Kikis-Papadakis, and F. Chaimala. weSPOT: A personal and social approach to inquiry-based learning. J.UCS: Journal of Universal Computer Science, 19(14):2093–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
14/wespot_a_personal_
and.

Musicante:2006:EWP

[MP06] M. A. Musicante and E. Potrich. Expressing workflow patterns for Web services: The case of PEWS. J.UCS: Journal of Universal Computer Science, 12(7):903–921, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_12_7/expressing_
workflow_patterns_for.

Manzino:2008:SFM

[MP08] C. Manzino and A. Pardo.

Shortcut fusion of monadic programs. J.UCS: Journal of Universal Computer Science, 14(21): 3431–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_21/shortcut_
fusion_of_monadic.

Muhlbacher:2009:DRD

[MP09] J. R. Mühlbacher and C. Praher. DS RBAC — dynamic sessions in role based access control. J.UCS: Journal of Universal Computer Science, 15(3):538–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_3/ds_rbac_dynamic.

Melia:2010:MDT

[MP10] M. Melia and C. Pahl.

Model-driven transformation and validation of adaptive educational hypermedia using CAVIAr. J.UCS: Journal of Universal Computer Science, 16(19):2862–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16_19/model_
driven_transformation_
and.

Mueller-Prothmann:2004:SSN

[MPF04] T. Mueller-Prothmann and I. Finke. SELaKT — social network analysis as a method for expert localisation and sustainable knowledge transfer. J.UCS: Journal of Universal Computer Science, 10(6):691–701, June 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_10_6/selakt_
                                        social_network_analysis.

Morgado:2016:BSA

[MPF^{+}16] L. Morgado, H. Paredes, B. Fonseca, P. Martins, A. Almeida, A. Vilela, F. Peixinho, and A. Santos. A bot spooler architecture to integrate virtual worlds with elearning management systems for corporate training. J.UCS: Journal of Universal Computer Science, 22(2):271–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_22_2/a_bot_
                                        spooler_architecture.

Marco:2013:CLT

[MPG13] F. A. Marco, V. M. Penichet, and J. A. Gallud. Collaborative elearning through drag & share in synchronous shared workspaces. J.UCS: Journal of Universal Computer Science, 19(7): 894–??, ???? 2013. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_7/collaborative_e_
learning_through.

Mutyam:2004:RTP

[MPK04] M. Mutyam, V. J. Prakash, and K. Krithivasan. Rewriting Tissue P systems. J.UCS: Journal of Universal Computer Science, 10(9):1250–1271, September 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
9/rewriting_tissue_p_
systems.

Modritscher:2011:APR

[MPL11] F. Mödritscher, Z. Petrushyna, and E. L.-C. Law. The application of pattern repositories for sharing PLE practices in networked communities. J.UCS: Journal of Universal Computer Science, 17 (10):1492–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_10/the_application_
of_pattern.

Mathrani:2012:KMI

[MPM12] A. Mathrani, D. Parsons, and S. Mathrani. Knowledge management initiatives in offshore software development: Vendors’ perspectives. J.UCS: Journal of Universal Computer Science, 18(19):2706–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_18_19/knowledge_
                                        management_initiatives_
                                        in.

Mayerwieser:1995:THS

[MPPS95] Wolfgang Mayerwieser, Karl C. Posch, Reinhard Posch, and Volker Schindler. Testing a highspeed data: Path the design of the RSAb crypto chip. J.UCS: Journal of Universal Computer Science, 1(11):728–743, November 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        1_11/testing_a_high_
                                        speed.

Montoto:2008:WLW

[MPR^{+}08] P. Montoto, A. Pan, J. Raposo, J. Losada, F. Bellas, and V. Carneiro. A workflow language for Web automation. J.UCS: Journal of Universal Computer Science, 14 (11):1838–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_11/a_workflow_language_
for.

Mateescu:1995:PPW

[MPRS95] A. Mateescu, Gh. Păun, G. Rozenberg, and A. Salomaa. Parikh prime words and GO-like territories. J.UCS: Journal of Universal Computer Science, 1(12): 790–810, December 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_1_12/parikh_
prime_words_and.

Moss:2005:AFC

[MR05] J. E. B. Moss and R. Rajwar. Atomicity as a firstclass system provision. J.UCS: Journal of Universal Computer Science, 11(5):651–660, May 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_11_5/atomicity_
as_a_first.

Memmel:2008:MBP

[MR08] T. Memmel and H. Reiterer. Model-based and prototyping-driven user interface specification to support collaboration and creativity. J.UCS: Journal of Universal Computer Science, 14(19):3217–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_14_19/model_
based_and_prototyping.

Murthy:2011:CCH

[MR11] V. Narasimha Murthy and A. G. Ramakrishnan. Choice of classifiers in hierarchical recognition of online handwritten Kannada and Tamil aksharas. J.UCS: Journal of Universal Computer Science, 17(1):94–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_1/choice_of_
classifiers_in.

Mellado:2012:OCI

[MR12] D. Mellado and D. G. Rosado. An overview of current information systems security challenges and innovations. J.UCS: Journal of Universal Computer Science, 18(12):1598–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.

org/jucs_18;   http:
//www.jucs.org/jucs_
18_12/an_overview_of_
current;   http://www.
jucs.org/jucsrssfeed-
issue.

Michelini:2014:UGK

[MR14] R. C. Michelini and R. P. Razzoli. Understanding the growth by KILT model and TYPUS metrics. J.UCS: Journal of Universal Computer Science, 20(6):924– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
6/understanding_the_
growth_by.

Medeiros:2014:CRR

[MRGF14] F. Medeiros, M. Ribeiro, R. Gheyi, and B. Fonseca. A catalogue of refactorings to remove incomplete annotations. J.UCS: Journal of Universal Computer Science, 20(5):746–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
20_5/a_catalogue_of_
refactorings.

Makrakis:1998:EEO

[MRK^{+}98] V. Makrakis, S. Retalis, A. Koutoumanos, N. Papaspyrou, and M. Skordalakis. Evaluating the effectiveness of an ODL hypermedia system and courseware at the National Technical University of Athens: a case study. J.UCS: Journal of Universal Computer Science, 4(3):259– 272, March 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_3/evaluating_
the_effectiveness_of;
internal&sk=0C220489.

Molina:2005:SSA

[MRO05] A. I. Molina, M. A. Redondo, and M. Ortega. A system to support asynchronous collaborative learning tasks using PDAs. J.UCS: Journal of Universal Computer Science, 11(9):1543– 1554, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
9/a_system_to_support_
asynchronous.

Molina:2008:CMD

[MROH08] A. I. Molina, M. A. Redondo, M. Ortega, and U. Hoppe. CIAM: a methodology for the development of groupware user interfaces. J.UCS: Journal of Universal Computer Science, 14(9):1435–??, ????

2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_14_9/ciam_a_
methodology_for.

Mayeh:2014:RAC

[MRP14] M. Mayeh, T. Ramayah, and S. Popa. The role of absorptive capacity in the usage of a complex information system: The case of the enterprise information system. J.UCS: Journal of Universal Computer Science, 20(6):826– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
6/the_role_of_absorptive.

Maurer:1994:JUN

[MS94] H. Maurer and K. Schmaranz.

J.UCS — the next generation in electronic journal publishing. J.UCS: Journal of Universal Computer Science, 0(0):118– 126, November 15, 1994. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_0_
0/j_ucs_the_next.

Maharaj:2000:SFM

[MS00] S. Maharaj and C. Shankland. A survey of formal methods applied to leader election in IEEE 1394.

J.UCS: Journal of Universal Computer Science, 6(11):1145–1163, November 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
11/a_survey_of_formal.

Moret:2001:AEN

[MS01] B. M. E. Moret and H. D. Shapiro. Algorithms and experiments: The new (and old) methodology. J.UCS: Journal of Universal Computer Science, 7(5):434–446, May 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_7_5/algorithms_
and_experiments_the.

Margenstern:2003:FTC

[MS03] M. Margenstern and G. Skordev. Fibonacci type coding for the regular rectangular tilings of the hyperbolic plane. J.UCS: Journal of Universal Computer Science, 9(5):398–422, May 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
5/fibonacci_type_coding_
for.

Moriyon:2005:TRC

[MS05] R. Moriyón and F. Saiz.

A tool for the reinforcement of conceptual learning: Description and use experiences. J.UCS: Journal of Universal Computer Science, 11(9):1482– 1493, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
9/a_tool_for_the.

Matturro:2010:MCM

[MS10] G. Matturro and A. Silva.

A model for capturing and managing software engineering knowledge and experience. J.UCS: Journal of Universal Computer Science, 16(3):479– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
3/a_model_for_capturing.

Muller:2011:EMA

[MS11] J. Müller and A. Stocker.

Enterprise microblogging for advanced knowledge sharing: The References@BT case study. J.UCS: Journal of Universal Computer Science, 17(4):532–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_4/enterprise_
microblogging_for_advanced.

Ma:2013:TMC

[MSA13] K. Ma, R. Sun, and A. Abraham. Toward a module-centralized and aspect-oriented monitoring framework in clouds. J.UCS: Journal of Universal Computer Science, 19(15):2241–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_19_15/toward_a_
                                        module_centralized.

Maurer:2003:FMM

[MSC03] H. Maurer, R. Stubenrauch, and D. G. Camhy. Foundations of MIR-ACLE: Multimedia information repository, a computer-supported language effort. J.UCS: Journal of Universal Computer Science, 9(4): 309–348, April 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_9_4/foundations_
                                        of_miracle_multimedia.

Moreno:1999:LIS

[MSF99] W. A. Moreno, J. R. Samson, Jr., and F. J. Falquez. Laser injection of soft faults for the validation of dependability design. J.UCS: Journal of Universal Computer Science, 5(10):712–729, October 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
5_10/laser_injection_
of_soft.

Matsumoto:2006:PNG

[MSHN06] M. Matsumoto, M. Saito, H. Haramoto, and T. Nishimura. Pseudorandom number generation: Impossibility and compromise. J.UCS: Journal of Universal Computer Science, 12(6):672–690, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_6/pseudorandom_number_
generation_impossibility.

Melliar-Smith:2007:AAW

[MSM07] P. M. Melliar-Smith and L. E. Moser. Achieving atomicity for Web services using commutativity of actions. J.UCS: Journal of Universal Computer Science, 13(8):1094– 1109, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_13_8/achieving_
atomicity_for_web.

Mecca:2014:MOG

[MSSV14] G. Mecca, M. Santomauro, D. Santoro, and E. Veltri. Middlewareoriented government interoperability frameworks: A comparison. J.UCS: Journal of Universal Computer Science, 20 (11):1543–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
20_11/middleware_oriented_
government_interoperability.

Mateescu:1995:LAS

[MSSY95] Alexandru Mateescu, Arto Salomaa, Kai Salomaa, and Sheng Yu. Lexical analysis with a simple finite-fuzzy-automaton model. J.UCS: Journal of Universal Computer Science, 1(5):292–311, May 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_1_5/lexical_
analysis_with_a.

Ma:2011:CW

[MSTW11] H. Ma, K.-D. Schewe, B. Thalheim, and Q. Wang. Cloud warehousing. J.UCS: Journal of Universal Computer Science, 17(8): 1183–??, ???? 2011. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_8/cloud_warehousing.

Ma:2012:CMS

[MSTW12] H. Ma, K.-D. Schewe, B. Thalheim, and Q. Wang. Conceptual modelling of services. J.UCS: Journal of Universal Computer Science, 18(17):2361–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_18;   http:
//www.jucs.org/jucs_
18_17/conceptual_modelling_
of_services.

Mauser:1999:ETC

[MT99] H. Mauser and E. Thurner.

Electronic throttle control — a dependability case study. J.UCS: Journal of Universal Computer Science, 5(10):730– 741, October 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_5_10/electronic_
throttle_control_a.

Maurer:2002:NPM

[MT02] H. Maurer and K. Tochtermann. On a new powerful model for knowledge management and its applications. J.UCS: Journal of Universal Computer Science, 8(1):85–96, January 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        1/on_a_new_powerful.

Maier:2008:OAK

[MTB^{+}08] R. Maier, S. Thalmann, F. Bayer, M. Krüger, H. Nitz, and A. Sandow. Optimizing assignment of knowledge workers to office space using knowledge management criteria: The flexible office case. J.UCS: Journal of Universal Computer Science, 14(4):508–525, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_4/optimizing_
                                        assignment_of_knowledge.

Murakawa:1997:IIF

[MTK97] Y. Murakawa, S. Tojo, and S. Kunifuji. Imperfect information flow of agents communication in arrow logic. J.UCS: Journal of Universal Computer Science, 3(11):1266– 1281, November 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_3_11/
                                        imperfect_information_
                                        flow_of.

Muelner:1995:WIR

[Mue95] H. Muelner. WAIS and information retrieval on the Internet. J.UCS: Journal of Universal Computer Science, 1(4):247– 250, April 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
4/wais_and_information_
retrieval.

Muehlbacher:2004:FHT

[Mue04] Joerg R. Muehlbacher.

Full hash table search using primitive roots of the prime residue group Z/p. J.UCS: Journal of Universal Computer Science, 10(9):1239–1249, September 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
10_9/full_hash_table_
search.

Mbale:2003:CSC

[MUF03] J. Mbale, D. Ursino, and X. X. Fei. Cyclical structure converter (CSC): a system for handling the interaction of structured and semi-structured data sources. J.UCS: Journal of Universal Computer Science, 9(5):423– 446, May 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
5/cyclical_structure_
converter_csc.

Muhlhauser:1996:IDE

[Müh96] M. Mühlhäuser. Interdisciplinary development of an electronic class and conference room. J.UCS: Journal of Universal Computer Science, 2(10): 694–710, October 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        interdisciplinary_development_
                                        of_an_electronic_class.

Muller:1998:FSIb

[Mul98a] J. Muller. Foreword to the special issue. J.UCS: Journal of Universal Computer Science, 4(2): 89, February 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs;   http://
                                        medoc.springer.de:8000/
                                        jucs/jucs_4_2/foreword_
                                        to_the_special;   internal&
                                        sk=015D6EC9/jucs_4_2.

Muller:1998:FSIa

[Mul98b]  J.-M. Muller.  Fore-

word to the special issue. J.UCS: Journal of Universal Computer Science, 4(1):1, January 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs;   http://
                                        medoc.springer.de:8000/

jucs/jucs_4_1/editorial;
internal&sk=015D6EC9/
jucs_4_1.

Mullins:2000:NAI

[Mul00] J. Mullins. Nondeterministic admissible interference. J.UCS: Journal of Universal Computer Science, 6(11):1054– 1070, November 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
11/nondeterministic_
admissible_interference.

Mulner:2002:SME

[Mül02] H. Mülner. Software for a multimedia encyclopedia. J.UCS: Journal of Universal Computer Science, 8(10):985–990, October 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
8_10/software_for_a_
multimedia.

Moench:2003:SOB

[MUSA03] E. Moench, M. Ullrich, H.-P. Schnurr, and J. Angele. SemanticMiner — ontology-based knowledge retrieval. J.UCS: Journal of Universal Computer Science, 9(7):682– 696, July 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
7/semantic_miner_ontology_
based.

Moreno:2015:PMA

[MV15] A. Ibáñez Moreno and A. Vermeulen. Profiling a MALL app for English oral practice: A case study. J.UCS: Journal of Universal Computer Science, 21(10):1339–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_10/profiling_
a_mall_app.

Martin-Vide:2000:UPA

[MVM00] C. Martín-Vide and V. Mitrana. Uniquely parsable accepting grammar systems. J.UCS: Journal of Universal Computer Science, 6(9):850– 860, September 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
9/uniquely_parsable_
accepting_grammar.

Martin-Valdivia:2012:LCN

[MVMRULS12] M.-T. Martín-Valdivia, A. Montejo-Ráez, A. Ureña-López, and M. R. Saleh. Learning to classify neutral examples from positive and negative opinions. J.UCS: Journal of Universal Computer Science, 18(16):2319–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_16/learning_
to_classify_neutral.

Martin-Vide:2002:PSS

[MVPP02] C. Martín-Vide, A. Paun, and G. Paun. On the power of P systems with symport rules. J.UCS: Journal of Universal Computer Science, 8(2):317–331, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_8_2/on_the_
power_of.

Madani:2010:PBR

[MWM10] S. Ahmad Madani, D. Weber, and S. Mahlknecht. Position-based routing protocol for low power wireless sensor networks. J.UCS: Journal of Universal Computer Science, 16(9):1215–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_9/position_bbased_
routing_protocol.

Munzer:2005:SGL

[MX05] S. Münzer and B. Xiao.

Small groups learning synchronously online at the workplace: The interaction of factors determining outcome and acceptance. J.UCS: Journal of Universal Computer Science, 11 (3):378–393, March 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_11_3/small_
groups_learning_synchronously.

Ma:2014:SCM

[MYC14] S.-P. Ma, C.-L. Yeh, and P.-C. Chen. Service composition management: A risk-driven approach. J.UCS: Journal of Universal Computer Science, 20(3):302– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
3/service_composition_
management_a.

Ma:2011:RAM

[MYCA11] K. Ma, B. Yang, Z. Chen, and A. Abraham. A relational approach to model transformation with QVT relations supporting model synchronization. J.UCS: Journal of Universal Computer Science, 17 (13):1863–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_13/a_relational_approach_
to.

Mori:2009:FCF

[MYT09] T. Mori, M. Yasugi, and Y. Tsujii. Finecomputable functions on the unit square and their integral. J.UCS: Journal of Universal Computer Science, 15(6):1264– ??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_6/fine_computable_
functions_on.

Miao:2006:PMT

[MYY06] Z. Miao, B. Yuan, and M. Yu. A pervasive multimodal tele-home healthcare system. J.UCS: Journal of Universal Computer Science, 12(1):99– 114, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
1/a_pervasive_multimodal_
tele.

Meduna:2012:CPG

[MZ12] A. Meduna and P. Zemek.

Controlled pure grammar systems. J.UCS: Journal of Universal Computer Science, 18(14):2024–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_18_
                                        14/controlled_pure_grammar_
                                        systems.

Nagy:2006:PEC

[Nag06] F. Nagy. Parameter estimation of the Cauchy distribution in information theory approach. J.UCS: Journal of Universal Computer Science, 12(9):1332–1344, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_12_9/parameter_
                                        estimation_of_the.

Na:2008:OCU

[NAK08] Y. J. Na, S. Abdullaev, and F. I. S. Ko. An optimization of CDN using efficient load distribution and RADS caching algorithm. J.UCS: Journal of Universal Computer Science, 14(14):2329–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        14/an_optimization_of_
                                        cdn.

Nalepa:2010:CKE

[Nal10] G. J. Nalepa. Collective knowledge engineering with Semantic Wikis. J.UCS: Journal of Universal Computer Science, 16(7):1006–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_7/collective_
knowledge_engineering_
with.

Navarro:2009:SBV

[Nav09] A. Navarro. A SWEBOKbased viewpoint of the Web engineering discipline. J.UCS: Journal of Universal Computer Science, 15(17):3169–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_17/a_swebok_
based_viewpoint.

Nieto:2006:IHA

[NBGS06] I. Nieto, J. A. Botía, and A. F. Gómez-Skarmeta. Information and hybrid architecture model of the OCP contextual information management system. J.UCS: Journal of Universal Computer Science, 12(3):357–366, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_3/information_and_
hybrid_architecture.

Nepomuceno-Chamorro:2004:JSM

[NC04] I. A. Nepomuceno-Chamorro.

A Java simulator for membrane computing. J.UCS: Journal of Universal Computer Science, 10(5):620–629, May 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_10_5/a_java_
simulator_for.

Nguyen:2016:CIT

[NCH16] N. T. Nguyen, I. Czarnowski, and D. Hwang. Computational intelligence tools for processing collective data. J.UCS: Journal of Universal Computer Science, 22(6):735–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_22;   http:
//www.jucs.org/jucs_
22_6/computational_intelligence_
tools_for.

Nedjah:2005:SHC

[Nd05] N. Nedjah and L. de Macedo Mourelle. Software/hardware co-design of efficient and secure cryptographic hardware. J.UCS: Journal of Universal Computer Science, 11(1):66–82, January 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_

1/software_hardware_
co_design.

Naldi:2008:NCM

[ND08] M. Naldi and G. D’Acquisto.

A normal copula model for the economic risk analysis of correlated failures in communications networks. J.UCS: Journal of Universal Computer Science, 14(5):786– 799, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_5/a_normal_copula_
model.

Navas-Delgado:2009:ESC

[NDAM09] I. Navas-Delgado and J. F.

Aldana-Montes. Extending SD-core for ontologybased data integration. J.UCS: Journal of Universal Computer Science, 15(17):3201–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_17/extending_
sd_core_for.

Neves:2008:TFE

[NdCFB08] L. A. Pereira Neves, J. M.

de Carvalho, J. Facon, and F. Bortolozzi. Tableform extraction with artefact removal. J.UCS: Journal of Universal Computer Science, 14(2): 252–265, ???? 2008.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_2/table_form_
                                        extraction_with.

Nedjah:2006:EH

[NdMM06a] N. Nedjah and L. de Macedo Mourelle. Evolvable hardware. J.UCS: Journal of Universal Computer Science, 12(4): 367–369, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        12;   http://www.jucs.
                                        org/jucs_12_4#;   http:
                                        //www.jucs.org/jucs_
                                        12_4/evolvable_hardware.

Nedjah:2006:POH

[NdMM06b] N. Nedjah and L. de Macedo Mourelle. Pareto-optimal hardware for substitution boxes. J.UCS: Journal of Universal Computer Science, 12(4):395–407, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        12_4/pareto_optimal_
                                        hardware_for.

Nedjah:2008:EOI

[NdMM08] N. Nedjah and L. de Macedo Mourelle. Evolutionary optimization for intelligent systems design. J.UCS: Journal of Universal Computer Science, 14(15):2453–??, ???? 2008. CODEN ????

ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_14;   http:
//www.jucs.org/jucs_
14_15#;   http://www.
jucs.org/jucs_14_15/
evolutionary_optimization_
for_intelligent.

Nedjah:2012:AMD

[NdMM12] N. Nedjah and L. de Macedo Mourelle.

Adaptive methodologies and designs for networkon-chip based systems. J.UCS: Journal of Universal Computer Science, 18(7):899–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_18;   http:
//www.jucs.org/jucs_
18_7/adaptive_methodologies_
and_designs;   http:/
/www.jucs.org/jucsrssfeed-
issue.

Negri:2005:PRL

[Neg05] S. Negri. Permutability of rules for linear lattices. J.UCS: Journal of Universal Computer Science, 11(12):1986–1995, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
12/permutability_of_
rules_for.

Neher:1998:ESI

[Neh98] M. Neher. Enclosing solutions of an inverse Sturm– Liouville problem for an impedance. J.UCS: Journal of Universal Computer Science, 4(2):178– 192, February 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_2/enclosing_
                                        solutions_of_an.

Nguyen:2005:PIK

[Ngu05]  N. T. Nguyen.  Pro-

cessing inconsistency of knowledge on semantic level. J.UCS: Journal of Universal Computer Science, 11(2):285– 302, February 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_11_2/processing_
                                        inconsistency_of_knowledge.

Nguyen:2009:RCN

[Ngu09] N. T. Nguyen. Rough classification — new approach and applications. J.UCS: Journal of Universal Computer Science, 15(13):2622–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_15_13/rough_
                                        classification_new_approach.

North:2003:BKM

[NH03] K. North and T. Hornung. The benefits of knowledge management

— results of the German award “Knowledge Manager 2002”. J.UCS: Journal of Universal Computer Science, 9(6):463– 471, June 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
9_6/the_benefits_of_
knowledge.

Nguyen:2009:KMA

[NH09] N. T. Nguyen and D.-S.

Huang. Knowledge management for autonomous systems and computational intelligence. J.UCS: Journal of Universal Computer Science, 15(4): 704–??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15;   http://www.jucs.
org/jucs_15_4#;   http:
//www.jucs.org/jucs_
15_4/knowledge_management_
for_autonomous.

Nakamura:2006:SDA

[NHH06] K. Nakamura, T. Higuchi, and N. Hirose. Sequential data assimilation: Information fusion of a numerical simulation and large scale observation data.

J.UCS: Journal of Universal Computer Science, 12(6):608–626, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        12_6/sequential_data_
                                        assimilation_information.

Nguyen:2003:AVF

[NI03] T. V. N. Nguyen and F. Irigoin. Alias verification for Fortran code optimization. J.UCS: Journal of Universal Computer Science, 9(3):270– 297, March 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        3/alias_verification_
                                        for_fortran.

Nielsen:1995:MFD

[NK95] Asger Munk Nielsen and Peter Kornerup. MSBfirst digit serial arithmetic. J.UCS: Journal of Universal Computer Science, 1(7):527– 547, July 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        1_7/msb_first_digit_
                                        serial.

Nunes:2009:ADM

[NKS^{+}09] C. Nunes, U. Kulesza, C. Sant’Anna, I. Nunes, A. Garcia, and C. Lucena. Assessment of the design modularity and stability of multiagent system product lines. J.UCS: Journal of Universal Computer Science, 15(11):2254–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
11/assessment_of_the_
design.

Nam:2010:TSI

[NL10] K. Nam and N. H. Lee.

Typology of service innovation from servicedominant logic perspective. J.UCS: Journal of Universal Computer Science, 16(13):1761–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_13/typology_
of_service_innovation.

Ning:2012:DPB

[NLLJ12] H. Ning, H. Liu, Q. Liu, and G. Ji. Directed path based authentication scheme for the Internet of Things. J.UCS: Journal of Universal Computer Science, 18(9):1112– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_

                                        9/directed_path_based_
                                        authentication.

Nguyen:2007:QSR

[NM07] D.-Q. Nguyen and P. Minet. Quality of service routing in a MANET with OLSR. J.UCS: Journal of Universal Computer Science, 13(1):56– 86, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_
                                        1/quality_of_service_
                                        routing.

Najera:2009:SMA

[NML09] P. Najera, F. Moyano, and J. Lopez. Security mechanisms and access control infrastructure for e-passports and general purpose edocuments. J.UCS: Journal of Universal Computer Science, 15(5):970– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15_
                                        5/security_mechanisms_
                                        and_access.

Nguyen:2007:CI

[NN07] N. T. Nguyen and T. Nishida. Communicative intelligence. J.UCS: Journal of Universal Computer Science, 13(2):133– 134, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13;
http://www.jucs.org/
jucs_13_2#;   http://
www.jucs.org/jucs_13_
2/communicative_intelligence.

Nguyen:2016:QMQ

[NNT16] L. T. Nguyen, N.-T.

Nguyen, and B. Trawiński. A quick method for querying top-k rules from class association rule set. J.UCS: Journal of Universal Computer Science, 22(6):822–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
22_6/a_quick_method_
for.

Nakaya:1998:FAS

[NO98] Y. Nakaya and S. Oishi.

Finding all solutions of nonlinear systems of equations using linear programming with guaranteed accuracy. J.UCS: Journal of Universal Computer Science, 4 (2):171–177, February 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_2/finding_
all_solutions_of.

Naranjo:2013:EUA

[NOGG^{+}13] J. Á. Muñoz Naranjo, P. Orduña, A. Gómez-Goiri, D. López de Ipiña, and L. González Casado. Enabling user access control in energy-constrained wireless smart environments. J.UCS: Journal of Universal Computer Science, 19(17):2490–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_17/enabling_
user_access_control.

Neyem:2008:ISO

[NOP08] A. Neyem, S. F. Ochoa, and J. A. Pino. Integrating service-oriented mobile units to support collaboration in ad-hoc scenarios. J.UCS: Journal of Universal Computer Science, 14(1):88– 122, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
1/integrating_service_
oriented_mobile.

Normark:2010:SUT

[Nør10] K. Nørmark. Systematic unit testing in a read-eval-print loop. J.UCS: Journal of Universal Computer Science, 16(2):296– ??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_

16_2/systematic_unit_
testing_in.

Nagy:2006:PLC

[NPB06] Péter Nagy, Istvan Pintér, and Mihály Bagány. Phasetransitionlike changes in human visual information processing. J.UCS: Journal of Universal Computer Science, 12(9):1345– 1357, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_9/phasetransition_
like_changes_in.

Nikolakopoulos:2009:PPN

[NPC^{+}09] I. G. Nikolakopoulos, C. Z. Patrikakis, A. Cimmino, M. Bauer, and H. Olesen. On the personalization of personal networks — service provision based on user profiles. J.UCS: Journal of Universal Computer Science, 15(12):2353–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
12/on_the_personalization_
of.

Newton:2008:CSO

[NR08a] J. Newton and C. Rhodes.

Custom specializers in object-oriented Lisp. J.UCS: Journal of Universal Computer Science, 14 (20):3370–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_20/custom_specializers_
in_object.

Nourollah:2008:LTA

[NR08b] A. Nourollah and M. Razzazi. A linear time approximation algorithm for ruler folding problem. J.UCS: Journal of Universal Computer Science, 14(4):566–574, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
4/a_linear_time_approximation.

Nagar:2012:RAB

[NR12] A. K. Nagar and T. Robinson. Recent advances in bio-inspired computing: Theory and applications. J.UCS: Journal of Universal Computer Science, 18(13):1757–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_18;   http:
//www.jucs.org/jucs_
18_13/recent_advances_
in_bio;   http://www.
jucs.org/jucsrssfeed-
issue.

Nguyen:2005:MTW

[NS05] N. T. Nguyen and J. Sobecki.

Modern technologies for Web-based adaptive systems. J.UCS: Journal of Universal Computer Science, 11(2):210–212, February 28, 2005. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
 http://www.jucs.org/
jucs_11_2/modern_technologies_
for_web.

Nagy:2006:MCG

[NS06] B. Nagy and L. Szegedi.

Membrane computing and graphical operating systems. J.UCS: Journal of Universal Computer Science, 12(9):1312–1331, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
9/membrane_computing_
and_graphical.

Niemann:2010:UBO

[NSF^{+}10] K. Niemann, M. Scheffel, M. Friedrich, U. Kirschenmann, H.-C. Schmitz, and M. Wolpers. Usagebased object similarity. J.UCS: Journal of Universal Computer Science, 16(16):2272–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_16_16/usage_
                                        based_object_similarity.

Navarro:2005:CCI

[NSFVH05] A. Navarro, J. L. Sierra, A. Fernández-Valmayor, and H. Hernanz. From Chasqui to Chasqui II: an evolution in the conceptualization of virtual objects. J.UCS: Journal of Universal Computer Science, 11(9):1518– 1529, ???? 2005. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        11_9/from_chasqui_to_
                                        chasqui.

Nuernberg:1996:DDL

[NSL96] P. Nuernberg, E. Schneider, and J. Leggett. Designing digital libraries for the hyperliterate age. J.UCS: Journal of Universal Computer Science, 2(9):610–622, September 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        designing_digital_libraries_
                                        for_the_hyperliterate_
                                        age.

Nait-Sidi-Moh:2012:MPE

[NSMBACBG12] A. Nait-Sidi-Moh, M. Bakhouya, W. Ait-Cheik-Bihi, and J. Gaber. Modeling and performance evaluation of a contractbased electronic signature process. J.UCS: Journal of Universal Computer Science, 18(5):676– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
5/modeling_and_performance_
evaluation.

Nguyen:2005:PSE

[NSNK05] L. Nguyen, R. Safavi-Naini, and K. Kurosawa. A provably secure and efficient verifiable shuffle based on a variant of the Paillier cryptosystem. J.UCS: Journal of Universal Computer Science, 11(6):986–1010, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
6/a_provably_secure_
and.

Nadeem:2005:TTA

[NuR05] A. Nadeem and M. Jaffar ur Rehman. TESTAF: a test automation framework for class testing using object-oriented formal specifications. J.UCS: Journal of Universal Computer Science, 11(6): 962–985, ???? 2005. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        11_6/testaf_a_test_automation.

Ngobeni:2012:MDF

[NVB12] S. Ngobeni, H. Venter, and I. Burke. The modelling of a digital forensic readiness approach for wireless local area networks. J.UCS: Journal of Universal Computer Science, 18(12):1721–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_18_
                                        12/the_modelling_of_a.

Novak:2004:SKC

[NW04] J. Novak and M. Wurst. Supporting knowledge creation and sharing in communities based on mapping implicit knowledge. J.UCS: Journal of Universal Computer Science, 10(3):235– 251, March 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_10_3/supporting_
                                        knowledge_creation_and.

Niu:2009:NST

[NWDX09] D. Niu, Y. Wang, C. Duan, and M. Xing. A new short-term power load forecasting model based on chaotic time series and SVM. J.UCS: Journal of Universal Computer Science, 15(13):2726–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
13/a_new_short_term.

Niu:2012:PSS

[NXSA12] Y. Niu, J. Xu, K. Subramanian, and R. Abdullah. P systems with shuffle operation and catalytic-like rules. J.UCS: Journal of Universal Computer Science, 18(13):1782–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_13/p_systems_
with_shuffle.

Nenadic:2005:RBC

[NZCG05] A. Nenadic, N. Zhang, B. Cheetham, and C. Goble. RSA-based certified delivery of E-goods using verifiable and recoverable signature encryption. J.UCS: Journal of Universal Computer Science, 11(1):175– 192, January 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_11_1/rsa_based_
certified_delivery.

Niu:2009:QII

[NZM09] Q. Niu, T. Zhou, and S. Ma. A quantuminspired immune algorithm for hybrid flow shop with makespan criterion. J.UCS: Journal of Universal Computer Science, 15(4):765– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15_
                                        4/a_quantum_inspired_
                                        immune.

Orduna:2014:LNE

[OAR^{+}14] P. Orduña, A. Almeida, S. Ros, D. López de Ipiña, and J. Garcia-Zubia. Leveraging non-explicit social communities for learning analytics in mobile remote laboratories. J.UCS: Journal of Universal Computer Science, 20(15):2043–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        20_15/leveraging_non_
                                        explicit_social.

Ottmann:1995:AF

[OB95] T. Ottmann and C. Bacher. Authoring on the fly. J.UCS: Journal of Universal Computer Science, 1(10):706–717, October 28, 1995. CODEN ????

ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/authoring_on_the_
fly.

Oliveira:2009:FSB

[OBO09] K. Oliveira, K. Breitman, and T. Oliveira. A flexible strategy-based model comparison approach: Bridging the syntactic and semantic gap. J.UCS: Journal of Universal Computer Science, 15(11):2225–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
11/a_flexible_strategy_
based.

Ortega:2010:KAO

[OCB^{+}10] A. Muñoz Ortega, J. M. Alcaraz Calero, J. A. Botía Blaya, G. Martínez Pérez, and F. J. Garcia Clemente. Knowledge authoring with ORE: Testing, debugging and validating knowledge rules in a Semantic Web framework. J.UCS: Journal of Universal Computer Science, 16(9):1234– ??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_9/knowledge_
authoring_with_ore.

Ortega-Cisneros:2007:DIA

[OCRPdlMG07] S. Ortega-Cisneros, J. J. Raygoza-Panduro, and A. de la Mora Gálvez. Design and implementation of the AMCC self-timed microprocessor in FPGAs. J.UCS: Journal of Universal Computer Science, 13(3):377–387, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_13_3/design_and_
                                        implementation_of.

Okeyo:2013:AMO

[OCW13] G. Okeyo, L. Chen, and H. Wang. An agentmediated ontology-based approach for composite activity recognition in smart homes. J.UCS: Journal of Universal Computer Science, 19 (17):2577–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_19_17/an_agent_
                                        mediated_ontology.

Olderog:2003:MRT

[OD03] E.-R. Olderog and H. Dierks. Moby/RT: a tool for specification and verification of real-time systems. J.UCS: Journal of Universal Computer Science, 9(2):88–105, February 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
2/mobyrt_a_tool_for.

Odlyzko:1994:TLG

[Odl94] A. Odlyzko. Tragic loss or good riddance? the impending demise of traditional scholarly journals. J.UCS: Journal of Universal Computer Science, 0(0):3–53, November 15, 1994. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_0_
0/tragic_loss_or_good.

Ordinez:2011:TED

[ODSO11] L. Ordinez, D. Donari, R. Santos, and J. Orozco. Time is not enough: Dealing with behavior in real-time systems. J.UCS: Journal of Universal Computer Science, 17(11):1572–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
11/time_is_not_enough.

Ozcinar:2016:SDS

[OEK16] Z. Ozcinar, N. Ekizoglu, and S. Kanbul. A study on developing a scale for determining the educational usage of mobile communication apps. J.UCS: Journal of Universal Computer Science, 22(1):146–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_22_1/a_study_on_
developing.

Ogata:2013:CWP

[OF13] K. Ogata and K. Futatsugi. Compositionally writing proof scores of invariants in the OTS/CafeOBJ method. J.UCS: Journal of Universal Computer Science, 19(6):771–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_6/compositionally_
writing_proof_scores.

Olivera:2008:BNO

[OFCB08] A. C. Olivera, M. Frutos, J. A. Carballido, and N. B. Brignole. Bus network optimization with a time-dependent hybrid algorithm. J.UCS: Journal of Universal Computer Science, 14(15):2512–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
15/bus_network_optimization_
with.

Oh:2016:AGI

[OHYJ16] K.-J. Oh, M.-D. Hong, U.-N. Yoon, and G.-S. Jo. Automatic generation of interactive cooking video with semantic annotation. J.UCS: Journal of Universal Computer Science, 22(6):742–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_22_6/automatic_
generation_of_interactive.

Oishi:1998:NVM

[Ois98] S. Oishi. Numerical verification method of existence of connecting orbits for continuous dynamical systems. J.UCS: Journal of Universal Computer Science, 4(2):193– 201, February 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_2/numerical_
verification_method_
of.

Oliveira:2008:CCR

[OJSB08] L. S. Oliveira, E. Justino, R. Sabourin, and F. Bortolozzi. Combining classifiers in the ROC-space for off-line signature verification. J.UCS: Journal of Universal Computer Science, 14(2):237–251, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_2/combining_
                                        classifiers_in_the.

Ogata:1998:SBP

[OK98] W. Ogata and K. Kurosawa. Some basic properties of general nonperfect secret sharing schemes. J.UCS: Journal of Universal Computer Science, 4(8):690– 704, August 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_8/some_
                                        basic_properties_of.

Ouimet:2008:TAS

[OL08] M. Ouimet and K. Lundqvist. The timed abstract state machine language: Abstract state machines for real-time system engineering. J.UCS: Journal of Universal Computer Science, 14(12):2007–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_12/the_timed_
                                        abstract_state.

Oliveira:2001:BCA

[Oli01] J. N. Oliveira. Bagatelle in C arranged for VDM SoLo. J.UCS: Journal of Universal Computer Science, 7(8):754–781, August 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
8/bagatelle_in_c_arranged.

Olivares:2012:RVA

[Oli12]  J. Olivares.  Recon-

figurable VBSME architecture using RB-SAD. J.UCS: Journal of Universal Computer Science, 18(2):264– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
/reconfigurable_vbsme_
architecture_using.

Oussalah:2010:TTM

[OMO10] M. Oussalah, Z. Messaoudi, and A. Ouldali. Track-to-track measurement fusion architectures and correlation analysis. J.UCS: Journal of Universal Computer Science, 16(1):37–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_1/track_to_track_
measurement.

Omar:1997:PPP

[ON97] J. Omar and J. M. Noras.

Prototyping on the PC with programmable hardware. J.UCS: Journal of Universal Computer Science, 3(2):86–119, February 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_3_
2/prototyping_on_the_
pc.

Oliva:2008:ALF

[ONRV08] E. Oliva, A. Natali, A. Ricci, and M. Viroli. An adaptation logic framework for Javabased component systems. J.UCS: Journal of Universal Computer Science, 14(13):2158–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_13/an_adaptation_
logic_framework.

Ou:2008:NRM

[OO08] C.-M. Ou and C.-R. Ou. Non-repudiation mechanism of agent-based mobile payment systems: Perspectives on wireless PKI. J.UCS: Journal of Universal Computer Science, 14(14):2309–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
14/non_repudiation_mechanism_
of.

OSullivan:2006:CCA

[OOHS06] T. O’Sullivan, J. O’Donoghue, J. Herbert, and R. Studdert. CAMMD: Contextaware mobile medical devices. J.UCS: Journal of Universal Computer Science, 12(1):45–58, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_12_1/cammd_
context_aware_mobile.

Ortego:2015:CEM

[OP15] R. Gil Ortego and J. A. Gil Pascual. Conceptual evaluation of massive open online courses through pathfinder associative networks. J.UCS: Journal of Universal Computer Science, 21(12):1577–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
12/conceptual_evaluation_
of_massive.

Ochoa:2009:ESP

[OPP09] S. F. Ochoa, J. A. Pino, and F. Poblete. Estimating software projects based on negotiation. J.UCS: Journal of Universal Computer Science, 15(9):1812–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_15_9/estimating_
                                        software_projects_based.

Olejar:1998:CPR

[OS98] D. Olejar and M. Stanek. On cryptographic properties of random Boolean functions. J.UCS: Journal of Universal Computer Science, 4(8):705– 717, August 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_8/on_
                                        cryptographic_properties_
                                        of.

Ozdamli:2016:DSS

[OT16] F. Ozdamli and T. Tavukcu. Determination of secondary school students’ attitudes towards tablet PC supported education. J.UCS: Journal of Universal Computer Science, 22(1):4–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_22_
                                        1/determination_of_secondary_
                                        school.

Ozdamli:2013:ECS

[Ozd13] F. Ozdamli. Effectiveness of cloud systems and social networks in improving self-directed learning abilities and developing positive seamless learning perceptions. J.UCS: Journal of Universal Computer Science, 19(5):602– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
5/effectiveness_of_cloud_
systems.

Papatheocharous:2012:SCM

[PA12] E. Papatheocharous and A. S. Andreou. Software cost modelling and estimation using artificial neural networks enhanced by input sensitivity analysis. J.UCS: Journal of Universal Computer Science, 18(14):2041–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_14/software_
cost_modelling_and.

Palmgren:2005:QSC

[Pal05]  E. Palmgren.  Quo-

tient spaces and coequalisers in formal topology. J.UCS: Journal of Universal Computer Science, 11(12):1996–2007, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_11_12/quotient_
spaces_and_coequalisers.

Pal:2015:EMC

[Pal15] S. Pal. Extending mobile cloud platforms using opportunistic networks: Survey, classification and open issues. J.UCS: Journal of Universal Computer Science, 21(12):1594–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_21_
                                        12/extending_mobile_
                                        cloud_platforms.

Papagiannakis:2013:MAI

[PANS13] H. Papagiannakis, M. Antona, S. Ntoa, and C. Stephanidis. A multimodal ambient intelligence environment for playful learning. J.UCS: Journal of Universal Computer Science, 19 (17):2617–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        19_17/a_multimodal_ambient_
                                        intelligence.

Park:2009:USC

[Par09] J. H. Park. USF-PAS: Study on core security technologies for ubiquitous security framework. J.UCS: Journal of Universal Computer Science, 15(5):1065–??, ???? 2009.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_5/usf_pas_study_on.

Paulson:1999:GTP

[Pau99]  L. C. Paulson.  A

generic tableau prover and its integration with Isabelle. J.UCS: Journal of Universal Computer Science, 5(3):73–87, March 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
3/a_generic_tableau_
prover.

Paun:2007:SNS

[Pau07] G. Paun. Spiking neural P systems with astrocytelike control. J.UCS: Journal of Universal Computer Science, 13(11): 1707–1721, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_11/spiking_neural_
p_systems.

Pauly:2009:RMR

[Pau09]  A. Pauly.  Represent-

ing measurement results. J.UCS: Journal of Universal Computer Science, 15(6):1280–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15_
                                        6/representing_measurement_
                                        results.

Pauly:2010:HIF

[Pau10] A. Pauly. How incomputable is finding Nash equilibria? J.UCS: Journal of Universal Computer Science, 16(18): 2686–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        16_18/how_incomputable_
                                        is_finding.

Paulin:2013:TSS

[Pau13] A. Paulin. Towards selfservice government — a study on the computability of legal eligibilities. J.UCS: Journal of Universal Computer Science, 19(12):1761–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_19_
                                        12/towards_self_service_
                                        government.

Pavec:1995:SAP

[Pav95] Raymond Pavec. Some algorithms providing rigorous bounds for the eigenvalues of a matrix. J.UCS: Journal of Universal Computer Science, 1(7):548– 558, July 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
1_7/some_algorithms_
providing_rigorous.

Pivec:2004:RAP

[PB04] M. Pivec and K. Baumann. The role of adaptation and personalisation in classroombased learning and in elearning. J.UCS: Journal of Universal Computer Science, 10(1):73–89, January 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
1/the_role_of_adaptation.

Park:2005:IQA

[PB05] J. Park and K. S. Barber. Information quality assurance by lazy exploration of information source combinations space in open multi-agent systems. J.UCS: Journal of Universal Computer Science, 11(1):193–209, January 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
1/information_quality_
assurance_by.

Panczyk:2014:SSA

[PB14] M. Pańczyk and H. Bielak. A self-stabilizing algorithm for locating the center of maximal outerplanar graphs. J.UCS: Journal of Universal Computer Science, 20(14): 1951–??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        20_14/a_self_stabilizing_
                                        algorithm.

Passos:2007:MRL

[PBB07] L. Teixeira Passos, M. A. S. Bigonha, and R. S. Bigonha. A methodology for removing LALR(k) conflicts. J.UCS: Journal of Universal Computer Science, 13(6):737– 752, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_
                                        6/a_methodology_for_
                                        removing.

Passos:2008:LPG

[PBB08] L. Teixeira Passos, M. A. S. Bigonha, and R. S. Bigonha. An LALR parser generator supporting conflict resolution. J.UCS: Journal of Universal Computer Science, 14(21):3447–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_21/an_lalr_parser_
generator.

Perez:2007:URR

[PBTW07] O. Pérez, Y. Berviller, C. Tanougast, and S. Weber. The use of runtime reconfiguration on FPGA circuits to increase the performance of the AES algorithm implementation. J.UCS: Journal of Universal Computer Science, 13(3):349– 362, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
3/the_use_of_runtime.

Putnik:2014:TVE

[PCC14] G. D. Putnik and M. M.

Cruz-Cunha. A taxonomy for virtual enterprises. J.UCS: Journal of Universal Computer Science, 20(6):859–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_20_6/a_taxonomy_
for_virtual.

Pham:2011:CAC

[PCKJ11] M. C. Pham, Y. Cao, R. Klamma, and M. Jarke. A clustering approach for collaborative filtering recommendation using social network analysis. J.UCS: Journal of Universal Computer Science, 17(4):583– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
4/a_clustering_approach_
for.

Poblet:2011:OOW

[PCLCC11] M. Poblet, P. Casanovas, J.-M. López-Cobo, and N. Casellas. ODR, ontologies, and Web 2.0. J.UCS: Journal of Universal Computer Science, 17(4):618–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_4/odr_ontologies_
and_web20.

Piedra:2015:SOE

[PCLT15] N. Piedra, J. Chicaiza, J. López, and E. Tovar. Seeking open educational resources to compose massive open online courses in engineering education an approach based on linked open data. J.UCS: Journal of Universal Computer Science, 21(5):679– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_

5/seeking_open_educational_
resources.

Pastor:2013:LSL

[PCS^{+}13] R. Pastor, A. C. Caminero, D. Sánchez, R. Hernández, S. Ros, A. Robles-Gómez, and L. Tobarra. Laboratories as a service (LaaS): Using cloud technologies in the field of education. J.UCS: Journal of Universal Computer Science, 19(14):2112–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
14/laboratories_as_a_
service.

Popescu:1999:LDC

[PD99] I. Popescu and M. Dumitrescu. Laha distribution: Computer generation and applications to life time modelling. J.UCS: Journal of Universal Computer Science, 5(8):471–481, August 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
8/laha_distribution_
computer_generation.

Pivec:2004:GBL

[PD04] M. Pivec and O. Dziabenko. Game-based learning in universities and lifelong learning:

“UniGame: Social skills and knowledge training” game concept. J.UCS: Journal of Universal Computer Science, 10(1): 14–26, January 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_10_1/game_based_
learning_in.

Pedrinaci:2010:TNW

[PD10] C. Pedrinaci and J. Domingue. Toward the next wave of services: Linked services for the Web of data. J.UCS: Journal of Universal Computer Science, 16(13):1694–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
13/toward_the_next_wave.

Portilla:2006:MAN

[PdCdlTR06] J. Portilla, A. de Castro, E. de la Torre, and T. Riesgo. A modular architecture for nodes in wireless sensor networks. J.UCS: Journal of Universal Computer Science, 12(3):328–339, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_12_3/a_modular_
architecture_for.

Panadero:2014:PWP

[PdlCBKN14] C. Fernández Panadero, V. de la Cruz Barquero, C. Delgado Kloos, and D. Morán Núñez. PhyMEL-WS: Physically experiencing the virtual world. insights into mixed reality and flow state on board a wheelchair simulator. J.UCS: Journal of Universal Computer Science, 20(12):1629–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_20_12/phymel_ws_
physically_experiencing.

Pereira:2004:TRM

[PdP^{+}04] F. M. Q. Pereira, M. T.

de Oliveira Valente, W. S. Pires, R. da Silva Bigonha, and M. A. da Silva Bigonha. Tactics for remote method invocation. J.UCS: Journal of Universal Computer Science, 10(7):824–842, July 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
7/tactics_for_remote_
method.

Pellas:2014:EIA

[Pel14] N. Pellas. Exploring interrelationships among high school students’ engagement factors in introductory programming courses via a 3D multi-user serious game created in Open Sim. J.UCS: Journal of Universal Computer Science, 20(12):1608–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_20_12/exploring_
interrelationships_among_
high.

Panach:2008:CIR

[PEPP08] J. I. Panach, S. España, I. Pederiva, and Ó. Pastor. Capturing interaction requirements in a model transformation technology based on MDA. J.UCS: Journal of Universal Computer Science, 14(9):1480–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_9/capturing_
interaction_requirements_
in.

Pettersson:2009:SFF

[Pet09]  U. Pettersson.  Suc-

cess and failure factors for KM: The utilization of knowledge in the Swedish armed forces. J.UCS: Journal of Universal Computer Science, 15(8):1735– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
8/success_and_failure_
factors.

Petrakis:2012:CCC

[Pet12] I. Petrakis. The contrapositive of countable choice for inhabited sets of naturals. J.UCS: Journal of Universal Computer Science, 18(20):2879–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
20/the_contrapositive_
of_countable.

Pinto:2011:MQA

[PF11] M. Pinto and L. Fuentes.

Modeling quality attributes with aspectoriented architectural templates. J.UCS: Journal of Universal Computer Science, 17(5):639– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
5/modeling_quality_attributes_
with.

Polcak:2015:CSB

[PF15] L. Polcák and B. Franková.

Clock-skew-based computer identification: Traps and pitfalls. J.UCS: Journal of Universal Computer Science, 21(9):1210– ??, ???? 2015. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_21_9/clock_skew_
                                        based_computer.

Panizo:2007:WFI

[PFS07] L. Panizo, R.-Á. Fernández, and L. Sánchez. A WebQuest framework to improve the study of deadlock and process synchronization. J.UCS: Journal of Universal Computer Science, 13(7):932– 937, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_13_7/a_webquest_
                                        framework_to.

Perez-Gonzalez:2015:PSP

[PGDD15] D. Pérez-González and R. Díaz-Díaz. Public services provided with ICT in the smart city environment: The case of Spanish cities. J.UCS: Journal of Universal Computer Science, 21(2):248– ??, ???? 2015. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        21_2/public_services_
                                        provided_with.

Perez-Gonzalez:2014:VCL

[PGSAP14] D. Perez-Gonzalez, P. Soto-Acosta, and S. Popa. A virtual campus for elearning inclusion: The case of SVC-G9. J.UCS: Journal of Universal Computer Science, 20 (2):240–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_20_2/a_virtual_
campus_for.

Pham:2009:EDR

[PGT09] D. H. Pham, G. Governatori, and S. Thakur. Extended defeasible reasoning for common goals in n-person argumentation games. J.UCS: Journal of Universal Computer Science, 15(13):2653–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_13/extended_
defeasible_reasoning_
for.

Petraschek:2008:SUA

[PHJ^{+}08] M. Petraschek, T. Hoeher, O. Jung, H. Hlavacs, and W. Gansterer. Security and usability aspects of man-in-the-middle attacks on ZRTP. J.UCS: Journal of Universal Computer Science, 14(5): 673–692, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_5/security_and_usability_
                                        aspects.

Piveta:2006:DBS

[PHPP06] E. Kessler Piveta, M. Hecht, M. Soares Pimenta, and R. T. Price. Detecting bad smells in AspectJ. J.UCS: Journal of Universal Computer Science, 12(7):811–827, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_12_7/detecting_
                                        bad_smells_in.

Pan:2004:SAM

[PI04] L. Pan and T.-O. Ishdorj. P systems with active membranes and separation rules. J.UCS: Journal of Universal Computer Science, 10(5):630– 649, May 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        10_5/p_systems_with_
                                        active.

Park:2015:DLA

[PJ15] Y. Park and I.-H. Jo. Development of the learning analytics dashboard to support students’ learning performance. J.UCS: Journal of Universal Computer Science, 21(1): 110–??, ???? 2015. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
21_1/development_of_
the_learning.

Pham:2012:BSP

[PJH12] X. H. Pham, J. J. Jung, and D. Hwang. Beating social pulse: Understanding information propagation via online social tagging systems. J.UCS: Journal of Universal Computer Science, 18(8):1022–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_8/beating_social_
pulse_understanding.

Pham:2013:IME

[PJN13] X. H. Pham, J. J. Jung, and N. T. Nguyen. Integrating multiple experts for correction process in interactive recommendation systems. J.UCS: Journal of Universal Computer Science, 19(4):581–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_4/integrating_multiple_
experts_for.

Pampin:2015:ERP

[PJO15] H. J. Corona Pampín, H. Jerbi, and M. P.

O’Mahony. Evaluating the relative performance of collaborative filtering recommender systems. J.UCS: Journal of Universal Computer Science, 21(13):1849–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
21_13/evaluating_the_
relative_performance.

Perez-Jimenez:2004:EFS

[PJRC04] M. J. Pérez-Jiménez and F. J. Romero-Campero. An efficient family of P systems for packing items into bins. J.UCS: Journal of Universal Computer Science, 10(5):650–670, May 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
5/an_efficient_family_
of.

Prechelt:1998:FVP

[PK98] L. Prechelt and Ch. Krämer. Functionality versus practicality: Employing existing tools for recovering structural design patterns. J.UCS: Journal of Universal Computer Science, 4(12):866–882, December 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_4_
12/functionality_versus_
practicality_employing.

Papazoglou:2008:IMA

[PKP08] P. M. Papazoglou, D. A.

Karras, and R. C. Papademetriou. An improved multi-agent simulation methodology for modelling and evaluating wireless communication systems resource allocation algorithms. J.UCS: Journal of Universal Computer Science, 14(7): 1061–1079, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_7/an_improved_multi_
agent.

Phuong:2009:DBO

[PKSR09] N. H. Q. Phuong, H.-J.

Kang, Y.-S. Suh, and Y.- S. Ro. A DCM based orientation estimation algorithm with an inertial measurement unit and a magnetic compass. J.UCS: Journal of Universal Computer Science, 15(4):859–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_15_4/a_dcm_
based_orientation.

Parapar:2014:CPL

[PLB14] J. Parapar, D. E. Losada, and A. Barreiro. Combining psycho-linguistic, content-based and chatbased features to detect predation in chatrooms. J.UCS: Journal of Universal Computer Science, 20(2):213–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_20_2/combining_
                                        psycho_linguistic_content.

Penalver:2013:DDC

[PLBG13] A. Peñalver, J. J. López, F. Botella, and J. A. Gallud. Defining distribution constraints in distributed user interfaces. J.UCS: Journal of Universal Computer Science, 19(6):831– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_19_
                                        6/defining_distribution_
                                        constraints_in.

Penichet:2008:ESF

[PLG^{+}08] V. M. Penichet, M. D. Lozano, J. A. Gallud, R. Tesoriero, M. L. Rodríguez, J. L. Garrido, M. Noguera, and M. V. Hurtado. Extending and supporting featured user interface models for the development of groupware applications. J.UCS: Journal of Universal Computer Science, 14(19): 3053–??, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_19/extending_and_
supporting_featured.

Preciado:2008:AWU

[PLSF08] J. C. Preciado, M. Linaje, and F. Sanchez-Figueroa. Adapting Web 1.0 user interfaces to Web 2.0 multidevice user interfaces using RUX-method. J.UCS: Journal of Universal Computer Science, 14(13):2239–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_13/adapting_
web_1.0_user.

Perera:2014:USM

[PMAM14] I. Perera, D. Meedeniya, C. Allison, and A. Miller. User support for managed immersive education: An evaluation of inworld training for Open-Sim. J.UCS: Journal of Universal Computer Science, 20(12):1690–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_20_
                                        12/user_support_for_
                                        managed.

Perrone:2009:AEE

[PMLL09] R. Perrone, R. Macedo, G. Lima, and V. Lima. An approach for estimating execution time probability distributions of component-based realtime systems. J.UCS: Journal of Universal Computer Science, 15 (11):2142–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        15_11/an_approach_for_
                                        estimating.

Prechelt:2002:FPA

[PMP02] L. Prechelt, G. Malpohl, and M. Philippsen. Finding plagiarisms among a set of programs with JPlag. J.UCS: Journal of Universal Computer Science, 8(11):1016–1038, November 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        11/finding_plagiarisms_
                                        among_a.

Paredes:2008:DCU

[PMRO08] M. Paredes, A. I. Molina, M. A. Redondo, and M. Ortega. Designing collaborative user interfaces for ubiquitous applications using CIAM: The AULA case study. J.UCS: Journal of Universal Computer Science, 14(16):2680–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_16/designing_
collaborative_user_interfaces.

Priss:2004:MLD

[PO04] U. Priss and L. J.

Old. Modelling lexical databases with formal concept analysis. J.UCS: Journal of Universal Computer Science, 10(8):967–984, August 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
8/modelling_lexical_
databases_with.

Piekarczyk:2011:HGG

[PO11] M. Piekarczyk and M. R.

Ogiela. Hierarchical graph-grammar model for secure and efficient handwritten signatures classification. J.UCS: Journal of Universal Computer Science, 17(6):926– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_

                                        6/hierarchical_graph_
                                        grammar_model.

Pires:2011:SCA

[POB11a] P. F. Pires, F. Oquendo, and A. P. Terra Bacelo. Software components, architectures and reuse. J.UCS: Journal of Universal Computer Science, 17(5):635–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_17;   http:
                                        //www.jucs.org/jucs_
                                        17_5/software_components_
                                        architectures_and;
                                        http://www.jucs.org/
                                        jucsrssfeed-issue.

Poblet:2011:RLG

[Pob11b] M. Poblet. Rule of law on the go: New developments of mobile governance. J.UCS: Journal of Universal Computer Science, 17(3):498– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_17_
                                        3/rule_of_law_on_the.

Pavelec:2008:UCA

[POJB08] D. Pavelec, L. S. Oliveira, E. Justino, and L. V. Batista. Using conjunctions and adverbs for author verification. J.UCS: Journal of Universal Computer Science, 14(18):2967–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_14_18/using_
conjunctions_and_adverbs.

Pool:2003:ACS

[Poo03] M. Pool. An applied calculus for spatial accessibility reasoning. J.UCS: Journal of Universal Computer Science, 9(9): 986–1007, September 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_9_9/an_applied_
calculus_for.

Popova:1995:FCI

[Pop95] E. Popova. On a formally correct implementation of IEEE computer arithmetic. J.UCS: Journal of Universal Computer Science, 1(7):560– 569, July 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
7/on_a_formally_correct.

Popova:1998:ASC

[Pop98]  E. D. Popova.  Al-

gebraic solutions to a class of interval equations. J.UCS: Journal of Universal Computer Science, 4(1):48–67, January 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_1/algebraic_
solutions_to_a.

Popel:2005:VMI

[Pop05] D. V. Popel. Visualization and manipulation of incomplete and uncertain dependencies by decision diagrams. J.UCS: Journal of Universal Computer Science, 11(11): 1849–1862, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_11/visualization_
and_manipulation_of.

Popescu:2007:BSF

[Pop07]  D.-R. Popescu.  Bal-

ance in systems of finite sets with applications. J.UCS: Journal of Universal Computer Science, 13(11):1755–1766, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
11/balance_in_systems_
of.

Paredes:2010:MSH

[POR10] P. Paredes, A. Ortigosa, and P. Rodriguez. A method for supporting heterogeneous-group formation through heuristics and visualization. J.UCS: Journal of Universal Computer Science, 16(19):2882–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_19/a_method_
for_supporting.

Posch:1998:PDA

[Pos98]  R. Posch.  Protecting

devices by active coating. J.UCS: Journal of Universal Computer Science, 4(7):652–668, July 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_7/protecting_
devices_by_active; internal&
sk=0C220489.

Posch:2001:WIE

[Pos01] R. Posch. Will Internet ever be secure? J.UCS: Journal of Universal Computer Science, 7(5): 447–455, May 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
5/will_internet_ever_
be.

Petre:1999:MAS

[PP99] I. Petre and L. Petre.

Mobile ambients and P - systems. J.UCS: Journal of Universal Computer Science, 5(9):588– 598, September 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
5_9/mobile_ambients_
and_p.

Pinz:1995:RAM

[PPG95] Axel Pinz, Manfred Prantl, and Harald Ganster. A robust affine matching algorithm using an exponentially decreasing distance function. J.UCS: Journal of Universal Computer Science, 1 (8):614–631, August 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/a_robust_affine_
matching_algorithm_using_
an_exponentially_decreasing_
distance_function.

Paun:2004:SBW

[PPJ04] G. Paun and M. J. Pérez-Jiménez. Second brainstorming week on membrane computing. J.UCS: Journal of Universal Computer Science, 10(5):499–501, May 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/free;

http://www.jucs.org/
jucs_10_5/second_brainstorming_
week_on.

Podobnik:2008:ABS

[PPJ08] V. Podobnik, A. Petric, and G. Jezic. An agentbased solution for dynamic supply chain management. J.UCS: Journal of Universal Computer Science, 14(7):1080– 1104, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_7/an_agent_based_
solution.

Pohjola:2011:PKS

[PPP^{+}11] M. V. Pohjola, P. Pohjola, S. Paavola, M. Bauters, and J. T. Tuomisto. Pragmatic knowledge services. J.UCS: Journal of Universal Computer Science, 17(3):472–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_3/pragmatic_
knowledge_services.

Pieprzyk:1999:RSF

[PQ99] J. Pieprzyk and Ch. Xin Qu. Rotation-symmetric functions and fast hashing. J.UCS: Journal of Universal Computer Science, 5(1):20–31, January 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_5_
                                        1/rotation_symmetric_
                                        functions_and.

Poizat:2006:FAD

[PR06] P. Poizat and J.-C. Royer. A formal architectural description language based on symbolic transition systems and temporal logic. J.UCS: Journal of Universal Computer Science, 12(12): 1741–1782, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_12_12/a_formal_
                                        architectural_description.

Pardede:2009:SXH

[PRAT09] E. Pardede, J. W. Rahayu, R. Kaur Aujla, and D. Taniar. SQL/ XML hierarchical query performance analysis in an XML-enabled database system. J.UCS: Journal of Universal Computer Science, 15(10):2058–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15_
                                        10/sqlxml_hierarchical_
                                        query_performance.

Paralic:2011:MKP

[PRB^{+}11] J. Paralič, C. Richter, F. Babič, J. Wagner, and M. Raček. Mirroring of knowledge practices based on user-defined patterns. J.UCS: Journal of Universal Computer Science, 17(10):1474–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_10/mirroring_
of_knowledge_practices.

Palomino-Ramirez:2013:LMC

[PRBLAP^{+}13]  L.  Palomino-Ramírez,

M. L. Bote-Lorenzo, J. I. Asensio-Pérez, L. Vignollet, and Y. A. Dimitriadis. LeadFlow4LD: a method for the computational representation of the learning flow and data flow in collaborative learning. J.UCS: Journal of Universal Computer Science, 19(6):805–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_6/lead_flow4LD_a_
method.

Perez:2013:MDD

[PRCCS13] J. Pérez, I. Ramos, J. A.

Carsí, and C. Costa-Soria. Model-driven development of aspectoriented software architectures. J.UCS: Journal of Universal Computer Science, 19(10):1433–??, ???? 2013. CODEN ????

ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_19_10/model_
                                        driven_development_of.

Perez-Rodriguez:2010:EMA

[PRCRARLN10] R. Perez-Rodriguez, M. Caeiro-Rodriguez, L. Anido-Rifon, and M. Llamas-Nistal. Execution model and authoring middleware enabling dynamic adaptation in educational scenarios scripted with Po-EML. J.UCS: Journal of Universal Computer Science, 16(19):2821–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_16_19/execution_
                                        model_and_authoring.

Prechelt:1997:WWN

[Pre97] L. Prechelt. Why we need an explicit forum for negative results. J.UCS: Journal of Universal Computer Science, 3(9):1074– 1083, September 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_3_9/why_
                                        we_need_an.

Preece:2004:EET

[Pre04]  J. Preece.  Etiquette,

empathy and trust in communities of practice:

Stepping-stones to social capital. J.UCS: Journal of Universal Computer Science, 10(3):294–302, March 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
3/etiquette_empathy_
and_trust.

Prechelt:2012:FNR

[Pre12] L. Prechelt. The forum for negative results (FNR) guest editorial. J.UCS: Journal of Universal Computer Science, 18(20):2748–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_20/the_forum_
for_negative.

Paun:1995:GBS

[PRS95] Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa. Grammars based on the shuffle operation. J.UCS: Journal of Universal Computer Science, 1(1):67–82, January 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_1_
1/grammars_based_on_
the_shuffle_operation/
html/abstract.html.

Palleja:2008:UOF

[PRT^{+}08] T. Pallejà, E. Rubión, M. Teixido, M. Tresanchez, A. Fernández del Viso, C. Rebate, and J. Palacin. Using the optical flow to implement a relative virtual mouse controlled by head movements. J.UCS: Journal of Universal Computer Science, 14(19):3127–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        19/using_the_optical_
                                        flow.

Platzner:1995:EPC

[PRW95] M. Platzner, B. Rinner, and R. Weiss. Exploiting parallelism in constraint satisfaction for qualitative simulation. J.UCS: Journal of Universal Computer Science, 1(12):811– 820, December 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        exploiting_parallelism_
                                        in_constraint_satisfaction_
                                        for_qualitative_simulation.

Panaitopol:1995:BHI

[PŞ95] Laurenţiu Panaitopol and Doru Ştefănescu. Bounds for heights of integer polynomial factors. J.UCS: Journal of Universal Computer Science, 1(8): 603–613, August 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/bounds_
for_heights_of_integer_
polynomial_factors.

Prinz:1997:TCT

[PS97] W. Prinz and A. Syri.

Two complementary tools for the cooperation in a ministerial environment. J.UCS: Journal of Universal Computer Science, 3(8):843–864, August 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs_3_8/two_complementary_
tools_for;   internal&
sk=05460486.

Power:2004:FMF

[PS04] J. F. Power and D. Sinclair. A formal model of Forth control words in the pi-calculus. J.UCS: Journal of Universal Computer Science, 10(9):1272– 1293, September 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
10_9/a_formal_model_
of.

Petho:2006:IHE

[PS06] A. Pethö and D. Sima.

Informatics in higher education. J.UCS: Journal of Universal Computer Science, 12(9):1085– 1086, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12;
 http://www.jucs.org/
jucs_12_9#;   http://
www.jucs.org/jucs_12_
9/informatics_in_higher_
education.

Pencole:2009:CBD

[PS09] Y. Pencolé and A. Subias. A chronicle-based diagnosability approach for discrete timed-event systems: Application to Web-services. J.UCS: Journal of Universal Computer Science, 15 (17):3246–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_17/a_chronicle_based_
diagnosability.

Predic:2012:LPA

[PS12] B. Predic and D. Stojanovic. Localized processing and analysis of accelerometer data in detecting traffic events and driver behaviour. J.UCS: Journal of Universal Computer Science, 18(9):1152–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_9/localized_
processing_and_analysis.

Pimentel:1998:SST

[PSF98] M. G. C. Pimentel, J. B.

Santos, Jr., and R. P. M. Fortes. Supporting structured teaching material in the WWW. J.UCS: Journal of Universal Computer Science, 4(11):825– 838, November 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_11/
supporting_structured_
teaching_material; internal&
sk=01237079.

Popa:2007:HLS

[PSS07] A. Popa, A. Sofronia, and G. Stefanescu. High-level structured interactive programs with registers and voices. J.UCS: Journal of Universal Computer Science, 13(11):1722–1754, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_13_11/high_
level_structured_interactive.

Park:2013:PPM

[PSS^{+}13] Y. Park, C. Sur, S. Shin, K.-H. Rhee, and C. Seo. A privacy preserving message delivery protocol using identity-hidden index in VDTNs. J.UCS: Journal of Universal Computer Science, 19(16): 2385–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_16/a_privacy_
preserving_message.

Paredes:2007:CCF

[PSVOVI07] M. Paredes, P. P. Sánchez-Villalón, M. Ortega, and J. Á. Velázquez-Iturbide. Collaborative composition in a foreign language with handheld computing and Web tools. J.UCS: Journal of Universal Computer Science, 13(7):948–958, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
7/collaborative_composition_
in_a.

Paraiso:2009:UEC

[PT09] E. Cabrera Paraiso and C. A. Tacla. Using embodied conversational assistants to interface users with multi-agent based CSCW applications: The WebAnima Agent. J.UCS: Journal of Universal Computer Science, 15(9):1991– ??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_9/using_embodied_
conversational_assistants.

Pirker:2010:TVT

[PT10] M. Pirker and R. Toegl.

Towards a virtual trusted platform. J.UCS: Journal of Universal Computer Science, 16(4):531– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
4/towards_a_virtual_
trusted.

Pais:2013:PAB

[PT13] J. Pais and A. Tasistro.

Proof assistant based on didactic considerations. J.UCS: Journal of Universal Computer Science, 19(11):1570–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_19_11/proof_
assistant_based_on.

PereiraeSilva:2009:ADI

[PTL^{+}09] G. Pereira e Silva, M. Thielo, R. Dueire Lins, B. Miro, and S. J. Simske. Automatically deciding if a document was scanned or photographed. J.UCS: Journal of Universal Computer Science, 15 (18):3364–??, ???? 2009.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        15_18/automatically_
                                        deciding_if_a.

Pan:2011:WSD

[PTL11] Y. Pan, Y. Tang, and S. Li. Web services discovery in a pay-as-you-go fashion. J.UCS: Journal of Universal Computer Science, 17(14):2029–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_17_
                                        14/web_services_discovery_
                                        in.

Perez-Toledano:2008:SDA

[PTNMC08] M. A. Pérez-Toledano, A. Navasa, J. M. Murillo, and C. Canal. A safe dynamic adaptation framework for aspectoriented software development. J.UCS: Journal of Universal Computer Science, 14(13): 2212–??, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_13/a_safe_dynamic_
                                        adaptation.

Palesi:2012:DRR

[PTO^{+}12] M. Palesi, R. Tornero, J. M. Orduñna, V. Catania, and D. Panno. Designing robust routing algorithms and mapping cores in networks-on-chip: a multi-objective evolutionary-based approach. J.UCS: Journal of Universal Computer Science, 18(7):937–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_7/designing_
robust_routing_algorithms.

Popova:1997:GBS

[PU97] E. D. Popova and Ch. P.

Ullrich. Generalizing BIAS specifications. J.UCS: Journal of Universal Computer Science, 3(1): 23–41, January 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_3_
1/generalizing_bias.

Puccetti:2010:SAX

[Puc10]  A. Puccetti.  Static

analysis of the XEN kernel using Frama-C. J.UCS: Journal of Universal Computer Science, 16(4):543–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_4/static_analysis_
of_the.

Pam:1995:CWH

[PV95] A. Pam and A. Vermeer. A comparison of WWW and Hyper-G. J.UCS: Journal of Universal Computer Science, 1(11):744–750, November 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/a_comparison_of_
                                        www_and_hyperg.

Petri:2016:HEE

[PvW16] G. Petri and C. Gresse von Wangenheim. How to evaluate educational games: a systematic literature review. J.UCS: Journal of Universal Computer Science, 22(7): 992–??, ???? 2016. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        22_7/how_to_evaluate_
                                        educational.

Paun:2000:SSS

[PY00] G. Paun and T. Yokomori. Simulating H systems by P systems. J.UCS: Journal of Universal Computer Science, 6(1):178– 193, January 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_6_1/simulating_
                                        h_systems_by.

Piveta:2003:AWS

[PZ03] E. K. Piveta and L. C.

Zancanella. Aspect weaving strategies. J.UCS: Journal of Universal Computer Science, 9 (8):970–983, August 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_9_8/aspect_
weaving_strategies.

Pecho:2009:APW

[PZDH09] P. Pecho, F. Zboril, Jr., M. Drahansky, and P. Hanacek. Agent platform for wireless sensor network with support for cryptographic protocols. J.UCS: Journal of Universal Computer Science, 15(5):992– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
5/agent_platform_for_
wireless.

Peng:2009:TLB

[PZJ09] T. Peng, Q. Zheng, and Y. Jin. Transmission latency based network friendly tree for peer-to-peer streaming. J.UCS: Journal of Universal Computer Science, 15(9):2011–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15_
                                        9/transmission_latency_
                                        based_network.

Padilla-Zea:2013:MEE

[PZLAS^{+}13] N. Padilla-Zea, J. R. López-Arcos, J. L. González Sánchez, F. L. Gutiérrez Vela, and A. Abad-Arranz. A method to evaluate emotions in educational video games for children. J.UCS: Journal of Universal Computer Science, 19(8):1066– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_19_
                                        8/a_method_to_evaluate.

Qian:2012:SES

[QC12] X. Qian and X. Che. Security-enhanced search engine design in Internet of Things. J.UCS: Journal of Universal Computer Science, 18(9):1218– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_18_
                                        9/security_enhanced_
                                        search_engine.

Qin:2012:ATA

[QF12] T. T. Qin and S. Fujita. Automatic tag attachment scheme based on text clustering for efficient file search in unstructured peer-to-peer file sharing systems. J.UCS: Journal of Universal Computer Science, 18(8):1032– ??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_8/automatic_
tag_attachment_scheme.

Quintana:2014:TRT

[QFB^{+}14] M. G. Badilla Quintana, M. À. Prats Fernández, M. Careaga Butter, J. C. Gacitúa, and C. Vásquez Carillo. Technological readiness for teaching practices in immersive learning environments open sim. J.UCS: Journal of Universal Computer Science, 20(15): 2024–??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
20_15/technological_
readiness_for_teaching.

Quintero:2014:CAM

[QGT^{+}14] R. Quintero, G. Guzman, M. Torres, R. Menchaca-Mendez, M. Moreno-Ibarra, and F. Mata. A compression algorithm for managing digital elevation models in mobile devices. J.UCS: Journal of Universal Computer Science, 20(10):1433–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        20_10/a_compression_
                                        algorithm_for.

Queiros:2012:OLS

[QL12] R. Queirós and J. P. Leal. Orchestration of E-learning services for automatic evaluation of programming exercises. J.UCS: Journal of Universal Computer Science, 18(11):1454–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_18_
                                        11/orchestration_of_
                                        elearning_services.

Queiros:2013:ELF

[QL13] R. Queirós and J. P. Leal. Ensemble — an e-learning framework. J.UCS: Journal of Universal Computer Science, 19(14): 2127–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_19_14/ensemble_
                                        an_elearning_framework.

Qazi:2011:AEO

[QQ11] N. Ikram Qazi and M. Abdul Qadir. Algorithms for the evaluation of ontologies for extended error taxonomy and their application on large ontologies. J.UCS: Journal of Universal Computer Science, 17(7):1005– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
7/algorithms_for_the_
evaluation.

Qin:2007:VEE

[QZ07] M. Qin and R. Zimmermann. VCA: An energy-efficient votingbased clustering algorithm for sensor networks. J.UCS: Journal of Universal Computer Science, 13(1):87–109, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_13_1/vca_an_
energy_efficient.

Queins:2000:LCC

[QZB^{+}00] S. Queins, G. Zimmermann, M. Becker, M. Kronenburg, Ch. Peper, R. Merz, and J. Schäfer. The light control case study: Problem description. J.UCS: Journal of Universal Computer Science, 6(7):586– 596, July 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_6_
                                        7/light_control_problem_
                                        description.

Qian:2011:SOM

[QZYL11] Z. Qian, S. Zhang, K. Yim, and S. Lu. Service oriented multimedia delivery system in pervasive environments. J.UCS: Journal of Universal Computer Science, 17(6):961– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_17_
                                        6/service_oriented_multimedia_
                                        delivery.

Rothe:2006:CCM

[RA06] J. Rothe and H. Arimura. Computational challenges of massive data sets and randomness in computation. J.UCS Special Issue on the First and Second Japanese-German Frontiers of Science Symposia. J.UCS: Journal of Universal Computer Science, 12(6):579–580, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12;
                                         http://www.jucs.org/
                                        jucs_12_6#;   http://
                                        www.jucs.org/jucs_12_

6/computational_challenges_
of_massive_data.

Rosa:2010:BPC

[RAC10] J. L. Garcia Rosa and J. M. Adan-Coello. Biologically plausible connectionist prediction of natural language thematic relations. J.UCS: Journal of Universal Computer Science, 16(21):3245–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
21/biologically_plausible_
connectionist_prediction.

Rada:1996:TWA

[Rad96] R. Rada. Teaching on the WWW: Assignment focus and information indexing. J.UCS: Journal of Universal Computer Science, 2(10):732–743, October 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
2_10/teaching_on_the_
www.

Radermacher:2001:KMS

[Rad01] F. J. Radermacher. Knowledge management in superorganisms. J.UCS: Journal of Universal Computer Science, 7(6): 507–516, June 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_7_
                                        6/knowledge_management_
                                        in_superorganisms.

Rady:2014:GES

[Rad14] M. Rady. Generating an excerpt of a service level agreement from a formal definition of nonfunctional aspects using OWL. J.UCS: Journal of Universal Computer Science, 20(3):366–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_20_3/generating_
                                        an_excerpt_of.

Rahonis:1999:STI

[Rah99] G. Rahonis. Splicing on trees: the iterated case. J.UCS: Journal of Universal Computer Science, 5(9):599–609, September 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_5_
                                        9/splicing_on_trees_
                                        the.

Rajamani:2007:SMT

[Raj07] S. K. Rajamani. Software is more than code. J.UCS: Journal of Universal Computer Science, 13(5):602–606, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_5/software_is_more_
than.

Ramhorst:2001:GTT

[Ram01] D. Ramhorst. A guided tour through the Siemens Business Services Knowledge Management Framework. J.UCS: Journal of Universal Computer Science, 7(7):610– 622, July 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
7/a_guided_tour_through.

Rico:2015:LLT

[RAS15] M. Rico, J. E. Agudo, and H. Sánchez. Language learning through handheld gaming: a case study of an English course with engineering students. J.UCS: Journal of Universal Computer Science, 21(10):1362–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
10/language_learning_
through_handheld.

Ratschan:2000:UPH

[Rat00] S. Ratschan. Uncertainty propagation in heterogeneous algebras for approximate quantified constraint solving. J.UCS:

Journal of Universal Computer Science, 6(9): 861–880, September 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_6_9/uncertainty_
                                        propagation_in_heterogeneous.

Rathjen:2005:CST

[Rat05]  M. Rathjen.  Con-

structive set theory and Brouwerian principles. J.UCS: Journal of Universal Computer Science, 11(12):2008–2033, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_11_
                                        12/constructive_set_
                                        theory_and.

Ras:2005:UWK

[RAWW05] E. Ras, G. Avram, P. Waterson, and S. Weibelzahl. Using Weblogs for knowledge sharing and learning in information spaces. J.UCS: Journal of Universal Computer Science, 11(3):394–409, March 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_11_3/using_
                                        weblogs_for_knowledge.

Rodrigues:2006:PSC

[RB06] N. F. Rodrigues and L. S. Barbosa. Program slicing by calculation. J.UCS: Journal of Universal Computer Science, 12(7):828– 848, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
7/program_slicing_by_
calculation.

Rodrigues:2007:HOL

[RB07] N. F. Rodrigues and L. S. Barbosa. Higherorder lazy functional slicing. J.UCS: Journal of Universal Computer Science, 13(6):854–873, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
6/higher_order_lazy_
functional.

Randolph:2008:PBC

[RB08] J. J. Randolph and R. Bednarik. Publication bias in the computer science education research literature. J.UCS: Journal of Universal Computer Science, 14(4):575– 589, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
4/publication_bias_in_
the.

Ribeiro:2006:GPA

[RBB06] P. R. Ribeiro, M. A. Barbosa, and L. S. Barbosa. Generic process algebra: a programming challenge. J.UCS: Journal of Universal Computer Science, 12(7):922–937, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        12_7/generic_process_
                                        algebra_a.

Reich:2002:OBS

[RBLR02] J. R. Reich, P. Brockhausen, T. Lau, and U. Reimer. Ontologybased skills management: Goals, opportunities and challenges. J.UCS: Journal of Universal Computer Science, 8(5):506– 515, May 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        5/onlogogy_based_skills_
                                        management.

Rodrigues:2007:PSD

[RC07] R. M. M. Rodrigues and J. M. P. Cardoso. On pipelining sequences of data-dependent loops. J.UCS: Journal of Universal Computer Science, 13(3):419–439, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_3/on_pipelining_sequences_
of.

Rifon:2010:MTT

[RC10] L. Anido Rifón and F. Corradini. Methodologies, technologies and tools enabling e-government. J.UCS: Journal of Universal Computer Science, 16(8):1055–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16;   http:
//www.jucs.org/jucs_
16_08/methodologies_
technologies_and_tools;
http://www.jucs.org/
jucs_16_8#.

Rius:2013:SPE

[RCGBS13] A. Rius, J. Conesa, E. García-Barriocanal, and M.-A. Sicília. Specifying patterns of educational settings by means of ontologies. J.UCS: Journal of Universal Computer Science, 19(3):353– ??, ???? 2013. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_3/specifying_
patterns_of_educational.

Rios:2011:OFM

[RdKO11] I. Rios, A. de Souza Britto Jr., A. Lameiras Koerich, and L. E. Soares Oliveira. An OCR free method for word spotting in printed documents: the evaluation of different feature sets. J.UCS: Journal of Universal Computer Science, 17(1):48– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
1/an_ocr_free_method.

Rubira:2008:SCA

[RdL08] C. M. F. Rubira and R. de Lemos. Software components, architectures and reuse. J.UCS: Journal of Universal Computer Science, 14(8):1179– 1181, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14;
 http://www.jucs.org/
jucs_14_8#;   http://
www.jucs.org/jucs_14_
8/software_components_
architectures_and.

Rebelsky:1996:EIW

[Reb96] S. A. Rebelsky. Evaluating and improving WWW-aided instruction. J.UCS: Journal of Universal Computer Science, 2(12):829–841, December 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
evaluating_and_improving_
www_aided_instruction.

Retalis:2008:CAL

[Ret08a]  S. Retalis.  Creating

adaptive e-learning board games for school settings using the ELG environment. J.UCS: Journal of Universal Computer Science, 14(17):2897–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
17/creating_adaptive_
elearning_board.

Rettinger:2008:BCC

[Ret08b] Robert Rettinger. Bloch’s constant is computable. J.UCS: Journal of Universal Computer Science, 14(6):896–907, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_6/blochs_constant_
is_computable.

Rex:1998:CDS

[Rex98]  G. Rex.  Component-

wise distance to singularity. J.UCS: Journal of Universal Computer Science, 4(1):68–81, January 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_1/componentwise_
distance_to_singularity.

Rega:2012:MPS

[RF12] I. Rega and F. Fanni. Measuring primary schools teachers’ perception of ICT through self-efficacy: a case study. J.UCS: Journal of Universal Computer Science, 18 (3):410–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_3/measuring_
primary_schools_teachers.

Rosado:2010:DSM

[RFMLP10] D. G. Rosado, E. Fernández-Medina, J. López, and M. Piattini. Developing a secure mobile grid system through a UML extension. J.UCS: Journal of Universal Computer Science, 16(17):2333–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
17/developing_a_secure_
mobile.

Rosu:2000:ECI

[RG00] G. Rosu and J. Goguen. On equational Craig interpolation. J.UCS: Journal of Universal Computer Science, 6(1):194– 200, January 28, 2000.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
1/on_equational_craig_
interpolation.

Rosner:1997:NLD

[RGHH97] D. Rösner, B. Grote, K. Hartmann, and B. Höfling. From natural language documents to sharable product knowledge: a knowledge engineering approach. J.UCS: Journal of Universal Computer Science, 3(8):955– 987, August 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs_3_8/from_natural_
language_documents;
internal&sk=05460486.

Reynoso:2010:RES

[RGP10] L. Reynoso, M. Genero, and M. Piattini. Refinement and extension of SMDM, a method for defining valid measures. J.UCS: Journal of Universal Computer Science, 16(21):3210–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_21/refinement_and_
extension_of_smdm.

Rizzardini:2015:MOO

[RGPK15] R. Hernández Rizzardini, F. J. García-Peñalvo, and C. Delgado Kloos. Massive open online courses: Combining methodologies and architecture for a success learning. J.UCS: Journal of Universal Computer Science, 21(5):636–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_21;   http:
                                        //www.jucs.org/jucs_
                                        21_5/massive_open_online_
                                        courses.

Rodriguez:2015:RSN

[RGRR15] A. Cañas Rodriguez, J. M. Santos Gago, L. E. Anido Rifón, and R. Pérez Rodríguez. A recommender system for non-traditional educational resources: A semantic approach. J.UCS: Journal of Universal Computer Science, 21(2): 306–??, ???? 2015. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        21_2/a_recommender_system_
                                        for.

Rademaker:2010:PPT

[RH10] A. Rademaker and E. H. Haeusler. Providing a proof-theoretical basis for explanation: a case study on UML and ALCQI reasoning. J.UCS: Journal of Universal Computer Science, 16(20):3016–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
20/providing_a_proof_
theoretical.

Rajaei:2015:AAS

[RHM15] M. Rajaei, M. S. Haghjoo, and E. K. Miyaneh. An anonymization algorithm for (α, β, γ, δ)-social network privacy considering data utility. J.UCS: Journal of Universal Computer Science, 21(2):268– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
2/an_anonymization_algorithm_
for.

Rhodes:2010:ULI

[Rho10]  C. Rhodes.  Using

Lisp implementation internals unportable but fun. J.UCS: Journal of Universal Computer Science, 16(2):315–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16_2/using_

                                        lisp_implementation_
                                        internals.

Richman:2005:CAM

[Ric05] F. Richman. Constructive aspects of Markov chains. J.UCS: Journal of Universal Computer Science, 11(12):2046–2055, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_11_
                                        12/constructive_aspects_
                                        of_markov.

Riekert:2002:ARI

[Rie02] W.-F. Riekert. Automated retrieval of information in the Internet by using thesauri and gazetteers as knowledge sources. J.UCS: Journal of Universal Computer Science, 8(6):581– 590, June 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        6/automated_retrieval_
                                        of_information.

Rihm:1998:IME

[Rih98] R. Rihm. Implicit methods for enclosing solutions of ODEs. J.UCS: Journal of Universal Computer Science, 4(2):202– 209, February 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_2/implicit_
methods_for_enclosing.

Rizzardini:2015:CIS

[Riz15] R. Hernández Rizzardini. Cloud interoperability service architecture for education environments. J.UCS: Journal of Universal Computer Science, 21(5):656–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_21_5/cloud_
interoperability_service_
architecture.

Rovan:2010:IPW

[RJB10] L. Rovan, T. Jagust, and M. Baranović. Integrating personal Web data through semantically enhanced Web portal. J.UCS: Journal of Universal Computer Science, 16(21):3278–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
21/integrating_personal_
web_data.

Rederlechner:1997:NCP

[RK97] B. Rederlechner and J. Keller. A note on correctness proofs for overflow detection logic in adders for dth complement numbers. J.UCS: Journal of Universal Computer Science, 3(10):1121–1125, October 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_3_10/
a_note_on_correctness.

Read:2015:RMA

[RKH15] T. Read and A. Kukulska-Hulme. The role of a mobile app for listening comprehension training in distance learning to sustain student motivation. J.UCS: Journal of Universal Computer Science, 21(10):1327–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
10/the_role_of_a.

Radhakrishna:2016:NST

[RKJ16] V. Radhakrishna, P. V. Kumar, and V. Janaki. A novel similar temporal system call pattern mining for efficient intrusion detection. J.UCS: Journal of Universal Computer Science, 22(4):475– ??, ???? 2016. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_

22_4/a_novel_similar_
temporal.

Ruan:2010:MCG

[RLL^{+}10] D. Ruan, J. Lu, E. Laes, G. Zhang, J. Ma, and G. Meskens. Multicriteria group decision support with linguistic variables in long-term scenarios for Belgian energy policy. J.UCS: Journal of Universal Computer Science, 16(1):103– ??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_1/multi_criteria_
group_decision.

Rizzardini:2013:CSI

[RLMS13] R. Hernandez Rizzardini, B. H. Linares, A. Mikroyannidis, and H.-C. Schmitz. Cloud services, interoperability and analytics within a ROLE-enabled personal learning environment. J.UCS: Journal of Universal Computer Science, 19(14):2054–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_14/cloud_services_
interoperability_and.

Rebelsky:1998:RVD

[RMF^{+}98] S. A. Rebelsky, F. Makedon, J. Ford, Ch. Owen, P. A. Gloor, and P. T. Metaxas. The roles of video in the design, use, and construction of interactive electronic conference proceedings. J.UCS: Journal of Universal Computer Science, 4(6):604– 628, June 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_6/the_
                                        roles_of_video;   internal&
                                        sk=0C220489.

Rebollo:2012:SRI

[RMFM12] O. Rebollo, D. Mellado, and E. Fernández-Medina. A systematic review of information security governance frameworks in the Cloud computing environment. J.UCS: Journal of Universal Computer Science, 18(6):798– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_18_
                                        6/a_systematic_review_
                                        of.

Rosado-Munoz:2008:ICG

[RMGCGCF08] A. Rosado-Muñoz, L. Gomez-Chova, L. Gomez-Chova, and J. Vila Francés. An IP Core and GUI for implementing multilayer perceptron with a fuzzy activation function on configurable logic devices. J.UCS: Journal of Universal Computer Science, 14(10):1678–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
10/an_ip_core_and.

Rodriguez:2009:HAA

[RMGT09] G. Rodríguez, M. J.

Martín, P. González, and J. Touriño. A heuristic approach for the automatic insertion of checkpoints in message-passing codes. J.UCS: Journal of Universal Computer Science, 15(14):2894–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
14/a_heuristic_approach_
for.

Rubbo:2008:IAR

[RMM^{+}08] F. Barden Rubbo, R. Machado, Á. Freitas Moreira, L. Ribeiro, and D. J. Nunes. On the interaction of advices and raw types in AspectJ. J.UCS: Journal of Universal Computer Science, 14(21):3534–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        21/on_the_interaction_
                                        of.

Ruiz-Martinez:2009:NFN

[RMMLBLGS09] A. Ruiz-Martínez, C. I. Marín-López, L. Baño-López, and A. F. Gómez-Skarmeta. A new fair non-repudiation protocol for secure negotiation and contract signing. J.UCS: Journal of Universal Computer Science, 15(3):555–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_15_3/a_new_fair_
                                        nonrepudiation.

Rupnik:2015:PDE

[RMZ15] B. Rupnik, D. Mongus, and B. Zalik. Point density evaluation of airborne LiDAR datasets. J.UCS: Journal of Universal Computer Science, 21(4):587–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_21_4/point_
                                        density_evaluation_of.

Reinhardt:2003:TTI

[RN03] K. Reinhardt and K. North. Transparency and transfer of individual competencies — a concept of integrative competence management. J.UCS: Journal of Universal Computer Science, 9 (12):1372–1380, December 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
12/transparency_and_
transfer_of.

Roberts:2006:PHM

[Rob06] J. Roberts. Pervasive health management and health management utilizing pervasive technologies: Synergy and issues. J.UCS: Journal of Universal Computer Science, 12(1):6–14, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_12_1/pervasive_
health_management_and.

Rojas:1996:CBN

[Roj96] R. Rojas. Conditional branching is not necessary for universal computation in von Neumann computers. J.UCS: Journal of Universal Computer Science, 2(11):756– 768, November 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_

                                        11/conditional_branching_
                                        is_not.

Rosu:1999:KEI

[Ros99]  G. Rosu.  Kan ex-

tensions of institutions. J.UCS: Journal of Universal Computer Science, 5(8):482–492, August 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_5_
                                        8/kan_extensions_of_
                                        institutions.

Rosaci:2005:EAO

[Ros05]  D. Rosaci.  Exploit-

ing agent ontologies in B2C virtual marketplaces. J.UCS: Journal of Universal Computer Science, 11(6):1011–1039, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_11_6/exploiting_
                                        agent_ontologies_in.

Ruokamo:1998:PPE

[RP98] H. Ruokamo and S. Pohjolainen. Pedagogical principles for evaluation of hypermedia-based learning environments in mathematics. J.UCS: Journal of Universal Computer Science, 4(3):292– 307, March 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_3/pedagogical_
principles_for_evaluation;
internal&sk=0C220489.

Ramos:2008:ESR

[RP08] M. A. Ramos and R. A. D.

Penteado. Embedded software revitalization through component mining and software product line techniques. J.UCS: Journal of Universal Computer Science, 14(8): 1211–1227, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_8/embedded_software_
revitalization_through.

Roman:2015:SBA

[RPCA15] N. Trevisan Roman, P. Piwek, A. M. Brito Rizzoni Carvalho, and A. Rossi Alvares. Sentiment and behaviour annotation in a corpus of dialogue summaries. J.UCS: Journal of Universal Computer Science, 21(4):561–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_4/sentiment_
and_behaviour_annotation.

Rieder:2011:PNB

[RPR11] R. Rieder, M. S. Pinho, and A. B. Raposo. A Petri nets based approach to specify individual and collaborative interaction in 3D virtual environments. J.UCS: Journal of Universal Computer Science, 17(2):243– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
2/a_petri_nets_based.

Riege:2006:CBH

[RR06a] T. Riege and J. Rothe. Completeness in the Boolean hierarchy: Exact-four-colorability, minimal graph uncolorability, and exact domatic number problems — a survey. J.UCS: Journal of Universal Computer Science, 12(5):551– 578, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
5/completenes_in_the_
boolean.

Riege:2006:IDR

[RR06b] T. Riege and J. Rothe. Improving deterministic and randomized exponentialtime algorithms for the satisfiability, the colorability, and the domatic number problem. J.UCS: Journal of Universal Computer Science, 12(6):725–745, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_12_6/improving_
deterministic_and_randomized.

Rechberger:2008:NRN

[RR08] C. Rechberger and V. Rijmen. New results on NMAC/HMAC when instantiated with popular hash functions. J.UCS: Journal of Universal Computer Science, 14(3): 347–376, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_3/new_results_on_
nmac.

Rampalli:2011:FCO

[RR11] R. Rampalli and A. G.

Ramakrishnan. Fusion of complementary online and offline strategies for recognition of handwritten Kannada characters. J.UCS: Journal of Universal Computer Science, 17(1):81–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_1/fusion_of_
complementary_online.

Ramalho:2003:XXL

[RRB03] F. Ramalho, J. Robin, and R. Barros. XOCL

— an XML language for specifying logical constraints in object oriented models. J.UCS: Journal of Universal Computer Science, 9(8):956– 969, August 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_9_8/xocl_an_xml_
language.

Rodriguez:2012:TPS

[RRM^{+}12] M. García Rodríguez, J. M. Alvarez Rodríguez, D. Berrueta Muñoz, L. Polo Paredes, J. E. Labra Gayo, and P. Ordoñez De Pablos. Towards a practical solution for data grounding in a Semantic Web services environment. J.UCS: Journal of Universal Computer Science, 18(11):1576–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_11/towards_a_
practical_solution.

Rabanal:2010:ARC

[RRR10] P. Rabanal, I. Rodríguez, and F. Rubio. Applying RFD to construct optimal quality-investment trees. J.UCS: Journal of Universal Computer Science, 16(14):1882–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/

jucs_16_14/applying_
rfd_to_construct.

Roberts:2000:PFB

[RS00] A. Roberts and A. Symvonis. Potential-function-based analysis of an offline heap construction algorithm. J.UCS: Journal of Universal Computer Science, 6(2):240– 255, February 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
2/potential_functions_
based_analysis.

Reif:2001:JUSa

[RS01a] W. Reif and G. Schellhorn. J.UCS special issue on tools for system design and verification — Part 1. J.UCS: Journal of Universal Computer Science, 7(1):1–2, January 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
7_1/tools_for_system_
design.

Reif:2001:JUSb

[RS01b] W. Reif and G. Schellhorn. J.UCS special issue on tools for system design and verification — Part 2. J.UCS: Journal of Universal Computer Science, 7(2):105– 106, February 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        7_2/tools_for_system_
                                        design.

Reinhardt:2002:FKC

[RS02] R. Reinhardt and B. Stattkus. Fostering knowledge communication: Concept and implementation. J.UCS: Journal of Universal Computer Science, 8(5):536– 545, May 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        5/fostering_knowledge_
                                        communication_concept.

Reimer:2003:PKM

[RS03] U. Reimer and Y. Sure. Professional knowledge management — experiences and visions — J.UCS special issue. J.UCS: Journal of Universal Computer Science, 9(7): 579–580, July 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/free;
                                         http://www.jucs.org/
                                        jucs_9_7/professional_
                                        knowledge_management_
                                        experiences.

Reeves:2005:CPP

[RS05] S. Reeves and D. Streader.

Constructing programs or processes. J.UCS: Journal of Universal Computer Science, 11(12):2034– 2045, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
12/constructing_programs_
or_processes.

Retalis:2011:VEC

[RS11] S. Retalis and P. B.

Sloep. Virtual environments for collaborative innovation and learning. J.UCS: Journal of Universal Computer Science, 17(12):1634–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_17;   http:
//www.jucs.org/jucs_
17_12/virtual_environments_
for_collaborative;
http://www.jucs.org/
jucsrssfeed-issue.

Rosado:2012:SIS

[RSFMJ12] D. G. Rosado, L. E.

Sánchez, E. Fernández-Medina, and J. Jürjens. Security in information systems: New challenges and opportunities. J.UCS: Journal of Universal Computer Science, 18(6):728–??, ???? 2012.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_18;   http:
//www.jucs.org/jucs_
18_6/security_in_information_
systems;   http://www.
jucs.org/jucsrssfeed-
issue.

Rizzardini:2013:CEE

[RSM^{+}13] R. Hernandez Rizzardini, H.-C. Schmitz, A. Mikroyannidis, C. Delgado Kloos, and L. Chao. Cloud education environment. J.UCS: Journal of Universal Computer Science, 19(14):2034–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_19;   http:
//www.jucs.org/jucs_
19_14/cloud_education_
environment.

Rijo:2014:DSS

[RSP^{+}14] R. Rijo, C. Silva, L. Pereira, D. Gonçalves, and M. Agostinho. Decision support system to diagnosis and classification of epilepsy in children. J.UCS: Journal of Universal Computer Science, 20(6):907– ??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
6/decision_support_system_
to.

Reif:2001:CER

[RSVR01] W. Reif, G. Schellhorn, T. Vollmer, and J. Ruf. Correctness of efficient real-time model checking. J.UCS: Journal of Universal Computer Science, 7(2):194– 209, February 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_7_2/correctness_
of_efficient_real.

Rinner:2004:RPE

[RSW04] B. Rinner, M. Schmid, and R. Weiss. A rapid prototyping environment for multi-DSP systems based on accurate performance prediction. J.UCS: Journal of Universal Computer Science, 10(2):120–144, February 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
2/a_rapid_prototyping_
environment.

Rastocny:2013:WSR

[RTB13] K. Rástocný, M. Tvarozek, and M. Bielikova. Web search results exploration via cluster-based viewes and zoom-based navigation. J.UCS: Journal of Universal Computer Science, 19(15):2320–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
15/web_search_results_
exploration.

Rothe:2001:CCS

[RTJ01] J. Rothe, H. Tews, and B. Jacobs. The coalgebraic class specification language CCSL. J.UCS: Journal of Universal Computer Science, 7(2):175–193, February 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
7_2/the_coalgebraic_
class_specification.

RauberDuBois:2005:MMC

[RTL05] A. Rauber Du Bois, P. Trinder, and H.-W. Loidl. mHaskell: Mobile computation in a purely functional language. J.UCS: Journal of Universal Computer Science, 11(7):1234–1254, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
7/....

RauberDuBois:2006:SMM

[RTL06] A. Rauber Du Bois, P. Trinder, and H.-W. Loidl. Strong mobility in Mobile Haskell. J.UCS: Journal of Universal Computer Science, 12(7):868–884, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_7/strong_mobility_
in_mobile.

Rudeanu:2004:DBF

[Rud04] S. Rudeanu. On the decomposition of Boolean functions via Boolean equations. J.UCS: Journal of Universal Computer Science, 10(9): 1294–1301, September 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_10_9/on_the_
decomposition_of.

Ruf:2001:RRT

[Ruf01] J. Ruf. RAVEN: Realtime analyzing and verification environment. J.UCS: Journal of Universal Computer Science, 7(1):89–104, January 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_

1/real_time_analyzing_
and.

Rushby:2007:AFM

[Rus07]  J. Rushby.  Auto-

mated formal methods enter the mainstream. J.UCS: Journal of Universal Computer Science, 13(5):650–660, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_13_5/automated_
formal_methods_enter.

Roa:2012:BAC

[RVC12] J. Roa, P. Villarreal, and O. Chiotti. Behavior alignment and control flow verification of process and service choreographies. J.UCS: Journal of Universal Computer Science, 18(17):2383–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_17/behavior_
alignment_and_control.

Rastogi:2012:ISS

[RvS12] R. Rastogi and R. von Solms. Information security service culture — information security for end-users. J.UCS: Journal of Universal Computer Science, 18(12): 1628–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_12/information_security_
service_culture.

Robin:2007:CPA

[RVW07] J. Robin, J. Vitorino, and A. Wolf. Constraint programming architectures: Review and a new proposal. J.UCS: Journal of Universal Computer Science, 13(6):701–720, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_13_6/constraint_
programming_architectures_
review.

Rettinger:2009:TCB

[RWZ09] R. Rettinger, K. Weihrauch, and N. Zhong. Topological complexity of blowup problems. J.UCS: Journal of Universal Computer Science, 15(6):1301– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
6/topological_complexity_
of_blowup.

Ruan:2009:MIC

[RY09] C. Ruan and S.-S. Yeo.

Modeling of an intelligent e-consent system in a healthcare domain. J.UCS: Journal of Universal Computer Science, 15(12):2429–??, ????

2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_15_12/modeling_
                                        of_an_intelligent.

Schellhorn:1997:RAA

[SA97] Gerhard Schellhorn and Wolfgang Ahrendt. Reasoning about abstract state machines: The WAM case study. J.UCS: Journal of Universal Computer Science, 3(4): 377–413, April 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs_3_4/wam_case;
                                         http://www.jucs.org/
                                        jucs_3_4/reasoning_about_
                                        abstract_state;   internal&
                                        sk=05460486.

Stefanutti:2003:SAP

[SA03] L. Stefanutti and D. Albert. Skill assessment in problem solving and simulated learning environments. J.UCS: Journal of Universal Computer Science, 9(12):1455–1468, December 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        12/skill_assessment_
                                        in_problem.

Silva:2009:DWA

[SA09] R. Silva and A. Andrade.

Development of a Web application for management of learning styles. J.UCS: Journal of Universal Computer Science, 15(7):1508–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_7/development_of_
a_web.

Salguero:2010:ISE

[SA10] A. Salguero and F. Araque.

Integration of similar evolving data sources for supporting decision making tasks. J.UCS: Journal of Universal Computer Science, 16(1):22– ??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_1/integration_of_
similar_evolving.

Solis:2011:SWB

[SA11] C. Solis and N. Ali.

A Semantic Wiki based on spatial hypertext. J.UCS: Journal of Universal Computer Science, 17(7):1043–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_7/a_semantic_wiki_
based.

Shahzad:2014:DCW

[SA14] B. Shahzad and E. Alwagait. Does a change in weekend days have an impact on social networking activity? J.UCS: Journal of Universal Computer Science, 20(15):2068–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_20_
                                        15/does_a_change_in.

Santos:2008:ISB

[SAA08] R. Santos, R. Azevedo, and G. Araujo. Instruction scheduling based on subgraph isomorphism for a high performance computer processor. J.UCS: Journal of Universal Computer Science, 14(21):3465–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        21/instruction_scheduling_
                                        based_on.

Stenz:1999:PTS

[SAB99] G. Stenz, W. Ahrendt, and B. Beckert. Proof transformations from searchoriented into interactionoriented tableau calculi. J.UCS: Journal of Universal Computer Science, 5(3):113–134, March 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
3/proof_trasformations_
from_search.

Safar:2008:SQR

[Saf08] M. Safar. Spatial queries in road networks based on PINE. J.UCS: Journal of Universal Computer Science, 14(4):590– 611, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
4/spatial_queries_in_
road.

Salah:2011:PES

[SAKAM11] K. Salah, M.-A.-R. Al-Khiaty, R. Ahmed, and A. Mahdi. Performance evaluation of Snort under Windows 7 and Windows Server 2008. J.UCS: Journal of Universal Computer Science, 17(11): 1605–??, ???? 2011. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_11/performance_evaluation_
of_snort.

Salomaa:2002:GCS

[Sal02]  A. Salomaa.  Genera-

tion of constants and synchronization of finite automata. J.UCS: Journal of Universal Computer Science, 8(2):332– 347, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_8_2/generation_
                                        of_constants_and.

Salomaa:2010:OCC

[Sal10] A. Salomaa. Ordered catenation closures and decompositions of languages related to a language of Derick Wood. J.UCS: Journal of Universal Computer Science, 16(5):821–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        16_5/ordered_catenation_
                                        closures_and.

Sarkar:2005:DEU

[Sar05] P. Sarkar. Domain extenders for UOWHF: a finite binary tree algorithm. J.UCS: Journal of Universal Computer Science, 11(6):1040–1053, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_11_6/domain_
                                        extenders_for_uowhf.

Satoh:2010:MAB

[Sat10] I. Satoh. Mobile agentbased context-aware services. J.UCS: Journal of Universal Computer Science, 16(15):1929–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
15/mobile_agent_based_
context.

Sapateiro:2011:DMC

[SBAZ11] C. Sapateiro, N. Baloian, P. Antunes, and G. Zurita. Developing a mobile collaborative tool for business continuity management. J.UCS: Journal of Universal Computer Science, 17(2):164–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_2/developing_
a_mobile_collaborative.

Sun:2015:JDC

[SBCD15] P. Sun, P. Bon, and S. Collart-Dutilleul. A joint development of coloured Petri nets and the B method in critical systems. J.UCS: Journal of Universal Computer Science, 21(12):1654–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
12/a_joint_development_
of.

Shashidhar:2003:AVT

[SBCJ03] K. C. Shashidhar, M. Bruynooghe, F. Catthoor, and G. Janssens. An automatic verification technique for loop and data reuse transformations based on geometric modeling of programs. J.UCS: Journal of Universal Computer Science, 9(3):248– 269, March 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
3/an_automatic_verification_
technique.

Sloep:2012:EIL

[SBG^{+}12] P. B. Sloep, A. J. Berlanga, W. Greller, S. Stoyanov, M. van der Klink, S. Retalis, and J. Hensgens. Educational innovation with learning networks: Tools and developments. J.UCS: Journal of Universal Computer Science, 18(1):44– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
1/educational_innovation_
with_learning.

Sabucedo:2014:ASA

[SBGI14] L. M. Álvarez Sabucedo, R. Soto Barreiros, J. M. Santos Gago, and M. J. Fernández Iglesias. An adaptive and social-aware recommendation algorithm for administration services. J.UCS: Journal of Universal Computer Science, 20(11):1523–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
11/an_adaptive_and_social.

Schirru:2010:ECU

[SBMD10] R. Schirru, S. Baumann, M. Memmel, and A. Dengel. Extraction of contextualized user interest profiles in social sharing platforms. J.UCS: Journal of Universal Computer Science, 16(16): 2196–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_16/extraction_of_
contextualized_user.

Saneifar:2015:TET

[SBPR15] H. Saneifar, S. Bonniol, P. Poncelet, and M. Roche. From terminology extraction to terminology validation: An approach adapted to log files. J.UCS: Journal of Universal Computer Science, 21(4):604– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_

                                        4/from_terminology_extraction_
                                        to.

Sie:2011:WIM

[SBRS11] R. L. Sie, M. Bitter-Rijpkema, and P. B. Sloep. What’s in it for me? Recommendation of peers in networked innovation. J.UCS: Journal of Universal Computer Science, 17(12):1659–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_17_
                                        12/whats_in_it_for.

Scullion:2015:UES

[SBS15] J. Scullion, G. Baxter, and M. Stansfield. UNITE: Enhancing students’ selfefficacy through the use of a 3D virtual world. J.UCS: Journal of Universal Computer Science, 21(12):1635–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_21_12/unite_
                                        enhancing_students_self.

Sessink:2004:SWB

[SBTH04] O. Sessink, R. Beeftink, J. Tramper, and R. Hartog. Securing Webbased exams. J.UCS: Journal of Universal Computer Science, 10 (2):145–157, February 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
2/securing_web_based_
exams.

Shen:2014:WSN

[SC14] C.-W. Shen and S.-H.

Chu. Web 2.0 and social networking services in municipal emergency management: A study of U.S. cities. J.UCS: Journal of Universal Computer Science, 20(15):1995–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
15/web_2.0_and_social.

Schmaranz:1996:PEP

[Sch96]  K. Schmaranz.  Pro-

fessional electronic publishing in Hyper-G: The next generation publishing solution on the Web. J.UCS: Journal of Universal Computer Science, 2(9):650–658, September 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
9/professional_electronic_
publishing_in.

Schneeweiss:1999:AFT

[Sch99] W. G. Schneeweiss. Advanced fault tree modeling. J.UCS: Journal of Universal Computer Science, 5(10):633–643, October 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        5_10/advanced_fault_
                                        tree_modeling.

Schellhorn:2001:VAR

[Sch01a] G. Schellhorn. Verification of ASM refinements using generalized forward simulation. J.UCS: Journal of Universal Computer Science, 7(11):952– 979, November 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        7_11/verification_of_
                                        asm_refinements.

Schinagl:2001:NLA

[Sch01b] W. Schinagl. New learning of adults in the information and knowledge society. J.UCS: Journal of Universal Computer Science, 7(7):623– 628, July 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        7_7/new_learning_of_
                                        adults.

Schmid:2001:CAS

[Sch01c] J. Schmid. Compiling abstract state machines to C++. J.UCS: Journal of Universal Computer Science, 7(11):1068– 1087, November 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_7_11/compiling_
abstract_state_machine.

Schmaranz:2002:SGD

[Sch02a]  K. Schmaranz.  On

second generation distributed component systems. J.UCS: Journal of Universal Computer Science, 8(1):97–116, January 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
1/on_second_generation_
distributed.

Schneider:2002:WES

[Sch02b]  K. Schneider.  What

to expect from software experience exploitation. J.UCS: Journal of Universal Computer Science, 8(6):570–580, June 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_8_6/what_to_
expect_from.

Schneider:2002:KAG

[Sch02c]  U. Schneider.  The

knowledge-attention-gap: Do we underestimate the problem of information overload in knowledge management. J.UCS: Journal of Universal Computer Science, 8(5): 482–490, May 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        5/the_knowledge_attention_
                                        gap.

Schwer:2002:RIG

[Sch02d] S. R. Schwer. Reasoning with intervals on granules. J.UCS: Journal of Universal Computer Science, 8(8):793–807, August 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        8/reasoning_with_intervals_
                                        on.

Schutt:2003:PNK

[Sch03] P. Schütt. The post-Nonaka knowledge management. J.UCS: Journal of Universal Computer Science, 9(6):451– 462, June 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        9_6/the_post_nonaka_
                                        knowledge.

Schewe:2005:FDC

[Sch05a] K.-D. Schewe. Functional dependencies with counting on trees. J.UCS: Journal of Universal Computer Science, 11(12): 2063–2075, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_12/functional_dependencies_
with_counting.

Schuster:2005:WCC

[Sch05b] P. Schuster. What is continuity, constructively? J.UCS: Journal of Universal Computer Science, 11(12):2076–2085, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_11_12/what_
is_continuity_constructively.

Schwichtenberg:2005:DPE

[Sch05c] H. Schwichtenberg. A direct proof of the equivalence between Brouwer’s Fan Theorem and König’s Lemma with a uniqueness hypothesis. J.UCS: Journal of Universal Computer Science, 11(12): 2086–2095, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_12/a_direct_proof_
of.

Schordan:2006:LVD

[Sch06] M. Schordan. The language of the visitor design pattern. J.UCS: Journal of Universal Computer Science, 12(7):849– 867, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12_
                                        7/the_language_of_the.

Schellhorn:2008:ARP

[Sch08a] G. Schellhorn. ASM refinement preserving invariants. J.UCS: Journal of Universal Computer Science, 14(12): 1929–??, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_12/asm_refinement_
                                        preserving_invariants.

Schroder:2008:RBF

[Sch08b] M. Schröder. On the relationship between filter spaces and weak limit spaces. J.UCS: Journal of Universal Computer Science, 14(6):996–1015, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_14_6/on_the_
                                        relationship_between.

Schewe:2009:LAS

[Sch09a] K.-D. Schewe. Logic, abstract state machines and databases. J.UCS: Journal of Universal Computer Science, 15(1):1–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15;
http://www.jucs.org/
jucs_15_1#;   http://
www.jucs.org/jucs_15_
1/logic_abstract_state_
machines.

Schmidt:2009:SDG

[Sch09b] A. U. Schmidt. On the superdistribution of digital goods. J.UCS: Journal of Universal Computer Science, 15(2):401–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_15_2/on_the_
superdistribution_of.

Schroder:2009:ETU

[Sch09c] M. Schröder. An effective Tietze–Urysohn theorem for QCB-spaces. J.UCS: Journal of Universal Computer Science, 15(6):1317–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_6/an_effective_tietze_
urysohn.

Schroder:2010:NCS

[Sch10] M. Schröder. A note on closed subsets in quasizero-dimensional qcb-spaces.

J.UCS: Journal of Universal Computer Science, 16(18):2711–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_18/a_note_on_
closed.

Seo:2009:HTB

[SCK^{+}09] D.-W. Seo, H.-U. Chae, B.-W. Kim, W.-H. Choi, and K.-H. Jo. Human tracking based on multiple view homography. J.UCS: Journal of Universal Computer Science, 15(13):2463–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_15_13/human_
tracking_based_on.

Stubenrauch:2003:AMW

[SCLM03] R. Stubenrauch, D. G. Camhy, J. A. Lennon, and H. Maurer. Applications of MIRACLE: Working with dynamic visual information. J.UCS: Journal of Universal Computer Science, 9(4):349– 368, April 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
9_4/applications_of_
miracle_working.

Santarosa:2013:OCP

[SCS13] L. M. Costi Santarosa, D. Conforto, and F. Chagas Schneider. One computer per student city — total UCA an all inclusive totality under discussion. J.UCS: Journal of Universal Computer Science, 19(11):1597–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
11/one_computer_per_
student.

Suarez-Cabal:2009:SCC

[SCT09] M. J. Suárez-Cabal and J. Tuya. Structural coverage criteria for testing SQL queries. J.UCS: Journal of Universal Computer Science, 15 (3):584–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_3/structural_
coverage_criteria_for.

Soursos:2008:DBP

[SCW08] S. Soursos, C. Courcoubetis, and R. Weber. Dynamic bandwidth pricing: Provision cost, market size, effective bandwidths and price games. J.UCS: Journal of Universal Computer Science, 14(5):766– 785, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
5/dynamic_bandwidth_
pricing_provision.

Simone:1997:ASC

[SD97] C. Simone and M. Divitini. Ariadne: Supporting coordination through a flexible use of the knowledge on work processes. J.UCS: Journal of Universal Computer Science, 3(8):865– 898, August 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs_3_8/ariadne_
supporting_coordination_
through; internal&sk=
05460486.

Sanchez:2013:DNS

[SdBC13] J. Sánchez and M. de Borba Campos. Development of navigation skills through audio haptic videogaming in learners who are blind. J.UCS: Journal of Universal Computer Science, 19(18):2677–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_18/development_of_
navigation_skills.

Sirjani:2005:MVC

[SdBM05] M. Sirjani, F. S. de Boer, and A. Movaghar. Modular verification of a component-based actor language. J.UCS: Journal of Universal Computer Science, 11(10):1695– 1717, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
10/modular_verificatin_
of_a.

Storme:1999:GTA

[SDJ99] L. Storme, A. De Vos, and G. Jacobs. Group theoretical aspects of reversible logic gates. J.UCS: Journal of Universal Computer Science, 5(5):307– 321, May 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
5/group_theoretical_
aspects.

Solar:2014:MGO

[SDLM14] M. Solar, F. Daniels, R. López, and L. Meijueiro. A model to guide the open government data implementation in public agencies. J.UCS: Journal of Universal Computer Science, 20(11):1564–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_20_
                                        11/a_model_to_guide.

Sert:2012:UAM

[SDÖ^{+}12] O. C. Sert, K. Dursun, T. Özyer, J. Jida, and R. Alhajj. The unification and assessment of multi-objective clustering results of categorical datasets with Hconfidence metric. J.UCS: Journal of Universal Computer Science, 18(4): 507–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        18_4/the_unification_
                                        and_assessment.

Santarosa:2009:EVE

[SdOB09] L. M. Costi Santarosa and L. de Oliveira Basso. Eduquito: Virtual environment for digital inclusion of people with special educational needs. J.UCS: Journal of Universal Computer Science, 15(7):1496–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        15_7/eduquito_virtual_
                                        environment_for.

Schmeil:2009:KSC

[SE09] A. Schmeil and M. J. Eppler. Knowledge sharing and collaborative learning in second life: a classification of virtual 3D group interaction scripts. J.UCS: Journal of Universal Computer Science, 15(3):665– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
3/knowledge_sharing_
and_collaborative.

Schmid:2013:DOC

[SEK13] K. Schmid, H. Eichelberger, and C. Kröher. Domain-oriented customization of service platforms: Combining product line engineering and service-oriented computing. J.UCS: Journal of Universal Computer Science, 19(2):233– ??, ???? 2013. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_2/domain_oriented_
customization_of.

Safar:2010:EOS

[SESMT10] M. Safar, N. El-Sayed, K. Mahdi, and D. Taniar. Entropy optimization of social networks using an evolutionary algorithm. J.UCS: Journal of Universal Computer Science, 16(6):983– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_16_
                                        6/entropy_optimization_
                                        of_social.

Smith:2000:IDR

[SF00] G. Smith and C. Fidge. Incremental development of real-time requirements: The light control case study. J.UCS: Journal of Universal Computer Science, 6(7):704– 730, July 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_6_7/incremental_
                                        development_of_real.

Savola:2012:RDS

[SFP12] R. M. Savola, C. Frühwirth, and A. Pietikäinen. Riskdriven security metrics in agile software development — an industrial pilot study. J.UCS: Journal of Universal Computer Science, 18(12):1679–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_18_12/risk_
                                        driven_security_metrics.

Sierra:2004:ADO

[SFVFMN04] J. L. Sierra, A. Fernandez-Valmayor, B. Fernandez-Manjon, and A. Navarro.

ADDS: a documentoriented approach for application development. J.UCS: Journal of Universal Computer Science, 10(9):1302–1324, September 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
10_9/adds_a_document_
oriented.

Shearer:1996:GCR

[SG96] J. Shearer and P. Gutmann. Government, cryptography, and the right to privacy. J.UCS: Journal of Universal Computer Science, 2(3):113– 146, March 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
government_cryptography_
and_the_right_to_privacy.

Schwotzer:2002:SSM

[SG02] T. Schwotzer and K. Geihs.

Shark: a system for management, synchronization and exchange of knowledge in mobile user groups. J.UCS: Journal of Universal Computer Science, 8(6):644– 651, June 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
6/shark_a_system_for.

Sanchez:2013:USQ

[SGB^{+}13] J. L. González Sánchez, R. García, J. M. Brunetti, R. Gil, and J. M. Gimeno. Using SWET–QUM to compare the quality in use of Semantic Web exploration tools. J.UCS: Journal of Universal Computer Science, 19(8):1025– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_19_
                                        8/using_SWET_QUM_to.

Sanchez-Gordon:2016:HCM

[SGLM16] S. Sanchez-Gordon and S. Lujan-Mora. How could MOOCs become accessible? The case of edX and the future of inclusive online learning. J.UCS: Journal of Universal Computer Science, 22(1):55–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        22_1/how_could_moocs_
                                        become.

Stein:2013:CFP

[SGS13] M. Stein and A. Geyer-Schulz. A comparison of five programming languages in a graph clustering scenario. J.UCS: Journal of Universal Computer Science, 19(3):428–??, ???? 2013.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_3/a_comparison_of_
five.

Schwan:1996:CLV

[SH96] S. Schwan and R. Hesse.

Communicating and learning in “virtual seminars”: The uses of spatial metaphors in interface design. J.UCS: Journal of Universal Computer Science, 2(6):503– 512, June 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
communicating_and_learning_
in_virtual_seminars_
the_uses_of_spatial_
methaphors_in_interface_
design.

Sanchez-Hernandez:2006:CFF

[SH06] J. Sánchez-Hernández.

Constructive failure in functional-logic programming: From theory to implementation. J.UCS: Journal of Universal Computer Science, 12 (11):1574–1593, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_11/constructive_failure_
in_functional.

Shakshuki:2009:AWB

[SH09] E. M. Shakshuki and R. Halliday. An agent for Web-based structured hypermedia algorithm explanation system. J.UCS: Journal of Universal Computer Science, 15(10):2078–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15_
                                        10/an_agent_for_web.

Salah:2010:ICB

[SH10] K. Salah and M. Hamawi. Impact of CPU-bound processes on IP forwarding of Linux and Windows XP. J.UCS: Journal of Universal Computer Science, 16(21):3299–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_16_21/impact_of_
                                        cpu_bound.

Smith:2011:DRU

[SH11] D. Smith and R. Harvey. Document retrieval using SIFT image features. J.UCS: Journal of Universal Computer Science, 17(1):3–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_17_

1/document_retrieval_
using_sift.

Shahzad:2011:OBU

[Sha11] S. K. Shahzad. Ontologybased user interface development: User experience elements pattern. J.UCS: Journal of Universal Computer Science, 17(7):1078–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_7/ontology_based_
user_interface.

Shearer:1996:ONO

[She96]  J. Shearer.  One net

one world — global citizenship and the Internet. J.UCS: Journal of Universal Computer Science, 2(12):842, December 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
12/one_net_one_world.

Sheeran:2005:HDF

[She05] M. Sheeran. Hardware design and functional programming: a perfect match. J.UCS: Journal of Universal Computer Science, 11(7):1135– 1158, ???? 2005. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_7/hardware_design_
and_functional.

Sandnes:2010:NEF

[SHH10] F. E. Sandnes, Y.-P. Huang, and Y.-M. Huang. Near eyes-free chauffeur computer interaction with chording and visual text mnemonics. J.UCS: Journal of Universal Computer Science, 16(10): 1311–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_10/near_eyes_free_
chauffeur.

Shim:2011:SAT

[Shi11]  K.-A. Shim.  Secu-

rity analysis of three password authentication schemes. J.UCS: Journal of Universal Computer Science, 17(11):1623–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
11/security_analysis_
of_three.

Scheibe:2010:PER

[SHK10] K. Scheibe, F. Huang, and R. Klette. Pose estimation of rotating sensors in the context of accurate 3D scene modeling. J.UCS: Journal of Universal Computer Science, 16(10):1269–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16_10/pose_
estimation_of_rotating.

Shum:1997:NCR

[Shu97] S. Shum. Negotiating the construction and reconstruction of organisational memories. J.UCS: Journal of Universal Computer Science, 3(8):899– 928, August 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs_3_8/negotiating_
the_construction_and;
internal&sk=05460486.

Shu:2010:CLO

[SHZ^{+}10] L. Shu, M. Hauswirth, Y. Zhang, J. Ma, G. Min, and Y. Wang. Cross layer optimization for data gathering in wireless multimedia sensor networks within expected network lifetime. J.UCS: Journal of Universal Computer Science, 16(10):1343–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16_10/cross_
layer_optimization_for.

Sato:2000:NCS

[SI00] Y. Sato and T. Ikegami.

Nonlinear computation with switching map systems. J.UCS: Journal of Universal Computer Science, 6(9):881– 905, September 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
9/nonlinear_computation_
with_switching.

Shah:2009:QPE

[SIIJ09] N. Shah, R. Iqbal, K. Iqbal, and A. James. A QoS perspective on exception diagnosis in service-oriented computing. J.UCS: Journal of Universal Computer Science, 15(9):1871– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
9/a_qos_perspective_
on.

Simovici:2007:MEP

[Sim07]  D. Simovici.  Metric-

entropy pairs on lattices. J.UCS: Journal of Universal Computer Science, 13(11):1767–1778, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_11/metric_entropy_
pairs_on.

Sinka:2006:PST

[Sin06] R. Sinka. Primary school teachers in the information society. J.UCS: Journal of Universal Computer Science, 12(9):1358– 1372, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_9/primary_school_
teachers_in.

Spasic:2013:MDF

[SJ13] A. Spasić and D. Janković.

Model-driven framework for design and production of low-budget stereoscopic TV content. J.UCS: Journal of Universal Computer Science, 19(1):78–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_19_1/model_
driven_framework_for.

Szemethy:2004:PMM

[SK04] T. Szemethy and G. Karsai. Platform modeling and model transformations for analysis. J.UCS: Journal of Universal Computer Science, 10(10):1383–1407, October 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_

                                        10/platform_modeling_
                                        and_model.

Safran:2008:SFW

[SK08] C. Safran and F. Kappe. Success factors in a Weblog community. J.UCS: Journal of Universal Computer Science, 14(4): 546–556, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_9/success_factors_
                                        in_a.

Sabouri:2013:MVR

[SK13] H. Sabouri and R. Khosravi. Modeling and verification of reconfigurable actor families. J.UCS: Journal of Universal Computer Science, 19(2):207–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_19_
                                        2/modeling_and_verification_
                                        of.

Sait:2008:PSS

[SKA08] S. M. Sait, K. S. Khan, and M. I. Ali. Parallel strategies for stochastic evolution. J.UCS: Journal of Universal Computer Science, 14(15): 2471–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_15/parallel_
strategies_for_stochastic.

Stern:2010:CRA

[SKH^{+}10] H. Stern, R. Kaiser, P. Hofmair, P. Kraker, S. N. Lindstaedt, and P. Scheir. Content recommendation in APOS-DLE using the associative network. J.UCS: Journal of Universal Computer Science, 16(16):2214–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
16/content_recommendation_
in_aposdle.

Suomi:2012:MVE

[SKH12] H. Suomi, K. Kilkki, and H. Hämmäinen. Modeling the value of endto-end multipath protocols. J.UCS: Journal of Universal Computer Science, 18(14):2071–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_14/modeling_
the_value_of.

Soikkeli:2014:CCF

[SKH14] T. Soikkeli, J. Karikoski, and H. Hämmäinen. Context classification framework for handset-based end user studies. J.UCS:

Journal of Universal Computer Science, 20 (14):1964–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
20_14/context_classification_
framework_for.

Slootmaker:2014:DSB

[SKHK14] A. Slootmaker, H. Kurvers, H. Hummel, and R. Koper. Developing scenario-based serious games for complex cognitive skills acquisition: Design, development and evaluation of the EMERGO platform. J.UCS: Journal of Universal Computer Science, 20(4):561–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_20_4/developing_
scenario_based_serious.

Skillicorn:1997:SPC

[Ski97] D. B. Skillicorn. Structured parallel computation in structured documents. J.UCS: Journal of Universal Computer Science, 3(1):42–68, January 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_3_
1/structured_parallel.

Skillicorn:2000:TCM

[Ski00] D. B. Skillicorn. Treeworld: a conceptual model for large-scale hypermedia. J.UCS: Journal of Universal Computer Science, 6(5):503– 516, May 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_6_5/treeworld_a_
conceptual_model.

Spaniol:2008:IUG

[SKL08] M. Spaniol, R. Klamma, and M. Lux. Imagesemantics: User-generated metadata, content based retrieval & beyond. J.UCS: Journal of Universal Computer Science, 14 (10):1792–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_10/imagesemantics_
user_generated_metadata.

Skordev:2008:SCS

[Sko08]  D. Skordev.  On the

subrecursive computability of several famous constants. J.UCS: Journal of Universal Computer Science, 14(6):861–875, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.

                                        org/jucs_14_6/on_the_
                                        subrecursive_computability.

Stathis:2008:ETB

[SKP08] P. Stathis, E. Kavallieratou, and N. Papamarkos. An evaluation technique for binarization algorithms. J.UCS: Journal of Universal Computer Science, 14(18): 3011–??, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_18/an_evaluation_
                                        technique_for.

Saito:2007:ACQ

[SKSN07] K. Saito, H. Kubota, Y. Sumi, and T. Nishida. Analysis of conversation quanta for conversational knowledge circulation. J.UCS: Journal of Universal Computer Science, 13(2):177– 185, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_
                                        2/analysis_of_conversation_
                                        quanta.

Schropfer:2009:ORA

[SKSP09] A. Schröpfer, F. Kerschbaum, C. Schütz, and R. Pibernik. Optimizations for risk-aware secure supply chain master planning. J.UCS: Journal of Universal Computer Science, 15(15):3019–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
15/optimizations_for_
risk_aware.

Schrum:1996:GCL

[SL96] L. Schrum and T. Lamb.

Groupware for collaborative learning: a research perspective on processes, opportunities, and obstacles. J.UCS: Journal of Universal Computer Science, 2(10):717– 731, October 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
groupware_for_collaborative_
learning.

Strohmaier:2005:IBP

[SL05] M. Strohmaier and S. N.

Lindstaedt. Integrating business processes and knowledge infrastructures. J.UCS: Journal of Universal Computer Science, 11(4):426–428, April 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/free;
http://www.jucs.org/
jucs_11_4/integrating_
business_processes_and.

Schummer:2009:UTP

[SL09] T. Schümmer and S. Lukosch. Understanding tools and practices for distributed pair programming. J.UCS: Journal of Universal Computer Science, 15 (16):3101–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        15_16/understanding_
                                        tools_and_practices.

Somprasertsri:2010:MFO

[SL10] G. Somprasertsri and P. Lalitrojwong. Mining feature-opinion in online customer reviews for opinion summarization. J.UCS: Journal of Universal Computer Science, 16(6):938–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        16_6/mining_feature_
                                        opinion_in.

Santos:2016:GBY

[SLD^{+}16] P. Santos, E. Lex, S. Dennerlein, D. Theiler, J. Cook, T. Treasure-Jones, D. Holley, M. Kerr, G. Attwell, and D. Kowald. Going beyond your personal learning network, using recommendations and trust through a multimedia question-answering service for decision-support: a case study in the healthcare. J.UCS: Journal of Universal Computer Science, 22(3):340– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_22_
3/going_beyond_your_
personal.

Sendin:2008:CEP

[SLJC08] M. Sendín, V. López-Jaquero, and C. A. Collazos. Collaborative explicit plasticity framework: a conceptual scheme for the generation of plastic and group-aware user interfaces. J.UCS: Journal of Universal Computer Science, 14(9):1447–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
9/collaborative_explicit_
plasticity_framework.

Schultes:2011:MUV

[SLK11] P. Schultes, F. Lehrer, and H. Kosch. Markup upon video — towards dynamic and interactive video annotations. J.UCS: Journal of Universal Computer Science, 17(4):605– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_17_
                                        4/markup_upon_video_
                                        towards.

Sagharichian:2016:CED

[SLN16] M. Sagharichian, M. A. Langouri, and H. Naderi. Calculating exact diameter metric of large static graphs. J.UCS: Journal of Universal Computer Science, 22(3):302– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_22_
                                        3/calculating_exact_
                                        diameter_metric.

Shen:2014:CCA

[SLPL14] W. Shen, W. Li, J. A. Pino, and J. Luo. Collaborative computing and applications. J.UCS: Journal of Universal Computer Science, 20(13): 1708–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_20;   http:
                                        //www.jucs.org/jucs_
                                        20_13/collaborative_
                                        computing_and_applications.

Synnes:1998:DEU

[SLPS98] K. Synnes, S. Lachapelle, P. Parnes, and D. Schefström. Distributed education using the mStar environment. J.UCS: Journal of Universal Computer Science, 4(10):807– 823, October 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_10/
distributed_education_
using_the.

Stanczuk:2008:TLP

[SLT08] W. Stańczuk, J. Lubacz, and E. Toczylowski. Trading links and paths on a communication bandwidth market. J.UCS: Journal of Universal Computer Science, 14(5): 642–652, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_5/trading_links_and_
paths.

Stefan:1996:CTC

[ŞM96] Gheorghe Ştefan and Mihaela Maliţa. Chaitin’s ToyLisp on a Connex Memory Machine. J.UCS: Journal of Universal Computer Science, 2(5): 410–426, May 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
5/chaitin_s_toylisp_
on.

Shearer:2002:DPI

[SM02a] J. Shearer and H. Maurer. Is democracy possible in the Internet? J.UCS: Journal of Universal Computer Science, 8(3): 396–407, March 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
3/is_democracy_possible_
in.

Surmann:2002:STT

[SM02b] H. Surmann and A. Morales. Scheduling tasks to a team of autonomous mobile service robots in indoor enviroments. J.UCS: Journal of Universal Computer Science, 8(8): 808–832, August 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
8_8/scheduling_tasks_
to_a.

Stumpf:2004:ISM

[SM04] S. Stumpf and J. Mc-Donnell. An investigation into sharing metadata: “I’m not thinking what you are thinking”. J.UCS: Journal of Universal Computer Science, 10(6):740–748, June 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
6/an_investigation_into_
sharing.

Sancho:2005:SWT

[SMFM05] P. Sancho, I. Martínez, and B. Fernández-Manjón. Semantic Web technologies applied to e-learning personalization in < eaula>. J.UCS: Journal of Universal Computer Science, 11(9):1470–1481, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
9/semantic_web_technologies_
applied.

Sanchez:2009:MGA

[SMGMT09] Ó. Sánchez, F. Molina, J. García-Molina, and A. Toval. ModelSec: a generative architecture for model-driven security. J.UCS: Journal of Universal Computer Science, 15(15):2957–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_15/modelsec_
a_generative_architecture.

Sioutas:2004:GRG

[SMK^{+}04] S. Sioutas, C. Makris, N. Kitsios, G. Lagogiannis, J. Tsaknakis, K. Tsichlas, and B. Vassiliadis. Geometric retrieval for grid points in the RAM model. J.UCS: Journal of Universal Computer Science, 10(9):1325– 1353, September 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_10_9/geometric_
retrieval_for_grid.

Scotton:2010:MSA

[SMMC10] J. Scotton, S. Moebs, J. McManis, and A. I. Cristea. Merging strategies for authoring QoEbased adaptive hypermedia. J.UCS: Journal of Universal Computer Science, 16(19):2756–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
19/merging_strategies_
for_authoring.

Sosevic:2013:SPL

[SMMM13] U. Sosević, I. Milenković, M. Milovanović, and M. Minović. Support platform for learning about multimodal biometrics. J.UCS: Journal of Universal Computer Science, 19(11):1684–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
11/support_platform_
for_learning.

Sielis:2011:CAR

[SMP^{+}11] G. A. Sielis, C. Mettouris, G. A. Papadopoulos, A. Tzanavari, R. M. Dols, and Q. Siebers. A context aware recommender system for creativity support tools. J.UCS: Journal of Universal Computer Science, 17(12):1743–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_12/a_context_
aware_recommender.

Sirjani:2005:MCA

[SMSdB05] M. Sirjani, A. Movaghar, A. Shali, and F. S. de Boer. Model checking, automated abstraction, and compositional verification of Rebeca models. J.UCS: Journal of Universal Computer Science, 11(6):1054– 1082, ???? 2005. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_6/model_checking_
automated_abstraction.

Sousa:2008:MDA

[SMV08] K. Sousa, H. Mendonça, and J. Vanderdonckt. A model-driven approach to align business processes with user interfaces. J.UCS: Journal of Universal Computer Science, 14(19):3236–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_19/a_model_driven_
                                        approach.

Smyth:2000:RBD

[Smy00] M. B. Smyth. Regionbased discrete geometry. J.UCS: Journal of Universal Computer Science, 6(4):447–459, April 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_6_4/region_
                                        based_discrete_geometry.

Stark:2001:LAS

[SN01] R. F. Stärk and S. Nanchen. A logic for abstract state machines. J.UCS: Journal of Universal Computer Science, 7(11):980– 1005, November 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_7_
                                        11/a_logic_for_abstract.

Sasaki:2007:HSC

[SNAF07] S. Sasaki, T. Nishihara, D. Ando, and M. Fujita. Hardware/software co-design and verification methodology from system level based on system dependence graph.

J.UCS: Journal of Universal Computer Science, 13(13):1972–2001, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_13_13/hardware_
software_codesign_and.

Sobecki:2005:CBH

[Sob05] J. Sobecki. Consensusbased hybrid adaptation of Web systems user interfaces. J.UCS: Journal of Universal Computer Science, 11(2):250– 270, February 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_2/consensus_based_
hybrid_adaptation.

Sorensen:1997:PIT

[SOO97] H. Sorensen, A. O’Riordan, and C. O’Riordan. Profiling with the INFOrmer text filtering agent. J.UCS: Journal of Universal Computer Science, 3(8): 988–1006, August 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://medoc.springer.
de:8000/jucs_3_8/profiling_
with_the_informer;
internal&sk=05460486.

Soufi:1999:TSR

[Sou99] L. Soufi. Type specification by regular expressions. J.UCS: Journal of Universal Computer Science, 5(9):622– 631, September 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_5_
                                        9/type_specification_
                                        by_regular.

Smailovic:2016:MSN

[SP16] V. Smailovic and V. Podobnik. Mining social networks for calculation of SmartSocial influence. J.UCS: Journal of Universal Computer Science, 22(3):394–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_22_3/mining_
                                        social_networks_for.

Spandl:2008:CTP

[Spa08] C. Spandl. Computability of topological pressure for sofic shifts with applications in statistical physics. J.UCS: Journal of Universal Computer Science, 14(6):876– 895, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        6/computability_of_topological_
                                        pressure.

Spitters:2005:CRO

[Spi05] B. Spitters. Constructive results on operator algebras. J.UCS: Journal of Universal Computer Science, 11(12):2096–2113, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
12/constructive_results_
on_operator.

Sanchez:2009:MSM

[SPRP09] L. E. Sánchez, A. Santos-Olmo Parra, D. G. Rosado, and M. Piattini. Managing security and its maturity in small and medium-sized enterprises. J.UCS: Journal of Universal Computer Science, 15(15):3038–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
15/managing_security_
and_its.

Schindel:2000:MTC

[SR00] W. D. Schindel and G. M.

Rogers. Methodologies and tools for continuous improvement of systems. J.UCS: Journal of Universal Computer Science, 6(3):289– 323, March 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_6_
                                        3/methodologies_and_
                                        tools_for.

Sabucedo:2010:LCE

[SR10] L. Álvarez Sabucedo and L. Anido Rifón. Locating and crawling eGovernment services. A light-weight semantic approach. J.UCS: Journal of Universal Computer Science, 16(8):1117– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_16_
                                        8/locating_and_crawling_
                                        egovernment.

Skyrme:2008:ELC

[SRI08] A. Skyrme, N. Rodriguez, and R. Ierusalimschy. Exploring Lua for concurrent programming. J.UCS: Journal of Universal Computer Science, 14(21):3556–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        21/exploring_lua_for_
                                        concurrent.

Sveda:2004:FSC

[SRR04] M. Sveda, C. Rattray, and J. W. Rozenblit. Formal specification of computer-based systems.

J.UCS: Journal of Universal Computer Science, 10(10):1354–1356, October 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/free;
http://www.jucs.org/
jucs_10_10/formal_specification_
of_computer.

Seres:2000:FRL

[SS00] S. Seres and M. Spivey.

Functional reading of logic programs. J.UCS: Journal of Universal Computer Science, 6(4):433– 446, April 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
4/functional_reading_
of_logic.

Scuglik:2003:AGC

[SS03] F. Scuglik and M. Sveda.

Automatically generated CSP specifications. J.UCS: Journal of Universal Computer Science, 9 (11):1277–1295, November 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
11/automatically_generated_
csp_specifications.

Settle:2007:DLS

[SS07] A. Settle and C. Settle. Distance learning and student satisfaction in Java programming courses. J.UCS: Journal of Universal Computer Science, 13(9):1270– 1286, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
9/distance_learning_
and_student.

Schrage:2008:BAP

[SS08] M. M. Schrage and S. D. Swierstra. Beyond ASCII — parsing programs with graphical presentations. J.UCS: Journal of Universal Computer Science, 14(21):3414–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
21/beyond_ascii_parsing_
programs.

Sali:2009:CCI

[SS09a] A. Sali and K.-D. Schewe. A characterisation of coincidence ideals for complex values. J.UCS: Journal of Universal Computer Science, 15(1):304– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
1/a_characterisation_
of_coincidence.

Selivanova:2009:CSO

[SS09b] S. Selivanova and V. Selivanov. Computing the solution operators of symmetric hyperbolic systems of PDE. J.UCS: Journal of Universal Computer Science, 15(6):1337– ??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_6/computing_
the_solution_operators.

Spruit:2015:DDW

[SS15] M. Spruit and C. Sacu.

DWCMM: The data warehouse capability maturity model. J.UCS: Journal of Universal Computer Science, 21(11):1508–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_11/dwcmm_the_
data_warehouse.

Sanchez-Santana:2013:TTC

[SSAB^{+}13] M.-A. Sanchez-Santana, J.-B. Aupet, M.-L. Betbeder, J.-C. Lapayre, and A. Camarena-Ibarrola. A tool for telediagnosis of cardiovascular diseases in a collaborative and adaptive approach. J.UCS: Journal of Universal Computer Science, 19(9):1275–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_19_9/a_tool_for_
                                        telediagnosis.

Stenzhorn:2008:ACO

[SSBS08] H. Stenzhorn, S. Schulz, M. Boeker, and B. Smith. Adapting clinical ontologies in real-world environments. J.UCS: Journal of Universal Computer Science, 14(22):3767–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        22/adapting_clinical_
                                        ontologies_in.

Stroele:2011:IWB

[SSdS^{+}11] V. Ströele, R. Silva, M. Ferreira de Souza, C. E. R. de Mello, J. M. Souza, G. Zimbrao, and J. Oliveira. Identifying workgroups in Brazilian scientific social networks. J.UCS: Journal of Universal Computer Science, 17(14):1951–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_17_
                                        14/identifying_workgroups_
                                        in_brazilian.

Sun:2010:TIT

[SSGS10] W. Sun, E. Song, P. C. Grabow, and D. M. Simmonds. Toward an integrated tool environment for static analysis of UML class and sequence models. J.UCS: Journal of Universal Computer Science, 16(17):2435–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_17/toward_an_
integrated_tool.

Schmidt:2011:TMI

[SSM11] B. Schmidt, T. Stoitsev, and M. Mühlhäuser. Task models for intentionaware systems. J.UCS: Journal of Universal Computer Science, 17 (10):1511–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_10/task_models_for_
intention.

Suryanarayanan:2012:EKN

[SSS12] M. G. Suryanarayanan, D. S. Sam, and S. Selvaraju. Establishing knowledge networks via analysis of research abstracts. J.UCS: Journal of Universal Computer Science, 18(8):993– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_

                                        8/establishing_knowledge_
                                        networks_via.

Shirali-Shahreza:2006:PAB

[SSSS06] M. H. Shirali-Shahreza and M. Shirali-Shahreza. Persian/Arabic baffletext CAPTCHA. J.UCS: Journal of Universal Computer Science, 12 (12):1783, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        12_12/persian_arabic_
                                        baffletext_captcha.

Silvasti:2010:ELX

[SSSS10] P. Silvasti, S. Sippu, and E. Soisalon-Soininen. Evaluating linear XPath expressions by patternmatching automata. J.UCS: Journal of Universal Computer Science, 16 (5):833–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_16_5/evaluating_
                                        linear_xpath_expressions.

Sanin:2007:OOS

[SST07] C. Sanin, E. Szczerbicki, and C. Toro. An OWL ontology of set of experience knowledge structure. J.UCS: Journal of Universal Computer Science, 13(2):209–223, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_2/an_owl_ontology_
of.

Studer:2002:MUF

[SSV02] R. Studer, Y. Sure, and R. Volz. Managing user focused access to distributed knowledge. J.UCS: Journal of Universal Computer Science, 8(6):662– 672, June 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_8_
6/managing_user_focused_
access.

Sambin:2005:MPR

[ST05] G. Sambin and G. Trentinaglia. On the meaning of positivity relations for regular formal spaces. J.UCS: Journal of Universal Computer Science, 11(12):2056–2062, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
12/on_the_meaning_of.

Staiger:2002:HLS

[Sta02] L. Staiger. How large is the set of disjunctive sequences? J.UCS: Journal of Universal Computer Science, 8(2):348– 362, February 28, 2002.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        2/how_large_is_the.

Staiger:2005:HML

[Sta05] L. Staiger. Hausdorff measure and Lukasiewicz languages. J.UCS: Journal of Universal Computer Science, 11(12):2114–2124, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_11_
                                        12/hausdorff_measure_
                                        and_lukasiewicz.

Soler:2009:DSD

[STBFM09] E. Soler, J. Trujillo, C. Blanco, and E. Fernández-Medina. Designing secure data warehouses by using MDA and QVT. J.UCS: Journal of Universal Computer Science, 15(8):1607–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_15_8/designing_
                                        secure_data_warehouses.

Stefanescu:1995:MPS

[Şte95] Cătălina Ştefănescu. A Markov process for sequential allocation. J.UCS: Journal of Universal Computer Science, 1(12): 821–827, December 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
a_markov_process_for_
sequential_allocation.

Stefanescu:1996:PCR

[Şte96] Doru Ştefănescu. Polynomials, constructivity and randomness. J.UCS: Journal of Universal Computer Science, 2(5): 396–409, May 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_2_5/polynomials_
constructivity_and_randomness.

Steggles:2000:SVR

[Ste00] L. J. Steggles. Specifying and verifying real-time systems using secondorder algebraic methods: a case study of the railroad crossing controller. J.UCS: Journal of Universal Computer Science, 6(4):460–473, April 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_6_4/specifying_
and_verifying_real.

Stefanescu:2005:NBP

[Şte05] Doru Ştefănescu. New bounds for positive roots of polynomials. J.UCS: Journal of Universal Computer Science, 11 (12):2125–2131, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_12/new_bounds_for_
positive.

Stewart:2008:ACO

[Ste08] C. Stewart. Authoring & culture in online education. J.UCS: Journal of Universal Computer Science, 14(17):2877–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_17/authoring_
and_culture_in.

Sancho:2012:MCE

[STFM12] P. Sancho, J. Torrente, and B. Fernández-Manjón. MareMonstrum: a contribution to empirical research about how the use of MUVEs may improve students’ motivation. J.UCS: Journal of Universal Computer Science, 18(18):2576–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_18/maremonstrum_a_
contribution_to.

Stolzenburg:1999:LDH

[Sto99] F. Stolzenburg. Loopdetection in hyper-tableaux by powerful model generation. J.UCS: Journal of Universal Computer Science, 5(3):135– 155, March 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
3/loop_detection_in_
hyper.

Stork:2002:WGK

[Sto02] H. G. Stork. Webs, grids and knowledge spaces: Programmes, projects and prospects. J.UCS: Journal of Universal Computer Science, 8(9):848– 868, September 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_8_9/webs_grids_
and_knowledge.

Stojanovic:2003:RLA

[Sto03] N. Stojanovic. On the role of the librarian agent in ontology-based knowledge management systems. J.UCS: Journal of Universal Computer Science, 9(7):697– 718, July 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
7/on_the_role_of.

Stroetmann:1997:CSP

[Str97]  K. Stroetmann.  The

constrained shortest path problem: a case study in using ASMs. J.UCS: Journal of Universal Computer Science, 3(4): 304–319, April 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs_3_4/shortest_
                                        path_problem;   http:
                                        //www.jucs.org/jucs_
                                        3_4/shortest_path_problem;
                                         internal&sk=05460486.

Sioutas:2007:EAM

[STVT07] S. Sioutas, K. Tsichlas, B. Vassiliadis, and D. Tsolis. Efficient access methods for temporal interval queries of video metadata. J.UCS: Journal of Universal Computer Science, 13(10):1411–1433, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_
                                        10/efficient_access_
                                        methods_for.

Schewe:2009:USU

[STW09] K.-D. Schewe, B. Thalheim, and Q. Wang. Updates, schema updates and validation of XML documents — using abstract state machines with automata-defined states. J.UCS: Journal of Universal Computer Science, 15(10):2028–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_10/updates_schema_
updates_and.

Sigmund:2001:SCS

[SU01] U. Sigmund and T. Ungerer. On speculation control in simultaneous multithreaded processors. J.UCS: Journal of Universal Computer Science, 7(9):848–868, September 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
9/on_speculation_control_
in.

Suen:2010:VRI

[Sue10] C. H. Suen. VIMM:

Runtime integrity measurement of a virtualized operating system. J.UCS: Journal of Universal Computer Science, 16(4):554– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
4/vimm_runtime_integrity_
measurement.

Seifert:2013:TRE

[SUKG13] C. Seifert, E. Ulbrich, R. Kern, and M. Granitzer. Text representation for efficient document annotation. J.UCS: Journal of Universal Computer Science, 19(3):383– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_19_
                                        3/text_representation_
                                        for_efficient.

Suthers:2001:TSS

[Sut01] D. D. Suthers. Towards a systematic study of representational guidance for collaborative learing discourse. J.UCS: Journal of Universal Computer Science, 7(3):254–277, March 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_7_3/towards_
                                        a_systematic_study.

Suzuki:2006:DMM

[Suz06] E. Suzuki. Data mining methods for discovering interesting exceptions from an unsupervised table. J.UCS: Journal of Universal Computer Science, 12(6):627– 653, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
6/data_mining_methods_
for.

Shearer:1995:SPI

[SV95] J. Shearer and A. Vermeer. Software patents and the Internet. J.UCS: Journal of Universal Computer Science, 1(5): 312–319, May 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
1_5/software_patents_
and_the.

Sitaraman:2005:LIC

[SV05] S. Sitaraman and S. Venkatesan. Low-intrusive consistent disk checkpointing: a tool for digital forensics. J.UCS: Journal of Universal Computer Science, 11(1):20–36, January 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
1/low_intrusive_consistent_
disk.

Slissenko:2008:STA

[SV08] A. Slissenko and P. Vasilyev. Simulation of timed abstract state machines with predicate logic model-checking. J.UCS: Journal of Universal Computer Science, 14 (12):1984–??, ???? 2008.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_12/simulation_of_
timed_abstract.

Safont:2015:VLS

[SVFR15] L. Vicent Safont, S. Villagrasa, D. Fonseca, and E. Redondo. Virtual learning scenarios for qualitative assessment in higher education 3D arts. J.UCS: Journal of Universal Computer Science, 21(8):1086–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
8/virtual_learning_scenarios_
for.

Santos:2015:TDO

[SVK^{+}15] J. L. Santos, K. Verbert, J. Klerkx, E. Duval, S. Charleer, and S. Ternier. Tracking data in open learning environments. J.UCS: Journal of Universal Computer Science, 21(7):976– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
7/tracking_data_in_open.

Svozil:1995:HPA

[Svo95] K. Svozil. Halting probability amplitude of quantum computers. J.UCS: Journal of Universal Computer Science, 1(3): 201–204, March 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
halting_probability_
amplitude_of_quantum_
computers.

Svozil:1996:QAI

[Svo96] K. Svozil. Quantum algorithmic information theory. J.UCS: Journal of Universal Computer Science, 2(5):311–346, May 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_2_5/quantum_
information_theory.

Spoelstra:2014:TPB

[SvRS14] H. Spoelstra, P. van Rosmalen, and P. Sloep. Toward project-based learning and team formation in open learning environments. J.UCS: Journal of Universal Computer Science, 20(1):57– ??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
20_1/toward_project_
based_learning.

Spoelstra:2013:TFP

[SvRvdV^{+}13] H. Spoelstra, P. van Rosmalen, E. van de Vrie, M. Obreza, and P. Sloep. A team formation and project-based learning support service for social learning networks. J.UCS: Journal of Universal Computer Science, 19(10):1474–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_19_
                                        10/a_team_formation_
                                        and.

Stragier:2010:HGP

[SVV10] J. Stragier, P. Verdegem, and G. Verleye. How is e-government progressing? A data driven approach to E-government monitoring. J.UCS: Journal of Universal Computer Science, 16(8):1075– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_16_
                                        8/how_is_egovernment_
                                        progressing.

Schmidt:2004:UCA

[SW04] A. Schmidt and C. Winterhalter. User context aware delivery of Elearning material: Approach and architecture. J.UCS: Journal of Universal Computer Science, 10(1):38–46, January 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
1/user_context_aware_
delivery.

Shin:2009:SCM

[SW09] C. Shin and W. Woo.

Service conflict management framework for multiuser inhabited smart home. J.UCS: Journal of Universal Computer Science, 15(12):2330–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
12/service_conflict_
management_framework.

Schewe:2010:XDT

[SW10] K.-D. Schewe and Q. Wang.

XML database transformations. J.UCS: Journal of Universal Computer Science, 16(20):3043–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
20/xml_database_transformations.

Sobolewski:2013:CDD

[SW13] P. Sobolewski and M. Woźniak.

Concept drift detection and model selection with simulated recurrence and ensembles of statistical detectors. J.UCS: Journal of Universal Computer Science, 19(4):462– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
4/concept_drift_detection_
and.

Swiatek:2007:PES

[Świ07] J. Światek. Parameter estimation of systems described by the relation with noisy observations. J.UCS: Journal of Universal Computer Science, 13(2):199–208, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_13_2/parameter_
estimation_of_systems.

Sun:2009:AAP

[SWY09] L. Sun, H. Wang, and J. Yong. Authorization algorithms for permissionrole assignments. J.UCS: Journal of Universal Computer Science, 15(9): 1782–??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_9/authorization_algorithms_
for_permission.

Salomaa:1999:SEL

[SY99] K. Salomaa and S. Yu.

Synchronization expressions and languages. J.UCS: Journal of Universal Computer Science, 5(9):610–621, September 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
5_9/synchronization_
expressions_and_languages.

Stroele:2012:MMA

[SZS12] V. Ströele, G. Zimbrão, and J. M. Souza. Modeling, mining and analysis of multi-relational scientific social network. J.UCS: Journal of Universal Computer Science, 18(8):1048–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_8/modeling_mining_
and_analysis.

She:2014:MIS

[SZWdP14] J. She, X. Zhang, W. Wang, and P. Ordóñez de Pablos. Mapping the impact of social media and mobile Internet on Chinese Academia’s performance: A case on telemedicine research 2005-2013. J.UCS: Journal of Universal Computer Science, 20 (15):2005–??, ???? 2014.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        20_15/mapping_the_impact_
                                        of.

Seberry:1995:RBP

[SZZ95] Jennifer Seberry, Xian-Mo Zhang, and Yuliang Zheng. The relationship between propagation characteristics and nonlinearity of cryptographic functions. J.UCS: Journal of Universal Computer Science, 1(2):136– 150, February 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        1_2/the_relationship_
                                        between_propagation.

Shen:2010:NMN

[SZZM10] W. Shen, C. Zhang, J. Zhang, and X. Ma. Newton method for nonlinear dynamic systems with adaptive time stepping. J.UCS: Journal of Universal Computer Science, 16(6):891– ??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_16_
                                        6/newton_method_for_
                                        nonlinear.

Tabakow:2007:UPI

[Tab07] I. Tabakow. Using place invariants and test point placement to isolate faults in discrete event systems. J.UCS: Journal of Universal Computer Science, 13(2):224–243, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_13_2/using_
place_invariants_and.

Takahashi:2003:RAP

[Tak03] K. Takahashi. Reasoning about propagation of properties over regions. J.UCS: Journal of Universal Computer Science, 9(9):1030–1045, September 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
9_9/reasoning_about_
propagation_of.

Takacs:2006:AEK

[Tak06] P. Takács. The additional examination of the Kudo– Mathuria time-release protocol. J.UCS: Journal of Universal Computer Science, 12(9):1373– 1384, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_9/the_additional_
examination_of%20.

Tobar:2008:WTA

[TAL08] L. M. Tobar, P. M. Latorre Andrés, and E. Lafuente Lapena. WebA: a tool for the assistance in design and evaluation of Websites. J.UCS: Journal of Universal Computer Science, 14(9):1496– ??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        9/weba_a_tool_for.

Tanter:2008:CAR

[Tan08] É. Tanter. Controlling aspect reentrancy. J.UCS: Journal of Universal Computer Science, 14 (21):3498–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_21/controlling_aspect_
                                        reentrancy.

Tarasov:2012:OBA

[Tar12] V. Tarasov. Ontologybased approach to competence profile management. J.UCS: Journal of Universal Computer Science, 18(20):2893–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_18_20/ontology_
                                        based_approach_to.

Tataram:2007:CRS

[Tat07] M. Tatarâm. Connectivity and reachability in signed networks. J.UCS: Journal of Universal Computer Science, 13(11):1779– 1790, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
11/connectivity_and_
reachability_in.

Ting:2016:AMS

[TB16] I.-H. Ting and B. Birregah. Analyzing and mining social networks for decision support. J.UCS: Journal of Universal Computer Science, 22(3):298–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_22;   http:
//www.jucs.org/jucs_
22_3/analyzing_and_mining_
social.

Traxler:2015:MTF

[TBL15] J. Traxler, E. Barcena, and J. García Laborda. Mobile technology for foreign language teaching: Building bridges between non-formal and formal scenarios. J.UCS: Journal of Universal Computer Science, 21(10): 1234–??, ???? 2015.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_21;   http:
//www.jucs.org/jucs_
21_10/mobile_technology_
for_foreign.

Tirelo:2008:DDS

[TBS08] F. Tirelo, R. S. Bigonha, and J. Saraiva. Disentangling denotational semantics definitions. J.UCS: Journal of Universal Computer Science, 14 (21):3592–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_21/disentangling_
denotational_semantics_
definitions.

Torre-Bastida:2015:DIE

[TBVRGLD15] A. I. Torre-Bastida, E. Villar-Rodriguez, S. Gil-Lopez, and J. Del Ser. Design and implementation of an extended corporate CRM-Database system with Big Data analytical functionalities. J.UCS: Journal of Universal Computer Science, 21(6):757–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_6/design_and_
implementation_of.

Thomas:2001:DKT

[TCK^{+}01] Jim Thomas, Paula Cowley, Olga Kuchar, Lucy Nowell, Judi Thompson, and Pak Chung Wong. Discovering knowledge through visual analysis. J.UCS: Journal of Universal Computer Science, 7(6):517–529, June 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_7_6/discovering_
knowledge_through_visual.

Tao:2003:TSG

[TCS^{+}03] F. Tao, L. Chen, N. Shadbolt, G. Pound, and S. Cox. Towards the semantic grid: Putting knowledge to work in design optimisation. J.UCS: Journal of Universal Computer Science, 9(6): 551–562, June 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_9_6/towards_the_
semantic_grid.

Ting:2012:UMU

[TCW12] I.-H. Ting, P. S. Chang, and S.-L. Wang. Understanding microblog users for social recommendation based on social networks analysis. J.UCS: Journal of Universal Computer Science, 18(4):554– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
4/understanding_microblog_
users_for.

Tochterman:1996:DFH

[TD96] K. Tochterman and G. Dittrich. The Dortmund family of hypermedia models

— concepts and their application. J.UCS: Journal of Universal Computer Science, 2(1):34–56, January 28, 1996. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_2_
1/the_dortmund_family_
of.

Taveira:2003:ARM

[Tddd03] Wendell Figueiredo Taveira, Marco Tulio de Oliveira Valente, Mariza Andrade da Silva Bigonha, and Roberto da Silva Bigonha. Asynchronous remote method invocation in Java. J.UCS: Journal of Universal Computer Science, 9(8):761– 775, August 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
8/asynchronous_remote_
method_invocation.

Toth:2006:SEN

[TE06] T. Tóth and F. Erdélyi.

Systems engineering: a new approach to complex IT-based technological systems in engineering education. J.UCS: Journal of Universal Computer Science, 12(9):1393– 1404, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
9/systems_engineering_
a_new%20.

Troyano:2007:IPT

[TEC^{+}07] J. A. Troyano, F. Enríquez, F. Cruz, J. M. Cañete-Valdeón, and F. J. Ortega. Improving the performance of a tagger generator in an information extraction application. J.UCS: Journal of Universal Computer Science, 13(9):1287– 1299, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_13_9/improving_
the_performance_of.

Tedre:2009:CE

[Ted09] M. Tedre. Computing as engineering. J.UCS: Journal of Universal Computer Science, 15(8):1642– ??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_15_8/computing_
                                        as_engineering.

Tran:2008:EAB

[TEK08] C. D. Tran, H. Ezzedine, and C. Kolski. Evaluation of agent-based interactive systems: Proposal of an electronic informer using Petri nets. J.UCS: Journal of Universal Computer Science, 14(19): 3202–??, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_19/evaluation_of_
                                        agent_based.

Tezer:2013:POR

[Tez13] M. Tezer. Parent opinions with regard to elementary school student’s use of the Internet. J.UCS: Journal of Universal Computer Science, 19(5):692– ??, ???? 2013. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        19_5/parent_opinions_
                                        with_regard.

Tanter:2009:SCS

[TF09] É. Tanter and J. Fabry. Supporting composition of structural aspects in an AOP kernel. J.UCS: Journal of Universal Computer Science, 15(3):620– ??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_3/supporting_
composition_of_structural.

Tentori:2006:QPQ

[TFG06] M. Tentori, J. Favela, and V. M. González. Quality of privacy (QoP) for the design of ubiquitous healthcare applications. J.UCS: Journal of Universal Computer Science, 12(3):252–269, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_12_3/quality_of_
privacy_for.

Terwilliger:2010:SSE

[TFMDM10] J. F. Terwilliger, R. J.

Fernández-Moctezuma, L. M. L. Delcambre, and D. Maier. Support for schema evolution in data stream management systems. J.UCS: Journal of Universal Computer Science, 16 (20):3073–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_20/support_for_schema_
evolution.

Terissi:2010:HPF

[TG10] L. D. Terissi and J. C. Gómez. 3D head pose and facial expression tracking using a single camera. J.UCS: Journal of Universal Computer Science, 16(6):903–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        16_6/3d_head_pose_and.

Trillo:2007:DSU

[TGEM07] R. Trillo, J. Gracia, M. Espinoza, and E. Mena. Discovering the semantics of user keywords. J.UCS: Journal of Universal Computer Science, 13(12):1908–1935, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_
                                        12/discovering_the_semantics_
                                        of.

Tesoriero:2010:CMD

[TGLP10] R. Tesoriero, J. A. Gallud, M. D. Lozano, and V. M. R. Penichet. CAUCE: Model-driven development of contextaware applications for ubiquitous computing environments. J.UCS: Journal of Universal Computer Science, 16(15): 2111–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_15/cauce_model_driven_
development.

Tan:1999:MMD

[TH99] E. C. Tan and M. H. Ho.

Matrix method to detect logic hazards in combinational circuits with EX-OR gates. J.UCS: Journal of Universal Computer Science, 5(11):765– 776, November 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
11/matrix_method_to_
detect.

Thalheim:2010:TTC

[Tha10] B. Thalheim. Towards a theory of conceptual modelling. J.UCS: Journal of Universal Computer Science, 16(20):3102–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_20/towards_a_
theory_of.

Thies:2000:CCP

[Thi00] S. Thies. Coffein: Construction and presentation of design knowledge. J.UCS: Journal of Universal Computer Science, 6(3):324–344, March 28, 2000. CODEN ????

ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_6_3/coffein_
construction_and_presentation.

Tran:2016:TSS

[THJ16] V. C. Tran, D. Hwang, and J. J. Jung. TwiSNER: Semi-supervised method for named entity recognition from text streams on Twitter. J.UCS: Journal of Universal Computer Science, 22(6):782– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_22_
6/twisner_semi_supervised_
method.

Trattner:2011:CEN

[THS11] C. Trattner, D. Helic, and M. Strohmaier. On the construction of efficiently navigable tag clouds using knowledge from structured Web content. J.UCS: Journal of Universal Computer Science, 17(4):565–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_17_4/on_the_
construction_of.

Tillian:2001:KMM

[Til01] B. Tillian. Knowledge management more effort

— more success? J.UCS:

Journal of Universal Computer Science, 7(7): 602–609, July 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
7/knowledge_management_
more_effort.

Tahera:2008:GNG

[TIL08] K. Tahera, R. N. Ibrahim, and P. B. Lochert. GADYM — a novel genetic algorithm in mechanical design problems. J.UCS: Journal of Universal Computer Science, 14 (15):2566–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_15/gadym_a_novel_
genetic.

Ting:2016:DSS

[Tin16] I.-H. Ting. Detection of the spiral of silence effect in social media. J.UCS: Journal of Universal Computer Science, 22(3):438–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_22_3/detection_
of_the_spiral.

Tran:2015:CES

[TJ15] Q. D. Tran and J. E. Jung.

CoCharNet: Extracting social networks using character co-occurrence in movies. J.UCS: Journal of Universal Computer Science, 21(6):796–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_6/cocharnet_
extracting_social_networks.

Taivan:2013:DCA

[TJS^{+}13] C. Taivan, R. José, B. Silva, I. Elhart, and J. Cardoso. Design considerations for application selection and control in multi-user public displays. J.UCS: Journal of Universal Computer Science, 19(17):2526–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
17/design_considerations_
for_application.

Tackenberg:2009:OSC

[TKD^{+}09] S. Tackenberg, B. Kausch, S. Duckwitz, C. M. Schlick, and S. Karahancer. Organizational simulation of complex process engineering projects in the chemical industry. J.UCS: Journal of Universal Computer Science, 15(9):1746–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_9/organizational_
simulation_of_complex.

Tick:2006:SOW

[TKF06] J. Tick, Z. Kovacs, and F. Friedler. Synthesis of optimal workflow structure. J.UCS: Journal of Universal Computer Science, 12(9):1385– 1392, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_12_9/synthesis_
of_optimal_workflow.

Ternier:2012:AAR

[TKK^{+}12] S. Ternier, R. Klemke, M. Kalz, P. van Ulzen, and M. Specht. AR-Learn: Augmented reality meets augmented virtuality. J.UCS: Journal of Universal Computer Science, 18(15):2143–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
15/ARLearn_augmented_
reality_meets.

Taniar:2011:CC

[TKP11] D. Taniar, I. Khalil, and E. Pardede. Cloud computing. J.UCS: Journal of Universal Computer Science, 17(8):1134–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_17;   http:
//www.jucs.org/jucs_
17_8/cloud_computing;
 http://www.jucs.org/
jucsrssfeed-issue.

Timbrell:2005:KIH

[TKSL05] G. Timbrell, S. Koller, N. Schefe, and S. Lindstaedt. A knowledge infrastructure hierarchy model for call-centre processes. J.UCS: Journal of Universal Computer Science, 11(4):546–564, April 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
4/a_knowledge_infrastructure_
hierarchy.

Tzitzikas:2007:RRS

[TKT07] Y. Tzitzikas, D. Kotzinos, and Y. Theoharis. On ranking RDF schema elements (and its application in visualization). J.UCS: Journal of Universal Computer Science, 13(12):1854–1880, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_12/on_ranking_rdf_
schema.

Teufl:2011:KER

[TL11] P. Teufl and G. Lackner. Knowledge extraction from RDF data with activation patterns. J.UCS: Journal of Universal Computer Science, 17(7):983–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_7/knowledge_
extraction_from_RDF.

Tsai:2011:AGA

[TLJ11] C.-F. Tsai, W. Li, and A. James. An adaptive genetic algorithm and application in a luggage design center. J.UCS: Journal of Universal Computer Science, 17(14):2029–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
14/an_adaptive_genetic_
algorithm.

Traonouez:2009:PMC

[TLR09] L.-M. Traonouez, D. Lime, and O. H. Roux. Parametric model-checking of stopwatch Petri nets. J.UCS: Journal of Universal Computer Science, 15(17):3273–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/

                                        jucs_15_17/parametric_
                                        model_checking_of.

Tissenbaum:2012:CDC

[TLS12] M. Tissenbaum, M. Lui, and J. D. Slotta. Codesigning collaborative smart classroom curriculum for secondary school science. J.UCS: Journal of Universal Computer Science, 18(3):327– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_18_
                                        3/co_designing_collaborative_
                                        smart.

Tochtermann:2000:KME

[TM00] K. Tochtermann and H. Maurer. Knowledge management and environmental informatics. J.UCS: Journal of Universal Computer Science, 6(5):517–536, May 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_6_
                                        5/knowledge_management_
                                        and_environmental.

Tochtermann:2001:JUSa

[TM01a] K. Tochtermann and H. Maurer. J.UCS special issue: I — Know ’01

— International Conference on Knowledge Management. J.UCS: Journal of Universal Computer Science, 7(6):456– 457, June 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
6/j_ucs_special_issue.

Tochtermann:2001:JUSb

[TM01b] K. Tochtermann and H. Maurer. J.UCS special issue: I — Know ’01 — International Conference on Knowledge Management. J.UCS: Journal of Universal Computer Science, 7(7):548–549, July 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_7_7/i_know_
international_conference.

Tochtermann:2002:JUSb

[TM02a] K. Tochtermann and H. Maurer. J.UCS special issue: I-Know ’02 — People-Oriented Knowledge Management. J.UCS: Journal of Universal Computer Science, 8(5):453–455, May 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/free;
http://www.jucs.org/
jucs_8_5/j_ucs_special_
issue.

Tochtermann:2002:JUSc

[TM02b] K. Tochtermann and H. Maurer. J.UCS special issue: I-Know ’02 — Technology-Oriented Knowledge Management. J.UCS: Journal of Universal Computer Science, 8(6):567–569, June 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/free;
http://www.jucs.org/
jucs_8_6/j_ucs_special_
issue.

Telmoudi:2009:MRM

[TNM09] A. J. Telmoudi, L. Nabli, and R. M’hiri. Modeling of robustness margins of the control of a predictive control-supervisory architecture. J.UCS: Journal of Universal Computer Science, 15(17): 3231–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_17/modeling_
of_robustness_margins.

Torres-Nino:2013:IAD

[TNRGCP^{+}13] J. Torres-Niño, A. Rodríguez-González, R. Colomo-Palacios, E. Jiménez-Domingo, and G. Alor-Hernandez. Improving accuracy of decision trees using clustering techniques. J.UCS: Journal of Universal Computer Science, 19(4):484– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
4/improving_accuracy_
of_decision.

Tochtermann:2002:JUSa

[Toc02] K. Tochtermann. J.UCS special issue: Hypermedia — State of the Art 2002. J.UCS: Journal of Universal Computer Science, 8(10):869–870, October 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/free;
http://www.jucs.org/
jucs_8_10/j_ucs_special_
issue.

Tochtermann:2003:JUS

[Toc03] K. Tochtermann. J.UCS special issue: I-Know ’03 — Hot Spots in Knowledge Management. J.UCS: Journal of Universal Computer Science, 9(6):447–450, June 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/free;
http://www.jucs.org/
jucs_9_6/hot_spots_in_
knowledge.

Tochtermann:2004:BSA

[Toc04] K. Tochtermann. Beyond the state-of-the-art of knowledge management. J.UCS: Journal of Universal Computer Science, 10(6):671–673, June 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/free;
                                        http://www.jucs.org/
                                        jucs_10_6/beyond_the_
                                        state_of.

Tomek:1995:MTC

[Tom95] I. Tomek. Microworlds for teaching concepts of object oriented programming. J.UCS: Journal of Universal Computer Science, 1(6):423–434, June 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/Microworlds_For_
                                        Teaching_Concepts_of_
                                        Object_Oriented_Programming.

Tomescu:1997:OHF

[Tom97] I. Tomescu. Optimum Huffman forests. J.UCS: Journal of Universal Computer Science, 3(7): 813–820, July 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs_3_7/optimum_
                                        huffman_forests;   internal&
                                        sk=05460486.

Tomek:2001:KMC

[Tom01] I. Tomek. Knowledge management and collaborative virtual environments. J.UCS: Journal of Universal Computer Science, 7(6):458– 471, June 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
6/knowledge_management_
and_collaborative.

Tome:2003:BMC

[Tom03] E. Tomé. Bad management and its consequences in a problematic European Union member (Portugal). J.UCS: Journal of Universal Computer Science, 9(12):1519, December 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
9_12/bad_management_
and_its.

Tovar:2012:ODP

[TPC^{+}12] E. Tovar, N. Piedra, J. Chicaiza, J. Lopez, and O. Martinez-Bonastre. OER development and promotion. outcomes of an international research projection the Open-CourseWare model. J.UCS: Journal of Universal Computer Science, 18(1): 123–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        18_1/oer_development_
                                        and_promotion.

Thackaberry:1998:EMC

[TR98] Ch. W. Thackaberry and R. Rada. Estimation metrics for courseware maintenance effort. J.UCS: Journal of Universal Computer Science, 4(3): 308–325, March 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_3/estimation_
                                        metrics_for_courseware;
                                        internal&sk=0C220489.

Thimm:2010:ISS

[TR10] H. Thimm and K. Boye Rasmussen. Information support services for intermediation tasks of collaborative networks. J.UCS: Journal of Universal Computer Science, 16 (13):1776–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        16_13/information_support_
                                        services_for.

Traore:2000:OPS

[Tra00] I. Traoré. An outline of PVS semantics for UML statecharts. J.UCS: Journal of Universal Computer Science, 6(11):1088– 1108, November 28, 2000.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
11/an_outline_of_pvs.

Trawinski:2013:EFS

[Tra13] B. Trawiński. Evolutionary fuzzy system ensemble approach to model real estate market based on data stream exploration. J.UCS: Journal of Universal Computer Science, 19(4):539– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
4/evolutionary_fuzzy_
system_ensemble.

Traetteberg:2008:IDM

[Træ08] H. Trætteberg. Integrating dialog modeling and domain modeling — the case of Diamodl and the Eclipse Modeling Framework. J.UCS: Journal of Universal Computer Science, 14(19):3265–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
19/integrating_dialog_
modeling_and.

Taweel:2006:CBR

[TRR06] A. Taweel, A. Rector, and J. Rogers. A collaborative biomedical research system. J.UCS: Journal of Universal Computer Science, 12(1):80– 98, ???? 2006. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
12_1/a_collaborative_
biomedical_research.

Trudel:2010:FCS

[Tru10] A. Trudel. Finding a consistent scenario to an interval algebra network containing possibly infinite intervals. J.UCS: Journal of Universal Computer Science, 16 (11):1425–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_11/finding_a_
consistent_scenario.

Treseler:2001:OSA

[TSCY01] H. Treseler, O. Stursberg, P. W. H. Chung, and S. Yang. An open software architecture for the verification of industrial controllers. J.UCS: Journal of Universal Computer Science, 7(1):37–53, January 28, 2001. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
7_1/an_open_software_
architecture.

Tsolis:2007:CIS

[TSDP07] D. K. Tsolis, S. Sioutas, L. Drossos, and T. S. Papatheodorou. A cultural information system providing e-commerce Web services, digital rights management and copyright protection. J.UCS: Journal of Universal Computer Science, 13(4):543– 571, ???? 2007. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_13_4/a_cultural_
information_system.

Taraghi:2015:TLA

[TSE^{+}15] B. Taraghi, A. Saranti, M. Ebner, V. Müller, and A. Großmann. Towards a learning-aware application guided by hierarchical classification of learner profiles. J.UCS: Journal of Universal Computer Science, 21(1):93– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
1/towards_a_learning_
aware.

Theng:1998:ADU

[TT98] Y. L. Theng and H. Thimbleby. Addressing design and usability issues in hypertext and on the World Wide Web by re-examining the “Lost in Hyperspace” problem. J.UCS: Journal of Universal Computer Science, 4(11):839–855, November 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_11/
addressing_design_and_
usability; internal&
sk=01237079.

Telford:2000:ETE

[TT00] A. Telford and D. Turner. Ensuring termination in ESFP. J.UCS: Journal of Universal Computer Science, 6(4):474– 488, April 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
4/ensuring_termination_
in_esfp.

Toledo:2008:LEA

[TT08] R. Toledo and É. Tanter. A lightweight and extensible AspectJ implementation. J.UCS: Journal of Universal Computer Science, 14(21):3517–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
21/a_lightweight_and_
extensible.

Toral:2009:MML

[TTB09] S. L. Toral, R. Martínez Torres, and F. Barrero. Modelling mailing list behaviour in open source projects: the case of ARM Embedded Linux. J.UCS: Journal of Universal Computer Science, 15(3):648– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
3/modelling_mailing_
list_behaviour.

Tomassetti:2013:MAS

[TTB13] F. Tomassetti, M. Torchiano, and L. Bazzani. MDD adoption in a small company: Risk management and stakeholders’ acceptance. J.UCS: Journal of Universal Computer Science, 19(2):186– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
2/mdd_adoption_in_a.

Tufekci:2013:ETO

[Tüf13] A. Tüfekci. Evaluation of Turkish ‘E-Okul’ system in terms of usability. J.UCS: Journal of Universal Computer Science, 19(5):639–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_19_5/evaluation_
                                        of_turkish_eokul.

Turner:2004:TFP

[Tur04] D. A. Turner. Total functional programming. J.UCS: Journal of Universal Computer Science, 10(7):751–768, July 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_10_7/total_
                                        functional_programming.

Thompson:2000:RCE

[TWH00] J. M. Thompson, M. W. Whalen, and M. P. E. Heimdahl. Requirements capture and evaluation in Nimbus: The light-control case study. J.UCS: Journal of Universal Computer Science, 6(7):731– 757, July 28, 2000. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_6_
                                        7/requirements_capture_
                                        and_evaluation.

Tao:2007:GAB

[TWW07] J. Tao, N. Wang, and X. Wang. Genetic algorithm based recurrent fuzzy neural network modeling of chemical processes. J.UCS: Journal of Universal Computer Science, 13(9):1332–1343, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
9/genetic_algorithm_
based_recurrent.

Tsuiki:2009:FTC

[TY09] H. Tsuiki and S. Yamada. On finite-time computability preserving conversions. J.UCS: Journal of Universal Computer Science, 15(6):1365– ??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_6/on_finite_time_
computability.

Tan:2009:MSH

[TYSY09] V. V. Tan, D.-S. Yoo, J.-C. Shin, and M.- J. Yi. A multiagent system for hierarchical control and monitoring. J.UCS: Journal of Universal Computer Science, 15(13):2485–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
13/a_multiagent_system_
for.

Urribarri:2013:GHL

[UCM13] D. K. Urribarri, S. M.

Castro, and S. R. Martig. Gyrolayout: a hyperbolic level-of-detail tree layout. J.UCS: Journal of Universal Computer Science, 19(1):132–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_1/gyrolayout_
a_hyperbolic_level.

Undrill:1997:SFD

[UDC97] P. E. Undrill, K. Delibasis, and G. G. Cameron. Stack filter design using a distributed parallel implementation of genetic algorithms. J.UCS: Journal of Universal Computer Science, 3(7):821– 834, July 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs_3_7/stack_
filter_design_using;
 internal&sk=05460486.

Uzunov:2012:ESD

[UFF12] A. V. Uzunov, E. B. Fernandez, and K. Falkner. Engineering security into distributed systems: a survey of methodologies. J.UCS: Journal of Universal Computer Science, 18(20):2920–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_

20/engineering_security_
into_distributed.

Uzunboylu:2015:DTU

[UHÖD15] H. Uzunboylu, Ç. Hürsen, G. Özütürk, and M. Demirok. Determination of Turkish university students’ attitudes for mobile integrated EFL classrooms in North Cyprus and scale development: ELLM-TAS. J.UCS: Journal of Universal Computer Science, 21(10):1283–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
10/determination_of_
turkish_university.

Ukkonen:2010:GPP

[Ukk10] E. Ukkonen. Geometric point pattern matching in the Knuth–Morris–Pratt way. J.UCS: Journal of Universal Computer Science, 16(14):1902–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_14/geometric_
point_pattern_matching.

Ulbrich:2004:HII

[UP04] A. Ulbrich and H. Pacnik. Human issues in implementing eLearning technology. J.UCS: Journal of Universal Computer Science, 10(1):1–3, January 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
 http://www.jucs.org/
jucs_10_1/human_issues_
in_implementing.

Urbieta:2013:AOA

[URG^{+}13] M. Urbieta, G. Rossi, S. E. Gordillo, A. Rodrigues, J. Araujo, and A. Moreira. An aspectoriented approach for spatial concerns in Web applications. J.UCS: Journal of Universal Computer Science, 19(1):110– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
1/an_aspect_oriented_
approach.

Laghari:2013:IRB

[uRLFH^{+}13] K. ur Rehman Laghari, T. H. Falk, M. Hyder, M. Haun, C. Hoene, and N. Crespi. An investigation into the relationship between perceived quality-of-experience and virtual acoustic environments: the case of 3D audio telephony. J.UCS: Journal of Universal Computer Science, 19 (12):1718–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_12/an_investigation_
into_the.

Uzunboylu:2016:VRT

[UT16] H. Uzunboylu and V. Tugun. Validity and reliability of tablet supported education attitude and usability scale. J.UCS: Journal of Universal Computer Science, 22(1):82–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
22_1/validity_and_reliability_
of.

Uustalu:2005:SC

[UV05] T. Uustalu and T. Vene.

Signals and comonads. J.UCS: Journal of Universal Computer Science, 11(7):1311–1326, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
7/signals_and_comonads.

Uzunboylu:2013:IIT

[Uzu13] H. Uzunboylu. Innovative instructional technologies. J.UCS: Journal of Universal Computer Science, 19(5):600–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_19;   http:
                                        //www.jucs.org/jucs_
                                        19_5/innovative_instructional_
                                        technologies.

Valkeapaa:2007:AFO

[VAH07] O. Valkeapää, O. Alm, and E. Hyvönen. An adaptable framework for ontology-based content creation on the Semantic Web. J.UCS: Journal of Universal Computer Science, 13(12):1835–1853, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_
                                        12/an_adaptable_framework_
                                        for.

Vaida:2000:NPO

[Vai00]  D. Vaida.  Notes on

partially-ordered structures in computer science: I. PA-ordered semirings and some related structures. J.UCS: Journal of Universal Computer Science, 6(1):201–211, January 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_6_
                                        1/notes_on_partially_
                                        ordered.

Vara:2012:TMD

[VAPM12] J. M. Vara, V. Andrikopoulos, M. P. Papazoglou, and E. Marcos. Towards modeldriven engineering support for service evolution. J.UCS: Journal of Universal Computer Science, 18(17):2364–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
18_17/towards_model_
driven_engineering.

VasudevanR:2005:NSS

[VAS05] A. Vasudevan R., A. Abraham, and S. Sanyal. A novel scheme for secured data transfer over computer networks. J.UCS: Journal of Universal Computer Science, 11(1):104–121, January 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
1/a_novel_scheme_for.

vandeRiet:1996:LTM

[vB96] R. P. van de Riet and J. F. M. Burg. Linguistic tools for modelling alter egos in cyberspace: Who is responsible? J.UCS: Journal of Universal Computer Science, 2(9):623– 636, September 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        2_9/linguistic_tools_
                                        for_modelling.

Vincini:2013:SIH

[VBB13] M. Vincini, D. Beneventano, and S. Bergamaschi. Semantic integration of heterogeneous data sources in the MOMIS data transformation system. J.UCS: Journal of Universal Computer Science, 19(13):1986–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_19_13/semantic_
                                        integration_of_heterogeneous.

vanBiljon:2008:CFM

[vBK08] J. van Biljon and P. Kotzé. Cultural factors in a mobile phone adoption and usage model. J.UCS: Journal of Universal Computer Science, 14 (16):2650–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_16/cultural_
                                        factors_in_a.

Vialardi:2008:IAC

[VBO08] C. Vialardi, J. Bravo, and A. Ortigosa. Improving AEH courses through log analysis. J.UCS: Journal of Universal Computer Science, 14(17): 2777–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_17/improving_
aeh_courses_through.

Vahey:2011:ICI

[VBP^{+}11] P. Vahey, J. Brecht, C. Patton, K. Rafanan, and B. Haugan Cheng. Investigating collaborative innovation in a virtual world task. J.UCS: Journal of Universal Computer Science, 17(12): 1638–??, ???? 2011. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_12/investigating_
collaborative_innovation_
in.

Vazquez-Briseno:2012:AMF

[VBVNHdDSL12] M. Vazquez-Briseno, P. Vincent, J. I. Nieto-Hipolito, and J. de Dios Sanchez-Lopez. Applying a modular framework to develop mobile applications and services. J.UCS: Journal of Universal Computer Science, 18(5):704– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
5/applying_a_modular_
framework.

Villalon:2013:DAS

[VC13] J. Villalon and R. A. Calvo. A decoupled architecture for scalability in text mining applications. J.UCS: Journal of Universal Computer Science, 19(3):406– ??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_19_
                                        3/a_decoupled_architecture_
                                        for.

Veit:2008:CTH

[VCB08] M. Veit, A. Capobianco, and D. Bechmann. Consequence of two-handed manipulation on speed, precision and perception on spatial input task in 3D modelling applications. J.UCS: Journal of Universal Computer Science, 14(19):3174–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_19/consequence_of_
                                        two-handed.

vanDalen:2005:HMO

[vD05] D. van Dalen. How the mathematical objects determine the mathematical principles. J.UCS: Journal of Universal Computer Science, 11(12): 2132–2141, ???? 2005.

CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_12/how_the_mathematical_
objects.

Volbracht:1998:CGE

[VDBNR98] S. Volbracht, G. Domik, D. Backe-Neuwald, and H-D. Rinkens. The city game: An example of a virtual environment for teaching spatial orientation. J.UCS: Journal of Universal Computer Science, 4(4):461– 465, April 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_4/the_
city_game_an;   internal&
sk=0C220489.

Verbert:2010:CAR

[VDLG10] K. Verbert, E. Duval, S. N. Lindstaedt, and D. Gillet. Contextaware recommender systems. J.UCS: Journal of Universal Computer Science, 16(16):2175–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_16;   http:
//www.jucs.org/jucs_
16_16/context_aware_
recommender_systems;
 http://www.jucs.org/
jucsrssfeed-issue.

Visaggio:2009:SMS

[VdR09] C. A. Visaggio and F. de Rosa. A system for managing security knowledge using case based reasoning and misuse cases. J.UCS: Journal of Universal Computer Science, 15(15): 3059–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_15_15/a_system_
                                        for_managing.

Vidal:2008:SBC

[VdSdMC08] M. Vidal, A. S. da Silva, E. S. de Moura, and J. M. Cavalcanti. Structurebased crawling in the hidden Web. J.UCS: Journal of Universal Computer Science, 14(11): 1857–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_11/structure_
                                        based_crawling_in.

Volbrach:1998:ETO

[VDSF98] S. Volbrach, G. Domik, K. Shahrbabaki, and G. Fels. An example of task oriented empirical evaluations of 3D-display modes. J.UCS: Journal of Universal Computer Science, 4(5):534– 546, May 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_5/an_
example_of_task;   internal&
sk=0C220489.

vanderVeer:2008:CEI

[vdV08] G. C. van der Veer. Cognitive ergonomics in interface design — discussion of a moving science. J.UCS: Journal of Universal Computer Science, 14(16):2614–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_16/cognitive_
ergonomics_in_interface_
design.

Velev:2004:TSF

[Vel04] M. N. Velev. Tuning SAT for formal verification and testing. J.UCS: Journal of Universal Computer Science, 10(12):1559–1561, December 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/free;
http://www.jucs.org/
jucs_10_12/tuning_sat_
for_formal.

Veldman:2005:PIV

[Vel05] W. Veldman. Perhaps the intermediate value theorem. J.UCS: Journal of Universal Computer Science, 11(12):2142–2158, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
12/perhaps_the_intermediate_
value.

Verna:2008:BMP

[Ver08] D. Verna. Binary methods programming: the Clos perspective. J.UCS: Journal of Universal Computer Science, 14(20): 3389–??, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_20/binary_methods_
programming_the.

Verna:2010:RVJ

[Ver10] D. Verna. Revisiting the visitor: the “Just Do It” pattern. J.UCS: Journal of Universal Computer Science, 16(2):246– ??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_2/revisiting_the_
visitor_the.

Vasilieva:2011:NQA

[VF11] A. Vasilieva and R. Freivalds. Nondeterministic query algorithms. J.UCS: Journal of Universal Computer Science, 17(6):859– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
6/nondeterministic_query_
algorithms.

Vasconcellos:2003:PTI

[VFC03] C. Vasconcellos, L. Figueiredo, and C. Camarao. Practical type inference for polymorphic recursion: an implementation in Haskell. J.UCS: Journal of Universal Computer Science, 9(8):873– 890, August 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
8/practical_type_inference_
for.

Vega-Gorgojo:2015:RLD

[VGAPGS^{+}15] G. Vega-Gorgojo, J. I.

Asensio-Pérez, E. Gómez-Sánchez, M. L. Bote-Lorenzo, J. A. Munoz-Cristobal, and A. Ruiz-Calleja. A review of linked data proposals in the learning domain. J.UCS: Journal of Universal Computer Science, 21(2):326– ??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
2/a_review_of_linked.

Vega-Gorgojo:2008:OSE

[VGBLGS^{+}08] G. Vega-Gorgojo, M. L. Bote-Lorenzo, E. Gómez-Sánchez, J. I. Asensio-Pérez, Y. A. Dimitriadis, and I. M. Jorrín-Abellán. Ontoolcole: Supporting educators in the semantic search of CSCL tools. J.UCS: Journal of Universal Computer Science, 14(1):27–58, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_1/ontoolcole_
                                        supporting_educators_
                                        in.

Valencia-Garcia:2016:NTO

[VGCPAH16] R. Valencia-García, R. Colomo-Palacios, and G. Alor-Hernández. New trends in opinion mining technologies in the industry. J.UCS: Journal of Universal Computer Science, 22(5):605–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_22;   http:
                                        //www.jucs.org/jucs_
                                        22_5/new_trends_in_opinion.

Vega-Gorgojo:2012:RLS

[VGGSBLAP12] G. Vega-Gorgojo, E. Gómez-Sánchez, M. L. Bote-Lorenzo, and J. I. Asensio-Pérez. RESTifying a legacy semantic search system: Experience and lessons learned. J.UCS: Journal of Universal Computer Science, 18 (2):286–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_2/restifying_
a_legacy_semantic.

Viedma:2003:SSC

[Vie03] J. M. Viedma. SCBS social capital benchmarking system — profiting from social capital when building network organisations. J.UCS: Journal of Universal Computer Science, 9(6):501–509, June 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_9_6/scbs_social_
capital_benchmarking.

Vita:2005:CSE

[Vît05] L. S. Vîta. On complements of sets and the Efremovic condition in pre-apartness spaces. J.UCS: Journal of Universal Computer Science, 11(12):2159–2164, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_

                                        12/on_complements_of_
                                        sets.

Vivet:1996:COL

[Viv96] M. Vivet. The classroom as ONE learning environment of the future. J.UCS: Journal of Universal Computer Science, 2(10):663–678, October 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_2_
                                        10/the_classroom_as_
                                        one.

Vavilin:2009:GBA

[VJ09] A. Vavilin and K.-H. Jo. Graph-based approach for robust road guidance sign recognition from differently exposed images. J.UCS: Journal of Universal Computer Science, 15(4):786–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_15_4/graph_
                                        based_approach_for.

Vujosevic-Janicic:2007:RGS

[VJTJ07] M. Vujosević-Janičić, J. Tomasević, and P. Janičić. Random k-GD-Sat model and its phase transition. J.UCS: Journal of Universal Computer Science, 13(4):572–591, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
13_4/random_k_gd_sat.

Virvou:2003:ESW

[VK03] M. Virvou and K. Kabassi.

Experimental studies within the software engineering process for intelligent assistance in a GUI. J.UCS: Journal of Universal Computer Science, 9(1):51–85, January 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
1/experimental_studies_
within_the.

vanderPoll:2004:ASE

[vKL04] J. A. van der Poll, P. Kotze, and W. A. Labuschagne. Automated support for enterprise information systems. J.UCS: Journal of Universal Computer Science, 10(11): 1519–1539, November 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
11/automated_support_
for_enterprise.

Volk:2015:LAE

[VKW15] H. Volk, K. Kellner, and D. Wohlhart. Learning analytics for English language teaching. J.UCS: Journal of Universal Computer Science, 21(1):156–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
1/learning_analytics_
for_english.

Vo:2014:SML

[VL14] G. N. Vo and R. Lai. A secure multi-layer edocument method for improving e-government processes. J.UCS: Journal of Universal Computer Science, 20(11): 1583–??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
20_11/a_secure_multi_
layer.

Vartiainen:2012:DOP

[VLE12] H. Vartiainen, A. Liljeström, and J. Enkenberg. Design-oriented pedagogy for technologyenhanced learning to cross over the borders between formal and informal environments. J.UCS: Journal of Universal Computer Science, 18(15):2097–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
15/design_oriented_pedagogy_
for.

Vleju:2014:AAC

[Vle14]  M. B. Vleju.  Au-

tomatic authentication to cloud-based services. J.UCS: Journal of Universal Computer Science, 20(3):385–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_20_3/automatic_
authentication_to_cloud.

Vidal:2013:CAR

[VM13] S. A. Vidal and C. Marcos. A catalog of aspect refactorings for Spring/ AOP. J.UCS: Journal of Universal Computer Science, 19(1):157–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_1/a_catalog_
of_aspect_refactorings.

Varajao:2014:EIS

[VMA14] J. Varajão, R. Martinho, and P. Soto Acosta. Enterprise information systems. J.UCS: Journal of Universal Computer Science, 20(6):822–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.

                                        org/jucs_20;   http:
                                        //www.jucs.org/jucs_
                                        20_6/enterprise_information_
                                        systems.

Valente:2008:SXB

[VMdCJ08] M. T. Valente, P. D. Mosses, and F. Heron de Carvalho Jr. SBLP 2008: XII Brazilian Symposium on Programming. J.UCS: Journal of Universal Computer Science, 14(21):3412–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14;
                                         http://www.jucs.org/
                                        jucs_14_21#;   http:
                                        //www.jucs.org/jucs_
                                        14_21/sblp_2008.

Vega:2014:UCV

[VMFO14] D. Vega, R. Meseguer, F. Freitag, and S. F. Ochoa. Understanding collaboration in volunteer computing systems. J.UCS: Journal of Universal Computer Science, 20(13):1738–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        20_13/understanding_
                                        collaboration_in_volunteer.

Viroli:2002:MAO

[VO02] M. Viroli and A. Omicini.

Modelling agents as observable sources. J.UCS: Journal of Universal Computer Science, 8(4): 423–452, April 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
8_4/modelling_agents_
as_observable.

Villegas:2010:EMC

[VO10] A. Villegas and A. Olivé.

Extending the methods for computing the importance of entity types in large conceptual schemas. J.UCS: Journal of Universal Computer Science, 16(20):3138–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_20/extending_
the_methods_for.

vonGudenberg:1998:JSC

[von98] J. W. von Gudenberg.

Java for scientific computing, pros and cons. J.UCS: Journal of Universal Computer Science, 4(1):11–15, January 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_1/java_
for_scientific_computing.

vonPlato:2005:CAS

[vP05] J. von Plato. A constructive approach to Sylvester’s conjecture. J.UCS: Journal of Universal Computer Science, 11(12):2165–2178, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        11_12/a_constructive_
                                        approach_to.

Valcarce:2015:DRP

[VPB15] D. Valcarce, J. Parapar, and Á. Barreiro. A distributed recommendation platform for Big Data. J.UCS: Journal of Universal Computer Science, 21(13):1810–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        21_13/a_distributed_
                                        recommendation_platform.

Vajda:2009:LAC

[VPF09] S. Vajda, T. Plötz, and G. A. Fink. Layout analysis for camera-based whiteboard notes. J.UCS: Journal of Universal Computer Science, 15 (18):3307–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_

15_18/layout_analysis_
for_camera.

Vega-Rodriguez:2005:CST

[VRGPSP05] M. A. Vega-Rodríguez, J. A. Gómez-Pulido, and J. M. Sánchez-Pérez. Case studies in teleeducation: Research and projects. J.UCS: Journal of Universal Computer Science, 11(9):1555– 1567, ???? 2005. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_9/case_studies_in_
teleeducation.

Vega-Rodriguez:2007:NAR

[VRGPSP07] M. A. Vega-Rodríguez, J. A. Gómez-Pulido, and J. M. Sánchez-Pérez. New advances in reconfigurable computing and its applications. J.UCS: Journal of Universal Computer Science, 13(3):345– 348, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13;
http://www.jucs.org/
jucs_13_3#;   http://
www.jucs.org/jucs_13_
3/new_advances_in_reconfigurable.

Veerubhotla:2005:GCF

[VSGP05] R. S. Veerubhotla, A. Saxena, V. P. Gulati, and A. K. Pujari. Gossip codes for fingerprinting: Construction, erasure analysis and pirate tracing. J.UCS: Journal of Universal Computer Science, 11(1):122–149, January 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_11_1/gossip_
codes_for_fingerprinting.

Volz:2003:PBI

[VSML03] R. Volz, R. Studer, A. Maedche, and B. Lauser. Pruning-based identification of domain ontologies. J.UCS: Journal of Universal Computer Science, 9(6):520–529, June 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_9_6/pruning_
based_identification_
of.

vanStaden:2011:CPC

[vSO11] W. van Staden and M. S. Olivier. On compound purposes and compound reasons for enabling privacy. J.UCS: Journal of Universal Computer Science, 17(3):426– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
3/on_compound_purposes_
and.

Villanueva:2013:FDW

[VTGA13] P. G. Villanueva, R. Tesoriero, J. A. Gallud, and A. H. Altalhi. A framework to develop Web applications based on RFID panels. J.UCS: Journal of Universal Computer Science, 19(12):1792–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
19_12/a_framework_to_
develop.

Vo:2016:USS

[VTHM16] B. Vo, T. Tran, T.-P.

Hong, and N. L. Minh. Using soft set theory for mining maximal association rules in text data. J.UCS: Journal of Universal Computer Science, 22(6):802–??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_22_6/using_soft_
set_theory.

Venkatesan:2012:APS

[VTRN12] A. P. Venkatesan, D. Thomas, T. Robinson, and A. K. Nagar. Array P system with shuffle on trajectories. J.UCS: Journal of Universal Computer Science, 18(13): 1802–??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        18_13/array_p_system_
                                        with.

Virginas:2008:IRE

[VUT^{+}08] B. Virginas, M. Ursu, E. Tsang, G. Owusu, and C. Voudouris. Intelligent resource exchanges: Solutions and pathways in a workforce allocation problem. J.UCS: Journal of Universal Computer Science, 14(14):2343–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        14/intelligent_resource_
                                        exchanges_solutions.

Varga:2006:CSL

[VV06] K. P. Varga and M. Várterész. Computer science, logic, informatics education. J.UCS: Journal of Universal Computer Science, 12(9):1405–1410, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12_
                                        9/computer_science_logic_
                                        informatics.

Vural:2012:CMS

[VV12] I. Vural and H. S. Venter. Combating mobile spam through botnet detection using artificial immune systems. J.UCS: Journal of Universal Computer Science, 18(6):750–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_6/combating_
mobile_spam_through.

Vargas-Vera:2015:FER

[VV15]  M. Vargas-Vera.  A

framework for extraction of relations from text using relational learning and similarity measures. J.UCS: Journal of Universal Computer Science, 21(11):1482–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_21_
11/a_framework_for_extraction.

Villanueva:2006:CAQ

[VVM^{+}06] F. J. Villanueva, D. Villa, F. Moya, J. Barba, F. Rincón, and J. C. López. Context-aware QoS provision for mobile ad-hoc network-based ambient intelligent environments. J.UCS: Journal of Universal Computer Science, 12(3):315– 327, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_

                                        3/context_aware_qos_
                                        provision.

vanZyl:2012:HEF

[vZdlH12] I. van Zyl and R. de la Harpe. AT-HOME 2.0 — an educational framework for home-based healthcare. J.UCS: Journal of Universal Computer Science, 18(3):429–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_18_3/at_home_2.
                                        0_an.

vanZijl:2011:DCA

[vZG11] L. van Zijl and J. Geldenhuys. Descriptional complexity of ambiguity in symmetric difference NFAs. J.UCS: Journal of Universal Computer Science, 17(6):874– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_17_
                                        6/descriptional_complexity_
                                        of_ambiguity.

Wactlar:2002:EVK

[Wac02] H. D. Wactlar. Extracting and visualizing knowledge from film and video archives. J.UCS: Journal of Universal Computer Science, 8(6):602– 612, June 28, 2002. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_8_6/extracting_
and_visualizing_knowledge.

Wang:1995:MPT

[Wan95]  Dongming Wang.  A

method for proving theorems in differential geometry and mechanics. J.UCS: Journal of Universal Computer Science, 1(9):658–673, September 28, 1995. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
a_method_for_proving_
theorems_in_differential_
geometry_and_mechanics.

Wastl:1998:LDK

[Was98a] R. Wastl. Linear derivations for keys of a database relation schema. J.UCS: Journal of Universal Computer Science, 4(12):883, December 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_4_12/linear_
derivations_for_keys.

Wastl:1998:NKR

[Was98b] R. Wastl. On the number of keys of a relational database schema. J.UCS: Journal of Universal Computer Science, 4(5):

547–559, May 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_5/on_
                                        the_number_of;   internal&
                                        sk=0C220489.

Watson:2002:FSA

[Wat02] B. W. Watson. A fast and simple algorithm for constructing minimal acyclic deterministic finite automata. J.UCS: Journal of Universal Computer Science, 8(2):363– 367, February 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_8_
                                        2/a_fast_and_simple.

Woodcock:2007:VGC

[WB07] J. Woodcock and R. Banach. The Verification Grand Challenge. J.UCS: Journal of Universal Computer Science, 13(5):661–668, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        13_5/the_verification_
                                        grand_challenge.

Wang:2012:PDE

[WBS12] J. Wang, C. Berger, and N. Szilas. Pedagogical design of an eTandem Chinese–French writing course. J.UCS: Journal of Universal Computer Science, 18(3):393– ??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
3/pedagogical_design_
of_an.

Wu:2008:RRS

[WC08] S. Wu and F. Crestani.

Ranking retrieval systems with partial relevance judgements. J.UCS: Journal of Universal Computer Science, 14(7):1020– 1030, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
7/ranking_retrieval_
systems_with.

Wang:2009:EDP

[WC09] J.-S. Wang and J.-C. Chiang. An efficient data preprocessing procedure for support vector clustering. J.UCS: Journal of Universal Computer Science, 15(4):705– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
4/an_efficient_data_
preprocessing.

Wu:2014:ELE

[WCH14] Y.-C. J. Wu, C.-L. Chang, and Y.-J. Hsieh. Enhancing learning experience of the disabled: An accessible tourism platform. J.UCS: Journal of Universal Computer Science, 20(15):2080–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_20_15/enhancing_
                                        learning_experience_
                                        of.

Westbomke:2002:TXB

[WD02] J. Westbomke and G. Dittrich. Towards an XMLbased implementation of structured hypermedia documents. J.UCS: Journal of Universal Computer Science, 8(10):944– 956, October 28, 2002. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_8_10/towards_an_
                                        xlm_based.

Weihrauch:2008:CMF

[Wei08] K. Weihrauch. The computable multi-functions on multi-represented sets are closed under programming. J.UCS: Journal of Universal Computer Science, 14(6):801– 844, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_6/the_computable_
multi_functions.

Weihrauch:2010:CST

[Wei10]  K. Weihrauch.  Com-

putable separation in topology, from T_{0} to T_{2}. J.UCS: Journal of Universal Computer Science, 16(18):2733–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
18/computable_separation_
in_topology.

Wang:2012:FOC

[WF12] Q. Wang and E. Fleury.

Fuzziness and overlapping communities in large-scale networks. J.UCS: Journal of Universal Computer Science, 18(4):457– ??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_4/fuzziness_
and_overlapping_communities.

Wilke:1998:MMU

[WFOC98] F. Z. Wilke, B. R. T.

Franciosi, P. W. Oliveira, and D. M. Claudio. Modelling the measurement uncertainty by intervals. J.UCS: Journal of Universal Computer Science, 4(1):82–88, January 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs/jucs_4_1/modelling_
                                        the_measurement_uncertainty.

Weihrauch:2009:ECT

[WG09] K. Weihrauch and T. Grubba. Elementary computable topology. J.UCS: Journal of Universal Computer Science, 15(6):1381– ??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_15_6/elementary_
                                        computable_topology.

Wang:2008:NML

[WH08] X.-F. Wang and D.-S. Huang. A novel multilayer level set method for image segmentation. J.UCS: Journal of Universal Computer Science, 14(14):2427–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_14_14/a_novel_multi_
                                        layer.

Wang:2009:PPN

[WHCL09] X. Wang, M. Hao, Y. Cheng, and R. Lei. PDE-PEDA: a new Pareto-based multiobjective optimization algorithm. J.UCS: Journal of Universal Computer Science, 15(4):722– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
4/pde_peda_a_new.

Westphal:2004:SSD

[WHD04] Brian T. Westphal, Frederick C. Harris, Jr., and Sergui M. Dascalu. Snippets: Support for dragand-drop programming in the Redwood Environment. J.UCS: Journal of Universal Computer Science, 10(7):859–871, July 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_10_
7/snippets_support_for_
drag.

Wiederhold:2008:DSI

[Wie08] G. Wiederhold. Determining software investment lag. J.UCS: Journal of Universal Computer Science, 14(22):3737–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
22/determining_software_
investment_lag%20.

Winkler:1997:TCE

[Win97a] J. F. H. Winkler. Type compatibility for extensible module types, their reference parameters, and their pointer types. J.UCS: Journal of Universal Computer Science, 3(2):120– 146, February 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_3_
                                        2/type_compatibility.

Winter:1997:MCA

[Win97b] K. Winter. Model checking for abstract state machines. J.UCS: Journal of Universal Computer Science, 3(5):689– 701, May 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //medoc.springer.de:
                                        8000/jucs_3_5/model_
                                        checking;   http://www.
                                        jucs.org/jucs_3_5/model_
                                        checking; internal&sk=
                                        05460486.

Witten:2008:SLW

[Wit08] I. H. Witten. Searching … in a Web. J.UCS: Journal of Universal Computer Science, 14 (10):1739–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_

14_10/searching_in_a_
web.

Woitsch:2005:POK

[WK05] R. Woitsch and D. Karagiannis. Process oriented knowledge management: a service based approach. J.UCS: Journal of Universal Computer Science, 11(4):565–588, April 28, 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
4/process_oriented_knowledge_
management.

Wulf:1999:IIO

[WKS^{+}99] V. Wulf, M. Krings, O. Stiemerling, G. Iacucci, M. Maidhof, R. Peters, P. Fuchs-Fronhofen, B. Nett, and J. Hinrichs. Improving interorganizational processes with integrated organization and technology development. J.UCS: Journal of Universal Computer Science, 5(6):339– 365, June 28, 1999. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
5_6/improving_inter_
organizational_processes.

Wells:2011:ASM

[WKSD^{+}11] W. H. Wells, W. Karwowski, S. Sala-Diakanda, K. Williams, T. Ahram, and J. A. Pharmer. Application of systems modeling language (SySML) for cognitive work analysis in systems engineering design process. J.UCS: Journal of Universal Computer Science, 17(9):1261–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
17_9/application_of_
systems_modeling.

Wible:2001:OWP

[WKTL01] D. Wible, C. H. Kuo, N. L. Tsao, and A. Liu. An online writing platform for language teachers. J.UCS: Journal of Universal Computer Science, 7(3):278–289, March 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
3/an_online_writing_
platform.

Wei:2005:ILW

[WKXL05] L. Wei, E. Keogh, X. Xi, and S. Lonardi. Integrating lite-weight but ubiquitous data mining into GUI operating systems. J.UCS: Journal of Universal Computer Science, 11(11):1820–1834, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
11/integrating_lite_
weight_but.

Wasser:2013:BPM

[WL13] A. Wasser and M. Lincoln. Business process management applications based on semantic process models: the ProcessGene Suite casestudy. J.UCS: Journal of Universal Computer Science, 19(13):1892–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
13/business_process_
management_applications.

Wang:2011:SEI

[WLKW11] X. Wang, P. E. Love, M. J.

Kim, and L. Wang. Studying the effects of information exchange channels in different communication modes on trust building in computer-mediated remote collaborative design. J.UCS: Journal of Universal Computer Science, 17(14):1971–??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_17_14/studying_
the_effects_of.

Wang:2009:DSM

[WLLL09] J. Wang, J. Lü, S. Lian, and G. Liu. On the design of secure multimedia authentication. J.UCS: Journal of Universal Computer Science, 15(2):426–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        15_2/on_the_design_of.

Wynn:2014:FCA

[WLtHN14] M. T. Wynn, W. Z. Low, A. H. M. ter Hofstede, and W. Nauta. A framework for cost-aware process management: Cost reporting and cost prediction. J.UCS: Journal of Universal Computer Science, 20(3):406– ??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        20_3/a_framework_for_
                                        cost.

Wu:2007:MLB

[WMJ^{+}07] C. Wu, M. Marchese, J. Jiang, A. Ivanyukovich, and Y. Liang. Machine learning-based keywords extraction for scientific literature. J.UCS: Journal of Universal Computer Science, 13(10):1471– 1483, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
10/machine_learning_
based_keywords.

Wu:2009:CBS

[WMSH09] W. Wu, Y. Mu, W. Susilo, and X. Huang. Certificatebased signatures revisited. J.UCS: Journal of Universal Computer Science, 15(8):1659–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
8/certificate_based_
signatures_revisited.

Werner:2009:SCA

[WO09] C. Werner and F. Oquendo.

Software components, architectures and reuse. J.UCS: Journal of Universal Computer Science, 15(11):2138–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_15;   http:
//www.jucs.org/jucs_
15_11#;   http://www.
jucs.org/jucs_15_11/
software_components_
architectures_and.

Wolff:1999:CIC

[Wol99a] G. Wolff. Computing as information compression by multiple alignment, unification and search. J.UCS: Journal of Universal Computer Science, 5(11):777– 815, November 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_5_11/computing_
as_information_compression.

Wolff:1999:PRI

[Wol99b] J. G. Wolff. Probabilistic reasoning as information compression by multiple alignment, unification and search: An introduction and overview. J.UCS: Journal of Universal Computer Science, 5(7):418–462, July 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_5_
7/probabilistic_reasoning_
as_information.

Wolff:2000:SPP

[Wol00] J. G. Wolff. Syntax, parsing and production of natural language in a framework of information compression by multiple alignment, unification and search. J.UCS: Journal of Universal Computer Science, 6(8):781–829, August 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
8/syntax_parsing_and_
production.

Won:2003:SKK

[WP03] M. Won and V. Pipek.

Sharing knowledge on knowledge — the eXact peripheral expertise awareness system. J.UCS: Journal of Universal Computer Science, 9(12):1388–1397, December 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_9_
12/sharing_knowledge_
on_knowledge.

Wendzel:2015:CME

[WP15] S. Wendzel and C. Palmer.

Creativity in mind: Evaluating and maintaining advances in network steganographic research. J.UCS: Journal of Universal Computer Science, 21(12):1684–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
21_12/creativity_in_
mind_evaluating.

Williams:1998:EUP

[WPL98] D. C. Williams, S. Pedersen, and M. Liu. An evaluation of the use of problem-based learning software by middle school students. J.UCS: Journal of Universal Computer Science, 4(4):466– 483, April 28, 1998. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_4/an_
evaluation_of_the;
internal&sk=0C220489.

Wurdel:2008:CDT

[WSF08] M. Wurdel, D. Sinnig, and P. Forbrig. CTML: Domain and task modeling for collaborative environments. J.UCS: Journal of Universal Computer Science, 14(19):3188–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_14_19/ctml_
domain_and_task.

Wang:2007:ISB

[WSL07] C. Wang, Y. Sun, and Y. Liang. An improved SVM based on similarity metric. J.UCS: Journal of Universal Computer Science, 13(10):1462–1470, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
10/an_improved_svm_based.

Wallace:2001:ASM

[WTA01] C. Wallace, G. Tremblay, and J. N. Amaral. An abstract state machine specification and verification of the location consistency memory model and cache protocol. J.UCS: Journal of Universal Computer Science, 7(11):1088– 1112, November 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
11/an_abstract_state_
machine.

Wurzinger:2010:ICL

[Wur10]  G. Wurzinger.  Infor-

mation consolidation in large bodies of information. J.UCS: Journal of Universal Computer Science, 16(21):3314–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_16_
21/information_consolidation_
in_large.

Wang:2015:EIS

[WWD15] H.-W. Wang, Y.-C. J.

Wu, and T.-P. Dong. Exploring the impacts of social networking on brand image and purchase intention in cyberspace. J.UCS: Journal of Universal Computer Science, 21(11): 1425–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_21_11/exploring_
                                        the_impacts_of.

Wang:2015:RSB

[WXZL15] X. Wang, L. Xu, S. Zhou, and J. K. Liu. A resolving set based algorithm for fault identification in wireless mesh networks. J.UCS: Journal of Universal Computer Science, 21(3):384–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        21_3/a_resolving_set_
                                        based.

Wang:2007:ACI

[WZC07] Y. Wang, Z. Zhang, and J. Cui. The architecture and circuital implementation scheme of a new cell neural network for analog signal processing. J.UCS: Journal of Universal Computer Science, 13(9):1344– 1353, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_
                                        9/the_architecture_and_
                                        circuital.

Weng:2009:JWR

[WZZ^{+}09] L. Weng, Y. Zhang, Y. Zhou, L. T. Yang, P. Tian, and M. Zhong. A joint Web resource recommendation method based on category tree and associate graph. J.UCS: Journal of Universal Computer Science, 15(12): 2387–??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_12/a_joint_web_resource.

Xie:2013:OCS

[XCJ13] C. Xie, H. Cai, and L. Jiang. Ontology combined structural and operational semantics for resource-oriented service composition. J.UCS: Journal of Universal Computer Science, 19 (13):1963–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_19_13/ontology_
combined_structural_
and.

Xiong:2009:NNB

[XHP^{+}09] N. Xiong, J. He, J. H.

Park, D. Cooley, and Y. Li. A neural network based vehicle classification system for pervasive smart road security. J.UCS: Journal of Universal Computer Science, 15(5):1119–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15_
                                        5/a_neural_network_based.

Xi:2003:DTP

[Xi03]  H. Xi.  Dependently

typed pattern matching. J.UCS: Journal of Universal Computer Science, 9(8):851–872, August 28, 2003. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        8/dependently_typed_
                                        pattern_matching.

Xu:2010:SVI

[XLMR10] Y. Xu, J. Liu, L. Martínez, and D. Ruan. Some views on information fusion and logic based approaches in decision making under uncertainty. J.UCS: Journal of Universal Computer Science, 16(1):3–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_16_1/some_views_
                                        on_information.

Xie:2013:ECP

[XMLC13] X. Xie, H. Ma, J. Li, and X. Chen. An efficient ciphertext-policy attribute-based access control towards revocation in cloud computing. J.UCS: Journal of Universal Computer Science, 19(16):2349–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_19_
16/an_efficient_ciphertext_
policy.

Xu:2010:HAP

[XMZbL10] M. Xu, C. Mu, Z. Zeng, and Z. b. Li. A heuristic approach to positive root isolation for multiple power sums. J.UCS: Journal of Universal Computer Science, 16(14): 1912–??, ???? 2010. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_14/a_heuristic_approach_
to.

Xia:2015:SPK

[XNKG15] Q. Xia, J. Ni, A. J. B. A.

Kanpogninge, and J. C. Gee. Searchable publickey encryption with data sharing in dynamic groups for mobile cloud storage. J.UCS: Journal of Universal Computer Science, 21(3):440–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_21_3/searchable_
public_key_encryption.

Xue:2009:TSM

[XWGS09] Y. Xue, C. Wang, H. H.

Ghenniwa, and W. Shen. A tree similarity measuring method and its application to ontology comparison. J.UCS: Journal of Universal Computer Science, 15(9):1766– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
9/a_tree_similarity_
measuring.

Xue:2000:ERW

[XZ00] K. Xue and M. Zimand. Extractors for the real world. J.UCS: Journal of Universal Computer Science, 6(1):212–225, January 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_6_
1/extractors_for_the_
real.

Xiang:2009:POW

[XZSS09] Y. Xiang, S. Zhang, Y. Shen, and M. Shi. Pattern-oriented workflow generation and optimization. J.UCS: Journal of Universal Computer Science, 15(9):1924–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
9/pattern_oriented_workflow_
generation.

Yan:2005:NAS

[Yan05] W. Yan. Network attack scenarios extraction and categorization by mining IDS alert streams. J.UCS: Journal of Universal Computer Science, 11(8):1367– 1382, ???? 2005. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
11_8/network_attack_
scenarios_extraction.

Yeung:2004:CCB

[Yeu04] W. L. Yeung. Checking consistency between UML class and state models based on CSP and B. J.UCS: Journal of Universal Computer Science, 10(11):1540–1558, November 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_10_11/checking_
consistency_between_
uml.

Yang:2014:EPE

[YJY14] P. Yang, P. Jin, and L. Yue. Exploiting the performance-energy tradeoffs for mobile database applications. J.UCS: Journal of Universal Computer Science, 20 (10):1488–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        20_10/exploiting_the_
                                        performance_energy.

Yassein:2016:FSB

[YKA16] M. B. Yassein, Y. Khamayseh, and M. AbuJazoh. Feature selection for black hole attacks. J.UCS: Journal of Universal Computer Science, 22(4):521– ??, ???? 2016. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_22_
                                        4/feature_selection_
                                        for_black.

Yim:2008:IPP

[YKD^{+}08] J. Yim, I. Ko, J. Do, J. Joo, and S. Jeong. Implementation of a prototype positioning system for LBS on Ucampus. J.UCS: Journal of Universal Computer Science, 14(14):2381–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_14_
                                        14/implementation_of_
                                        a_prototype.

Yang:2007:CFS

[YLL^{+}07] J. Yang, T. Li, S. Liu, T. Wang, D. Wang, and G. Liang. Computer forensics system based on artificial immune systems. J.UCS: Journal of Universal Computer Science, 13(9):1354–1365, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
9/computer_forensics_
system_based.

Yang:2014:MCO

[YLW^{+}14] M. Yang, B. Liu, W. Wang, J. Luo, and X. Shen. Maximum capacity overlapping channel assignment based on max-cut in 802.11 wireless mesh networks. J.UCS: Journal of Universal Computer Science, 20(13):1855–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
13/maximum_apacity_overlapping_
channel.

Yu:2010:SOP

[YLZW10] Z. Yu, C. Li, X. Zhou, and H. Wang. A service-oriented platform for ubiquitous personalized multimedia provisioning. J.UCS: Journal of Universal Computer Science, 16(10): 1291–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/
                                        jucs_16_10/a_service_
                                        oriented_platform.

Yu:2015:RAS

[YMA15] Y. Yu, Y. Mu, and G. Ateniese. Recent advances in security and privacy in Big Data. J.UCS: Journal of Universal Computer Science, 21(3):365– ??, ???? 2015. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_21;   http:
                                        //www.jucs.org/jucs_
                                        21_3/recent_advances_
                                        in_security.

Yang:2008:FAS

[YMP08] M. Yang, G. J. Michaelson, and R. J. Pooley. Formal action semantics for a UML action language. J.UCS: Journal of Universal Computer Science, 14(21):3608–??, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        14_21/formal_action_
                                        semantics_for.

Yong:2011:SPP

[Yon11]  J. Yong.  Security

and privacy preservation for mobile E-learning via digital identity attributes. J.UCS: Journal of Universal Computer Science, 17(2):296– ??, ???? 2011. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_17_
2/security_and_privacy_
preservation.

Yang:2005:FDA

[YS05] J. Yang and U. Speidel. A fast T -decomposition algorithm. J.UCS: Journal of Universal Computer Science, 11(6):1083–1101, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

URL http://www.jucs.
org/jucs_11_6/a_fast_
t_decomposition.

Yong:2009:CTA

[YSP09] J. Yong, W. Shen, and J. A. Pino. CSCWD technologies, applications and challenges. J.UCS: Journal of Universal Computer Science, 15(9):1744– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15;
http://www.jucs.org/
jucs_15_9#;   http://
www.jucs.org/jucs_15_
9/cscwd_technologies_
applications_and.

Yasugi:2005:SCF

[YTM05] M. Yasugi, Y. Tsujii, and T. Mori. Sequential computability of a function. effective fine space and limiting recursion. J.UCS: Journal of Universal Computer Science, 11(12): 2179, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
12/sequential_computability_
of_a.

You:2013:IS

[YW13] I. You and E. Weippl. Information security. J.UCS: Journal of Universal Computer Science, 19 (16):2347–??, ???? 2013. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/#
;   http://www.jucs.
org/jucs_19;   http:
//www.jucs.org/jucs_
19_16/information_security.

Yang:2008:FSD

[YWD08] G. Yang, D. S. Wong, and X. Deng. Formal security definition and efficient construction for roaming with a privacy-preserving extension. J.UCS: Journal of Universal Computer Science, 14(3):441– 462, ???? 2008. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
14_3/formal_security_
definition_and.

Yang:2010:DMU

[YX10] L. Yang and Y. Xu.

Decision making with uncertainty information based on lattice-valued fuzzy concept lattice. J.UCS: Journal of Universal Computer Science, 16(1):159–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_1/decision_making_
with_uncertainty.

Yu:2009:MWD

[YYZ^{+}09] Z. Yu, Z. Yu, X. Zhou, D. Zhang, and Y. Nakamura. Meeting warmingup: Detecting common interests and conflicts among participants before a meeting. J.UCS: Journal of Universal Computer Science, 15(12): 2311–??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
15_12/meeting_warming_
up_detecting.

Zurita:2008:UPM

[ZAB^{+}08] G. Zurita, P. Antunes, N. Baloian, L. Carriço, F. Baytelman, and M. Sá. Using PDAs in meetings: Patterns, architecture and components. J.UCS: Journal of Universal Computer Science, 14(1):123–147, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_14_1/using_pdas_
in_meetings.

Zurita:2007:MSE

[ZABB07] G. Zurita, P. Antunes, N. Baloian, and F. Baytelman. Mobile sensemaking: Exploring proximity and mobile applications in the classroom. J.UCS: Journal of Universal Computer Science, 13(10):1434–1448, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
10/mobile_sensemaking_
exploring_proximity.

Zygmunt:2012:KPA

[ZBKK12] A. Zygmunt, P. Bródka, P. Kazienko, and J. Koźlak. Key person analysis in social communities within the blogosphere. J.UCS: Journal of Universal Computer Science, 18 (4):577–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_18_4/key_person_
analysis_in.

Zhang:2009:MEP

[ZC09] J. Zhang and K.-W. Chau.

Multilayer ensemble pruning via novel multi-sub-swarm particle swarm optimization. J.UCS: Journal of Universal Computer Science, 15(4):840– ??, ???? 2009. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_15_4/multilayer_
ensemble_pruning_via.

Zhu:2009:DSA

[ZD09] D. Zhu and H. Dreher.

Discovering semantic aspects of socially constructed knowledge hierarchy to boost the relevance of Web searching. J.UCS: Journal of Universal Computer Science, 15(8):1685–??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_15_
8/discovering_semantic_
aspects_of.

Zouari:2014:MCS

[ZDE14] M. Zouari, C. Diop, and E. Exposito. Multilevel and coordinated self-management in autonomic systems based on service bus. J.UCS: Journal of Universal Computer Science, 20(3):431– ??, ???? 2014. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
20_3/multilevel_and_
coordinated_self.

Zou:2010:DMT

[ZDI10] J. Wenhui Zou, X. Deng, and M. Li (II). Detecting market trends by ignoring it, some days. J.UCS: Journal of Universal Computer Science, 16(5):852–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_5/detecting_
market_trends_by.

Zwantschko:1998:TPP

[ZFS98] B. Zwantschko, D. Freismuth, and K. Schmaranz. Telematic platform for patient oriented services. J.UCS: Journal of Universal Computer Science, 4(11):856, November 28, 1998. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs/jucs_4_11/
telematic_platform_for_

patient; internal&sk=
01237079.

Zimmermann:1997:CCC

[ZG97] W. Zimmermann and T. Gaul. On the construction of correct compiler back-ends: An ASMapproach. J.UCS: Journal of Universal Computer Science, 3(5):504– 567, May 28, 1997. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//medoc.springer.de:
8000/jucs_3_5/correct_
compiler;   http://www.
jucs.org/jucs_3_5/correct_
compiler; internal&sk=
05460486.

Zanaty:2005:CFS

[ZG05] E. A. Zanaty and M. R.

Girgis. Collect the fitted surfaces into complex based on C^{0} continuity. J.UCS: Journal of Universal Computer Science, 11(6):1102– 1114, ???? 2005. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_11_
6/collect_the_fitted_
surfaces.

Zaupa:2008:SOP

[ZGC^{+}08] F. Zaupa, I. M. S.

Gimenes, D. Cowan, P. Alencar, and C. J. P. Lucena. A serviceoriented process to develop Web applications. J.UCS: Journal of Universal Computer Science, 14(8):1368–1387, ???? 2008. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_14_
8/a_service_oriented_
process.

Zgrzywa:2007:CDD

[Zgr07] M. Zgrzywa. Consensus determining with dependencies of attributes with interval values. J.UCS: Journal of Universal Computer Science, 13(2): 329–344, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

ohttp://www.jucs.org/
jucs_13_2/consensus_
determining_with_dependencies.

Zhou:2006:NSD

[ZHG^{+}06] H. Zhou, K. M. Hou, L. Gineste, C. De Vaulx, and J. Ponsonnaille. A new system dedicated to real-time cardiac arrhythmias tele-assistance and monitoring. J.UCS: Journal of Universal Computer Science, 12(1):30– 44, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_12_
1/a_new_system_dedicated.

Zima:2001:DDS

[Zim01] H. P. Zima. Data distribution specification for high performance computing. J.UCS: Journal of Universal Computer Science, 7(8):736–753, August 28, 2001. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

/www.jucs.org/jucs_7_
8/data_distribution_
specification_for.

Zulkernain:2010:MII

[ZMAS10] S. Zulkernain, P. Madiraju, S. I. Ahamed, and K. Stamm. A mobile intelligent interruption management system. J.UCS: Journal of Universal Computer Science, 16(15):2060–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

http://www.jucs.org/
jucs_16_15/a_mobile_
intelligent_interruption.

Zhang:2012:IT

[ZNX^{+}12] D. Zhang, H. Ning, K. Xu, F. Lin, and L. Yang. Internet of Things. J.UCS: Journal of Universal Computer Science, 18(9):1069– ??, ???? 2012. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL

                                        http://www.jucs.org/#
                                        ;   http://www.jucs.
                                        org/jucs_18;   http:
                                        //www.jucs.org/jucs_
                                        18_9/internet_of_things;
                                         http://www.jucs.org/
                                        jucsrssfeed-issue.

Zouari:2006:SCR

[Zou06] B. Zouari. A structure causality relation for liveness characterisation in Petri nets. J.UCS: Journal of Universal Computer Science, 12(2):214– 232, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_12_
                                        2/a_structure_causality_
                                        relation.

Zuck:2003:VMT

[ZPFG03] L. Zuck, A. Pnueli, Y. Fang, and B. Goldberg. VOC: a methodology for the translation validation of optimizing compilers. J.UCS: Journal of Universal Computer Science, 9(3):223– 247, March 28, 2003. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_9_
                                        3/voc_a_methodology_
                                        for.

Zhang:2015:MAA

[ZQQ15] G. Zhang, J. Qin, and S. Qazi. Multi-authority attribute-based encryption scheme from lattices. J.UCS: Journal of Universal Computer Science, 21(3):483–??, ???? 2015. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
21_3/multi_authority_
attribute_based.

Zenou:2004:GLT

[ZS04] E. Zenou and M. Samuelides.

Galois lattice theory for probabilistic visual landmarks. J.UCS: Journal of Universal Computer Science, 10(8):1014–1033, August 28, 2004. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
10_8/galois_lattice_
theory_for.

Zheng:2014:AIR

[ZSG14] C. Zheng, W. Shen, and H. Ghenniwa. An adaptive intent resolving scheme for service discovery and integration. J.UCS: Journal of Universal Computer Science, 20(13):1791–??, ???? 2014. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_20_
13/an_adaptive_intent_
resolving.

Zhao:2009:DDW

[ZSK09] J. Zhao, K.-D. Schewe, and H. Koehler. Dynamic data warehouse design with abstract state machines. J.UCS: Journal of Universal Computer Science, 15(1):355– ??, ???? 2009. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_15_
                                        1/dynamic_data_warehouse_
                                        design.

Zambrano:2015:VSA

[ZTN^{+}15] A. Zambrano, C. Toro, M. Nieto, R. Sotaquira, C. Sanín, and E. Szczerbicki. Video semantic analysis framework based on run-time production rules — towards cognitive vision. J.UCS: Journal of Universal Computer Science, 21(6):856– ??, ???? 2015. CO-DEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

                                        //www.jucs.org/jucs_
                                        21_6/video_semantic_
                                        analysis_framework.

Zeng:2007:MIG

[ZTX^{+}07] T. Zeng, C. Tang, Y. Xiang, P. Chen, and Y. Liu. A model of immune gene expression programming for rule mining. J.UCS: Journal of Universal Computer Science, 13(10):1484–1497, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_13_
10/a_model_of_immune.

ZadJabbari:2010:OBA

[ZWH10] B. ZadJabbari, P. Wongthongtham, and F. Khadeer Hussain. Ontology based approach in knowledge sharing measurement. J.UCS: Journal of Universal Computer Science, 16(6):956–??, ???? 2010. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_
16_6/ontology_based_
approach_in.

Zhang:1995:GCG

[ZZ95] Xian-Mo Zhang and Yuliang Zheng. GAC — the criterion for global avalanche characteristics of cryptographic functions. J.UCS: Journal of Universal Computer Science, 1(5):320–337, May 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/gac_the_
criterion_for_global_
avalanche_characteristics_
of_cryptographic_functions.

Zhang:1996:DCC

[ZZ96] Xian-Mo Zhang and Yuliang Zheng. On the difficulty of constructing cryptographically strong substitution boxes. J.UCS: Journal of Universal Computer Science, 2(3): 147–162, March 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:/

                                        /www.jucs.org/jucs_2_
                                        3/on_the_difficulty_
                                        of.

Zheng:2000:ONB

[ZZ00] Y. Zheng and X. Zhang. The k-th-order nonhomomorphicity of S-boxes. J.UCS: Journal of Universal Computer Science, 6(8):830–848, August 28, 2000. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic).

                                        URL http://www.jucs.
                                        org/jucs_6_8/the_kth_
                                        order_nonhomomorphicity.

Zamfirescu:2007:HTG

[ZZ07] C. Zamfirescu and T. Zamfirescu. Hamiltonicity of topological grid graphs. J.UCS: Journal of Universal Computer Science, 13(11):1791–1800, ???? 2007. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

                                        www.jucs.org/jucs_13_

11/hamiltonicity_of_
topological_grid.

Zhang:2012:NMA

[ZZH^{+}12] G. Zhang, F. Zhou, X. Huang, J. Cheng, M. Gheorghe, F. Ipate, and R. Lefticaru. A novel membrane algorithm based on particle swarm optimization for solving broadcasting problems. J.UCS: Journal of Universal Computer Science, 18(13):1821–??, ???? 2012. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://

www.jucs.org/jucs_18_
13/a_novel_membrane_
algorithm.